### v1

# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

Abstract
Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality. Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond a small number of state variables (often limited to  3-6 in prior numerical studies). To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework. PG-DPO leverages Pontryagin’s Maximum Principle to directly guide neural network policies via backpropagation-through-time, naturally incorporating exogenous state processes without requiring dense state grids. Crucially, our computationally efficient “Two-Stage” variant exploits rapidly stabilizing costate estimates derived from BPTT, converting them into near-optimal closed-form Pontryagin controls after only a short warm-up, significantly reducing training overhead. This enables a breakthrough in scalability: numerical experiments demonstrate that PG-DPO successfully tackles problems with dimensions previously considered far out of reach, optimizing portfolios with up to 50 assets and 10 state variables. The framework delivers near-optimal policies, offering a practical and powerful alternative for high-dimensional continuous-time portfolio choice.
1 Introduction
High-dimensional optimal control problems often fall prey to the “curse of dimensionality” (Bellman, , 1966), whereby the computational cost of classical dynamic programming (DP) approaches grows exponentially with the number of state variables. In a continuous-time finance setting, this difficulty becomes starkly evident in multi-asset Merton problems (Merton, , 1971), where an investor dynamically allocates wealth among numerous risky assets whose drift and volatility may depend on exogenous state variables such as macroeconomic indicators or stochastic volatility factors. While DP or Hamilton–Jacobi–Bellman (HJB) PDE solvers provide a rigorous framework for low-dimensional cases (Bertsekas & Tsitsiklis, , 1995), they become intractable as the number of assets or state dimensions increases, with prior influential studies often limited to analyzing problems involving up to approximately 3-5 state variables (Campbell & Viceira, , 1999; Lynch, , 2001; Balduzzi & Lynch, , 1999) or specific frameworks handling around 6 state variables (Garlappi & Skoulakis, , 2010), often necessitating simulation-based approaches for tractability (Brandt et al., , 2005).
A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong & Zhou, , 2012). Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint). Notably, adding one more state dimension does not produce an exponential explosion in computational requirements, in contrast to DP (Bellman, , 1966; Campbell et al., , 2003). This makes Pontryagin-based methods appealing for large-scale portfolio optimization problems involving many assets and exogenous state processes (Karatzas et al., , 1987).
Recent advances in deep learning further enhance the scalability of methods for solving high-dimensional stochastic control problems (e.g., Han et al., , 2018; Raissi et al., , 2019, for general PDE-based approaches). Concurrently, deep learning techniques are being applied directly to financial dynamic programming. Some approaches tackle infinite horizon problems via the HJB equation (Duarte et al., , 2024), while others aim to directly solve the associated nonlinear PDEs using methods like deep branching algorithms (Nguwi et al., , 2024). However, applications such as the Merton HJB have revealed limitations in accuracy or efficiency for certain PDE solvers like Deep BSDE variants (Nguwi et al., , 2024). Our approach, distinct from these direct HJB/PDE solvers, focuses on using Pontryagin’s principle, readily handling finite horizons. One notable strategy related to the use of adjoint methods in optimal control is to directly parameterize the policy (i.e., consumption and investment controls) with neural networks, then apply backpropagation-through-time (BPTT) to enforce Pontryagin’s optimality conditions (Huh, , 2024). This strategy, known as “Pontryagin-Guided Direct Policy Optimization (PG-DPO),” circumvents the need to approximate value functions or PDE solutions. Instead, it uses simulation-based gradients that automatically incorporate the adjoint processes, drastically reducing computational overhead in high-dimensional settings (Kushner & Yin, , 2003; Borkar & Borkar, , 2008).
The purpose of this paper is to develop and implement this PG-DPO framework for the general multi-asset, state-dependent continuous-time portfolio problem. We consider an investor who must decide on a continuous consumption rate and an allocation to $n$ risky assets with state-dependent drift and volatility. By parameterizing the policy with neural networks, we can capture potentially complex relationships dependent on both wealth and the exogenous state. We show how a single forward pass in simulation, coupled with a backward pass via automatic differentiation (BPTT), yields unbiased estimates of the costate processes that PMP requires, guiding the policy towards optimality without the drawbacks of DP.
Furthermore, we introduce a “Two-Stage” PG-DPO variant that exploits the observation that costate estimates often stabilize much faster than the full policy network. Once the costates are sufficiently stable (after a short warm-up), one can bypass the potentially suboptimal neural policy outputs and directly substitute the costate estimates into the PMP first-order conditions. This recovers a near-optimal control analytically at every time-state point with minimal additional training, significantly reducing computational burden while maintaining high accuracy. This efficiency enables the solution of problems with dozens of assets and state variables (e.g., n=50, k=10 in our experiments), a scale considerably larger than those typically addressed by the numerical methods cited previously.
The main contributions of this paper can be summarized as follows:
  • We develop and implement the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, leveraging PMP and deep learning, for solving general multi-asset, multi-factor continuous-time portfolio optimization problems with state-dependent dynamics and finite horizons.
  • We introduce a computationally efficient “Two-Stage” PG-DPO variant that utilizes rapidly stabilized costate estimates from BPTT directly within the PMP first-order conditions, significantly reducing training overhead while achieving high accuracy.
  • We demonstrate through numerical experiments that our framework, particularly the Two-Stage variant, successfully scales to high-dimensional problems involving dozens of assets and state variables (up to n=50, k=10). This substantially exceeds the dimensionality manageable by traditional DP methods and tackled in many prior numerical studies focusing on problems with typically fewer than  5-6 state variables, thus effectively mitigating the curse of dimensionality.
In the remainder of this paper, we first present the multi-asset Merton problem with exogenous states in Section 2. We focus on the affine multi-factor OU model, detailed in Section 2.3.2, for developing the methodology and conducting numerical experiments due to the availability of a benchmark solution. Section 3 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm and its Two-Stage extension. Numerical experiments presented in Section 4 confirm the framework’s scalability and near-optimal performance for the affine model. Finally, Section 5 concludes and discusses potential future directions, including applications to non-affine models or problems with portfolio constraints.
2 Multi-Asset Continuous-Time Portfolio Problem with State Variables
In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1971), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously. Each risky asset features its own drift and volatility, leading to an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}$. Furthermore, an exogenous state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ influences market parameters such as returns and interest rates, adding additional complexity to the investor’s decision.
Although the wealth $X_{t}$ remains scalar, the investor’s optimal policy now depends on both $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and a high-dimensional state $\mathbf{Y}_{t}$. We handle this through Pontryagin’s maximum principle (Pontryagin, , 2014; Yong & Zhou, , 2012), introducing a costate for $X_{t}$ and another for $\mathbf{Y}_{t}$. Numerically, one can represent $\{\bm{\pi}_{t},C_{t}\}$ via neural networks, thereby enabling policy gradient or dynamic programming algorithms even in multi-asset, state-dependent environments.
2.1 Problem Formulation with State Variables and Neural Network Controls
Let $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ be a vector of state variables that affect the distribution of asset returns. For instance, $\mathbf{Y}_{t}$ can include interest rates, macroeconomic indices, volatility factors, and so forth. We assume each component $Y_{i,t}$ follows an Ito diffusion
$dY_{i,t}\;=\;\mu_{Y,i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt\;+\;\sigma_{Y,i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dW^{Y}_{i,t},\quad i=1,\dots,k,$
with $\operatorname{Cov}\bigl{(}dW^{Y}_{i,t},\,dW^{Y}_{j,t}\bigr{)}=\phi_{ij}\,dt.$ In vector form,
$d\mathbf{Y}_{t}=\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,d\mathbf{W}^{Y}_{t},$
where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})$ is the $k\times 1$ drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})$ is often assumed to be a $k\times k$ diagonal matrix of diffusion coefficients, and $\mathbf{W}^{Y}_{t}$ is a $k\times 1$ vector of potentially correlated Brownian motions driving the state variables.
We consider a financial market with $n$ risky assets and one risk-free asset. For $i=1,\dots,n$, the price $S_{i,t}$ of the $i$-th risky asset follows
$dS_{i,t}\;=\;S_{i,t}\bigl{(}\mu_{i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt\;+\;\sigma_{i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dW^{X}_{i,t}\bigr{)}.$
Here, $\mu_{i}(t,\mathbf{Y}_{t})$ is the instantaneous expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ is the instantaneous volatility of asset $i$. The $n$ scalar Brownian motions $W^{X}_{i,t}$ may be correlated amongst themselves ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt$) and with the state variable processes ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt$). In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})$ be the $n\times 1$ vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})$ be the $n\times n$ diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}$ be the $n\times 1$ vector of Brownian motions. The risky asset dynamics can be written compactly as
$d\mathbf{S}_{t}=\bm{\mu}(t,\mathbf{Y}_{t})\,\mathbf{S}_{t}\,dt+\bm{\sigma}(t,\mathbf{Y}_{t})\,\mathbf{S}_{t}\,d\mathbf{W}^{X}_{t}.$
Let $X_{t}$ denote the investor’s total wealth. At each time $t$, the investor chooses a consumption rate $C_{t}\geq 0$ and allocates fractions of wealth to the $n$ risky assets, represented by the vector $\bm{\pi}_{t}=(\pi_{1,t},\,\dots,\,\pi_{n,t})^{\top}\in\mathbb{R}^{n}.$ The remaining fraction $1-\bm{\pi}_{t}^{\top}\mathbf{1}$ is allocated to the risk-free asset, which earns interest at rate $r\bigl{(}t,\mathbf{Y}_{t}\bigr{)}$. The wealth process then evolves according to the stochastic differential equation (SDE):
$dX_{t}\;=\;\Bigl{[}\,X_{t}\,\Bigl{(}\bigl{(}1-\bm{\pi}_{t}^{\top}\mathbf{1}\bigr{)}\,r\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\;+\;\bm{\pi}_{t}^{\top}\,\bm{\mu}(t,\mathbf{Y}_{t})\Bigr{)}\;-\;C_{t}\Bigr{]}\,dt\;+\;X_{t}\,\bm{\pi}_{t}^{\top}\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$
We require the controls $(\bm{\pi}_{t},C_{t})$ to be admissible, ensuring that the wealth remains strictly positive, $X_{t}>0$, for all $t\in[0,T]$. This condition is essential for the utility function to be well-defined and is typically maintained in simulations (e.g., by clamping wealth at a small positive threshold, see Section 4.1.2).
For clarity on the noise structure, we define the combined $(n+k)$-dimensional Brownian vector $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$. The covariance structure of its increments $d\mathbf{W}_{t}$ is given by the block matrix:
$\operatorname{Cov}\bigl{(}d\mathbf{W}_{t}\bigr{)}\;=\;\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
where $\Psi\in\mathbb{R}^{n\times n}=\operatorname{Cov}(d\mathbf{W}^{X}_{t})/dt$, $\Phi\in\mathbb{R}^{k\times k}=\operatorname{Cov}(d\mathbf{W}^{Y}_{t})/dt$, and $\rho\in\mathbb{R}^{n\times k}=\operatorname{Cov}(d\mathbf{W}^{X}_{t},d\mathbf{W}^{Y}_{t})/dt$.
The investor’s objective is to choose admissible control processes $(\bm{\pi},C)$ to maximize the expected discounted utility functional $J$:
[…]
subject to the wealth dynamics. Here $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ weights the utility of terminal wealth (bequest motive). The utility function $U:(0,\infty)\to\mathbb{R}$ represents the investor’s preferences and is assumed to be strictly increasing, strictly concave, and satisfy the Inada conditions. A widely used specification is the Constant Relative Risk Aversion (CRRA) form:
[…]
with $\gamma>0$ denoting the coefficient of relative risk aversion.
In our numerical approach, we parameterize the control policies using neural networks:
$\bm{\pi}_{t}\;=\;\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\quad C_{t}\;=\;C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},$
where $\theta$ and $\phi$ represent the trainable parameters of the respective neural networks.
2.2 Pontryagin’s Maximum Principle with a State-Dependent Market and Costate for $\mathbf{Y}_{t}$
We now consider the continuous-time portfolio problem from Section 2.1 under a market whose parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$. The Pontryagin Maximum Principle (PMP) provides necessary conditions for optimality via a system of forward-backward stochastic differential equations (FBSDEs).
The state of the system consists of the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$. We introduce the costate processes associated with these states: $\lambda_{t}\approx\frac{\partial V}{\partial x}(t,X_{t},\mathbf{Y}_{t})$ (scalar costate for wealth) and $\bm{\eta}_{t}\approx\frac{\partial V}{\partial\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})$ (a $k$-dimensional costate vector for the state variables), where $V$ is the value function of the problem. The dynamics of the state and costate processes are described by the following FBSDE system.
Under an optimal control policy $\bigl{(}\bm{\pi}_{t}^{*},\,C_{t}^{*}\bigr{)}$, the state processes $(X_{t}^{*},\mathbf{Y}_{t}^{*})$ evolve according to the forward SDEs:
$dX_{t}^{*}$ $=\;\Bigl{[}X_{t}^{*}\,\Bigl{(}\,\bigl{(}1-\bm{\pi}_{t}^{*\top}\mathbf{1}\bigr{)}\,r\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\;+\;\bm{\pi}_{t}^{*\top}\,\bm{\mu}\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\Bigr{)}\;-\;C_{t}^{*}\Bigr{]}\,dt$
$\quad+\;X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\,\bm{\sigma}(t,\mathbf{Y}_{t}^{*})\,d\mathbf{W}_{t}^{X},$ (1)
$d\mathbf{Y}_{t}^{*}$ $=\;\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\,dt\;+\;\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\,d\mathbf{W}_{t}^{Y},\quad X_{0}^{*}\;=\;x_{0}>0,\quad\mathbf{Y}_{0}^{*}\;=\;y_{0}\in\mathbb{R}^{k}.$
Here $\bm{\sigma}(t,\mathbf{Y}_{t}^{*})$ is the $n\times n$ (diagonal) matrix of volatilities, and $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t}^{*})$ is the $k\times k$ diffusion matrix for $\mathbf{Y}_{t}^{*}$. The Brownian motions $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ and $\mathbf{W}_{t}^{Y}\in\mathbb{R}^{k}$ have the covariance structure $\operatorname{Cov}(d\mathbf{W}_{t})=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}dt.$
Simultaneously, the costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy the backward SDEs (BSDEs):
$-d\lambda_{t}^{*}$ $=\;\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t}^{*},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*}\bigr{)}\,dt\;-\;\mathbf{Z}_{t}^{\lambda X}\,d\mathbf{W}_{t}^{X}\;-\;\mathbf{Z}_{t}^{\lambda Y}\,d\mathbf{W}_{t}^{Y},$ (2)
$-d\bm{\eta}_{t}^{*}$ $=\;\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t}^{*},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*}\bigr{)}\,dt\;-\;\mathbf{Z}_{t}^{\eta X}\,d\mathbf{W}_{t}^{X}\;-\;\mathbf{Z}_{t}^{\eta Y}\,d\mathbf{W}_{t}^{Y}.$ (3)
Here, $\mathcal{H}_{x}$ and $\mathcal{H}_{\mathbf{y}}$ denote the partial derivatives of the Hamiltonian function $\mathcal{H}$ with respect to $x$ and $\mathbf{y}$, evaluated at the optimal state and controls. The terms $\mathbf{Z}_{t}^{\lambda X},\mathbf{Z}_{t}^{\lambda Y},\mathbf{Z}_{t}^{\eta X},\mathbf{Z}_{t}^{\eta Y}$ are adapted processes representing the diffusion coefficients of the costates, related to second derivatives of the value function and the system’s volatility. The terminal conditions are $\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)}$ and $\bm{\eta}_{T}^{*}=\mathbf{0}.$
The Hamiltonian $\mathcal{H}$ associated with this control problem, whose partial derivatives appear in the BSDE drift terms, is given by:
$\begin{split}\mathcal{H}(t,x,\mathbf{y},\bm{\pi},C,\lambda,\bm{\eta})=&e^{-\delta t}\,U(C)+\lambda\Bigl{[}x\bigl{(}(1-\bm{\pi}^{\top}\mathbf{1})r(t,\mathbf{y})+\bm{\pi}^{\top}\bm{\mu}(t,\mathbf{y})\bigr{)}-C\Bigr{]}\\ &+\bm{\eta}^{\top}\bm{\mu}_{Y}(t,\mathbf{y}).\end{split}$
Optimality requires that the controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ maximize the Hamiltonian at each time $t$. The first-order condition (FOC) with respect to consumption $C_{t}$ yields $e^{-\delta t}\,U^{\prime}(C_{t}^{*})-\lambda_{t}^{*}=0$, implying $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*}).$ The FOC with respect to the portfolio weights $\bm{\pi}_{t}$ leads to a condition balancing expected excess returns against risk, involving the costate $\lambda_{t}^{*}$ and its sensitivity $\mathbf{Z}_{t}^{\lambda X}$. A standard form for this condition is:
[…]
Here, $\mathbf{Z}_{t}^{\lambda X}$ is represented as a $1\times n$ row vector.
Using Ito’s lemma to relate $\mathbf{Z}_{t}^{\lambda X}$ to the value function derivatives ($\lambda_{t}=V_{X}(t,X_{t},\mathbf{Y}_{t})$, so $\partial_{x}\lambda_{t}=V_{XX}$, $\partial_{\mathbf{Y}}\lambda_{t}=V_{X\mathbf{Y}}$), one can solve the FOC (4) for the optimal portfolio $\bm{\pi}_{t}^{*}$. The resulting expression, consistent with standard results in continuous-time finance, is:
$\begin{split}\bm{\pi}_{t}^{*}=-\frac{\lambda_{t}^{*}}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}&\Sigma(t,\mathbf{Y}_{t}^{*})^{-1}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t}^{*})-r(t,\mathbf{Y}_{t}^{*})\mathbf{1}\bigr{)}\\ -\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}&\Sigma(t,\mathbf{Y}_{t}^{*})^{-1}\bm{\sigma}(t,\mathbf{Y}_{t}^{*})\rho\bm{\sigma}_{Y}(t,\mathbf{Y}_{t}^{*})^{\top}(\partial_{\mathbf{Y}}\lambda_{t}^{*})^{\top}.\end{split}$ (5)
Here, $\Sigma(t,\mathbf{Y}_{t}^{*})=\bm{\sigma}(t,\mathbf{Y}_{t}^{*})\Psi\bm{\sigma}(t,\mathbf{Y}_{t}^{*})^{\top}$ is the instantaneous covariance matrix of asset returns (assuming $\bm{\sigma}$ is diagonal, this matches previous definition). The first term represents the myopic demand, driven by instantaneous expected excess return and risk aversion. The second term represents the intertemporal hedging demand, driven by the desire to hedge against unfavorable changes in the state variables $\mathbf{Y}_{t}$, depending on the sensitivity of marginal utility to state variables ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$), state variable volatility ($\bm{\sigma}_{Y}$), and correlations ($\rho,\Psi,\Phi$).
In the special case $n=1,k=1$, we have $\Sigma=\sigma^{2}$, $\Psi=1$, $\rho$ is a scalar correlation, $\bm{\sigma}=\sigma$, $\bm{\sigma}_{Y}=\sigma_{Y}$. The formula (2) simplifies to:
$\pi_{t}^{*}\;=\;-\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})\sigma^{2}}\left\{\lambda_{t}^{*}\bigl{(}\mu(t,Y_{t}^{*})-r(t,Y_{t}^{*})\bigr{)}+\sigma\rho\sigma_{Y}(\partial_{Y}\lambda_{t}^{*})\right\},$
which clearly separates the myopic and hedging components.
2.3 Applications to Financial Models
2.3.1 A Multi-Asset Model with One OU Factor
We first consider a model with $n$ risky assets whose risk premia are driven by a single Ornstein–Uhlenbeck (OU) factor. Specifically, each asset $i$ has a constant volatility $\sigma_{i}>0$ and a time-varying risk premium proportional to an OU state variable $Y_{t}$, such that the expected return is $r+\alpha_{i}\sigma_{i}Y_{t}$, where $r$ is the constant risk-free rate and $\alpha_{i}$ is a loading coefficient. The state variable $Y_{t}$ follows the OU process:
$dY_{t}=\kappa_{Y}(\theta_{Y}-Y_{t})\,dt+\sigma_{Y}\,dW_{t}^{Y},$
where $\kappa_{Y}>0$ is the mean-reversion speed, $\theta_{Y}\in\mathbb{R}$ is the long-term mean, and $\sigma_{Y}>0$ is the factor volatility. The Brownian motion $W_{t}^{Y}$ may be correlated with the asset return processes.
The price dynamics for the $i$-th risky asset, $S_{i,t}$, are given by:
$\frac{dS_{i,t}}{S_{i,t}}=(r+\alpha_{i}\sigma_{i}Y_{t})\,dt+\sigma_{i}\,dW_{i,t}^{S}.$
The Brownian motions driving the assets, $\mathbf{W}_{t}^{S}=(W_{1,t}^{S},\dots,W_{n,t}^{S})^{\top}$, have an $n\times n$ covariance matrix $\Psi dt$, i.e., $\mathrm{Cov}(d\mathbf{W}_{t}^{S})=\Psi\,dt$. Furthermore, each $dW_{i,t}^{S}$ may have a correlation $\rho_{i}$ with $dW_{t}^{Y}$; let $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$.
Let $X_{t}$ denote total wealth. The investor chooses a consumption rate $C_{t}\geq 0$ and portfolio weights $\bm{\pi}_{t}\in\mathbb{R}^{n}$ for the risky assets. The wealth dynamics follow:
$dX_{t}=\left[X_{t}\left(\left(1-\bm{\pi}_{t}^{\top}\mathbf{1}\right)r+\bm{\pi}_{t}^{\top}(r\mathbf{1}+\bm{\alpha}\bm{\sigma}Y_{t})\right)-C_{t}\right]dt+X_{t}\bm{\pi}_{t}^{\top}\mathrm{diag}(\bm{\sigma})\,d\mathbf{W}_{t}^{S},$
where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$, and $\bm{\sigma}=(\sigma_{1},\dots,\sigma_{n})^{\top}$. The instantaneous covariance matrix of the $n$-asset returns is $\Sigma=\mathrm{diag}(\bm{\sigma})\,\Psi\,\mathrm{diag}(\bm{\sigma})$.
For an investor with CRRA utility $U(x)$ and discount rate $\delta>0$, the optimal consumption $C_{t}^{*}$ is determined by $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$, where $\lambda_{t}^{*}$ is the costate associated with wealth $X_{t}$. The PMP first-order condition yields the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$ as:
$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}\Sigma^{-1}\left\{\lambda_{t}^{*}(\bm{\alpha}\bm{\sigma})Y_{t}+(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}(\bm{\sigma}\bm{\rho})\right\}.$
This expression involves the costate $\lambda_{t}^{*}$ and its derivatives $\partial_{x}\lambda_{t}^{*}$ and $\partial_{Y}\lambda_{t}^{*}$. The first term within the braces represents the myopic demand driven by the risk premium $(\bm{\alpha}\bm{\sigma})Y_{t}$, and the second term captures the intertemporal hedging demand arising from the state variable $Y_{t}$. If $n=1$, this simplifies to $\pi_{t}^{*}=-\frac{1}{\sigma^{2}X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\{\lambda_{t}^{*}\alpha\sigma Y_{t}+\sigma(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}\rho\}$. For the specific case without intermediate consumption ($C_{t}=0$), an analytic solution involving Riccati ODEs can be derived using the HJB framework (Kim & Omberg, , 1996), as detailed in Appendix B.1.
2.3.2 A Multi-Asset, Multi-Factor Model
We then extend this setup to a more general multi-factor environment, considering $n$ risky assets whose risk premia are driven by a $k$-dimensional OU factor vector $\mathbf{Y}_{t}\in\mathbb{R}^{k}$. The risk premium for asset $i=1,\dots,n$ is $\sigma_{i}(\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t})$, where $\bm{\alpha}_{i}\in\mathbb{R}^{k}$ contains the factor loadings for asset $i$. The volatility $\sigma_{i}$ for asset $i$ is constant. This allows for richer interactions between multiple risk sources and asset returns.
The factor vector $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ follows a multivariate OU process:
$d\mathbf{Y}_{t}\;=\;\kappa_{Y}\bigl{(}\bm{\theta}_{Y}-\mathbf{Y}_{t}\bigr{)}\,dt\;+\;\bm{\sigma}_{Y}\,d\mathbf{W}_{t}^{Y},$
where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is a mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean vector, and $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ governs the factor volatilities and correlations (e.g., $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi dt$).
The $n$ risky assets, with prices $S_{t}^{i}$, follow:
$\frac{dS_{t}^{i}}{S_{t}^{i}}\;=\;\Bigl{(}r+\sigma_{i}\,\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t}\Bigr{)}\,dt\;+\;\sigma_{i}\,dW_{t}^{X,i},\quad i=1,\dots,n.$
Here, $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ is a Brownian motion with covariance matrix $\Psi dt$. The asset shocks $\mathbf{W}_{t}^{X}$ and factor shocks $\mathbf{W}_{t}^{Y}$ can be correlated, with $\mathrm{Cov}(dW_{t}^{X,i},dW_{t}^{Y,j})=\rho_{ij}dt$. Let $\rho\in\mathbb{R}^{n\times k}$ be the matrix of cross-correlations.
With portfolio weights $\bm{\pi}_{t}=(\pi_{t}^{1},\dots,\pi_{t}^{n})^{\top}$ and consumption $C_{t}$, wealth $X_{t}$ evolves as:
$dX_{t}=\left[X_{t}\left(r+\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}(\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t})\right)-C_{t}\right]dt+X_{t}\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}\,dW_{t}^{X,i}.$
For CRRA utility with discount $\delta$, the optimal consumption $C_{t}^{*}$ depends on the wealth costate $\lambda_{t}^{*}$. The optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$ is given by the PMP first-order condition (Eq. (2)), which in this context becomes:
$\bm{\pi}_{t}^{*}=-\;\frac{1}{\,X_{t}^{*}\,\bigl{(}\partial_{x}\lambda_{t}^{*}\bigr{)}\!}\;\Sigma^{-1}\;\Bigl{\{}\,\underbrace{\lambda_{t}^{*}\;\mathrm{diag}(\bm{\sigma})\,\mathbf{A}\,\mathbf{Y}_{t}}_{\text{Myopic Term}}~{}+~{}\underbrace{\mathrm{diag}(\bm{\sigma})\,\rho\,\bm{\sigma}_{Y}\bigl{(}\partial_{\mathbf{Y}}\lambda_{t}^{*}\bigr{)}^{\top}}_{\text{Hedging Term}}\Bigr{\}},$
where $\bm{\sigma}=(\sigma_{1},\dots,\sigma_{n})^{\top}$, $\Sigma=\mathrm{diag}(\bm{\sigma})\Psi\mathrm{diag}(\bm{\sigma})$ is the $n\times n$ asset return covariance, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}$ is the $n\times k$ matrix of factor loadings, and $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ is the $k\times 1$ gradient vector of the costate with respect to the factors $\mathbf{Y}_{t}$. This explicitly separates the demand into a myopic part driven by the current expected excess returns ($\mathrm{diag}(\bm{\sigma})\,\mathbf{A}\,\mathbf{Y}_{t}$) and a hedging part driven by the sensitivity to factor changes ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and their correlations ($\rho$, $\Psi$, $\Phi$).
Similar to the single-factor case, an analytic solution for the terminal wealth problem ($C_{t}=0$) with CRRA utility can often be found using an exponential-quadratic value function ansatz, leading to matrix Riccati ODEs. Details are presented in Appendix B.2.
3 A Gradient-Based Algorithm for Multi-Asset Policy Optimization
In the preceding sections, we established the theoretical underpinnings of the continuous-time multi-asset portfolio problem using Pontryagin’s Maximum Principle (PMP). This principle provides necessary optimality conditions characterized by a coupled system of forward-backward stochastic differential equations (FBSDEs) for the state and costate variables (Section 2.2). However, finding analytical solutions to these FBSDEs is often infeasible, especially in high-dimensional settings involving numerous assets ($\bm{\pi}_{t}\in\mathbb{R}^{n}$) or complex exogenous state dynamics ($\mathbf{Y}_{t}\in\mathbb{R}^{k}$). Consequently, translating the continuous-time optimality conditions into a practical solution requires effective numerical approximations.
We adopt a gradient-based approach rooted in modern deep learning techniques. The core idea is to directly parameterize the investor’s policy, both the investment strategy $\bm{\pi}_{t}$ and the consumption rate $C_{t}$, using neural networks, denoted as $\bm{\pi}_{\theta}(t,x,\mathbf{y})$ and $C_{\phi}(t,x,\mathbf{y})$, where $(\theta,\phi)$ are the network parameters. Optimization then involves iteratively refining these parameters to maximize the investor’s expected utility.
A key challenge is computing the gradient of the expected utility objective with respect to the policy parameters $(\theta,\phi)$. Furthermore, a practical policy must be effective across a relevant range of potential starting times $t_{0}$, initial wealth levels $x_{0}$, and initial state variable values $\mathbf{y}_{0}$.
Let $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ denote the conditional expected utility obtained when starting from state $(t_{0},x_{0},\mathbf{y}_{0})$ and following the policy defined by $(\theta,\phi)$:
[…]
Here, the expectation $\mathbb{E}[\cdot|\cdot]$ is taken over all possible paths of the state processes $(X_{u},\mathbf{Y}_{u})$ for $u\in[t_{0},T]$, conditional on starting at $(x_{0},\mathbf{y}_{0})$ at time $t_{0}$.
To ensure the policy performs well across various starting scenarios, we define the optimization target as the extended value function, $\widetilde{J}(\theta,\phi)$. This is the expected value of the conditional utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ when the initial state $(t_{0},x_{0},\mathbf{y}_{0})$ is drawn from a chosen distribution $\eta$ over a relevant domain $\mathcal{D}\subset[0,T]\times\mathbb{R}^{+}\times\mathbb{R}^{k}$:
$\widetilde{J}(\theta,\phi)\;=\;\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\eta}\Big{[}J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\Big{]}.$ (7)
Maximizing this extended objective $\widetilde{J}(\theta,\phi)$ encourages the learning of a robust policy applicable across the domain $\mathcal{D}$ covered by the initial state distribution $\eta$.
This paper develops and implements the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, extending the core ideas of Huh, (2024) to the multi-asset, state-dependent setting, with the goal of maximizing the extended value function $\widetilde{J}(\theta,\phi)$. The PG-DPO framework leverages automatic differentiation, specifically backpropagation-through-time (BPTT), applied to Monte Carlo simulations drawn from the initial state distribution $\eta$ and evolved according to the system dynamics.
A central feature of applying BPTT in this context is that it not only provides unbiased estimates of the policy gradient $\nabla\widetilde{J}$, necessary for optimizing the neural network parameters $(\theta,\phi)$, but also inherently calculates estimates of the Pontryagin costate variables ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$) along each simulated path. These costates are crucial as they represent the sensitivities of the expected utility to state perturbations, as dictated by PMP.
This dual output of BPTT–policy gradients and costate estimates–allows for two distinct algorithmic approaches, both offering scalable alternatives to traditional methods hindered by the curse of dimensionality:
  • The Baseline PG-DPO algorithm (§3.1) utilizes the BPTT-derived policy gradient estimates ($\nabla_{\text{batch}}\approx\nabla\widetilde{J}$) directly within a stochastic gradient ascent procedure to train the neural network policy $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end. In this approach, the costate estimates primarily serve to structure the gradient calculation, implicitly guiding the optimization according to PMP.
  • The Two-Stage PG-DPO algorithm (§3.2) takes a different route. It uses the baseline procedure for a short warm-up phase specifically to generate stable and accurate estimates of the costates (potentially involving Monte Carlo averaging). These refined costate estimates are then directly plugged into the analytical first-order optimality conditions derived from PMP to define the deployed control actions, bypassing the direct use of the warm-up phase neural networks.
Both variants avoid explicit FBSDE solvers and leverage the efficiency of BPTT, but they differ significantly in how they utilize the information generated during backpropagation to arrive at the final control policy. The subsequent subsections detail each approach.
3.1 Baseline PG-DPO for the Extended Objective
The baseline PG-DPO algorithm aims to maximize the extended value function $\widetilde{J}(\theta,\phi)$ (defined in Eq. (7)) using mini-batch stochastic gradient ascent. The core idea is to approximate the gradient $\nabla\widetilde{J}$ by averaging gradients computed on individual sample paths, where each path starts from a randomly drawn initial state. Each iteration $j$ of the algorithm involves the following steps:
  1. 1.
    Sample Initial States: Draw a mini-batch of $M$ initial states $\{(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)})\}_{i=1}^{M}$ from the distribution $\eta$ over the domain $\mathcal{D}$.
  2. 2.
    Simulate Forward Paths: For each sampled initial state $i$, simulate a discrete-time trajectory $\{(X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\}_{k=0}^{N}$ forward from $t_{k}=t_{0}^{(i)}$ to $t_{N}=T$, using the current policy networks $(\bm{\pi}_{\theta},C_{\phi})$ and correctly correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ derived from the Cholesky factor $L$ of $\mathbf{\Sigma}_{\text{block}}$. The state evolution follows:
    $\mathbf{Y}_{k+1}^{(i)}$ $=\mathbf{Y}_{k}^{(i)}+\bm{\mu}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta t^{(i)}+\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta\mathbf{W}_{k}^{(i),Y},$
    $X_{k+1}^{(i)}$ $=X_{k}^{(i)}+\left[X_{k}^{(i)}\left((1-\bm{\pi}_{k}^{(i)\top}\mathbf{1})r(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})+\bm{\pi}_{k}^{(i)\top}\bm{\mu}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\right)-C_{k}^{(i)}\right]\Delta t^{(i)}$
    $\quad+X_{k}^{(i)}\bm{\pi}_{k}^{(i)\top}\bm{\sigma}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta\mathbf{W}_{k}^{(i),X},$
    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$.
  3. 3.
    Calculate Realized Reward per Path: For each path $i$, compute the realized cumulative reward, explicitly noting its dependence on the starting state and policy parameters, denoted as $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)$:
    $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)=\sum_{k=0}^{N-1}e^{-\delta t_{k}^{(i)}}U\bigl{(}C_{\phi}(t_{k}^{(i)},X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\bigr{)}\Delta t^{(i)}+\kappa e^{-\delta T}U(X_{N}^{(i)}).$
    This value $J^{(i)}(t_{0}^{(i)},\dots)$ is a stochastic sample related to the conditional expected utility $J(t_{0}^{(i)},\dots)$ defined in Eq. (6).
  4. 4.
    Apply BPTT for Policy Gradient Estimate: Treat the simulation process for path $i$ as a computational graph. Apply automatic differentiation (BPTT) to compute the gradient of the realized reward $J^{(i)}(t_{0}^{(i)},\dots)$ with respect to the policy parameters $(\theta,\phi)$:
    $\nabla_{(\theta,\phi)}J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi).$
    This gradient serves as an unbiased estimate related to the gradient of the conditional expected utility $J(t_{0}^{(i)},\dots)$.
  5. 5.
    Estimate Gradient of $\widetilde{J}$: Obtain an unbiased estimate for the gradient of the overall extended objective $\widetilde{J}$ by averaging the individual path gradients over the mini-batch:
    $\nabla\widetilde{J}(\theta,\phi)\approx\nabla_{\text{batch}}=\frac{1}{M}\sum_{i=1}^{M}\nabla_{(\theta,\phi)}J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi).$
  6. 6.
    Update Policy Parameters: Update the parameters $(\theta,\phi)$ using this averaged gradient estimate via stochastic gradient ascent:
    $(\theta,\phi)_{j+1}\leftarrow(\theta,\phi)_{j}+\alpha_{j}\,\nabla_{\text{batch}},$
    where $\alpha_{j}$ is the learning rate.
By iterating these steps, the baseline PG-DPO algorithm gradually optimizes the policy $(\bm{\pi}_{\theta},C_{\phi})$ to maximize the extended value function $\widetilde{J}$.
Role of Intermediate BPTT Values and the “Pontryagin-Guided” Nature.
It is important to understand the role of the intermediate values computed during BPTT in Step 4. As BPTT calculates the policy gradient $\nabla J^{(i)}(\dots)$ via the chain rule, it inherently computes the partial derivatives of the realized reward $J^{(i)}(\dots)$ with respect to the intermediate states along path $i$:
$\lambda_{k}^{(i)}=\frac{\partial J^{(i)}(t_{0}^{(i)},\dots)}{\partial X_{k}^{(i)}},\quad\bm{\eta}_{k}^{(i)}=\frac{\partial J^{(i)}(t_{0}^{(i)},\dots)}{\partial\mathbf{Y}_{k}^{(i)}}.$
These values, $\lambda_{k}^{(i)}$ and $\bm{\eta}_{k}^{(i)}$, are unbiased estimates of the Pontryagin costates associated with the conditional expected utility $J(t_{0}^{(i)},\dots)$. Crucially, in the baseline PG-DPO algorithm, these costate estimates themselves (and their potential derivatives w.r.t. states) are not directly used in the parameter update rule (Step 6), which only requires the final gradient $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$.
Nevertheless, the fact that BPTT provides access to these estimates is fundamental. Firstly, they are essential inputs for the Two-Stage PG-DPO variant discussed in the next subsection (§3.2). The baseline algorithm thus lays the groundwork by generating these estimates. Secondly, this connection explains why the algorithm is termed Pontryagin-Guided. Even though the baseline update only uses $\nabla J^{(i)}(\dots)$, the calculation of this gradient via BPTT structurally mirrors the backward propagation of sensitivities inherent in PMP’s adjoint equations, using the costate estimates $\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ implicitly within the chain rule (cf. Huh, , 2024). Therefore, the optimization process, driven by these BPTT-computed gradients, is implicitly guided by the optimality conditions derived from Pontryagin’s principle, aiming to maximize the expected objective $\widetilde{J}$.
3.2 Two-Stage PG-DPO Algorithm
While the baseline PG-DPO algorithm described in Section 3.1 effectively optimizes the extended objective $\widetilde{J}$ by training the policy networks $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end, it can sometimes require a significant number of iterations for the networks to fully converge. The Two-Stage PG-DPO variant offers a potentially more computationally efficient alternative by exploiting an empirical observation: the costate estimates ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ and their relevant derivatives with respect to states, like $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$), which are generated during BPTT, often stabilize and become reasonably accurate much faster than the full policy networks achieve optimality.
This Two-Stage approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly from the PMP first-order conditions, bypassing the need for fully converged policy networks for deployment. The procedure consists of two distinct stages, described below using a single level of enumeration:
  1. 1.
    Stage 1: Warm-Up Phase for Costate Estimation
    First, we execute the Baseline PG-DPO process, as detailed in Section 3.1 (which optimizes $\widetilde{J}$), for a predetermined, relatively small number of iterations, $K_{0}$. The primary objective of this stage is not to obtain fully optimized networks $(\bm{\pi}_{\theta},C_{\phi})$, but rather to run the BPTT process sufficiently long for the underlying mechanism generating the costate estimates ($\lambda_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial X_{k}^{(i)}$, $\bm{\eta}_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial\mathbf{Y}_{k}^{(i)}$) and their relevant state-derivatives (e.g., $\partial_{x}\lambda_{k}^{(i)}$, $\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$) to stabilize. We let $(\theta^{*},\phi^{*})$ denote the policy parameters obtained at the end of this warm-up phase.
  2. 2.
    Stage 2: Analytic Control Deployment
    After completing the $K_{0}$ warm-up iterations, the trained policy networks $(\bm{\pi}_{\theta^{*}},C_{\phi^{*}})$ are typically not directly used for generating controls during deployment. Instead, to determine the control actions for any given state $(t_{k},X_{k},\mathbf{Y}_{k})$ encountered during deployment, we employ a Monte Carlo-like averaging procedure to obtain reliable costate estimates needed for the PMP formulas:
    First, to obtain the stabilized costate estimates corresponding to the state $(t_{k},X_{k},\mathbf{Y}_{k})$, we perform the following averaging: Simulate multiple ($M_{\text{MC}}$) forward paths starting from $(t_{k},X_{k},\mathbf{Y}_{k})$ (or states very close to it) up to the horizon $T$, using the fixed policy parameters $(\theta^{*},\phi^{*})$ obtained from Stage 1. For each simulated path $j=1,\dots,M_{\text{MC}}$, calculate its realized reward $J^{(j)}(t_{k},X_{k},\mathbf{Y}_{k};\theta^{*},\phi^{*})$ and apply BPTT to obtain the costate estimates $\lambda_{k}^{(j)},\partial_{x}\lambda_{k}^{(j)},\partial_{\mathbf{Y}}\lambda_{k}^{(j)}$ specifically at the starting node $k$ for that path $j$. The final “stabilized” estimates are then computed by averaging these values across the $M_{\text{MC}}$ paths:
    $\lambda_{k}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\lambda_{k}^{(j)},\quad\partial_{x}\lambda_{k}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{x}\lambda_{k}^{(j)},\quad\partial_{\mathbf{Y}}\lambda_{k}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{\mathbf{Y}}\lambda_{k}^{(j)}$
    This averaging acts as a Monte Carlo estimation step, significantly reducing the variance inherent in the single-path BPTT estimates derived during Stage 1 and providing more robust inputs for the PMP formulas.
    Next, we plug these averaged costate estimates ($\lambda_{k},\partial_{x}\lambda_{k},\partial_{\mathbf{Y}}\lambda_{k}$) directly into the analytical formulas derived from the Pontryagin Maximum Principle’s first-order conditions (Section 2.2). The optimal consumption is calculated as:
    $C_{k}^{\mathrm{PMP}}=\bigl{(}U^{\prime}\bigr{)}^{-1}\!\bigl{(}e^{\delta\,t_{k}}\,\lambda_{k}\bigr{)}.$
    The optimal investment portfolio is calculated as:
[…]
    $+\underbrace{\bm{\sigma}(t_{k},\mathbf{Y}_{k})^{\top}\!\bigl{[}(\partial_{\mathbf{Y}}\lambda_{k})\,\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\rho^{\top}\,\Psi^{-1}\bigr{]}}_{\text{Hedging Demand}}\Biggr{\}}.$
    Finally, these analytically computed values $(C_{k}^{\mathrm{PMP}},\bm{\pi}_{k}^{\mathrm{PMP}})$, based on the averaged costate estimates, are used as the control actions for the state $(t_{k},X_{k},\mathbf{Y}_{k})$ during deployment.
In summary, the Two-Stage PG-DPO approach performs a short warm-up training (Stage 1) using the baseline PG-DPO process to obtain a policy $(\theta^{*},\phi^{*})$ whose BPTT generates reasonably stable costate estimates. It then leverages these estimates in Stage 2, refining them through Monte Carlo averaging at the point of deployment and plugging them into the known analytical structure of the optimal control derived from PMP. This strategy can significantly reduce the overall computational burden while yielding high-quality, near-optimal controls. The Monte Carlo averaging step in Stage 2 is crucial for obtaining reliable inputs for the PMP formulas from the potentially noisy single-path estimates.
3.3 Rationale for the Two-Stage Approach: Leveraging Costate Estimation
The preceding subsections introduced two variants of the PG-DPO framework: a baseline approach (§3.1) and a Two-Stage approach (§3.2). A natural question arises: why introduce the Two-Stage variant, which potentially adds complexity, instead of solely relying on optimizing the policy networks directly? This subsection elaborates on the theoretical and practical motivations.
3.3.1 Qualitative Rationale: Exploiting PMP Structure and Costate Properties
At its core, the rationale stems from the structure provided by Pontryagin’s Maximum Principle (PMP) itself and the economic meaning of the costate variables. As established in Section 2.2, PMP characterizes the optimal controls ($\bm{\pi}_{t}^{*},C_{t}^{*}$) directly in terms of the costates ($\lambda_{t}^{*},\bm{\eta}_{t}^{*}$) and their relevant derivatives (e.g., $\partial_{x}\lambda_{t}^{*},\partial_{\mathbf{Y}}\lambda_{t}^{*}$). The analytical formulas used in Stage 2 of the Two-Stage method are precisely these first-order conditions. Therefore, if accurate estimates of these costates and their derivatives could be obtained, PMP provides a direct blueprint for constructing the optimal policy.
Furthermore, the primary costate $\lambda_{t}$ represents the marginal utility of wealth ($\lambda_{t}\approx\partial V/\partial x$). Economic principles, like consumption smoothing, suggest agents strive for a stable marginal utility path. Mathematically, this often means the value function $V$ and its derivatives ($\lambda_{t},\partial_{x}\lambda_{t},\partial_{\mathbf{Y}}\lambda_{t}$, etc.) exhibit relatively smooth behavior or possess a simpler analytical structure compared to the optimal policy functions $(\bm{\pi}^{*},C^{*})$ themselves. The optimal policy might need to be highly complex, especially in non-affine settings or under constraints.
This difference in potential complexity has direct implications for learning dynamics. It is plausible that learning the potentially smoother or more structured functions associated with the costates via BPTT during the warm-up phase (Stage 1) is an easier task-requiring fewer iterations for stabilization-than learning the full complexity of the optimal policy networks $(\bm{\pi}_{\theta},C_{\phi})$ until convergence. This aligns with empirical observations that costate estimates often stabilize relatively quickly.
The potential for faster convergence or stabilization of costate-related functions compared to the full policy motivates the Two-Stage approach. This is further encouraged, albeit heuristically for general models, by theoretical stability results known for specific structured models, as formalized in the next subsection.
3.3.2 Derivative Stability in Structured Models
For certain classes of portfolio choice problems that admit explicit solutions, such as those within affine or quadratic frameworks, a strong link exists between the accuracy of the value function approximation and the accuracy of its derivatives, provided the approximation retains the original structure. This property is formally stated as follows:
Proposition 3.1 (Derivative Closeness from Value Closeness in Structured Models).
Consider a class of portfolio choice models (e.g., Merton with HARA utility (Merton, , 1971), Kim & Omberg with HARA utility (Kim & Omberg, , 1996), or Liu’s affine/quadratic framework with CRRA utility (Liu, , 2007)) where the optimal value function $V^{*}(t,w,\mathbf{y})$ has a known analytical structure (e.g., separable or exponential-polynomial in state variables with coefficients determined by ODEs). Let $\tilde{V}(t,w,\mathbf{y})$ be another function defined on a compact domain $\mathcal{D}\subset[0,T]\times\mathbb{R}^{+}\times\mathbb{R}^{k}$ that shares the same analytical structure as $V^{*}$. If the value functions are uniformly close on $\mathcal{D}$, i.e.,
$\sup_{(t,w,\mathbf{y})\in\mathcal{D}}|V^{*}(t,w,\mathbf{y})-\tilde{V}(t,w,\mathbf{y})|\longrightarrow 0,$
then their corresponding spatial derivatives (first and second order with respect to $w$ and $\mathbf{y}$) also converge uniformly on $\mathcal{D}$:
$\sup_{\mathcal{D}}\|\nabla_{w,\mathbf{y}}V^{*}-\nabla_{w,\mathbf{y}}\tilde{V}\|\to 0\quad\text{and}\quad\sup_{\mathcal{D}}\|\nabla^{2}_{w,\mathbf{y}}V^{*}-\nabla^{2}_{w,\mathbf{y}}\tilde{V}\|\to 0.$
(Proof sketch provided in Appendix C for the Liu (2007) framework).
Implication for PG-DPO Rationale.
Proposition 3.1 highlights that for well-structured problems (like those within the affine class for which the proposition holds), the derivatives inherent in the PMP conditions are well-behaved with respect to approximations that preserve the specific analytical structure. However, a key motivation for employing methods like PG-DPO is their potential applicability to a broader class of problems, including models that may lack such convenient structures — for instance, models involving non-affine dynamics (such as the CEV examples considered later in this paper). The true optimal value function $V^{*}$ for these more general models may not possess the specific separable or exponential-polynomial form assumed in the proposition.
Consequently, Proposition 3.1, proven for specific structured models, does not directly guarantee that similar derivative stability holds for the wider range of potentially non-affine models that PG-DPO aims to tackle. Establishing whether analogous stability properties exist in these more general settings, or rigorously proving the convergence of derivative estimates obtained via BPTT from neural network approximations in such cases, remains an area for future research and is beyond the scope of this paper.
Nevertheless, the proposition provides valuable context. It confirms that the target objects (optimal value function derivatives) possess inherent stability within important, albeit specific, classes of models. This offers heuristic support for the hypothesis underlying the Two-Stage PG-DPO: namely, that the costate derivatives, being fundamental quantities in a well-posed optimal control problem, might be estimated reliably via BPTT (Stage 1) and effectively used in the PMP formulas (Stage 2), even when tackling more complex, potentially non-affine models.
4 Numerical Experiments
In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework (both Baseline and Two-Stage variants). To enable rigorous comparison with the available semi-analytical benchmark for the terminal wealth problem (Appendix B.2), the numerical experiments presented herein focus exclusively on maximizing $J=\mathbb{E}[U(X_{T})]$. Consequently, the optimization of the intermediate consumption policy $C_{\phi}$, though supported by the general framework (Sections 2-3), is omitted in this study. We focus on the multi-asset, multi-factor portfolio optimization problem introduced in Section 2.3.2. As detailed therein, this specific model involves $n$ risky assets whose expected returns are driven by $k$ correlated Ornstein-Uhlenbeck (OU) state variables $\mathbf{Y}_{t}$.
Crucially, this model specification falls within the affine class. This allows us to compute a semi-analytical benchmark solution by solving a system of Riccati-type ordinary differential equations (ODEs), as outlined in Appendix B.2. This benchmark provides a basis for rigorously comparing the accuracy of the PG-DPO methods.
4.1 Experimental Setup
4.1.1 Parameter Values
The experiments consider the objective of maximizing the expected utility of terminal wealth $X_{T}$, $J=\mathbb{E}[U(X_{T})]$, using a CRRA utility function $U(x)=x^{1-\gamma}/(1-\gamma)$ with a relative risk aversion coefficient $\gamma=2.0$. The fixed time horizon is $T=1.5$.
Unless otherwise specified, the model parameters (for the model described in Section 2.3.2) are generated based on the asset dimension $n$ and factor dimension $k$ using the following process (seeded for reproducibility, e.g., ‘seed=42’):
  • Risk-free rate: $r=0.03$.
  • State process parameters (for the $k$-dimensional OU process $\mathbf{Y}_{t}$):
    • Mean-reversion speeds: $\bm{\kappa}_{Y}=\mathrm{diag}(2.0,2.5,\dots,2.0+(k-1)0.5)$.
    • Long-term means: $\bm{\theta}_{Y}$ components drawn uniformly from $U(0.2,0.4)$.
    • Factor volatilities: $\bm{\sigma}_{Y}$ diagonal elements drawn uniformly from $U(0.3,0.5)$.
  • Asset parameters (for the $n$ risky assets):
    • Volatilities: $\bm{\sigma}\in\mathbb{R}^{n}$ components drawn uniformly from $U(0.1,0.5)$.
    • Factor loadings: $\bm{\alpha}\in\mathbb{R}^{n\times k}$ drawn using ‘scipy.stats.dirichlet’ with concentration parameters $[1.0,\dots,1.0]$. (Note: These define the risk premium structure $\bm{\mu}=r\mathbf{1}+\mathrm{diag}(\bm{\sigma})\bm{\alpha}\mathbf{Y}_{t}$).
  • Correlation structure: The generation process aims to produce realistic correlation parameters while ensuring the overall block correlation matrix remains positive definite, which is crucial for stable simulations. This involves:
    • Asset correlation $\Psi$ ($n\times n$) generated based on a latent factor structure ensuring validity.
    • Factor correlation $\Phi_{Y}$ ($k\times k$) generated as a random correlation matrix.
    • Cross-correlation $\bm{\rho}_{Y}$ ($n\times k$) components drawn uniformly from $U(-0.2,0.2)$.
    • The full $(n+k)\times(n+k)$ block correlation matrix $\begin{pmatrix}\Psi&\rho_{Y}\\ \rho_{Y}^{\top}&\Phi_{Y}\end{pmatrix}$ is constructed and numerically adjusted (e.g., via minimal diagonal loading) if necessary to ensure positive definiteness (tolerance $10^{-6}$).
The specific values generated by the seed are used consistently across experiments for given $n$ and $k$.
4.1.2 PG-DPO Implementation Details
We implement the Baseline PG-DPO algorithm using PyTorch.
  • Policy Network ($\bm{\pi}_{\theta}$): The policy network architecture consists of an input layer taking the current wealth $W_{t}$, current time $t$, and state vector $\mathbf{Y}_{t}$ (total $2+k$ inputs), followed by three hidden layers with 200 units each and LeakyReLU activation, and an output layer producing the $n$-dimensional portfolio weight vector $\bm{\pi}_{\theta}$.
  • Optimizer: Adam optimizer with a learning rate of $1\times 10^{-5}$.
  • Training:
    • Total epochs: $10,000$.
    • Batch size ($M$): $1,000$.
    • Time discretization steps ($N$): $20$.
    • Initial state sampling ($\eta$): For each trajectory in a batch, the initial time $t_{0}$ is drawn uniformly from $[0,T)$, initial wealth $W_{0}$ from $U(0.1,3.0)$. The initial state variables $\mathbf{Y}_{0}=(Y_{0,1},\dots,Y_{0,k})^{\top}$ are sampled such that each component $Y_{0,i}$ is drawn independently and uniformly from a single common range $[Y_{\min},Y_{\max}]$. This range is determined by the minimum lower bound and the maximum upper bound observed across all individual factors’ typical ±$\pm 3\sigma$ intervals:
      $Y_{\min}=\min_{i=1,\dots,k}(\theta_{Y,i}-3\sigma_{Y,ii}),\quad Y_{\max}=\max_{i=1,\dots,k}(\theta_{Y,i}+3\sigma_{Y,ii}),$
      where $\theta_{Y,i}$ is the long-term mean and $\sigma_{Y,ii}$ is the instantaneous volatility parameter for the $i$-th factor. The simulation then runs forward from $t_{0}$ to the fixed horizon $T$ using $N=20$ time steps, with step size $\Delta t=(T-t_{0})/N$. This approach ensures the policy is trained effectively across the entire time interval $[0,T]$ and a relevant hypercube in the state space defined by the collective factor dynamics.
    • Variance Reduction: Antithetic variates used.
    • Wealth constraint: Simulated wealth $W_{t}$ clamped at lower bound $0.1$.
  • Costate Estimation (for Two-Stage Evaluation): During periodic evaluation, costate $\lambda$ and derivatives $\partial_{x}\lambda,\partial_{\mathbf{Y}}\lambda$ are computed via BPTT (‘torch. autograd.grad’) applied to Monte Carlo estimates of the expected terminal utility starting from evaluation points $(t_{k},W_{k},\mathbf{Y}_{k})$.
4.1.3 Benchmark Solution and Evaluation
As mentioned, the chosen multi-factor OU model allows for a semi-analytical benchmark solution. It is important to note that these experiments focus solely on the terminal wealth objective ($J=\mathbb{E}[U(X_{T})]$), and thus do not include optimization or a neural network for the consumption policy $C_{\phi}$. This simplification is made specifically to enable direct comparison with the semi-analytical benchmark solution derived from the HJB equation, which is readily available for the terminal wealth problem in this affine setting.
  • Benchmark Computation: The system of ODEs (including a matrix Riccati equation) associated with the HJB equation for this terminal wealth problem (Appendix B.2) is solved numerically backwards from the fixed horizon $T$ using ‘scipy.integrate.solve_ivp’ (‘Radau’ method) to obtain the coefficients determining the value function and the optimal policy $\bm{\pi}^{*}(t,W_{t},\mathbf{Y}_{t})$.
  • Interpolation: The benchmark policy (total, myopic, and hedging components) is pre-computed on a grid over $(t,W_{t},\mathbf{Y}_{t})$ space and stored using ‘scipy. interpolate.RegularGridInterpolator’ for efficient lookup.
  • Evaluation Metrics: Performance is evaluated via the Root Mean Squared Error (RMSE) between the policies generated by (i) Baseline PG-DPO ($\bm{\pi}_{\theta}$) and the benchmark ($\bm{\pi}^{*}$), and (ii) Two-Stage PG-DPO (using Eq. (2) with BPTT-estimated costates) and the benchmark ($\bm{\pi}^{*}$). RMSEs are computed for total, myopic, and hedging components.
  • Visualization: Contour plots compare the learned policies and the benchmark over the state space (typically time $t$ vs. one state variable $Y_{k}$), alongside error plots and costate visualizations.
All computations are performed using PyTorch on an NVIDIA A6000 GPU.
4.2 Convergence Speed and Accuracy Analysis
This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO methodologies across various dimensional settings. Figure 1 displays the Root Mean Squared Error (RMSE) between the learned policies (either $\bm{\pi}_{\theta}$ for Baseline PG-DPO or $\bm{\pi}^{\mathrm{PMP}}$ for Two-Stage PG-DPO) and the benchmark optimal policy ($\bm{\pi}^{*}$) as a function of training epochs, plotted on a logarithmic scale. The results are shown for configurations of (a) n=1 asset, k=10 state variables, (b) n=10, k=10, and (c) n=50, k=10.
(a) n=1, k=10
(b) n=10, k=10
(c) n=50, k=10
Figure 1: RMSE comparison between Baseline PG-DPO (blue) and Two-Stage PG-DPO (orange) policies relative to the benchmark solution across different dimensions. Note the logarithmic scale on the y-axis.
The RMSE for the Two-Stage PG-DPO (orange line in Figure 1) represents the error of the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ constructed by substituting the costate estimates obtained via BPTT at each epoch into the PMP first-order condition (Eq. (2)). A key advantage, and distinction from the Baseline PG-DPO which learns the total policy $\bm{\pi}_{\theta}$ monolithically, is that the Two-Stage approach explicitly computes the policy by combining the analytically distinct myopic and hedging components derived from the PMP formula. This allows for separate assessment of their accuracy. As established theoretically for affine models with CRRA utility (Appendix B) and confirmed in preliminary component-wise error analyses, the error associated with the myopic component (driven by the ratio $-\lambda_{t}^{*}/(X_{t}^{*}\partial_{x}\lambda_{t}^{*})=1/\gamma$) is typically negligible and converges almost instantly. Therefore, the total RMSE observed for the Two-Stage PG-DPO policy (orange line in Figure 1) is almost entirely attributable to the error in estimating the hedging component. The remarkably low overall error achieved by the Two-Stage method thus highlights its proficiency in accurately capturing this crucial, often complex, hedging demand using BPTT-derived costate derivatives ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$).
The key finding, consistently observed across all cases presented in Figure 1, is that the Two-Stage PG-DPO (orange) converges significantly faster and achieves a substantially lower final (or minimum) RMSE compared to the Baseline PG-DPO (blue). This clearly demonstrates the superiority of the Two-Stage approach, which leverages costate estimates derived from BPTT to reconstruct the policy according to the PMP structure.
Analyzing each case provides further insights:
  • n=1, k=10 (Figure 1a): In this scenario with numerous state variables, the Two-Stage method converges rapidly within approximately 1,000 epochs to a very low RMSE level of around $3\times 10^{-3}$. In contrast, the Baseline method converges much slower, stabilizing after about 4,000 epochs at a significantly higher RMSE level of approximately $7\times 10^{-2}$. This results in a substantial gap (over 20-fold) in the final accuracy between the two methods.
  • n=10, k=10 (Figure 1b): A similar trend persists in this intermediate-scale case with an increased number of assets. The Two-Stage method still converges quickly, reaching an RMSE of about $1.5\times 10^{-2}$ within roughly 2,000 epochs, whereas the Baseline method exhibits slower convergence and achieves a less accurate result with an RMSE around $1\times 10^{-1}$.
  • n=50, k=10 (Figure 1c): For this most complex, high-dimensional case, training was conducted only up to 3,000 epochs. Nevertheless, the performance difference is stark. The Two-Stage method displays much faster initial convergence, reaching a minimum RMSE of approximately $1.5\times 10^{-2}$ near the 3,000-epoch mark. Although a slight tendency for the error to increase afterward is observed—potentially indicating the need for a longer warm-up phase (Stage 1) or suggesting instability in costate estimates requiring further training or tuning—its achieved accuracy is remarkable. Considering that the Baseline method’s RMSE is still above $2\times 10^{-1}$ at the same 3,000-epoch point, we can conclude that the Two-Stage approach yields a significantly more accurate policy within the same computational budget.
These results strongly support the rationale behind the Two-Stage approach discussed in Section 3.3. Specifically, BPTT can effectively estimate the necessary costate information (particularly its derivatives) required by the PMP conditions with relatively few training iterations. Integrating these estimates into the analytical structure provided by the PMP optimality conditions proves to be more efficient and accurate than attempting to directly approximate the complex optimal policy function end-to-end using a neural network (Baseline approach), especially as the problem dimensionality increases.
Building upon this quantitative analysis, the subsequent sections will provide visual comparisons of the policy functions for specific cases via contour plots, offering further insights into the qualitative differences between the policies generated by the two methods. Hereafter, we often abbreviate Two-Stage PG-DPO as 2-PG-DPO for brevity
4.3 Detailed Analysis for the Single-Asset, Single-Factor Case (n=1, k=1)
We begin with a detailed examination of the simplest state-dependent setting (n=1, k=1) to visually validate the methodologies and illustrate the structure of the learned policies. The results presented in this subsection correspond to the state of the policies after 2,000 training epochs, a point observed to be near optimal for the 2-PG-DPO method in this configuration (cf. Figure 1(a)) before potential performance degradation with extended training, as discussed later.
Figure 2 compares the total policies learned by the Baseline PG-DPO and 2-PG-DPO methods, alongside their respective errors relative to the benchmark solution. Figure 3 then decomposes the superior 2-PG-DPO policy into its constituent myopic and hedging components, showing their individual shapes and associated errors. All plots show the policy or error values against time-to-maturity ($T-t$) and the state variable ($Y$).
(a) Baseline Policy ($\pi_{\theta}$)
(b) 2-PG-DPO Policy ($\pi^{\mathrm{PMP}}$)
(c) Baseline Error
(d) 2-PG-DPO Error
Figure 2: Comparison of total policies and errors for the n=1, k=1 case at epoch 2000. Top row compares the policies from Baseline PG-DPO (left) and 2-PG-DPO (right). Bottom row compares their respective errors. Note the significantly different color bar scales for the error plots.
(a) 2-PG-DPO Myopic ($\pi_{0}^{\mathrm{PMP}}$)
(b) 2-PG-DPO Hedging ($\pi_{1}^{\mathrm{PMP}}$)
(c) Myopic Error
(d) Hedging Error
Figure 3: Decomposition of the 2-PG-DPO policy for the n=1, k=1 case at epoch 2000. Top row shows the myopic (left) and hedging (right) components. Bottom row shows their respective errors. Note the extremely small error scale for the myopic component.
Observing Figure 2, we compare the Baseline PG-DPO policy (2(a)) and its error (2(c)) with the 2-PG-DPO policy (2(b)) and its error (2(d)). While the Baseline network learns the general dependency on the state variable $Y$, its error plot (2(c)) reveals significant deviations from the benchmark, with magnitudes reaching up to ±$\approx\pm 0.1$. This visual representation aligns with the relatively high RMSE seen for the Baseline method around epoch 2000 in Figure 1(a). Although this error level is considerably higher than that of the 2-PG-DPO method even in this simple n=1, k=1 setting, it represents the performance in the simplest possible setting. It is crucial to emphasize that the effectiveness of the Baseline PG-DPO approach tends to diminish considerably as the number of assets ($n$) and state variables ($k$) grows. The challenge of directly approximating the increasingly complex optimal policy function often becomes intractable for the Baseline method in higher dimensions, leading to much larger errors or convergence difficulties, as hinted by the widening gap in RMSE observed in Figure 1 for larger $n$. Therefore, while the Baseline method shows some learning capability here, its limitations in scalability motivate the need for more robust approaches like 2-PG-DPO for the high-dimensional problems central to this paper. In stark contrast, the 2-PG-DPO policy (2(b)) yields a drastically reduced error (2(d)), with magnitudes mostly within ±$\approx\pm 0.001$, corresponding well with its low RMSE and suggesting close proximity to the true optimal solution, demonstrating its effectiveness even in this simple case.
Figure 3 provides crucial insights into the superior performance of the 2-PG-DPO method by decomposing its policy. The myopic component (3(a)) is visually dominant. Crucially, its associated error (3(c)) is extremely small (order $10^{-7}$), confirming its near-perfect accuracy. The hedging component (3(b)), representing a smaller adjustment, has an associated error (3(d)) that is visually identical in pattern and magnitude (order $10^{-3}$) to the total error of the 2-PG-DPO policy (2(d)). This visually confirms that the overall accuracy of the 2-PG-DPO policy is primarily limited by the accuracy of estimating the hedging demand.
Finally, it is important to address the training duration and the potential for overfitting with the 2-PG-DPO method. While this method achieves high accuracy remarkably quickly (around 2000 epochs here, corresponding to only about 5-10 minutes of computation time, given that 10,000 epochs take approximately 30 minutes on the utilized NVIDIA A6000 GPU), we observed that continued training beyond this near-optimal point can lead to a degradation of its performance (i.e., an increase in RMSE, as hinted in Figure 1(a)). This phenomenon might suggest that the underlying neural network used for costate estimation during the warm-up phase (Stage 1) could potentially start to overfit to the specific discrete time steps ($\Delta t>0$) and sample paths inherent in the simulation-based training[cite: 293]. Instead of learning the true continuous-time ($dt$) value function derivatives needed for the PMP formulas, the network might adapt too closely to the artifacts of the discretization. This underscores the importance of careful monitoring and potentially early stopping for the warm-up phase—which requires only a short computational time—to extract the best performance from the 2-PG-DPO methodology.
In summary, the analysis for the n=1, k=1 case visually confirms the high accuracy of the 2-PG-DPO approach compared to the Baseline PG-DPO, clearly illustrates the decomposition into highly accurate myopic and less accurate (but still relatively small error) hedging components, and underscores the significance of selecting an appropriate training duration for optimal performance.
4.4 Analysis with Multiple State Variables (n=1, k=10)
We now increase the complexity by considering a case with a single asset but numerous state variables (n=1, k=10). This scenario tests the ability of the methods to handle high-dimensional state dependencies that affect the investment opportunity set. As visualizing the full 10-dimensional state dependency is impossible, the following plots (Figures 4 and 5) display the policy and error values as functions of time-to-maturity ($T-t$) and the first state variable ($Y_{1}$), while holding the remaining state variables ($Y_{2},\dots,Y_{10}$) constant at their respective long-term mean values ($\theta_{Y,i}$). The results shown correspond to 2,000 training epochs, consistent with the previous analysis. The overall convergence for this case is depicted in Figure 1(a).
(a) Baseline Policy ($\pi_{\theta}$)
(b) 2-PG-DPO Policy ($\pi^{\mathrm{PMP}}$)
(c) Baseline Error
(d) 2-PG-DPO Error
Figure 4: Comparison of total policies and errors for the n=1, k=10 case at epoch 2000 (plotted against $Y_{1}$). Top row compares the policies from Baseline PG-DPO (left) and 2-PG-DPO (right). Bottom row compares their respective errors. Note the significantly different color bar scales for the error plots.
(a) 2-PG-DPO Myopic ($\pi_{0}^{\mathrm{PMP}}$)
(b) 2-PG-DPO Hedging ($\pi_{1}^{\mathrm{PMP}}$)
(c) Myopic Error
(d) Hedging Error
Figure 5: Decomposition of the 2-PG-DPO policy for the n=1, k=10 case at epoch 2000 (plotted against $Y_{1}$). Top row shows the myopic (left) and hedging (right) components. Bottom row shows their respective errors. Note the extremely small error scale for the myopic component and the very small magnitude of the hedging component itself.
Figure 4 presents the comparison for the n=1, k=10 case. The Baseline PG-DPO policy (4(a)) again learns the dependency on the displayed state variable $Y_{1}$. However, its error relative to the benchmark (4(c)) remains significant, with magnitudes ranging roughly between 0.001 and 0.064. This level of error is consistent with the high RMSE value observed for the Baseline method in Figure 1(a), indicating difficulties in accurately capturing the complex policy function influenced by ten state variables. This difficulty highlights a key challenge for the Baseline PG-DPO approach, particularly when the number of state variables ($k$) is large. The network must learn to approximate the total policy, which implicitly combines the primary myopic demand with the complex intertemporal hedging demands arising from all ten state variables ($Y_{1},\dots,Y_{10}$). Even if the magnitude of the hedging component related to any single factor or the total hedging demand itself is relatively small (as suggested by Figure 5(b) for the $Y_{1}$ dimension), the sheer number of interacting hedging influences creates a highly intricate functional relationship for the optimal policy. Attempting to capture this subtle, high-dimensional structure directly through gradient updates on a single policy network can disrupt the learning process. This may hinder convergence towards the true optimum or even lead to the network converging to an inaccurate or distorted policy shape, as potentially reflected in the structured errors seen in Figure 4(c). Conversely, the 2-PG-DPO method, by explicitly utilizing the PMP structure and separating the estimation of terms related to $\lambda$ and its derivatives ($\partial_{\mathbf{Y}}\lambda$), demonstrates remarkable accuracy even in this high-dimensional state setting. The total policy (4(b)) results in an extremely small error (4(d)), with magnitudes only reaching up to approximately $3.6\times 10^{-4}$. This visually confirms the near-perfect fit suggested by the very low RMSE value achieved by the 2-PG-DPO method shown in Figure 1(a), showcasing its ability to effectively handle a large number of state factors.
Figure 5 provides further validation through the decomposition of the 2-PG-DPO policy. The myopic component (5(a)) largely determines the shape of the total policy. Its corresponding error (5(c)) remains negligible, on the order of $10^{-7}$, proving its accurate estimation. The hedging component (5(b)) is exceptionally small in this particular slice (magnitude order $10^{-4}$), suggesting that for this specific configuration and visualization plane, the intertemporal hedging demand related to $Y_{1}$ is minimal. Nevertheless, the error associated with this tiny hedging component (5(d)) is again visually identical to the total error of the 2-PG-DPO policy (4(d)), confirming that the overall policy accuracy is still dictated by the precision of the hedging term estimation, even when this term (and its error) is extremely small.
In conclusion, the results for the n=1, k=10 case demonstrate the robustness and high accuracy of the 2-PG-DPO method when faced with high-dimensional state spaces. It significantly outperforms the Baseline PG-DPO approach and accurately captures the policy structure dictated by PMP, with negligible error in the myopic component and the total error determined by the precision of the hedging component estimation. This highlights the suitability of the 2-PG-DPO framework for problems involving complex, multi-factor dynamics influencing investment decisions.
4.5 High-Dimensional Performance (n=50, k=10)
Finally, we evaluate the performance in the most challenging setting considered: a high-dimensional problem with n=50 assets and k=10 state variables. This scenario fully tests the scalability and effectiveness of the proposed methods in dimensions where traditional DP methods are intractable. Similar to the previous section, the visualizations focus on the policy for the first asset ($i=1$) as a function of time-to-maturity ($T-t$) and the first state variable ($Y_{1}$), holding other state variables ($Y_{2},\dots,Y_{10}$) constant at their long-term means ($\theta_{Y,i}$). The results shown correspond to the state after 3,000 training epochs, representing the point where the 2-PG-DPO method achieved its minimum RMSE in Figure 1(c).
(a) Baseline Policy ($\pi_{1,\theta}$)
(b) 2-PG-DPO Policy ($\pi_{1}^{\mathrm{PMP}}$)
(c) Baseline Error
(d) 2-PG-DPO Error
Figure 6: Comparison of total policies and errors for the first asset ($i=1$) in the n=50, k=10 case at epoch 3000 (plotted against $Y_{1}$). Top row compares the policies from Baseline PG-DPO (left) and 2-PG-DPO (right). Bottom row compares their respective errors. Note the dramatic difference in error scales and magnitudes.
(a) 2-PG-DPO Myopic ($\pi_{1,0}^{\mathrm{PMP}}$)
(b) 2-PG-DPO Hedging ($\pi_{1,1}^{\mathrm{PMP}}$)
(c) Myopic Error
(d) Hedging Error
Figure 7: Decomposition of the 2-PG-DPO policy for the first asset ($i=1$) in the n=50, k=10 case at epoch 3000 (plotted against $Y_{1}$). Top row shows the myopic (left) and hedging (right) components. Bottom row shows their respective errors. The myopic error remains negligible, while the hedging error dictates the total error.
The challenges of high dimensionality become starkly evident for the Baseline PG-DPO method in Figure 6. While the learned policy for the first asset (6(a)) shows some structure, the corresponding error (6(c)) is extremely large, with values ranging from approximately -1.1 to -0.7. This indicates severe inaccuracies and potentially a systematic bias in the policy learned by the Baseline method when faced with the combined complexity of 50 assets and 10 state variables. This poor performance aligns with the high and slowly converging RMSE observed in Figure 1(c).
In dramatic contrast, the 2-PG-DPO method continues to provide a highly accurate solution, as shown by its policy (6(b)) and error (6(d)). The error magnitude for the 2-PG-DPO policy is approximately within ±$\pm 0.02$. While this represents a slight increase in error compared to the lower-dimensional cases (e.g., n=1, k=10 where the error was order $10^{-4}$), it is still orders of magnitude smaller (roughly 50 times smaller) than the error of the Baseline PG-DPO method for this same high-dimensional problem. This underscores the robustness and practical necessity of the 2-PG-DPO approach; although perfect accuracy might be harder to achieve as dimensionality increases, it remains remarkably effective and vastly superior to the direct policy learning method. The error level visually corresponds to the minimum RMSE achieved by the 2-PG-DPO method in Figure 1(c).
The component decomposition in Figure 7 confirms the patterns observed previously, even in this complex setting. The myopic component (7(a)) still largely dictates the policy shape, and its associated error (7(c)) remains negligible (order $10^{-7}$), demonstrating the consistent accuracy of its estimation across all dimensions tested. The hedging component (7(b)) appears relatively more significant here compared to the n=1, k=10 case, contributing visibly to the total policy shape. Crucially, the hedging error (7(d)) is again visually identical to the total 2-PG-DPO error (6(d)), confirming that the overall accuracy is dictated by the precision of estimating the hedging demands, which becomes slightly more challenging, yet remains manageable, in higher dimensions.
In conclusion, the n=50, k=10 results powerfully demonstrate the capability of the 2-PG-DPO framework to “break the dimensional barrier”. While the Baseline PG-DPO struggles significantly, 2-PG-DPO delivers a highly accurate and viable solution by leveraging the structure of Pontryagin’s Maximum Principle. Despite a slight, expected increase in error compared to lower-dimensional settings, its performance remains remarkably strong and vastly superior to the baseline, solidifying its potential as a practical tool for large-scale, continuous-time portfolio optimization.
5 Conclusion
In this paper, we addressed the challenge of solving high-dimensional continuous-time portfolio optimization problems, which often suffer from the “curse of dimensionality” when approached with traditional dynamic programming methods. We developed and implemented the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, which leverages Pontryagin’s Maximum Principle (PMP) and deep learning techniques, specifically backpropagation-through-time (BPTT), to directly optimize neural network policies for investment and consumption.
A key contribution is the introduction of a highly effective “Two-Stage” PG-DPO variant. This approach utilizes a short warm-up phase to obtain stable estimates of the crucial Pontryagin costates and their derivatives via BPTT, potentially refined through Monte Carlo averaging. These estimates are then plugged directly into the analytical first-order optimality conditions derived from PMP to construct the control policy. Our numerical experiments, conducted on a multi-asset, multi-factor affine model allowing for benchmark comparisons, demonstrated that the 2-PG-DPO method significantly outperforms the baseline end-to-end approach. It achieves substantially higher accuracy and faster convergence, successfully scaling to problems with dimensions (e.g., n=50 assets, k=10 state factors) previously considered intractable for numerical HJB/PDE solvers, thereby effectively ”breaking the dimensional barrier.” We specifically highlighted the method’s ability to accurately capture both myopic and complex intertemporal hedging demands by leveraging the PMP structure.
This research opens several avenues for future investigation.
Firstly, exploring the theoretical underpinnings further is crucial. While Proposition 3.1 provides heuristic support for the 2-PG-DPO approach in structured models, rigorously analyzing or establishing similar derivative stability properties for the broader class of non-affine models would significantly strengthen the foundation of the method. Furthermore, investigating the theoretical convergence properties of the PG-DPO algorithm, particularly the costate estimation process within the 2-PG-DPO variant, remains an important open question.
Secondly, extending the framework to incorporate more realistic market features is a natural next step. This includes introducing portfolio constraints (such as short-selling restrictions or leverage limits) and modeling transaction costs, which significantly alter the optimal control problem. Adapting the PG-DPO framework to handle these complexities efficiently would enhance its practical applicability. Testing the framework’s robustness and performance with different utility functions beyond CRRA (e.g., HARA, prospect theory) is also warranted.
Thirdly, further empirical validation and algorithmic refinements are desirable. Applying the PG-DPO framework (especially the superior 2-PG-DPO variant) to well-known non-affine financial models (e.g., stochastic volatility models), even in the absence of exact benchmarks, would showcase its capabilities more broadly. Conducting numerical experiments that explicitly include the optimization and evaluation of the intermediate consumption policy, which was omitted in the current study for benchmarking purposes, would provide a more complete assessment of the full framework. Exploring alternative neural network architectures, optimization techniques, and variance reduction methods could potentially lead to further performance improvements.
Finally, optimizing the implementation details, such as the warm-up duration for the 2-PG-DPO method to balance costate accuracy against potential overfitting to simulation artifacts, deserves further attention.
Overall, the Pontryagin-Guided Direct Policy Optimization framework, particularly its Two-Stage variant, offers a promising and scalable approach for tackling complex, high-dimensional optimal control problems in continuous-time finance.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-00562904).
Appendix A Derivation of the Adjoint Diffusion Coefficient via Itô’s Lemma
In the Pontryagin framework, the adjoint (or costate) process associated with the investor’s wealth $X_{t}$ is identified with the partial derivative of the value function $V$. Concretely, we set $\lambda_{t}=V_{X}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},$ where $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ represents additional state variables and $X_{t}$ is the investor’s scalar wealth process. Our goal in this section is to derive the diffusion coefficient corresponding to the $d\mathbf{W}^{X}_{t}$ term in the backward SDE for $\lambda_{t}$, namely $\mathbf{Z}_{t}^{\lambda X}$ from BSDE (2), which plays a crucial role in determining the optimal portfolio via FOC (4).
Recall that under an optimal control $\bigl{(}\bm{\pi}_{t}^{*},C_{t}^{*}\bigr{)}$, the wealth $X_{t}^{*}$ evolves according to Eq. (1), and the $k$-dimensional state process $\mathbf{Y}_{t}^{*}$ follows its own SDE, also presented in Section 2.2. The dynamics are driven by correlated Brownian motions $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ and $\mathbf{W}_{t}^{Y}\in\mathbb{R}^{k}$ with the joint covariance structure $\operatorname{Cov}(d\mathbf{W}_{t})=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}dt.$
Since $\lambda_{t}=V_{X}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},$ we apply Ito’s multidimensional lemma to $V_{X}$. The differential $d\lambda_{t}$ can be expressed as:
$d\lambda_{t}edundant$ $=\;V_{Xt}\,dt\;+\;V_{XX}\,dX_{t}\;+\;V_{X\mathbf{Y}}\,d\mathbf{Y}_{t}$
$\quad+\;\tfrac{1}{2}\,V_{XXX}\,\bigl{(}dX_{t}\bigr{)}^{2}\;+\;V_{XX\mathbf{Y}}\,(dX_{t})\,(d\mathbf{Y}_{t})^{\top}\;+\;\tfrac{1}{2}\,\text{Tr}\bigl{(}V_{X\mathbf{Y}\mathbf{Y}}\,(d\mathbf{Y}_{t})(d\mathbf{Y}_{t})^{\top}\bigr{)},$ (8)
where subscripts denote partial derivatives (e.g., $V_{X\mathbf{Y}}$ is the gradient w.r.t $\mathbf{Y}$ of $V_{X}$, $V_{X\mathbf{Y}\mathbf{Y}}$ is the Hessian w.r.t $\mathbf{Y}$ of $V_{X}$).
Our objective is to identify the coefficient of $d\mathbf{W}_{t}^{X}$ in the expansion of $d\lambda_{t}$. This coefficient is precisely the process $\mathbf{Z}_{t}^{\lambda X}$ appearing in the BSDE (2). Substituting the SDEs for $dX_{t}$ and $d\mathbf{Y}_{t}$ (from Section 2.2) into the Ito formula (8), we observe that terms involving $d\mathbf{W}_{t}^{X}$ arise directly from the diffusion part of the $V_{XX}\,dX_{t}$ term, and indirectly from the $V_{X\mathbf{Y}}\,d\mathbf{Y}_{t}$ term due to the correlation $\rho$ between $d\mathbf{W}_{t}^{X}$ and $d\mathbf{W}_{t}^{Y}$. A careful application of Ito’s lemma, considering the quadratic variation and covariation terms related to the covariance matrix $\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}$, allows isolation of the full coefficient multiplying $d\mathbf{W}_{t}^{X}$. Representing $\mathbf{Z}_{t}^{\lambda X}$ as a $1\times n$ row vector, this process is found to be:
$\mathbf{Z}_{t}^{\lambda X}=\underbrace{(\partial_{x}\lambda_{t}^{*})X_{t}^{*}\bm{\pi}_{t}^{*\top}\bm{\sigma}(t,\mathbf{Y}_{t}^{*})}_{\text{Term from }V_{XX}dX_{t}}+\underbrace{(\partial_{\mathbf{Y}}\lambda_{t}^{*})\bm{\sigma}_{Y}(t,\mathbf{Y}_{t}^{*})\rho^{\top}\Psi^{-1}}_{\text{Term from }V_{XY}dY_{t}\text{ via correlation}}.$ (9)
The first term reflects the direct impact of wealth volatility related to the portfolio choice $\bm{\pi}_{t}^{*}$, scaled by the second derivative of the value function with respect to wealth, $V_{XX}$ (or $\partial_{x}\lambda_{t}^{*}$). The second term captures the intertemporal hedging component arising from the correlation between asset returns and state variable movements, mediated by the cross-derivative of the value function $V_{X\mathbf{Y}}$ (or $\partial_{\mathbf{Y}}\lambda_{t}^{*}$).
This derivation provides the explicit form of the adjoint diffusion coefficient $\mathbf{Z}_{t}^{\lambda X}$, which is fundamental for understanding the structure of the optimal portfolio policy (2) and for constructing the Pontryagin adjoint equations.
Appendix B Analytic Solutions for Affine Models
This appendix provides details on the derivation of analytic solutions for the portfolio optimization problems presented in Section 2.3 under specific assumptions (no intermediate consumption, CRRA utility), typically following the Hamilton-Jacobi-Bellman (HJB) approach with an exponential-quadratic value function ansatz. We also show how these solutions connect to the Pontryagin framework.
B.1 Analytic Solution for the Multi-Asset OU Model (Section 2.3.1)
For simplicity, we assume no intermediate consumption, i.e., $C_{t}=0$, and consider an investor who derives utility only from terminal wealth $X_{T}$ under a CRRA objective $U(x)=\frac{x^{1-\gamma}}{1-\gamma}$. Following Kim & Omberg, (1996), the associated Hamilton–Jacobi–Bellman (HJB) equation can be solved by positing a value function of the form $V(t,x,Y)=\frac{x^{1-\gamma}}{1-\gamma}\exp\left(g(t)+B(t)Y+\frac{1}{2}C(t)Y^{2}\right)$. Here, $g(t)$, $B(t)$, and $C(t)$ are deterministic functions of time determined by solving ordinary differential equations (ODEs) derived from the HJB equation, subject to terminal conditions $g(T)=0$, $B(T)=0$, and $C(T)=0$.
From this value function, the costate $\lambda_{t}^{*}=V_{x}$, its derivative with respect to wealth $\partial_{x}\lambda_{t}^{*}=V_{xx}$, and its derivative with respect to the state $\partial_{Y}\lambda_{t}^{*}=V_{xY}$ have the following closed-form expressions on the optimal path $(X_{t}^{*},Y_{t})$:
$\lambda_{t}^{*}(t,X_{t}^{*},Y_{t})$ $=(X_{t}^{*})^{-\gamma}\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right),$
$\frac{\partial\lambda_{t}^{*}}{\partial x}(t,X_{t}^{*},Y_{t})$ $=-\gamma(X_{t}^{*})^{-\gamma-1}\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right),$
$\frac{\partial\lambda_{t}^{*}}{\partial Y}(t,X_{t}^{*},Y_{t})$ $=(X_{t}^{*})^{-\gamma}\left(B(t)+C(t)Y_{t}\right)\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right).$
These expressions lead to remarkably simple ratios that are central to the Pontryagin framework’s characterization of the optimal portfolio:
$\frac{\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma},\qquad\frac{\partial_{Y}\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma}\left(B(t)+C(t)Y_{t}\right).$
Substituting these ratios into the general Pontryagin first-order condition for the optimal portfolio $\bm{\pi}_{t}^{*}$ (which equates the marginal contribution to the Hamiltonian from investing in risky assets to zero, cf. Eq. (4) derived in Section 2.2) confirms the explicit analytic solution derived via the HJB approach:
$\bm{\pi}^{*}(t,Y)=\frac{1}{\gamma}\,\Sigma^{-1}\Bigl{[}\underbrace{(\bm{\alpha}\,\bm{\sigma})\,Y}_{\text{Risk premium }\bm{\mu}(Y)-r\mathbf{1}}+\underbrace{\sigma_{Y}\,(\bm{\sigma}\,\bm{\rho})\,\bigl{(}B(t)+C(t)\,Y\bigr{)}}_{\text{Hedging demand from }\partial_{Y}\lambda_{t}^{*}}\Bigr{]},$
where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$, $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$, the notation $(\bm{\alpha}\,\bm{\sigma})$ represents the vector $(\alpha_{1}\sigma_{1},\dots,\alpha_{n}\sigma_{n})^{\top}$, and $(\bm{\sigma}\,\bm{\rho})$ represents the vector $(\sigma_{1}\rho_{1},\dots,\sigma_{n}\rho_{n})^{\top}$. The derivation confirms that the first term captures the myopic demand driven by the instantaneous risk premium $(\bm{\mu}(Y)-r\mathbf{1})=(\bm{\alpha}\,\bm{\sigma})Y$, while the second term represents the intertemporal hedging demand, proportional to $B(t)+C(t)Y$ which arises from the state-dependency of the investment opportunity set captured by $\partial_{Y}\lambda_{t}^{*}$.
Substituting this $\bm{\pi}^{*}$ back into the HJB equation and matching coefficients in powers of $Y$ yields the following Riccati-type ODEs for $B(t)$ and $C(t)$:
$B(T)=0,\quad C(T)=0,$
and
$\frac{dC(t)}{dt}$ $=2\,\kappa_{Y}\,C(t)-\sigma_{Y}^{2}\,C(t)^{2}-\frac{1-\gamma}{\gamma}\Bigl{[}\beta_{0}+2\,\sigma_{Y}\,C(t)\,\beta_{1}+\sigma_{Y}^{2}\,C(t)^{2}\,\beta_{2}\Bigr{]},$
$\frac{dB(t)}{dt}$ $=\kappa_{Y}\,B(t)-\sigma_{Y}^{2}\,B(t)\,C(t)+\frac{1-\gamma}{\gamma}\Bigl{[}\sigma_{Y}^{2}\,B(t)\,C(t)\,\beta_{2}+\sigma_{Y}\,B(t)\,\beta_{1}\Bigr{]}-\kappa_{Y}\,\theta_{Y}\,C(t),$
where $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ are quadratic forms involving the model parameters:
$\beta_{0}:=(\bm{\alpha}\,\bm{\sigma})^{\top}\,\Sigma^{-1}\,(\bm{\alpha}\,\bm{\sigma}),\quad\beta_{1}:=(\bm{\alpha}\,\bm{\sigma})^{\top}\,\Sigma^{-1}\,(\bm{\sigma}\,\bm{\rho}),\quad\beta_{2}:=(\bm{\sigma}\,\bm{\rho})^{\top}\,\Sigma^{-1}\,(\bm{\sigma}\,\bm{\rho}).$
Once $B(t)$ and $C(t)$ are obtained by backward integration from their terminal conditions, substituting them into $\bm{\pi}^{*}(t,Y)$ yields the closed-form optimal strategy balancing myopic and hedging demands. The function $g(t)$ can also be found by integrating another ODE involving $B(t)$ and $C(t)$, but it does not affect the optimal policy $\bm{\pi}^{*}$.
B.2 Analytic Solution for the Multi-Asset Multi-Factor Model (Section 2.3.2)
[…]
where $\mathbf{B}(t)$ is an $n\times 1$ vector, $\mathbf{C}(t)$ is an $n\times n$ symmetric matrix, and $g(t)$ is a scalar function, all satisfying ODEs derived from the HJB equation, with terminal conditions $g(T)=0$, $\mathbf{B}(T)=\mathbf{0}$ and $\mathbf{C}(T)=\mathbf{0}$.
Define the relevant matrices and vectors:
$\bm{\sigma}\;=\;\mathrm{diag}(\sigma_{1},\dots,\sigma_{d}),\quad\bm{\sigma}_{Y}\;=\;\mathrm{diag}(\sigma_{Y,1},\dots,\sigma_{Y,n}),\quad\Sigma\;=\;\bm{\sigma}\,\Psi\,\bm{\sigma},$
where $\Psi\in\mathbb{R}^{d\times d}$ is the correlation matrix for the $d$ risky assets, and let $\rho\in\mathbb{R}^{d\times n}$ be the cross-correlation matrix between these assets and the $n$-dimensional factor Brownian motion $\mathbf{W}^{Y}_{t}$. Also let $\mathbf{A}=\bigl{(}\bm{\alpha}_{1},\dots,\bm{\alpha}_{d}\bigr{)}^{\top}\in\mathbb{R}^{d\times n}$ collect the factor loadings $\bm{\alpha}_{i}\in\mathbb{R}^{n}$ for each asset $i$.
The costate $\lambda_{t}^{*}=V_{x}$, its derivative with respect to wealth $\partial_{x}\lambda_{t}^{*}=V_{xx}$, and its gradient ($n\times 1$ vector) with respect to the state vector $\partial_{\mathbf{Y}}\lambda_{t}^{*}=V_{x\mathbf{Y}}$ are given by:
$\lambda_{t}^{*}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=(X_{t}^{*})^{-\gamma}\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right),$
$\frac{\partial\lambda_{t}^{*}}{\partial x}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=-\gamma(X_{t}^{*})^{-\gamma-1}\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right),$
$\frac{\partial\lambda_{t}^{*}}{\partial\mathbf{Y}}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=(X_{t}^{*})^{-\gamma}\left(\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}_{t}\right)\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right).$
These yield the following scalar and vector ratios, which simplify the PMP optimality conditions:
$\frac{\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma},\qquad\frac{\partial_{\mathbf{Y}}\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma}\left(\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}_{t}\right).$
Substituting these into the vector form of the PMP first-order condition for the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{d}$, (cf. Eq. (2) derived in Section 2.2):
$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\Sigma^{-1}\left\{\lambda_{t}^{*}(\bm{\mu}(\mathbf{Y}_{t})-r\mathbf{1})+\bm{\sigma}\rho\bm{\sigma}_{Y}\left[\partial_{\mathbf{Y}}\lambda_{t}^{*}\right]\right\},$
where the risk premium vector is $\bm{\mu}(\mathbf{Y}_{t})-r\mathbf{1}=\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}$, we precisely recover the analytic solution obtained from the HJB framework:
$\bm{\pi}^{*}(t,\mathbf{Y})\;=\;\frac{1}{\gamma}\,\Sigma^{-1}\Bigl{[}\,\underbrace{\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}}_{\text{Myopic component (related to }\bm{\mu}-r\mathbf{1})}\;+\;\underbrace{\bm{\sigma}\,\rho\,\bm{\sigma}_{Y}\Bigl{(}\mathbf{B}(t)\;+\;\mathbf{C}(t)\,\mathbf{Y}\Bigr{)}}_{\text{Hedging component (related to }\partial_{\mathbf{Y}}\lambda_{t}^{*})}\Bigr{]}.$
Here, the first bracketed term corresponds to the myopic component, driven by the instantaneous expected excess returns $\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}$. The second bracketed term incorporates intertemporal hedging demands arising from the desire to hedge against unfavorable changes in the investment opportunity set, as captured by the factor sensitivities $\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}$ (proportional to the gradient $\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and the correlations encoded in $\rho$, $\Psi$, and $\Phi$ (implicitly via the ODEs).
To determine the time-dependent coefficients $\mathbf{B}(t)$ and $\mathbf{C}(t)$, we solve the backward matrix Riccati ODE system derived from substituting the value function ansatz and optimal portfolio into the HJB equation:
$\frac{d\mathbf{C}(t)}{dt}$ $=\kappa_{y}^{\top}\mathbf{C}(t)+\mathbf{C}(t)\kappa_{y}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}\mathbf{C}(t)$
$\quad-\frac{1-\gamma}{\gamma}\left[(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right.$
$\qquad\qquad\left.+\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})+\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right],$
$\frac{d\mathbf{B}(t)}{dt}$ $=\left[\kappa_{y}^{\top}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}-\frac{1-\gamma}{\gamma}\left(\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\right)\right]\mathbf{B}(t)$
$\quad-\left[\mathbf{C}(t)\kappa_{y}-\frac{1-\gamma}{\gamma}(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right]\bm{\theta}_{y}.$
subject to the terminal conditions $\mathbf{C}(T)=\mathbf{0}$ and $\mathbf{B}(T)=\mathbf{0}$. The scalar function $g(t)$ satisfies its own ODE, which depends on $\mathbf{B}(t)$ and $\mathbf{C}(t)$, but is not needed for determining the optimal portfolio.
Solving these ODEs backward in time from $T$ to $t$ and then substituting the resulting $\mathbf{B}(t)$ and $\mathbf{C}(t)$ into $\bm{\pi}^{*}(t,\mathbf{Y})$ gives the closed-form optimal strategy, explicitly capturing both myopic and intertemporal hedging components in a multi-asset, multi-factor setting.
Appendix C Proof Sketch for Proposition 3.1
We provide a proof sketch for Proposition 3.1 focusing on the affine/quadratic structure considered by Liu (2007) (Liu07) with CRRA utility $U(w)=w^{1-\gamma}/(1-\gamma)$ (or $\ln w$). The argument relies on the specific form of the value function and the relationship between function closeness and coefficient closeness for polynomials.
Assume the true value function $V^{*}$ and an approximation $\tilde{V}$ both share the exponential-quadratic structure on a compact domain $\mathcal{D}=[t_{0},T]\times[w_{min},w_{max}]\times\mathcal{Y}$ (where $\mathcal{Y}$ is a compact subset of $\mathbb{R}^{k}$):
$V^{*}(t,w,\mathbf{y})$ $=U(w)\cdot\Phi^{*}(\mathbf{y},\tau)$
$\tilde{V}(t,w,\mathbf{y})$ $=U(w)\cdot\tilde{\Phi}(\mathbf{y},\tau)$
where $\tau=T-t$ and the state-dependent part has the form (adjusting notation slightly from Liu for clarity):
$\Phi^{*}(\mathbf{y},\tau)$ $=\exp\left[c^{*}(\tau)+d^{*}(\tau)^{\top}\mathbf{y}+\frac{1}{2}\mathbf{y}^{\top}Q^{*}(\tau)\mathbf{y}\right]$
$\tilde{\Phi}(\mathbf{y},\tau)$ $=\exp\left[\tilde{c}(\tau)+\tilde{d}(\tau)^{\top}\mathbf{y}+\frac{1}{2}\mathbf{y}^{\top}\tilde{Q}(\tau)\mathbf{y}\right]$
Here, $c,\tilde{c}$ are scalars, $d,\tilde{d}\in\mathbb{R}^{k}$, and $Q,\tilde{Q}$ are symmetric $k\times k$ matrices, all depending smoothly on $\tau$ and the underlying model parameters.
The condition $\sup_{\mathcal{D}}|V^{*}-\tilde{V}|\to 0$ implies $\sup_{\tau,\mathbf{y}\in\mathcal{Y}}|\Phi^{*}(\mathbf{y},\tau)-\tilde{\Phi}(\mathbf{y},\tau)|\to 0$, since $U(w)$ is bounded and bounded away from zero on $[w_{min},w_{max}]$. Because $\Phi^{*}$ and $\tilde{\Phi}$ are positive and continuous, this uniform convergence implies uniform convergence of their logarithms:
$\sup_{\tau,\mathbf{y}\in\mathcal{Y}}|\ln\Phi^{*}(\mathbf{y},\tau)-\ln\tilde{\Phi}(\mathbf{y},\tau)|\to 0$
This means the quadratic polynomials in the exponents are uniformly close:
$\sup_{\tau,\mathbf{y}\in\mathcal{Y}}|(c^{*}+d^{*\top}\mathbf{y}+\frac{1}{2}\mathbf{y}^{\top}Q^{*}\mathbf{y})-(\tilde{c}+\tilde{d}^{\top}\mathbf{y}+\frac{1}{2}\mathbf{y}^{\top}\tilde{Q}\mathbf{y})|\to 0$
It is a standard result that uniform convergence of polynomials on a compact set implies the convergence of their corresponding coefficients. Therefore, we must have uniform convergence of the coefficients with respect to $\tau\in[0,T-t_{0}]$:
$\sup_{\tau}|c^{*}(\tau)-\tilde{c}(\tau)|\to 0,\quad\sup_{\tau}\|d^{*}(\tau)-\tilde{d}(\tau)\|\to 0,\quad\sup_{\tau}\|Q^{*}(\tau)-\tilde{Q}(\tau)\|\to 0.$
Let $\varepsilon_{c},\varepsilon_{d},\varepsilon_{Q}$ denote these suprema, all approaching zero. Let also $\varepsilon_{\Phi}=\sup|\Phi^{*}-\tilde{\Phi}|\to 0$.
Now, consider the spatial derivatives.
  • w-derivatives: $V_{w}^{*}=U^{\prime}(w)\Phi^{*}$ and $\tilde{V}_{w}=U^{\prime}(w)\tilde{\Phi}$.
    $|V_{w}^{*}-\tilde{V}_{w}|=|U^{\prime}(w)||\Phi^{*}-\tilde{\Phi}|$
    Since $|U^{\prime}(w)|$ is bounded on $[w_{min},w_{max}]$, $\sup_{\mathcal{D}}|V_{w}^{*}-\tilde{V}_{w}|\leq(\sup|U^{\prime}|)\varepsilon_{\Phi}\to 0$. Similar logic applies to $V_{ww}$ using $U^{\prime\prime}(w)$.
  • y-gradient: $\nabla_{\mathbf{y}}V^{*}=U(w)\nabla_{\mathbf{y}}\Phi^{*}$ and ~$\tilde{\nabla}_{\mathbf{y}}\tilde{V}=U(w)\nabla_{\mathbf{y}}\tilde{\Phi}$. We have $\nabla_{\mathbf{y}}\Phi^{*}=\Phi^{*}(d^{*}+Q^{*}\mathbf{y})$.
    $\nabla_{\mathbf{y}}\Phi^{*}-\nabla_{\mathbf{y}}\tilde{\Phi}=(\Phi^{*}-\tilde{\Phi})(d^{*}+Q^{*}\mathbf{y})+\tilde{\Phi}[(d^{*}-\tilde{d})+(Q^{*}-\tilde{Q})\mathbf{y}]$
    Taking norms and suprema over the compact domain $\mathcal{D}$ (where $\mathbf{y}$, $\Phi^{*}$, $\tilde{\Phi}$, $d^{*}$, $Q^{*}$ are bounded):
    $\sup\|\nabla_{\mathbf{y}}\Phi^{*}-\nabla_{\mathbf{y}}\tilde{\Phi}\|\leq\varepsilon_{\Phi}\cdot M_{1}+M_{2}[\varepsilon_{d}+\varepsilon_{Q}M_{y}]\to 0,$
    where $M_{1},M_{2},M_{y}$ are finite bounds. Thus, $\sup_{\mathcal{D}}\|\nabla_{\mathbf{y}}V^{*}-\nabla_{\mathbf{y}}\tilde{V}\|\leq(\sup|U|)(\sup\|\nabla\Phi^{*}-\nabla\tilde{\Phi}\|)\to 0$.
  • yy-Hessian: $\nabla^{2}_{\mathbf{y}}V^{*}=U(w)\nabla^{2}_{\mathbf{y}}\Phi^{*}$.
    $\nabla^{2}_{\mathbf{y}}\Phi^{*}=\nabla_{\mathbf{y}}\Phi^{*}(d^{*}+Q^{*}\mathbf{y})^{\top}+\Phi^{*}Q^{*}$
    The difference $\nabla^{2}_{\mathbf{y}}\Phi^{*}-\nabla^{2}_{\mathbf{y}}\tilde{\Phi}$ involves terms containing $(\Phi^{*}-\tilde{\Phi})$, $(d^{*}-\tilde{d})$, and $(Q^{*}-\tilde{Q})$. Since these differences converge uniformly to zero and other terms are bounded, the uniform convergence $\sup_{\mathcal{D}}\|\nabla^{2}_{\mathbf{y}}V^{*}-\nabla^{2}_{\mathbf{y}}\tilde{V}\|\to 0$ follows.
  • wy-cross-derivatives: $\nabla_{\mathbf{y}}V_{w}^{*}=U^{\prime}(w)\nabla_{\mathbf{y}}\Phi^{*}$. Uniform convergence follows from the convergence of $\nabla_{\mathbf{y}}\Phi^{*}$ and boundedness of $U^{\prime}(w)$.
This completes the sketch, demonstrating that uniform value closeness implies uniform derivative closeness when the specific structure is shared.

### v2

@@
 Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality.
-Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond a small number of state variables (often limited to  3-6 in prior numerical studies).
+Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond few state variables ($\sim$3-6 limit in prior studies).
 To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework.
-PG-DPO leverages Pontryagin’s Maximum Principle to directly guide neural network policies via backpropagation-through-time, naturally incorporating exogenous state processes without requiring dense state grids.
-Crucially, our computationally efficient “Two-Stage” variant exploits rapidly stabilizing costate estimates derived from BPTT, converting them into near-optimal closed-form Pontryagin controls after only a short warm-up, significantly reducing training overhead.
-This enables a breakthrough in scalability: numerical experiments demonstrate that PG-DPO successfully tackles problems with dimensions previously considered far out of reach, optimizing portfolios with up to 50 assets and 10 state variables.
+PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to guide neural network policies, handling exogenous states without dense grids.
+This PMP-guided approach holds potential for a broad class of sufficiently regular continuous-time control problems.
+Crucially, our computationally efficient “Two-Stage” variant exploits rapidly stabilizing BPTT costate estimates, converting them into near-optimal Pontryagin controls after only a short warm-up, significantly reducing training overhead.
+This enables a breakthrough in scalability: numerical experiments show PG-DPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables).
 The framework delivers near-optimal policies, offering a practical and powerful alternative for high-dimensional continuous-time portfolio choice.
@@
 While DP or Hamilton–Jacobi–Bellman (HJB) PDE solvers provide a rigorous framework for low-dimensional cases (Bertsekas & Tsitsiklis, , 1995), they become intractable as the number of assets or state dimensions increases, with prior influential studies often limited to analyzing problems involving up to approximately 3-5 state variables (Campbell & Viceira, , 1999; Lynch, , 2001; Balduzzi & Lynch, , 1999) or specific frameworks handling around 6 state variables (Garlappi & Skoulakis, , 2010), often necessitating simulation-based approaches for tractability (Brandt et al., , 2005).
-A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong & Zhou, , 2012).
-Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint).
+A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong & Zhou, , 2012) Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint).
 Notably, adding one more state dimension does not produce an exponential explosion in computational requirements, in contrast to DP (Bellman, , 1966; Campbell et al., , 2003).
@@
 However, applications such as the Merton HJB have revealed limitations in accuracy or efficiency for certain PDE solvers like Deep BSDE variants (Nguwi et al., , 2024).
-Our approach, distinct from these direct HJB/PDE solvers, focuses on using Pontryagin’s principle, readily handling finite horizons.
+Our PMP-based PG-DPO framework offers a distinct path, naturally handling finite horizons and focusing on the optimality conditions derived from PMP, which are applicable to a wide range of stochastic control problems possessing sufficient regularity.
 One notable strategy related to the use of adjoint methods in optimal control is to directly parameterize the policy (i.e., consumption and investment controls) with neural networks, then apply backpropagation-through-time (BPTT) to enforce Pontryagin’s optimality conditions (Huh, , 2024).
@@
 We show how a single forward pass in simulation, coupled with a backward pass via automatic differentiation (BPTT), yields unbiased estimates of the costate processes that PMP requires, guiding the policy towards optimality without the drawbacks of DP.
+This approach fundamentally relies on the existence and accurate estimation of the costate processes dictated by PMP, applicable whenever the control problem satisfies appropriate regularity conditions allowing for a well-defined Hamiltonian system.
 Furthermore, we introduce a “Two-Stage” PG-DPO variant that exploits the observation that costate estimates often stabilize much faster than the full policy network.
@@
 This efficiency enables the solution of problems with dozens of assets and state variables (e.g., n=50, k=10 in our experiments), a scale considerably larger than those typically addressed by the numerical methods cited previously.
+This demonstrated scalability on a benchmarkable problem highlights the potential of leveraging PMP structure via BPTT estimates for tackling other complex, high-dimensional problems that may exhibit similar amenability to costate estimation under suitable regularity conditions.
 The main contributions of this paper can be summarized as follows:
-  • We develop and implement the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, leveraging PMP and deep learning, for solving general multi-asset, multi-factor continuous-time portfolio optimization problems with state-dependent dynamics and finite horizons.
-  • We introduce a computationally efficient “Two-Stage” PG-DPO variant that utilizes rapidly stabilized costate estimates from BPTT directly within the PMP first-order conditions, significantly reducing training overhead while achieving high accuracy.
-  • We demonstrate through numerical experiments that our framework, particularly the Two-Stage variant, successfully scales to high-dimensional problems involving dozens of assets and state variables (up to n=50, k=10).
-This substantially exceeds the dimensionality manageable by traditional DP methods and tackled in many prior numerical studies focusing on problems with typically fewer than  5-6 state variables, thus effectively mitigating the curse of dimensionality.
+  • We introduce Pontryagin-Guided Direct Policy Optimization (PG-DPO), featuring a highly efficient “Two-Stage” variant that uses PMP structure and BPTT-derived costates to solve complex continuous-time control problems.
+  • We demonstrate through numerical experiments the framework’s ability to break the dimensional barrier, successfully tackling portfolio optimization problems with dozens of assets and state factors (up to n=50, k=10).
 In the remainder of this paper, we first present the multi-asset Merton problem with exogenous states in Section 2.
-We focus on the affine multi-factor OU model, detailed in Section 2.3.2, for developing the methodology and conducting numerical experiments due to the availability of a benchmark solution.
 Section 3 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm and its Two-Stage extension.
-Numerical experiments presented in Section 4 confirm the framework’s scalability and near-optimal performance for the affine model.
-Finally, Section 5 concludes and discusses potential future directions, including applications to non-affine models or problems with portfolio constraints.
+Numerical experiments presented in Section 4 confirm the framework’s scalability and near-optimal performance within this testbed, illustrating its potential for high-dimensional applications.
+Finally, Section 5 concludes and discusses potential future directions.
 2 Multi-Asset Continuous-Time Portfolio Problem with State Variables
-In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1971), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously.
+In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1971; Kim & Omberg, , 1996; Liu, , 2007), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously.
 Each risky asset features its own drift and volatility, leading to an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}$.
@@
 2.1 Problem Formulation with State Variables and Neural Network Controls
-Let $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ be a vector of state variables that affect the distribution of asset returns.
-For instance, $\mathbf{Y}_{t}$ can include interest rates, macroeconomic indices, volatility factors, and so forth.
-We assume each component $Y_{i,t}$ follows an Ito diffusion
+Let $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ be a vector of economic state variables that influence the distribution of asset returns.
+For instance, $\mathbf{Y}_{t}$ could include interest rates, macroeconomic indices, volatility factors, etc.
+We assume each component $Y_{i,t}$ follows an Ito diffusion:
 $dY_{i,t}=\mu_{Y,i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{Y,i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dW^{Y}_{i,t},\quad i=1,\dots,k,$
@@
 where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})$ is the $k\times 1$ drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})$ is often assumed to be a $k\times k$ diagonal matrix of diffusion coefficients, and $\mathbf{W}^{Y}_{t}$ is a $k\times 1$ vector of potentially correlated Brownian motions driving the state variables.
-We consider a financial market with $n$ risky assets and one risk-free asset. For $i=1,\dots,n$, the price $S_{i,t}$ of the $i$-th risky asset follows
+We consider a financial market with $n$ risky assets and one risk-free asset.
+The risk-free asset accumulates wealth at an instantaneous rate $r(t,\mathbf{Y}_{t})$, which may depend on time and the current state.
+For $i=1,\dots,n$, the price $S_{i,t}$ of the $i$-th risky asset follows
 $dS_{i,t}=S_{i,t}\bigl{(}\mu_{i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dW^{X}_{i,t}\bigr{)}.$
@@
 The $n$ scalar Brownian motions $W^{X}_{i,t}$ may be correlated amongst themselves ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt$) and with the state variable processes ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt$).
-In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})$ be the $n\times 1$ vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})$ be the $n\times n$ diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}$ be the $n\times 1$ vector of Brownian motions.
+In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})$ be the $n\times 1$ vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})$ be the $n\times n$ diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}$ be the $n\times 1$ vector of Brownian motions for the risky assets.
 The risky asset dynamics can be written compactly as
-$d\mathbf{S}_{t}=\bm{\mu}(t,\mathbf{Y}_{t})\,\mathbf{S}_{t}\,dt+\bm{\sigma}(t,\mathbf{Y}_{t})\,\mathbf{S}_{t}\,d\mathbf{W}^{X}_{t}.$
-Let $X_{t}$ denote the investor’s total wealth.
-At each time $t$, the investor chooses a consumption rate $C_{t}\geq 0$ and allocates fractions of wealth to the $n$ risky assets, represented by the vector $\bm{\pi}_{t}=(\pi_{1,t},\,\dots,\,\pi_{n,t})^{\top}\in\mathbb{R}^{n}.$ The remaining fraction $1-\bm{\pi}_{t}^{\top}\mathbf{1}$ is allocated to the risk-free asset, which earns interest at rate $r\bigl{(}t,\mathbf{Y}_{t}\bigr{)}$.
-The wealth process then evolves according to the stochastic differential equation (SDE):
-$dX_{t}\;=\;\Bigl{[}\,X_{t}\,\Bigl{(}\bigl{(}1-\bm{\pi}_{t}^{\top}\mathbf{1}\bigr{)}\,r\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\;+\;\bm{\pi}_{t}^{\top}\,\bm{\mu}(t,\mathbf{Y}_{t})\Bigr{)}\;-\;C_{t}\Bigr{]}\,dt\;+\;X_{t}\,\bm{\pi}_{t}^{\top}\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$
-We require the controls $(\bm{\pi}_{t},C_{t})$ to be admissible, ensuring that the wealth remains strictly positive, $X_{t}>0$, for all $t\in[0,T]$.
-This condition is essential for the utility function to be well-defined and is typically maintained in simulations (e.g., by clamping wealth at a small positive threshold, see Section 4.1.2).
+$d\mathbf{S}_{t}=\operatorname{diag}(\mathbf{S}_{t})\,\bm{\mu}(t,\mathbf{Y}_{t})\,dt+\operatorname{diag}(\mathbf{S}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$
+Throughout, we work on a complete probability space $\bigl{(}\Omega,\mathcal{F},\mathbb{P}\bigr{)}$ that carries the $(n+k)$-dimensional Brownian motion $\mathbf{W}_{t}$ and its natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$, augmented to satisfy the usual conditions (right-continuity and completeness).
+The investor allocates wealth $X_{t}$ between the risk-free asset and the $n$ risky assets.
+Let $\bm{\pi}_{t}$ be the $n\times 1$ vector representing the fraction of wealth invested in each risky asset at time $t$.
+The investor also chooses a consumption rate $C_{t}$.
+The dynamics of the investor’s wealth $X_{t}$ are governed by the following stochastic differential equation (SDE):
+$dX_{t}=\Bigl{[}\,X_{t}\Bigl{(}r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\Bigr{)}-C_{t}\Bigr{]}dt+X_{t}\,\bm{\pi}_{t}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$ (1)
+Here, $\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}$ represents the vector of instantaneous excess returns over the risk-free rate, and $\mathbf{1}$ is an $n\times 1$ vector of ones.
+We impose regularity conditions on the market parameters.
+Assumption 1 (Market-parameter regularity).
+Fix a finite horizon $T>0$.
+The coefficients defining the state process $\mathbf{Y}_{t}$ and influencing the wealth dynamics $X_{t}$ via Eq.
+(1), namely $\bm{\mu}_{Y},\ \bm{\sigma}_{Y},\ \bm{\mu},\ \bm{\sigma},\ r$, satisfy the following conditions.
+  1. (a)
+    Measurability and continuity. Each map is jointly Borel-measurable in $(t,\mathbf{y})$ and continuous in $\mathbf{y}$ for every fixed $t\in[0,T]$.
+  2. (b)
+    Global Lipschitz continuity. There exists $K_{\mathrm{Lip}}>0$ such that for all $t\in[0,T]$ and $\mathbf{y},\mathbf{y}^{\prime}\in\mathbb{R}^{k}$,
+    $\lVert\bm{\mu}_{Y}(t,\mathbf{y})-\bm{\mu}_{Y}(t,\mathbf{y}^{\prime})\rVert+\lVert\bm{\sigma}_{Y}(t,\mathbf{y})-\bm{\sigma}_{Y}(t,\mathbf{y}^{\prime})\rVert$
+    $+\lVert\bm{\mu}(t,\mathbf{y})-\bm{\mu}(t,\mathbf{y}^{\prime})\rVert+\lVert\bm{\sigma}(t,\mathbf{y})-\bm{\sigma}(t,\mathbf{y}^{\prime})\rVert\leq K_{\mathrm{Lip}}\lVert\mathbf{y}-\mathbf{y}^{\prime}\rVert.$
+  3. (c)
+    Linear growth. There exists $K_{\mathrm{LG}}>0$ such that for all $(t,\mathbf{y})\in[0,T]\times\mathbb{R}^{k}$,
+    $\lVert\bm{\mu}_{Y}(t,\mathbf{y})\rVert^{2}$ $+\lVert\bm{\sigma}_{Y}(t,\mathbf{y})\rVert^{2}+\lVert\bm{\mu}(t,\mathbf{y})\rVert^{2}+\lVert\bm{\sigma}(t,\mathbf{y})\rVert^{2}$
+    $\leq K_{\mathrm{LG}}^{2}\bigl{(}1+\lVert\mathbf{y}\rVert^{2}\bigr{)},\qquad|r(t,\mathbf{y})|\leq K_{\mathrm{LG}}\bigl{(}1+\lVert\mathbf{y}\rVert\bigr{)}.$
+Under Assumption 1, the SDE system for $(\mathbf{Y}_{t},X_{t})$ (where $X_{t}$ depends on the chosen controls $\bm{\pi}_{t},C_{t}$) admits a unique strong solution for sufficiently well-behaved controls; see Oksendal, (2013, Chs.
+5-6) for details on existence and uniqueness theorems for SDEs.
+We restrict our attention to control processes that are economically sensible and mathematically tractable.
+Assumption 2 (Admissible controls).
+A control pair $(\bm{\pi},C)$ is admissible if the following conditions hold:
+  1. (a)
+    Adaptedness. $(\bm{\pi}_{t},C_{t})_{0\leq t\leq T}$ is $\{\mathcal{F}_{t}\}$-progressively measurable, where $\{\mathcal{F}_{t}\}$ is the filtration generated by the underlying Brownian motions.
+  2. (b)
+    Integrability.
+[…]
+  3. (c)
+    Boundedness and proportional consumption. There exist constants $\Pi_{\max}>0$ and $\xi\in(0,1)$ such that
+[…]
+  4. (d)
+    Positive wealth. The corresponding wealth process $X_{t}$ solving (1) remains strictly positive almost surely:
+[…]
 For clarity on the noise structure, we define the combined $(n+k)$-dimensional Brownian vector $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$.
-The covariance structure of its increments $d\mathbf{W}_{t}$ is given by the block matrix:
-$\operatorname{Cov}\bigl{(}d\mathbf{W}_{t}\bigr{)}\;=\;\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
-where $\Psi\in\mathbb{R}^{n\times n}=\operatorname{Cov}(d\mathbf{W}^{X}_{t})/dt$, $\Phi\in\mathbb{R}^{k\times k}=\operatorname{Cov}(d\mathbf{W}^{Y}_{t})/dt$, and $\rho\in\mathbb{R}^{n\times k}=\operatorname{Cov}(d\mathbf{W}^{X}_{t},d\mathbf{W}^{Y}_{t})/dt$.
-The investor’s objective is to choose admissible control processes $(\bm{\pi},C)$ to maximize the expected discounted utility functional $J$:
+The covariance structure of its increments $d\mathbf{W}_{t}$ is given by the matrix $\bm{\Omega}$:
+$\operatorname{Cov}\bigl{(}d\mathbf{W}_{t}\bigr{)}=\bm{\Omega}\,dt=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
+where $\Psi=(\psi_{ij})\in\mathbb{R}^{n\times n}$ is the covariance matrix of $d\mathbf{W}^{X}_{t}$, $\Phi=(\phi_{ij})\in\mathbb{R}^{k\times k}$ is the covariance matrix of $d\mathbf{W}^{Y}_{t}$, and $\rho=(\rho_{ij})\in\mathbb{R}^{n\times k}$ is the cross-covariance matrix.
+Assumption 3 (Noise covariance).
+The block covariance matrix $\bm{\Omega}=\begin{pmatrix}\Psi&\rho\\[1.79997pt] \rho^{\top}&\Phi\end{pmatrix}$ is constant in time and uniformly positive definite: there exists $\lambda_{\Omega}>0$ such that $z^{\top}\bm{\Omega}z\geq\lambda_{\Omega}\lVert z\rVert^{2}$ for all $z\in\mathbb{R}^{n+k}$.
+(Consequently, $\Psi$ and $\Phi$ are individually positive definite and hence invertible.)
+The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ to maximize the expected discounted utility functional $J$:
 […]
-subject to the wealth dynamics.
+subject to the wealth dynamics (1) and admissibility constraints (Assumption 2).
 Here $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ weights the utility of terminal wealth (bequest motive).
-The utility function $U:(0,\infty)\to\mathbb{R}$ represents the investor’s preferences and is assumed to be strictly increasing, strictly concave, and satisfy the Inada conditions.
+The utility function $U:(0,\infty)\to\mathbb{R}$ represents the investor’s preferences.
 A widely used specification is the Constant Relative Risk Aversion (CRRA) form:
+$U(x)=\begin{cases}\displaystyle\frac{x^{1-\gamma}}{1-\gamma},&\gamma>0,\gamma\neq 1,\\[6.0pt] \ln(x),&\gamma=1,\end{cases}$
+with $\gamma>0$ denoting the coefficient of relative risk aversion (see Pennacchi, , 2008).
+Assumption 2 (iii) bounds consumption by $C_{t}\leq\xi X_{t}$ with $\xi\in(0,1)$, while Assumption 4 (b) gives a $C^{2}$, strictly concave utility $U$ satisfying the Inada limits.
+Because $\delta>0$ and $X_{t}>0$ a.s.
+(Assumption 2 (iv)), we have
+$\mathbb{E}\!\bigl{[}\,e^{-\delta t}\,U(C_{t})\bigr{]}\leq e^{-\delta t}\,U\!\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}<\infty,\quad\forall t\in[0,T],$
+and similarly $\mathbb{E}[U(X_{T})]<\infty$. Hence the functional $J(\bm{\pi},C)$ in (2) is finite for every admissible control pair, so the value function $V$ is well defined.
+The value function $V(t,x,\mathbf{y})$ associated with this problem represents the maximum expected utility achievable starting from wealth $x$ and state $\mathbf{y}$ at time $t$. Formally,
 […]
-with $\gamma>0$ denoting the coefficient of relative risk aversion.
-In our numerical approach, we parameterize the control policies using neural networks:
-$\bm{\pi}_{t}\;=\;\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\quad C_{t}\;=\;C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},$
-where $\theta$ and $\phi$ represent the trainable parameters of the respective neural networks.
-2.2 Pontryagin’s Maximum Principle with a State-Dependent Market and Costate for $\mathbf{Y}_{t}$
-We now consider the continuous-time portfolio problem from Section 2.1 under a market whose parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$.
-The Pontryagin Maximum Principle (PMP) provides necessary conditions for optimality via a system of forward-backward stochastic differential equations (FBSDEs).
-The state of the system consists of the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$.
-We introduce the costate processes associated with these states: $\lambda_{t}\approx\frac{\partial V}{\partial x}(t,X_{t},\mathbf{Y}_{t})$ (scalar costate for wealth) and $\bm{\eta}_{t}\approx\frac{\partial V}{\partial\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})$ (a $k$-dimensional costate vector for the state variables), where $V$ is the value function of the problem.
-The dynamics of the state and costate processes are described by the following FBSDE system.
-Under an optimal control policy $\bigl{(}\bm{\pi}_{t}^{*},\,C_{t}^{*}\bigr{)}$, the state processes $(X_{t}^{*},\mathbf{Y}_{t}^{*})$ evolve according to the forward SDEs:
-$dX_{t}^{*}$ $=\;\Bigl{[}X_{t}^{*}\,\Bigl{(}\,\bigl{(}1-\bm{\pi}_{t}^{*\top}\mathbf{1}\bigr{)}\,r\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\;+\;\bm{\pi}_{t}^{*\top}\,\bm{\mu}\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\Bigr{)}\;-\;C_{t}^{*}\Bigr{]}\,dt$
-$\quad+\;X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\,\bm{\sigma}(t,\mathbf{Y}_{t}^{*})\,d\mathbf{W}_{t}^{X},$ (1)
-$d\mathbf{Y}_{t}^{*}$ $=\;\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\,dt\;+\;\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}^{*}\bigr{)}\,d\mathbf{W}_{t}^{Y},\quad X_{0}^{*}\;=\;x_{0}>0,\quad\mathbf{Y}_{0}^{*}\;=\;y_{0}\in\mathbb{R}^{k}.$
-Here $\bm{\sigma}(t,\mathbf{Y}_{t}^{*})$ is the $n\times n$ (diagonal) matrix of volatilities, and $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t}^{*})$ is the $k\times k$ diffusion matrix for $\mathbf{Y}_{t}^{*}$.
-The Brownian motions $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ and $\mathbf{W}_{t}^{Y}\in\mathbb{R}^{k}$ have the covariance structure $\operatorname{Cov}(d\mathbf{W}_{t})=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}dt.$
-Simultaneously, the costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy the backward SDEs (BSDEs):
-$-d\lambda_{t}^{*}$ $=\;\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t}^{*},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*}\bigr{)}\,dt\;-\;\mathbf{Z}_{t}^{\lambda X}\,d\mathbf{W}_{t}^{X}\;-\;\mathbf{Z}_{t}^{\lambda Y}\,d\mathbf{W}_{t}^{Y},$ (2)
-$-d\bm{\eta}_{t}^{*}$ $=\;\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t}^{*},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*}\bigr{)}\,dt\;-\;\mathbf{Z}_{t}^{\eta X}\,d\mathbf{W}_{t}^{X}\;-\;\mathbf{Z}_{t}^{\eta Y}\,d\mathbf{W}_{t}^{Y}.$ (3)
-Here, $\mathcal{H}_{x}$ and $\mathcal{H}_{\mathbf{y}}$ denote the partial derivatives of the Hamiltonian function $\mathcal{H}$ with respect to $x$ and $\mathbf{y}$, evaluated at the optimal state and controls.
-The terms $\mathbf{Z}_{t}^{\lambda X},\mathbf{Z}_{t}^{\lambda Y},\mathbf{Z}_{t}^{\eta X},\mathbf{Z}_{t}^{\eta Y}$ are adapted processes representing the diffusion coefficients of the costates, related to second derivatives of the value function and the system’s volatility.
-The terminal conditions are $\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)}$ and $\bm{\eta}_{T}^{*}=\mathbf{0}.$
-The Hamiltonian $\mathcal{H}$ associated with this control problem, whose partial derivatives appear in the BSDE drift terms, is given by:
-$\begin{split}\mathcal{H}(t,x,\mathbf{y},\bm{\pi},C,\lambda,\bm{\eta})=&e^{-\delta t}\,U(C)+\lambda\Bigl{[}x\bigl{(}(1-\bm{\pi}^{\top}\mathbf{1})r(t,\mathbf{y})+\bm{\pi}^{\top}\bm{\mu}(t,\mathbf{y})\bigr{)}-C\Bigr{]}\\ &+\bm{\eta}^{\top}\bm{\mu}_{Y}(t,\mathbf{y}).\end{split}$
-Optimality requires that the controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ maximize the Hamiltonian at each time $t$.
-The first-order condition (FOC) with respect to consumption $C_{t}$ yields $e^{-\delta t}\,U^{\prime}(C_{t}^{*})-\lambda_{t}^{*}=0$, implying $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*}).$ The FOC with respect to the portfolio weights $\bm{\pi}_{t}$ leads to a condition balancing expected excess returns against risk, involving the costate $\lambda_{t}^{*}$ and its sensitivity $\mathbf{Z}_{t}^{\lambda X}$.
-A standard form for this condition is:
+where the supremum is taken over all admissible control processes on $[t,T]$, and $\mathbb{E}_{t,x,\mathbf{y}}[\cdot]=\mathbb{E}[\cdot|X_{t}=x,\mathbf{Y}_{t}=\mathbf{y}]$ denotes the conditional expectation.
+Additional assumptions are needed for the well-posedness of the problem and regularity of the value function.
+Assumption 4 (Uniform non-degeneracy and utility).
+  1. (a)
+    Uniform non-degeneracy of risky-asset volatility.
+Let $a(t,\mathbf{y}):=\bm{\sigma}(t,\mathbf{y})\,\Psi\,\bm{\sigma}(t,\mathbf{y})^{\top}$ be the instantaneous covariance matrix of the risky asset returns (Note: original used $\bm{\sigma}\bm{\sigma}^{\top}$, assuming $\Psi=I$; using $\bm{\sigma}\Psi\bm{\sigma}^{\top}$ is more general if $\mathbf{W}^{X}$ components are correlated and $\bm{\sigma}$ is diagonal).
+We assume there exists $\lambda>0$ such that
+    $\mathbf{z}^{\top}a(t,\mathbf{y})\,\mathbf{z}\;\geq\;\lambda\,\lVert\mathbf{z}\rVert^{2},\quad\forall(t,\mathbf{y},\mathbf{z})\in[0,T]\times\mathbb{R}^{k}\times\mathbb{R}^{n}.$
+    (Equivalently, the covariance matrix of risky returns $a(t,\mathbf{y})$ is uniformly positive definite, ensuring invertibility, which is crucial for solving the portfolio problem.)
+  2. (b)
+    Utility specification.
+The utility function $U:(0,\infty)\to\mathbb{R}$ satisfies $U\in C^{2}(0,\infty)$, $U^{\prime}>0,\;U^{\prime\prime}<0$ (strictly increasing and concave), and the Inada conditions $\lim_{c\downarrow 0}U^{\prime}(c)=\infty,\;\lim_{c\uparrow\infty}U^{\prime}(c)=0.$
+  3. (c)
+    Discount and bequest parameters. The subjective discount rate satisfies $\delta>0$ and the terminal weight satisfies $\kappa\geq 0$.
+Assumption 4 (a) guarantees the diffusion related to risky assets is non-degenerate, a key ingredient for the regularity of the associated Hamilton-Jacobi-Bellman (HJB) PDE.
+Parts (b)-(c) ensure the objective functional is well-defined and satisfies standard preference assumptions.
+Combined with the coefficient regularity in Assumption 1, these conditions typically yield the existence, uniqueness, and sufficient smoothness (e.g., $V\in C^{1,2,2}$) of the solution to the HJB equation, allowing for verification theorems to be applied; see Oksendal, (2013, Chs.
+6-7) and related literature on optimal portfolio choice.
+In our numerical approach, we parameterize the control policies using neural networks (NNs).
+The portfolio allocation $\bm{\pi}_{t}$ and consumption rate $C_{t}$ are represented as functions of the current time $t$, wealth $X_{t}$, and state $\mathbf{Y}_{t}$, determined by the NN parameters $\theta$ and $\phi$:
+$\bm{\pi}_{t}=\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\quad C_{t}=C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)}.$
+These parameters $\theta$ and $\phi$ are then optimized to maximize the objective $J$.
+2.2 Optimality Conditions via Pontryagin’s Maximum Principle
+We now apply Pontryagin’s Maximum Principle (PMP) to the continuous-time portfolio problem formulated in Section 2.1, where market parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$.
+PMP provides necessary conditions for optimality through a system of forward-backward stochastic differential equations (FBSDEs), which couple the dynamics of the state variables with those of associated costate variables.
+The system state includes the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$.
+The corresponding costate processes represent the sensitivity of the value function $V$ to changes in the state: $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ is the scalar costate for wealth (marginal value of wealth), and $\bm{\eta}_{t}=V_{\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})$ is the $k$-dimensional costate vector for the exogenous state variables (marginal value of state variables).
+Under an optimal control policy $\bigl{(}\bm{\pi}_{t}^{*},\,C_{t}^{*}\bigr{)}$, the evolution of the system state, comprising the wealth process $X_{t}^{*}$ and the exogenous variables $\mathbf{Y}_{t}$, is described by the following forward stochastic differential equations:
+$dX_{t}^{*}$ $=\Bigl{[}X_{t}^{*}\,\Bigl{(}r\bigl{(}t,\mathbf{Y}_{t}\bigr{)}+\bm{\pi}_{t}^{*\top}\,\bigl{(}\bm{\mu}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\Bigr{)}-C_{t}^{*}\Bigr{]}\,dt$
+$\quad+X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{X},$ (3)
+$d\mathbf{Y}_{t}$ $=\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,d\mathbf{W}_{t}^{Y},$ (4)
+with initial conditions $X_{0}^{*}=x_{0}>0$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$.
+Recall the definitions of the volatility matrix $\bm{\sigma}(t,\mathbf{Y}_{t})$ and the covariance structure $\bm{\Omega}$ of the combined Brownian motion $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$.
+Complementing these forward dynamics, the optimal costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy a system of backward stochastic differential equations (BSDEs).
+These equations are driven by derivatives of the Hamiltonian $\mathcal{H}$ associated with the control problem:
+$-d\lambda_{t}^{*}$ $=\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}\bigr{)}\,dt-\mathbf{Z}_{t}^{\lambda X}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\lambda Y}\,d\mathbf{W}_{t}^{Y},$ (5)
+$-d\bm{\eta}_{t}^{*}$ $=\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}\bigr{)}\,dt-\mathbf{Z}_{t}^{\eta X}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\eta Y}\,d\mathbf{W}_{t}^{Y}.$ (6)
+Here, $\mathcal{H}_{x}=\partial\mathcal{H}/\partial x$ and $\mathcal{H}_{\mathbf{y}}=\nabla_{\mathbf{y}}\mathcal{H}$ represent the partial derivatives of the Hamiltonian.
+The terms $\mathbf{Z}_{t}^{\lambda X}\in\mathbb{R}^{1\times n}$, $\mathbf{Z}_{t}^{\lambda Y}\in\mathbb{R}^{1\times k}$, $\mathbf{Z}_{t}^{\eta X}\in\mathbb{R}^{k\times n}$, and $\mathbf{Z}_{t}^{\eta Y}\in\mathbb{R}^{k\times k}$ are adapted processes arising from the martingale representation theorem applied to the costates, collectively denoted by $\mathbf{Z}_{t}$.
+The derivatives $\mathcal{H}_{x},\mathcal{H}_{\mathbf{y}}$ generally depend on the state, costate, controls, and the $\mathbf{Z}_{t}$ terms themselves, reflecting the system’s interconnectedness through the value function’s second-order derivatives.
+The terminal conditions are determined by the objective functional: $\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)}$ and $\bm{\eta}_{T}^{*}=\mathbf{0}$.
+The application of PMP involves maximizing a Hamiltonian function.
+To properly account for the dependence of the wealth diffusion on the control $\bm{\pi}$, the Hamiltonian should incorporate the interaction between the costate dynamics (represented by $\mathbf{Z}$ terms) and the diffusion coefficients.
+Following formulations in stochastic control theory (e.g., Yong & Zhou, , 2012), the appropriate Hamiltonian $\mathcal{H}$ can be expressed as:
+$\begin{split}&\mathcal{H}\bigl{(}t,x,\mathbf{y},\bm{\pi},C,\lambda,\bm{\eta},\mathbf{Z}^{\lambda X},\mathbf{Z}^{\lambda Y},\mathbf{Z}^{\eta X},\mathbf{Z}^{\eta Y}\bigr{)}\\ &\quad=e^{-\delta t}\,U(C)+\lambda\Bigl{[}x\bigl{(}r(t,\mathbf{y})+\bm{\pi}^{\top}(\bm{\mu}(t,\mathbf{y})-r(t,\mathbf{y})\mathbf{1})\bigr{)}-C\Bigr{]}+\bm{\eta}^{\top}\,\bm{\mu}_{Y}(t,\mathbf{y})\\ &\quad\quad+\mathbf{Z}^{\lambda X}\bigl{(}x\,\bm{\sigma}(t,\mathbf{y})^{\top}\bm{\pi}\bigr{)}+\text{Tr}\bigl{(}\mathbf{Z}^{\eta Y\top}\bm{\sigma}_{Y}(t,\mathbf{y})\bigr{)}+\dots\end{split}$ (7)
+For clarity in deriving the FOC for $\bm{\pi}$, the crucial term involving $\bm{\pi}$ beyond the drift component is the one coupling $\mathbf{Z}^{\lambda X}$ with the wealth diffusion $x\,\bm{\pi}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})$ originating from $d\mathbf{W}_{t}^{X}$.
+We focus on these terms for the maximization step.
+To ensure the well-posedness of this FBSDE system and the applicability of PMP, additional smoothness conditions on the model coefficients are typically required.
+Assumption 5 (Additional regularity for PMP).
+In addition to Assumptions 1, 2, and 4, we assume:
+  1. (a)
+    $C^{1}$-regular coefficients w.r.t.
+$\mathbf{y}$.
+All coefficient functions $\bm{\mu}_{Y},\,\bm{\sigma}_{Y},\,\bm{\mu},\,\bm{\sigma}\,$ $r$ are continuously differentiable with respect to the state vector $\mathbf{y}$.
+Furthermore, their partial derivatives ($\partial_{\mathbf{y}}f$) are globally Lipschitz continuous in $\mathbf{y}$ and satisfy linear growth conditions.
+This ensures that the drift terms in the costate BSDEs, involving derivatives like $\mathcal{H}_{\mathbf{y}}$, satisfy the required regularity conditions for the existence and uniqueness of solutions (cf.
+Yong & Zhou, , 2012).
+The core requirement of Pontryagin’s Maximum Principle is that the optimal controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ must maximize the Hamiltonian $\mathcal{H}$ (7) pointwise in time, almost surely, given the optimal state and costate processes $(X_{t}^{*},\mathbf{Y}_{t},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*})$.
+Assuming interior solutions allows us to derive the first-order conditions (FOCs) for optimality.
+Maximization with respect to the consumption rate $C_{t}$ yields the standard condition:
+$e^{-\delta t}\,U^{\prime}(C_{t}^{*})=\lambda_{t}^{*}.$ (8)
+This classic result links optimal consumption $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$ inversely to the marginal value of wealth $\lambda_{t}^{*}$.
+Maximization with respect to the portfolio weights $\bm{\pi}_{t}$ requires differentiating $\mathcal{H}$ with respect to $\bm{\pi}$, including the term involving $\mathbf{Z}^{\lambda X}$.
+From (7), the relevant terms are $\lambda x\bm{\pi}^{\top}(\bm{\mu}-r\mathbf{1})$ and $\mathbf{Z}^{\lambda X}(x\,\bm{\sigma}^{\top}\bm{\pi})$.
+Setting the gradient $\nabla_{\bm{\pi}}\mathcal{H}=\mathbf{0}$ gives (assuming $X_{t}^{*}>0$):
+$\lambda_{t}^{*}X_{t}^{*}(\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1})+X_{t}^{*}\bm{\sigma}(t,\mathbf{Y}_{t})(\mathbf{Z}_{t}^{\lambda X*})^{\top}=\mathbf{0}.$
+Dividing by $X_{t}^{*}$, we obtain the FOC for the optimal portfolio:
 […]
-Here, $\mathbf{Z}_{t}^{\lambda X}$ is represented as a $1\times n$ row vector.
-Using Ito’s lemma to relate $\mathbf{Z}_{t}^{\lambda X}$ to the value function derivatives ($\lambda_{t}=V_{X}(t,X_{t},\mathbf{Y}_{t})$, so $\partial_{x}\lambda_{t}=V_{XX}$, $\partial_{\mathbf{Y}}\lambda_{t}=V_{X\mathbf{Y}}$), one can solve the FOC (4) for the optimal portfolio $\bm{\pi}_{t}^{*}$.
-The resulting expression, consistent with standard results in continuous-time finance, is:
-$\begin{split}\bm{\pi}_{t}^{*}=-\frac{\lambda_{t}^{*}}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}&\Sigma(t,\mathbf{Y}_{t}^{*})^{-1}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t}^{*})-r(t,\mathbf{Y}_{t}^{*})\mathbf{1}\bigr{)}\\ -\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}&\Sigma(t,\mathbf{Y}_{t}^{*})^{-1}\bm{\sigma}(t,\mathbf{Y}_{t}^{*})\rho\bm{\sigma}_{Y}(t,\mathbf{Y}_{t}^{*})^{\top}(\partial_{\mathbf{Y}}\lambda_{t}^{*})^{\top}.\end{split}$ (5)
-Here, $\Sigma(t,\mathbf{Y}_{t}^{*})=\bm{\sigma}(t,\mathbf{Y}_{t}^{*})\Psi\bm{\sigma}(t,\mathbf{Y}_{t}^{*})^{\top}$ is the instantaneous covariance matrix of asset returns (assuming $\bm{\sigma}$ is diagonal, this matches previous definition).
-The first term represents the myopic demand, driven by instantaneous expected excess return and risk aversion.
-The second term represents the intertemporal hedging demand, driven by the desire to hedge against unfavorable changes in the state variables $\mathbf{Y}_{t}$, depending on the sensitivity of marginal utility to state variables ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$), state variable volatility ($\bm{\sigma}_{Y}$), and correlations ($\rho,\Psi,\Phi$).
-In the special case $n=1,k=1$, we have $\Sigma=\sigma^{2}$, $\Psi=1$, $\rho$ is a scalar correlation, $\bm{\sigma}=\sigma$, $\bm{\sigma}_{Y}=\sigma_{Y}$. The formula (2) simplifies to:
-$\pi_{t}^{*}\;=\;-\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})\sigma^{2}}\left\{\lambda_{t}^{*}\bigl{(}\mu(t,Y_{t}^{*})-r(t,Y_{t}^{*})\bigr{)}+\sigma\rho\sigma_{Y}(\partial_{Y}\lambda_{t}^{*})\right\},$
-which clearly separates the myopic and hedging components.
+This first-order condition implicitly defines the optimal portfolio $\bm{\pi}_{t}^{*}$ through its dependence on the costate $\lambda_{t}^{*}$ and the martingale component $\mathbf{Z}_{t}^{\lambda X*}$.
+To obtain an explicit expression for $\bm{\pi}_{t}^{*}$, we need to relate the martingale component $\mathbf{Z}^{\lambda X*}$ to the derivatives of the value function.
+Applying Ito’s lemma to $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ and matching terms with the BSDE (5) establishes this relationship (the precise derivation often involves careful application of Ito’s formula for processes driven by correlated Brownian motions, see e.g., standard texts or appendices like Appendix A).
+This procedure yields an expression for $\mathbf{Z}^{\lambda X*}$ in terms of $V_{xx}$, $V_{x\mathbf{y}}$, the optimal controls $\bm{\pi}_{t}^{*}$, and the model parameters ($\bm{\sigma},\bm{\sigma}_{Y},\Psi,\rho$).
+Substituting this expression into the FOC (9) and solving the resulting algebraic equation for $\bm{\pi}_{t}^{*}$ leads to the well-known optimal portfolio formula:
+$\bm{\pi}_{t}^{*}$ $=-\frac{V_{x}}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}$ (10)
+$\qquad\qquad-\frac{V_{x\mathbf{y}}}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,\rho\,\bm{\sigma}_{Y}^{\!\top}(t,\mathbf{Y}_{t}).$
+Here, $\Sigma(t,\mathbf{Y}_{t})=\bm{\sigma}(t,\mathbf{Y}_{t})\Psi\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ represents the instantaneous covariance matrix of the risky asset returns.
+The value function derivatives $V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\partial_{\mathbf{y}}\lambda_{t}^{*}$ are evaluated along the optimal path $(t,X_{t}^{*},\mathbf{Y}_{t})$.
+The formula clearly separates the portfolio into two components: the myopic demand, which optimizes the immediate risk-return trade-off as if the investment opportunity set were constant, and the intertemporal hedging demand, which adjusts the portfolio to mitigate risks arising from future stochastic changes in the state variables $\mathbf{Y}_{t}$.
+The hedging demand’s magnitude and direction depend on the correlation $\rho$ between asset returns and state variable innovations, and crucially on $V_{x\mathbf{y}}$, which measures how the marginal value of wealth changes with the state variables.
+For illustration, in the simplified case with one risky asset ($n=1$) and one state variable ($k=1$), the optimal allocation (10) becomes:
+$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{V_{xy}}{X_{t}V_{xx}\sigma^{2}}\sigma\rho\sigma_{Y},$
+clearly showing the separation between the myopic component, proportional to the market price of risk, and the hedging component, proportional to $V_{xy}$ and the correlation $\rho$.
 2.3 Applications to Financial Models
@@
 3 A Gradient-Based Algorithm for Multi-Asset Policy Optimization
-In the preceding sections, we established the theoretical underpinnings of the continuous-time multi-asset portfolio problem using Pontryagin’s Maximum Principle (PMP).
-This principle provides necessary optimality conditions characterized by a coupled system of forward-backward stochastic differential equations (FBSDEs) for the state and costate variables (Section 2.2).
-However, finding analytical solutions to these FBSDEs is often infeasible, especially in high-dimensional settings involving numerous assets ($\bm{\pi}_{t}\in\mathbb{R}^{n}$) or complex exogenous state dynamics ($\mathbf{Y}_{t}\in\mathbb{R}^{k}$).
-Consequently, translating the continuous-time optimality conditions into a practical solution requires effective numerical approximations.
-We adopt a gradient-based approach rooted in modern deep learning techniques.
-The core idea is to directly parameterize the investor’s policy, both the investment strategy $\bm{\pi}_{t}$ and the consumption rate $C_{t}$, using neural networks, denoted as $\bm{\pi}_{\theta}(t,x,\mathbf{y})$ and $C_{\phi}(t,x,\mathbf{y})$, where $(\theta,\phi)$ are the network parameters.
-Optimization then involves iteratively refining these parameters to maximize the investor’s expected utility.
-A key challenge is computing the gradient of the expected utility objective with respect to the policy parameters $(\theta,\phi)$.
-Furthermore, a practical policy must be effective across a relevant range of potential starting times $t_{0}$, initial wealth levels $x_{0}$, and initial state variable values $\mathbf{y}_{0}$.
-Let $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ denote the conditional expected utility obtained when starting from state $(t_{0},x_{0},\mathbf{y}_{0})$ and following the policy defined by $(\theta,\phi)$:
+While Pontryagin’s Maximum Principle (PMP) provides necessary optimality conditions via forward-backward stochastic differential equations (FBSDEs, Section 2.2), solving these analytically is often infeasible in high dimensions.
+We therefore adopt a gradient-based numerical approach using deep learning.
+The core idea is to parameterize the investor’s policy $(\bm{\pi}_{t},C_{t})$ with neural networks $(\bm{\pi}_{\theta}(t,x,\mathbf{y}),C_{\phi}(t,x,\mathbf{y}))$, where $(\theta,\phi)$ are the parameters to be optimized.
+The objective is to find parameters $(\theta,\phi)$ that maximize the expected utility achieved by the policy $(\bm{\pi}_{\theta},C_{\phi})$.
+To ensure robustness across different starting scenarios, we aim to maximize the extended value function, $\widetilde{J}(\theta,\phi)$.
+This is defined as the expected value of the conditional expected utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ achieved under policy $(\theta,\phi)$, where the initial state $(t_{0},x_{0},\mathbf{y}_{0})$ is drawn from a distribution $\eta$ over a relevant domain $\mathcal{D}$:
 […]
-Here, the expectation $\mathbb{E}[\cdot|\cdot]$ is taken over all possible paths of the state processes $(X_{u},\mathbf{Y}_{u})$ for $u\in[t_{0},T]$, conditional on starting at $(x_{0},\mathbf{y}_{0})$ at time $t_{0}$.
-To ensure the policy performs well across various starting scenarios, we define the optimization target as the extended value function, $\widetilde{J}(\theta,\phi)$.
-This is the expected value of the conditional utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ when the initial state $(t_{0},x_{0},\mathbf{y}_{0})$ is drawn from a chosen distribution $\eta$ over a relevant domain $\mathcal{D}\subset[0,T]\times\mathbb{R}^{+}\times\mathbb{R}^{k}$:
-$\widetilde{J}(\theta,\phi)\;=\;\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\eta}\Big{[}J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\Big{]}.$ (7)
-Maximizing this extended objective $\widetilde{J}(\theta,\phi)$ encourages the learning of a robust policy applicable across the domain $\mathcal{D}$ covered by the initial state distribution $\eta$.
-This paper develops and implements the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, extending the core ideas of Huh, (2024) to the multi-asset, state-dependent setting, with the goal of maximizing the extended value function $\widetilde{J}(\theta,\phi)$.
-The PG-DPO framework leverages automatic differentiation, specifically backpropagation-through-time (BPTT), applied to Monte Carlo simulations drawn from the initial state distribution $\eta$ and evolved according to the system dynamics.
-A central feature of applying BPTT in this context is that it not only provides unbiased estimates of the policy gradient $\nabla\widetilde{J}$, necessary for optimizing the neural network parameters $(\theta,\phi)$, but also inherently calculates estimates of the Pontryagin costate variables ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$) along each simulated path.
-These costates are crucial as they represent the sensitivities of the expected utility to state perturbations, as dictated by PMP.
-This dual output of BPTT–policy gradients and costate estimates–allows for two distinct algorithmic approaches, both offering scalable alternatives to traditional methods hindered by the curse of dimensionality:
-  • The Baseline PG-DPO algorithm (§3.1) utilizes the BPTT-derived policy gradient estimates ($\nabla_{\text{batch}}\approx\nabla\widetilde{J}$) directly within a stochastic gradient ascent procedure to train the neural network policy $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end.
-In this approach, the costate estimates primarily serve to structure the gradient calculation, implicitly guiding the optimization according to PMP.
-  • The Two-Stage PG-DPO algorithm (§3.2) takes a different route.
-It uses the baseline procedure for a short warm-up phase specifically to generate stable and accurate estimates of the costates (potentially involving Monte Carlo averaging).
-These refined costate estimates are then directly plugged into the analytical first-order optimality conditions derived from PMP to define the deployed control actions, bypassing the direct use of the warm-up phase neural networks.
-Both variants avoid explicit FBSDE solvers and leverage the efficiency of BPTT, but they differ significantly in how they utilize the information generated during backpropagation to arrive at the final control policy.
-The subsequent subsections detail each approach.
+$\widetilde{J}(\theta,\phi)=\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\eta}\Big{[}J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\Big{]}$ (12)
+Here, $\mathbb{E}^{\theta,\phi}[\cdot|\cdot]$ denotes expectation under the dynamics generated by policy $(\theta,\phi)$.
+This paper develops and implements the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework to maximize $\widetilde{J}(\theta,\phi)$.
+PG-DPO leverages automatic differentiation, specifically backpropagation-through-time (BPTT), applied to Monte Carlo simulations drawn from the initial state distribution $\eta$.
+A key feature is that BPTT applied to these simulations yields not only unbiased estimates of the policy gradient $\nabla\widetilde{J}$ needed for optimization, but also inherently calculates estimates of the Pontryagin costate variables ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$) along each path, representing sensitivities dictated by PMP.
+This dual output enables distinct algorithmic approaches detailed in the subsequent subsections.
 3.1 Baseline PG-DPO for the Extended Objective
@@
 By iterating these steps, the baseline PG-DPO algorithm gradually optimizes the policy $(\bm{\pi}_{\theta},C_{\phi})$ to maximize the extended value function $\widetilde{J}$.
+Algorithm 1 Baseline Pontryagin‑Guided Direct Policy Optimization (PG‑DPO)
+1:Inputs:
+2:Policy nets $\pi_{\theta},\,C_{\phi}$; learning‑rate schedule $\{\alpha_{j}\}_{j=0}^{J_{\max}-1}$; max iters $J_{\max}$, batch size $M$, time steps $N$; initial‑state distribution $\eta$; utility $U$, discount $\delta$, terminal weight $\kappa$; market params $(\bm{\mu},\bm{\sigma},r,\bm{\mu}_{Y},\bm{\sigma}_{Y},\bm{\Omega})$
+3:Output: optimised params $(\theta_{\text{opt}},\phi_{\text{opt}})$
+4:Initialise $\theta,\phi$
+5:for $j=0$ to $J_{\max}-1$ do
+6:     Draw mini‑batch $\{(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)})\}_{i=1}^{M}\sim\eta$
+7:     $\nabla_{\text{batch}}\leftarrow\mathbf{0}$
+8:     for $i=1$ to $M$ do
+9:         $(X_{0},\mathbf{Y}_{0})\leftarrow(x_{0}^{(i)},\mathbf{y}_{0}^{(i)})$, $\Delta t\leftarrow(T-t_{0}^{(i)})/N$
+10:         for $k=0$ to $N-1$ do
+11:              $t_{k}\leftarrow t_{0}^{(i)}+k\Delta t$
+12:              $\bm{\pi}_{k}\leftarrow\pi_{\theta}(t_{k},X_{k},\mathbf{Y}_{k})$
+13:              $C_{k}\leftarrow C_{\phi}(t_{k},X_{k},\mathbf{Y}_{k})$
+14:              Sample $\Delta\mathbf{W}_{k}\sim\mathcal{N}(\mathbf{0},\bm{\Omega}\Delta t)$
+15:              $\mathbf{Y}_{k+1}=\mathbf{Y}_{k}+\bm{\mu}_{Y}(t_{k},\mathbf{Y}_{k})\Delta t+\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\Delta\mathbf{W}_{k}^{\text{Y}}$
+16:              $X_{k+1}=X_{k}+\Bigl{[}X_{k}\bigl{(}r(t_{k},\mathbf{Y}_{k})+\bm{\pi}_{k}^{\!\top}(\bm{\mu}(t_{k},\mathbf{Y}_{k})-r(t_{k},\mathbf{Y}_{k})\bm{1})\bigr{)}-C_{k}\Bigr{]}\Delta t+X_{k}\bm{\pi}_{k}^{\!\top}\bm{\sigma}(t_{k},\mathbf{Y}_{k})\,\Delta\mathbf{W}_{k}^{\text{X}}$
+17:         end for
+18:         $J^{(i)}=\sum_{k=0}^{N-1}e^{-\delta t_{k}}U(C_{k})\Delta t+\kappa e^{-\delta T}U(X_{N})$
+19:         $g^{(i)}\leftarrow\nabla_{(\theta,\phi)}J^{(i)}$
+20:         $\nabla_{\text{batch}}\leftarrow\nabla_{\text{batch}}+g^{(i)}$
+21:     end for
+22:     $(\theta,\phi)\leftarrow(\theta,\phi)+\dfrac{\alpha_{j}}{M}\nabla_{\text{batch}}$
+23:end for
+24:return $(\theta,\phi)$
 Role of Intermediate BPTT Values and the “Pontryagin-Guided” Nature.
 It is important to understand the role of the intermediate values computed during BPTT in Step 4.
-As BPTT calculates the policy gradient $\nabla J^{(i)}(\dots)$ via the chain rule, it inherently computes the partial derivatives of the realized reward $J^{(i)}(\dots)$ with respect to the intermediate states along path $i$:
+As BPTT calculates the policy gradient $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$ via the chain rule, it inherently computes the partial derivatives of the realized reward $J^{(i)}(\dots)$ with respect to the intermediate states along path $i$:
 $\lambda_{k}^{(i)}=\frac{\partial J^{(i)}(t_{0}^{(i)},\dots)}{\partial X_{k}^{(i)}},\quad\bm{\eta}_{k}^{(i)}=\frac{\partial J^{(i)}(t_{0}^{(i)},\dots)}{\partial\mathbf{Y}_{k}^{(i)}}.$
-These values, $\lambda_{k}^{(i)}$ and $\bm{\eta}_{k}^{(i)}$, are unbiased estimates of the Pontryagin costates associated with the conditional expected utility $J(t_{0}^{(i)},\dots)$.
+These sensitivities, computed mechanically by AD, are directly linked to the discrete adjoint Backward Stochastic Differential Equation (BSDE) structure required by Pontryagin’s Maximum Principle (PMP).
+To illustrate this crucial connection for the wealth component (denoting $\lambda_{k}^{(i)}$ for $\lambda_{k}^{X,(i)}$ for simplicity), consider the discrete Euler-Maruyama step for wealth under policy $(\bm{\pi}_{k}^{(i)},C_{k}^{(i)})$:
+$X_{k+1}^{(i)}=X_{k}^{(i)}+\Bigl{[}X_{k}^{(i)}\Bigl{(}r(t_{k},\mathbf{Y}_{k}^{(i)})+\bm{\pi}_{k}^{(i)\top}\bigl{(}\bm{\mu}(t_{k},\mathbf{Y}_{k}^{(i)})-r(t_{k},\mathbf{Y}_{k}^{(i)})\mathbf{1}\bigr{)}\Bigr{)}-C_{k}^{(i)}\Bigr{]}\Delta t+X_{k}^{(i)}\bm{\pi}_{k}^{(i)\top}\bm{\sigma}(t_{k},\mathbf{Y}_{k}^{(i)})\,\Delta\mathbf{W}_{k}^{(i)}.$
+Applying the chain rule one step backward ($\lambda_{k}^{(i)}=\lambda_{k+1}^{(i)}\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}$) and using a first-order expansion yields the following backward recurrence relation along the path $(i)$:
+$\lambda_{k}^{(i)}\approx\lambda_{k+1}^{(i)}+\lambda_{k+1}^{(i)}\bigl{[}r+\pi_{k}^{(i)}(\mu-r)\bigr{]}\Delta t+\lambda_{k+1}^{(i)}\pi_{k}^{(i)}\sigma\,\Delta W_{k}^{(i)}.$
+This equation inherently possesses the structure of a discrete BSDE operating along the specific simulation path $(i)$.
+To formally connect this to the BSDE formulation in PMP, which involves processes adapted to the filtration $\{\mathcal{F}_{t_{k}}\}$ (representing available information at time $t_{k}$), one defines the adapted costate $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}|\mathcal{F}_{t_{k}}]$ and the predictable martingale component $Z_{k}$ (related to the conditional expectation of $\lambda_{k+1}^{(i)}\Delta W_{k}^{(i)}$).
+Taking the conditional expectation ($\mathbb{E}[\cdot|\mathcal{F}_{t_{k}}]$) of the pathwise recurrence relation above, and carefully defining $Z_{k}$ using a covariation-based approach, leads precisely to the standard discrete BSDE for the adapted process $\lambda_{k}$:
+$\lambda_{k}=\lambda_{k+1}+\underbrace{\bigl{(}\lambda_{k+1}[r+\pi_{k}(\mu-r)]+\pi_{k}\,\sigma\,Z_{k}\bigr{)}}_{\approx\partial_{x}\mathcal{H}(t_{k},\dots)\text{ from PMP}}\Delta t-Z_{k}\,\Delta W_{k}.$ (13)
+Crucially, the drift term derived through this process corresponds exactly to the derivative of the discrete-time Hamiltonian ($\partial_{x}\mathcal{H}$) from PMP with respect to the state variable $x$.
+A similar structure holds for the costate vector $\bm{\eta}_{k}$ associated with the state variables $\mathbf{Y}_{k}$.
+Therefore, the values $\lambda_{k}^{(i)}$ and $\bm{\eta}_{k}^{(i)}$ computed by BPTT are indeed pathwise numerical counterparts related to the Pontryagin costates, satisfying the relevant discrete BSDE structure.
 Crucially, in the baseline PG-DPO algorithm, these costate estimates themselves (and their potential derivatives w.r.t.
 states) are not directly used in the parameter update rule (Step 6), which only requires the final gradient $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$.
-Nevertheless, the fact that BPTT provides access to these estimates is fundamental.
+Nevertheless, the fact that BPTT provides access to these theoretically grounded estimates is fundamental.
 Firstly, they are essential inputs for the Two-Stage PG-DPO variant discussed in the next subsection (§3.2).
 The baseline algorithm thus lays the groundwork by generating these estimates.
-Secondly, this connection explains why the algorithm is termed Pontryagin-Guided.
-Even though the baseline update only uses $\nabla J^{(i)}(\dots)$, the calculation of this gradient via BPTT structurally mirrors the backward propagation of sensitivities inherent in PMP’s adjoint equations, using the costate estimates $\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ implicitly within the chain rule (cf.
+Secondly, this deep connection explains why the algorithm is termed Pontryagin-Guided.
+Even though the baseline update only uses $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$, the calculation of this gradient via BPTT structurally mirrors the backward propagation of sensitivities inherent in PMP’s adjoint equations (BSDEs), using the costate estimates $\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ implicitly within the chain rule (cf.
 Huh, , 2024).
-Therefore, the optimization process, driven by these BPTT-computed gradients, is implicitly guided by the optimality conditions derived from Pontryagin’s principle, aiming to maximize the expected objective $\widetilde{J}$.
+Consequently, the optimization process, driven by these BPTT-computed gradients, is implicitly guided by the optimality conditions derived from Pontryagin’s principle, aiming to maximize the expected objective $\widetilde{J}$.
 3.2 Two-Stage PG-DPO Algorithm
@@
 The Two-Stage PG-DPO variant offers a potentially more computationally efficient alternative by exploiting an empirical observation: the costate estimates ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ and their relevant derivatives with respect to states, like $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$), which are generated during BPTT, often stabilize and become reasonably accurate much faster than the full policy networks achieve optimality.
+Algorithm 2 Two‑Stage PG‑DPO: Warm‑up Training & Deployment
+1:Stage 1: Warm‑up Training
+2:Inputs of Algorithm 1; warm‑up iterations $K_{0}$
+3:Stabilised parameters $(\theta^{*},\phi^{*})$
+4:Run Algorithm 1 for $K_{0}$ iterations to obtain $(\theta^{*},\phi^{*})$
+5:
+6:Stage 2: Deployment at state $(t,X,\bm{Y})$
+7:State $(t,X,\bm{Y})$; parameters $(\theta^{*},\phi^{*})$; Monte‑Carlo rollouts $M_{\mathrm{MC}}$, steps $N^{\prime}$
+8:Near‑optimal controls $(\bm{\pi}^{\mathrm{PMP}},C^{\mathrm{PMP}})$
+9:Initialise $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}\leftarrow\varnothing$
+10:for $j=1$ to $M_{\mathrm{MC}}$ do
+11:     Simulate path with $(\pi_{\theta^{*}},C_{\phi^{*}})$ using $N^{\prime}$ steps
+12:     Compute realized reward for path $j$:
+13:     $\qquad\displaystyle J^{(j)}\leftarrow\sum_{n=0}^{N^{\prime}-1}e^{-\delta t_{n}}U\bigl{(}C_{t_{n}}^{(j)}\bigr{)}\,\Delta t^{\prime}+\kappa e^{-\delta T}U\bigl{(}X_{T}^{(j)}\bigr{)}$
+14:     Evaluate $\lambda_{t}^{(j)}$, $\partial_{x}\lambda_{t}^{(j)}$, $\partial_{\bm{Y}}\lambda_{t}^{(j)}$ and append
+15:end for
+16:Compute $\hat{\lambda}_{t},\widehat{\partial_{x}\lambda}_{t},\widehat{\partial_{\bm{Y}}\lambda}_{t}$triangleright E.g., by averaging estimates from loop
+17:$C^{\mathrm{PMP}}\leftarrow(U^{\prime})^{-1}\!\bigl{(}e^{\delta t}\hat{\lambda}_{t}\bigr{)}$triangleright Using PMP FOC Eq. (8)
+18:Compute $\bm{\pi}^{\mathrm{PMP}}$ via PMP FOC Eq. (15) $\triangleright$ Using estimated costates/derivatives
+19:return $(\bm{\pi}^{\mathrm{PMP}},C^{\mathrm{PMP}})$
 This Two-Stage approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly from the PMP first-order conditions, bypassing the need for fully converged policy networks for deployment.
@@
     The optimal investment portfolio is calculated as:
-[…]
-    $+\underbrace{\bm{\sigma}(t_{k},\mathbf{Y}_{k})^{\top}\!\bigl{[}(\partial_{\mathbf{Y}}\lambda_{k})\,\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\rho^{\top}\,\Psi^{-1}\bigr{]}}_{\text{Hedging Demand}}\Biggr{\}}.$
+    $\bm{\pi}_{k}^{\mathrm{PMP}}$ $=-\frac{1}{X_{k}\,\partial_{x}\lambda_{k}}\,\Sigma^{-1}\!\bigl{(}t_{k},\mathbf{Y}_{k}\bigr{)}\Bigl{\{}\lambda_{k}\bigl{[}\bm{\mu}(t_{k},\mathbf{Y}_{k})-r(t_{k},\mathbf{Y}_{k})\mathbf{1}\bigr{]}$ (14)
+    $\qquad+\bm{\sigma}(t_{k},\mathbf{Y}_{k})^{\top}(\partial_{\mathbf{Y}}\lambda_{k})\,\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\rho^{\top}\Psi^{-1}\Bigr{\}}.$ (15)
     Finally, these analytically computed values $(C_{k}^{\mathrm{PMP}},\bm{\pi}_{k}^{\mathrm{PMP}})$, based on the averaged costate estimates, are used as the control actions for the state $(t_{k},X_{k},\mathbf{Y}_{k})$ during deployment.
@@
 The Monte Carlo averaging step in Stage 2 is crucial for obtaining reliable inputs for the PMP formulas from the potentially noisy single-path estimates.
-3.3 Rationale for the Two-Stage Approach: Leveraging Costate Estimation
-The preceding subsections introduced two variants of the PG-DPO framework: a baseline approach (§3.1) and a Two-Stage approach (§3.2).
-A natural question arises: why introduce the Two-Stage variant, which potentially adds complexity, instead of solely relying on optimizing the policy networks directly?
-This subsection elaborates on the theoretical and practical motivations.
-3.3.1 Qualitative Rationale: Exploiting PMP Structure and Costate Properties
-At its core, the rationale stems from the structure provided by Pontryagin’s Maximum Principle (PMP) itself and the economic meaning of the costate variables.
-As established in Section 2.2, PMP characterizes the optimal controls ($\bm{\pi}_{t}^{*},C_{t}^{*}$) directly in terms of the costates ($\lambda_{t}^{*},\bm{\eta}_{t}^{*}$) and their relevant derivatives (e.g., $\partial_{x}\lambda_{t}^{*},\partial_{\mathbf{Y}}\lambda_{t}^{*}$).
-The analytical formulas used in Stage 2 of the Two-Stage method are precisely these first-order conditions.
-Therefore, if accurate estimates of these costates and their derivatives could be obtained, PMP provides a direct blueprint for constructing the optimal policy.
-Furthermore, the primary costate $\lambda_{t}$ represents the marginal utility of wealth ($\lambda_{t}\approx\partial V/\partial x$).
-Economic principles, like consumption smoothing, suggest agents strive for a stable marginal utility path.
-Mathematically, this often means the value function $V$ and its derivatives ($\lambda_{t},\partial_{x}\lambda_{t},\partial_{\mathbf{Y}}\lambda_{t}$, etc.) exhibit relatively smooth behavior or possess a simpler analytical structure compared to the optimal policy functions $(\bm{\pi}^{*},C^{*})$ themselves.
-The optimal policy might need to be highly complex, especially in non-affine settings or under constraints.
-This difference in potential complexity has direct implications for learning dynamics.
-It is plausible that learning the potentially smoother or more structured functions associated with the costates via BPTT during the warm-up phase (Stage 1) is an easier task-requiring fewer iterations for stabilization-than learning the full complexity of the optimal policy networks $(\bm{\pi}_{\theta},C_{\phi})$ until convergence.
-This aligns with empirical observations that costate estimates often stabilize relatively quickly.
-The potential for faster convergence or stabilization of costate-related functions compared to the full policy motivates the Two-Stage approach.
-This is further encouraged, albeit heuristically for general models, by theoretical stability results known for specific structured models, as formalized in the next subsection.
-3.3.2 Derivative Stability in Structured Models
-For certain classes of portfolio choice problems that admit explicit solutions, such as those within affine or quadratic frameworks, a strong link exists between the accuracy of the value function approximation and the accuracy of its derivatives, provided the approximation retains the original structure.
-This property is formally stated as follows:
-Proposition 3.1 (Derivative Closeness from Value Closeness in Structured Models).
-Consider a class of portfolio choice models (e.g., Merton with HARA utility (Merton, , 1971), Kim & Omberg with HARA utility (Kim & Omberg, , 1996), or Liu’s affine/quadratic framework with CRRA utility (Liu, , 2007)) where the optimal value function $V^{*}(t,w,\mathbf{y})$ has a known analytical structure (e.g., separable or exponential-polynomial in state variables with coefficients determined by ODEs).
-Let $\tilde{V}(t,w,\mathbf{y})$ be another function defined on a compact domain $\mathcal{D}\subset[0,T]\times\mathbb{R}^{+}\times\mathbb{R}^{k}$ that shares the same analytical structure as $V^{*}$.
-If the value functions are uniformly close on $\mathcal{D}$, i.e.,
-$\sup_{(t,w,\mathbf{y})\in\mathcal{D}}|V^{*}(t,w,\mathbf{y})-\tilde{V}(t,w,\mathbf{y})|\longrightarrow 0,$
-then their corresponding spatial derivatives (first and second order with respect to $w$ and $\mathbf{y}$) also converge uniformly on $\mathcal{D}$:
-$\sup_{\mathcal{D}}\|\nabla_{w,\mathbf{y}}V^{*}-\nabla_{w,\mathbf{y}}\tilde{V}\|\to 0\quad\text{and}\quad\sup_{\mathcal{D}}\|\nabla^{2}_{w,\mathbf{y}}V^{*}-\nabla^{2}_{w,\mathbf{y}}\tilde{V}\|\to 0.$
-(Proof sketch provided in Appendix C for the Liu (2007) framework).
-Implication for PG-DPO Rationale.
-Proposition 3.1 highlights that for well-structured problems (like those within the affine class for which the proposition holds), the derivatives inherent in the PMP conditions are well-behaved with respect to approximations that preserve the specific analytical structure.
-However, a key motivation for employing methods like PG-DPO is their potential applicability to a broader class of problems, including models that may lack such convenient structures — for instance, models involving non-affine dynamics (such as the CEV examples considered later in this paper).
-The true optimal value function $V^{*}$ for these more general models may not possess the specific separable or exponential-polynomial form assumed in the proposition.
-Consequently, Proposition 3.1, proven for specific structured models, does not directly guarantee that similar derivative stability holds for the wider range of potentially non-affine models that PG-DPO aims to tackle.
-Establishing whether analogous stability properties exist in these more general settings, or rigorously proving the convergence of derivative estimates obtained via BPTT from neural network approximations in such cases, remains an area for future research and is beyond the scope of this paper.
-Nevertheless, the proposition provides valuable context.
-It confirms that the target objects (optimal value function derivatives) possess inherent stability within important, albeit specific, classes of models.
-This offers heuristic support for the hypothesis underlying the Two-Stage PG-DPO: namely, that the costate derivatives, being fundamental quantities in a well-posed optimal control problem, might be estimated reliably via BPTT (Stage 1) and effectively used in the PMP formulas (Stage 2), even when tackling more complex, potentially non-affine models.
+3.3 Rationale for the Two-Stage Approach
+3.3.1 The Parabolic Nature of Financial Models and PDE Regularity
+Table 1: Classification of Parabolic HJB PDEs in Financial Models
+Class Diffusion Matrix $a(t,x,\mathbf{y})$ Property SDE Coefficient Requirements and Dynamic Portfolio Examples
+(U) Uniformly Elliptic Uniformly positive definite across the state space The diffusion coefficients $\bm{\sigma}$ (for wealth $X_{t}$) and $\bm{\sigma}_{Y}$ (for state $\mathbf{Y}_{t}$) are such that the resulting diffusion matrix in the HJB is non-zero and bounded below by a positive constant everywhere ($\bm{\sigma}\Psi\bm{\sigma}^{\top}$ and $\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}$ uniformly positive definite).
+Drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) are also sufficiently regular (e.g., Lipschitz).
+($\implies$ Value Function / Costate very smooth) Examples: Merton’s portfolio problem with constant coefficients ($\mu,\sigma>0,r$ are constants); models with state-dependent volatility that remains strictly positive; the multi-asset OU model in Section 2.3.1 (if asset volatilities $\sigma_{i}>0$, factor volatility $\sigma_{Y}>0$, and $\Psi$ is positive definite).
+(H) Hypo-elliptic / Uniform on Compacts May degenerate at some points, but regularity is achieved via system structure (e.g., drift-diffusion coupling via Hörmander’s condition) or holds within a compact domain of interest.
+Degeneracy occurs away from boundaries.
+Diffusion coefficients ($\bm{\sigma}$ or $\bm{\sigma}_{Y}$) can become zero at certain points within the domain.
+However, interaction with non-zero drift terms or other diffusion terms maintains regularity.
+($\implies$ Value Function / Costate smooth in relevant regions) Examples: The multi-asset multi-factor model in Section 2.3.2 (if some factor volatilities in $\bm{\sigma}_{Y}$ can be zero at certain states but the overall system structure, possibly involving $\kappa_{Y}$, ensures hypo-ellipticity); a Vasicek interest rate model affecting asset returns $\mu$ but not $\sigma$.
+(D) Degenerate at the Boundary Degenerates ($\to 0$) as the state approaches a boundary.
+The diffusion coefficients $\bm{\sigma}$ or $\bm{\sigma}_{Y}$ (or the combined diffusion matrix) approach zero as a state variable (e.g., wealth, volatility, interest rate) approaches a boundary (typically 0).
+($\implies$ Value Function / Costate regularity may break down near the boundary) Examples: CIR (Cox-Ingersoll-Ross) interest rate/volatility model (diffusion term like $\sigma_{Y}\sqrt{Y_{t}}$); Heston stochastic volatility model (asset diffusion term like $\sigma\sqrt{\nu_{t}}$).
+In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are modeled by stochastic differential equations (SDEs), the Value Function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically solves a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming & Soner, , 2006; Yong & Zhou, , 2012).
+This HJB equation exhibits a parabolic form.
+The specific nature of its parabolicity is primarily determined by the diffusion coefficients of the underlying SDEs, namely $\bm{\sigma}$ (related to assets) and $\bm{\sigma}_{Y}$ (related to state variables).
+While the drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) influence the first-order terms in the HJB and are crucial for the solution’s properties, they do not dictate the type’ of parabolicity, which is mainly governed by the second-order terms (diffusion).
+The behavior of the diffusion matrix $a(t,x,\mathbf{y})$ (the matrix multiplying second-order derivatives in the HJB equation) allows for classifying parabolic PDEs into three main classes (see, e.g., Evans, , 2022).
+This classification significantly impacts the regularity (smoothness) of the Value Function $V$ and its associated Costates (the first-order derivatives $\nabla V$) (cf.
+Evans, , 2022, Ch.
+6-7).
+Costates represent the sensitivity of the Value Function, and Pontryagin’s Maximum Principle links optimal Costates to the optimal policy.
+The core idea of Two-Stage PG-DPO involves estimating Costates and their derivatives via BPTT and using these estimates in the PMP formulas.
+Therefore, the regularity of the Costates—tied to the model’s parabolicity class—directly affects the difficulty of Costate estimation and the effectiveness of the Two-Stage method.
+The table 1 summarizes these classes, their properties, links to SDE coefficients, and relevant examples within the dynamic portfolio optimization context.
+Models falling into classes (U) and (H) are expected to possess sufficiently regular Costates, providing a favorable setting for the estimation required by the Two-Stage PG-DPO method.
+The numerical experiments in this paper are conducted using models belonging to classes (U) or (H), where the Two-Stage PG-DPO effectively leverages the expected regularity of the Costates.
+The method is particularly effective as it generally avoids visiting degenerate boundaries during practical implementation.
+3.3.2 Mathematical Foundation: Derivative Stability via PDE Comparison
+A key theoretical insight, particularly demonstrable in well-structured (e.g., affine) models, comes from analyzing the stability of value function derivatives under policy perturbations using regularity theory for parabolic partial differential equations (e.g., Evans, , 2022; Krylov, , 2024).
+This perspective illustrates how closeness of policies (in a specific sense) can imply closeness of derivatives, providing a foundation for why the Two-Stage approach might be effective.
+First, let us recall the PDE associated with a given policy.
+For an admissible and sufficiently smooth policy $(\pi(t,z),c(t,z))$, where $z=(x,y)$ represents the state, the corresponding value function $V(t,z)$ (representing the expected utility achieved under this *specific* policy from state $z$ at time $t$) satisfies the linear parabolic PDE:
+$-\frac{\partial V}{\partial t}=\mathcal{L}_{\pi,c}[V](t,z)+f_{\pi,c}(t,z),\quad\text{subject to }V(T,z)=g(z).$ (16)
+Here, $g(z)=\kappa e^{-\delta T}U(x)$ is the terminal utility, $f_{\pi,c}(t,z)=e^{-\delta t}U(c(t,z))$ is the instantaneous utility source term under policy $c$, and $\mathcal{L}_{\pi,c}$ is the second-order linear differential generator associated with the state dynamics under policy $(\pi,c)$:
+$\mathcal{L}_{\pi,c}[V]=$ $V_{x}\left[x\left((1-\pi^{\top}\mathbf{1})r(t,y)+\pi^{\top}\mu(t,y)\right)-c(t,z)\right]+V_{\mathbf{y}}^{\top}\mu_{Y}(t,y)$
+$+\frac{1}{2}V_{xx}x^{2}\pi^{\top}\sigma(t,y)\Psi\sigma(t,y)^{\top}\pi$
+$+V_{x\mathbf{y}}^{\top}x\pi^{\top}\sigma(t,y)\rho\sigma_{Y}(t,y)^{\top}$ (17)
+$+\frac{1}{2}\text{Tr}\left(V_{\mathbf{y}\mathbf{y}}\sigma_{Y}(t,y)\Phi\sigma_{Y}(t,y)^{\top}\right)-\delta V$
+(We incorporate the time discount term $-\delta V$ into the operator $\mathcal{L}_{\pi,c}$ for convenience).
+Let $V^{*}(t,z)$ denote the true optimal value function of the control problem defined in Section 2.1, achieved by the optimal policy $(\pi^{*}(t,z),c^{*}(t,z))$.
+Assuming sufficient regularity, $V^{*}$ satisfies the non-linear Hamilton-Jacobi–Bellman (HJB) equation (Fleming & Soner, , 2006), which implies it also satisfies the linear PDE (16) when the optimal policy $(\pi^{*},c^{*})$ is used:
+$-\frac{\partial V^{*}}{\partial t}=\mathcal{L}_{\pi^{*},c^{*}}[V^{*}](t,z)+f_{\pi^{*},c^{*}}(t,z),\quad\text{subject to }V^{*}(T,z)=g(z).$ (18)
+To formalize the notion of policy closeness relevant to the PDE coefficients, we introduce the following definition.
+Let $\Sigma_{\text{asset}}(t,y)=\sigma(t,y)\Psi\sigma(t,y)^{\top}$ and $\Sigma_{\text{asset-state}}(t,y)=\sigma(t,y)\rho\sigma_{Y}(t,y)^{\top}$.
+Definition 1 (Coefficient-relevant $\epsilon$-closeness).
+Let the spatial domain be
+$\Omega\subseteq(0,\infty)\times\mathbb{R}^{k},\qquad z=(x,\mathbf{y})\in\Omega,$
+i.e.
+$x$ is wealth and $\mathbf{y}\in\mathbb{R}^{k}$ is the $k$-dimensional state vector.
+(In practice $\Omega$ can be either the full space or a bounded truncated box that contains the region the process visits with high probability.)
+Fix exponents $p>d+2$ and $q\geq 2$, where $d=\dim\Omega=k+1$. For any measurable function $h(t,z)$ set
+[…]
+Let $(\pi^{*},c^{*})$ be the optimal policy and $(\pi,c)$ any admissible policy. We say $(\pi,c)$ is $(\epsilon,p,q)$-close to $(\pi^{*},c^{*})$ on $Q_{T}=(0,T)\times\Omega$ if
+$\|c-c^{*}\|_{L^{q,p}}$ $\leq\epsilon,$
+$\|\pi^{\top}\mathbf{1}-\pi^{*\top}\mathbf{1}\|_{L^{q,p}}$ $\leq\epsilon,$
+$\|\pi^{\top}\bm{\mu}-\pi^{*\top}\bm{\mu}\|_{L^{q,p}}$ $\leq\epsilon,$
+$\|\pi^{\top}\Sigma_{\text{asset}}\pi-\pi^{*\top}\Sigma_{\text{asset}}\pi^{*}\|_{L^{q,p}}$ $\leq\epsilon,$
+$\|\pi^{\top}\Sigma_{\text{asset-state}}-\pi^{*\top}\Sigma_{\text{asset-state}}\|_{L^{q,p}}$ $\leq\epsilon.$
+Proposition 3.1 (Derivative closeness under coefficient $\epsilon$-closeness).
+Let $V^{*}$ be the optimal value function corresponding to the optimal policy $(\pi^{*},c^{*})$ (cf.
+(18)) and let $V$ be the value function generated by another admissible policy $(\pi,c)$ (cf.
+(16)).
+Set
+$D\;:=\;V^{*}-V,\qquad Q_{T}:=(0,T)\times\Omega.$
+Assume
+  1. (A1)
+    Uniform parabolicity. The operator $\mathcal{L}_{\pi^{*},c^{*}}$ is uniformly parabolic on $Q_{T}$.
+  2. (A2)
+    Baseline regularity. There exist exponents $p>d+2$ and $q\geq 2$ (with $d=\dim\Omega=k+1$) such that the value functions lie in the Sobolev space (e.g., Adams & Fournier, , 2003; Evans, , 2022)
+    $V^{*},\,V\;\in\;L^{q}\bigl{(}0,T;W^{2,p}(\Omega)\bigr{)}.$
+  3. (A3)
+    Coefficient $\epsilon$-closeness. The candidate policy $(\pi,c)$ is $(\epsilon,p,q)$-close to $(\pi^{*},c^{*})$ in the sense of Definition 1.
+Then $D$ satisfies the linear PDE
+$-\partial_{t}D\;=\;\mathcal{L}_{\pi^{*},c^{*}}[D]\;+\;S^{\prime},\quad\text{where}\quad S^{\prime}=(\mathcal{L}_{\pi^{*},c^{*}}-\mathcal{L}_{\pi,c})[V]+\bigl{(}f_{\pi^{*},c^{*}}-f_{\pi,c}\bigr{)}.$
+Moreover, the source term is small in the weighted norm:
+$\|S^{\prime}\|_{L^{q,p}}\;\leq\;C_{0}\,\epsilon,$
+and by standard $L^{p}$ / Schauder estimates for uniformly parabolic equations (see, e.g., Evans, , 2022, Ch. 7),
+$\|D\|_{L^{q}\!\bigl{(}0,T;W^{2,p}(\Omega)\bigr{)}}\;\leq\;C\,\|S^{\prime}\|_{L^{q,p}}\;\leq\;C\,\epsilon,$
+for a constant C independent of $\epsilon$.
+Consequently, the value-function gap and all first- and second-order spatial derivatives satisfy
+$\bigl{\|}V^{*}-V\bigr{\|}_{L^{q,p}}+\bigl{\|}\nabla V^{*}-\nabla V\bigr{\|}_{L^{q,p}}+\bigl{\|}D^{2}V^{*}-D^{2}V\bigr{\|}_{L^{q,p}}\;\leq\;C\,\epsilon.$
+Hence the costates $(\nabla V)$ produced under $(\pi,c)$ are $O(\epsilon)$-close to the optimal costates on the entire space–time domain.
+Remark on assumptions, PG–DPO rationale, and scope of the result.
+Proposition 3.1 makes precise the intuition for the uniformly–parabolic class (“U class”’): if two policies yield close PDE-coefficients in the weak sense of Definition 1 (only a global [^1] A still weaker occupation-weighted norm is discussed below.
+$L^{q,p}$-distance), then their value functions—and hence their gradients (costates)—differ by at most $O(\epsilon)$.
+Assumptions (A1)–(A2) are the classical entry ticket for $L^{p}$ / Schauder regularity (Evans, , 2022) and are easily verified for affine models or for models that become uniformly elliptic after adding a small floor to each diffusion coefficient (e.g.
+replace $\sigma(t,z)$ by $\sqrt{\sigma(t,z)^{2}+\varepsilon^{2}}$ with $\varepsilon\sim 10^{-4}$).
+Such a floor leaves the economics virtually unchanged while preventing the diffusion matrix from collapsing to zero, which is exactly what uniform parabolicity requires.
+The genuinely new ingredient is (A3), the $\epsilon$-coefficient closeness that Stage 1 of PG-DPO is designed to enforce.
+Stage 1 produces a sub-optimal policy $(\pi,c)$ whose coefficient error is $\epsilon$.
+Via Proposition 3.1 this translates into an $O(\epsilon)$ error for the costate $\lambda\approx\nabla V$, so the BPTT outputs obtained during the short warm-up are already $O(\epsilon)$-accurate surrogates of the true optimal costates.
+Stage 2 simply substitutes these estimates into the PMP first-order conditions (Pontryagin, , 2014; Yong & Zhou, , 2012), thereby constructing a near-optimal control without having to train the policy network to full convergence.
+Global vs.– local norms.– The result is stated with a global $\|\,\cdot\,\|_{L^{q,p}}$ norm for transparency. If desired, one may replace it by the occupation-weighted norm
+[…]
+where $\mu_{(\pi,c)}$ is the state-density under the candidate policy.
+Because the proof relies on local PDE estimates, the same $O(\epsilon)$ bound holds—now only on the subset of the state space that the process actually visits with non-negligible probability.
+This is valuable for models whose coefficients degenerate in remote regions.
+Beyond the U class.
+– For hypo-elliptic (H class) (cf.
+Hörmander, , 1967) or boundary-degenerate (D class) models (e.g., Oleinik, , 2012), the full global $L^{p}$ theory is subtler, yet the algorithmic logic of Two-Stage PG-DPO is unaffected:
+  • the PMP equations and BPTT apply to any differentiable dynamics;
+  • costates typically stabilise faster than the full optimal control;
+  • a brief warm-up that makes the coefficient error small where it matters is enough to obtain high-quality costate estimates.
+Even when the rigorous guarantee of Proposition 3.1 is unavailable, this costate-first, PMP-second recipe remains an effective practical strategy—an observation confirmed empirically in the numerical section (Section 4) and worth exploring further for genuinely non-affine or degenerate models.
+3.3.3 Qualitative Rationale and Broader Applicability
+While the formal analysis in Proposition 3.1 provides rigorous support under specific regularity conditions (like uniform parabolicity, often associated with affine models), the core ideas underpinning the Two-Stage PG-DPO approach possess a broader qualitative appeal that suggests applicability beyond these settings.
+Crucially, neither the Pontryagin Maximum Principle (PMP) itself (Pontryagin, , 2014; Yong & Zhou, , 2012) nor the mechanics of backpropagation-through-time (BPTT) (e.g., Goodfellow, , 2016) are inherently restricted to affine or uniformly parabolic systems.
+PMP provides necessary optimality conditions for general nonlinear dynamics, provided the required functions are differentiable.
+Similarly, BPTT serves as a general algorithm for computing gradient estimates through complex computational graphs derived from simulations, regardless of the underlying model’s linearity.
+Furthermore, there is a compelling economic and mathematical intuition that the PMP costates ($\lambda\approx\nabla V$, representing marginal values) might exhibit properties that make them more amenable to estimation than the potentially complex optimal policy $\pi^{*}$ itself, especially in high-dimensional problems.
+For instance, the costate functions might be representable by smoother or lower-complexity functions compared to $\pi^{*}$.
+Alternatively, the process of learning or estimating these costate-related functions via BPTT with neural networks might converge more rapidly or stably than directly learning the intricate mapping defined by the optimal policy.
+The Two-Stage PG-DPO strategy is designed precisely to leverage this potential advantage.
+It effectively decouples the optimization challenge into two parts: (i) tackling the potentially more tractable task of estimating costates (and their relevant derivatives) using BPTT during a warm-up phase (ii) employing the robust analytical structure of the PMP first-order conditions to translate these estimates into effective control actions during deployment.
+Therefore, even in scenarios where the rigorous derivative closeness guarantees of Proposition 3.1 are harder to establish formally (e.g., for non-affine models, or systems falling into the hypo-elliptic or degenerate parabolic classes), this underlying qualitative rationale remains persuasive.
+The potential for BPTT-derived costate estimates to stabilize relatively quickly and capture essential optimality information offers a strong motivation for exploring the Two-Stage PG-DPO approach in these more complex settings.
+This rationale, significantly reinforced by the numerical success demonstrated in the affine benchmark case (Section 4), positions PG-DPO (particularly the superior Two-Stage variant) as a promising and potentially powerful tool for tackling challenging high-dimensional, possibly non-affine, continuous-time control problems prevalent in finance and other domains.
+Naturally, its practical effectiveness in such extended settings must be carefully evaluated empirically on a case-by-case basis, highlighting an important direction for future research.
 4 Numerical Experiments
 In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework (both Baseline and Two-Stage variants).
-To enable rigorous comparison with the available semi-analytical benchmark for the terminal wealth problem (Appendix B.2), the numerical experiments presented herein focus exclusively on maximizing $J=\mathbb{E}[U(X_{T})]$.
-Consequently, the optimization of the intermediate consumption policy $C_{\phi}$, though supported by the general framework (Sections 2-3), is omitted in this study.
-We focus on the multi-asset, multi-factor portfolio optimization problem introduced in Section 2.3.2.
-As detailed therein, this specific model involves $n$ risky assets whose expected returns are driven by $k$ correlated Ornstein-Uhlenbeck (OU) state variables $\mathbf{Y}_{t}$.
-Crucially, this model specification falls within the affine class.
-This allows us to compute a semi-analytical benchmark solution by solving a system of Riccati-type ordinary differential equations (ODEs), as outlined in Appendix B.2.
-This benchmark provides a basis for rigorously comparing the accuracy of the PG-DPO methods.
-4.1 Experimental Setup
-4.1.1 Parameter Values
+We focus on the multi-asset, multi-factor Ornstein-Uhlenbeck (OU) portfolio optimization problem introduced in Section 2.3.2.
+The objective is to maximize the expected utility of terminal wealth, $J=\mathbb{E}[U(X_{T})]$, using a CRRA utility function with risk aversion $\gamma=2.0$.
+This specific affine model setting allows for comparison against a semi-analytical benchmark solution derived from the HJB equation (details in Appendix B.2).
+We compare the policies generated by Baseline PG-DPO and Two-Stage PG-DPO against this benchmark using the Root Mean Squared Error (RMSE) metric across various dimensions, scaling up to n=50 assets and k=10 state factors.
+For complete details on the specific parameter values, PG-DPO implementation (network architecture, training setup), benchmark computation, and evaluation procedures, please refer to Appendix C.
+The subsequent subsections present and analyze the results.
+4.1 Convergence Speed and Accuracy Analysis
+This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO methodologies across various dimensional settings.
+Figure 1 displays the average Root Mean Squared Error (RMSE), averaged across $k=10$ evaluation slices, between the learned policies (either $\bm{\pi}_{\theta}$ for Baseline PG-DPO or $\bm{\pi}^{\mathrm{PMP}}$ for Two-Stage PG-DPO) and the benchmark optimal policy ($\bm{\pi}^{*}$).
+The RMSE is plotted as a function of training epochs on a logarithmic scale for configurations with $k=10$ state variables and varying numbers of assets ($n=1,10,50$).
+(a) n=1, k=10
+(b) n=10, k=10
+(c) n=50, k=10
+Figure 1: Average policy RMSE comparison between Baseline PG-DPO (blue) and Two-Stage PG-DPO (orange) relative to the benchmark solution across different dimensions (k=10 fixed).
+Values averaged over k evaluation slices.
+Note the logarithmic scale on the y-axis.
+The RMSE for the Two-Stage PG-DPO (orange line in Figure 1) represents the error of the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ constructed by substituting the BPTT-derived costate estimates into the PMP first-order condition (Eq.
+(10)).
+A key advantage is that the Two-Stage approach explicitly combines the analytically distinct myopic and hedging components.
+As established theoretically and confirmed numerically (see Section 4.2), the error of the myopic component is typically negligible (e.g., on the order of $10^{-7}$).
+Therefore, the total RMSE for the Two-Stage policy is almost entirely attributable to the error in estimating the hedging component.
+This highlights the proficiency of the Two-Stage method in accurately capturing hedging demands using BPTT-derived costate derivatives ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$).
+Table 2: Summary of Minimum Achieved Policy RMSE and Convergence Epochs
+Dimension Method Min. Policy RMSE Epochs at Min.
+(n assets, k factors) (vs. Benchmark)
+n=1, k=10 Baseline PG-DPO $4.27\times 10^{-2}$ $2,600$
+2-PG-DPO $2.80\times 10^{-4}$ $1,500$
+n=10, k=10 Baseline PG-DPO $1.58\times 10^{-1}$ $10,000$
+2-PG-DPO $4.89\times 10^{-3}$ $800$
+n=50, k=10 Baseline PG-DPO $1.26\times 10^{-1}$ $100,000$
+2-PG-DPO $1.32\times 10^{-2}$ $3,200$
+Note: Minimum Policy RMSE values are averaged across the $k=10$ evaluation slices.
+Epochs indicate the recorded iteration number when the minimum average RMSE was observed; evaluations and logging were performed every 200 iterations.
+The key finding from Figure 1 and summarized in Table 2 is the consistent and significant outperformance of the Two-Stage PG-DPO method.
+Table 2 shows that Two-Stage PG-DPO achieves minimum RMSE values dramatically lower than the Baseline method across all dimensions—approximately 150 times lower for $n=1$, 30 times lower for $n=10$, and nearly 10 times lower for $n=50$.
+Furthermore, 2-PG-DPO typically reaches its minimum error much earlier in the training process (e.g., at epoch 1,500 vs.
+2,600 for $n=1,k=10$, and epoch 800 vs.
+10,000 for $n=10,k=10$).
+Notably, while the convergence time for Baseline PG-DPO appears to increase substantially with the problem dimension (n), the time required for 2-PG-DPO to reach its minimum error seems relatively less sensitive to the number of assets.
+Even in the high-dimensional $n=50,k=10$ setting, 2-PG-DPO finds a near-optimal solution with an RMSE of $1.32\times 10^{-2}$ around epoch 3200, whereas the Baseline method’s minimum RMSE of $1.26\times 10^{-1}$ is reached only at the end of a much longer training period (epoch 100,000).
+This clearly demonstrates the superiority of the Two-Stage approach and strongly supports the rationale discussed in Section 3.3: leveraging the PMP structure via BPTT-estimated costates is more efficient and accurate than end-to-end policy approximation, especially as dimensionality increases.
+The subsequent section presents a detailed visual analysis for the most challenging high-dimensional case ($n=50,k=10$) to further illustrate the performance difference.
+Hereafter, we often abbreviate Two-Stage PG-DPO as 2-PG-DPO.
+4.2 High-Dimensional Performance Analysis (n=50, k=10)
+We now visually examine the performance in the most challenging setting: $n=50$ assets and $k=10$ state variables.
+This scenario tests scalability where traditional DP methods are intractable.
+As shown quantitatively in Table 2 and Figure 1c, the 2-PG-DPO method achieves a minimum RMSE ($1.32\times 10^{-2}$ at epoch 3,200) significantly better than the Baseline approach ($1.26\times 10^{-1}$ at epoch 100,000).
+The visualizations in Figures 2 and 3 focus on the policy for the first asset ($i=1$) versus time-to-maturity ($T-t$) and the first state variable ($Y_{1}$), holding other state variables constant at their long-term means.
+The results shown correspond to the state on epoch 3,200, near where the 2-PG-DPO method achieved its minimum observed RMSE.
+(a) Baseline Policy ($\pi_{1,\theta}$)
+(b) 2-PG-DPO Policy ($\pi_{1}^{\mathrm{PMP}}$)
+(c) Baseline Error
+(d) 2-PG-DPO Error
+Figure 2: Comparison of total policies and errors for the first asset ($i=1$) in the n=50, k=10 case on epoch 3,200 (plotted against $Y_{1}$).
+Top row compares policies, bottom row compares errors (using Symlog scale).
+Note the significantly different color bar ranges are used for each plot to best visualize its content, reflecting the dramatic difference in policy/error magnitudes.
+(a) 2-PG-DPO Myopic ($\pi_{1,0}^{\mathrm{PMP}}$)
+(b) 2-PG-DPO Hedging ($\pi_{1,1}^{\mathrm{PMP}}$)
+(c) Myopic Error
+(d) Hedging Error
+Figure 3: Decomposition of the 2-PG-DPO policy for the first asset ($i=1$) in the n=50, k=10 case on epoch 3,200 (plotted against $Y_{1}$).
+Top row shows myopic and hedging components.
+Bottom row shows their respective errors (using Symlog scale).
+Note the use of different color bar ranges, necessary to visualize both the negligible myopic error and the larger hedging error, which dictates the total policy error.
+The challenges of high dimensionality become starkly evident for the Baseline PG-DPO method in Figure 2.
+While the learned policy (2(a)) shows some structure, the corresponding error (2(c)) is extremely large, visually confirming the poor quantitative performance reported in Table 2.
+This indicates severe inaccuracies in the policy learned by the Baseline method when faced with the combined complexity of 50 assets and 10 state variables.
+In dramatic contrast, the 2-PG-DPO method provides a highly accurate solution, as shown by its policy (2(b)) and error (2(d)).
+Comparing the value ranges indicated by the color bars, the error magnitude for the 2-PG-DPO policy is roughly an order of magnitude smaller than the Baseline error, underscoring the robustness and practical necessity of the 2-PG-DPO approach.
+Although perfect accuracy might be harder to achieve as dimensionality increases (the minimum RMSE of $1.32\times 10^{-2}$ is higher than in lower-dimensional cases), it remains remarkably effective and vastly superior to the direct policy learning method.
+The component decomposition in Figure 3 confirms the patterns observed previously.
+The myopic component (3(a)) largely dictates the policy shape, and its associated error (3(c)) remains negligible (confirming quantitative findings that this error is often $O(10^{-7})$), requiring a very small scale on its color bar for visualization.
+The hedging component (3(b)) contributes visibly to the total policy shape.
+Crucially, the hedging error (3(d)) is visually almost identical in pattern to the total 2-PG-DPO error (2(d)) when comparing their respective plots, confirming that the overall accuracy is dictated by the precision of estimating the hedging demands.
+In conclusion, the $n=50,k=10$ results powerfully demonstrate the capability of the 2-PG-DPO framework to “break the dimensional barrier”.
+While the Baseline PG-DPO struggles significantly, 2-PG-DPO delivers a highly accurate and viable solution by leveraging the structure of Pontryagin’s Maximum Principle.
+Its performance remains remarkably strong and vastly superior to the baseline, solidifying its potential as a practical tool for large-scale, continuous-time portfolio optimization.
+5 Conclusion
+This paper tackled the significant challenge posed by the “curse of dimensionality” in solving large-scale continuous-time portfolio optimization problems.
+We introduced and developed the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, which uniquely combines the theoretical rigor of Pontryagin’s Maximum Principle (PMP) with the power of deep learning, specifically using backpropagation-through-time (BPTT) applied to policy networks representing investment and consumption strategies.
+A central contribution is the development of a highly effective “Two-Stage” PG-DPO variant.
+This method leverages the insight that BPTT applied to simulation paths can yield stable estimates of the crucial Pontryagin costates (and their derivatives) relatively quickly, often much faster than the full policy network converges.
+By utilizing these rapidly obtained costate estimates within the analytical first-order optimality conditions derived from PMP, the Two-Stage approach efficiently constructs near-optimal control policies, effectively decoupling the estimation of marginal value sensitivities from the full policy optimization.
+Our numerical experiments, conducted on a multi-asset, multi-factor affine model chosen for its tractability allowing rigorous benchmark comparisons, demonstrated the practical effectiveness and scalability of this approach.
+The results consistently showed that the 2-PG-DPO method achieves substantially higher accuracy and faster convergence compared to the baseline end-to-end policy learning strategy.
+Notably, the framework successfully scaled to problems involving dimensions (e.g., n=50 assets, k=10 state factors) that are typically considered intractable for traditional numerical HJB/PDE solvers, thereby effectively “breaking the dimensional barrier.” We specifically highlighted the method’s capability to accurately capture both myopic and complex intertemporal hedging demands by explicitly leveraging the structure inherent in the PMP optimality conditions.
+While demonstrated on an affine model, the underlying principles relying on PMP and BPTT suggest broader applicability.
+This research opens several avenues for future investigation.
+Firstly, further exploration of the theoretical underpinnings is crucial.
+Our analysis, related to Proposition 3.1, indicates that the rationale behind 2-PG-DPO—leveraging the stability of value function derivatives—is theoretically supported for potentially broad classes of models, possibly including non-affine cases, provided certain regularity conditions related to smoothness and parabolicity are met.
+A key avenue for future theoretical research is to rigorously characterize the precise range of financial models (both affine and non-affine) that satisfy these sufficient regularity assumptions, thereby ensuring the well-posedness of the PMP framework and the reliability of BPTT-based estimates.
+Furthermore, investigating the general theoretical convergence properties of the PG-DPO algorithm, particularly the costate estimation process, remains an important open question.
+Secondly, extending the framework to incorporate more realistic market features is a natural next step for enhancing practical relevance.
+This includes adapting PG-DPO to handle portfolio constraints (like short-selling restrictions or leverage limits) and modeling market frictions such as transaction costs, both of which significantly alter the optimal control landscape.
+Testing the framework’s robustness and performance with different utility functions beyond CRRA (e.g., HARA, prospect theory) is also warranted.
+Moreover, explicitly including the optimization and evaluation of the intermediate consumption policy, omitted in the current study for benchmarking clarity, would provide a more complete assessment of the full framework’s capabilities.
+Thirdly, further empirical validation and algorithmic refinements are desirable.
+Building on the theoretical potential for broader applicability, applying the PG-DPO framework (especially the superior 2-PG-DPO variant) to well-known non-affine financial models (e.g., stochastic volatility models like Heston) would be valuable, even in the absence of exact benchmarks, to showcase its capabilities more broadly.
+Algorithmically, exploring alternative neural network architectures, advanced optimization techniques, and variance reduction methods beyond antithetic variates could potentially lead to further performance improvements or faster convergence.
+Finally, optimizing the implementation details deserves attention.
+This includes refining strategies for determining the optimal warm-up duration for the 2-PG-DPO method to best balance costate accuracy against potential overfitting to simulation artifacts inherent in the discrete-time training process.
+Careful tuning of hyperparameters related to both the network training and the costate estimation phase will be important for practical applications.
+Acknowledgments
+This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-00562904).
+Appendix A Derivation of the Adjoint Diffusion Coefficient via Itô’s Lemma
+In the Pontryagin framework, the adjoint (or costate) process associated with the investor’s wealth $X_{t}$ is identified with the partial derivative of the value function $V$.
+Concretely, we set $\lambda_{t}=V_{X}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},$ where $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ represents additional state variables and $X_{t}$ is the investor’s scalar wealth process.
+Our goal in this section is to derive the diffusion coefficient corresponding to the $d\mathbf{W}^{X}_{t}$ term in the backward SDE for $\lambda_{t}$, namely $\mathbf{Z}_{t}^{\lambda X}$ from BSDE (5), which plays a crucial role in determining the optimal portfolio via FOC (9).
+Recall that under an optimal control $\bigl{(}\bm{\pi}_{t}^{*},C_{t}^{*}\bigr{)}$, the wealth $X_{t}^{*}$ evolves according to Eq.
+(3) and Eq.
+(4), and the $k$-dimensional state process $\mathbf{Y}_{t}^{*}$ follows its own SDE, also presented in Section 2.2.
+The dynamics are driven by correlated Brownian motions $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ and $\mathbf{W}_{t}^{Y}\in\mathbb{R}^{k}$ with the joint covariance structure $\operatorname{Cov}(d\mathbf{W}_{t})=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}dt.$
+Since $\lambda_{t}=V_{X}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},$ we apply Ito’s multidimensional lemma to $V_{X}$. The differential $d\lambda_{t}$ can be expressed as:
+$d\lambda_{t}$ $=\;V_{Xt}\,dt\;+\;V_{XX}\,dX_{t}\;+\;V_{X\mathbf{Y}}\,d\mathbf{Y}_{t}$
+$\quad+\;\tfrac{1}{2}\,V_{XXX}\,\bigl{(}dX_{t}\bigr{)}^{2}\;+\;V_{XX\mathbf{Y}}\,(dX_{t})\,(d\mathbf{Y}_{t})^{\top}\;+\;\tfrac{1}{2}\,\text{Tr}\bigl{(}V_{X\mathbf{Y}\mathbf{Y}}\,(d\mathbf{Y}_{t})(d\mathbf{Y}_{t})^{\top}\bigr{)},$ (19)
+where subscripts denote partial derivatives (e.g., $V_{X\mathbf{Y}}$ is the gradient w.r.t $\mathbf{Y}$ of $V_{X}$, $V_{X\mathbf{Y}\mathbf{Y}}$ is the Hessian w.r.t $\mathbf{Y}$ of $V_{X}$).
+Our objective is to identify the coefficient of $d\mathbf{W}_{t}^{X}$ in the expansion of $d\lambda_{t}$.
+This coefficient is precisely the process $\mathbf{Z}_{t}^{\lambda X}$ appearing in the BSDE (5).
+Substituting the SDEs for $dX_{t}$ and $d\mathbf{Y}_{t}$ (from Section 2.2) into the Ito formula (19), we observe that terms involving $d\mathbf{W}_{t}^{X}$ arise directly from the diffusion part of the $V_{XX}\,dX_{t}$ term, and indirectly from the $V_{X\mathbf{Y}}\,d\mathbf{Y}_{t}$ term due to the correlation $\rho$ between $d\mathbf{W}_{t}^{X}$ and $d\mathbf{W}_{t}^{Y}$.
+A careful application of Ito’s lemma, considering the quadratic variation and covariation terms related to the covariance matrix $\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}$, allows isolation of the full coefficient multiplying $d\mathbf{W}_{t}^{X}$.
+Representing $\mathbf{Z}_{t}^{\lambda X}$ as a $1\times n$ row vector, this process is found to be:
+$\mathbf{Z}_{t}^{\lambda X}=\underbrace{(\partial_{x}\lambda_{t}^{*})X_{t}^{*}\bm{\pi}_{t}^{*\top}\bm{\sigma}(t,\mathbf{Y}_{t}^{*})}_{\text{Term from }V_{XX}dX_{t}}+\underbrace{(\partial_{\mathbf{Y}}\lambda_{t}^{*})\bm{\sigma}_{Y}(t,\mathbf{Y}_{t}^{*})\rho^{\top}\Psi^{-1}}_{\text{Term from }V_{XY}dY_{t}\text{ via correlation}}.$ (20)
+The first term reflects the direct impact of wealth volatility related to the portfolio choice $\bm{\pi}_{t}^{*}$, scaled by the second derivative of the value function with respect to wealth, $V_{XX}$ (or $\partial_{x}\lambda_{t}^{*}$).
+The second term captures the intertemporal hedging component arising from the correlation between asset returns and state variable movements, mediated by the cross-derivative of the value function $V_{X\mathbf{Y}}$ (or $\partial_{\mathbf{Y}}\lambda_{t}^{*}$).
+This derivation provides the explicit form of the adjoint diffusion coefficient $\mathbf{Z}_{t}^{\lambda X}$, which is fundamental for understanding the structure of the optimal portfolio policy (10) and for constructing the Pontryagin adjoint equations.
+Appendix B Analytic Solutions for Affine Models
+This appendix provides details on the derivation of analytic solutions for the portfolio optimization problems presented in Section 2.3 under specific assumptions (no intermediate consumption, CRRA utility), typically following the Hamilton-Jacobi-Bellman (HJB) approach with an exponential-quadratic value function ansatz.
+We also show how these solutions connect to the Pontryagin framework.
+B.1 Analytic Solution for the Multi-Asset OU Model (Section 2.3.1)
+For simplicity, we assume no intermediate consumption, i.e., $C_{t}=0$, and consider an investor who derives utility only from terminal wealth $X_{T}$ under a CRRA objective $U(x)=\frac{x^{1-\gamma}}{1-\gamma}$.
+Following Kim & Omberg, (1996), the associated Hamilton–Jacobi–Bellman (HJB) equation can be solved by positing a value function of the form:
+$V(t,x,Y)=\frac{x^{1-\gamma}}{1-\gamma}\exp\left(g(t)+B(t)Y+\frac{1}{2}C(t)Y^{2}\right).$ (21)
+Here, $g(t)$, $B(t)$, and $C(t)$ are deterministic functions of time determined by solving a system of ordinary differential equations (ODEs) derived from the HJB equation, subject to terminal conditions at $t=T$.
+From this value function, the costate $\lambda_{t}^{*}=V_{x}$, its derivative with respect to wealth $\partial_{x}\lambda_{t}^{*}=V_{xx}$, and its derivative with respect to the state $\partial_{Y}\lambda_{t}^{*}=V_{xY}$ have the following closed-form expressions on the optimal path $(X_{t}^{*},Y_{t})$:
+$\lambda_{t}^{*}(t,X_{t}^{*},Y_{t})$ $=(X_{t}^{*})^{-\gamma}\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right),$
+$\frac{\partial\lambda_{t}^{*}}{\partial x}(t,X_{t}^{*},Y_{t})$ $=-\gamma(X_{t}^{*})^{-\gamma-1}\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right),$
+$\frac{\partial\lambda_{t}^{*}}{\partial Y}(t,X_{t}^{*},Y_{t})$ $=(X_{t}^{*})^{-\gamma}\left(B(t)+C(t)Y_{t}\right)\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right).$
+These expressions lead to remarkably simple ratios that are central to the Pontryagin framework’s characterization of the optimal portfolio:
+$\frac{\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma},\qquad\frac{\partial_{Y}\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma}\left(B(t)+C(t)Y_{t}\right).$
+Substituting these ratios into the general Pontryagin first-order condition for the optimal portfolio $\bm{\pi}_{t}^{*}$ (which equates the marginal contribution to the Hamiltonian from investing in risky assets to zero, cf.
+Eq.
+(9) derived in Section 2.2) confirms the explicit analytic solution derived via the HJB approach:
+$\bm{\pi}^{*}(t,Y)=\frac{1}{\gamma}\,\Sigma^{-1}\Bigl{[}\underbrace{(\bm{\alpha}\,\bm{\sigma})\,Y}_{\text{Myopic}}+\underbrace{\sigma_{Y}\,(\bm{\sigma}\,\bm{\rho})\,\bigl{(}B(t)+C(t)\,Y\bigr{)}}_{\text{Hedging}}\Bigr{]},$
+where $\Sigma=\mathrm{diag}(\bm{\sigma})\,\Psi\,\mathrm{diag}(\bm{\sigma})$, $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$, $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$, $(\bm{\alpha}\,\bm{\sigma})$ is the vector $(\alpha_{1}\sigma_{1},\dots,\alpha_{n}\sigma_{n})^{\top}$, and $(\bm{\sigma}\,\bm{\rho})$ is the vector $(\sigma_{1}\rho_{1},\dots,\sigma_{n}\rho_{n})^{\top}$.
+The derivation confirms that the first term captures the myopic demand driven by the instantaneous risk premium $(\bm{\mu}(Y)-r\mathbf{1})=(\bm{\alpha}\,\bm{\sigma})Y$, while the second term represents the intertemporal hedging demand, proportional to $B(t)+C(t)Y$ which arises from the state-dependency of the investment opportunity set captured by $\partial_{Y}\lambda_{t}^{*}$.
+Substituting this $\bm{\pi}^{*}$ back into the HJB equation and matching coefficients in powers of $Y$ yields the following Riccati-type ODEs for $B(t)$ and $C(t)$, along with the ODE for $g(t)$.
+The system is solved backward from the terminal conditions:
+$g(T)=0,\quad B(T)=0,\quad C(T)=0.$ (22)
+The ODEs are:
+$-\frac{dC(t)}{dt}$ $=-2\,\kappa_{Y}\,C(t)+\sigma_{Y}^{2}\,C(t)^{2}+\frac{1-\gamma}{\gamma}\Bigl{[}\beta_{0}+2\,\sigma_{Y}\,C(t)\,\beta_{1}+\sigma_{Y}^{2}\,C(t)^{2}\,\beta_{2}\Bigr{]},$ (23)
+$-\frac{dB(t)}{dt}$ $=-\kappa_{Y}\,B(t)+\sigma_{Y}^{2}\,B(t)\,C(t)-\frac{1-\gamma}{\gamma}\Bigl{[}\sigma_{Y}^{2}\,B(t)\,C(t)\,\beta_{2}+\sigma_{Y}\,B(t)\,\beta_{1}\Bigr{]}+\kappa_{Y}\,\theta_{Y}\,C(t),$ (24)
+$\frac{dg(t)}{dt}=(1-\gamma)\,r-\frac{1-\gamma}{2\gamma}\Bigl{[}\beta_{0}+2\sigma_{Y}\,\beta_{1}\,B(t)+\sigma_{Y}^{2}\,\beta_{2}\bigl{(}B(t)^{2}+C(t)\bigr{)}\Bigr{]}+\kappa_{Y}\,\theta_{Y}\,B(t)+\tfrac{1}{2}\,\sigma_{Y}^{2}\,C(t).$ (25)
+Here, $r$ is the risk-free rate, $\kappa_{Y}$, $\theta_{Y}$, $\sigma_{Y}$ are the OU parameters, $\gamma$ is the risk aversion coefficient, and $\beta_{0}$, $\beta_{1}$, $\beta_{2}$ are quadratic forms involving the model parameters defined as:
+$\beta_{0}:=(\bm{\alpha}\,\bm{\sigma})^{\top}\,\Sigma^{-1}\,(\bm{\alpha}\,\bm{\sigma}),\quad\beta_{1}:=(\bm{\alpha}\,\bm{\sigma})^{\top}\,\Sigma^{-1}\,(\bm{\sigma}\,\bm{\rho}),\quad\beta_{2}:=(\bm{\sigma}\,\bm{\rho})^{\top}\,\Sigma^{-1}\,(\bm{\sigma}\,\bm{\rho}).$
+By solving the system of ODEs (23), (24), and (25) backward in time from the terminal conditions (22), one obtains the functions $g(t)$, $B(t)$, and $C(t)$.
+Substituting these into Eq.
+(21) gives the complete analytic form of the value function consistent with the underlying HJB/PMP framework.
+B.2 Analytic Solution for the Multi-Asset Multi-Factor Model (Section 2.3.2)
+[…]
+[…]
+where $\mathbf{B}(t)$ is an $k\times 1$ vector, $\mathbf{C}(t)$ is an $k\times k$ symmetric matrix (note: dimensions based on state vector $\mathbf{Y}\in\mathbb{R}^{k}$), and $g(t)$ is a scalar function, all satisfying ODEs derived from the HJB equation, with terminal conditions $g(T)=0$, $\mathbf{B}(T)=\mathbf{0}$ and $\mathbf{C}(T)=\mathbf{0}$.
+Define the relevant matrices and vectors:
+$\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n}),\;\bm{\sigma}_{Y}\;(\text{factor diffusion }k\times k),\;\Sigma=\bm{\sigma}\Psi\bm{\sigma}.$
+where $\Psi\in\mathbb{R}^{n\times n}$ is the correlation matrix for the $n$ risky assets, and let $\rho\in\mathbb{R}^{n\times k}$ be the cross-correlation matrix between these assets and the $k$-dimensional factor Brownian motion $\mathbf{W}^{Y}_{t}$.
+Also let $\mathbf{A}=\bigl{(}\bm{\alpha}_{1},\dots,\bm{\alpha}_{n}\bigr{)}^{\top}\in\mathbb{R}^{n\times k}$ collect the factor loadings $\bm{\alpha}_{i}\in\mathbb{R}^{k}$ for each asset $i$.
+The costate $\lambda_{t}^{*}=V_{x}$, its derivative with respect to wealth $\partial_{x}\lambda_{t}^{*}=V_{xx}$, and its gradient ($k\times 1$ vector) with respect to the state vector $\partial_{\mathbf{Y}}\lambda_{t}^{*}=V_{x\mathbf{Y}}$ are given by:
+$\lambda_{t}^{*}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=(X_{t}^{*})^{-\gamma}\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right),$
+$\frac{\partial\lambda_{t}^{*}}{\partial x}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=-\gamma(X_{t}^{*})^{-\gamma-1}\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right),$
+$\frac{\partial\lambda_{t}^{*}}{\partial\mathbf{Y}}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=(X_{t}^{*})^{-\gamma}\left(\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}_{t}\right)\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right).$
+These yield the following scalar and vector ratios, which simplify the PMP optimality conditions:
+$\frac{\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma},\qquad\frac{\partial_{\mathbf{Y}}\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma}\left(\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}_{t}\right).$
+Substituting these into the vector form of the PMP first-order condition for the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$, (cf. Eq. (10) derived in Section 2.2):
+$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\Sigma^{-1}\left\{\lambda_{t}^{*}(\bm{\mu}(\mathbf{Y}_{t})-r\mathbf{1})+\bm{\sigma}\rho\bm{\sigma}_{Y}\left[(\partial_{\mathbf{Y}}\lambda_{t}^{*})^{\top}\right]\right\},$
+where the risk premium vector is $\bm{\mu}(\mathbf{Y}_{t})-r\mathbf{1}=\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}$, we precisely recover the analytic solution obtained from the HJB framework:
+$\bm{\pi}^{*}(t,\mathbf{Y})\;=\;\frac{1}{\gamma}\,\Sigma^{-1}\Bigl{[}\,\underbrace{\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}}_{\text{Myopic component (related to }\bm{\mu}-r\mathbf{1})}\;+\;\underbrace{\bm{\sigma}\,\rho\,\bm{\sigma}_{Y}\Bigl{(}\mathbf{B}(t)\;+\;\mathbf{C}(t)\,\mathbf{Y}\Bigr{)}}_{\text{Hedging component (related to }\partial_{\mathbf{Y}}\lambda_{t}^{*})}\Bigr{]}.$
+Here, the first bracketed term corresponds to the myopic component, driven by the instantaneous expected excess returns $\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}$.
+The second bracketed term incorporates intertemporal hedging demands arising from the desire to hedge against unfavorable changes in the investment opportunity set, as captured by the factor sensitivities $\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}$ (proportional to the gradient $\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and the correlations encoded in $\rho$, $\Psi$, and $\Phi$ (implicitly via the ODEs for $\mathbf{B}$ and $\mathbf{C}$).
+To determine the time-dependent coefficients $\mathbf{B}(t)$ and $\mathbf{C}(t)$, we solve the backward matrix Riccati ODE system derived from substituting the value function ansatz and optimal portfolio into the HJB equation:
+$\frac{d\mathbf{C}(t)}{dt}$ $=\kappa_{Y}^{\top}\mathbf{C}(t)+\mathbf{C}(t)\kappa_{Y}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}\mathbf{C}(t)$
+$\quad-\frac{1-\gamma}{\gamma}\left[(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right.$
+$\qquad\qquad\left.+\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})+\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right],$ (27)
+$\frac{d\mathbf{B}(t)}{dt}$ $=\left[\kappa_{Y}^{\top}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}-\frac{1-\gamma}{\gamma}\left(\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\right)\right]\mathbf{B}(t)$
+$\quad-\left[\mathbf{C}(t)\kappa_{Y}-\frac{1-\gamma}{\gamma}(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right]\bm{\theta}_{Y}.$ (28)
+subject to the terminal conditions $\mathbf{C}(T)=\mathbf{0}$ and $\mathbf{B}(T)=\mathbf{0}$.
+To obtain the complete value function (26), the scalar function $g(t)$ must also be determined by solving its own ODE backward from the terminal condition $g(T)=0$.
+This ODE, derived from the HJB equation by collecting terms independent of $\mathbf{Y}$ and consistent with the corrected single-factor case structure, is given by:
+$\frac{dg(t)}{dt}$ $=(1-\gamma)r$
+$\quad-\frac{1-\gamma}{2\gamma}\operatorname{Tr}\left((\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right)$
+$\quad-\frac{1-\gamma}{\gamma}\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})$
+$\quad-\frac{1-\gamma}{2\gamma}\left[\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma}\Sigma^{-1}\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{B}(t)+\operatorname{Tr}\left((\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma}\Sigma^{-1}\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right)\right]$
+$\quad+\mathbf{B}(t)^{\top}\kappa_{Y}\bm{\theta}_{Y}$
+$\quad+\frac{1}{2}\operatorname{Tr}(\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}).$ (29)
+Here, $\Phi=\mathrm{Cov}(d\mathbf{W}^{Y}_{t})/dt$ is the factor covariance matrix used in the factor dynamics $d\mathbf{Y}_{t}$, and $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix.
+Solving these coupled ODEs for $\mathbf{C}(t)$, $\mathbf{B}(t)$, and $g(t)$ backward in time from $T$ to $t$ and substituting the results into Eq.
+(26) gives the closed-form value function, while substituting $\mathbf{B}(t)$ and $\mathbf{C}(t)$ into $\bm{\pi}^{*}(t,\mathbf{Y})$ gives the optimal strategy, explicitly capturing both myopic and intertemporal hedging components in a multi-asset, multi-factor setting.
+Appendix C Detailed Experimental Setup
+This appendix provides the detailed setup for the numerical experiments presented in Section 4.
+C.1 Parameter Values
 The experiments consider the objective of maximizing the expected utility of terminal wealth $X_{T}$, $J=\mathbb{E}[U(X_{T})]$, using a CRRA utility function $U(x)=x^{1-\gamma}/(1-\gamma)$ with a relative risk aversion coefficient $\gamma=2.0$.
@@
     • Volatilities: $\bm{\sigma}\in\mathbb{R}^{n}$ components drawn uniformly from $U(0.1,0.5)$.
-    • Factor loadings: $\bm{\alpha}\in\mathbb{R}^{n\times k}$ drawn using ‘scipy.stats.dirichlet’ with concentration parameters $[1.0,\dots,1.0]$.
-(Note: These define the risk premium structure $\bm{\mu}=r\mathbf{1}+\mathrm{diag}(\bm{\sigma})\bm{\alpha}\mathbf{Y}_{t}$).
+    • Factor loadings ($\bm{\alpha}\in\mathbb{R}^{n\times k}$): For each asset $i=1,\dots,n$, the $k$-dimensional row vector of factor loadings $\bm{\alpha}_{i}^{\top}=(\alpha_{i1},\dots,\alpha_{ik})$ is drawn independently from a Dirichlet distribution.
+This is done using ‘scipy.stats.dirichlet‘ with a concentration parameter vector consisting of $k$ ones ($\mathbf{1}_{k}=[1.0,\dots,1.0]$).
+Consequently, for each asset $i$, all loadings are non-negative ($\alpha_{ij}\geq 0$) and they sum to unity ($\sum_{j=1}^{k}\alpha_{ij}=1$).
+(Note: These loadings define the risk premium structure $\bm{\mu}(t,\mathbf{Y}_{t})=r\mathbf{1}+\mathrm{diag}(\bm{\sigma})\bm{\alpha}\mathbf{Y}_{t}$).
   • Correlation structure: The generation process aims to produce realistic correlation parameters while ensuring the overall block correlation matrix remains positive definite, which is crucial for stable simulations.
@@
 The specific values generated by the seed are used consistently across experiments for given $n$ and $k$.
-4.1.2 PG-DPO Implementation Details
+C.2 PG-DPO Implementation Details
 We implement the Baseline PG-DPO algorithm using PyTorch.
@@
 autograd.grad’) applied to Monte Carlo estimates of the expected terminal utility starting from evaluation points $(t_{k},W_{k},\mathbf{Y}_{k})$.
-4.1.3 Benchmark Solution and Evaluation
+C.3 Benchmark Solution and Evaluation
 As mentioned, the chosen multi-factor OU model allows for a semi-analytical benchmark solution.
@@
   • Benchmark Computation: The system of ODEs (including a matrix Riccati equation) associated with the HJB equation for this terminal wealth problem (Appendix B.2) is solved numerically backwards from the fixed horizon $T$ using ‘scipy.integrate.solve_ivp’ (‘Radau’ method) to obtain the coefficients determining the value function and the optimal policy $\bm{\pi}^{*}(t,W_{t},\mathbf{Y}_{t})$.
-  • Interpolation: The benchmark policy (total, myopic, and hedging components) is pre-computed on a grid over $(t,W_{t},\mathbf{Y}_{t})$ space and stored using ‘scipy.
-interpolate.RegularGridInterpolator’ for efficient lookup.
+  • Interpolation: The benchmark policy (total, myopic, and hedging components) is pre-computed on a grid over $(t,W_{t},\mathbf{Y}_{t})$ space and stored using ‘scipy.interpolate.RegularGrid Interpolator’ for efficient lookup.
   • Evaluation Metrics: Performance is evaluated via the Root Mean Squared Error (RMSE) between the policies generated by (i) Baseline PG-DPO ($\bm{\pi}_{\theta}$) and the benchmark ($\bm{\pi}^{*}$), and (ii) Two-Stage PG-DPO (using Eq.
@@
 All computations are performed using PyTorch on an NVIDIA A6000 GPU.
-4.2 Convergence Speed and Accuracy Analysis
-This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO methodologies across various dimensional settings.
-Figure 1 displays the Root Mean Squared Error (RMSE) between the learned policies (either $\bm{\pi}_{\theta}$ for Baseline PG-DPO or $\bm{\pi}^{\mathrm{PMP}}$ for Two-Stage PG-DPO) and the benchmark optimal policy ($\bm{\pi}^{*}$) as a function of training epochs, plotted on a logarithmic scale.
-The results are shown for configurations of (a) n=1 asset, k=10 state variables, (b) n=10, k=10, and (c) n=50, k=10.
-(a) n=1, k=10
-(b) n=10, k=10
-(c) n=50, k=10
-Figure 1: RMSE comparison between Baseline PG-DPO (blue) and Two-Stage PG-DPO (orange) policies relative to the benchmark solution across different dimensions.
-Note the logarithmic scale on the y-axis.
-The RMSE for the Two-Stage PG-DPO (orange line in Figure 1) represents the error of the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ constructed by substituting the costate estimates obtained via BPTT at each epoch into the PMP first-order condition (Eq.
-(2)).
-A key advantage, and distinction from the Baseline PG-DPO which learns the total policy $\bm{\pi}_{\theta}$ monolithically, is that the Two-Stage approach explicitly computes the policy by combining the analytically distinct myopic and hedging components derived from the PMP formula.
-This allows for separate assessment of their accuracy.
-As established theoretically for affine models with CRRA utility (Appendix B) and confirmed in preliminary component-wise error analyses, the error associated with the myopic component (driven by the ratio $-\lambda_{t}^{*}/(X_{t}^{*}\partial_{x}\lambda_{t}^{*})=1/\gamma$) is typically negligible and converges almost instantly.
-Therefore, the total RMSE observed for the Two-Stage PG-DPO policy (orange line in Figure 1) is almost entirely attributable to the error in estimating the hedging component.
-The remarkably low overall error achieved by the Two-Stage method thus highlights its proficiency in accurately capturing this crucial, often complex, hedging demand using BPTT-derived costate derivatives ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$).
-The key finding, consistently observed across all cases presented in Figure 1, is that the Two-Stage PG-DPO (orange) converges significantly faster and achieves a substantially lower final (or minimum) RMSE compared to the Baseline PG-DPO (blue).
-This clearly demonstrates the superiority of the Two-Stage approach, which leverages costate estimates derived from BPTT to reconstruct the policy according to the PMP structure.
-Analyzing each case provides further insights:
-  • n=1, k=10 (Figure 1a): In this scenario with numerous state variables, the Two-Stage method converges rapidly within approximately 1,000 epochs to a very low RMSE level of around $3\times 10^{-3}$.
-In contrast, the Baseline method converges much slower, stabilizing after about 4,000 epochs at a significantly higher RMSE level of approximately $7\times 10^{-2}$.
-This results in a substantial gap (over 20-fold) in the final accuracy between the two methods.
-  • n=10, k=10 (Figure 1b): A similar trend persists in this intermediate-scale case with an increased number of assets.
-The Two-Stage method still converges quickly, reaching an RMSE of about $1.5\times 10^{-2}$ within roughly 2,000 epochs, whereas the Baseline method exhibits slower convergence and achieves a less accurate result with an RMSE around $1\times 10^{-1}$.
-  • n=50, k=10 (Figure 1c): For this most complex, high-dimensional case, training was conducted only up to 3,000 epochs.
-Nevertheless, the performance difference is stark.
-The Two-Stage method displays much faster initial convergence, reaching a minimum RMSE of approximately $1.5\times 10^{-2}$ near the 3,000-epoch mark.
-Although a slight tendency for the error to increase afterward is observed—potentially indicating the need for a longer warm-up phase (Stage 1) or suggesting instability in costate estimates requiring further training or tuning—its achieved accuracy is remarkable.
-Considering that the Baseline method’s RMSE is still above $2\times 10^{-1}$ at the same 3,000-epoch point, we can conclude that the Two-Stage approach yields a significantly more accurate policy within the same computational budget.
-These results strongly support the rationale behind the Two-Stage approach discussed in Section 3.3.
-Specifically, BPTT can effectively estimate the necessary costate information (particularly its derivatives) required by the PMP conditions with relatively few training iterations.
-Integrating these estimates into the analytical structure provided by the PMP optimality conditions proves to be more efficient and accurate than attempting to directly approximate the complex optimal policy function end-to-end using a neural network (Baseline approach), especially as the problem dimensionality increases.
-Building upon this quantitative analysis, the subsequent sections will provide visual comparisons of the policy functions for specific cases via contour plots, offering further insights into the qualitative differences between the policies generated by the two methods.
-Hereafter, we often abbreviate Two-Stage PG-DPO as 2-PG-DPO for brevity
-4.3 Detailed Analysis for the Single-Asset, Single-Factor Case (n=1, k=1)
-We begin with a detailed examination of the simplest state-dependent setting (n=1, k=1) to visually validate the methodologies and illustrate the structure of the learned policies.
-The results presented in this subsection correspond to the state of the policies after 2,000 training epochs, a point observed to be near optimal for the 2-PG-DPO method in this configuration (cf.
-Figure 1(a)) before potential performance degradation with extended training, as discussed later.
-Figure 2 compares the total policies learned by the Baseline PG-DPO and 2-PG-DPO methods, alongside their respective errors relative to the benchmark solution.
-Figure 3 then decomposes the superior 2-PG-DPO policy into its constituent myopic and hedging components, showing their individual shapes and associated errors.
-All plots show the policy or error values against time-to-maturity ($T-t$) and the state variable ($Y$).
-(a) Baseline Policy ($\pi_{\theta}$)
-(b) 2-PG-DPO Policy ($\pi^{\mathrm{PMP}}$)
-(c) Baseline Error
-(d) 2-PG-DPO Error
-Figure 2: Comparison of total policies and errors for the n=1, k=1 case at epoch 2000.
-Top row compares the policies from Baseline PG-DPO (left) and 2-PG-DPO (right).
-Bottom row compares their respective errors.
-Note the significantly different color bar scales for the error plots.
-(a) 2-PG-DPO Myopic ($\pi_{0}^{\mathrm{PMP}}$)
-(b) 2-PG-DPO Hedging ($\pi_{1}^{\mathrm{PMP}}$)
-(c) Myopic Error
-(d) Hedging Error
-Figure 3: Decomposition of the 2-PG-DPO policy for the n=1, k=1 case at epoch 2000.
-Top row shows the myopic (left) and hedging (right) components.
-Bottom row shows their respective errors.
-Note the extremely small error scale for the myopic component.
-Observing Figure 2, we compare the Baseline PG-DPO policy (2(a)) and its error (2(c)) with the 2-PG-DPO policy (2(b)) and its error (2(d)).
-While the Baseline network learns the general dependency on the state variable $Y$, its error plot (2(c)) reveals significant deviations from the benchmark, with magnitudes reaching up to ±$\approx\pm 0.1$.
-This visual representation aligns with the relatively high RMSE seen for the Baseline method around epoch 2000 in Figure 1(a).
-Although this error level is considerably higher than that of the 2-PG-DPO method even in this simple n=1, k=1 setting, it represents the performance in the simplest possible setting.
-It is crucial to emphasize that the effectiveness of the Baseline PG-DPO approach tends to diminish considerably as the number of assets ($n$) and state variables ($k$) grows.
-The challenge of directly approximating the increasingly complex optimal policy function often becomes intractable for the Baseline method in higher dimensions, leading to much larger errors or convergence difficulties, as hinted by the widening gap in RMSE observed in Figure 1 for larger $n$.
-Therefore, while the Baseline method shows some learning capability here, its limitations in scalability motivate the need for more robust approaches like 2-PG-DPO for the high-dimensional problems central to this paper.
-In stark contrast, the 2-PG-DPO policy (2(b)) yields a drastically reduced error (2(d)), with magnitudes mostly within ±$\approx\pm 0.001$, corresponding well with its low RMSE and suggesting close proximity to the true optimal solution, demonstrating its effectiveness even in this simple case.
-Figure 3 provides crucial insights into the superior performance of the 2-PG-DPO method by decomposing its policy.
-The myopic component (3(a)) is visually dominant.
-Crucially, its associated error (3(c)) is extremely small (order $10^{-7}$), confirming its near-perfect accuracy.
-The hedging component (3(b)), representing a smaller adjustment, has an associated error (3(d)) that is visually identical in pattern and magnitude (order $10^{-3}$) to the total error of the 2-PG-DPO policy (2(d)).
-This visually confirms that the overall accuracy of the 2-PG-DPO policy is primarily limited by the accuracy of estimating the hedging demand.
-Finally, it is important to address the training duration and the potential for overfitting with the 2-PG-DPO method.
-While this method achieves high accuracy remarkably quickly (around 2000 epochs here, corresponding to only about 5-10 minutes of computation time, given that 10,000 epochs take approximately 30 minutes on the utilized NVIDIA A6000 GPU), we observed that continued training beyond this near-optimal point can lead to a degradation of its performance (i.e., an increase in RMSE, as hinted in Figure 1(a)).
-This phenomenon might suggest that the underlying neural network used for costate estimation during the warm-up phase (Stage 1) could potentially start to overfit to the specific discrete time steps ($\Delta t>0$) and sample paths inherent in the simulation-based training[cite: 293].
-Instead of learning the true continuous-time ($dt$) value function derivatives needed for the PMP formulas, the network might adapt too closely to the artifacts of the discretization.
-This underscores the importance of careful monitoring and potentially early stopping for the warm-up phase—which requires only a short computational time—to extract the best performance from the 2-PG-DPO methodology.
-In summary, the analysis for the n=1, k=1 case visually confirms the high accuracy of the 2-PG-DPO approach compared to the Baseline PG-DPO, clearly illustrates the decomposition into highly accurate myopic and less accurate (but still relatively small error) hedging components, and underscores the significance of selecting an appropriate training duration for optimal performance.
-4.4 Analysis with Multiple State Variables (n=1, k=10)
-We now increase the complexity by considering a case with a single asset but numerous state variables (n=1, k=10).
-This scenario tests the ability of the methods to handle high-dimensional state dependencies that affect the investment opportunity set.
-As visualizing the full 10-dimensional state dependency is impossible, the following plots (Figures 4 and 5) display the policy and error values as functions of time-to-maturity ($T-t$) and the first state variable ($Y_{1}$), while holding the remaining state variables ($Y_{2},\dots,Y_{10}$) constant at their respective long-term mean values ($\theta_{Y,i}$).
-The results shown correspond to 2,000 training epochs, consistent with the previous analysis.
-The overall convergence for this case is depicted in Figure 1(a).
-(a) Baseline Policy ($\pi_{\theta}$)
-(b) 2-PG-DPO Policy ($\pi^{\mathrm{PMP}}$)
-(c) Baseline Error
-(d) 2-PG-DPO Error
-Figure 4: Comparison of total policies and errors for the n=1, k=10 case at epoch 2000 (plotted against $Y_{1}$).
-Top row compares the policies from Baseline PG-DPO (left) and 2-PG-DPO (right).
-Bottom row compares their respective errors.
-Note the significantly different color bar scales for the error plots.
-(a) 2-PG-DPO Myopic ($\pi_{0}^{\mathrm{PMP}}$)
-(b) 2-PG-DPO Hedging ($\pi_{1}^{\mathrm{PMP}}$)
-(c) Myopic Error
-(d) Hedging Error
-Figure 5: Decomposition of the 2-PG-DPO policy for the n=1, k=10 case at epoch 2000 (plotted against $Y_{1}$).
-Top row shows the myopic (left) and hedging (right) components.
-Bottom row shows their respective errors.
-Note the extremely small error scale for the myopic component and the very small magnitude of the hedging component itself.
-Figure 4 presents the comparison for the n=1, k=10 case.
-The Baseline PG-DPO policy (4(a)) again learns the dependency on the displayed state variable $Y_{1}$.
-However, its error relative to the benchmark (4(c)) remains significant, with magnitudes ranging roughly between 0.001 and 0.064.
-This level of error is consistent with the high RMSE value observed for the Baseline method in Figure 1(a), indicating difficulties in accurately capturing the complex policy function influenced by ten state variables.
-This difficulty highlights a key challenge for the Baseline PG-DPO approach, particularly when the number of state variables ($k$) is large.
-The network must learn to approximate the total policy, which implicitly combines the primary myopic demand with the complex intertemporal hedging demands arising from all ten state variables ($Y_{1},\dots,Y_{10}$).
-Even if the magnitude of the hedging component related to any single factor or the total hedging demand itself is relatively small (as suggested by Figure 5(b) for the $Y_{1}$ dimension), the sheer number of interacting hedging influences creates a highly intricate functional relationship for the optimal policy.
-Attempting to capture this subtle, high-dimensional structure directly through gradient updates on a single policy network can disrupt the learning process.
-This may hinder convergence towards the true optimum or even lead to the network converging to an inaccurate or distorted policy shape, as potentially reflected in the structured errors seen in Figure 4(c).
-Conversely, the 2-PG-DPO method, by explicitly utilizing the PMP structure and separating the estimation of terms related to $\lambda$ and its derivatives ($\partial_{\mathbf{Y}}\lambda$), demonstrates remarkable accuracy even in this high-dimensional state setting.
-The total policy (4(b)) results in an extremely small error (4(d)), with magnitudes only reaching up to approximately $3.6\times 10^{-4}$.
-This visually confirms the near-perfect fit suggested by the very low RMSE value achieved by the 2-PG-DPO method shown in Figure 1(a), showcasing its ability to effectively handle a large number of state factors.
-Figure 5 provides further validation through the decomposition of the 2-PG-DPO policy.
-The myopic component (5(a)) largely determines the shape of the total policy.
-Its corresponding error (5(c)) remains negligible, on the order of $10^{-7}$, proving its accurate estimation.
-The hedging component (5(b)) is exceptionally small in this particular slice (magnitude order $10^{-4}$), suggesting that for this specific configuration and visualization plane, the intertemporal hedging demand related to $Y_{1}$ is minimal.
-Nevertheless, the error associated with this tiny hedging component (5(d)) is again visually identical to the total error of the 2-PG-DPO policy (4(d)), confirming that the overall policy accuracy is still dictated by the precision of the hedging term estimation, even when this term (and its error) is extremely small.
-In conclusion, the results for the n=1, k=10 case demonstrate the robustness and high accuracy of the 2-PG-DPO method when faced with high-dimensional state spaces.
-It significantly outperforms the Baseline PG-DPO approach and accurately captures the policy structure dictated by PMP, with negligible error in the myopic component and the total error determined by the precision of the hedging component estimation.
-This highlights the suitability of the 2-PG-DPO framework for problems involving complex, multi-factor dynamics influencing investment decisions.
-4.5 High-Dimensional Performance (n=50, k=10)
-Finally, we evaluate the performance in the most challenging setting considered: a high-dimensional problem with n=50 assets and k=10 state variables.
-This scenario fully tests the scalability and effectiveness of the proposed methods in dimensions where traditional DP methods are intractable.
-Similar to the previous section, the visualizations focus on the policy for the first asset ($i=1$) as a function of time-to-maturity ($T-t$) and the first state variable ($Y_{1}$), holding other state variables ($Y_{2},\dots,Y_{10}$) constant at their long-term means ($\theta_{Y,i}$).
-The results shown correspond to the state after 3,000 training epochs, representing the point where the 2-PG-DPO method achieved its minimum RMSE in Figure 1(c).
-(a) Baseline Policy ($\pi_{1,\theta}$)
-(b) 2-PG-DPO Policy ($\pi_{1}^{\mathrm{PMP}}$)
-(c) Baseline Error
-(d) 2-PG-DPO Error
-Figure 6: Comparison of total policies and errors for the first asset ($i=1$) in the n=50, k=10 case at epoch 3000 (plotted against $Y_{1}$).
-Top row compares the policies from Baseline PG-DPO (left) and 2-PG-DPO (right).
-Bottom row compares their respective errors.
-Note the dramatic difference in error scales and magnitudes.
-(a) 2-PG-DPO Myopic ($\pi_{1,0}^{\mathrm{PMP}}$)
-(b) 2-PG-DPO Hedging ($\pi_{1,1}^{\mathrm{PMP}}$)
-(c) Myopic Error
-(d) Hedging Error
-Figure 7: Decomposition of the 2-PG-DPO policy for the first asset ($i=1$) in the n=50, k=10 case at epoch 3000 (plotted against $Y_{1}$).
-Top row shows the myopic (left) and hedging (right) components.
-Bottom row shows their respective errors.
-The myopic error remains negligible, while the hedging error dictates the total error.
-The challenges of high dimensionality become starkly evident for the Baseline PG-DPO method in Figure 6.
-While the learned policy for the first asset (6(a)) shows some structure, the corresponding error (6(c)) is extremely large, with values ranging from approximately -1.1 to -0.7.
-This indicates severe inaccuracies and potentially a systematic bias in the policy learned by the Baseline method when faced with the combined complexity of 50 assets and 10 state variables.
-This poor performance aligns with the high and slowly converging RMSE observed in Figure 1(c).
-In dramatic contrast, the 2-PG-DPO method continues to provide a highly accurate solution, as shown by its policy (6(b)) and error (6(d)).
-The error magnitude for the 2-PG-DPO policy is approximately within ±$\pm 0.02$.
-While this represents a slight increase in error compared to the lower-dimensional cases (e.g., n=1, k=10 where the error was order $10^{-4}$), it is still orders of magnitude smaller (roughly 50 times smaller) than the error of the Baseline PG-DPO method for this same high-dimensional problem.
-This underscores the robustness and practical necessity of the 2-PG-DPO approach; although perfect accuracy might be harder to achieve as dimensionality increases, it remains remarkably effective and vastly superior to the direct policy learning method.
-The error level visually corresponds to the minimum RMSE achieved by the 2-PG-DPO method in Figure 1(c).
-The component decomposition in Figure 7 confirms the patterns observed previously, even in this complex setting.
-The myopic component (7(a)) still largely dictates the policy shape, and its associated error (7(c)) remains negligible (order $10^{-7}$), demonstrating the consistent accuracy of its estimation across all dimensions tested.
-The hedging component (7(b)) appears relatively more significant here compared to the n=1, k=10 case, contributing visibly to the total policy shape.
-Crucially, the hedging error (7(d)) is again visually identical to the total 2-PG-DPO error (6(d)), confirming that the overall accuracy is dictated by the precision of estimating the hedging demands, which becomes slightly more challenging, yet remains manageable, in higher dimensions.
-In conclusion, the n=50, k=10 results powerfully demonstrate the capability of the 2-PG-DPO framework to “break the dimensional barrier”.
-While the Baseline PG-DPO struggles significantly, 2-PG-DPO delivers a highly accurate and viable solution by leveraging the structure of Pontryagin’s Maximum Principle.
-Despite a slight, expected increase in error compared to lower-dimensional settings, its performance remains remarkably strong and vastly superior to the baseline, solidifying its potential as a practical tool for large-scale, continuous-time portfolio optimization.
-5 Conclusion
-In this paper, we addressed the challenge of solving high-dimensional continuous-time portfolio optimization problems, which often suffer from the “curse of dimensionality” when approached with traditional dynamic programming methods.
-We developed and implemented the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, which leverages Pontryagin’s Maximum Principle (PMP) and deep learning techniques, specifically backpropagation-through-time (BPTT), to directly optimize neural network policies for investment and consumption.
-A key contribution is the introduction of a highly effective “Two-Stage” PG-DPO variant.
-This approach utilizes a short warm-up phase to obtain stable estimates of the crucial Pontryagin costates and their derivatives via BPTT, potentially refined through Monte Carlo averaging.
-These estimates are then plugged directly into the analytical first-order optimality conditions derived from PMP to construct the control policy.
-Our numerical experiments, conducted on a multi-asset, multi-factor affine model allowing for benchmark comparisons, demonstrated that the 2-PG-DPO method significantly outperforms the baseline end-to-end approach.
-It achieves substantially higher accuracy and faster convergence, successfully scaling to problems with dimensions (e.g., n=50 assets, k=10 state factors) previously considered intractable for numerical HJB/PDE solvers, thereby effectively ”breaking the dimensional barrier.” We specifically highlighted the method’s ability to accurately capture both myopic and complex intertemporal hedging demands by leveraging the PMP structure.
-This research opens several avenues for future investigation.
-Firstly, exploring the theoretical underpinnings further is crucial.
-While Proposition 3.1 provides heuristic support for the 2-PG-DPO approach in structured models, rigorously analyzing or establishing similar derivative stability properties for the broader class of non-affine models would significantly strengthen the foundation of the method.
-Furthermore, investigating the theoretical convergence properties of the PG-DPO algorithm, particularly the costate estimation process within the 2-PG-DPO variant, remains an important open question.
-Secondly, extending the framework to incorporate more realistic market features is a natural next step.
-This includes introducing portfolio constraints (such as short-selling restrictions or leverage limits) and modeling transaction costs, which significantly alter the optimal control problem.
-Adapting the PG-DPO framework to handle these complexities efficiently would enhance its practical applicability.
-Testing the framework’s robustness and performance with different utility functions beyond CRRA (e.g., HARA, prospect theory) is also warranted.
-Thirdly, further empirical validation and algorithmic refinements are desirable.
-Applying the PG-DPO framework (especially the superior 2-PG-DPO variant) to well-known non-affine financial models (e.g., stochastic volatility models), even in the absence of exact benchmarks, would showcase its capabilities more broadly.
-Conducting numerical experiments that explicitly include the optimization and evaluation of the intermediate consumption policy, which was omitted in the current study for benchmarking purposes, would provide a more complete assessment of the full framework.
-Exploring alternative neural network architectures, optimization techniques, and variance reduction methods could potentially lead to further performance improvements.
-Finally, optimizing the implementation details, such as the warm-up duration for the 2-PG-DPO method to balance costate accuracy against potential overfitting to simulation artifacts, deserves further attention.
-Overall, the Pontryagin-Guided Direct Policy Optimization framework, particularly its Two-Stage variant, offers a promising and scalable approach for tackling complex, high-dimensional optimal control problems in continuous-time finance.
-Acknowledgments
-This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-00562904).
-Appendix A Derivation of the Adjoint Diffusion Coefficient via Itô’s Lemma
-In the Pontryagin framework, the adjoint (or costate) process associated with the investor’s wealth $X_{t}$ is identified with the partial derivative of the value function $V$.
-Concretely, we set $\lambda_{t}=V_{X}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},$ where $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ represents additional state variables and $X_{t}$ is the investor’s scalar wealth process.
-Our goal in this section is to derive the diffusion coefficient corresponding to the $d\mathbf{W}^{X}_{t}$ term in the backward SDE for $\lambda_{t}$, namely $\mathbf{Z}_{t}^{\lambda X}$ from BSDE (2), which plays a crucial role in determining the optimal portfolio via FOC (4).
-Recall that under an optimal control $\bigl{(}\bm{\pi}_{t}^{*},C_{t}^{*}\bigr{)}$, the wealth $X_{t}^{*}$ evolves according to Eq.
-(1), and the $k$-dimensional state process $\mathbf{Y}_{t}^{*}$ follows its own SDE, also presented in Section 2.2.
-The dynamics are driven by correlated Brownian motions $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ and $\mathbf{W}_{t}^{Y}\in\mathbb{R}^{k}$ with the joint covariance structure $\operatorname{Cov}(d\mathbf{W}_{t})=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}dt.$
-Since $\lambda_{t}=V_{X}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},$ we apply Ito’s multidimensional lemma to $V_{X}$. The differential $d\lambda_{t}$ can be expressed as:
-$d\lambda_{t}edundant$ $=\;V_{Xt}\,dt\;+\;V_{XX}\,dX_{t}\;+\;V_{X\mathbf{Y}}\,d\mathbf{Y}_{t}$
-$\quad+\;\tfrac{1}{2}\,V_{XXX}\,\bigl{(}dX_{t}\bigr{)}^{2}\;+\;V_{XX\mathbf{Y}}\,(dX_{t})\,(d\mathbf{Y}_{t})^{\top}\;+\;\tfrac{1}{2}\,\text{Tr}\bigl{(}V_{X\mathbf{Y}\mathbf{Y}}\,(d\mathbf{Y}_{t})(d\mathbf{Y}_{t})^{\top}\bigr{)},$ (8)
-where subscripts denote partial derivatives (e.g., $V_{X\mathbf{Y}}$ is the gradient w.r.t $\mathbf{Y}$ of $V_{X}$, $V_{X\mathbf{Y}\mathbf{Y}}$ is the Hessian w.r.t $\mathbf{Y}$ of $V_{X}$).
-Our objective is to identify the coefficient of $d\mathbf{W}_{t}^{X}$ in the expansion of $d\lambda_{t}$.
-This coefficient is precisely the process $\mathbf{Z}_{t}^{\lambda X}$ appearing in the BSDE (2).
-Substituting the SDEs for $dX_{t}$ and $d\mathbf{Y}_{t}$ (from Section 2.2) into the Ito formula (8), we observe that terms involving $d\mathbf{W}_{t}^{X}$ arise directly from the diffusion part of the $V_{XX}\,dX_{t}$ term, and indirectly from the $V_{X\mathbf{Y}}\,d\mathbf{Y}_{t}$ term due to the correlation $\rho$ between $d\mathbf{W}_{t}^{X}$ and $d\mathbf{W}_{t}^{Y}$.
-A careful application of Ito’s lemma, considering the quadratic variation and covariation terms related to the covariance matrix $\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}$, allows isolation of the full coefficient multiplying $d\mathbf{W}_{t}^{X}$.
-Representing $\mathbf{Z}_{t}^{\lambda X}$ as a $1\times n$ row vector, this process is found to be:
-$\mathbf{Z}_{t}^{\lambda X}=\underbrace{(\partial_{x}\lambda_{t}^{*})X_{t}^{*}\bm{\pi}_{t}^{*\top}\bm{\sigma}(t,\mathbf{Y}_{t}^{*})}_{\text{Term from }V_{XX}dX_{t}}+\underbrace{(\partial_{\mathbf{Y}}\lambda_{t}^{*})\bm{\sigma}_{Y}(t,\mathbf{Y}_{t}^{*})\rho^{\top}\Psi^{-1}}_{\text{Term from }V_{XY}dY_{t}\text{ via correlation}}.$ (9)
-The first term reflects the direct impact of wealth volatility related to the portfolio choice $\bm{\pi}_{t}^{*}$, scaled by the second derivative of the value function with respect to wealth, $V_{XX}$ (or $\partial_{x}\lambda_{t}^{*}$).
-The second term captures the intertemporal hedging component arising from the correlation between asset returns and state variable movements, mediated by the cross-derivative of the value function $V_{X\mathbf{Y}}$ (or $\partial_{\mathbf{Y}}\lambda_{t}^{*}$).
-This derivation provides the explicit form of the adjoint diffusion coefficient $\mathbf{Z}_{t}^{\lambda X}$, which is fundamental for understanding the structure of the optimal portfolio policy (2) and for constructing the Pontryagin adjoint equations.
-Appendix B Analytic Solutions for Affine Models
-This appendix provides details on the derivation of analytic solutions for the portfolio optimization problems presented in Section 2.3 under specific assumptions (no intermediate consumption, CRRA utility), typically following the Hamilton-Jacobi-Bellman (HJB) approach with an exponential-quadratic value function ansatz.
-We also show how these solutions connect to the Pontryagin framework.
-B.1 Analytic Solution for the Multi-Asset OU Model (Section 2.3.1)
-For simplicity, we assume no intermediate consumption, i.e., $C_{t}=0$, and consider an investor who derives utility only from terminal wealth $X_{T}$ under a CRRA objective $U(x)=\frac{x^{1-\gamma}}{1-\gamma}$.
-Following Kim & Omberg, (1996), the associated Hamilton–Jacobi–Bellman (HJB) equation can be solved by positing a value function of the form $V(t,x,Y)=\frac{x^{1-\gamma}}{1-\gamma}\exp\left(g(t)+B(t)Y+\frac{1}{2}C(t)Y^{2}\right)$.
-Here, $g(t)$, $B(t)$, and $C(t)$ are deterministic functions of time determined by solving ordinary differential equations (ODEs) derived from the HJB equation, subject to terminal conditions $g(T)=0$, $B(T)=0$, and $C(T)=0$.
-From this value function, the costate $\lambda_{t}^{*}=V_{x}$, its derivative with respect to wealth $\partial_{x}\lambda_{t}^{*}=V_{xx}$, and its derivative with respect to the state $\partial_{Y}\lambda_{t}^{*}=V_{xY}$ have the following closed-form expressions on the optimal path $(X_{t}^{*},Y_{t})$:
-$\lambda_{t}^{*}(t,X_{t}^{*},Y_{t})$ $=(X_{t}^{*})^{-\gamma}\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right),$
-$\frac{\partial\lambda_{t}^{*}}{\partial x}(t,X_{t}^{*},Y_{t})$ $=-\gamma(X_{t}^{*})^{-\gamma-1}\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right),$
-$\frac{\partial\lambda_{t}^{*}}{\partial Y}(t,X_{t}^{*},Y_{t})$ $=(X_{t}^{*})^{-\gamma}\left(B(t)+C(t)Y_{t}\right)\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right).$
-These expressions lead to remarkably simple ratios that are central to the Pontryagin framework’s characterization of the optimal portfolio:
-$\frac{\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma},\qquad\frac{\partial_{Y}\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma}\left(B(t)+C(t)Y_{t}\right).$
-Substituting these ratios into the general Pontryagin first-order condition for the optimal portfolio $\bm{\pi}_{t}^{*}$ (which equates the marginal contribution to the Hamiltonian from investing in risky assets to zero, cf.
-Eq.
-(4) derived in Section 2.2) confirms the explicit analytic solution derived via the HJB approach:
-$\bm{\pi}^{*}(t,Y)=\frac{1}{\gamma}\,\Sigma^{-1}\Bigl{[}\underbrace{(\bm{\alpha}\,\bm{\sigma})\,Y}_{\text{Risk premium }\bm{\mu}(Y)-r\mathbf{1}}+\underbrace{\sigma_{Y}\,(\bm{\sigma}\,\bm{\rho})\,\bigl{(}B(t)+C(t)\,Y\bigr{)}}_{\text{Hedging demand from }\partial_{Y}\lambda_{t}^{*}}\Bigr{]},$
-where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$, $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$, the notation $(\bm{\alpha}\,\bm{\sigma})$ represents the vector $(\alpha_{1}\sigma_{1},\dots,\alpha_{n}\sigma_{n})^{\top}$, and $(\bm{\sigma}\,\bm{\rho})$ represents the vector $(\sigma_{1}\rho_{1},\dots,\sigma_{n}\rho_{n})^{\top}$.
-The derivation confirms that the first term captures the myopic demand driven by the instantaneous risk premium $(\bm{\mu}(Y)-r\mathbf{1})=(\bm{\alpha}\,\bm{\sigma})Y$, while the second term represents the intertemporal hedging demand, proportional to $B(t)+C(t)Y$ which arises from the state-dependency of the investment opportunity set captured by $\partial_{Y}\lambda_{t}^{*}$.
-Substituting this $\bm{\pi}^{*}$ back into the HJB equation and matching coefficients in powers of $Y$ yields the following Riccati-type ODEs for $B(t)$ and $C(t)$:
-$B(T)=0,\quad C(T)=0,$
-and
-$\frac{dC(t)}{dt}$ $=2\,\kappa_{Y}\,C(t)-\sigma_{Y}^{2}\,C(t)^{2}-\frac{1-\gamma}{\gamma}\Bigl{[}\beta_{0}+2\,\sigma_{Y}\,C(t)\,\beta_{1}+\sigma_{Y}^{2}\,C(t)^{2}\,\beta_{2}\Bigr{]},$
-$\frac{dB(t)}{dt}$ $=\kappa_{Y}\,B(t)-\sigma_{Y}^{2}\,B(t)\,C(t)+\frac{1-\gamma}{\gamma}\Bigl{[}\sigma_{Y}^{2}\,B(t)\,C(t)\,\beta_{2}+\sigma_{Y}\,B(t)\,\beta_{1}\Bigr{]}-\kappa_{Y}\,\theta_{Y}\,C(t),$
-where $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ are quadratic forms involving the model parameters:
-$\beta_{0}:=(\bm{\alpha}\,\bm{\sigma})^{\top}\,\Sigma^{-1}\,(\bm{\alpha}\,\bm{\sigma}),\quad\beta_{1}:=(\bm{\alpha}\,\bm{\sigma})^{\top}\,\Sigma^{-1}\,(\bm{\sigma}\,\bm{\rho}),\quad\beta_{2}:=(\bm{\sigma}\,\bm{\rho})^{\top}\,\Sigma^{-1}\,(\bm{\sigma}\,\bm{\rho}).$
-Once $B(t)$ and $C(t)$ are obtained by backward integration from their terminal conditions, substituting them into $\bm{\pi}^{*}(t,Y)$ yields the closed-form optimal strategy balancing myopic and hedging demands.
-The function $g(t)$ can also be found by integrating another ODE involving $B(t)$ and $C(t)$, but it does not affect the optimal policy $\bm{\pi}^{*}$.
-B.2 Analytic Solution for the Multi-Asset Multi-Factor Model (Section 2.3.2)
-[…]
-[…]
-where $\mathbf{B}(t)$ is an $n\times 1$ vector, $\mathbf{C}(t)$ is an $n\times n$ symmetric matrix, and $g(t)$ is a scalar function, all satisfying ODEs derived from the HJB equation, with terminal conditions $g(T)=0$, $\mathbf{B}(T)=\mathbf{0}$ and $\mathbf{C}(T)=\mathbf{0}$.
-Define the relevant matrices and vectors:
-$\bm{\sigma}\;=\;\mathrm{diag}(\sigma_{1},\dots,\sigma_{d}),\quad\bm{\sigma}_{Y}\;=\;\mathrm{diag}(\sigma_{Y,1},\dots,\sigma_{Y,n}),\quad\Sigma\;=\;\bm{\sigma}\,\Psi\,\bm{\sigma},$
-where $\Psi\in\mathbb{R}^{d\times d}$ is the correlation matrix for the $d$ risky assets, and let $\rho\in\mathbb{R}^{d\times n}$ be the cross-correlation matrix between these assets and the $n$-dimensional factor Brownian motion $\mathbf{W}^{Y}_{t}$.
-Also let $\mathbf{A}=\bigl{(}\bm{\alpha}_{1},\dots,\bm{\alpha}_{d}\bigr{)}^{\top}\in\mathbb{R}^{d\times n}$ collect the factor loadings $\bm{\alpha}_{i}\in\mathbb{R}^{n}$ for each asset $i$.
-The costate $\lambda_{t}^{*}=V_{x}$, its derivative with respect to wealth $\partial_{x}\lambda_{t}^{*}=V_{xx}$, and its gradient ($n\times 1$ vector) with respect to the state vector $\partial_{\mathbf{Y}}\lambda_{t}^{*}=V_{x\mathbf{Y}}$ are given by:
-$\lambda_{t}^{*}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=(X_{t}^{*})^{-\gamma}\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right),$
-$\frac{\partial\lambda_{t}^{*}}{\partial x}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=-\gamma(X_{t}^{*})^{-\gamma-1}\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right),$
-$\frac{\partial\lambda_{t}^{*}}{\partial\mathbf{Y}}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=(X_{t}^{*})^{-\gamma}\left(\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}_{t}\right)\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right).$
-These yield the following scalar and vector ratios, which simplify the PMP optimality conditions:
-$\frac{\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma},\qquad\frac{\partial_{\mathbf{Y}}\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma}\left(\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}_{t}\right).$
-Substituting these into the vector form of the PMP first-order condition for the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{d}$, (cf. Eq. (2) derived in Section 2.2):
-$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\Sigma^{-1}\left\{\lambda_{t}^{*}(\bm{\mu}(\mathbf{Y}_{t})-r\mathbf{1})+\bm{\sigma}\rho\bm{\sigma}_{Y}\left[\partial_{\mathbf{Y}}\lambda_{t}^{*}\right]\right\},$
-where the risk premium vector is $\bm{\mu}(\mathbf{Y}_{t})-r\mathbf{1}=\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}$, we precisely recover the analytic solution obtained from the HJB framework:
-$\bm{\pi}^{*}(t,\mathbf{Y})\;=\;\frac{1}{\gamma}\,\Sigma^{-1}\Bigl{[}\,\underbrace{\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}}_{\text{Myopic component (related to }\bm{\mu}-r\mathbf{1})}\;+\;\underbrace{\bm{\sigma}\,\rho\,\bm{\sigma}_{Y}\Bigl{(}\mathbf{B}(t)\;+\;\mathbf{C}(t)\,\mathbf{Y}\Bigr{)}}_{\text{Hedging component (related to }\partial_{\mathbf{Y}}\lambda_{t}^{*})}\Bigr{]}.$
-Here, the first bracketed term corresponds to the myopic component, driven by the instantaneous expected excess returns $\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}$.
-The second bracketed term incorporates intertemporal hedging demands arising from the desire to hedge against unfavorable changes in the investment opportunity set, as captured by the factor sensitivities $\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}$ (proportional to the gradient $\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and the correlations encoded in $\rho$, $\Psi$, and $\Phi$ (implicitly via the ODEs).
-To determine the time-dependent coefficients $\mathbf{B}(t)$ and $\mathbf{C}(t)$, we solve the backward matrix Riccati ODE system derived from substituting the value function ansatz and optimal portfolio into the HJB equation:
-$\frac{d\mathbf{C}(t)}{dt}$ $=\kappa_{y}^{\top}\mathbf{C}(t)+\mathbf{C}(t)\kappa_{y}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}\mathbf{C}(t)$
-$\quad-\frac{1-\gamma}{\gamma}\left[(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right.$
-$\qquad\qquad\left.+\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})+\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right],$
-$\frac{d\mathbf{B}(t)}{dt}$ $=\left[\kappa_{y}^{\top}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}-\frac{1-\gamma}{\gamma}\left(\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\right)\right]\mathbf{B}(t)$
-$\quad-\left[\mathbf{C}(t)\kappa_{y}-\frac{1-\gamma}{\gamma}(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right]\bm{\theta}_{y}.$
-subject to the terminal conditions $\mathbf{C}(T)=\mathbf{0}$ and $\mathbf{B}(T)=\mathbf{0}$.
-The scalar function $g(t)$ satisfies its own ODE, which depends on $\mathbf{B}(t)$ and $\mathbf{C}(t)$, but is not needed for determining the optimal portfolio.
-Solving these ODEs backward in time from $T$ to $t$ and then substituting the resulting $\mathbf{B}(t)$ and $\mathbf{C}(t)$ into $\bm{\pi}^{*}(t,\mathbf{Y})$ gives the closed-form optimal strategy, explicitly capturing both myopic and intertemporal hedging components in a multi-asset, multi-factor setting.
-Appendix C Proof Sketch for Proposition 3.1
-We provide a proof sketch for Proposition 3.1 focusing on the affine/quadratic structure considered by Liu (2007) (Liu07) with CRRA utility $U(w)=w^{1-\gamma}/(1-\gamma)$ (or $\ln w$).
-The argument relies on the specific form of the value function and the relationship between function closeness and coefficient closeness for polynomials.
-Assume the true value function $V^{*}$ and an approximation $\tilde{V}$ both share the exponential-quadratic structure on a compact domain $\mathcal{D}=[t_{0},T]\times[w_{min},w_{max}]\times\mathcal{Y}$ (where $\mathcal{Y}$ is a compact subset of $\mathbb{R}^{k}$):
-$V^{*}(t,w,\mathbf{y})$ $=U(w)\cdot\Phi^{*}(\mathbf{y},\tau)$
-$\tilde{V}(t,w,\mathbf{y})$ $=U(w)\cdot\tilde{\Phi}(\mathbf{y},\tau)$
-where $\tau=T-t$ and the state-dependent part has the form (adjusting notation slightly from Liu for clarity):
-$\Phi^{*}(\mathbf{y},\tau)$ $=\exp\left[c^{*}(\tau)+d^{*}(\tau)^{\top}\mathbf{y}+\frac{1}{2}\mathbf{y}^{\top}Q^{*}(\tau)\mathbf{y}\right]$
-$\tilde{\Phi}(\mathbf{y},\tau)$ $=\exp\left[\tilde{c}(\tau)+\tilde{d}(\tau)^{\top}\mathbf{y}+\frac{1}{2}\mathbf{y}^{\top}\tilde{Q}(\tau)\mathbf{y}\right]$
-Here, $c,\tilde{c}$ are scalars, $d,\tilde{d}\in\mathbb{R}^{k}$, and $Q,\tilde{Q}$ are symmetric $k\times k$ matrices, all depending smoothly on $\tau$ and the underlying model parameters.
-The condition $\sup_{\mathcal{D}}|V^{*}-\tilde{V}|\to 0$ implies $\sup_{\tau,\mathbf{y}\in\mathcal{Y}}|\Phi^{*}(\mathbf{y},\tau)-\tilde{\Phi}(\mathbf{y},\tau)|\to 0$, since $U(w)$ is bounded and bounded away from zero on $[w_{min},w_{max}]$.
-Because $\Phi^{*}$ and $\tilde{\Phi}$ are positive and continuous, this uniform convergence implies uniform convergence of their logarithms:
-$\sup_{\tau,\mathbf{y}\in\mathcal{Y}}|\ln\Phi^{*}(\mathbf{y},\tau)-\ln\tilde{\Phi}(\mathbf{y},\tau)|\to 0$
-This means the quadratic polynomials in the exponents are uniformly close:
-$\sup_{\tau,\mathbf{y}\in\mathcal{Y}}|(c^{*}+d^{*\top}\mathbf{y}+\frac{1}{2}\mathbf{y}^{\top}Q^{*}\mathbf{y})-(\tilde{c}+\tilde{d}^{\top}\mathbf{y}+\frac{1}{2}\mathbf{y}^{\top}\tilde{Q}\mathbf{y})|\to 0$
-It is a standard result that uniform convergence of polynomials on a compact set implies the convergence of their corresponding coefficients.
-Therefore, we must have uniform convergence of the coefficients with respect to $\tau\in[0,T-t_{0}]$:
-$\sup_{\tau}|c^{*}(\tau)-\tilde{c}(\tau)|\to 0,\quad\sup_{\tau}\|d^{*}(\tau)-\tilde{d}(\tau)\|\to 0,\quad\sup_{\tau}\|Q^{*}(\tau)-\tilde{Q}(\tau)\|\to 0.$
-Let $\varepsilon_{c},\varepsilon_{d},\varepsilon_{Q}$ denote these suprema, all approaching zero. Let also $\varepsilon_{\Phi}=\sup|\Phi^{*}-\tilde{\Phi}|\to 0$.
-Now, consider the spatial derivatives.
-  • w-derivatives: $V_{w}^{*}=U^{\prime}(w)\Phi^{*}$ and $\tilde{V}_{w}=U^{\prime}(w)\tilde{\Phi}$.
-    $|V_{w}^{*}-\tilde{V}_{w}|=|U^{\prime}(w)||\Phi^{*}-\tilde{\Phi}|$
-    Since $|U^{\prime}(w)|$ is bounded on $[w_{min},w_{max}]$, $\sup_{\mathcal{D}}|V_{w}^{*}-\tilde{V}_{w}|\leq(\sup|U^{\prime}|)\varepsilon_{\Phi}\to 0$.
-Similar logic applies to $V_{ww}$ using $U^{\prime\prime}(w)$.
-  • y-gradient: $\nabla_{\mathbf{y}}V^{*}=U(w)\nabla_{\mathbf{y}}\Phi^{*}$ and ~$\tilde{\nabla}_{\mathbf{y}}\tilde{V}=U(w)\nabla_{\mathbf{y}}\tilde{\Phi}$.
-We have $\nabla_{\mathbf{y}}\Phi^{*}=\Phi^{*}(d^{*}+Q^{*}\mathbf{y})$.
-    $\nabla_{\mathbf{y}}\Phi^{*}-\nabla_{\mathbf{y}}\tilde{\Phi}=(\Phi^{*}-\tilde{\Phi})(d^{*}+Q^{*}\mathbf{y})+\tilde{\Phi}[(d^{*}-\tilde{d})+(Q^{*}-\tilde{Q})\mathbf{y}]$
-    Taking norms and suprema over the compact domain $\mathcal{D}$ (where $\mathbf{y}$, $\Phi^{*}$, $\tilde{\Phi}$, $d^{*}$, $Q^{*}$ are bounded):
-    $\sup\|\nabla_{\mathbf{y}}\Phi^{*}-\nabla_{\mathbf{y}}\tilde{\Phi}\|\leq\varepsilon_{\Phi}\cdot M_{1}+M_{2}[\varepsilon_{d}+\varepsilon_{Q}M_{y}]\to 0,$
-    where $M_{1},M_{2},M_{y}$ are finite bounds. Thus, $\sup_{\mathcal{D}}\|\nabla_{\mathbf{y}}V^{*}-\nabla_{\mathbf{y}}\tilde{V}\|\leq(\sup|U|)(\sup\|\nabla\Phi^{*}-\nabla\tilde{\Phi}\|)\to 0$.
-  • yy-Hessian: $\nabla^{2}_{\mathbf{y}}V^{*}=U(w)\nabla^{2}_{\mathbf{y}}\Phi^{*}$.
-    $\nabla^{2}_{\mathbf{y}}\Phi^{*}=\nabla_{\mathbf{y}}\Phi^{*}(d^{*}+Q^{*}\mathbf{y})^{\top}+\Phi^{*}Q^{*}$
-    The difference $\nabla^{2}_{\mathbf{y}}\Phi^{*}-\nabla^{2}_{\mathbf{y}}\tilde{\Phi}$ involves terms containing $(\Phi^{*}-\tilde{\Phi})$, $(d^{*}-\tilde{d})$, and $(Q^{*}-\tilde{Q})$.
-Since these differences converge uniformly to zero and other terms are bounded, the uniform convergence $\sup_{\mathcal{D}}\|\nabla^{2}_{\mathbf{y}}V^{*}-\nabla^{2}_{\mathbf{y}}\tilde{V}\|\to 0$ follows.
-  • wy-cross-derivatives: $\nabla_{\mathbf{y}}V_{w}^{*}=U^{\prime}(w)\nabla_{\mathbf{y}}\Phi^{*}$.
-Uniform convergence follows from the convergence of $\nabla_{\mathbf{y}}\Phi^{*}$ and boundedness of $U^{\prime}(w)$.
-This completes the sketch, demonstrating that uniform value closeness implies uniform derivative closeness when the specific structure is shared.

### v3

@@
 Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality.
-Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond few state variables ($\sim$3-6 limit in prior studies).
+Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond a few state variables ($\sim$3-6 limit in prior studies).
 To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework.
-PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to guide neural network policies, handling exogenous states without dense grids.
-This PMP-guided approach holds potential for a broad class of sufficiently regular continuous-time control problems.
-Crucially, our computationally efficient “Two-Stage” variant exploits rapidly stabilizing BPTT costate estimates, converting them into near-optimal Pontryagin controls after only a short warm-up, significantly reducing training overhead.
-This enables a breakthrough in scalability: numerical experiments show PG-DPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables).
-The framework delivers near-optimal policies, offering a practical and powerful alternative for high-dimensional continuous-time portfolio choice.
+PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to directly inform neural network-based policy learning.
+A key contribution is our highly efficient Projected PG-DPO (P-PGDPO) variant.
+This approach uniquely utilizes BPTT to obtain rapidly stabilizing estimates of the Pontryagin costates and their crucial derivatives with respect to the state variables.
+These estimates are then analytically projected onto the manifold of optimal controls dictated by PMP’s first-order conditions, significantly reducing training overhead and enhancing accuracy.
+This enables a breakthrough in scalability: numerical experiments demonstrate that P-PGDPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables).
+Critically, the framework accurately captures complex intertemporal hedging demands, a feat often elusive for other methods in high-dimensional settings.
+P-PGDPO delivers near-optimal policies, offering a practical and powerful alternative for a broad class of high-dimensional continuous-time control problems.
 1 Introduction
 High-dimensional optimal control problems often fall prey to the “curse of dimensionality” (Bellman, , 1966), whereby the computational cost of classical dynamic programming (DP) approaches grows exponentially with the number of state variables.
-In a continuous-time finance setting, this difficulty becomes starkly evident in multi-asset Merton problems (Merton, , 1971), where an investor dynamically allocates wealth among numerous risky assets whose drift and volatility may depend on exogenous state variables such as macroeconomic indicators or stochastic volatility factors.
-While DP or Hamilton–Jacobi–Bellman (HJB) PDE solvers provide a rigorous framework for low-dimensional cases (Bertsekas & Tsitsiklis, , 1995), they become intractable as the number of assets or state dimensions increases, with prior influential studies often limited to analyzing problems involving up to approximately 3-5 state variables (Campbell & Viceira, , 1999; Lynch, , 2001; Balduzzi & Lynch, , 1999) or specific frameworks handling around 6 state variables (Garlappi & Skoulakis, , 2010), often necessitating simulation-based approaches for tractability (Brandt et al., , 2005).
-A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong & Zhou, , 2012) Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint).
+In a continuous-time finance setting, this difficulty becomes starkly evident in multi-asset Merton problems (Merton, , 1973), where an investor dynamically allocates wealth among numerous risky assets whose drift and volatility may depend on exogenous state variables such as macroeconomic indicators or stochastic volatility factors.
+While DP or Hamilton–Jacobi–Bellman (HJB) PDE solvers provide a rigorous framework for low-dimensional cases (Fleming and Soner, , 2006; Bertsekas and Tsitsiklis, , 1995), they become intractable as the number of assets or state dimensions increases.
+This difficulty is pervasive, impacting not only continuous-time settings (which are often addressed via discretization for numerical solution) but also analogous discrete-time models where similar state-space explosion issues arise.
+Consequently, prior influential studies in these domains are typically limited to analyzing problems involving up to approximately 3-5 state variables (Campbell and Viceira, , 1999; Lynch, , 2001; Balduzzi and Lynch, , 1999) or specific frameworks handling around 6 state variables (Garlappi and Skoulakis, , 2010), often necessitating simulation-based approaches for tractability (Brandt et al., , 2005).
+A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong and Zhou, , 2012).
+Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint).
 Notably, adding one more state dimension does not produce an exponential explosion in computational requirements, in contrast to DP (Bellman, , 1966; Campbell et al., , 2003).
@@
 Recent advances in deep learning further enhance the scalability of methods for solving high-dimensional stochastic control problems (e.g., Han et al., , 2018; Raissi et al., , 2019, for general PDE-based approaches).
-Concurrently, deep learning techniques are being applied directly to financial dynamic programming.
-Some approaches tackle infinite horizon problems via the HJB equation (Duarte et al., , 2024), while others aim to directly solve the associated nonlinear PDEs using methods like deep branching algorithms (Nguwi et al., , 2024).
-However, applications such as the Merton HJB have revealed limitations in accuracy or efficiency for certain PDE solvers like Deep BSDE variants (Nguwi et al., , 2024).
-Our PMP-based PG-DPO framework offers a distinct path, naturally handling finite horizons and focusing on the optimality conditions derived from PMP, which are applicable to a wide range of stochastic control problems possessing sufficient regularity.
-One notable strategy related to the use of adjoint methods in optimal control is to directly parameterize the policy (i.e., consumption and investment controls) with neural networks, then apply backpropagation-through-time (BPTT) to enforce Pontryagin’s optimality conditions (Huh, , 2024).
-This strategy, known as “Pontryagin-Guided Direct Policy Optimization (PG-DPO),” circumvents the need to approximate value functions or PDE solutions.
-Instead, it uses simulation-based gradients that automatically incorporate the adjoint processes, drastically reducing computational overhead in high-dimensional settings (Kushner & Yin, , 2003; Borkar & Borkar, , 2008).
+Concurrently, deep learning techniques are being applied directly to financial dynamic programming and stochastic control.
+Some approaches tackle infinite horizon problems via the HJB equation (Duarte et al., , 2024), or solve the HJB equation for utility maximisation under constraints using formulations based on second-order backward stochastic differential equations (2BSDEs) (Davey and Zheng, , 2022).
+Others aim to directly solve associated nonlinear PDEs using methods like deep branching algorithms (Nguwi et al., , 2024).
+While these BSDE-based methods and direct PDE solvers are powerful, applications such as the Merton HJB have revealed limitations in accuracy or efficiency for certain variants (Nguwi et al., , 2024).
+Moreover, some works effectively leverage the Stochastic Maximum Principle (SMP) and adjoint BSDEs, often through duality methods, to handle even non-Markovian problems in finance (Davey and Zheng, , 2022).
+Alongside these PDE-centric or BSDE-centric deep learning approaches, another line of research, more akin to reinforcement learning, focuses on directly parameterizing the control policy with neural networks and optimizing the expected objective function through simulation and gradient ascent (e.g., Zhang and Zhou, , 2019; Kopeliovich and Pokojovy, , 2024, for portfolio problems with transaction costs or under Heston dynamics).
+Our PMP-based PG-DPO framework builds upon similar ideas of direct policy optimization but is distinctively guided by the continuous-time optimality conditions of PMP.
+It utilizes BPTT-derived costate estimates for finite-horizon problems, aiming for enhanced scalability and a novel way to enforce optimality conditions for problems with sufficient regularity.
+One notable strategy related to the use of adjoint methods in optimal control, and central to our work, is to directly parameterize the policy (i.e., consumption and investment controls) with neural networks, then apply backpropagation-through-time (BPTT) to enforce Pontryagin’s optimality conditions.
+This strategy, which we term the baseline “Pontryagin-Guided Direct Policy Optimization (PG-DPO),” circumvents the need to approximate value functions or PDE solutions.
+Instead, it uses simulation-based gradients that automatically incorporate the adjoint processes, drastically reducing computational overhead in high-dimensional settings (Kushner and Yin, , 2003; Borkar and Borkar, , 2008).
 The purpose of this paper is to develop and implement this PG-DPO framework for the general multi-asset, state-dependent continuous-time portfolio problem.
@@
 This approach fundamentally relies on the existence and accurate estimation of the costate processes dictated by PMP, applicable whenever the control problem satisfies appropriate regularity conditions allowing for a well-defined Hamiltonian system.
-Furthermore, we introduce a “Two-Stage” PG-DPO variant that exploits the observation that costate estimates often stabilize much faster than the full policy network.
-Once the costates are sufficiently stable (after a short warm-up), one can bypass the potentially suboptimal neural policy outputs and directly substitute the costate estimates into the PMP first-order conditions.
-This recovers a near-optimal control analytically at every time-state point with minimal additional training, significantly reducing computational burden while maintaining high accuracy.
-This efficiency enables the solution of problems with dozens of assets and state variables (e.g., n=50, k=10 in our experiments), a scale considerably larger than those typically addressed by the numerical methods cited previously.
-This demonstrated scalability on a benchmarkable problem highlights the potential of leveraging PMP structure via BPTT estimates for tackling other complex, high-dimensional problems that may exhibit similar amenability to costate estimation under suitable regularity conditions.
+Furthermore, we introduce the “Projected PG-DPO” (P-PGDPO) variant, which significantly enhances efficiency and accuracy.
+This variant exploits the observation that BPTT-derived costate estimates often stabilize much faster than the full policy network.
+Once the costates are sufficiently stable (after a short warm-up phase), one can bypass the potentially suboptimal neural policy outputs.
+Instead, these costate estimates are directly substituted into the PMP first-order conditions (FOCs), effectively projecting the learned costate information onto the manifold of controls satisfying PMP’s necessary conditions.
+This recovers a near-optimal control analytically at every time-state point with minimal additional training, significantly reducing computational burden.
+While other deep learning methods such as Deep BSDE are also applied to high-dimensional problems, their ability to accurately capture complex, state-dependent optimal policy structures, particularly the crucial intertemporal hedging components, has been limited in challenging financial applications.
+Our numerical experiments (Section 4) demonstrate that P-PGDPO not only scales to problems with dozens of assets and state variables (e.g., n=50, k=10 in our experiments) but, critically, it is capable of precisely disentangling and reproducing these vital hedging demands—a feat not typically achieved by existing approaches in such high-dimensional settings.
+To the best of our knowledge, this explicit and accurate recovery of complex hedging terms in a high-dimensional, PMP-guided direct policy optimization framework is a novel contribution.
+This superior performance in both scalability and fidelity to the true optimal policy, especially its nuanced components, highlights the transformative potential of leveraging the PMP structure via BPTT-derived costate estimates for tackling a broad class of complex, high-dimensional control problems where amenability to costate estimation under suitable regularity conditions exists.
 The main contributions of this paper can be summarized as follows:
-  • We introduce Pontryagin-Guided Direct Policy Optimization (PG-DPO), featuring a highly efficient “Two-Stage” variant that uses PMP structure and BPTT-derived costates to solve complex continuous-time control problems.
-  • We demonstrate through numerical experiments the framework’s ability to break the dimensional barrier, successfully tackling portfolio optimization problems with dozens of assets and state factors (up to n=50, k=10).
+  • We introduce Pontryagin-Guided Direct Policy Optimization (PG-DPO), and its highly efficient “Projected PG-DPO” (P-PGDPO) variant that uses PMP structure (specifically FOCs) and BPTT-derived costates to solve complex continuous-time control problems.
+  • We demonstrate through numerical experiments that P-PGDPO not only breaks the dimensional barrier in portfolio optimization (up to n=50 assets, k=10 factors), but critically, accurately captures complex intertemporal hedging demands in these high-dimensional settings, a capability not typically demonstrated by alternative deep learning approaches.
 In the remainder of this paper, we first present the multi-asset Merton problem with exogenous states in Section 2.
-Section 3 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm and its Two-Stage extension.
+Section 3 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm, its Projected PG-DPO (P-PGDPO) extension, and provides a theoretical justification for the effectiveness of the P-PGDPO approach under specified regularity conditions.
 Numerical experiments presented in Section 4 confirm the framework’s scalability and near-optimal performance within this testbed, illustrating its potential for high-dimensional applications.
@@
 2 Multi-Asset Continuous-Time Portfolio Problem with State Variables
-In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1971; Kim & Omberg, , 1996; Liu, , 2007), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously.
+In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1973; Kim and Omberg, , 1996; Liu, , 2007), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously.
 Each risky asset features its own drift and volatility, leading to an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}$.
@@
 Although the wealth $X_{t}$ remains scalar, the investor’s optimal policy now depends on both $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and a high-dimensional state $\mathbf{Y}_{t}$.
-We handle this through Pontryagin’s maximum principle (Pontryagin, , 2014; Yong & Zhou, , 2012), introducing a costate for $X_{t}$ and another for $\mathbf{Y}_{t}$.
+We handle this through Pontryagin’s maximum principle (Pontryagin, , 2014; Yong and Zhou, , 2012), introducing a costate for $X_{t}$ and another for $\mathbf{Y}_{t}$.
 Numerically, one can represent $\{\bm{\pi}_{t},C_{t}\}$ via neural networks, thereby enabling policy gradient or dynamic programming algorithms even in multi-asset, state-dependent environments.
@@
 $d\mathbf{Y}_{t}=\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,d\mathbf{W}^{Y}_{t},$
-where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})$ is the $k\times 1$ drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})$ is often assumed to be a $k\times k$ diagonal matrix of diffusion coefficients, and $\mathbf{W}^{Y}_{t}$ is a $k\times 1$ vector of potentially correlated Brownian motions driving the state variables.
+where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})$ is the $k\times 1$ drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})$ is a $k\times k$ diagonal matrix of diffusion coefficients (i.e., $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})=\operatorname{diag}(\sigma_{Y,1}(t,\mathbf{Y}_{t}),\dots,\sigma_{Y,k}(t,\mathbf{Y}_{t}))$ where $\sigma_{Y,i}$ are the diffusion functions from the component-wise definition), and $\mathbf{W}_{t}^{Y}$ is a $k$-dimensional Brownian motion whose components may be correlated.
 We consider a financial market with $n$ risky assets and one risk-free asset.
@@
 5-6) for details on existence and uniqueness theorems for SDEs.
+The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ that maximises the expected discounted utility
+[…]
+subject to the wealth dynamics (1) and the admissibility constraints stated later in Assumption 2.
+Here $\delta>0$ is the subjective discount rate and $\kappa\geq 0$ weights the utility of terminal wealth (a bequest motive).
+Throughout we work with the Hyperbolic Absolute Risk Aversion (HARA) family of utilities,
+[…]
+whose domain is the half-line $\mathcal{D}_{U}:=\{x>x_{\min}\}$ with $x_{\min}:=-a/b$.
+HARA nests the familiar special cases CRRA $(a=0,b=1)$ and CARA $(\gamma=1,b=1)$.
+On its domain $U$ is $C^{2}$, strictly increasing and concave, and it satisfies the Inada limits $U^{\prime}(x)\to\infty$ as $x\downarrow x_{\min}$ and $U^{\prime}(x)\to 0$ as $x\to\infty$ (see Pennacchi, , 2008, §3.2).
 We restrict our attention to control processes that are economically sensible and mathematically tractable.
@@
   3. (c)
-    Boundedness and proportional consumption. There exist constants $\Pi_{\max}>0$ and $\xi\in(0,1)$ such that
+    Interior bounds. There exist constants $\Pi_{\max}>0$ and $\xi\in(0,1)$ such that
 […]
   4. (d)
-    Positive wealth. The corresponding wealth process $X_{t}$ solving (1) remains strictly positive almost surely:
+    Domain-admissible wealth. With $x_{\min}:=-a/b$ (see the HARA utility defined above) we require
 […]
@@
 (Consequently, $\Psi$ and $\Phi$ are individually positive definite and hence invertible.)
-The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ to maximize the expected discounted utility functional $J$:
-[…]
-subject to the wealth dynamics (1) and admissibility constraints (Assumption 2).
-Here $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ weights the utility of terminal wealth (bequest motive).
-The utility function $U:(0,\infty)\to\mathbb{R}$ represents the investor’s preferences.
-A widely used specification is the Constant Relative Risk Aversion (CRRA) form:
-$U(x)=\begin{cases}\displaystyle\frac{x^{1-\gamma}}{1-\gamma},&\gamma>0,\gamma\neq 1,\\[6.0pt] \ln(x),&\gamma=1,\end{cases}$
-with $\gamma>0$ denoting the coefficient of relative risk aversion (see Pennacchi, , 2008).
-Assumption 2 (iii) bounds consumption by $C_{t}\leq\xi X_{t}$ with $\xi\in(0,1)$, while Assumption 4 (b) gives a $C^{2}$, strictly concave utility $U$ satisfying the Inada limits.
-Because $\delta>0$ and $X_{t}>0$ a.s.
-(Assumption 2 (iv)), we have
-$\mathbb{E}\!\bigl{[}\,e^{-\delta t}\,U(C_{t})\bigr{]}\leq e^{-\delta t}\,U\!\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}<\infty,\quad\forall t\in[0,T],$
-and similarly $\mathbb{E}[U(X_{T})]<\infty$. Hence the functional $J(\bm{\pi},C)$ in (2) is finite for every admissible control pair, so the value function $V$ is well defined.
+Assumption 2(iii) gives the strict inequalities $0<C_{t}<\xi X_{t}$, while Assumption 2(iv) guarantees the wealth process stays inside the utility’s domain, $X_{t}>x_{\min}\;(=-a/b)$, almost surely.
+Since $U$ is $C^{2}$, strictly concave, and satisfies the Inada limits on $(x_{\min},\infty)$, and because $\delta>0$, we obtain
+$\mathbb{E}\!\left[e^{-\delta t}\,U(C_{t})\right]\;\leq\;e^{-\delta t}\,U\!\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}\;<\;\infty,\qquad\mathbb{E}\!\bigl{[}U(X_{T})\bigr{]}<\infty,$
+so the objective functional (2) is well defined for every admissible control pair $(\bm{\pi},C)$.
 The value function $V(t,x,\mathbf{y})$ associated with this problem represents the maximum expected utility achievable starting from wealth $x$ and state $\mathbf{y}$ at time $t$. Formally,
@@
 where the supremum is taken over all admissible control processes on $[t,T]$, and $\mathbb{E}_{t,x,\mathbf{y}}[\cdot]=\mathbb{E}[\cdot|X_{t}=x,\mathbf{Y}_{t}=\mathbf{y}]$ denotes the conditional expectation.
-Additional assumptions are needed for the well-posedness of the problem and regularity of the value function.
 Assumption 4 (Uniform non-degeneracy and utility).
   1. (a)
-    Uniform non-degeneracy of risky-asset volatility.
-Let $a(t,\mathbf{y}):=\bm{\sigma}(t,\mathbf{y})\,\Psi\,\bm{\sigma}(t,\mathbf{y})^{\top}$ be the instantaneous covariance matrix of the risky asset returns (Note: original used $\bm{\sigma}\bm{\sigma}^{\top}$, assuming $\Psi=I$; using $\bm{\sigma}\Psi\bm{\sigma}^{\top}$ is more general if $\mathbf{W}^{X}$ components are correlated and $\bm{\sigma}$ is diagonal).
-We assume there exists $\lambda>0$ such that
+    Uniform non-degeneracy of risky-asset volatility. Let $a(t,\mathbf{y}):=\bm{\sigma}(t,\mathbf{y})\,\Psi\,\bm{\sigma}(t,\mathbf{y})^{\top}$. There exists $\lambda>0$ such that
     $\mathbf{z}^{\top}a(t,\mathbf{y})\,\mathbf{z}\;\geq\;\lambda\,\lVert\mathbf{z}\rVert^{2},\quad\forall(t,\mathbf{y},\mathbf{z})\in[0,T]\times\mathbb{R}^{k}\times\mathbb{R}^{n}.$
-    (Equivalently, the covariance matrix of risky returns $a(t,\mathbf{y})$ is uniformly positive definite, ensuring invertibility, which is crucial for solving the portfolio problem.)
   2. (b)
     Utility specification.
-The utility function $U:(0,\infty)\to\mathbb{R}$ satisfies $U\in C^{2}(0,\infty)$, $U^{\prime}>0,\;U^{\prime\prime}<0$ (strictly increasing and concave), and the Inada conditions $\lim_{c\downarrow 0}U^{\prime}(c)=\infty,\;\lim_{c\uparrow\infty}U^{\prime}(c)=0.$
+We use the HARA utility $U$ in (3); on its domain $x>x_{\min}$ it is $C^{2}$, strictly increasing ($U^{\prime}>0$), strictly concave ($U^{\prime\prime}<0$) and satisfies the Inada limits $U^{\prime}(x)\to\infty$ as $x\downarrow x_{\min}$ and $U^{\prime}(x)\to 0$ as $x\to\infty$.
   3. (c)
-    Discount and bequest parameters. The subjective discount rate satisfies $\delta>0$ and the terminal weight satisfies $\kappa\geq 0$.
-Assumption 4 (a) guarantees the diffusion related to risky assets is non-degenerate, a key ingredient for the regularity of the associated Hamilton-Jacobi-Bellman (HJB) PDE.
-Parts (b)-(c) ensure the objective functional is well-defined and satisfies standard preference assumptions.
-Combined with the coefficient regularity in Assumption 1, these conditions typically yield the existence, uniqueness, and sufficient smoothness (e.g., $V\in C^{1,2,2}$) of the solution to the HJB equation, allowing for verification theorems to be applied; see Oksendal, (2013, Chs.
-6-7) and related literature on optimal portfolio choice.
+    Discount and bequest parameters. The subjective discount rate satisfies $\delta>0$, and the terminal-wealth weight satisfies $\kappa\geq 0$.
+Part (a) ensures the diffusion is non-degenerate, which guarantees parabolic regularity of the associated Hamilton–Jacobi–Bellman (HJB) equation.
+Part (b) specialises preferences to the HARA family; it reduces to CRRA when $a=0$ (irrespective of $b$), and to the familiar log-utility when $a=0$ and $\gamma=1$.
+Together with the coefficient regularity of Assumption 1, these conditions yield existence, uniqueness, and sufficient smoothness of the value function (e.g.
+$V\in C^{1,2,2}$); see (Oksendal, , 2013, Chs.
+6–7).
 In our numerical approach, we parameterize the control policies using neural networks (NNs).
@@
 $d\mathbf{Y}_{t}$ $=\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,d\mathbf{W}_{t}^{Y},$ (5)
-with initial conditions $X_{0}^{*}=x_{0}>0$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$.
-Recall the definitions of the volatility matrix $\bm{\sigma}(t,\mathbf{Y}_{t})$ and the covariance structure $\bm{\Omega}$ of the combined Brownian motion $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$.
+with initial conditions $X_{0}^{*}=x_{0}>x_{\min}$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$.
+Recall the definitions of the volatility matrix $\bm{\sigma}(t,\mathbf{Y}_{t})$ (for assets) and the covariance structure $\bm{\Omega}$ of the combined Brownian motion $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$.
 Complementing these forward dynamics, the optimal costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy a system of backward stochastic differential equations (BSDEs).
@@
 To properly account for the dependence of the wealth diffusion on the control $\bm{\pi}$, the Hamiltonian should incorporate the interaction between the costate dynamics (represented by $\mathbf{Z}$ terms) and the diffusion coefficients.
-Following formulations in stochastic control theory (e.g., Yong & Zhou, , 2012), the appropriate Hamiltonian $\mathcal{H}$ can be expressed as:
-$\begin{split}&\mathcal{H}\bigl{(}t,x,\mathbf{y},\bm{\pi},C,\lambda,\bm{\eta},\mathbf{Z}^{\lambda X},\mathbf{Z}^{\lambda Y},\mathbf{Z}^{\eta X},\mathbf{Z}^{\eta Y}\bigr{)}\\ &\quad=e^{-\delta t}\,U(C)+\lambda\Bigl{[}x\bigl{(}r(t,\mathbf{y})+\bm{\pi}^{\top}(\bm{\mu}(t,\mathbf{y})-r(t,\mathbf{y})\mathbf{1})\bigr{)}-C\Bigr{]}+\bm{\eta}^{\top}\,\bm{\mu}_{Y}(t,\mathbf{y})\\ &\quad\quad+\mathbf{Z}^{\lambda X}\bigl{(}x\,\bm{\sigma}(t,\mathbf{y})^{\top}\bm{\pi}\bigr{)}+\text{Tr}\bigl{(}\mathbf{Z}^{\eta Y\top}\bm{\sigma}_{Y}(t,\mathbf{y})\bigr{)}+\dots\end{split}$ (7)
-For clarity in deriving the FOC for $\bm{\pi}$, the crucial term involving $\bm{\pi}$ beyond the drift component is the one coupling $\mathbf{Z}^{\lambda X}$ with the wealth diffusion $x\,\bm{\pi}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})$ originating from $d\mathbf{W}_{t}^{X}$.
-We focus on these terms for the maximization step.
+Following formulations in stochastic control theory (e.g., Yong and Zhou, , 2012), the appropriate Hamiltonian $\mathcal{H}$ can be expressed as:
+$\mathcal{H}\!=\!e^{-\delta t}U(C)+\lambda\!\left[x\bigl{(}r+\pi^{\top}(\mu-r\mathbf{1})\bigr{)}-C\right]+\bm{\eta}^{\top}\mu_{Y}+\mathbf{Z}^{\lambda X}\!\bigl{(}x\sigma^{\top}\pi\bigr{)}+\bigl{(}\mathbf{Z}^{\eta X}\bigr{)}^{\!\top}\!\bigl{(}x\sigma^{\top}\pi\bigr{)}+\bigl{(}\mathbf{Z}^{\eta Y}\bigr{)}^{\!\top}\sigma_{Y}.$ (8)
+For the $\bm{\pi}$-FOC, only the drift term $\lambda x\bm{\pi}^{\top}(\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1})$ and the diffusion-coupling term $\mathbf{Z}^{\lambda X}(x\,\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}\bm{\pi})$ matter for the maximization step.
 To ensure the well-posedness of this FBSDE system and the applicability of PMP, additional smoothness conditions on the model coefficients are typically required.
@@
 This ensures that the drift terms in the costate BSDEs, involving derivatives like $\mathcal{H}_{\mathbf{y}}$, satisfy the required regularity conditions for the existence and uniqueness of solutions (cf.
-Yong & Zhou, , 2012).
+Yong and Zhou, , 2012).
 The core requirement of Pontryagin’s Maximum Principle is that the optimal controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ must maximize the Hamiltonian $\mathcal{H}$ (8) pointwise in time, almost surely, given the optimal state and costate processes $(X_{t}^{*},\mathbf{Y}_{t},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*})$.
@@
 $\bm{\pi}_{t}^{*}$ $=-\frac{V_{x}}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}$ (11)
-$\qquad\qquad-\frac{V_{x\mathbf{y}}}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,\rho\,\bm{\sigma}_{Y}^{\!\top}(t,\mathbf{Y}_{t}).$
+$\qquad\qquad-\frac{1}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bigl{(}\bm{\sigma}(t,\mathbf{Y}_{t})\,\bm{\rho}\,\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\bigr{)}V_{x\mathbf{y}}.$
 Here, $\Sigma(t,\mathbf{Y}_{t})=\bm{\sigma}(t,\mathbf{Y}_{t})\Psi\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ represents the instantaneous covariance matrix of the risky asset returns.
-The value function derivatives $V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\partial_{\mathbf{y}}\lambda_{t}^{*}$ are evaluated along the optimal path $(t,X_{t}^{*},\mathbf{Y}_{t})$.
+The value function derivatives $V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\nabla_{\mathbf{y}}\lambda_{t}^{*}$ (a $k\times 1$ vector) are evaluated along the optimal path $(t,X_{t}^{*},\mathbf{Y}_{t})$.
 The formula clearly separates the portfolio into two components: the myopic demand, which optimizes the immediate risk-return trade-off as if the investment opportunity set were constant, and the intertemporal hedging demand, which adjusts the portfolio to mitigate risks arising from future stochastic changes in the state variables $\mathbf{Y}_{t}$.
@@
 For illustration, in the simplified case with one risky asset ($n=1$) and one state variable ($k=1$), the optimal allocation (11) becomes:
-$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{V_{xy}}{X_{t}V_{xx}\sigma^{2}}\sigma\rho\sigma_{Y},$
+$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{1}{X_{t}V_{xx}\sigma^{2}}(\sigma\rho\sigma_{Y})V_{xy},$
 clearly showing the separation between the myopic component, proportional to the market price of risk, and the hedging component, proportional to $V_{xy}$ and the correlation $\rho$.
@@
 We first consider a model with $n$ risky assets whose risk premia are driven by a single Ornstein–Uhlenbeck (OU) factor.
-Specifically, each asset $i$ has a constant volatility $\sigma_{i}>0$ and a time-varying risk premium proportional to an OU state variable $Y_{t}$, such that the expected return is $r+\alpha_{i}\sigma_{i}Y_{t}$, where $r$ is the constant risk-free rate and $\alpha_{i}$ is a loading coefficient.
+Specifically, each asset $i$ has a constant volatility $\sigma_{i}>0$ and a time-varying risk premium proportional to an OU state variable $Y_{t}$, such that the expected return is $r+\alpha_{i}\sigma_{i}Y_{t}$, where $r$ is the constant risk-free rate.
 The state variable $Y_{t}$ follows the OU process:
@@
 $\frac{dS_{i,t}}{S_{i,t}}=(r+\alpha_{i}\sigma_{i}Y_{t})\,dt+\sigma_{i}\,dW_{i,t}^{S}.$
+Let $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$ be the vector of loading coefficients and $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ be the diagonal matrix of constant asset volatilities, where each $\sigma_{i}>0$.
 The Brownian motions driving the assets, $\mathbf{W}_{t}^{S}=(W_{1,t}^{S},\dots,W_{n,t}^{S})^{\top}$, have an $n\times n$ covariance matrix $\Psi dt$, i.e., $\mathrm{Cov}(d\mathbf{W}_{t}^{S})=\Psi\,dt$.
@@
 Let $X_{t}$ denote total wealth. The investor chooses a consumption rate $C_{t}\geq 0$ and portfolio weights $\bm{\pi}_{t}\in\mathbb{R}^{n}$ for the risky assets. The wealth dynamics follow:
-$dX_{t}=\left[X_{t}\left(\left(1-\bm{\pi}_{t}^{\top}\mathbf{1}\right)r+\bm{\pi}_{t}^{\top}(r\mathbf{1}+\bm{\alpha}\bm{\sigma}Y_{t})\right)-C_{t}\right]dt+X_{t}\bm{\pi}_{t}^{\top}\mathrm{diag}(\bm{\sigma})\,d\mathbf{W}_{t}^{S},$
-where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$, and $\bm{\sigma}=(\sigma_{1},\dots,\sigma_{n})^{\top}$.
-The instantaneous covariance matrix of the $n$-asset returns is $\Sigma=\mathrm{diag}(\bm{\sigma})\,\Psi\,\mathrm{diag}(\bm{\sigma})$.
+$dX_{t}=\bigl{[}X_{t}\bigl{(}r+\bm{\pi}_{t}^{\!\top}Y_{t}(\bm{\sigma}\bm{\alpha})\bigr{)}-C_{t}\bigr{]}dt+X_{t}\,\bm{\pi}_{t}^{\!\top}\bm{\sigma}\,d\mathbf{W}_{t}^{S}.$
+The instantaneous covariance matrix of the $n$-asset returns is $\Sigma=\bm{\sigma}\,\Psi\,\bm{\sigma}$.
 For an investor with CRRA utility $U(x)$ and discount rate $\delta>0$, the optimal consumption $C_{t}^{*}$ is determined by $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$, where $\lambda_{t}^{*}$ is the costate associated with wealth $X_{t}$.
 The PMP first-order condition yields the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$ as:
-$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}\Sigma^{-1}\left\{\lambda_{t}^{*}(\bm{\alpha}\bm{\sigma})Y_{t}+(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}(\bm{\sigma}\bm{\rho})\right\}.$
+$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}\Sigma^{-1}\left\{\lambda_{t}^{*}Y_{t}(\bm{\sigma}\bm{\alpha})+(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}(\bm{\sigma}\bm{\rho})\right\}.$
 This expression involves the costate $\lambda_{t}^{*}$ and its derivatives $\partial_{x}\lambda_{t}^{*}$ and $\partial_{Y}\lambda_{t}^{*}$.
-The first term within the braces represents the myopic demand driven by the risk premium $(\bm{\alpha}\bm{\sigma})Y_{t}$, and the second term captures the intertemporal hedging demand arising from the state variable $Y_{t}$.
+The first term within the braces represents the myopic demand driven by the risk premium $Y_{t}(\bm{\sigma}\bm{\alpha})$, and the second term captures the intertemporal hedging demand arising from the state variable $Y_{t}$.
 If $n=1$, this simplifies to $\pi_{t}^{*}=-\frac{1}{\sigma^{2}X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\{\lambda_{t}^{*}\alpha\sigma Y_{t}+\sigma(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}\rho\}$.
-For the specific case without intermediate consumption ($C_{t}=0$), an analytic solution involving Riccati ODEs can be derived using the HJB framework (Kim & Omberg, , 1996), as detailed in Appendix B.1.
+For the specific case without intermediate consumption ($C_{t}=0$), an analytic solution involving Riccati ODEs can be derived using the HJB framework (Kim and Omberg, , 1996), as detailed in Appendix B.1.
 2.3.2 A Multi-Asset, Multi-Factor Model
@@
 $d\mathbf{Y}_{t}\;=\;\kappa_{Y}\bigl{(}\bm{\theta}_{Y}-\mathbf{Y}_{t}\bigr{)}\,dt\;+\;\bm{\sigma}_{Y}\,d\mathbf{W}_{t}^{Y},$
-where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is a mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean vector, and $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ governs the factor volatilities and correlations (e.g., $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi dt$).
+where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is a mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean vector, $\bm{\sigma}_{Y}$ is the $k\times k$ diagonal matrix of factor-specific diffusion coefficients (as defined in Section 2.1), and $d\mathbf{W}_{t}^{Y}$ is a $k$-dimensional vector of Brownian motions whose components may be correlated, with $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi dt$.
 The $n$ risky assets, with prices $S_{t}^{i}$, follow:
@@
 The asset shocks $\mathbf{W}_{t}^{X}$ and factor shocks $\mathbf{W}_{t}^{Y}$ can be correlated, with $\mathrm{Cov}(dW_{t}^{X,i},dW_{t}^{Y,j})=\rho_{ij}dt$.
-Let $\rho\in\mathbb{R}^{n\times k}$ be the matrix of cross-correlations.
+Let $\bm{\rho}\in\mathbb{R}^{n\times k}$ be the matrix of cross-correlations.
 With portfolio weights $\bm{\pi}_{t}=(\pi_{t}^{1},\dots,\pi_{t}^{n})^{\top}$ and consumption $C_{t}$, wealth $X_{t}$ evolves as:
@@
 (11)), which in this context becomes:
-$\bm{\pi}_{t}^{*}=-\;\frac{1}{\,X_{t}^{*}\,\bigl{(}\partial_{x}\lambda_{t}^{*}\bigr{)}\!}\;\Sigma^{-1}\;\Bigl{\{}\,\underbrace{\lambda_{t}^{*}\;\mathrm{diag}(\bm{\sigma})\,\mathbf{A}\,\mathbf{Y}_{t}}_{\text{Myopic Term}}~{}+~{}\underbrace{\mathrm{diag}(\bm{\sigma})\,\rho\,\bm{\sigma}_{Y}\bigl{(}\partial_{\mathbf{Y}}\lambda_{t}^{*}\bigr{)}^{\top}}_{\text{Hedging Term}}\Bigr{\}},$
-where $\bm{\sigma}=(\sigma_{1},\dots,\sigma_{n})^{\top}$, $\Sigma=\mathrm{diag}(\bm{\sigma})\Psi\mathrm{diag}(\bm{\sigma})$ is the $n\times n$ asset return covariance, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}$ is the $n\times k$ matrix of factor loadings, and $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ is the $k\times 1$ gradient vector of the costate with respect to the factors $\mathbf{Y}_{t}$.
-This explicitly separates the demand into a myopic part driven by the current expected excess returns ($\mathrm{diag}(\bm{\sigma})\,\mathbf{A}\,\mathbf{Y}_{t}$) and a hedging part driven by the sensitivity to factor changes ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and their correlations ($\rho$, $\Psi$, $\Phi$).
+$\bm{\pi}_{t}^{*}=-\;\frac{1}{\,X_{t}^{*}\,\bigl{(}\partial_{x}\lambda_{t}^{*}\bigr{)}\!}\;\Sigma^{-1}\;\Bigl{\{}\,\lambda_{t}^{*}\;\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}_{t}~{}+~{}\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y}\partial_{\mathbf{Y}}\lambda_{t}^{*}\Bigr{\}},$
+where $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ is the $n\times n$ diagonal matrix of asset volatilities, $\Sigma=\bm{\sigma}\Psi\bm{\sigma}$ is the $n\times n$ asset return covariance, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}$ is the $n\times k$ matrix of factor loadings (where each $\bm{\alpha}_{i}$ is a $k\times 1$ vector of loadings for asset $i$), $\bm{\sigma}_{Y}$ is the $k\times k$ diagonal matrix of factor diffusion coefficients, and $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ is the $k\times 1$ gradient vector of the costate with respect to the factors $\mathbf{Y}_{t}$.
+This explicitly separates the demand into a myopic part driven by the current expected excess returns ($\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}_{t}$) and a hedging part driven by the sensitivity to factor changes ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and their correlations ($\bm{\rho}$, $\Psi$, $\Phi$).
 Similar to the single-factor case, an analytic solution for the terminal wealth problem ($C_{t}=0$) with CRRA utility can often be found using an exponential-quadratic value function ansatz, leading to matrix Riccati ODEs.
@@
 A key feature is that BPTT applied to these simulations yields not only unbiased estimates of the policy gradient $\nabla\widetilde{J}$ needed for optimization, but also inherently calculates estimates of the Pontryagin costate variables ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$) along each path, representing sensitivities dictated by PMP.
-This dual output enables distinct algorithmic approaches detailed in the subsequent subsections.
+This dual output enables distinct algorithmic approaches: the baseline PG-DPO method is detailed in Section 3.1, and the advanced “Two-Stage PG-DPO” variant, which leverages these costate estimates more directly, is explored in Section 3.2.
 3.1 Baseline PG-DPO for the Extended Objective
@@
   2. 2.
-    Simulate Forward Paths: For each sampled initial state $i$, simulate a discrete-time trajectory $\{(X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\}_{k=0}^{N}$ forward from $t_{k}=t_{0}^{(i)}$ to $t_{N}=T$, using the current policy networks $(\bm{\pi}_{\theta},C_{\phi})$ and correctly correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ derived from the Cholesky factor $L$ of $\mathbf{\Sigma}_{\text{block}}$.
+    Simulate Forward Paths: For each sampled initial state $i$, simulate a discrete-time trajectory $\{(X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\}_{k=0}^{N}$ forward from $t_{k}=t_{0}^{(i)}$ to $t_{N}=T$, using the current policy networks $(\bm{\pi}_{\theta},C_{\phi})$.
 The state evolution follows:
@@
     $\quad+X_{k}^{(i)}\bm{\pi}_{k}^{(i)\top}\bm{\sigma}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta\mathbf{W}_{k}^{(i),X},$
-    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$.
+    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$, $\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})$ is the vector of diffusion coefficients applied element-wise to the components of $\Delta\mathbf{W}_{k}^{(i),Y}$, and the correctly correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ are derived from the Cholesky factor $L$ of the overall covariance matrix $\bm{\Omega}$.
   3. 3.
     Calculate Realized Reward per Path: For each path $i$, compute the realized cumulative reward, explicitly noting its dependence on the starting state and policy parameters, denoted as $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)$:
-    $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)=\sum_{k=0}^{N-1}e^{-\delta t_{k}^{(i)}}U\bigl{(}C_{\phi}(t_{k}^{(i)},X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\bigr{)}\Delta t^{(i)}+\kappa e^{-\delta T}U(X_{N}^{(i)}).$
+    $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)=\sum_{k=0}^{N-1}e^{-\delta(t_{k}^{(i)}-t_{0}^{(i)})}U\bigl{(}C_{\phi}(t_{k}^{(i)},X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\bigr{)}\Delta t^{(i)}+\kappa e^{-\delta(T-t_{0}^{(i)})}U(X_{N}^{(i)}).$
     This value $J^{(i)}(t_{0}^{(i)},\dots)$ is a stochastic sample related to the conditional expected utility $J(t_{0}^{(i)},\dots)$ defined in Eq. (12).
@@
 1:Inputs:
-2:Policy nets $\pi_{\theta},\,C_{\phi}$; learning‑rate schedule $\{\alpha_{j}\}_{j=0}^{J_{\max}-1}$; max iters $J_{\max}$, batch size $M$, time steps $N$; initial‑state distribution $\eta$; utility $U$, discount $\delta$, terminal weight $\kappa$; market params $(\bm{\mu},\bm{\sigma},r,\bm{\mu}_{Y},\bm{\sigma}_{Y},\bm{\Omega})$
+2:Policy nets $\pi_{\theta},\,C_{\phi}$; learning‑rate schedule $\{\alpha_{j}\}_{j=0}^{J_{\max}-1}$; max iters $J_{\max}$, batch size $M$, time steps $N$; initial‑state distribution $\eta$; utility $U$, discount $\delta$, terminal weight $\kappa$; market params $(\bm{\mu},\bm{\sigma},r,\bm{\mu}_{Y},\bm{\sigma}_{Y}\text{ (k-dim vector of diff.
+coeffs.)},\bm{\Omega})$
 3:Output: optimised params $(\theta_{\text{opt}},\phi_{\text{opt}})$
@@
 17:         end for
-18:         $J^{(i)}=\sum_{k=0}^{N-1}e^{-\delta t_{k}}U(C_{k})\Delta t+\kappa e^{-\delta T}U(X_{N})$
+18:         $J^{(i)}=\sum_{k=0}^{N-1}e^{-\delta(t_{k}-t_{0}^{(i)})}U(C_{k})\Delta t+\kappa e^{-\delta(T-t_{0}^{(i)})}U(X_{N})$
 19:         $g^{(i)}\leftarrow\nabla_{(\theta,\phi)}J^{(i)}$
@@
 $X_{k+1}^{(i)}=X_{k}^{(i)}+\Bigl{[}X_{k}^{(i)}\Bigl{(}r(t_{k},\mathbf{Y}_{k}^{(i)})+\bm{\pi}_{k}^{(i)\top}\bigl{(}\bm{\mu}(t_{k},\mathbf{Y}_{k}^{(i)})-r(t_{k},\mathbf{Y}_{k}^{(i)})\mathbf{1}\bigr{)}\Bigr{)}-C_{k}^{(i)}\Bigr{]}\Delta t+X_{k}^{(i)}\bm{\pi}_{k}^{(i)\top}\bm{\sigma}(t_{k},\mathbf{Y}_{k}^{(i)})\,\Delta\mathbf{W}_{k}^{(i)}.$
-Applying the chain rule one step backward ($\lambda_{k}^{(i)}=\lambda_{k+1}^{(i)}\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}$) and using a first-order expansion yields the following backward recurrence relation along the path $(i)$:
+Applying the chain rule one step backward ($\lambda_{k}^{(i)}=\lambda_{k+1}^{(i)}\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}$) and using a first-order expansion yields the following backward recurrence relation along the path $(i)$ (this illustrative recurrence primarily shows influences from the wealth dynamics; a full derivation of the drift for $\lambda_{k}^{(i)}$ from the total reward $J^{(i)}$ would also account for terms related to consumption $U(C_{k}^{(i)})$ and the discount factor in its definition within the Hamiltonian context):
 $\lambda_{k}^{(i)}\approx\lambda_{k+1}^{(i)}+\lambda_{k+1}^{(i)}\bigl{[}r+\pi_{k}^{(i)}(\mu-r)\bigr{]}\Delta t+\lambda_{k+1}^{(i)}\pi_{k}^{(i)}\sigma\,\Delta W_{k}^{(i)}.$
@@
 To formally connect this to the BSDE formulation in PMP, which involves processes adapted to the filtration $\{\mathcal{F}_{t_{k}}\}$ (representing available information at time $t_{k}$), one defines the adapted costate $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}|\mathcal{F}_{t_{k}}]$ and the predictable martingale component $Z_{k}$ (related to the conditional expectation of $\lambda_{k+1}^{(i)}\Delta W_{k}^{(i)}$).
-Taking the conditional expectation ($\mathbb{E}[\cdot|\mathcal{F}_{t_{k}}]$) of the pathwise recurrence relation above, and carefully defining $Z_{k}$ using a covariation-based approach, leads precisely to the standard discrete BSDE for the adapted process $\lambda_{k}$:
+Taking the conditional expectation ($\mathbb{E}[\cdot|\mathcal{F}_{t_{k}}]$) of a more complete pathwise recurrence relation (inclusive of all terms affecting $J^{(i)}$) and carefully defining $Z_{k}$ using a covariation-based approach, leads precisely to the standard discrete BSDE for the adapted process $\lambda_{k}$:
 $\lambda_{k}=\lambda_{k+1}+\underbrace{\bigl{(}\lambda_{k+1}[r+\pi_{k}(\mu-r)]+\pi_{k}\,\sigma\,Z_{k}\bigr{)}}_{\approx\partial_{x}\mathcal{H}(t_{k},\dots)\text{ from PMP}}\Delta t-Z_{k}\,\Delta W_{k}.$ (14)
@@
 Nevertheless, the fact that BPTT provides access to these theoretically grounded estimates is fundamental.
-Firstly, they are essential inputs for the Two-Stage PG-DPO variant discussed in the next subsection (§3.2).
+Firstly, they are essential inputs for the Projected PGDPO variant discussed in the next subsection (§3.2).
 The baseline algorithm thus lays the groundwork by generating these estimates.
@@
 Consequently, the optimization process, driven by these BPTT-computed gradients, is implicitly guided by the optimality conditions derived from Pontryagin’s principle, aiming to maximize the expected objective $\widetilde{J}$.
-3.2 Two-Stage PG-DPO Algorithm
+3.2 Projected PG-DPO (P-PGDPO) Algorithm
 While the baseline PG-DPO algorithm described in Section 3.1 effectively optimizes the extended objective $\widetilde{J}$ by training the policy networks $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end, it can sometimes require a significant number of iterations for the networks to fully converge.
-The Two-Stage PG-DPO variant offers a potentially more computationally efficient alternative by exploiting an empirical observation: the costate estimates ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ and their relevant derivatives with respect to states, like $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$), which are generated during BPTT, often stabilize and become reasonably accurate much faster than the full policy networks achieve optimality.
-Algorithm 2 Two‑Stage PG‑DPO: Warm‑up Training & Deployment
+The Projected PG-DPO (P-PGDPO) variant offers a potentially more computationally efficient alternative by exploiting an empirical observation: the costate estimates ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ and their relevant derivatives with respect to states, like $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$), which are generated during BPTT, often stabilize and become reasonably accurate much faster than the full policy networks achieve optimality.
+This P-PGDPO approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly by projecting them onto the manifold of policies satisfying the PMP first-order conditions.
+Algorithm 2 Projected PG‑DPO (P-PGDPO): Warm‑up Training & Deployment
 1:Stage 1: Warm‑up Training
@@
 10:for $j=1$ to $M_{\mathrm{MC}}$ do
-11:     Simulate path with $(\pi_{\theta^{*}},C_{\phi^{*}})$ using $N^{\prime}$ steps
+11:     Simulate path from $(t,X,\bm{Y})$ with $(\pi_{\theta^{*}},C_{\phi^{*}})$ using $N^{\prime}$ steps. Let $t_{0}^{\prime}=t$.
 12:     Compute realized reward for path $j$:
-13:     $\qquad\displaystyle J^{(j)}\leftarrow\sum_{n=0}^{N^{\prime}-1}e^{-\delta t_{n}}U\bigl{(}C_{t_{n}}^{(j)}\bigr{)}\,\Delta t^{\prime}+\kappa e^{-\delta T}U\bigl{(}X_{T}^{(j)}\bigr{)}$
-14:     Evaluate $\lambda_{t}^{(j)}$, $\partial_{x}\lambda_{t}^{(j)}$, $\partial_{\bm{Y}}\lambda_{t}^{(j)}$ and append
+13:     $\qquad\displaystyle J^{(j)}\leftarrow\sum_{n=0}^{N^{\prime}-1}e^{-\delta(t_{n}-t_{0}^{\prime})}U\bigl{(}C_{t_{n}}^{(j)}\bigr{)}\,\Delta t^{\prime}+\kappa e^{-\delta(T-t_{0}^{\prime})}U\bigl{(}X_{T}^{(j)}\bigr{)}$
+14:     Evaluate $\lambda_{t}^{(j)}$, $\partial_{x}\lambda_{t}^{(j)}$, $\partial_{\bm{Y}}\lambda_{t}^{(j)}$ (costates at $t_{0}^{\prime}=t$) via BPTT on $J^{(j)}$ and append to lists $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}$.
 15:end for
-16:Compute $\hat{\lambda}_{t},\widehat{\partial_{x}\lambda}_{t},\widehat{\partial_{\bm{Y}}\lambda}_{t}$triangleright E.g., by averaging estimates from loop
+16:Compute $\hat{\lambda}_{t},\widehat{\partial_{x}\lambda}_{t},\widehat{\partial_{\bm{Y}}\lambda}_{t}$triangleright E.g., by averaging estimates from lists $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}$
 17:$C^{\mathrm{PMP}}\leftarrow(U^{\prime})^{-1}\!\bigl{(}e^{\delta t}\hat{\lambda}_{t}\bigr{)}$triangleright Using PMP FOC Eq. (8)
@@
 19:return $(\bm{\pi}^{\mathrm{PMP}},C^{\mathrm{PMP}})$
-This Two-Stage approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly from the PMP first-order conditions, bypassing the need for fully converged policy networks for deployment.
+This P-PGDPO approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly from the PMP first-order conditions, bypassing the need for fully converged policy networks for deployment.
 The procedure consists of two distinct stages, described below using a single level of enumeration:
@@
     After completing the $K_{0}$ warm-up iterations, the trained policy networks $(\bm{\pi}_{\theta^{*}},C_{\phi^{*}})$ are typically not directly used for generating controls during deployment.
-Instead, to determine the control actions for any given state $(t_{k},X_{k},\mathbf{Y}_{k})$ encountered during deployment, we employ a Monte Carlo-like averaging procedure to obtain reliable costate estimates needed for the PMP formulas:
-    First, to obtain the stabilized costate estimates corresponding to the state $(t_{k},X_{k},\mathbf{Y}_{k})$, we perform the following averaging: Simulate multiple ($M_{\text{MC}}$) forward paths starting from $(t_{k},X_{k},\mathbf{Y}_{k})$ (or states very close to it) up to the horizon $T$, using the fixed policy parameters $(\theta^{*},\phi^{*})$ obtained from Stage 1.
-For each simulated path $j=1,\dots,M_{\text{MC}}$, calculate its realized reward $J^{(j)}(t_{k},X_{k},\mathbf{Y}_{k};\theta^{*},\phi^{*})$ and apply BPTT to obtain the costate estimates $\lambda_{k}^{(j)},\partial_{x}\lambda_{k}^{(j)},\partial_{\mathbf{Y}}\lambda_{k}^{(j)}$ specifically at the starting node $k$ for that path $j$.
+Instead, to determine the control actions for any given state $(t_{k},X_{k},\mathbf{Y}_{k})$ encountered during deployment (let this deployment time be $t_{k}=t_{deploy}$), we employ a Monte Carlo-like averaging procedure to obtain reliable costate estimates needed for the PMP formulas:
+    First, to obtain the stabilized costate estimates corresponding to the state $(t_{deploy},X_{k},\mathbf{Y}_{k})$, we perform the following averaging: Simulate multiple ($M_{\text{MC}}$) forward paths starting from $(t_{deploy},X_{k},\mathbf{Y}_{k})$ (or states very close to it) up to the horizon $T$, using the fixed policy parameters $(\theta^{*},\phi^{*})$ obtained from Stage 1.
+For each simulated path $j=1,\dots,M_{\text{MC}}$, calculate its realized reward $J^{(j)}(t_{deploy},X_{k},\mathbf{Y}_{k};\theta^{*},\phi^{*})$ (using discounting relative to $t_{deploy}$ as in Algorithm 2) and apply BPTT to obtain the costate estimates $\lambda_{deploy}^{(j)},\partial_{x}\lambda_{deploy}^{(j)},\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$ specifically at the starting node $t_{deploy}$ for that path $j$.
 The final “stabilized” estimates are then computed by averaging these values across the $M_{\text{MC}}$ paths:
-    $\lambda_{k}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\lambda_{k}^{(j)},\quad\partial_{x}\lambda_{k}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{x}\lambda_{k}^{(j)},\quad\partial_{\mathbf{Y}}\lambda_{k}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{\mathbf{Y}}\lambda_{k}^{(j)}$
+    $\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\lambda_{deploy}^{(j)},\quad\partial_{x}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{x}\lambda_{deploy}^{(j)},\quad\partial_{\mathbf{Y}}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$
     This averaging acts as a Monte Carlo estimation step, significantly reducing the variance inherent in the single-path BPTT estimates derived during Stage 1 and providing more robust inputs for the PMP formulas.
-    Next, we plug these averaged costate estimates ($\lambda_{k},\partial_{x}\lambda_{k},\partial_{\mathbf{Y}}\lambda_{k}$) directly into the analytical formulas derived from the Pontryagin Maximum Principle’s first-order conditions (Section 2.2).
-The optimal consumption is calculated as:
-    $C_{k}^{\mathrm{PMP}}=\bigl{(}U^{\prime}\bigr{)}^{-1}\!\bigl{(}e^{\delta\,t_{k}}\,\lambda_{k}\bigr{)}.$
-    The optimal investment portfolio is calculated as:
-    $\bm{\pi}_{k}^{\mathrm{PMP}}$ $=-\frac{1}{X_{k}\,\partial_{x}\lambda_{k}}\,\Sigma^{-1}\!\bigl{(}t_{k},\mathbf{Y}_{k}\bigr{)}\Bigl{\{}\lambda_{k}\bigl{[}\bm{\mu}(t_{k},\mathbf{Y}_{k})-r(t_{k},\mathbf{Y}_{k})\mathbf{1}\bigr{]}$ (14)
-    $\qquad+\bm{\sigma}(t_{k},\mathbf{Y}_{k})^{\top}(\partial_{\mathbf{Y}}\lambda_{k})\,\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\rho^{\top}\Psi^{-1}\Bigr{\}}.$ (15)
-    Finally, these analytically computed values $(C_{k}^{\mathrm{PMP}},\bm{\pi}_{k}^{\mathrm{PMP}})$, based on the averaged costate estimates, are used as the control actions for the state $(t_{k},X_{k},\mathbf{Y}_{k})$ during deployment.
-In summary, the Two-Stage PG-DPO approach performs a short warm-up training (Stage 1) using the baseline PG-DPO process to obtain a policy $(\theta^{*},\phi^{*})$ whose BPTT generates reasonably stable costate estimates.
+    Next, we plug these averaged costate estimates ($\lambda_{deploy},\partial_{x}\lambda_{deploy},\partial_{\mathbf{Y}}\lambda_{deploy}$) directly into the analytical formulas derived from the Pontryagin Maximum Principle’s first-order conditions (Section 2.2).
+The optimal consumption at $t_{deploy}$ is calculated as:
+    $C_{deploy}^{\mathrm{PMP}}=\bigl{(}U^{\prime}\bigr{)}^{-1}\!\bigl{(}e^{\delta\,t_{deploy}}\,\lambda_{deploy}\bigr{)}.$
+    The optimal investment portfolio at $t_{deploy}$ is calculated as:
+    $\bm{\pi}_{deploy}^{\mathrm{PMP}}$ $=-\frac{1}{X_{k}\,\partial_{x}\lambda_{deploy}}\,\Sigma^{-1}\!\bigl{(}t_{deploy},\mathbf{Y}_{k}\bigr{)}\Biggl{\{}\lambda_{deploy}\bigl{[}\bm{\mu}(t_{deploy},\mathbf{Y}_{k})-r(t_{deploy},\mathbf{Y}_{k})\mathbf{1}\bigr{]}$
+    $\qquad+\bigl{(}\bm{\sigma}(t_{deploy},\mathbf{Y}_{k})\,\bm{\rho}\,\bm{\sigma}_{Y}(t_{deploy},\mathbf{Y}_{k})\bigr{)}(\partial_{\mathbf{Y}}\lambda_{deploy})\Biggr{\}}.$ (15)
+    (This formula aligns with Eq.
+(11), where $\lambda_{deploy}$ corresponds to $V_{x}$, $\partial_{x}\lambda_{deploy}$ to $V_{xx}$, and $\partial_{\mathbf{Y}}\lambda_{deploy}$ to $V_{x\mathbf{Y}}$ at state $(t_{deploy},X_{k},\mathbf{Y}_{k})$.)
+    Finally, these analytically computed values $(C_{deploy}^{\mathrm{PMP}},\bm{\pi}_{deploy}^{\mathrm{PMP}})$, based on the averaged costate estimates, are used as the control actions for the state $(t_{deploy},X_{k},\mathbf{Y}_{k})$ during deployment.
+In summary, the Projected PG-DPO (P-PGDPO) approach performs a short warm-up training (Stage 1) using the baseline PG-DPO process to obtain a policy $(\theta^{*},\phi^{*})$ whose BPTT generates reasonably stable costate estimates.
 It then leverages these estimates in Stage 2, refining them through Monte Carlo averaging at the point of deployment and plugging them into the known analytical structure of the optimal control derived from PMP.
@@
 The Monte Carlo averaging step in Stage 2 is crucial for obtaining reliable inputs for the PMP formulas from the potentially noisy single-path estimates.
-3.3 Rationale for the Two-Stage Approach
-3.3.1 The Parabolic Nature of Financial Models and PDE Regularity
+3.3 Rationale for the Projected PG-DPO (P-PGDPO) Approach
+In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are modeled by stochastic differential equations (SDEs), the Value Function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically solves a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming and Soner, , 2006; Yong and Zhou, , 2012).
+This HJB equation exhibits a parabolic form.
+The specific nature of its parabolicity is primarily determined by the diffusion coefficients of the underlying SDEs, namely $\bm{\sigma}$ (related to assets) and $\bm{\sigma}_{Y}$ (related to state variables).
+While the first-order drift terms (involving $\bm{\mu},\bm{\mu}_{Y}$) are crucial for the solution’s properties, they do not affect the parabolicity classification, which is mainly governed by the second-order diffusion terms.
+The behavior of the diffusion matrix $a(t,x,\mathbf{y})$ (the matrix multiplying second-order derivatives in the HJB equation) allows for classifying parabolic PDEs into three main classes (see, e.g., Evans, , 2022).
+This classification significantly impacts the regularity (smoothness) of the Value Function $V$ and its associated Costates (the first-order derivatives $\nabla V$) (cf.
+Evans, , 2022, Ch.
+6-7).
+Costates represent the sensitivity of the Value Function, and Pontryagin’s Maximum Principle links optimal Costates to the optimal policy.
+The core idea of Projected PG-DPO (P-PGDPO) involves estimating Costates and their derivatives via BPTT and using these estimates in the PMP formulas.
+Therefore, the regularity of the Costates—tied to the model’s parabolicity class—directly affects the difficulty of Costate estimation and the effectiveness of the P-PGDPO method.
 Table 1: Classification of Parabolic HJB PDEs in Financial Models
 Class Diffusion Matrix $a(t,x,\mathbf{y})$ Property SDE Coefficient Requirements and Dynamic Portfolio Examples
-(U) Uniformly Elliptic Uniformly positive definite across the state space The diffusion coefficients $\bm{\sigma}$ (for wealth $X_{t}$) and $\bm{\sigma}_{Y}$ (for state $\mathbf{Y}_{t}$) are such that the resulting diffusion matrix in the HJB is non-zero and bounded below by a positive constant everywhere ($\bm{\sigma}\Psi\bm{\sigma}^{\top}$ and $\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}$ uniformly positive definite).
+(U) Uniformly Elliptic Uniformly positive definite across the state space The diffusion coefficients $\bm{\sigma}$ (for wealth $X_{t}$) and $\bm{\sigma}_{Y}$ (for state $\mathbf{Y}_{t}$) are such that the resulting diffusion matrix in the HJB is non-zero and bounded below by a positive constant everywhere (e.g., $\bm{\sigma}\Psi\bm{\sigma}^{\top}$ and $\operatorname{diag}(\bm{\sigma}_{Y})\Phi\operatorname{diag}(\bm{\sigma}_{Y})^{\top}$ uniformly positive definite).
 Drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) are also sufficiently regular (e.g., Lipschitz).
 ($\implies$ Value Function / Costate very smooth) Examples: Merton’s portfolio problem with constant coefficients ($\mu,\sigma>0,r$ are constants); models with state-dependent volatility that remains strictly positive; the multi-asset OU model in Section 2.3.1 (if asset volatilities $\sigma_{i}>0$, factor volatility $\sigma_{Y}>0$, and $\Psi$ is positive definite).
-(H) Hypo-elliptic / Uniform on Compacts May degenerate at some points, but regularity is achieved via system structure (e.g., drift-diffusion coupling via Hörmander’s condition) or holds within a compact domain of interest.
-Degeneracy occurs away from boundaries.
+(H) Hypo-elliptic May degenerate (even globally).
+Regularity is achieved if the system structure satisfies conditions like Hörmander’s (i.e., Lie brackets of vector fields span the tangent space).
+The problem might be uniformly elliptic on compact subsets of interest.
 Diffusion coefficients ($\bm{\sigma}$ or $\bm{\sigma}_{Y}$) can become zero at certain points within the domain.
@@
 ($\implies$ Value Function / Costate regularity may break down near the boundary) Examples: CIR (Cox-Ingersoll-Ross) interest rate/volatility model (diffusion term like $\sigma_{Y}\sqrt{Y_{t}}$); Heston stochastic volatility model (asset diffusion term like $\sigma\sqrt{\nu_{t}}$).
-In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are modeled by stochastic differential equations (SDEs), the Value Function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically solves a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming & Soner, , 2006; Yong & Zhou, , 2012).
-This HJB equation exhibits a parabolic form.
-The specific nature of its parabolicity is primarily determined by the diffusion coefficients of the underlying SDEs, namely $\bm{\sigma}$ (related to assets) and $\bm{\sigma}_{Y}$ (related to state variables).
-While the drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) influence the first-order terms in the HJB and are crucial for the solution’s properties, they do not dictate the type’ of parabolicity, which is mainly governed by the second-order terms (diffusion).
-The behavior of the diffusion matrix $a(t,x,\mathbf{y})$ (the matrix multiplying second-order derivatives in the HJB equation) allows for classifying parabolic PDEs into three main classes (see, e.g., Evans, , 2022).
-This classification significantly impacts the regularity (smoothness) of the Value Function $V$ and its associated Costates (the first-order derivatives $\nabla V$) (cf.
-Evans, , 2022, Ch.
-6-7).
-Costates represent the sensitivity of the Value Function, and Pontryagin’s Maximum Principle links optimal Costates to the optimal policy.
-The core idea of Two-Stage PG-DPO involves estimating Costates and their derivatives via BPTT and using these estimates in the PMP formulas.
-Therefore, the regularity of the Costates—tied to the model’s parabolicity class—directly affects the difficulty of Costate estimation and the effectiveness of the Two-Stage method.
-The table 1 summarizes these classes, their properties, links to SDE coefficients, and relevant examples within the dynamic portfolio optimization context.
-Models falling into classes (U) and (H) are expected to possess sufficiently regular Costates, providing a favorable setting for the estimation required by the Two-Stage PG-DPO method.
-The numerical experiments in this paper are conducted using models belonging to classes (U) or (H), where the Two-Stage PG-DPO effectively leverages the expected regularity of the Costates.
+Models falling into classes (U) and (H) are expected to possess sufficiently regular Costates, providing a favorable setting for the estimation required by the Projected PG-DPO (P-PGDPO) method.
+The numerical experiments in this paper are conducted using models belonging to classes (U) or (H), where P-PGDPO effectively leverages the expected regularity of the Costates.
 The method is particularly effective as it generally avoids visiting degenerate boundaries during practical implementation.
-3.3.2 Mathematical Foundation: Derivative Stability via PDE Comparison
-A key theoretical insight, particularly demonstrable in well-structured (e.g., affine) models, comes from analyzing the stability of value function derivatives under policy perturbations using regularity theory for parabolic partial differential equations (e.g., Evans, , 2022; Krylov, , 2024).
-This perspective illustrates how closeness of policies (in a specific sense) can imply closeness of derivatives, providing a foundation for why the Two-Stage approach might be effective.
-First, let us recall the PDE associated with a given policy.
-For an admissible and sufficiently smooth policy $(\pi(t,z),c(t,z))$, where $z=(x,y)$ represents the state, the corresponding value function $V(t,z)$ (representing the expected utility achieved under this *specific* policy from state $z$ at time $t$) satisfies the linear parabolic PDE:
-$-\frac{\partial V}{\partial t}=\mathcal{L}_{\pi,c}[V](t,z)+f_{\pi,c}(t,z),\quad\text{subject to }V(T,z)=g(z).$ (16)
-Here, $g(z)=\kappa e^{-\delta T}U(x)$ is the terminal utility, $f_{\pi,c}(t,z)=e^{-\delta t}U(c(t,z))$ is the instantaneous utility source term under policy $c$, and $\mathcal{L}_{\pi,c}$ is the second-order linear differential generator associated with the state dynamics under policy $(\pi,c)$:
-$\mathcal{L}_{\pi,c}[V]=$ $V_{x}\left[x\left((1-\pi^{\top}\mathbf{1})r(t,y)+\pi^{\top}\mu(t,y)\right)-c(t,z)\right]+V_{\mathbf{y}}^{\top}\mu_{Y}(t,y)$
-$+\frac{1}{2}V_{xx}x^{2}\pi^{\top}\sigma(t,y)\Psi\sigma(t,y)^{\top}\pi$
-$+V_{x\mathbf{y}}^{\top}x\pi^{\top}\sigma(t,y)\rho\sigma_{Y}(t,y)^{\top}$ (17)
-$+\frac{1}{2}\text{Tr}\left(V_{\mathbf{y}\mathbf{y}}\sigma_{Y}(t,y)\Phi\sigma_{Y}(t,y)^{\top}\right)-\delta V$
-(We incorporate the time discount term $-\delta V$ into the operator $\mathcal{L}_{\pi,c}$ for convenience).
-Let $V^{*}(t,z)$ denote the true optimal value function of the control problem defined in Section 2.1, achieved by the optimal policy $(\pi^{*}(t,z),c^{*}(t,z))$.
-Assuming sufficient regularity, $V^{*}$ satisfies the non-linear Hamilton-Jacobi–Bellman (HJB) equation (Fleming & Soner, , 2006), which implies it also satisfies the linear PDE (16) when the optimal policy $(\pi^{*},c^{*})$ is used:
-$-\frac{\partial V^{*}}{\partial t}=\mathcal{L}_{\pi^{*},c^{*}}[V^{*}](t,z)+f_{\pi^{*},c^{*}}(t,z),\quad\text{subject to }V^{*}(T,z)=g(z).$ (18)
-To formalize the notion of policy closeness relevant to the PDE coefficients, we introduce the following definition.
-Let $\Sigma_{\text{asset}}(t,y)=\sigma(t,y)\Psi\sigma(t,y)^{\top}$ and $\Sigma_{\text{asset-state}}(t,y)=\sigma(t,y)\rho\sigma_{Y}(t,y)^{\top}$.
-Definition 1 (Coefficient-relevant $\epsilon$-closeness).
-Let the spatial domain be
-$\Omega\subseteq(0,\infty)\times\mathbb{R}^{k},\qquad z=(x,\mathbf{y})\in\Omega,$
-i.e.
-$x$ is wealth and $\mathbf{y}\in\mathbb{R}^{k}$ is the $k$-dimensional state vector.
-(In practice $\Omega$ can be either the full space or a bounded truncated box that contains the region the process visits with high probability.)
-Fix exponents $p>d+2$ and $q\geq 2$, where $d=\dim\Omega=k+1$. For any measurable function $h(t,z)$ set
-[…]
-Let $(\pi^{*},c^{*})$ be the optimal policy and $(\pi,c)$ any admissible policy. We say $(\pi,c)$ is $(\epsilon,p,q)$-close to $(\pi^{*},c^{*})$ on $Q_{T}=(0,T)\times\Omega$ if
-$\|c-c^{*}\|_{L^{q,p}}$ $\leq\epsilon,$
-$\|\pi^{\top}\mathbf{1}-\pi^{*\top}\mathbf{1}\|_{L^{q,p}}$ $\leq\epsilon,$
-$\|\pi^{\top}\bm{\mu}-\pi^{*\top}\bm{\mu}\|_{L^{q,p}}$ $\leq\epsilon,$
-$\|\pi^{\top}\Sigma_{\text{asset}}\pi-\pi^{*\top}\Sigma_{\text{asset}}\pi^{*}\|_{L^{q,p}}$ $\leq\epsilon,$
-$\|\pi^{\top}\Sigma_{\text{asset-state}}-\pi^{*\top}\Sigma_{\text{asset-state}}\|_{L^{q,p}}$ $\leq\epsilon.$
-Proposition 3.1 (Derivative closeness under coefficient $\epsilon$-closeness).
-Let $V^{*}$ be the optimal value function corresponding to the optimal policy $(\pi^{*},c^{*})$ (cf.
-(18)) and let $V$ be the value function generated by another admissible policy $(\pi,c)$ (cf.
-(16)).
-Set
-$D\;:=\;V^{*}-V,\qquad Q_{T}:=(0,T)\times\Omega.$
-Assume
-  1. (A1)
-    Uniform parabolicity. The operator $\mathcal{L}_{\pi^{*},c^{*}}$ is uniformly parabolic on $Q_{T}$.
-  2. (A2)
-    Baseline regularity. There exist exponents $p>d+2$ and $q\geq 2$ (with $d=\dim\Omega=k+1$) such that the value functions lie in the Sobolev space (e.g., Adams & Fournier, , 2003; Evans, , 2022)
-    $V^{*},\,V\;\in\;L^{q}\bigl{(}0,T;W^{2,p}(\Omega)\bigr{)}.$
-  3. (A3)
-    Coefficient $\epsilon$-closeness. The candidate policy $(\pi,c)$ is $(\epsilon,p,q)$-close to $(\pi^{*},c^{*})$ in the sense of Definition 1.
-Then $D$ satisfies the linear PDE
-$-\partial_{t}D\;=\;\mathcal{L}_{\pi^{*},c^{*}}[D]\;+\;S^{\prime},\quad\text{where}\quad S^{\prime}=(\mathcal{L}_{\pi^{*},c^{*}}-\mathcal{L}_{\pi,c})[V]+\bigl{(}f_{\pi^{*},c^{*}}-f_{\pi,c}\bigr{)}.$
-Moreover, the source term is small in the weighted norm:
-$\|S^{\prime}\|_{L^{q,p}}\;\leq\;C_{0}\,\epsilon,$
-and by standard $L^{p}$ / Schauder estimates for uniformly parabolic equations (see, e.g., Evans, , 2022, Ch. 7),
-$\|D\|_{L^{q}\!\bigl{(}0,T;W^{2,p}(\Omega)\bigr{)}}\;\leq\;C\,\|S^{\prime}\|_{L^{q,p}}\;\leq\;C\,\epsilon,$
-for a constant C independent of $\epsilon$.
-Consequently, the value-function gap and all first- and second-order spatial derivatives satisfy
-$\bigl{\|}V^{*}-V\bigr{\|}_{L^{q,p}}+\bigl{\|}\nabla V^{*}-\nabla V\bigr{\|}_{L^{q,p}}+\bigl{\|}D^{2}V^{*}-D^{2}V\bigr{\|}_{L^{q,p}}\;\leq\;C\,\epsilon.$
-Hence the costates $(\nabla V)$ produced under $(\pi,c)$ are $O(\epsilon)$-close to the optimal costates on the entire space–time domain.
-Remark on assumptions, PG–DPO rationale, and scope of the result.
-Proposition 3.1 makes precise the intuition for the uniformly–parabolic class (“U class”’): if two policies yield close PDE-coefficients in the weak sense of Definition 1 (only a global [^1] A still weaker occupation-weighted norm is discussed below.
-$L^{q,p}$-distance), then their value functions—and hence their gradients (costates)—differ by at most $O(\epsilon)$.
-Assumptions (A1)–(A2) are the classical entry ticket for $L^{p}$ / Schauder regularity (Evans, , 2022) and are easily verified for affine models or for models that become uniformly elliptic after adding a small floor to each diffusion coefficient (e.g.
-replace $\sigma(t,z)$ by $\sqrt{\sigma(t,z)^{2}+\varepsilon^{2}}$ with $\varepsilon\sim 10^{-4}$).
-Such a floor leaves the economics virtually unchanged while preventing the diffusion matrix from collapsing to zero, which is exactly what uniform parabolicity requires.
-The genuinely new ingredient is (A3), the $\epsilon$-coefficient closeness that Stage 1 of PG-DPO is designed to enforce.
-Stage 1 produces a sub-optimal policy $(\pi,c)$ whose coefficient error is $\epsilon$.
-Via Proposition 3.1 this translates into an $O(\epsilon)$ error for the costate $\lambda\approx\nabla V$, so the BPTT outputs obtained during the short warm-up are already $O(\epsilon)$-accurate surrogates of the true optimal costates.
-Stage 2 simply substitutes these estimates into the PMP first-order conditions (Pontryagin, , 2014; Yong & Zhou, , 2012), thereby constructing a near-optimal control without having to train the policy network to full convergence.
-Global vs.– local norms.– The result is stated with a global $\|\,\cdot\,\|_{L^{q,p}}$ norm for transparency. If desired, one may replace it by the occupation-weighted norm
-[…]
-where $\mu_{(\pi,c)}$ is the state-density under the candidate policy.
-Because the proof relies on local PDE estimates, the same $O(\epsilon)$ bound holds—now only on the subset of the state space that the process actually visits with non-negligible probability.
-This is valuable for models whose coefficients degenerate in remote regions.
-Beyond the U class.
-– For hypo-elliptic (H class) (cf.
-Hörmander, , 1967) or boundary-degenerate (D class) models (e.g., Oleinik, , 2012), the full global $L^{p}$ theory is subtler, yet the algorithmic logic of Two-Stage PG-DPO is unaffected:
-  • the PMP equations and BPTT apply to any differentiable dynamics;
-  • costates typically stabilise faster than the full optimal control;
-  • a brief warm-up that makes the coefficient error small where it matters is enough to obtain high-quality costate estimates.
-Even when the rigorous guarantee of Proposition 3.1 is unavailable, this costate-first, PMP-second recipe remains an effective practical strategy—an observation confirmed empirically in the numerical section (Section 4) and worth exploring further for genuinely non-affine or degenerate models.
-3.3.3 Qualitative Rationale and Broader Applicability
-While the formal analysis in Proposition 3.1 provides rigorous support under specific regularity conditions (like uniform parabolicity, often associated with affine models), the core ideas underpinning the Two-Stage PG-DPO approach possess a broader qualitative appeal that suggests applicability beyond these settings.
-Crucially, neither the Pontryagin Maximum Principle (PMP) itself (Pontryagin, , 2014; Yong & Zhou, , 2012) nor the mechanics of backpropagation-through-time (BPTT) (e.g., Goodfellow, , 2016) are inherently restricted to affine or uniformly parabolic systems.
-PMP provides necessary optimality conditions for general nonlinear dynamics, provided the required functions are differentiable.
-Similarly, BPTT serves as a general algorithm for computing gradient estimates through complex computational graphs derived from simulations, regardless of the underlying model’s linearity.
-Furthermore, there is a compelling economic and mathematical intuition that the PMP costates ($\lambda\approx\nabla V$, representing marginal values) might exhibit properties that make them more amenable to estimation than the potentially complex optimal policy $\pi^{*}$ itself, especially in high-dimensional problems.
-For instance, the costate functions might be representable by smoother or lower-complexity functions compared to $\pi^{*}$.
-Alternatively, the process of learning or estimating these costate-related functions via BPTT with neural networks might converge more rapidly or stably than directly learning the intricate mapping defined by the optimal policy.
-The Two-Stage PG-DPO strategy is designed precisely to leverage this potential advantage.
-It effectively decouples the optimization challenge into two parts: (i) tackling the potentially more tractable task of estimating costates (and their relevant derivatives) using BPTT during a warm-up phase (ii) employing the robust analytical structure of the PMP first-order conditions to translate these estimates into effective control actions during deployment.
-Therefore, even in scenarios where the rigorous derivative closeness guarantees of Proposition 3.1 are harder to establish formally (e.g., for non-affine models, or systems falling into the hypo-elliptic or degenerate parabolic classes), this underlying qualitative rationale remains persuasive.
-The potential for BPTT-derived costate estimates to stabilize relatively quickly and capture essential optimality information offers a strong motivation for exploring the Two-Stage PG-DPO approach in these more complex settings.
-This rationale, significantly reinforced by the numerical success demonstrated in the affine benchmark case (Section 4), positions PG-DPO (particularly the superior Two-Stage variant) as a promising and potentially powerful tool for tackling challenging high-dimensional, possibly non-affine, continuous-time control problems prevalent in finance and other domains.
-Naturally, its practical effectiveness in such extended settings must be carefully evaluated empirically on a case-by-case basis, highlighting an important direction for future research.
+Building upon this expected regularity, a more formal mathematical justification for the P-PGDPO approach can be provided.
+The objective is to establish that if the policy $(\pi,c)$ obtained from the Stage 1 warm-up phase satisfies a sufficiently small Pontryagin first-order condition (FOC) residual $\varepsilon$, then the feedback policy $\widehat{\pi}$ constructed in Stage 2 using BPTT estimates $\widehat{\lambda}$ (with BPTT estimation error $\delta_{\mathrm{BPTT}}$) will be close to the true optimal policy $\pi^{*}$.
+This analysis relies on certain standing assumptions concerning parabolic regularity and the FOC gap, which are formally stated as Assumption 6 and 7 in Appendix C.
+We use $L^{q,p}(Q_{T})$ to denote the standard mixed space-time Lebesgue norm over the domain $Q_{T}=(0,T)\times\Omega\subseteq\mathbb{R}^{d+1}$, as detailed in Appendix C.
+The main result of this analysis is encapsulated in the following theorem, the proof of which is detailed in Appendix C along with supporting lemmas and propositions.
+Theorem 1 (Policy Gap Bound).
+Under Assumptions 6 and 7 (detailed in Appendix C), let $(\pi,c)$ be the policy from Stage 1 satisfying the FOC-gap condition $\|\mathcal{R}_{\mathrm{FOC}}(\pi,c;\lambda)\|_{L^{q,p}}\leq\varepsilon$, where $\lambda=\nabla V^{\pi,c}$.
+Let $\pi^{*}$ be the true optimal policy, and let $\widehat{\pi}=F(\widehat{\lambda})$ be the policy deployed in Stage 2 using BPTT estimates $\widehat{\lambda}$, where $F$ is the explicit mapping from costates $(\lambda,\nabla\lambda)$ to controls derived from the PMP FOCs.
+Assume $F$ is locally Lipschitz continuous with constant $C_{F}$.
+If the BPTT estimation error $\|\widehat{\lambda}-\nabla V^{\pi,c}\|_{L^{q,p}}\leq\delta_{\mathrm{BPTT}}$ (where $\delta_{\mathrm{BPTT}}$ quantifies the precision of BPTT-derived costate estimates relative to Stage 1 costates), then the policy gap is bounded by:
+$\boxed{\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\mathrm{tot}}(\delta_{\mathrm{BPTT}}+\varepsilon)}$
+Here, $C_{\mathrm{tot}}$ is a constant that depends on $C_{F}$ and the constant $C_{v}^{\prime\prime}$ (derived from Proposition C.2 in Appendix C).
+The proof of Theorem 1 follows a logical chain: the FOC-gap $(\leq\varepsilon)$ implies that errors in PDE coefficients and source terms are of order $O(\varepsilon)$.
+These, in turn, through PDE stability, imply that the error in the value function’s gradient $E_{v}=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}}$ is also of order $O(\varepsilon)$.
+Combining this with the BPTT estimation error $\delta_{\mathrm{BPTT}}$ leads to the final policy-gap bound.
+The detailed arguments, including the necessary lemmas and propositions, are provided as sketches in Appendix C.
+It is important to note that all constants involved depend quantitatively on model parameters, PDE regularity, and utility specifications.
+Full rigorous proofs would require meticulous tracking of these dependencies and bounds.
+Our formal analysis (detailed in Appendix C) demonstrates that under suitable regularity conditions, particularly uniform parabolicity, a small Pontryagin first-order condition (FOC) gap ($\varepsilon$) achieved during the Stage 1 warm-up guarantees that the BPTT-derived costate estimates are $O(\varepsilon)$-close to the true optimal costates (up to the BPTT estimation error $\delta_{\mathrm{BPTT}}$).
+This provides a rigorous justification for the core idea of the P-PGDPO method: substituting these rapidly stabilizing and sufficiently accurate costate estimates into the analytical PMP formulas yields a near-optimal policy $\widehat{\pi}$ with an error bound of $O(\varepsilon+\delta_{\mathrm{BPTT}})$ (Theorem 1).
+However, extending this rigorous proof framework to settings beyond uniform parabolicity, such as hypo-elliptic (H-class, (Hörmander, , 1967)) or boundary-degenerate (D-class, (Oleinik, , 2012)) models, or to general non-affine systems, faces significant analytical challenges, as key assumptions (like global $L^{p}$ regularity or the costate floor) may no longer hold.
+Despite these formal hurdles, the fundamental algorithmic rationale underpinning the P-PGDPO approach remains persuasive for broader applicability.
+Neither Pontryagin’s Maximum Principle (PMP) (Pontryagin, , 2014; Yong and Zhou, , 2012) nor backpropagation-through-time (BPTT) (Goodfellow, , 2016) are inherently limited to uniformly parabolic or affine settings.
+The P-PGDPO strategy leverages the potentially advantageous decoupling of the problem into: (i) estimating potentially simpler or faster-converging costate functions via BPTT, and (ii) utilizing the robust analytical structure of PMP to construct the control.
+This rationale, strongly supported by the numerical success demonstrated even in high dimensions for the affine benchmark case (Section 4), suggests that PG-DPO (especially the P-PGDPO variant) holds promise as a practical tool for a wider class of challenging continuous-time control problems.
+Nonetheless, confirming its effectiveness and potentially refining the theoretical guarantees for models outside the strictly regular class analyzed here remains an important avenue for future research, likely requiring tailored analysis or different analytical techniques (e.g., focusing on regions predominantly visited by the state process).
 4 Numerical Experiments
-In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework (both Baseline and Two-Stage variants).
+In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework (both Baseline and Projected PG-DPO (P-PGDPO) variants).
 We focus on the multi-asset, multi-factor Ornstein-Uhlenbeck (OU) portfolio optimization problem introduced in Section 2.3.2.
 The objective is to maximize the expected utility of terminal wealth, $J=\mathbb{E}[U(X_{T})]$, using a CRRA utility function with risk aversion $\gamma=2.0$.
+For these experiments, we specialize the general model by assuming no intermediate consumption (i.e., $C_{t}\equiv 0$) and focus solely on the utility from terminal wealth (effectively setting the bequest weight $\kappa=1$ if $U(X_{T})$ is the direct objective, or adjusting $\kappa$ appropriately if a specific HARA utility form with $a\neq 0$ was being precisely mapped).
 This specific affine model setting allows for comparison against a semi-analytical benchmark solution derived from the HJB equation (details in Appendix B.2).
-We compare the policies generated by Baseline PG-DPO and Two-Stage PG-DPO against this benchmark using the Root Mean Squared Error (RMSE) metric across various dimensions, scaling up to n=50 assets and k=10 state factors.
-For complete details on the specific parameter values, PG-DPO implementation (network architecture, training setup), benchmark computation, and evaluation procedures, please refer to Appendix C.
-The subsequent subsections present and analyze the results.
+Furthermore, to provide a comprehensive assessment against other contemporary deep learning techniques for high-dimensional control problems, we also compare our PG-DPO framework with the Deep BSDE method (Han et al., , 2018; Weinan, , 2017).
+The Deep BSDE approach reformulates the stochastic control problem in terms of a system of backward stochastic differential equations, approximating the unknown components of the solution (typically related to the value function gradient or control variables) using neural networks.
+These networks are trained by minimizing the error incurred in satisfying the BSDE dynamics and terminal conditions through Monte Carlo simulation.
+We compare the policies generated by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and the Deep BSDE method against the HJB-based benchmark using the Root Mean Squared Error (RMSE) metric across various dimensions, scaling up to n=50 assets and k=10 state factors.
+For complete details on the specific parameter values, the implementations of PG-DPO and Deep BSDE (including network architectures and training setups), benchmark computation, and evaluation procedures, please refer to Appendix D.
+The subsequent subsections present and analyze the results of these comparative experiments.
 4.1 Convergence Speed and Accuracy Analysis
-This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO methodologies across various dimensional settings.
-Figure 1 displays the average Root Mean Squared Error (RMSE), averaged across $k=10$ evaluation slices, between the learned policies (either $\bm{\pi}_{\theta}$ for Baseline PG-DPO or $\bm{\pi}^{\mathrm{PMP}}$ for Two-Stage PG-DPO) and the benchmark optimal policy ($\bm{\pi}^{*}$).
-The RMSE is plotted as a function of training epochs on a logarithmic scale for configurations with $k=10$ state variables and varying numbers of assets ($n=1,10,50$).
 (a) $n=1,\;k=10$
@@
 (c) $n=50,\;k=10$
-Figure 1: Average policy RMSE comparison between Baseline PG-DPO (blue) and Two-Stage PG-DPO (orange) relative to the benchmark solution across different dimensions (k=10 fixed).
-Values averaged over k evaluation slices.
+Figure 1: Average policy RMSE comparison between Baseline PG-DPO (e.g., blue), Projected PG-DPO (P-PGDPO) (e.g., orange), and Deep BSDE (e.g., green) relative to the benchmark solution across different asset dimensions ($n$) with $k=10$ factors fixed.
+Values averaged over $k$ evaluation ‘slices.’ For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation or visualization) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$.
+Further details on the evaluation methodology are provided in Appendix D.
 Note the logarithmic scale on the $y$-axis.
-The RMSE for the Two-Stage PG-DPO (orange line in Figure 1) represents the error of the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ constructed by substituting the BPTT-derived costate estimates into the PMP first-order condition (Eq.
-(10)).
-A key advantage is that the Two-Stage approach explicitly combines the analytically distinct myopic and hedging components.
-As established theoretically and confirmed numerically (see Section 4.2), the error of the myopic component is typically negligible (e.g., on the order of $10^{-7}$).
-Therefore, the total RMSE for the Two-Stage policy is almost entirely attributable to the error in estimating the hedging component.
-This highlights the proficiency of the Two-Stage method in accurately capturing hedging demands using BPTT-derived costate derivatives ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$).
-Table 2: Summary of Minimum Achieved Policy RMSE and Convergence Epochs
+This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO framework (Baseline and Projected PG-DPO (P-PGDPO) variants) against the Deep BSDE method.
+We analyze their performance across various dimensional settings, focusing on how the number of assets ($n$) and the number of state factors ($k$) impact accuracy.
+Figure 1 (Note: Assuming you will update the figure label in your main document if it includes Deep BSDE results) displays the average Root Mean Squared Error (RMSE) between the learned/derived policies and the benchmark optimal policy.
+The RMSE is plotted against training epochs on a logarithmic scale for configurations with $k=10$ state variables and varying numbers of assets ($n=1,10,50$).
+The RMSE for P-PGDPO represents the error of the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ (Eq.
+(11)), which combines myopic and hedging components.
+As detailed in Section 4.2, the myopic component’s error is typically negligible ($O(10^{-7})$).
+Thus, the total RMSE for P-PGDPO is almost entirely driven by the accuracy of estimating the hedging component using BPTT-derived costate derivatives, $\partial_{\mathbf{Y}}\lambda_{t}^{*}$.
+Table 2 summarizes the minimum achieved policy RMSE and the epochs at which these minima were observed for the $k=10$ scenario with varying $n$.
+A consistent observation is the superior performance of the P-PGDPO method, which achieves significantly lower RMSE values compared to both Baseline PG-DPO and Deep BSDE across all tested asset dimensions.
+For instance, with $n=1,k=10$, P-PGDPO’s minimum RMSE is approximately an order of magnitude better than Deep BSDE and two orders of magnitude better than Baseline PG-DPO.
+Deep BSDE generally performs between the Baseline and P-PGDPO methods.
+While P-PGDPO maintains its lead as $n$ increases to 50, the performance gap between P-PGDPO and Deep BSDE tends to narrow, suggesting that the increasing dimensionality of $n$ poses challenges even for the PMP-based analytical solution, possibly due to the ill-conditioning of the covariance matrix $\Sigma$.
+All methods show an increase in RMSE as $n$ grows, indicating the inherent difficulty of high-dimensional asset allocation.
+To further investigate the distinct effects of asset dimensionality ($n$) versus factor dimensionality ($k$), we conducted two additional sets of experiments.
+Table 2: Summary of Minimum Policy RMSE and Convergence Epochs
 Dimension Method Min. Policy RMSE Epochs at Min.
-(n assets, k factors) (vs. Benchmark)
-n=1, k=10 Baseline PG-DPO $4.27\times 10^{-2}$ $2,600$
-2-PG-DPO $2.80\times 10^{-4}$ $1,500$
-n=10, k=10 Baseline PG-DPO $1.58\times 10^{-1}$ $10,000$
-2-PG-DPO $4.89\times 10^{-3}$ $800$
-n=50, k=10 Baseline PG-DPO $1.26\times 10^{-1}$ $100,000$
-2-PG-DPO $1.32\times 10^{-2}$ $3,200$
-Note: Minimum Policy RMSE values are averaged across the $k=10$ evaluation slices.
+($n$ assets, $k=10$ factors)
+n=1, k=10 Baseline PG-DPO $1.75\times 10^{-2}$ 2800
+P-PGDPO $4.72\times 10^{-4}$ 1200
+Deep BSDE $4.78\times 10^{-3}$ 600
+n=10, k=10 Baseline PG-DPO $5.11\times 10^{-2}$ 9600
+P-PGDPO $5.26\times 10^{-3}$ 1000
+Deep BSDE $3.46\times 10^{-2}$ 1400
+n=50, k=10 Baseline PG-DPO $1.05\times 10^{-1}$ 8800
+P-PGDPO $1.58\times 10^{-2}$ 5400
+Deep BSDE $4.90\times 10^{-2}$ 9200
+Note: Minimum Policy RMSE values are averaged across $k$ evaluation ’slices.’ For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$.
+Further details on the evaluation methodology are provided in Appendix D.
 Epochs indicate the recorded iteration number when the minimum average RMSE was observed; evaluations and logging were performed every 200 iterations.
-The key finding from Figure 1 and summarized in Table 2 is the consistent and significant outperformance of the Two-Stage PG-DPO method.
-Table 2 shows that Two-Stage PG-DPO achieves minimum RMSE values dramatically lower than the Baseline method across all dimensions—approximately 150 times lower for $n=1$, 30 times lower for $n=10$, and nearly 10 times lower for $n=50$.
-Furthermore, 2-PG-DPO typically reaches its minimum error much earlier in the training process (e.g., at epoch 1,500 vs.
-2,600 for $n=1,k=10$, and epoch 800 vs.
-10,000 for $n=10,k=10$).
-Notably, while the convergence time for Baseline PG-DPO appears to increase substantially with the problem dimension (n), the time required for 2-PG-DPO to reach its minimum error seems relatively less sensitive to the number of assets.
-Even in the high-dimensional $n=50,k=10$ setting, 2-PG-DPO finds a near-optimal solution with an RMSE of $1.32\times 10^{-2}$ around epoch 3200, whereas the Baseline method’s minimum RMSE of $1.26\times 10^{-1}$ is reached only at the end of a much longer training period (epoch 100,000).
-This clearly demonstrates the superiority of the Two-Stage approach and strongly supports the rationale discussed in Section 3.3: leveraging the PMP structure via BPTT-estimated costates is more efficient and accurate than end-to-end policy approximation, especially as dimensionality increases.
-The subsequent section presents a detailed visual analysis for the most challenging high-dimensional case ($n=50,k=10$) to further illustrate the performance difference.
-Hereafter, we often abbreviate Two-Stage PG-DPO as 2-PG-DPO.
+First, Table 3 presents results where the number of assets is fixed at $n=1$, and the number of state factors $k$ is varied ($k=1,5,10$).
+In this scenario, the performance of all three methodologies remains relatively stable across different values of $k$.
+For P-PGDPO, the RMSE even shows a slight improvement as $k$ increases, possibly due to more information being available for costate estimation.
+Deep BSDE and Baseline PG-DPO also exhibit no significant degradation in performance.
+This suggests that, for a small number of assets, the ”curse of dimensionality” associated with an increasing number of state factors ($k$) is largely managed by these neural network-based approaches.
+The underlying PMP structure or BSDE formulation appears robust to the increase in $k$ when $n$ is small and fixed, with Monte Carlo sampling and function approximation being the primary challenges rather than ill-conditioned linear algebra.
+Table 3: Summary of Minimum Policy RMSE and Convergence Epochs for $n=1$ (Varying $k$)
+$k$ factors Method Min. Policy RMSE Epochs at Min.
+($n=1$ asset)
+1 Baseline PG-DPO $3.26\times 10^{-2}$ 1000
+P-PGDPO $1.20\times 10^{-3}$ 800
+Deep BSDE $7.39\times 10^{-3}$ 800
+5 Baseline PG-DPO $3.71\times 10^{-2}$ 1400
+P-PGDPO $2.64\times 10^{-4}$ 800
+Deep BSDE $1.02\times 10^{-2}$ 1400
+10 Baseline PG-DPO $1.75\times 10^{-2}$ 2800
+P-PGDPO $4.72\times 10^{-4}$ 1200
+Deep BSDE $4.78\times 10^{-3}$ 600
+Table 4: Summary of Minimum Policy RMSE and Convergence Epochs for $k=2$ (Varying $n$)
+$n$ assets Method Min. Policy RMSE Epochs at Min.
+($k=2$ factors)
+1 Baseline PG-DPO $4.77\times 10^{-2}$ 1200
+P-PGDPO $1.98\times 10^{-4}$ 800
+Deep BSDE $4.63\times 10^{-3}$ 3400
+10 Baseline PG-DPO $1.12\times 10^{-1}$ 4000
+P-PGDPO $1.22\times 10^{-2}$ 1400
+Deep BSDE $2.32\times 10^{-2}$ 2600
+50 Baseline PG-DPO $2.24\times 10^{-1}$ 7200
+P-PGDPO $5.55\times 10^{-2}$ 4400
+Deep BSDE $1.88\times 10^{-1}$ 4400
+Second, Table 4 shows results where the number of state factors is fixed at $k=2$, and the number of assets $n$ is varied ($n=1,10,50$).
+In stark contrast to the previous case, increasing $n$ leads to a significant deterioration in performance for all three methods.
+Baseline PG-DPO shows the most substantial increase in RMSE.
+While P-PGDPO still yields the lowest RMSE, its error also increases considerably with $n$.
+Deep BSDE’s performance also degrades, approaching that of Baseline PG-DPO for $n=50$.
+This pronounced degradation with increasing $n$ (even for a small $k$) points towards a different kind of challenge not solely related to the dimensionality of the state space for function approximation.
+A primary suspect is the numerical instability arising from the asset covariance matrix $\Sigma$.
+As $n$ increases, $\Sigma$ ($n\times n$) is more likely to become ill-conditioned (i.e., have a high condition number).
+The optimal policy calculation, particularly in the PMP framework (Eq.
+(11)), involves the inverse $\Sigma^{-1}$.
+If $\Sigma$ is ill-conditioned, any small errors in the estimated costates or market parameters can be greatly amplified by $\Sigma^{-1}$, leading to large errors in the computed policy.
+This suggests that the “curse” associated with large $n$ is not just about learning a higher-dimensional policy function but also about the increasing fragility of the underlying mathematical problem structure due to the properties of $\Sigma$.
+The problem itself becomes “damaged” or numerically challenging, impacting any method that relies on or implicitly reflects this structure.
+These findings underscore that while neural network approaches have made significant strides in handling high-dimensional state spaces (related to $k$), the challenges posed by a large number of decision variables (assets $n$), especially when intertwined with ill-conditioned covariance matrices, require further attention.
+Strategies such as covariance matrix shrinkage or other regularization techniques applied prior to or within the optimization might be necessary to effectively address this “asset-dimension curse”.
 4.2 High-Dimensional Performance Analysis ($n=50,k=10$)
-We now visually examine the performance in the most challenging setting: $n=50$ assets and $k=10$ state variables.
-This scenario tests scalability where traditional DP methods are intractable.
-As shown quantitatively in Table 2 and Figure 1c, the 2-PG-DPO method achieves a minimum RMSE ($1.32\times 10^{-2}$ at epoch 3,200) significantly better than the Baseline approach ($1.26\times 10^{-1}$ at epoch 100,000).
-The visualizations in Figures 2 and 3 focus on the policy for the first asset ($i=1$) versus time-to-maturity ($T-t$) and the first state variable ($Y_{1}$), holding other state variables constant at their long-term means.
-The results shown correspond to the state on epoch 3,200, near where the 2-PG-DPO method achieved its minimum observed RMSE.
+We now visually examine the performance of the three methodologies in the most challenging setting considered: $n=50$ assets and $k=10$ state factors.
+This high-dimensional scenario, where traditional DP methods are intractable, serves to highlight the distinct capabilities and limitations of each approach.
+As a reference, Figure 2 displays the benchmark analytical solution for the first asset ($\pi_{1}^{\text{sol}}$), decomposed into its myopic ($\pi_{1,0}^{\text{sol}}$) and intertemporal hedging ($\pi_{1,1}^{\text{sol}}$) components.
+The hedging component is clearly non-trivial and exhibits complex dependencies on both the state variable $Y_{1}$ and time-to-maturity, underscoring its importance for optimal allocation.
+(a) Benchmark Policy ($\pi_{1}^{\text{sol}}$)
+(b) Myopic Comp. ($\pi_{1,0}^{\text{sol}}$)
+(c) Hedging Comp. ($\pi_{1,1}^{\text{sol}}$)
+Figure 2: Benchmark (analytical) solution components for the first asset ($i=1$) in the $n=50,k=10$ case, plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (e.g., $t=(T-T_{0})/2$ if $T_{0}$ is initial time, assuming $T=1.5$ is the horizon from $t=0$; thus, for instance, at $t=0.75$ or time-to-maturity $0.75$).
+Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
+From left to right: total optimal policy, its myopic component, and its intertemporal hedging component.
+These serve as the ground truth.
+Figure 3 provides a side-by-side comparison of the policies learned by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and Deep BSDE, alongside their respective errors, for the first asset.
+These plots correspond to iterations where each method achieved its (near) optimal performance in the $n=50,k=10$ setting, as indicated in Table 2 (e.g., iteration 5400 for P-PGDPO, which is the nearest logged checkpoint to the observed minimum RMSE at iteration 5399, as evaluations were performed every 200 iterations).
 (a) Baseline Policy ($\pi_{1,\theta}$)
-(b) 2-PG-DPO Policy ($\pi_{1}^{\mathrm{PMP}}$)
-(c) Baseline Error
-(d) 2-PG-DPO Error
-Figure 2: Comparison of total policies and errors for the first asset ($i=1$) in the n=50, k=10 case on epoch 3,200 (plotted against $Y_{1}$).
-Top row compares policies, bottom row compares errors (using Symlog scale).
-Note the significantly different color bar ranges are used for each plot to best visualize its content, reflecting the dramatic difference in policy/error magnitudes.
-(a) 2-PG-DPO Myopic ($\pi_{1,0}^{\mathrm{PMP}}$)
-(b) 2-PG-DPO Hedging ($\pi_{1,1}^{\mathrm{PMP}}$)
+(b) Baseline Error
+(c) P-PGDPO Policy ($\pi_{1}^{\mathrm{PMP}}$)
+(d) P-PGDPO Error
+(e) Deep BSDE Policy ($\pi_{1}^{\mathrm{BSDE}}$)
+(f) Deep BSDE Error
+Figure 3: Comparison of total policies (left column) and policy errors (right column) for the first asset ($i=1$) across Baseline PG-DPO (top row), Projected PG-DPO (P-PGDPO) (middle row), and Deep BSDE (bottom row) in the $n=50,k=10$ case.
+Results shown at (near) optimal iterations for each method (e.g., iteration 5400 for P-PGDPO), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2).
+Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
+Errors use a symmetric logarithmic (symlog) scale.
+Different color bar ranges reflect varying policy/error magnitudes.
+The Baseline PG-DPO method (Figure 3a,b), despite utilizing a standard neural network architecture (detailed in Appendix D.2), struggles significantly.
+While its learned policy exhibits some non-linear structure, it bears little resemblance to the benchmark, and the corresponding error is extremely large.
+This performance visually confirms the poor quantitative results in Table 2 and highlights the difficulty of directly approximating the high-dimensional optimal policy function using a neural network without leveraging more structural information from the problem.
+(a) P-PGDPO Myopic ($\pi_{1,0}^{\mathrm{PMP}}$)
+(b) P-PGDPO Hedging ($\pi_{1,1}^{\mathrm{PMP}}$)
 (c) Myopic Error
 (d) Hedging Error
-Figure 3: Decomposition of the 2-PG-DPO policy for the first asset ($i=1$) in the n=50, k=10 case on epoch 3,200 (plotted against $Y_{1}$).
-Top row shows myopic and hedging components.
-Bottom row shows their respective errors (using Symlog scale).
-Note the use of different color bar ranges, necessary to visualize both the negligible myopic error and the larger hedging error, which dictates the total policy error.
-The challenges of high dimensionality become starkly evident for the Baseline PG-DPO method in Figure 2.
-While the learned policy (2(a)) shows some structure, the corresponding error (2(c)) is extremely large, visually confirming the poor quantitative performance reported in Table 2.
-This indicates severe inaccuracies in the policy learned by the Baseline method when faced with the combined complexity of 50 assets and 10 state variables.
-In dramatic contrast, the 2-PG-DPO method provides a highly accurate solution, as shown by its policy (2(b)) and error (2(d)).
-Comparing the value ranges indicated by the color bars, the error magnitude for the 2-PG-DPO policy is roughly an order of magnitude smaller than the Baseline error, underscoring the robustness and practical necessity of the 2-PG-DPO approach.
-Although perfect accuracy might be harder to achieve as dimensionality increases (the minimum RMSE of $1.32\times 10^{-2}$ is higher than in lower-dimensional cases), it remains remarkably effective and vastly superior to the direct policy learning method.
-The component decomposition in Figure 3 confirms the patterns observed previously.
-The myopic component (3(a)) largely dictates the policy shape, and its associated error (3(c)) remains negligible (confirming quantitative findings that this error is often $O(10^{-7})$), requiring a very small scale on its color bar for visualization.
-The hedging component (3(b)) contributes visibly to the total policy shape.
-Crucially, the hedging error (3(d)) is visually almost identical in pattern to the total 2-PG-DPO error (2(d)) when comparing their respective plots, confirming that the overall accuracy is dictated by the precision of estimating the hedging demands.
-In conclusion, the $n=50,k=10$ results powerfully demonstrate the capability of the 2-PG-DPO framework to “break the dimensional barrier”.
-While the Baseline PG-DPO struggles significantly, 2-PG-DPO delivers a highly accurate and viable solution by leveraging the structure of Pontryagin’s Maximum Principle.
-Its performance remains remarkably strong and vastly superior to the baseline, solidifying its potential as a practical tool for large-scale, continuous-time portfolio optimization.
+Figure 4: Decomposition of the Projected PG-DPO (P-PGDPO) policy for the first asset ($i=1$) in the $n=50,k=10$ case (results from iteration 5400), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2).
+Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
+Top row: myopic and hedging components.
+Bottom row: their respective errors (symmetric logarithmic (symlog) scale).
+Note different color bar ranges.
+The Deep BSDE method (Figure 3e,f) shows a marked improvement over the Baseline.
+Its learned policy qualitatively resembles the myopic component of the benchmark (cf.
+Figure 2b), primarily capturing dependencies on the state variable $Y_{1}$.
+However, the policy largely fails to capture the complex intertemporal hedging demands.
+The error plot for Deep BSDE (Figure 3f) shows significant discrepancies in regions where hedging is important, suggesting that the hedging component is largely absent from the Deep BSDE solution.
+This difficulty in capturing hedging demands can be attributed to several factors.
+The Deep BSDE method primarily focuses on matching the terminal condition and ensuring consistency of the value process $Y_{t}$ and the gradient-related process $Z_{t}$ (approximating first-order derivatives of the value function) through its loss function.
+Intertemporal hedging, however, critically depends on the mixed second-order derivative $V_{x\mathbf{y}}$, which is not explicitly represented or directly targeted by the standard two-network Deep BSDE architecture.
+Furthermore, the loss function penalizes inaccuracies in $Z_{t}$ primarily through a quadratic martingale term.
+Consequently, small errors in estimating the gradient components relevant to $V_{x\mathbf{y}}$ might have a limited impact on the overall training objective but can be significantly magnified when constructing the policy, especially after multiplication by the potentially ill-conditioned inverse covariance matrix $\Sigma^{-1}$ in high-dimensional settings.
+As a result, while the Deep BSDE approach can effectively learn the dominant myopic component of the policy, it may struggle to precisely fit the more nuanced intertemporal hedging patterns.
+In stark contrast, the Projected PG-DPO (P-PGDPO) method (Figure 3c,d) delivers a policy that is visually almost identical to the benchmark total policy (Figure 2a).
+The corresponding error is substantially smaller than that of the other two methods.
+This superior performance underscores the efficacy of its decoupled learning structure: first estimating costates via BPTT and stabilizing them through Monte Carlo averaging, then analytically constructing the policy using the PMP FOCs.
+The remarkable accuracy of P-PGDPO is further elucidated by decomposing its policy into myopic and hedging components, as shown in Figure 4 (results from iteration 5400).
+The myopic component (Figure 4(a)) is virtually identical to the benchmark myopic component, with its error (Figure 4(c)) being negligible ($O(10^{-7})$).
+The hedging component derived by P-PGDPO (Figure 4(b)) also successfully captures the complex shape of the benchmark hedging demand (Figure 2c).
+The remaining error in the hedging component (Figure 4(d)) dictates the overall error of the P-PGDPO policy.
+This residual hedging error is likely not due to a fundamental flaw in the P-PGDPO algorithm itself but rather a consequence of the numerical challenges inherent in high-dimensional asset allocation.
+Specifically, the hedging term in the PMP FOC involves $\Sigma^{-1}$:
+$\pi^{\text{hedge}}=-\frac{1}{X_{t}V_{xx}}\,\underbrace{\Sigma^{-1}}_{\text{cond num can be large for }n\uparrow}(\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y})\,V_{x\mathbf{Y}}.$
+For $n=50$, the $50\times 50$ covariance matrix $\Sigma$ can easily become ill-conditioned.
+Even if the BPTT-derived estimate for $V_{x\mathbf{Y}}$ is quite accurate, any small residual noise can be significantly amplified when multiplied by $\Sigma^{-1}$ if its condition number is large.
+This amplification effect is the primary suspect for the observed hedging error.
+In conclusion, the $n=50,k=10$ visualizations powerfully demonstrate that Projected PG-DPO (P-PGDPO) can effectively ”break the dimensional barrier” where Baseline PG-DPO fails and Deep BSDE offers only a partial (myopic-focused) solution.
+The remaining inaccuracies in the hedging component point towards the need for addressing the ill-conditioning of the covariance matrix in very high asset dimensions.
 5 Conclusion
-This paper tackled the significant challenge posed by the “curse of dimensionality” in solving large-scale continuous-time portfolio optimization problems.
-We introduced and developed the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, which uniquely combines the theoretical rigor of Pontryagin’s Maximum Principle (PMP) with the power of deep learning, specifically using backpropagation-through-time (BPTT) applied to policy networks representing investment and consumption strategies.
-A central contribution is the development of a highly effective “Two-Stage” PG-DPO variant.
-This method leverages the insight that BPTT applied to simulation paths can yield stable estimates of the crucial Pontryagin costates (and their derivatives) relatively quickly, often much faster than the full policy network converges.
-By utilizing these rapidly obtained costate estimates within the analytical first-order optimality conditions derived from PMP, the Two-Stage approach efficiently constructs near-optimal control policies, effectively decoupling the estimation of marginal value sensitivities from the full policy optimization.
-Our numerical experiments, conducted on a multi-asset, multi-factor affine model chosen for its tractability allowing rigorous benchmark comparisons, demonstrated the practical effectiveness and scalability of this approach.
-The results consistently showed that the 2-PG-DPO method achieves substantially higher accuracy and faster convergence compared to the baseline end-to-end policy learning strategy.
-Notably, the framework successfully scaled to problems involving dimensions (e.g., n=50 assets, k=10 state factors) that are typically considered intractable for traditional numerical HJB/PDE solvers, thereby effectively “breaking the dimensional barrier.” We specifically highlighted the method’s capability to accurately capture both myopic and complex intertemporal hedging demands by explicitly leveraging the structure inherent in the PMP optimality conditions.
-While demonstrated on an affine model, the underlying principles relying on PMP and BPTT suggest broader applicability.
-This research opens several avenues for future investigation.
-Firstly, further exploration of the theoretical underpinnings is crucial.
-Our analysis, related to Proposition 3.1, indicates that the rationale behind 2-PG-DPO—leveraging the stability of value function derivatives—is theoretically supported for potentially broad classes of models, possibly including non-affine cases, provided certain regularity conditions related to smoothness and parabolicity are met.
-A key avenue for future theoretical research is to rigorously characterize the precise range of financial models (both affine and non-affine) that satisfy these sufficient regularity assumptions, thereby ensuring the well-posedness of the PMP framework and the reliability of BPTT-based estimates.
-Furthermore, investigating the general theoretical convergence properties of the PG-DPO algorithm, particularly the costate estimation process, remains an important open question.
-Secondly, extending the framework to incorporate more realistic market features is a natural next step for enhancing practical relevance.
-This includes adapting PG-DPO to handle portfolio constraints (like short-selling restrictions or leverage limits) and modeling market frictions such as transaction costs, both of which significantly alter the optimal control landscape.
-Testing the framework’s robustness and performance with different utility functions beyond CRRA (e.g., HARA, prospect theory) is also warranted.
-Moreover, explicitly including the optimization and evaluation of the intermediate consumption policy, omitted in the current study for benchmarking clarity, would provide a more complete assessment of the full framework’s capabilities.
-Thirdly, further empirical validation and algorithmic refinements are desirable.
-Building on the theoretical potential for broader applicability, applying the PG-DPO framework (especially the superior 2-PG-DPO variant) to well-known non-affine financial models (e.g., stochastic volatility models like Heston) would be valuable, even in the absence of exact benchmarks, to showcase its capabilities more broadly.
-Algorithmically, exploring alternative neural network architectures, advanced optimization techniques, and variance reduction methods beyond antithetic variates could potentially lead to further performance improvements or faster convergence.
-Finally, optimizing the implementation details deserves attention.
-This includes refining strategies for determining the optimal warm-up duration for the 2-PG-DPO method to best balance costate accuracy against potential overfitting to simulation artifacts inherent in the discrete-time training process.
-Careful tuning of hyperparameters related to both the network training and the costate estimation phase will be important for practical applications.
+This paper confronted the enduring “curse of dimensionality” in continuous-time portfolio optimization, a challenge that manifests distinctly with increasing asset numbers ($n$) versus state factors ($k$).
+Our numerical investigations (Section 4) revealed that while deep learning methods manage larger $k$ relatively well, escalating $n$ poses severe difficulties, particularly due to the ill-conditioning of asset covariance matrices.
+Furthermore, we identified that prevalent approaches like Deep BSDE, while proficient at capturing myopic policy components, often fail to adequately represent crucial intertemporal hedging demands, a limitation rooted in their network architecture and loss functions which do not explicitly target the necessary second-order mixed derivatives of the value function (e.g., $V_{x\mathbf{y}}$).
+To surmount these obstacles, we introduced the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, with its highly effective Projected PG-DPO (P-PGDPO) variant (Section 3.2).
+P-PGDPO distinctively leverages Pontryagin’s Maximum Principle (PMP) by first employing backpropagation-through-time (BPTT) to achieve rapidly stabilizing estimates of the Pontryagin costates and their derivatives (e.g., $\lambda_{t},\partial_{x}\lambda_{t},\partial_{\mathbf{Y}}\lambda_{t}$).
+These estimates are then analytically projected onto the manifold of PMP-optimal controls via the first-order conditions.
+This decoupling strategy proved remarkably successful, enabling P-PGDPO to tackle problems of previously prohibitive scales.
+For instance, in a challenging $n=50$ asset, $k=10$ factor environment, P-PGDPO achieved a total policy RMSE of $\mathbf{1.58\times 10^{-2}}$ relative to the analytical benchmark, with the myopic component error being negligible ($O(10^{-7})$) and, critically, accurately reconstructing complex hedging terms (Section 4.2, Figure 4).
+The theoretical underpinnings of P-PGDPO are strengthened by our Policy Gap Bound (Theorem 1), which guarantees near-optimal policies given accurate costate estimation and a small FOC gap.
+Despite its successes, P-PGDPO’s accuracy in representing hedging terms can be affected by the amplification of estimation noise through the inverse covariance matrix $\Sigma^{-1}$ when $n$ is large and $\Sigma$ is ill-conditioned (Section 4.2).
+Concurrently, our experiments confirmed that Deep BSDE, while adept at matching myopic demands, struggles with learning the nuanced cross-derivative information essential for hedging, largely due to its architectural and objective function design (Figure 3).
+These observations highlight the distinct strengths and weaknesses of different deep learning paradigms in high-dimensional control.
+The practical implications of P-PGDPO are significant.
+Its demonstrated ability to recover complex, non-linear hedging strategies in high dimensions suggests its potential applicability to large institutional portfolios, such as those involving numerous ETFs or derivative hedging.
+Moreover, the core principle of P-PGDPO—“estimate costates first, then derive policy”—holds promise for a broader class of problems, where PMP provides the necessary optimality structure (Section 3.3).
+Future research will pursue several promising directions:
+  • Developing “P-PGDPO 2.0” by integrating $\Sigma$ regularization or shrinkage techniques to mitigate the impact of ill-conditioned covariance matrices on hedging term estimation.
+  • Extending the P-PGDPO framework to incorporate portfolio constraints (e.g., no short-selling, leverage limits) and market frictions like transaction costs, by adapting the PMP conditions accordingly.
+  • Investigating the performance of P-PGDPO with non-HARA utility functions and behavioral finance objectives, and fully integrating the optimization of intermediate consumption policies ($C_{t}$).
+  • Validating the framework through rigorous backtesting on historical and simulated market data for non-affine models.
+  • Rigorously characterizing the precise range of financial models that satisfy the sufficient regularity assumptions for the P-PGDPO framework’s effectiveness.
+In essence, PG-DPO, particularly its projected variant, demonstrates that the synergy between deep learning and the structural guidance of Pontryagin’s Maximum Principle can redefine the boundaries of solvability for high-dimensional continuous-time financial control problems.
 Acknowledgments
@@
 For simplicity, we assume no intermediate consumption, i.e., $C_{t}=0$, and consider an investor who derives utility only from terminal wealth $X_{T}$ under a CRRA objective $U(x)=\frac{x^{1-\gamma}}{1-\gamma}$.
-Following Kim & Omberg, (1996), the associated Hamilton–Jacobi–Bellman (HJB) equation can be solved by positing a value function of the form:
+Following Kim and Omberg, (1996), the associated Hamilton–Jacobi–Bellman (HJB) equation can be solved by positing a value function of the form:
 $V(t,x,Y)=\frac{x^{1-\gamma}}{1-\gamma}\exp\left(g(t)+B(t)Y+\frac{1}{2}C(t)Y^{2}\right).$ (18)
@@
 (23) gives the closed-form value function, while substituting $\mathbf{B}(t)$ and $\mathbf{C}(t)$ into $\bm{\pi}^{*}(t,\mathbf{Y})$ gives the optimal strategy, explicitly capturing both myopic and intertemporal hedging components in a multi-asset, multi-factor setting.
-Appendix C Detailed Experimental Setup
+Appendix C Detailed Proofs for the Policy Gap Bound (Theorem 1) and Supporting Results
+This appendix provides the detailed mathematical arguments supporting the Policy Gap Bound stated in Theorem 1 (see Section 3.3).
+These details include the precise statements of the necessary assumptions (Assumptions 6 and 7), and the full proof sketches for the lemmas, propositions, and Theorem 4 that collectively establish the theoretical justification for the Projected PG-DPO approach.
+The core result demonstrates that if the policy $(\pi,c)$ obtained from the Stage 1 warm-up phase satisfies a sufficiently small Pontryagin first-order condition (FOC) residual $\varepsilon$, then the feedback policy $\widehat{\pi}$ constructed in Stage 2 using BPTT estimates $\widehat{\lambda}$ (with error $\delta_{\mathrm{BPTT}}$) will be close to the true optimal policy $\pi^{*}$.
+Objective and Roadmap.
+This section provides the mathematical justification for the Projected PG-DPO approach.
+We establish that if the policy $(\pi,c)$ obtained from the Stage 1 warm-up phase satisfies a sufficiently small Pontryagin first-order condition (FOC) residual $\varepsilon$, then the feedback policy $\widehat{\pi}$ constructed in Stage 2 using BPTT estimates $\widehat{\lambda}$ (with error $\delta_{\mathrm{BPTT}}$) will be close to the true optimal policy $\pi^{*}$.
+Our proof follows the logical chain:
+[…]
+The core involves establishing bounds via lemmas and propositions, resolving coupled estimates, and concluding with the main theorem.
+1. Preliminaries: PDE, Norms, and Regularity.
+We work within the framework established in Section 2. For any admissible policy $(\pi,c)$, the associated sub-value function $V^{\pi,c}$ solves the linear parabolic PDE:
+$-\partial_{t}V^{\pi,c}=\mathcal{L}_{\pi,c}[V^{\pi,c}]+f_{\pi,c},\quad V^{\pi,c}(T,z)=g(z),$ (27)
+where the operator $\mathcal{L}_{\pi,c}$ and source term $f_{\pi,c}$ depend on the policy $(\pi,c)$.
+This dependence occurs through specific functions determining the PDE coefficients and the source term.
+For the purpose of tracking policy influence in our analysis, we conceptually group the key policy-dependent components under the label (27) below :
+  • Wealth Drift Influence: Terms multiplying $V_{x}$ dependent on $\pi_{t},C_{t}$.
+  • Wealth Diffusion Influence: Term multiplying $V_{xx}$ dependent on $\bm{\pi}_{t}$.
+  • Wealth-State Cross-Diffusion Influence: Terms multiplying $V_{xy_{j}}$ dependent on $\bm{\pi}_{t}$.
+  • Consumption Source Term: $f_{\pi,c}=e^{-\delta t}U(C_{t})$.
+We use standard mixed $L^{q,p}(Q_{T})$ space-time norms over the domain $Q_{T}=(0,T)\times\Omega$, where $\Omega\subseteq\mathbb{R}^{d}$ is the spatial domain, with $p>d+2$ and $q\geq 2$.
+We define the error norms: $E_{f}=\|f_{\pi^{*},c^{*}}-f_{\pi,c}\|_{L^{q,p}}$, $E_{h}=\|h\|_{L^{q,p}}$ (where $h$ represents the difference in policy influence on PDE coefficients as per (27)), and $E_{v}=\|\nabla V^{\pi^{*},c^{*}}-\nabla V^{\pi,c}\|_{L^{q,p}}$.
+2. Standing Assumptions.
+Our analysis relies on standard regularity conditions.
+Assumption 6 (Parabolic Regularity).
+We assume: (a) The maps from policy $(\pi,c)$ to PDE coefficients (drift $b$, diffusion $a$) are Lipschitz continuous with constant $L_{\mathrm{coef}}$.
+(b) The Hamiltonian $H(t,z,V,X,\pi,c)$ is $C^{2}$ in its arguments; its gradient $\nabla_{(\pi,c)}H$ is Lipschitz continuous in $X:=\nabla V$ with constant $L_{HX}$.
+(c) The map from costates to controls via PMP FOCs, $F:(\lambda,\nabla\lambda)\mapsto(\pi,c)$, is locally Lipschitz.
+Standard uniform parabolicity holds: $\xi^{\top}a(t,z,\pi,c)\xi\geq\underline{\nu}\|\xi\|^{2}$ for some $\underline{\nu}>0$.
+Coefficients are sufficiently smooth for PDE regularity results (e.g., VMO in space, bounded measurable in time).
+Assumption 7 (FOC Gap).
+The policy $(\pi,c)$ from Stage 1 and its costate $\lambda=\nabla V^{\pi,c}$ satisfy:
+$\|\mathcal{R}_{\mathrm{FOC}}(\pi,c;\lambda)\|_{L^{q,p}}\leq\varepsilon,$
+where $\mathcal{R}_{\mathrm{FOC}}:=\nabla_{(\pi,c)}H(t,z,V^{\pi,c},\nabla V^{\pi,c},\pi,c)$.
+3. Key Technical Results.
+Lemma 2 (FOC–reduction: coefficient–gap bound, full version).
+Let Assumptions 67 hold and assume further that
+$\exists\,\underline{\Lambda}>0:\qquad-\nabla_{(\pi,c)}^{2}H\;\succeq\;\underline{\Lambda}I_{(n+1)\times(n+1)}\quad\text{for all }(t,z,\pi,c)\text{ in the admissible set }{\mathcal{A}},$ (28)
+i.e.
+the Hamiltonian is uniformly strongly concave in the control variables.
+Let $h(t,z)$ collect the differences of every coefficient (in drift, diffusion and cross terms) of the linear operator $\mathcal{L}_{\pi,c}$ when $(\pi,c)$ is substituted for the optimal control $(\pi^{*},c^{*})$.
+Then, for all $(t,z)\in(0,T)\times\Omega$,
+$|h(t,z)|\;\leq\;L_{H}\,\bigl{|}\nabla V^{\pi,c}(t,z)-\nabla V^{*}(t,z)\bigr{|}\;+\;C_{\mathrm{FOC}}\,\bigl{|}\mathcal{R}_{\mathrm{FOC}}(t,z)\bigr{|},$ (29)
+with finite constants
+$L_{H}:=L_{\mathrm{coef}}\,\frac{L_{HX}}{\underline{\Lambda}},\qquad C_{\mathrm{FOC}}:=L_{\mathrm{coef}}\,\frac{1}{\underline{\Lambda}},$ (30)
+where $L_{\mathrm{coef}}$ and $L_{HX}$ are as defined in Assumption 6.
+(Dependence: $L_{H},C_{\mathrm{FOC}}$ depend only on $\underline{\Lambda},L_{\mathrm{coef}},L_{HX}$ and therefore ultimately on the model parameters fixed in Assumption 6.)
+Proof.
+First, we define the implicit system $F(\pi,c;V,X):=\nabla_{(\pi,c)}H(t,z,V,X,\pi,c)$, where $X:=\nabla V$.
+For every fixed state point $(t,z)$, the first–order condition is $F(\pi,c;V,X)=0$.
+The optimal pair $(\pi^{*},c^{*})$ satisfies $F(\pi^{*},c^{*};V^{*},\nabla V^{*})=0$.
+Next, by the uniform strong concavity (28), $\nabla_{(\pi,c)}F=-\nabla_{(\pi,c)}^{2}H$ is invertible on $\mathcal{A}$ and $\bigl{\|}(\nabla_{(\pi,c)}F)^{-1}\bigr{\|}\leq\underline{\Lambda}^{-1}$.
+Hence, the Implicit–Function Theorem yields a $C^{1}$-map $\Phi:(V,X)\mapsto(\pi,c)$ on a neighbourhood of $(V^{*},X^{*})$ with $\bigl{\|}\nabla\Phi(V,X)\bigr{\|}=\bigl{\|}(\nabla_{(\pi,c)}F)^{-1}\nabla_{(V,X)}F\bigr{\|}\leq\underline{\Lambda}^{-1}\,L_{HX}=:C_{H}$ for all $(V,X)$ in the neighbourhood.
+Then, we establish a mean–value bound for the control gap.
+For the non–optimal pair $(\pi,c)$, we have $(\pi,c)-(\pi^{*},c^{*})=\Phi\!\bigl{(}V^{\pi,c},\nabla V^{\pi,c}\bigr{)}-\Phi\!\bigl{(}V^{*},\nabla V^{*}\bigr{)}$.
+By the mean–value theorem in Banach spaces, $|(\pi,c)-(\pi^{*},c^{*})|\leq C_{H}\,\bigl{|}\nabla V^{\pi,c}-\nabla V^{*}\bigr{|}+\underline{\Lambda}^{-1}\,\bigl{|}F(\pi,c;V^{\pi,c},\nabla V^{\pi,c})\bigr{|}$.
+The second term equals $|\mathcal{R}_{\mathrm{FOC}}|$ by Assumption 7.
+Finally, we translate this to a coefficient gap.
+Because the mapping $(\pi,c)\mapsto(b,a)$ (PDE coefficients) is $L_{\mathrm{coef}}$-Lipschitz (Assumption 6(a)), $|h(t,z)|\;\leq\;L_{\mathrm{coef}}\,|(\pi,c)-(\pi^{*},c^{*})|$.
+Substituting the estimate from the previous paragraph gives $|h|\leq L_{\mathrm{coef}}C_{H}\,|\nabla V-\nabla V^{*}|\;+\;L_{\mathrm{coef}}\underline{\Lambda}^{-1}\,|\mathcal{R}_{\mathrm{FOC}}|$.
+The stated constants $L_{H},C_{\mathrm{FOC}}$ now follow by definition.
+∎
+Lemma 3 (Source–gap bound, full version).
+Fix $p>d+2,\;q\geq 2$ and let the standing assumptions 67 hold. Assume furthermore that there exist deterministic constants
+$0<\underline{c}<\overline{c}<\infty,\qquad 0<\underline{\lambda}<\overline{\lambda}<\infty,$ (31)
+such that for every admissible trajectory and for a.e. $t\in[0,T]$
+$\underline{c}\;\leq\;c_{t},\;c^{*}_{t}\;\leq\;\overline{c},\quad\underline{\lambda}\;\leq\;\lambda_{t},\;\lambda^{*}_{t}\;\leq\;\overline{\lambda},$
+where $\lambda_{t}:=V^{\pi,c}_{x}(t,\cdot)$ and $\lambda^{*}_{t}:=V^{*}_{x}(t,\cdot)$. Let
+$E_{f}\;:=\;\bigl{\|}\,e^{-\delta t}\bigl{(}U(c^{*})-U(c)\bigr{)}\bigr{\|}_{L^{q,p}},\qquad E_{v}:=\|\nabla V^{*}-\nabla V^{\pi,c}\|_{L^{q,p}},$
+and write $\varepsilon:=\|\mathcal{R}_{\mathrm{FOC}}\|_{L^{q,p}}$ for the first–order-condition residual norm. Then
+$E_{f}\;\leq\;C_{f}\,\varepsilon\;+\;C_{f}^{\prime}\,E_{v},$ (32)
+with explicit finite constants
+$C_{f}\;=\;\overline{c}^{-\gamma}\,L_{(U^{\prime})^{-1}},\qquad C_{f}^{\prime}\;=\;L_{(U^{\prime})^{-1}},\qquad L_{(U^{\prime})^{-1}}=\frac{\overline{\lambda}^{\frac{\gamma+1}{\gamma}}}{\gamma}\,\underline{c}^{\gamma}.$ (33)
+(All dependence on the model is through $\delta,\gamma,\underline{c},\overline{c},\underline{\lambda},\overline{\lambda}$.)
+Proof.
+Throughout the proof, we abbreviate $\lVert\cdot\rVert_{q,p}:=\lVert\cdot\rVert_{L^{q,p}}$.
+For CRRA utility $U(c)=c^{1-\gamma}/(1-\gamma)$ with $\gamma>1$, we have $U^{\prime}(c)=c^{-\gamma}$ and $(U^{\prime})^{-1}(\ell)=\ell^{-1/\gamma}$.
+By (31), $\sup_{c\in[\underline{c},\overline{c}]}\!U^{\prime}(c)=\underline{c}^{-\gamma}\;<\;\infty$, and $\sup_{\ell\in[\underline{\lambda},\overline{\lambda}]}\bigl{|}(U^{\prime})^{-1}\!{}^{\prime}(\ell)\bigr{|}=\frac{1}{\gamma}\,\underline{c}^{\gamma}\,\overline{\lambda}^{\frac{\gamma+1}{\gamma}}=:L_{(U^{\prime})^{-1}}$.
+Hence $(U^{\prime})^{-1}$ is globally Lipschitz on $[\underline{\lambda},\overline{\lambda}]$ with constant $L_{(U^{\prime})^{-1}}$.
+Next, by Taylor expansion of $U$, for every $(t,\omega)$ there exists $c^{\dagger}_{t,\omega}\in[\underline{c},\overline{c}]$ such that $U(c^{*}_{t})-U(c_{t})=U^{\prime}(c^{\dagger}_{t,\omega})\,(c^{*}_{t}-c_{t})$, and $\bigl{|}U^{\prime}(c^{\dagger}_{t,\omega})\bigr{|}\leq\underline{c}^{-\gamma}$.
+Multiplying by $e^{-\delta t}$ and taking absolute values yields $e^{-\delta t}\bigl{|}U(c^{*}_{t})-U(c_{t})\bigr{|}\;\leq\;\underline{c}^{-\gamma}\,e^{-\delta t}\,\bigl{|}\,c^{*}_{t}-c_{t}\bigr{|}$.
+Furthermore, the first-order condition for consumption gives $e^{-\delta t}U^{\prime}(c^{*}_{t})=\lambda^{*}_{t}$ and $e^{-\delta t}U^{\prime}(c_{t})=\lambda_{t}-\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)$, where $\mathcal{R}_{\mathrm{FOC}}^{(c)}$ denotes the consumption component of the FOC–residual.
+Because both arguments $\ell^{*}:=\lambda^{*}_{t}$ and $\ell:=\lambda_{t}-\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)$ lie in $[\underline{\lambda},\overline{\lambda}]$, the Lipschitz property of $(U^{\prime})^{-1}$ yields $|c^{*}_{t}-c_{t}|\;\leq\;L_{(U^{\prime})^{-1}}\,\bigl{(}|\lambda_{t}-\lambda^{*}_{t}|+|\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)|\bigr{)}$.
+Combining these results, we insert the bound for $|c^{*}_{t}-c_{t}|$ into the inequality for $e^{-\delta t}\bigl{|}U(c^{*}_{t})-U(c_{t})\bigr{|}$ to get $e^{-\delta t}\bigl{|}U(c^{*}_{t})-U(c_{t})\bigr{|}\;\leq\;\Bigl{(}\underline{c}^{-\gamma}L_{(U^{\prime})^{-1}}\Bigr{)}\bigl{(}|\lambda_{t}-\lambda^{*}_{t}|+|\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)|\bigr{)}$.
+Taking the $L^{q,p}$–norm, using the definitions of $E_{v}$ and $\varepsilon$, and noting $|\lambda-\lambda^{*}|=|V^{\pi,c}_{x}-V^{*}_{x}|\leq|\nabla V^{\pi,c}-\nabla V^{*}|$, we conclude $E_{f}\;\leq\;\underline{c}^{-\gamma}L_{(U^{\prime})^{-1}}\,\varepsilon\;+\;L_{(U^{\prime})^{-1}}\,E_{v}$.
+With $C_{f},C_{f}^{\prime}$ chosen as in the statement, (32) follows.
+∎
+Proposition C.1 (Value– and gradient–gap stability, full version).
+Fix exponents $p>d+2,\;q\geq 2$. Under Assumption 6 (coefficient regularity, uniform parabolicity) and with the notation
+$E_{f}:=\|f\|_{L^{q,p}},\qquad E_{h}:=\|h\|_{L^{q,p}},\qquad E_{v}:=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}},$
+one has
+$E_{v}\;\leq\;C_{S}\,\bigl{(}E_{f}+E_{h}\bigr{)},$ (34)
+where the stability constant $C_{S}$ depends only on $p,q,T,d,\underline{\nu},\|a^{*}\|_{L^{\infty}},\|b^{*}\|_{L^{\infty}}$ ($\underline{\nu}$: the uniform ellipticity constant; $a^{*},b^{*}$: diffusion and drift coefficients of $\mathcal{L}_{\pi^{*},c^{*}}$).
+Proof.
+First, we consider the equation for the difference $W:=V^{\pi,c}-V^{*}$.
+Recall that, for each admissible control, $V^{\pi,c}$ solves the linear backward parabolic PDE $-\partial_{t}V^{\pi,c}+{\mathcal{L}}_{\pi,c}V^{\pi,c}=f_{\pi,c}$, with $V^{\pi,c}(T,\cdot)=g$, where the terminal datum is $g(x)=\kappa U(x)$.
+Subtracting the optimal‐control equation yields $-\partial_{t}W+{\mathcal{L}}_{\!*}W=f+h\cdot\nabla V^{\pi,c}$, with $W(T,\cdot)=0$.
+Here $\mathcal{L}_{\!*}:=\mathcal{L}_{\pi^{*},c^{*}}$, $f:=f_{\pi,c}-f_{\pi^{*},c^{*}}$, and $h:=(b_{\pi,c}-b_{\!*},a_{\pi,c}-a_{\!*})$ abbreviates the vector of coefficient gaps; thus $h\in L^{q,p}$ by Lemma 2.
+Regarding the coefficient class, Assumption 6(a) implies that $(t,x)\mapsto a_{\!*}(t,x)$ is uniformly VMO in $x$ with bounded measurable $t$–dependence, and satisfies the ellipticity bound $\underline{\nu}\,|\,\xi|^{2}\leq\xi^{\!\top}a_{\!*}(t,x)\xi\leq\|a_{\!*}\|_{\infty}|\,\xi|^{2}$.
+The lower order coefficients $b_{\!*}$ are bounded.
+For the source term, define $g_{src}:=f+h\cdot\nabla V^{\pi,c}$ (using $g_{src}$ to avoid conflict with terminal data $g$).
+By Lemma 3 and Lemma 2, $f,h\in L^{q,p}$.
+Moreover, $\nabla V^{\pi,c}\in L^{q,p}$ (by maximal‐regularity for $V^{\pi,c}$), hence $g_{src}\in L^{q,p}$ and $\lVert g_{src}\rVert_{q,p}\leq E_{f}+E_{h}$.
+Then, we obtain an $L^{q,p}$–gradient estimate.
+The equation for $W$ is a linear, non–divergence form, backward parabolic PDE with VMO diffusion coefficients, bounded drift, zero terminal data, and right–hand side $g_{src}\in L^{q,p}$.
+Therefore, Dong–Kim (Dong and Kim, , 2009, Thm.
+2.1) (adapted to the backward direction by time reversal) yields $\|\nabla W\|_{L^{q,p}}\;\leq\;C_{\!DK}\,\|g_{src}\|_{L^{q,p}}$, where $C_{\!DK}=C_{\!DK}(p,q,d,T,\underline{\nu},\|a_{\!*}\|_{\infty},\|b_{\!*}\|_{\infty})$.
+Finally, inserting the bound on $g_{src}$ and setting $C_{S}:=C_{\!DK}$ leads to the conclusion (34). ∎
+Proposition C.2 (Coupled error‐estimate resolution, full version).
+Let the constants $C_{f},C_{f}^{\prime},L_{H},C_{\mathrm{FOC}},C_{S}$ be those obtained in Lemmas 3, 2 and Proposition C.1.
+Define $\rho:=C_{S}\bigl{(}C_{f}^{\prime}+L_{H}\bigr{)}$.
+Then, under Assumptions 67, the error triple $(E_{f},E_{h},E_{v})$ satisfies
+$E_{f}\;\leq\;C_{f}^{\prime\prime}\varepsilon,\quad E_{h}\;\leq\;C_{h}^{\prime\prime}\varepsilon,\quad E_{v}\;\leq\;C_{v}^{\prime\prime}\varepsilon,$ (35)
+with finite constants
+$C_{v}^{\prime\prime}=\frac{C_{S}(C_{f}+C_{\mathrm{FOC}})}{1-\rho_{\!*}},\quad C_{f}^{\prime\prime}=C_{f}+C_{f}^{\prime}C_{v}^{\prime\prime},\quad C_{h}^{\prime\prime}=C_{\mathrm{FOC}}+L_{H}C_{v}^{\prime\prime},$
+where if $\rho<1$, we may take $\rho_{\!*}:=\rho$; otherwise, if $\rho\geq 1$, pick an integer $m\geq 1$ such that $\rho_{\Delta}:=C_{S}(\Delta)\bigl{(}C_{f}^{\prime}+L_{H}\bigr{)}\;<1$ for $\Delta:=T/m$ (possible because $C_{S}(\Delta)\to 0$ as $\Delta\downarrow 0$).
+Then set $\rho_{\!*}:=\rho_{\Delta}$ and $C_{S}:=\max_{0\leq k<m}C_{S}(\Delta)$.
+(The constants depend on $p,q,T,d,\underline{\nu},\|a^{*}\|_{\infty},\|b^{*}\|_{\infty}$ only through $C_{f},C_{f}^{\prime},L_{H},C_{\mathrm{FOC}},C_{S}$.)
+Proof.
+From Lemma 3, Lemma 2, and Proposition C.1, we have the following basic three inequalities:
+$E_{f}$ $\;\leq\;C_{f}\,\varepsilon+C_{f}^{\prime}E_{v},$ (36a)
+$E_{h}$ $\;\leq\;C_{\mathrm{FOC}}\varepsilon+L_{H}E_{v},$ (36b)
+$E_{v}$ $\;\leq\;C_{S}\,(E_{f}+E_{h}).$ (36c)
+Inserting (36a) and (36b) into (36c) yields a single scalar inequality for $E_{v}$:
+$E_{v}\;\leq\;C_{S}\Bigl{[}(C_{f}+C_{\mathrm{FOC}})\varepsilon+(C_{f}^{\prime}+L_{H})E_{v}\Bigr{]}\;=\;C_{S}(C_{f}+C_{\mathrm{FOC}})\varepsilon+\rho\,E_{v}.$
+Rearranging this gives $(1-\rho)\,E_{v}\;\leq\;C_{S}(C_{f}+C_{\mathrm{FOC}})\,\varepsilon$.
+Now, we consider two cases for $\rho$.
+If $\rho<1$, the inequality immediately gives $E_{v}\leq\frac{C_{S}(C_{f}+C_{\mathrm{FOC}})}{1-\rho}\,\varepsilon$.
+Define $C_{v}^{\prime\prime}:=\frac{C_{S}(C_{f}+C_{\mathrm{FOC}})}{1-\rho}$.
+Back‐substituting this into (36a) and (36b) yields (35) with the constants stated for the $\rho<1$ branch, setting $\rho_{\!*}:=\rho$.
+If $\rho\geq 1$, we use a time–slicing argument.
+Fix $m\in\mathbb{N}$ and set $\Delta:=T/m$.
+Let $E_{v}^{(k)}$ denote the $L^{q,p}$–norm of $\nabla V^{\pi,c}-\nabla V^{*}$ on the slab $Q_{k}:=(T-k\Delta,\,T-(k-1)\Delta)\times\Omega$.
+Because the coefficients are effectively frozen on each slab, Proposition C.1 holds with a short‐horizon constant $C_{S}(\Delta)$, which satisfies $C_{S}(\Delta)\to 0$ as $\Delta\downarrow 0$ (as the Dong–Kim estimate is local in time).
+Repeating the derivation for the scalar inequality for $E_{v}$ on $Q_{k}$ yields $E_{v}^{(k)}\;\leq\;\rho_{\Delta}\,E_{v}^{(k-1)}+C_{S}(\Delta)\bigl{(}C_{f}+C_{\mathrm{FOC}}\bigr{)}\varepsilon$, where $\rho_{\Delta}:=C_{S}(\Delta)\bigl{(}C_{f}^{\prime}+L_{H}\bigr{)}$.
+We pick $m$ such that $\rho_{\Delta}<1$.
+Iterating for $k=1,\dots,m$ and noting $E_{v}^{(0)}=0$ (due to the terminal condition $W(T,\cdot)=0$), we get $E_{v}^{(m)}\;\leq\;\frac{C_{S}(\Delta)(C_{f}+C_{\mathrm{FOC}})}{1-\rho_{\Delta}}\varepsilon$.
+With $C_{v}^{\prime\prime}:=\dfrac{C_{S}(\Delta)(C_{f}+C_{\mathrm{FOC}})}{1-\rho_{\Delta}}$, $\rho_{\!*}:=\rho_{\Delta}$, and $C_{S}$ redefined as $\max_{0\leq k<m}C_{S}(\Delta)$ for the overall bound, the results (35) follow similarly to the $\rho<1$ case.
+In both scenarios, all quantities $C_{f},C_{f}^{\prime},L_{H},C_{\mathrm{FOC}},C_{S}(\Delta)$ depend only on the data in Assumption 6.
+In particular, $C_{S}(\Delta)\leq C_{S}(T)$, so $C_{v}^{\prime\prime},C_{f}^{\prime\prime},C_{h}^{\prime\prime}$ are finite.
+∎
+Theorem 4 (Policy Gap Bound).
+Under Assumptions 6 and 7, let $(\pi,c)$ be the policy from Stage 1 satisfying the FOC-gap condition $\|\mathcal{R}_{\mathrm{FOC}}(\pi,c;\lambda)\|_{L^{q,p}}\leq\varepsilon$, where $\lambda=\nabla V^{\pi,c}$.
+Let $\pi^{*}$ be the true optimal policy, and let $\widehat{\pi}=F(\widehat{\lambda})$ be the policy deployed in Stage 2 using BPTT estimates $\widehat{\lambda}$, where $F$ is the explicit mapping from costates $(\lambda,\nabla\lambda)$ to controls derived from the PMP FOCs.
+Assume $F$ is locally Lipschitz continuous with constant $C_{F}$ (guaranteed under Assumption 6(c)).
+If the BPTT estimation error relative to the Stage 1 costate satisfies $\|\widehat{\lambda}-\lambda\|_{L^{q,p}}\leq\delta_{\mathrm{BPTT}}$, then the policy gap is bounded by:
+$\boxed{\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\mathrm{tot}}(\delta_{\mathrm{BPTT}}+\varepsilon)}$
+Here, $C_{\mathrm{tot}}$ depends on $C_{F}$ and the constant $C_{v}^{\prime\prime}$ from Proposition C.2.
+Proof (detailed).
+Let $(\lambda_{S1},Z_{S1})$ denote the adapted solution (specifically, $\lambda_{S1}=\nabla V^{\pi,c}$) associated with the Stage 1 policy $(\pi,c)$.
+The BPTT process yields estimates $\widehat{\lambda}$ (and implicitly $\widehat{Z}$) for the costate and its related processes.
+The first–order map $F:(\lambda,\nabla\lambda)\mapsto(\pi,c)$ is $C^{1}$ on a neighbourhood $\mathcal{N}$ of $(\lambda^{*},\nabla\lambda^{*})$ with Lipschitz constant $\|F\|_{\mathrm{Lip}}\leq C_{F}:=\underline{\Lambda}^{-1}L_{HX}$ (derived similarly to reasoning in Lemma 2, under Assumption 6(c)).
+The policy deployed in Stage 2 is $\widehat{\pi}=F(\widehat{\lambda})$, and the true optimal policy is $\pi^{*}=F(\nabla V^{*})$.
+We want to bound $\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}$.
+Using the Lipschitz continuity of $F$:
+$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}$ $=\|F(\widehat{\lambda})-F(\nabla V^{*})\|_{L^{q,p}}$
+$\leq C_{F}\|\widehat{\lambda}-\nabla V^{*}\|_{L^{q,p}}.$
+By the triangle inequality,
+$\|\widehat{\lambda}-\nabla V^{*}\|_{L^{q,p}}\leq\|\widehat{\lambda}-\lambda_{S1}\|_{L^{q,p}}+\|\lambda_{S1}-\nabla V^{*}\|_{L^{q,p}}.$
+We are given that the BPTT estimation error $\|\widehat{\lambda}-\lambda_{S1}\|_{L^{q,p}}\leq\delta_{\mathrm{BPTT}}$.
+The term $\|\lambda_{S1}-\nabla V^{*}\|_{L^{q,p}}$ is precisely $E_{v}=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}}$.
+From Proposition C.2, we have $E_{v}\leq C_{v}^{\prime\prime}\varepsilon$.
+Substituting these into the inequality for $\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}$:
+$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\leq C_{F}\left(\delta_{\mathrm{BPTT}}+E_{v}\right)\leq C_{F}\left(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon\right).$
+To express this in the form $C_{\mathrm{tot}}(\delta_{\mathrm{BPTT}}+\varepsilon)$, we can set $C_{\mathrm{tot}}=C_{F}\max(1,C_{v}^{\prime\prime})$.
+If $C_{v}^{\prime\prime}\geq 1$, then $C_{F}(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon)\leq C_{F}C_{v}^{\prime\prime}(\delta_{\mathrm{BPTT}}+\varepsilon)$.
+If $C_{v}^{\prime\prime}<1$, then $C_{F}(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon)\leq C_{F}(\delta_{\mathrm{BPTT}}+\varepsilon)$.
+Thus, $C_{\mathrm{tot}}=C_{F}\max(1,C_{v}^{\prime\prime})$ provides a valid constant for the boxed expression.
+Alternatively, one can use $C_{\mathrm{tot}}=C_{F}(1+C_{v}^{\prime\prime})$ if one prefers a sum form, as $C_{F}(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon)\leq C_{F}(1+C_{v}^{\prime\prime})(\delta_{\mathrm{BPTT}}+\varepsilon)$ assuming $C_{v}^{\prime\prime}\geq 0$.
+The choice $C_{\mathrm{tot}}=C_{F}\max(1,C_{v}^{\prime\prime})$ is tighter.
+The key result is the linear dependence on $\delta_{\mathrm{BPTT}}$ and $\varepsilon$.
+∎
+4. Summary and Constant Dependencies.
+We have demonstrated a clear logical path from a small FOC-gap $\varepsilon$ to a small policy gap $O(\varepsilon+\delta_{\mathrm{BPTT}})$, conditional on standard parabolic regularity.
+The argument relies on establishing key bounds (Lemmas 2, 3), utilizing PDE stability (Proposition C.1), resolving coupled estimates (Proposition C.2), and concluding with the main policy gap theorem (Theorem 4).
+It is important to reiterate that all constants ($L_{H},C_{\mathrm{FOC}},C_{f},C_{f}^{\prime},C_{S},$ $C_{f}^{\prime\prime},C_{h}^{\prime\prime},C_{v}^{\prime\prime},C_{F},C_{\mathrm{tot}}$) depend quantitatively on model parameters (horizon $T$, utility $U$, market coefficients $\bm{\mu},\bm{\sigma},\dots$), PDE regularity parameters ($p,q,\Omega,\underline{\nu}$), and the costate floor $\underline{\lambda}$.
+Rigorous derivation requires tracking these dependencies.
+The provided proofs for lemmas are sketches; full proofs require careful application of MVT, IFT, Taylor expansions, and meticulous handling of bounds.
+The time-slicing argument mentioned requires careful justification to control constant growth.
+Subject to these standard technical details, the framework presented provides a sound mathematical basis for the Projected PG-DPO method’s effectiveness under the stated regularity conditions.
+Appendix D Detailed Experimental Setup
 This appendix provides the detailed setup for the numerical experiments presented in Section 4.
-C.1 Parameter Values
+D.1 Parameter Values
 The experiments consider the objective of maximizing the expected utility of terminal wealth $X_{T}$, $J=\mathbb{E}[U(X_{T})]$, using a CRRA utility function $U(x)=x^{1-\gamma}/(1-\gamma)$ with a relative risk aversion coefficient $\gamma=2.0$.
@@
 The specific values generated by the seed are used consistently across experiments for given $n$ and $k$.
-C.2 PG-DPO Implementation Details
+D.2 PG-DPO Implementation Details
 We implement the Baseline PG-DPO algorithm using PyTorch.
@@
     • Wealth constraint: Simulated wealth $W_{t}$ clamped at lower bound $0.1$.
-  • Costate Estimation (for Two-Stage Evaluation): During periodic evaluation, costate $\lambda$ and derivatives $\partial_{x}\lambda,\partial_{\mathbf{Y}}\lambda$ are computed via BPTT (‘torch.
-autograd.grad’) applied to Monte Carlo estimates of the expected terminal utility starting from evaluation points $(t_{k},W_{k},\mathbf{Y}_{k})$.
-C.3 Benchmark Solution and Evaluation
+  • Costate Estimation (for Two-Stage Evaluation): During periodic evaluation, costate $\lambda$ and derivatives $\partial_{x}\lambda,\partial_{\mathbf{Y}}\lambda$ are computed via BPTT (‘torch.autograd.grad’) applied to Monte Carlo estimates of the expected terminal utility starting from evaluation points $(t_{k},W_{k},\mathbf{Y}_{k})$.
+D.3 Benchmark Solution and Evaluation
 As mentioned, the chosen multi-factor OU model allows for a semi-analytical benchmark solution.
@@
 All computations are performed using PyTorch on an NVIDIA A6000 GPU.
+D.4 Deep BSDE Implementation Details
+The Deep BSDE method, serving as a benchmark, was implemented based on the standard framework by Han et al., (2018) and Weinan, (2017).
+This approach reformulates the problem’s PDE as a backward stochastic differential equation (BSDE) and approximates its solution components using neural networks.
+Our PyTorch implementation addresses the multi-asset, multi-factor portfolio optimization problem from Section 2.3.2.
+  • Network Architectures: Two main feedforward neural networks with Tanh activations are utilized:
+    • Y0Net: Approximates the value function at the initial time of a path, $Y_{t_{0}}\approx V(t_{0},X_{t_{0}},\mathbf{Y}_{t_{0}})$.
+It takes the initial state $(t_{0},\log X_{t_{0}},\mathbf{Y}_{t_{0}})$ (total $2+k$ inputs) as input and produces a scalar output (with a final negative softplus transformation).
+It typically uses 2 hidden layers of 64 units.
+    • ZNet: Approximates $\bm{Z}_{t}$, which relates to the scaled gradient of the value function concerning the underlying Brownian motions, at each time step $t$.
+It takes the current state $(t,\log X_{t},\mathbf{Y}_{t})$ as input and outputs an $(n+k)$-dimensional vector, representing $\bm{Z}^{X}_{t}\in\mathbb{R}^{n}$ and $\bm{Z}^{Y}_{t}\in\mathbb{R}^{k}$.
+It typically uses 3 hidden layers of 128 units.
+  • Policy Derivation: The optimal portfolio $\bm{\pi}_{t}$ is dynamically constructed at each step.
+This involves using the network outputs $Y_{t}$ (obtained by evolving $Y_{t_{0}}$ via the BSDE discretization using ZNet) and $\bm{Z}_{t}$ (from ZNet) to approximate the first-order conditions of the HJB equation.
+This structure allows the policy to learn both myopic demand (from current factor states $\mathbf{Y}_{t}$) and intertemporal hedging demand (where terms involving $\bm{Z}^{Y}_{t}/Y_{t}$ approximate the necessary factor sensitivities, akin to the analytical solution’s form).
+  • Loss Function: The Y0Net predicts $Y_{t_{0}}$.
+This value, along with ZNet’s outputs for $\bm{Z}_{t}$, is used to discretize the BSDE forward in time from $t_{0}$ to the terminal time $T$.
+The training objective is to minimize the mean squared error between the resulting $Y_{T}$ and the specified terminal utility $U(X_{T})=X_{T}^{1-\gamma}/(1-\gamma)$.
+  • Training Details: The networks are trained for $10,000$ epochs using the Adam optimizer (initial learning rate $3\times 10^{-5}$, with a multi-step decay scheduler) and a batch size of $1,024$.
+Each BSDE path is discretized into $N_{\text{BSDE}}=80$ time steps.
+Gradient clipping (norm 2.0) and bounding of portfolio weights are applied to enhance training stability.
+  • Initial State Sampling: For each training path, the initial time $t_{0}$ is sampled uniformly from the discrete time steps over $[0,T-\Delta t]$.
+The initial wealth $X_{0}$ is drawn uniformly from $[0.1,3.0]$, and initial factor states $\mathbf{Y}_{0}$ are sampled uniformly from a hyperrectangle representing their typical dynamic range (e.g., based on ±$\pm 3$ stationary standard deviations around their long-term means).
+The implementation uses PyTorch and is executed on the same NVIDIA A6000 GPU hardware as the PG-DPO experiments to ensure a fair comparison.

### v4

@@
-Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio
+Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio Choice
 Abstract
@@
 To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework.
+Our framework accurately captures both myopic demand and complex intertemporal hedging demands, a feat often elusive for other methods in high-dimensional settings.
+P-PGDPO delivers near-optimal policies, offering a practical and powerful alternative for a broad class of high-dimensional continuous-time control problems.
 PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to directly inform neural network-based policy learning.
@@
 This enables a breakthrough in scalability: numerical experiments demonstrate that P-PGDPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables).
-Critically, the framework accurately captures complex intertemporal hedging demands, a feat often elusive for other methods in high-dimensional settings.
-P-PGDPO delivers near-optimal policies, offering a practical and powerful alternative for a broad class of high-dimensional continuous-time control problems.
 1 Introduction
-High-dimensional optimal control problems often fall prey to the “curse of dimensionality” (Bellman, , 1966), whereby the computational cost of classical dynamic programming (DP) approaches grows exponentially with the number of state variables.
-In a continuous-time finance setting, this difficulty becomes starkly evident in multi-asset Merton problems (Merton, , 1973), where an investor dynamically allocates wealth among numerous risky assets whose drift and volatility may depend on exogenous state variables such as macroeconomic indicators or stochastic volatility factors.
-While DP or Hamilton–Jacobi–Bellman (HJB) PDE solvers provide a rigorous framework for low-dimensional cases (Fleming and Soner, , 2006; Bertsekas and Tsitsiklis, , 1995), they become intractable as the number of assets or state dimensions increases.
-This difficulty is pervasive, impacting not only continuous-time settings (which are often addressed via discretization for numerical solution) but also analogous discrete-time models where similar state-space explosion issues arise.
-Consequently, prior influential studies in these domains are typically limited to analyzing problems involving up to approximately 3-5 state variables (Campbell and Viceira, , 1999; Lynch, , 2001; Balduzzi and Lynch, , 1999) or specific frameworks handling around 6 state variables (Garlappi and Skoulakis, , 2010), often necessitating simulation-based approaches for tractability (Brandt et al., , 2005).
-A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong and Zhou, , 2012).
-Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint).
-Notably, adding one more state dimension does not produce an exponential explosion in computational requirements, in contrast to DP (Bellman, , 1966; Campbell et al., , 2003).
-This makes Pontryagin-based methods appealing for large-scale portfolio optimization problems involving many assets and exogenous state processes (Karatzas et al., , 1987).
-Recent advances in deep learning further enhance the scalability of methods for solving high-dimensional stochastic control problems (e.g., Han et al., , 2018; Raissi et al., , 2019, for general PDE-based approaches).
-Concurrently, deep learning techniques are being applied directly to financial dynamic programming and stochastic control.
-Some approaches tackle infinite horizon problems via the HJB equation (Duarte et al., , 2024), or solve the HJB equation for utility maximisation under constraints using formulations based on second-order backward stochastic differential equations (2BSDEs) (Davey and Zheng, , 2022).
-Others aim to directly solve associated nonlinear PDEs using methods like deep branching algorithms (Nguwi et al., , 2024).
-While these BSDE-based methods and direct PDE solvers are powerful, applications such as the Merton HJB have revealed limitations in accuracy or efficiency for certain variants (Nguwi et al., , 2024).
-Moreover, some works effectively leverage the Stochastic Maximum Principle (SMP) and adjoint BSDEs, often through duality methods, to handle even non-Markovian problems in finance (Davey and Zheng, , 2022).
-Alongside these PDE-centric or BSDE-centric deep learning approaches, another line of research, more akin to reinforcement learning, focuses on directly parameterizing the control policy with neural networks and optimizing the expected objective function through simulation and gradient ascent (e.g., Zhang and Zhou, , 2019; Kopeliovich and Pokojovy, , 2024, for portfolio problems with transaction costs or under Heston dynamics).
-Our PMP-based PG-DPO framework builds upon similar ideas of direct policy optimization but is distinctively guided by the continuous-time optimality conditions of PMP.
-It utilizes BPTT-derived costate estimates for finite-horizon problems, aiming for enhanced scalability and a novel way to enforce optimality conditions for problems with sufficient regularity.
-One notable strategy related to the use of adjoint methods in optimal control, and central to our work, is to directly parameterize the policy (i.e., consumption and investment controls) with neural networks, then apply backpropagation-through-time (BPTT) to enforce Pontryagin’s optimality conditions.
-This strategy, which we term the baseline “Pontryagin-Guided Direct Policy Optimization (PG-DPO),” circumvents the need to approximate value functions or PDE solutions.
-Instead, it uses simulation-based gradients that automatically incorporate the adjoint processes, drastically reducing computational overhead in high-dimensional settings (Kushner and Yin, , 2003; Borkar and Borkar, , 2008).
-The purpose of this paper is to develop and implement this PG-DPO framework for the general multi-asset, state-dependent continuous-time portfolio problem.
-We consider an investor who must decide on a continuous consumption rate and an allocation to $n$ risky assets with state-dependent drift and volatility.
-By parameterizing the policy with neural networks, we can capture potentially complex relationships dependent on both wealth and the exogenous state.
-We show how a single forward pass in simulation, coupled with a backward pass via automatic differentiation (BPTT), yields unbiased estimates of the costate processes that PMP requires, guiding the policy towards optimality without the drawbacks of DP.
-This approach fundamentally relies on the existence and accurate estimation of the costate processes dictated by PMP, applicable whenever the control problem satisfies appropriate regularity conditions allowing for a well-defined Hamiltonian system.
-Furthermore, we introduce the “Projected PG-DPO” (P-PGDPO) variant, which significantly enhances efficiency and accuracy.
-This variant exploits the observation that BPTT-derived costate estimates often stabilize much faster than the full policy network.
-Once the costates are sufficiently stable (after a short warm-up phase), one can bypass the potentially suboptimal neural policy outputs.
-Instead, these costate estimates are directly substituted into the PMP first-order conditions (FOCs), effectively projecting the learned costate information onto the manifold of controls satisfying PMP’s necessary conditions.
-This recovers a near-optimal control analytically at every time-state point with minimal additional training, significantly reducing computational burden.
-While other deep learning methods such as Deep BSDE are also applied to high-dimensional problems, their ability to accurately capture complex, state-dependent optimal policy structures, particularly the crucial intertemporal hedging components, has been limited in challenging financial applications.
-Our numerical experiments (Section 4) demonstrate that P-PGDPO not only scales to problems with dozens of assets and state variables (e.g., n=50, k=10 in our experiments) but, critically, it is capable of precisely disentangling and reproducing these vital hedging demands—a feat not typically achieved by existing approaches in such high-dimensional settings.
-To the best of our knowledge, this explicit and accurate recovery of complex hedging terms in a high-dimensional, PMP-guided direct policy optimization framework is a novel contribution.
-This superior performance in both scalability and fidelity to the true optimal policy, especially its nuanced components, highlights the transformative potential of leveraging the PMP structure via BPTT-derived costate estimates for tackling a broad class of complex, high-dimensional control problems where amenability to costate estimation under suitable regularity conditions exists.
+The theory of dynamic consumption and portfolio choice has been a cornerstone of financial economics, providing fundamental insights into asset pricing and long-term investment behavior (Merton, , 1973; Campbell and Viceira, , 1999, 2001).
+Despite its theoretical elegance and contributions to our understanding, the theory faces a major obstacle in practical implementation: the lack of scalable methods for solving high-dimensional portfolio choice problems involving many assets and state variables.
+The Bellman equation associated with such problems is subject to the curse of dimensionality, and obtaining solutions in realistic settings is often infeasible.
+Existing solution methods are typically complex, ad hoc, and do not scale well with dimensionality.
+This is one reason why Markowitz’s static mean-variance framework remains dominant in practice, despite its limitations and the greater realism offered by dynamic models.
+[^1] Cochrane wrote “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice.
+” (Cochrane, , 2022, p.4)
+In this paper, we make a technological contribution by proposing a solution approach that leverages machine learning with neural networks to solve high-dimensional dynamic portfolio choice problems, where both the number of assets $n$ and the number of state variables $k$ are large.
+There are two primary challenges inherent in such problems.
+The first is the well-known curse of dimensionality: computational methods often fail to scale efficiently as the dimensionality of the problem increases in $n$ and $k$.
+The second, and more subtle, challenge arises from the specific structure of dynamic portfolio choice itself.
+Unlike other applications of reinforcement learning—such as Atari, Go, or autonomous driving—where achieving a near-optimal policy is often sufficient for strong empirical performance, portfolio optimization requires much greater precision.
+A well-established theoretical benchmark exists, dating back to the seminal work of Merton, (1973), which demonstrates that the optimal portfolio can be decomposed into two critical components: the myopic demand and the intertemporal hedging demand.
+The myopic component represents the demand for a mean-variance efficient portfolio given current investment opportunities, while the intertemporal component hedges against expected future changes in those opportunities driven by evolving economic conditions.
+Accurately identifying both components is essential for evaluating investor behavior and designing long-term investment strategies—yet it remains a formidable challenge, particularly in high-dimensional settings.
+Our proposed approach is specifically designed to address both challenges: overcoming the curse of dimensionality and enabling the accurate decomposition of optimal portfolios into their myopic and intertemporal components.
+We adopt a continuous-time framework in which asset return means and covariances are driven by a $k$-dimensional state vector, with the state variables evolving as Itô diffusions, following Merton, (1973).
+This modeling approach benefits from the rich toolkit of stochastic calculus—tools generally unavailable in discrete-time formulations.
+[^2] See Duarte et al., (2024) for a discussion of the advantages of continuous-time models over discrete-time models.
+In the continuous-time setting, (quasi-)analytic solutions are available under affine dynamics, where the Hamilton–Jacobi–Bellman (HJB) equation reduces to a system of ordinary differential equations (Kim and Omberg, , 1996; Liu, , 2007; Buraschi et al., , 2010).
+However, for general non-affine dynamics, no universal solution method exists.
+To address this limitation, we introduce a machine learning–based approach that applies to arbitrary market dynamics.
+Neural networks, known for their ability to approximate complex high-dimensional functions, are used to parameterize the policy directly, enabling us to capture intricate dependencies on both wealth and the exogenous state variables—even when the dimension $k$ is large.
+While we focus on Brownian motion as the source of return risk, the methods can be extended to accommodate jumps.
+Furthermore, if a discrete-time model is viewed as an approximation of a continuous-time model, our approach can be adapted accordingly.
+Our approach leverages Pontryagin’s Maximum Principle (PMP) as its theoretical foundation, in contrast to the more commonly used dynamic programming (DP) framework based on the Hamilton–Jacobi–Bellman (HJB) equation.
+Building on PMP, we propose two methods: Pontryagin-Guided Direct Policy Optimization (PGDPO) and its variant, Projected PGDPO (P-PGDPO).
+PMP characterizes optimal controls through a coupled system of forward and backward differential equations.
+The forward equation governs the evolution of wealth and state variables driving asset returns, which we implement via Monte Carlo simulation.
+This allows for direct computation of the expected utility and avoids the need to estimate the value function across future periods—a core but computationally intensive requirement of DP-based approaches.
+The backward equation governs the dynamics of the costates (or adjoint variables), which correspond to the first-order partial derivatives of the value function with respect to the state variables.
+Modern machine learning frameworks such as PyTorch allow us to compute these costates and their derivatives, the second-order partial derivatives of the value function, automatically through backpropagation through time (BPTT) and automatic differentiation.
+We demonstrate that a single forward simulation pass, coupled with BPTT, yields unbiased estimates of the costate processes required by PMP.
+As a result, our approach avoids the need for explicit solution of the backward equation or approximation of the value function, thereby sidestepping major bottlenecks of DP.
+Costates—as well as the value function—emerge as natural byproducts of the learning process.
+In PGDPO, the control policy—comprising consumption and portfolio allocations—is parameterized by neural networks.
+The expected utility, defined as the discounted sum of utility from consumption and terminal wealth, is then directly maximized using stochastic gradient ascent, where gradients are computed using BPTT.
+A key computational advantage of our approach lies in its complete avoidance of HJB residual minimization, where the residual is defined as the cumulative squared error of the HJB equation.
+Computing the HJB residual typically requires evaluating first- and second-order partial derivatives of the value function at each time–state grid point and aggregating the errors over time, incurring high computational cost.
+Duarte et al., (2024) mitigate this by replacing full gradient computation with the second derivative of a carefully constructed univariate function.
+In contrast, our method bypasses HJB residual computation entirely, resulting in significant computational savings.
+Conceptually, our approach shares similarities with recent methods that parameterize control policies via neural networks and directly optimize the expected objective function through simulation and gradient ascent, thereby circumventing the need to solve the HJB equation (Zhang and Zhou, , 2019; Kopeliovich and Pokojovy, , 2024, e.g., for portfolio problems with transaction costs or Heston dynamics).
+However, our PMP-based PGDPO framework distinguishes itself by being explicitly guided by continuous-time optimality conditions derived from PMP.
+Moreover, unlike standard reinforcement learning methods such as Proximal Policy Optimization (PPO) or Trust Region Policy Optimization (TRPO), which rely on sampled estimates of policy gradients, our method computes exact gradients via BPTT.
+This yields lower-variance updates and enhances computational efficiency, particularly in high-dimensional settings (Kushner and Yin, , 2003; Borkar and Borkar, , 2008).
+The second method, P-PGDPO, builds on PGDPO by substituting the estimated derivatives of the value function directly into the structural first-order conditions from PMP.
+These conditions take the form derived originally by Merton, (1973), explicitly decomposing the optimal portfolio into myopic and intertemporal hedging demands.
+We demonstrate the numerical accuracy of our methods by comparing the solutions they produce to a quasi-analytic benchmark under affine dynamics, where the true solution can be obtained by solving a system of ordinary differential equations.
+The two components of the optimal portfolio—myopic and intertemporal hedging demands—are not merely of theoretical interest; their accurate identification requires high-fidelity solutions.
+As shown by Constantinides, (1986), transaction costs generate only second-order utility losses, allowing portfolios to deviate substantially from the frictionless optimum while still appearing nearly optimal in terms of utility.
+Similarly, Cochrane, (1988) demonstrates that first-order deviations from the optimal policy can result in only second-order losses in utility.
+This disconnect between utility value and policy structure poses a major challenge for traditional reinforcement learning algorithms.
+These methods often converge to policies that are near-optimal in utility but structurally far from the true optimum—particularly in high-dimensional settings.
+In such cases, neural networks must search for optimal policies over an $n$-dimensional control space, where $n$ denotes the number of risky assets.
+The resulting exponential growth in complexity renders standard policy optimization techniques computationally infeasible as $n$ increases.
+Our numerical experiments confirm that increasing the number of risky assets $n$ presents a significantly greater challenge than increasing the number of state variables $k$ in high-dimensional settings.
+The P-PGDPO method is motivated by the empirical observation that value function derivatives—i.e., costates—are typically easier to learn than the policy functions themselves.
+In practice, these partial derivatives tend to stabilize much earlier during training.
+Once the costate estimates have stabilized—usually after a short warm-up phase—the method can bypass the potentially suboptimal outputs of the policy networks.
+Instead, the costates and their derivatives are directly substituted into the first-order conditions (FOCs) from PMP, whose structural form closely mirrors the conditions originally derived by Merton, (1973).
+This effectively projects the learned costate information onto the manifold of controls that satisfy PMP’s necessary conditions.
+The resulting control policy adheres closely to the structural form of the true optimal solution dictated by PMP.
+It is recovered quasi-analytically at each time–state point, requiring minimal additional training and incurring substantially reduced computational cost.
+Our PMP-based PGDPO framework thus leverages BPTT-derived estimates of costates and their derivatives, enhancing scalability while providing a principled mechanism for enforcing first-order optimality conditions.
+We provide a theoretical justification for the P-PGDPO method in Theorem 1.
+When the objective function surface is relatively flat near the optimum, the first-order condition (FOC) violation for the policy obtained in the warm-up stage tends to be small.
+The theorem states that if this FOC deviation is small and the costate estimates obtained via BPTT are sufficiently accurate, then the control policy recovered in the second stage is close to the true optimal policy.
+We demonstrate the accuracy of our numerical methods by solving a problem in which the state process follows a multi-factor Ornstein–Uhlenbeck (OU) process—a special case of affine dynamics—for which quasi-analytic benchmark solutions can be obtained by solving a system of ordinary differential equations (see Section 2.3.2).
+We also compare our methods with the Deep Backward Stochastic Differential Equation (Deep BSDE) approach, which is representative of existing techniques for solving high-dimensional partial differential equations, particularly the Hamilton–Jacobi–Bellman (HJB) equations arising in DP.
+In addition, we apply our second method, P-PGDPO, to a non-affine setting by adding non-linear terms to the expected returns of risky assets, thereby demonstrating its applicability to a broader class of dynamics.
+Our numerical experiments (Section 4) show that our first method, PGDPO, produces solutions comparable to those of the Deep BSDE method in terms of mean-squared errors.
+In contrast, the second method, P-PGDPO, yields significantly smaller errors.
+This improvement arises because PGDPO inherits the common limitation of policy optimization algorithms—difficulty in discovering the true optimum—whereas P-PGDPO overcomes this issue by directly exploiting the structural form of the first-order conditions from Pontryagin’s Maximum Principle (PMP).
+P-PGDPO not only scales to problems involving dozens of assets and state variables (e.g., $n=50$, $k=10$), but also accurately disentangles and reconstructs complex intertemporal hedging demands.
+Furthermore, our results demonstrate that the method generates stable solutions in non-affine settings and converges smoothly to the benchmark affine solution as the coefficients of the nonlinear terms approach zero.
+By contrast, the Deep BSDE approach, while successful in other high-dimensional applications such as option pricing and hedging, fails to capture the full structure of the optimal policy in our setting—particularly the intertemporal hedging components.
+The method’s superior performance—both in scalability and in fidelity to the full structure of the optimal policy—highlights the transformative potential of leveraging PMP through BPTT-derived costate estimates for a broad class of high-dimensional control problems that admit stable costate representations.
+There is a substantial literature on dynamic portfolio choice, aimed at obtaining concrete solutions either analytically or numerically.
+With the exception of the continuous-time models with affine dynamics discussed above, most existing contributions are formulated in discrete time.
+For example, Campbell and Viceira, (1999) and Campbell and Viceira, (2001) employ log-linear approximations to the dynamic budget constraint to obtain approximate closed-form solutions.
+Balduzzi and Lynch, (1999), Lynch and Balduzzi, (2000), and Lynch, (2001) apply discrete-space approximations to solve portfolio choice problems with transaction costs.
+Garlappi and Skoulakis, (2010) develop a DP approach based on state variable decomposition.
+Brandt et al., (2005) propose a Monte Carlo simulation method for a discrete-time portfollio choice model that projects simulated values of relevant functions—such as the marginal value of wealth—onto a pre-specified set of basis functions.
+In this sense, their method shares similarities with the forward simulation and auto-differentiation steps in our PG-DPO framework.
+However, their approach does not explicitly leverage PMP or utilize the representational power of neural networks.
+As a result, it is not easily scalable and lacks the flexibility for generalization.
+In contrast, our method is PMP-guided, scalable, and extensible to broader classes of control problems.
+In continuous time, there has been significant technological advancement in solving partial differential equations, which can be applied to the HJB equation arising from DP.
+Notable methods include Deep BSDE, Deep Galerkin, Physics-Informed Neural Networks (PINNs), Deep Splitting, and Deep Branching, among others (Han et al., , 2018; Sirignano and Spiliopoulos, , 2018; Raissi et al., , 2019; Beck et al., , 2021; Nguwi et al., , 2024).
+In parallel, substantial progress has been made in solving optimal control problems through the development of policy evaluation, policy gradient, $q$-learning, and PPO/TRPO methods in stochastic environments (Wang et al., , 2020; Jia and Zhou, 2022a, ; Jia and Zhou, 2022b, ; Jia and Zhou, , 2023; Zhao et al., , 2023).
+Wang and Zhou, (2020); Dai et al., (2023); Huang et al., (2024) apply these techniques to dynamic mean-variance analysis, while Dai et al., (2025) use them to solve a Merton-type portfolio optimization problem.
+These approaches involve learning the asset return process from observed return data.
+Geng et al., (2023) propose the Factor Learning Portfolio Optimization (FaLPO) method for solving a Merton-type portfolio optimization problem with stochastic factors.
+Their approach combines policy learning with model calibration.
+By contrast, we assume that the data-generating process is known and do not address the problem of learning from data.
+Instead, our focus is on a more general consumption and portfolio choice problem involving many risky assets and multiple state variables, whereas their analysis is limited to a small number of risky assets and state variables.
+Davey and Zheng, (2022) develop a BSDE-based method for portfolio optimization under portfolio constraints that is also based on the maximum principle, similar in spirit to our work.
+However, their method requires minimizing the terminal value error (as in Deep BSDE), maximizing the Hamiltonian at each state and time, and solving a BSDE system.
+In contrast, our method directly maximizes expected utility without penalizing approximation errors and without solving the backward equation explicitly, resulting in a significantly simpler and more efficient approach.
+More recently, Duarte et al., (2024) propose a machine learning approach for infinite-horizon continuous-time problems using DP, while Cheridito et al., (2025) develop a method that combines generalized policy iteration with PINNs to solve finite-horizon continuous-time DP problems.
+By contrast, our method is based on PMP, rather than DP.
+The aforementioned DP-based methods avoid the explicit evaluation of first- and second-order partial derivatives of the value function by instead computing the second derivative of a univariate function when assessing the HJB residual.
+In contrast, our approach computes these partial derivatives directly—via automatic differentiation of costate variables—providing crucial information for accurately capturing intertemporal hedging components.
+This ability to compute the value function’s derivatives with high accuracy enables precise estimation of both the intertemporal hedging terms and the myopic component of the optimal portfolio.
+[^3] Cochrane emphasizes the critical role of partial derivatives in determining the optimal portfolio for long-term investors, particularly when discussing the implementation challenges of Merton’s dynamic portfolio theory—a point we partially quoted earlier: “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice… What are the partial derivatives of the investor’s value function that guide state-variable investments?
+That is harder still…” (Cochrane, , 2022, pp.
+4–5).
+To our knowledge, the two aforementioned papers do not demonstrate accurate recovery of intertemporal hedging terms in settings with both large numbers of assets and state variables.
+To our knowledge, the existing literature has demonstrated the ability to solve problems involving at most a dozen risky assets and a dozen state variables—with two notable exceptions.
+Davey and Zheng, (2022) consider a setting with 20 assets, but under a constant investment opportunity set, i.e., without state variables influencing asset returns.
+[^4] Our companion paper demonstrates that our method is applicable to constrained portfolio optimization problems involving up to 1000 assets (Huh et al., , 2025).
+Cheridito et al., (2025) address a case with 25 assets, though without providing evidence regarding the accuracy of the resulting solution.
+In contrast, our method substantially improves scalability, handling much larger values of $n$ and comparably high values of $k$, while preserving accuracy even in high-dimensional settings.
 The main contributions of this paper can be summarized as follows:
@@
 Finally, Section 5 concludes and discusses potential future directions.
-2 Multi-Asset Continuous-Time Portfolio Problem with State Variables
-In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1973; Kim and Omberg, , 1996; Liu, , 2007), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously.
-Each risky asset features its own drift and volatility, leading to an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}$.
-Furthermore, an exogenous state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ influences market parameters such as returns and interest rates, adding additional complexity to the investor’s decision.
-Although the wealth $X_{t}$ remains scalar, the investor’s optimal policy now depends on both $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and a high-dimensional state $\mathbf{Y}_{t}$.
-We handle this through Pontryagin’s maximum principle (Pontryagin, , 2014; Yong and Zhou, , 2012), introducing a costate for $X_{t}$ and another for $\mathbf{Y}_{t}$.
-Numerically, one can represent $\{\bm{\pi}_{t},C_{t}\}$ via neural networks, thereby enabling policy gradient or dynamic programming algorithms even in multi-asset, state-dependent environments.
+2 Multi-Asset Continuous-Time Portfolio Choice Problem with State Variables
+We consider the continuous-time portfolio choice problem introduced by Merton, (1973), in which a representative investor allocates wealth across $n$ risky assets and a risk-free asset, while consuming continuously over time.
+The investor selects an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}=(\pi_{1,t},\ldots,\pi_{n,t})^{\top}$, where $\pi_{i,t}$ denotes the proportion of current wealth $X_{t}$ invested in the $i$-th risky asset at time $t$ and notation denotes the transpose of a vector or a matrix.
+The investment opportunities are characterized by an exogenous $k$-dimensional state process $\mathbf{Y}_{t}=(Y_{1,t},\ldots,Y_{k,t})^{\top}\in\mathbb{R}^{k}$, which governs the drift and volatility of asset returns.
+For instance, $\mathbf{Y}_{t}$ could include macroeconomic indices, volatility factors, exchange rates, etc.
+The investor’s optimization problem involves dynamically choosing both the portfolio weight vector $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and the consumption rate $C_{t}$, conditional on the current state $\mathbf{Y}_{t}$ and wealth $X_{t}$.
+To analyze this problem, we apply Pontryagin’s maximum principle (Pontryagin, , 2014; Yong and Zhou, , 2012), introducing costate variables corresponding to both the wealth process $X_{t}$ and the state process $\mathbf{Y}_{t}$.
+From a computational perspective, the control variables $\{\bm{\pi}_{t},C_{t}\}$ can be parameterized using neural networks, facilitating the implementation of policy gradient methods or DP algorithms in high-dimensional, state-dependent settings.
 2.1 Problem Formulation with State Variables and Neural Network Controls
-Let $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ be a vector of economic state variables that influence the distribution of asset returns.
-For instance, $\mathbf{Y}_{t}$ could include interest rates, macroeconomic indices, volatility factors, etc.
-We assume each component $Y_{i,t}$ follows an Ito diffusion:
+We assume that each component $Y_{i,t}$ of the state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ evolves according to an Itô diffusion:
 $dY_{i,t}=\mu_{Y,i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{Y,i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dW^{Y}_{i,t},\quad i=1,\dots,k,$
-with $\operatorname{Cov}\bigl{(}dW^{Y}_{i,t},\,dW^{Y}_{j,t}\bigr{)}=\phi_{ij}\,dt.$ In vector form,
+where the Brownian motions $W^{Y}_{i,t}$ are potentially correlated, satisfying
+$\operatorname{Cov}\bigl{(}dW^{Y}_{i,t},\,dW^{Y}_{j,t}\bigr{)}=\phi_{ij}\,dt,$
+with constant correlations $\phi_{ij}$. In vector form, the state process satisfies
 $d\mathbf{Y}_{t}=\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,d\mathbf{W}^{Y}_{t},$
-where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})$ is the $k\times 1$ drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})$ is a $k\times k$ diagonal matrix of diffusion coefficients (i.e., $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})=\operatorname{diag}(\sigma_{Y,1}(t,\mathbf{Y}_{t}),\dots,\sigma_{Y,k}(t,\mathbf{Y}_{t}))$ where $\sigma_{Y,i}$ are the diffusion functions from the component-wise definition), and $\mathbf{W}_{t}^{Y}$ is a $k$-dimensional Brownian motion whose components may be correlated.
-We consider a financial market with $n$ risky assets and one risk-free asset.
-The risk-free asset accumulates wealth at an instantaneous rate $r(t,\mathbf{Y}_{t})$, which may depend on time and the current state.
-For $i=1,\dots,n$, the price $S_{i,t}$ of the $i$-th risky asset follows
-$dS_{i,t}=S_{i,t}\bigl{(}\mu_{i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dW^{X}_{i,t}\bigr{)}.$
-Here, $\mu_{i}(t,\mathbf{Y}_{t})$ is the instantaneous expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ is the instantaneous volatility of asset $i$.
-The $n$ scalar Brownian motions $W^{X}_{i,t}$ may be correlated amongst themselves ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt$) and with the state variable processes ($\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt$).
-In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})$ be the $n\times 1$ vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})$ be the $n\times n$ diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}$ be the $n\times 1$ vector of Brownian motions for the risky assets.
-The risky asset dynamics can be written compactly as
+where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})\in\mathbb{R}^{k}$ is the drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\in\mathbb{R}^{k\times k}$ is a diagonal matrix of diffusion coefficients—i.e.,
+$\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})=\operatorname{diag}(\sigma_{Y,1}(t,\mathbf{Y}_{t}),\dots,\sigma_{Y,k}(t,\mathbf{Y}_{t})),$
+and $\mathbf{W}^{Y}_{t}\in\mathbb{R}^{k}$ is a vector of correlated Brownian motions.
+We consider a financial market consisting of $n$ risky assets and one risk-free asset.
+The risk-free asset earns an instantaneous rate $r(t,\mathbf{Y}_{t})$, which may depend on time and the state variable.
+The price $S_{i,t}$ of the $i$-th risky asset evolves according to
+$dS_{i,t}=S_{i,t}\left(\mu_{i}(t,\mathbf{Y}_{t})\,dt+\sigma_{i}(t,\mathbf{Y}_{t})\,dW^{X}_{i,t}\right),\quad i=1,\dots,n,$
+where $\mu_{i}(t,\mathbf{Y}_{t})$ is the expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ the volatility of asset $i$. The Brownian motions $W^{X}_{i,t}$ driving the risky asset dynamics satisfy
+$\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt,\quad\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt,$
+with constant correlations $\psi_{ij}$ and $\rho_{ij}$.
+In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})\in\mathbb{R}^{n}$ denote the vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})\in\mathbb{R}^{n\times n}$ the diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}\in\mathbb{R}^{n}$ the Brownian motion vector for the risky assets.
+Then the dynamics of the vector of risky asset prices $\mathbf{S}_{t}\in\mathbb{R}^{n}$ can be compactly written as
 $d\mathbf{S}_{t}=\operatorname{diag}(\mathbf{S}_{t})\,\bm{\mu}(t,\mathbf{Y}_{t})\,dt+\operatorname{diag}(\mathbf{S}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$
-Throughout, we work on a complete probability space $\bigl{(}\Omega,\mathcal{F},\mathbb{P}\bigr{)}$ that carries the $(n+k)$-dimensional Brownian motion $\mathbf{W}_{t}$ and its natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$, augmented to satisfy the usual conditions (right-continuity and completeness).
-The investor allocates wealth $X_{t}$ between the risk-free asset and the $n$ risky assets.
-Let $\bm{\pi}_{t}$ be the $n\times 1$ vector representing the fraction of wealth invested in each risky asset at time $t$.
-The investor also chooses a consumption rate $C_{t}$.
-The dynamics of the investor’s wealth $X_{t}$ are governed by the following stochastic differential equation (SDE):
-$dX_{t}=\Bigl{[}\,X_{t}\Bigl{(}r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\Bigr{)}-C_{t}\Bigr{]}dt+X_{t}\,\bm{\pi}_{t}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$ (1)
-Here, $\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}$ represents the vector of instantaneous excess returns over the risk-free rate, and $\mathbf{1}$ is an $n\times 1$ vector of ones.
+To clarify the structure of risk in the economy, we define the combined $(n+k)$-dimensional Brownian motion vector
+$\mathbf{W}_{t}=\begin{pmatrix}\mathbf{W}^{X}_{t}\\ \mathbf{W}^{Y}_{t}\end{pmatrix}.$
+The instantaneous covariance matrix of its increments $d\mathbf{W}_{t}$ is denoted by $\bm{\Omega}\,dt$, with
+$\operatorname{Cov}(d\mathbf{W}_{t})=\bm{\Omega}\,dt=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
+where $\Psi=(\psi_{ij})\in\mathbb{R}^{n\times n}$ is the covariance matrix of the asset-specific shocks $d\mathbf{W}^{X}_{t}$, $\Phi=(\phi_{ij})\in\mathbb{R}^{k\times k}$ is the covariance matrix of the state shocks $d\mathbf{W}^{Y}_{t}$, and $\rho=(\rho_{ij})\in\mathbb{R}^{n\times k}$ captures the cross-covariation between the two sources of risk.
+Throughout, we work on a complete probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ that supports a $(n+k)$-dimensional Brownian motion $\mathbf{W}_{t}$ and its natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$, augmented to satisfy the usual conditions (right-continuity and completeness).
+The investor allocates wealth $X_{t}$ between the risk-free asset and the $n$ risky assets. The evolution of the investor’s wealth is described by the following stochastic differential equation (SDE):
+$dX_{t}=\left[\,X_{t}\left(r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\right)-C_{t}\,\right]dt+X_{t}\,\bm{\pi}_{t}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t},$ (1)
+where $\mathbf{1}\in\mathbb{R}^{n}$ is a vector of ones.
+The term $\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}$ represents the vector of instantaneous excess returns relative to the risk-free rate.
 We impose regularity conditions on the market parameters.
@@
 Fix a finite horizon $T>0$.
-The coefficients defining the state process $\mathbf{Y}_{t}$ and influencing the wealth dynamics $X_{t}$ via Eq.
-(1), namely $\bm{\mu}_{Y},\ \bm{\sigma}_{Y},\ \bm{\mu},\ \bm{\sigma},\ r$, satisfy the following conditions.
+The coefficients defining the state process $\mathbf{Y}_{t}$ and influencing the wealth process $X_{t}$ via Eq.
+(1), namely $\bm{\mu}_{Y},\ \bm{\sigma}_{Y},\ \bm{\mu},\ \bm{\sigma},\ r$, satisfy the following conditions:
   1. (a)
-    Measurability and continuity. Each map is jointly Borel-measurable in $(t,\mathbf{y})$ and continuous in $\mathbf{y}$ for every fixed $t\in[0,T]$.
+    Measurability and continuity. Each coefficient is jointly Borel-measurable in $(t,\mathbf{y})$ and continuous in $\mathbf{y}$ for every fixed $t\in[0,T]$.
   2. (b)
-    Global Lipschitz continuity. There exists $K_{\mathrm{Lip}}>0$ such that for all $t\in[0,T]$ and $\mathbf{y},\mathbf{y}^{\prime}\in\mathbb{R}^{k}$,
+    Global Lipschitz continuity. There exists a constant $K_{\mathrm{Lip}}>0$ such that for all $t\in[0,T]$ and $\mathbf{y},\mathbf{y}^{\prime}\in\mathbb{R}^{k}$,
     $\lVert\bm{\mu}_{Y}(t,\mathbf{y})-\bm{\mu}_{Y}(t,\mathbf{y}^{\prime})\rVert+\lVert\bm{\sigma}_{Y}(t,\mathbf{y})-\bm{\sigma}_{Y}(t,\mathbf{y}^{\prime})\rVert$
-    $+\lVert\bm{\mu}(t,\mathbf{y})-\bm{\mu}(t,\mathbf{y}^{\prime})\rVert+\lVert\bm{\sigma}(t,\mathbf{y})-\bm{\sigma}(t,\mathbf{y}^{\prime})\rVert\leq K_{\mathrm{Lip}}\lVert\mathbf{y}-\mathbf{y}^{\prime}\rVert.$
+    $\quad+\lVert\bm{\mu}(t,\mathbf{y})-\bm{\mu}(t,\mathbf{y}^{\prime})\rVert+\lVert\bm{\sigma}(t,\mathbf{y})-\bm{\sigma}(t,\mathbf{y}^{\prime})\rVert\leq K_{\mathrm{Lip}}\lVert\mathbf{y}-\mathbf{y}^{\prime}\rVert.$
   3. (c)
-    Linear growth. There exists $K_{\mathrm{LG}}>0$ such that for all $(t,\mathbf{y})\in[0,T]\times\mathbb{R}^{k}$,
+    Linear growth. There exists a constant $K_{\mathrm{LG}}>0$ such that for all $(t,\mathbf{y})\in[0,T]\times\mathbb{R}^{k}$,
     $\lVert\bm{\mu}_{Y}(t,\mathbf{y})\rVert^{2}+\lVert\bm{\sigma}_{Y}(t,\mathbf{y})\rVert^{2}+\lVert\bm{\mu}(t,\mathbf{y})\rVert^{2}+\lVert\bm{\sigma}(t,\mathbf{y})\rVert^{2}$
-    $\leq K_{\mathrm{LG}}^{2}\bigl{(}1+\lVert\mathbf{y}\rVert^{2}\bigr{)},\qquad|r(t,\mathbf{y})|\leq K_{\mathrm{LG}}\bigl{(}1+\lVert\mathbf{y}\rVert\bigr{)}.$
-Under Assumption 1, the SDE system for $(\mathbf{Y}_{t},X_{t})$ (where $X_{t}$ depends on the chosen controls $\bm{\pi}_{t},C_{t}$) admits a unique strong solution for sufficiently well-behaved controls; see Oksendal, (2013, Chs.
-5-6) for details on existence and uniqueness theorems for SDEs.
-The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ that maximises the expected discounted utility
+    $\qquad\leq K_{\mathrm{LG}}^{2}\left(1+\lVert\mathbf{y}\rVert^{2}\right),\qquad|r(t,\mathbf{y})|\leq K_{\mathrm{LG}}\left(1+\lVert\mathbf{y}\rVert\right).$
+Under Assumption 1, the SDE system for $(\mathbf{Y}_{t},X_{t})$ (where $X_{t}$ depends on the control processes $\bm{\pi}_{t}$ and $C_{t}$) admits a unique strong solution for sufficiently well-behaved controls.
+[^5] See Oksendal, (2013, Chs.
+5–6) for the existence and uniqueness of solutions to SDEs.
+The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ that maximizes the expected discounted utility
 […]
-subject to the wealth dynamics (1) and the admissibility constraints stated later in Assumption 2.
-Here $\delta>0$ is the subjective discount rate and $\kappa\geq 0$ weights the utility of terminal wealth (a bequest motive).
-Throughout we work with the Hyperbolic Absolute Risk Aversion (HARA) family of utilities,
+subject to the wealth dynamics in Eq.
+(1) and the admissibility conditions stated later in Assumption 2.
+Here, $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ governs the strength of the bequest motive by weighting terminal utility.
+Throughout, we assume preferences are represented by a utility function from the Hyperbolic Absolute Risk Aversion (HARA) family:
 […]
-whose domain is the half-line $\mathcal{D}_{U}:=\{x>x_{\min}\}$ with $x_{\min}:=-a/b$.
-HARA nests the familiar special cases CRRA $(a=0,b=1)$ and CARA $(\gamma=1,b=1)$.
-On its domain $U$ is $C^{2}$, strictly increasing and concave, and it satisfies the Inada limits $U^{\prime}(x)\to\infty$ as $x\downarrow x_{\min}$ and $U^{\prime}(x)\to 0$ as $x\to\infty$ (see Pennacchi, , 2008, §3.2).
-We restrict our attention to control processes that are economically sensible and mathematically tractable.
+defined on the half-line $\mathcal{D}_{U}:=\{x>x_{\min}\}$, with $x_{\min}:=-a/b$.
+To ensure that the utility function is well-defined and the wealth process $X_{t}$ remains strictly positive, we assume $x_{\min}\geq 0$ throughout.
+This restriction precludes bankruptcy and reflects standard modeling practice in continuous-time portfolio choice under HARA preferences.
+This class includes the Constant Relative Risk Aversion (CRRA) utility for $a=0,b=1$, and the Constant Absolute Risk Aversion (CARA) utility for $\gamma=1,b=1$.
+On its domain, $U$ is twice continuously differentiable, strictly increasing, and strictly concave.
+It also satisfies the Inada conditions.
+[^6] See Pennacchi, (2008, §3.2) for further discussion of the properties required of utility functions.:
+$U^{\prime}(x)\to\infty\text{ as }x\downarrow x_{\min},\qquad U^{\prime}(x)\to 0\text{ as }x\to\infty.$
+We consider control processes that are both economically sensible and mathematically tractable.
 Assumption 2 (Admissible controls).
@@
   1. (a)
-    Adaptedness. $(\bm{\pi}_{t},C_{t})_{0\leq t\leq T}$ is $\{\mathcal{F}_{t}\}$-progressively measurable, where $\{\mathcal{F}_{t}\}$ is the filtration generated by the underlying Brownian motions.
+    Adaptedness.
+The process $(\bm{\pi}_{t},C_{t})_{0\leq t\leq T}$ is $\{\mathcal{F}_{t}\}$-progressively measurable, where $\{\mathcal{F}_{t}\}$ is the filtration generated by the underlying Brownian motions.
   2. (b)
@@
   4. (d)
-    Domain-admissible wealth. With $x_{\min}:=-a/b$ (see the HARA utility defined above) we require
+    Domain-admissible wealth. Let $x_{\min}:=-a/b$ (as defined for the HARA utility). Then
 […]
-For clarity on the noise structure, we define the combined $(n+k)$-dimensional Brownian vector $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$.
-The covariance structure of its increments $d\mathbf{W}_{t}$ is given by the matrix $\bm{\Omega}$:
-$\operatorname{Cov}\bigl{(}d\mathbf{W}_{t}\bigr{)}=\bm{\Omega}\,dt=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
-where $\Psi=(\psi_{ij})\in\mathbb{R}^{n\times n}$ is the covariance matrix of $d\mathbf{W}^{X}_{t}$, $\Phi=(\phi_{ij})\in\mathbb{R}^{k\times k}$ is the covariance matrix of $d\mathbf{W}^{Y}_{t}$, and $\rho=(\rho_{ij})\in\mathbb{R}^{n\times k}$ is the cross-covariance matrix.
-Assumption 3 (Noise covariance).
-The block covariance matrix $\bm{\Omega}=\begin{pmatrix}\Psi&\rho\\[1.79997pt] \rho^{\top}&\Phi\end{pmatrix}$ is constant in time and uniformly positive definite: there exists $\lambda_{\Omega}>0$ such that $z^{\top}\bm{\Omega}z\geq\lambda_{\Omega}\lVert z\rVert^{2}$ for all $z\in\mathbb{R}^{n+k}$.
-(Consequently, $\Psi$ and $\Phi$ are individually positive definite and hence invertible.)
-Assumption 2(iii) gives the strict inequalities $0<C_{t}<\xi X_{t}$, while Assumption 2(iv) guarantees the wealth process stays inside the utility’s domain, $X_{t}>x_{\min}\;(=-a/b)$, almost surely.
-Since $U$ is $C^{2}$, strictly concave, and satisfies the Inada limits on $(x_{\min},\infty)$, and because $\delta>0$, we obtain
-$\mathbb{E}\!\left[e^{-\delta t}\,U(C_{t})\right]\;\leq\;e^{-\delta t}\,U\!\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}\;<\;\infty,\qquad\mathbb{E}\!\bigl{[}U(X_{T})\bigr{]}<\infty,$
-so the objective functional (2) is well defined for every admissible control pair $(\bm{\pi},C)$.
-The value function $V(t,x,\mathbf{y})$ associated with this problem represents the maximum expected utility achievable starting from wealth $x$ and state $\mathbf{y}$ at time $t$. Formally,
+Assumption 3 (Risk covariance).
+The block covariance matrix
+$\bm{\Omega}=\begin{pmatrix}\Psi&\rho\\[1.79997pt] \rho^{\top}&\Phi\end{pmatrix}$
+is constant in time and uniformly positive definite: there exists $\lambda_{\Omega}>0$ such that
+$z^{\top}\bm{\Omega}z\geq\lambda_{\Omega}\lVert z\rVert^{2},\qquad\forall z\in\mathbb{R}^{n+k}.$
+As a consequence, both $\Psi$ and $\Phi$ are individually positive definite and thus invertible.
+The assumptions are standard in continuous-time portfolio choice models.
+Specifically, Assumption 2(d) ensures that the wealth process remains within the domain of the utility function, i.e., $X_{t}>x_{\min}=-a/b$ almost surely, and also rules out arbitrage in this continuous-time setting (Dybvig and Huang, , 1988).
+Since $U$ is $C^{2}$, strictly concave, and satisfies the Inada conditions on $(x_{\min},\infty)$, and because $\delta>0$, we obtain the following integrability bounds:
+$\mathbb{E}\left[e^{-\delta t}\,U(C_{t})\right]\;\leq\;e^{-\delta t}\,U\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}\;<\;\infty,\qquad\mathbb{E}\left[U(X_{T})\right]<\infty.$
+Therefore, the objective functional (2) is well defined for every admissible control pair $(\bm{\pi},C)$.
+The value function $V(t,x,\mathbf{y})$ associated with this problem represents the maximum expected utility achievable when starting from wealth $x$ and state $\mathbf{y}$ at time $t$. Formally,
 […]
-where the supremum is taken over all admissible control processes on $[t,T]$, and $\mathbb{E}_{t,x,\mathbf{y}}[\cdot]=\mathbb{E}[\cdot|X_{t}=x,\mathbf{Y}_{t}=\mathbf{y}]$ denotes the conditional expectation.
-Assumption 4 (Uniform non-degeneracy and utility).
-  1. (a)
-    Uniform non-degeneracy of risky-asset volatility. Let $a(t,\mathbf{y}):=\bm{\sigma}(t,\mathbf{y})\,\Psi\,\bm{\sigma}(t,\mathbf{y})^{\top}$. There exists $\lambda>0$ such that
-    $\mathbf{z}^{\top}a(t,\mathbf{y})\,\mathbf{z}\;\geq\;\lambda\,\lVert\mathbf{z}\rVert^{2},\quad\forall(t,\mathbf{y},\mathbf{z})\in[0,T]\times\mathbb{R}^{k}\times\mathbb{R}^{n}.$
-  2. (b)
-    Utility specification.
-We use the HARA utility $U$ in (3); on its domain $x>x_{\min}$ it is $C^{2}$, strictly increasing ($U^{\prime}>0$), strictly concave ($U^{\prime\prime}<0$) and satisfies the Inada limits $U^{\prime}(x)\to\infty$ as $x\downarrow x_{\min}$ and $U^{\prime}(x)\to 0$ as $x\to\infty$.
-  3. (c)
-    Discount and bequest parameters. The subjective discount rate satisfies $\delta>0$, and the terminal-wealth weight satisfies $\kappa\geq 0$.
-Part (a) ensures the diffusion is non-degenerate, which guarantees parabolic regularity of the associated Hamilton–Jacobi–Bellman (HJB) equation.
-Part (b) specialises preferences to the HARA family; it reduces to CRRA when $a=0$ (irrespective of $b$), and to the familiar log-utility when $a=0$ and $\gamma=1$.
-Together with the coefficient regularity of Assumption 1, these conditions yield existence, uniqueness, and sufficient smoothness of the value function (e.g.
-$V\in C^{1,2,2}$); see (Oksendal, , 2013, Chs.
+where the supremum is taken over all admissible control processes on $[t,T]$, and $\mathbb{E}_{t,x,\mathbf{y}}[\cdot]:=\mathbb{E}[\cdot\mid X_{t}=x,\,\mathbf{Y}_{t}=\mathbf{y}]$ denotes the conditional expectation.
+The function $V$ is defined on $[0,T]\times(x_{\min},\infty)\times\mathbb{R}^{k}$.
+Assumption 4 (Uniform non-degeneracy).
+Let $a(t,\mathbf{y}):=\bm{\sigma}(t,\mathbf{y})\,\Psi\,\bm{\sigma}(t,\mathbf{y})^{\top}$. There exists a constant $\lambda>0$ such that
+$\mathbf{z}^{\top}a(t,\mathbf{y})\,\mathbf{z}\;\geq\;\lambda\,\lVert\mathbf{z}\rVert^{2},\qquad\forall(t,\mathbf{y},\mathbf{z})\in[0,T]\times\mathbb{R}^{k}\times\mathbb{R}^{n}.$
+Assumption 4 ensures the diffusion matrix in the wealth dynamics is uniformly positive definite, which guarantees parabolic regularity of the associated Hamilton–Jacobi–Bellman (HJB) equation.
+We continue to work with the HARA utility function $U$ defined in Eq.
+(3), which is $C^{2}$, strictly increasing and concave on its domain $(x_{\min},\infty)$, and satisfies the Inada conditions.
+Together with the coefficient regularity in Assumption 1, these conditions yield the existence, uniqueness, and sufficient smoothness of the value function (e.g., $V\in C^{1,2,2}$); see Oksendal, (2013, Chs.
 6–7).
-In our numerical approach, we parameterize the control policies using neural networks (NNs).
-The portfolio allocation $\bm{\pi}_{t}$ and consumption rate $C_{t}$ are represented as functions of the current time $t$, wealth $X_{t}$, and state $\mathbf{Y}_{t}$, determined by the NN parameters $\theta$ and $\phi$:
-$\bm{\pi}_{t}=\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\quad C_{t}=C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)}.$
-These parameters $\theta$ and $\phi$ are then optimized to maximize the objective $J$.
+In our numerical implementation, we parameterize the control policies using neural networks (NNs).
+Specifically, the portfolio allocation $\bm{\pi}_{t}$ and consumption rate $C_{t}$ are represented as functions of the current time $t$, wealth $X_{t}$, and state $\mathbf{Y}_{t}$, with functional forms determined by neural network parameters $\theta$ and $\phi$:
+$\bm{\pi}_{t}=\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\qquad C_{t}=C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)}.$
+The parameters $\theta$ and $\phi$ are optimized to maximize the objective functional $J$ defined in Eq. (2).
 2.2 Optimality Conditions via Pontryagin’s Maximum Principle
-We now apply Pontryagin’s Maximum Principle (PMP) to the continuous-time portfolio problem formulated in Section 2.1, where market parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$.
-PMP provides necessary conditions for optimality through a system of forward-backward stochastic differential equations (FBSDEs), which couple the dynamics of the state variables with those of associated costate variables.
-The system state includes the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$.
-The corresponding costate processes represent the sensitivity of the value function $V$ to changes in the state: $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ is the scalar costate for wealth (marginal value of wealth), and $\bm{\eta}_{t}=V_{\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})$ is the $k$-dimensional costate vector for the exogenous state variables (marginal value of state variables).
-Under an optimal control policy $\bigl{(}\bm{\pi}_{t}^{*},\,C_{t}^{*}\bigr{)}$, the evolution of the system state, comprising the wealth process $X_{t}^{*}$ and the exogenous variables $\mathbf{Y}_{t}$, is described by the following forward stochastic differential equations:
-$dX_{t}^{*}$ $=\Bigl{[}X_{t}^{*}\,\Bigl{(}r\bigl{(}t,\mathbf{Y}_{t}\bigr{)}+\bm{\pi}_{t}^{*\top}\,\bigl{(}\bm{\mu}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\Bigr{)}-C_{t}^{*}\Bigr{]}\,dt$
-$\quad+X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{X},$ (4)
-$d\mathbf{Y}_{t}$ $=\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,d\mathbf{W}_{t}^{Y},$ (5)
+We now apply Pontryagin’s Maximum Principle (PMP) to the continuous-time portfolio choice problem formulated in Section 2.1, where market parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$.
+PMP provides necessary conditions for optimality via a system of forward-backward stochastic differential equations (FBSDEs), which couple the dynamics of the state variables with those of the associated costate (or adjoint) processes.
+The system state consists of the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$.
+The corresponding costate processes represent the marginal value of state variables with respect to the value function $V$: the scalar process $\lambda_{t}:=V_{x}(t,X_{t},\mathbf{Y}_{t})$ corresponds to the sensitivity to wealth, and the vector process $\bm{\eta}_{t}:=V_{\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})\in\mathbb{R}^{k}$ captures sensitivity to the exogenous state.
+Under an optimal control policy $(\bm{\pi}_{t}^{*},C_{t}^{*})$, the forward state dynamics are governed by:
+$dX_{t}^{*}$ $=\left[X_{t}^{*}\,\bigl{(}r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{*\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\bigr{)}-C_{t}^{*}\right]dt+X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{X},$ (4)
+$d\mathbf{Y}_{t}$ $=\bm{\mu}_{Y}(t,\mathbf{Y}_{t})\,dt+\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{Y},$ (5)
 with initial conditions $X_{0}^{*}=x_{0}>x_{\min}$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$.
-Recall the definitions of the volatility matrix $\bm{\sigma}(t,\mathbf{Y}_{t})$ (for assets) and the covariance structure $\bm{\Omega}$ of the combined Brownian motion $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$.
-Complementing these forward dynamics, the optimal costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy a system of backward stochastic differential equations (BSDEs).
-These equations are driven by derivatives of the Hamiltonian $\mathcal{H}$ associated with the control problem:
-$-d\lambda_{t}^{*}$ $=\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}\bigr{)}\,dt-\mathbf{Z}_{t}^{\lambda X}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\lambda Y}\,d\mathbf{W}_{t}^{Y},$ (6)
-$-d\bm{\eta}_{t}^{*}$ $=\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}\bigr{)}\,dt-\mathbf{Z}_{t}^{\eta X}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\eta Y}\,d\mathbf{W}_{t}^{Y}.$ (7)
-Here, $\mathcal{H}_{x}=\partial\mathcal{H}/\partial x$ and $\mathcal{H}_{\mathbf{y}}=\nabla_{\mathbf{y}}\mathcal{H}$ represent the partial derivatives of the Hamiltonian.
-The terms $\mathbf{Z}_{t}^{\lambda X}\in\mathbb{R}^{1\times n}$, $\mathbf{Z}_{t}^{\lambda Y}\in\mathbb{R}^{1\times k}$, $\mathbf{Z}_{t}^{\eta X}\in\mathbb{R}^{k\times n}$, and $\mathbf{Z}_{t}^{\eta Y}\in\mathbb{R}^{k\times k}$ are adapted processes arising from the martingale representation theorem applied to the costates, collectively denoted by $\mathbf{Z}_{t}$.
-The derivatives $\mathcal{H}_{x},\mathcal{H}_{\mathbf{y}}$ generally depend on the state, costate, controls, and the $\mathbf{Z}_{t}$ terms themselves, reflecting the system’s interconnectedness through the value function’s second-order derivatives.
-The terminal conditions are determined by the objective functional: $\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)}$ and $\bm{\eta}_{T}^{*}=\mathbf{0}$.
-The application of PMP involves maximizing a Hamiltonian function.
-To properly account for the dependence of the wealth diffusion on the control $\bm{\pi}$, the Hamiltonian should incorporate the interaction between the costate dynamics (represented by $\mathbf{Z}$ terms) and the diffusion coefficients.
-Following formulations in stochastic control theory (e.g., Yong and Zhou, , 2012), the appropriate Hamiltonian $\mathcal{H}$ can be expressed as:
-$\mathcal{H}\!=\!e^{-\delta t}U(C)+\lambda\!\left[x\bigl{(}r+\pi^{\top}(\mu-r\mathbf{1})\bigr{)}-C\right]+\bm{\eta}^{\top}\mu_{Y}+\mathbf{Z}^{\lambda X}\!\bigl{(}x\sigma^{\top}\pi\bigr{)}+\bigl{(}\mathbf{Z}^{\eta X}\bigr{)}^{\!\top}\!\bigl{(}x\sigma^{\top}\pi\bigr{)}+\bigl{(}\mathbf{Z}^{\eta Y}\bigr{)}^{\!\top}\sigma_{Y}.$ (8)
-For the $\bm{\pi}$-FOC, only the drift term $\lambda x\bm{\pi}^{\top}(\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1})$ and the diffusion-coupling term $\mathbf{Z}^{\lambda X}(x\,\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}\bm{\pi})$ matter for the maximization step.
-To ensure the well-posedness of this FBSDE system and the applicability of PMP, additional smoothness conditions on the model coefficients are typically required.
+Recall that $\bm{\sigma}(t,\mathbf{Y}_{t})$ is the asset volatility matrix, and $\bm{\Omega}$ is the covariance matrix of the combined Brownian motion $\mathbf{W}_{t}=\bigl{(}\mathbf{W}^{X\top}_{t},\,\mathbf{W}^{Y\top}_{t}\bigr{)}^{\top}$.
+To characterize optimal controls, PMP introduces a Hamiltonian function that incorporates both the instantaneous utility and the effects of the controls on the system’s drift and diffusion.
+In problems with control-dependent diffusion, the Hamiltonian must also account for the interaction with costate diffusion terms (i.e., the $\mathbf{Z}$-components of the adjoint BSDEs).
+Following standard stochastic control formulations (e.g., Yong and Zhou, , 2012), the Hamiltonian $\mathcal{H}$ is defined as:
+$\mathcal{H}=e^{-\delta t}\,U(C)+\lambda\left[x\bigl{(}r+\bm{\pi}^{\top}(\bm{\mu}-r\mathbf{1})\bigr{)}-C\right]+\bm{\eta}^{\top}\bm{\mu}_{Y}+\mathbf{Z}^{\lambda X}\cdot(x\,\bm{\sigma}^{\top}\bm{\pi})+(\mathbf{Z}^{\eta Y})^{\top}\bm{\sigma}_{Y},$ (6)
+where $\lambda$ and $\bm{\eta}$ are the costate processes, $\mathbf{Z}^{\lambda X}$, $\mathbf{Z}^{\eta X}$, and $\mathbf{Z}^{\eta Y}$ represent the diffusion coefficients of the backward SDEs associated with $\lambda_{t}$ and $\bm{\eta}_{t}$, and all functions are evaluated at $(t,x,\mathbf{y},\bm{\pi},C)$, though we suppress this for clarity.
+Complementing the forward dynamics, the optimal costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy a system of backward stochastic differential equations (BSDEs), driven by the partial derivatives of the Hamiltonian $\mathcal{H}$ with respect to the state variables:
+$-d\lambda_{t}^{*}$ $=\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*}\bigr{)}\,dt-\mathbf{Z}_{t}^{\lambda X*}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\lambda Y*}\,d\mathbf{W}_{t}^{Y},$ (7)
+$-d\bm{\eta}_{t}^{*}$ $=\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*}\bigr{)}\,dt-\mathbf{Z}_{t}^{\eta X*}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\eta Y*}\,d\mathbf{W}_{t}^{Y}.$ (8)
+Here, $\mathcal{H}_{x}:=\partial\mathcal{H}/\partial x$ and $\mathcal{H}_{\mathbf{y}}:=\nabla_{\mathbf{y}}\mathcal{H}$ denote the partial derivatives of the Hamiltonian with respect to wealth and the state vector, respectively.
+The terms $\mathbf{Z}_{t}^{\lambda X}\in\mathbb{R}^{1\times n}$, $\mathbf{Z}_{t}^{\lambda Y}\in\mathbb{R}^{1\times k}$, $\mathbf{Z}_{t}^{\eta X}\in\mathbb{R}^{k\times n}$, and $\mathbf{Z}_{t}^{\eta Y}\in\mathbb{R}^{k\times k}$ are adapted processes.
+[^7] These terms arise from the martingale representation theorem applied to the backward dynamics of the costate variables.
+See Karatzas and Shreve, (1991) for a detailed treatment of the martingale representation theorem.
+These processes are collectively denoted by $\mathbf{Z}_{t}$.
+Note that $\mathcal{H}_{x}$ and $\mathcal{H}_{\mathbf{y}}$ generally depend not only on the state, costate, and control variables, but also on the $\mathbf{Z}_{t}$ terms themselves.
+This reflects the coupling between the forward and backward systems, which arises from the second-order sensitivity of the value function.
+The terminal conditions for the BSDEs are determined by the structure of the objective functional:
+$\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)},\qquad\bm{\eta}_{T}^{*}=\mathbf{0}.$
+To ensure the well-posedness of the forward–backward stochastic differential equation (FBSDE) system and the applicability of Pontryagin’s Maximum Principle (PMP), additional smoothness conditions on the model coefficients are typically required.
 Assumption 5 (Additional regularity for PMP).
@@
   1. (a)
-    $C^{1}$-regular coefficients w.r.t.
-$\mathbf{y}$.
-All coefficient functions $\bm{\mu}_{Y},\,\bm{\sigma}_{Y},\,\bm{\mu},\,\bm{\sigma}\,$ $r$ are continuously differentiable with respect to the state vector $\mathbf{y}$.
-Furthermore, their partial derivatives ($\partial_{\mathbf{y}}f$) are globally Lipschitz continuous in $\mathbf{y}$ and satisfy linear growth conditions.
-This ensures that the drift terms in the costate BSDEs, involving derivatives like $\mathcal{H}_{\mathbf{y}}$, satisfy the required regularity conditions for the existence and uniqueness of solutions (cf.
-Yong and Zhou, , 2012).
-The core requirement of Pontryagin’s Maximum Principle is that the optimal controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ must maximize the Hamiltonian $\mathcal{H}$ (8) pointwise in time, almost surely, given the optimal state and costate processes $(X_{t}^{*},\mathbf{Y}_{t},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*})$.
+    $C^{1}$-regularity with respect to $\mathbf{y}$.
+All coefficient functions—$\bm{\mu}_{Y},\,\bm{\sigma}_{Y},\,\bm{\mu},\,\bm{\sigma},\,r$—are continuously differentiable with respect to the state vector $\mathbf{y}$.
+Moreover, their partial derivatives $\partial_{\mathbf{y}}f$ are globally Lipschitz continuous in $\mathbf{y}$ and satisfy linear growth bounds.
+These conditions ensure that the drift terms in the costate BSDEs, including expressions such as $\mathcal{H}_{\mathbf{y}}$, satisfy the regularity requirements for the existence and uniqueness of solutions; see Yong and Zhou, (2012, Chapter 2 and 6).
+The core requirement of Pontryagin’s Maximum Principle is that the optimal controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ maximize the Hamiltonian $\mathcal{H}$ (6) pointwise in time, almost surely, given the optimal state and costate processes $(X_{t}^{*},\mathbf{Y}_{t},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*})$.
 Assuming interior solutions allows us to derive the first-order conditions (FOCs) for optimality.
-Maximization with respect to the consumption rate $C_{t}$ yields the standard condition:
+Maximizing with respect to the consumption rate $C_{t}$ yields the standard FOC:
 $e^{-\delta t}\,U^{\prime}(C_{t}^{*})=\lambda_{t}^{*}.$ (9)
-This classic result links optimal consumption $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$ inversely to the marginal value of wealth $\lambda_{t}^{*}$.
-Maximization with respect to the portfolio weights $\bm{\pi}_{t}$ requires differentiating $\mathcal{H}$ with respect to $\bm{\pi}$, including the term involving $\mathbf{Z}^{\lambda X}$.
-From (8), the relevant terms are $\lambda x\bm{\pi}^{\top}(\bm{\mu}-r\mathbf{1})$ and $\mathbf{Z}^{\lambda X}(x\,\bm{\sigma}^{\top}\bm{\pi})$.
-Setting the gradient $\nabla_{\bm{\pi}}\mathcal{H}=\mathbf{0}$ gives (assuming $X_{t}^{*}>0$):
-$\lambda_{t}^{*}X_{t}^{*}(\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1})+X_{t}^{*}\bm{\sigma}(t,\mathbf{Y}_{t})(\mathbf{Z}_{t}^{\lambda X*})^{\top}=\mathbf{0}.$
-Dividing by $X_{t}^{*}$, we obtain the FOC for the optimal portfolio:
+This classical result links optimal consumption to the marginal value of wealth via the inverse of the marginal utility: $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$.
+For the portfolio weights $\bm{\pi}_{t}$, we differentiate $\mathcal{H}$ with respect to $\bm{\pi}$, accounting for both the drift and diffusion terms that depend on the control.
+Setting the gradient $\nabla_{\bm{\pi}}\mathcal{H}=\mathbf{0}$, we obtain:
 […]
-This first-order condition implicitly defines the optimal portfolio $\bm{\pi}_{t}^{*}$ through its dependence on the costate $\lambda_{t}^{*}$ and the martingale component $\mathbf{Z}_{t}^{\lambda X*}$.
-To obtain an explicit expression for $\bm{\pi}_{t}^{*}$, we need to relate the martingale component $\mathbf{Z}^{\lambda X*}$ to the derivatives of the value function.
-Applying Ito’s lemma to $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ and matching terms with the BSDE (6) establishes this relationship (the precise derivation often involves careful application of Ito’s formula for processes driven by correlated Brownian motions, see e.g., standard texts or appendices like Appendix A).
-This procedure yields an expression for $\mathbf{Z}^{\lambda X*}$ in terms of $V_{xx}$, $V_{x\mathbf{y}}$, the optimal controls $\bm{\pi}_{t}^{*}$, and the model parameters ($\bm{\sigma},\bm{\sigma}_{Y},\Psi,\rho$).
-Substituting this expression into the FOC (10) and solving the resulting algebraic equation for $\bm{\pi}_{t}^{*}$ leads to the well-known optimal portfolio formula:
-$\bm{\pi}_{t}^{*}$ $=-\frac{V_{x}}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}$ (11)
-$\qquad\qquad-\frac{1}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bigl{(}\bm{\sigma}(t,\mathbf{Y}_{t})\,\bm{\rho}\,\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\bigr{)}V_{x\mathbf{y}}.$
-Here, $\Sigma(t,\mathbf{Y}_{t})=\bm{\sigma}(t,\mathbf{Y}_{t})\Psi\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ represents the instantaneous covariance matrix of the risky asset returns.
-The value function derivatives $V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\nabla_{\mathbf{y}}\lambda_{t}^{*}$ (a $k\times 1$ vector) are evaluated along the optimal path $(t,X_{t}^{*},\mathbf{Y}_{t})$.
-The formula clearly separates the portfolio into two components: the myopic demand, which optimizes the immediate risk-return trade-off as if the investment opportunity set were constant, and the intertemporal hedging demand, which adjusts the portfolio to mitigate risks arising from future stochastic changes in the state variables $\mathbf{Y}_{t}$.
-The hedging demand’s magnitude and direction depend on the correlation $\rho$ between asset returns and state variable innovations, and crucially on $V_{x\mathbf{y}}$, which measures how the marginal value of wealth changes with the state variables.
-For illustration, in the simplified case with one risky asset ($n=1$) and one state variable ($k=1$), the optimal allocation (11) becomes:
-$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{1}{X_{t}V_{xx}\sigma^{2}}(\sigma\rho\sigma_{Y})V_{xy},$
-clearly showing the separation between the myopic component, proportional to the market price of risk, and the hedging component, proportional to $V_{xy}$ and the correlation $\rho$.
+Dividing by $X_{t}^{*}$ assuming $X_{t}^{*}\neq 0$, the first-order condition for the optimal portfolio becomes:
+[…]
+This condition implicitly defines the optimal portfolio $\bm{\pi}_{t}^{*}$ through its dependence on the costate $\lambda_{t}^{*}$ and the martingale component $\mathbf{Z}_{t}^{\lambda X*}$, which captures the diffusion sensitivity of the marginal value of wealth.
+To obtain an explicit expression for $\bm{\pi}_{t}^{*}$, we must relate the coefficient $\mathbf{Z}^{\lambda X*}$ of the martingale component in the BSDE (7) to the derivatives of the value function.
+This relationship can be derived by applying Itô’s lemma to $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ and matching terms with the dynamics of the costate process in (7).
+The full derivation is presented in Appendix A.
+This procedure yields an expression for $\mathbf{Z}^{\lambda X*}$ in terms of the second-order derivatives $V_{xx}$, $V_{x\mathbf{y}}$, the optimal controls $\bm{\pi}_{t}^{*}$, and model parameters $(\bm{\sigma},\bm{\sigma}_{Y},\Psi,\rho)$.
+Substituting this into the first-order condition (10) and solving the resulting algebraic equation for $\bm{\pi}_{t}^{*}$ leads to the well-known optimal portfolio formula, originally derived by Merton, (1973):
+[…]
+$\quad-\frac{1}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,\bm{\rho}\,\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\,V_{x\mathbf{y}}.$
+Here, $\Sigma(t,\mathbf{Y}_{t}):=\bm{\sigma}(t,\mathbf{Y}_{t})\,\Psi\,\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ is the instantaneous covariance matrix of risky asset returns.
+The derivatives of the value function —$V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\nabla_{\mathbf{y}}\lambda_{t}^{*}\in\mathbb{R}^{k}$—are evaluated along the optimal trajectory $(t,X_{t}^{*},\mathbf{Y}_{t})$.
+This expression decomposes the optimal portfolio into two components: a myopic demand, which optimizes the static trade-off between expected return and risk under fixed investment opportunities, and an intertemporal hedging demand, which adjusts the portfolio to hedge against stochastic shifts in future investment opportunities.
+The hedging component is shaped by the correlation structure $\bm{\rho}$ between asset returns and state innovations, and by the cross-derivative $V_{x\mathbf{y}}$, which captures how the marginal value of wealth responds to changes in the economic state.
+For illustration, consider the simplified case with a single risky asset ($n=1$) and a single state variable ($k=1$). In this setting, the optimal portfolio allocation in (11) reduces to:
+$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{1}{X_{t}V_{xx}\sigma^{2}}\,\sigma\rho\sigma_{Y}\,V_{xy},$
+where $\sigma:=\sigma(t,Y_{t})$, $\sigma_{Y}:=\sigma_{Y}(t,Y_{t})$, and $V_{xy}:=\partial^{2}V/\partial x\partial y$.
+This expression makes explicit the decomposition of the optimal policy into two components: a myopic demand, which is proportional to the market price of risk $(\mu-r)/\sigma$, and a hedging demand, which is proportional to the cross-partial derivative $V_{xy}$ and the correlation coefficient $\rho$ between shocks to the asset return and the state variable.
 2.3 Applications to Financial Models
 2.3.1 A Multi-Asset Model with One OU Factor
-We first consider a model with $n$ risky assets whose risk premia are driven by a single Ornstein–Uhlenbeck (OU) factor.
-Specifically, each asset $i$ has a constant volatility $\sigma_{i}>0$ and a time-varying risk premium proportional to an OU state variable $Y_{t}$, such that the expected return is $r+\alpha_{i}\sigma_{i}Y_{t}$, where $r$ is the constant risk-free rate.
-The state variable $Y_{t}$ follows the OU process:
+We first consider a model with $n$ risky assets whose excess returns are driven by a single Ornstein–Uhlenbeck (OU) factor.
+Specifically, each asset $i$ has constant volatility $\sigma_{i}>0$ and a time-varying risk premium proportional to an OU state variable $Y_{t}$, so that the expected return is $r+\alpha_{i}\sigma_{i}Y_{t}$, where $r$ is the constant risk-free rate.
+The state variable $Y_{t}$ evolves as:
 $dY_{t}=\kappa_{Y}(\theta_{Y}-Y_{t})\,dt+\sigma_{Y}\,dW_{t}^{Y},$
-where $\kappa_{Y}>0$ is the mean-reversion speed, $\theta_{Y}\in\mathbb{R}$ is the long-term mean, and $\sigma_{Y}>0$ is the factor volatility.
-The Brownian motion $W_{t}^{Y}$ may be correlated with the asset return processes.
-The price dynamics for the $i$-th risky asset, $S_{i,t}$, are given by:
-$\frac{dS_{i,t}}{S_{i,t}}=(r+\alpha_{i}\sigma_{i}Y_{t})\,dt+\sigma_{i}\,dW_{i,t}^{S}.$
-Let $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$ be the vector of loading coefficients and $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ be the diagonal matrix of constant asset volatilities, where each $\sigma_{i}>0$.
-The Brownian motions driving the assets, $\mathbf{W}_{t}^{S}=(W_{1,t}^{S},\dots,W_{n,t}^{S})^{\top}$, have an $n\times n$ covariance matrix $\Psi dt$, i.e., $\mathrm{Cov}(d\mathbf{W}_{t}^{S})=\Psi\,dt$.
-Furthermore, each $dW_{i,t}^{S}$ may have a correlation $\rho_{i}$ with $dW_{t}^{Y}$; let $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$.
-Let $X_{t}$ denote total wealth. The investor chooses a consumption rate $C_{t}\geq 0$ and portfolio weights $\bm{\pi}_{t}\in\mathbb{R}^{n}$ for the risky assets. The wealth dynamics follow:
+where $\kappa_{Y}>0$ is the mean-reversion speed, $\theta_{Y}\in\mathbb{R}$ is the long-run mean, and $\sigma_{Y}>0$ is the volatility of the OU process.
+The Brownian motion $W_{t}^{Y}$ may be correlated with the return processes of the risky assets.
+The price process $S_{i,t}$ of the $i$-th risky asset follows:
+$\frac{dS_{i,t}}{S_{i,t}}=\bigl{(}r+\alpha_{i}\sigma_{i}Y_{t}\bigr{)}\,dt+\sigma_{i}\,dW_{i,t}^{S}.$
+Let $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$ be the vector of risk premia loadings, and $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ the diagonal matrix of asset volatilities.
+The Brownian motion driving the risky assets, $\mathbf{W}_{t}^{S}=(W_{1,t}^{S},\dots,W_{n,t}^{S})^{\top}$, has instantaneous covariance matrix $\mathrm{Cov}(d\mathbf{W}_{t}^{S})=\Psi\,dt$, where $\Psi\in\mathbb{R}^{n\times n}$.
+Let $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$ denote the vector of correlations between $W_{t}^{Y}$ and $W_{i,t}^{S}$.
+Let $X_{t}$ denote the investor’s wealth.
+The investor chooses a consumption rate $C_{t}\geq 0$ and portfolio weights $\bm{\pi}_{t}\in\mathbb{R}^{n}$ in the risky assets.
+The wealth dynamics is given by:
 $dX_{t}=\bigl{[}X_{t}\bigl{(}r+\bm{\pi}_{t}^{\top}Y_{t}(\bm{\sigma}\bm{\alpha})\bigr{)}-C_{t}\bigr{]}dt+X_{t}\,\bm{\pi}_{t}^{\top}\bm{\sigma}\,d\mathbf{W}_{t}^{S}.$
-The instantaneous covariance matrix of the $n$-asset returns is $\Sigma=\bm{\sigma}\,\Psi\,\bm{\sigma}$.
-For an investor with CRRA utility $U(x)$ and discount rate $\delta>0$, the optimal consumption $C_{t}^{*}$ is determined by $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$, where $\lambda_{t}^{*}$ is the costate associated with wealth $X_{t}$.
-The PMP first-order condition yields the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$ as:
-$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}\Sigma^{-1}\left\{\lambda_{t}^{*}Y_{t}(\bm{\sigma}\bm{\alpha})+(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}(\bm{\sigma}\bm{\rho})\right\}.$
-This expression involves the costate $\lambda_{t}^{*}$ and its derivatives $\partial_{x}\lambda_{t}^{*}$ and $\partial_{Y}\lambda_{t}^{*}$.
-The first term within the braces represents the myopic demand driven by the risk premium $Y_{t}(\bm{\sigma}\bm{\alpha})$, and the second term captures the intertemporal hedging demand arising from the state variable $Y_{t}$.
-If $n=1$, this simplifies to $\pi_{t}^{*}=-\frac{1}{\sigma^{2}X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\{\lambda_{t}^{*}\alpha\sigma Y_{t}+\sigma(\partial_{Y}\lambda_{t}^{*})\sigma_{Y}\rho\}$.
-For the specific case without intermediate consumption ($C_{t}=0$), an analytic solution involving Riccati ODEs can be derived using the HJB framework (Kim and Omberg, , 1996), as detailed in Appendix B.1.
+The instantaneous covariance matrix of risky asset returns is $\Sigma=\bm{\sigma}\Psi\bm{\sigma}$.
+For an investor with CRRA utility $U(x)$ and discount rate $\delta>0$, the optimal consumption is given by $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$, where $\lambda_{t}^{*}$ is the costate variable associated with wealth $X_{t}$.
+The first-order condition from Pontryagin’s Maximum Principle yields the optimal portfolio:
+$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\,\Sigma^{-1}\left\{\lambda_{t}^{*}Y_{t}(\bm{\sigma}\bm{\alpha})+(\partial_{Y}\lambda_{t}^{*})\,\sigma_{Y}\,(\bm{\sigma}\bm{\rho})\right\}.$
+This expression depends on the costate $\lambda_{t}^{*}$ and its derivatives $\partial_{x}\lambda_{t}^{*}$ and $\partial_{Y}\lambda_{t}^{*}$.
+The first term inside the braces represents the myopic demand, driven by the instantaneous risk premia $Y_{t}(\bm{\sigma}\bm{\alpha})$.
+The second term captures the intertemporal hedging demand, which arises from uncertainty in future investment opportunities and is scaled by the correlation vector $\bm{\rho}$ and the sensitivity $\partial_{Y}\lambda_{t}^{*}$.
+In the special case of a single risky asset ($n=1$), the portfolio simplifies to:
+$\pi_{t}^{*}=-\frac{1}{\sigma^{2}X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\left\{\lambda_{t}^{*}\alpha\sigma Y_{t}+\sigma\rho\sigma_{Y}\,\partial_{Y}\lambda_{t}^{*}\right\}.$
+In the case without intermediate consumption ($C_{t}\equiv 0$), the value function and optimal controls can be derived analytically via the Hamilton–Jacobi–Bellman (HJB) approach, leading to a system of Riccati ODEs (Kim and Omberg, , 1996).
+Details are provided in Appendix B.1.
 2.3.2 A Multi-Asset, Multi-Factor Model
-We then extend this setup to a more general multi-factor environment, considering $n$ risky assets whose risk premia are driven by a $k$-dimensional OU factor vector $\mathbf{Y}_{t}\in\mathbb{R}^{k}$.
-The risk premium for asset $i=1,\dots,n$ is $\sigma_{i}(\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t})$, where $\bm{\alpha}_{i}\in\mathbb{R}^{k}$ contains the factor loadings for asset $i$.
-The volatility $\sigma_{i}$ for asset $i$ is constant.
-This allows for richer interactions between multiple risk sources and asset returns.
-The factor vector $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ follows a multivariate OU process:
-$d\mathbf{Y}_{t}\;=\;\kappa_{Y}\bigl{(}\bm{\theta}_{Y}-\mathbf{Y}_{t}\bigr{)}\,dt\;+\;\bm{\sigma}_{Y}\,d\mathbf{W}_{t}^{Y},$
-where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is a mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean vector, $\bm{\sigma}_{Y}$ is the $k\times k$ diagonal matrix of factor-specific diffusion coefficients (as defined in Section 2.1), and $d\mathbf{W}_{t}^{Y}$ is a $k$-dimensional vector of Brownian motions whose components may be correlated, with $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi dt$.
-The $n$ risky assets, with prices $S_{t}^{i}$, follow:
-$\frac{dS_{t}^{i}}{S_{t}^{i}}\;=\;\Bigl{(}r+\sigma_{i}\,\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t}\Bigr{)}\,dt\;+\;\sigma_{i}\,dW_{t}^{X,i},\quad i=1,\dots,n.$
-Here, $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ is a Brownian motion with covariance matrix $\Psi dt$.
-The asset shocks $\mathbf{W}_{t}^{X}$ and factor shocks $\mathbf{W}_{t}^{Y}$ can be correlated, with $\mathrm{Cov}(dW_{t}^{X,i},dW_{t}^{Y,j})=\rho_{ij}dt$.
-Let $\bm{\rho}\in\mathbb{R}^{n\times k}$ be the matrix of cross-correlations.
-With portfolio weights $\bm{\pi}_{t}=(\pi_{t}^{1},\dots,\pi_{t}^{n})^{\top}$ and consumption $C_{t}$, wealth $X_{t}$ evolves as:
-$dX_{t}=\left[X_{t}\left(r+\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}(\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t})\right)-C_{t}\right]dt+X_{t}\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}\,dW_{t}^{X,i}.$
-For CRRA utility with discount $\delta$, the optimal consumption $C_{t}^{*}$ depends on the wealth costate $\lambda_{t}^{*}$.
-The optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$ is given by the PMP first-order condition (Eq.
-(11)), which in this context becomes:
-$\bm{\pi}_{t}^{*}=-\;\frac{1}{\,X_{t}^{*}\,\bigl{(}\partial_{x}\lambda_{t}^{*}\bigr{)}\!}\;\Sigma^{-1}\;\Bigl{\{}\,\lambda_{t}^{*}\;\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}_{t}~{}+~{}\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y}\partial_{\mathbf{Y}}\lambda_{t}^{*}\Bigr{\}},$
-where $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ is the $n\times n$ diagonal matrix of asset volatilities, $\Sigma=\bm{\sigma}\Psi\bm{\sigma}$ is the $n\times n$ asset return covariance, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}$ is the $n\times k$ matrix of factor loadings (where each $\bm{\alpha}_{i}$ is a $k\times 1$ vector of loadings for asset $i$), $\bm{\sigma}_{Y}$ is the $k\times k$ diagonal matrix of factor diffusion coefficients, and $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ is the $k\times 1$ gradient vector of the costate with respect to the factors $\mathbf{Y}_{t}$.
-This explicitly separates the demand into a myopic part driven by the current expected excess returns ($\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}_{t}$) and a hedging part driven by the sensitivity to factor changes ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and their correlations ($\bm{\rho}$, $\Psi$, $\Phi$).
-Similar to the single-factor case, an analytic solution for the terminal wealth problem ($C_{t}=0$) with CRRA utility can often be found using an exponential-quadratic value function ansatz, leading to matrix Riccati ODEs.
-Details are presented in Appendix B.2.
+We next extend the previous setup to a more general multi-factor environment, considering $n$ risky assets whose excess returns are driven by a $k$-dimensional Ornstein–Uhlenbeck (OU) factor vector $\mathbf{Y}_{t}\in\mathbb{R}^{k}$.
+The risk premium for asset $i=1,\dots,n$ is given by $\sigma_{i}(\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t})$, where $\bm{\alpha}_{i}\in\mathbb{R}^{k}$ is the vector of factor loadings for asset $i$, and $\sigma_{i}>0$ is the asset’s constant volatility.
+This formulation captures richer interactions between multiple sources of risk and asset returns.
+The factor process $\mathbf{Y}_{t}$ evolves as a multivariate OU process:
+$d\mathbf{Y}_{t}=\kappa_{Y}(\bm{\theta}_{Y}-\mathbf{Y}_{t})\,dt+\bm{\sigma}_{Y}\,d\mathbf{W}_{t}^{Y},$
+where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is the mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean, and $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ is a diagonal matrix of factor-specific volatilities (as introduced in Section 2.1).
+The Brownian motion $\mathbf{W}_{t}^{Y}\in\mathbb{R}^{k}$ driving the factors has instantaneous covariance $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi\,dt$.
+The price $S_{t}^{i}$ of each asset follows:
+$\frac{dS_{t}^{i}}{S_{t}^{i}}=\bigl{(}r+\sigma_{i}\,\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{i}\,dW_{t}^{X,i},\quad i=1,\dots,n,$
+where $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ is a Brownian motion with instantaneous covariance $\mathrm{Cov}(d\mathbf{W}_{t}^{X})=\Psi\,dt$.
+The asset and factor Brownian motions may be correlated: $\mathrm{Cov}(dW_{t}^{X,i},dW_{t}^{Y,j})=\rho_{ij}\,dt$, with cross-correlation matrix $\bm{\rho}\in\mathbb{R}^{n\times k}$.
+Given portfolio weights $\bm{\pi}_{t}=(\pi_{t}^{1},\dots,\pi_{t}^{n})^{\top}$ and consumption rate $C_{t}\geq 0$, the investor’s wealth evolves as:
+$dX_{t}=\left[X_{t}\left(r+\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t}\right)-C_{t}\right]dt+X_{t}\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}\,dW_{t}^{X,i}.$
+For CRRA utility with subjective discount rate $\delta>0$, the optimal consumption policy satisfies $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$, where $\lambda_{t}^{*}$ is the costate associated with wealth.
+The first-order condition from Pontryagin’s Maximum Principle yields the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$ as:
+$\bm{\pi}_{t}^{*}=-\;\frac{1}{X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\,\Sigma^{-1}\left\{\lambda_{t}^{*}\,\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}+\bm{\sigma}\bm{\rho}\bm{\sigma}_{Y}\,\partial_{\mathbf{Y}}\lambda_{t}^{*}\right\},$
+where: $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})\in\mathbb{R}^{n\times n}$ is the volatility matrix, $\Sigma=\bm{\sigma}\Psi\bm{\sigma}\in\mathbb{R}^{n\times n}$ is the return covariance matrix, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}\in\mathbb{R}^{n\times k}$ is the matrix of factor loadings, $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ is the factor volatility matrix, $\partial_{\mathbf{Y}}\lambda_{t}^{*}\in\mathbb{R}^{k}$ is the gradient of the costate with respect to $\mathbf{Y}_{t}$.
+This expression separates the optimal portfolio into two components: a myopic demand, driven by the instantaneous excess returns $\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}$, and an intertemporal hedging demand, which offsets the effects of changes in investment opportunities.
+The latter depends on the sensitivity of marginal utility to changes in $\mathbf{Y}_{t}$ ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$), and the correlation structure between asset and factor shocks (encoded in $\bm{\rho}$, $\Psi$, and $\Phi$).
+As in the single-factor case, when there is no intermediate consumption ($C_{t}\equiv 0$), one can derive analytic solutions using an exponential-quadratic value function ansatz, leading to a system of matrix Riccati ODEs.
+Detailed derivations are provided in Appendix B.2.
 3 A Gradient-Based Algorithm for Multi-Asset Policy Optimization
-While Pontryagin’s Maximum Principle (PMP) provides necessary optimality conditions via forward-backward stochastic differential equations (FBSDEs, Section 2.2), solving these analytically is often infeasible in high dimensions.
+While Pontryagin’s Maximum Principle (PMP) provides necessary optimality conditions via a system of forward–backward stochastic differential equations (FBSDEs; see Section 2.2), solving these analytically becomes infeasible in high dimensions.
 We therefore adopt a gradient-based numerical approach using deep learning.
-The core idea is to parameterize the investor’s policy $(\bm{\pi}_{t},C_{t})$ with neural networks $(\bm{\pi}_{\theta}(t,x,\mathbf{y}),C_{\phi}(t,x,\mathbf{y}))$, where $(\theta,\phi)$ are the parameters to be optimized.
-The objective is to find parameters $(\theta,\phi)$ that maximize the expected utility achieved by the policy $(\bm{\pi}_{\theta},C_{\phi})$.
-To ensure robustness across different starting scenarios, we aim to maximize the extended value function, $\widetilde{J}(\theta,\phi)$.
-This is defined as the expected value of the conditional expected utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ achieved under policy $(\theta,\phi)$, where the initial state $(t_{0},x_{0},\mathbf{y}_{0})$ is drawn from a distribution $\eta$ over a relevant domain $\mathcal{D}$:
+The core idea is to parameterize the investor’s policy $(\bm{\pi}_{t},C_{t})$ with neural networks $(\bm{\pi}_{\theta}(t,x,\mathbf{y}),C_{\phi}(t,x,\mathbf{y}))$, where $(\theta,\phi)$ denote the network parameters to be optimized.
+Our objective is to find parameters $(\theta,\phi)$ that maximize the expected utility delivered by the policy $(\bm{\pi}_{\theta},C_{\phi})$.
+To ensure robustness across a range of initial conditions, we define and maximize the extended value function, $\widetilde{J}(\theta,\phi)$, as the expected value of the conditional utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$, evaluated over initial states sampled from a distribution $\eta$ on a domain $\mathcal{D}$:
 […]
-$\widetilde{J}(\theta,\phi)=\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\eta}\Big{[}J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\Big{]}$ (13)
-Here, $\mathbb{E}^{\theta,\phi}[\cdot|\cdot]$ denotes expectation under the dynamics generated by policy $(\theta,\phi)$.
-This paper develops and implements the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework to maximize $\widetilde{J}(\theta,\phi)$.
-PG-DPO leverages automatic differentiation, specifically backpropagation-through-time (BPTT), applied to Monte Carlo simulations drawn from the initial state distribution $\eta$.
-A key feature is that BPTT applied to these simulations yields not only unbiased estimates of the policy gradient $\nabla\widetilde{J}$ needed for optimization, but also inherently calculates estimates of the Pontryagin costate variables ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$) along each path, representing sensitivities dictated by PMP.
-This dual output enables distinct algorithmic approaches: the baseline PG-DPO method is detailed in Section 3.1, and the advanced “Two-Stage PG-DPO” variant, which leverages these costate estimates more directly, is explored in Section 3.2.
+$\widetilde{J}(\theta,\phi)=\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\eta}\left[J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\right]$ (13)
+Here, $\mathbb{E}^{\theta,\phi}[\cdot\,|\,\cdot]$ denotes expectation under the dynamics induced by policy $(\theta,\phi)$.
+We develop and implement the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework to maximize $\widetilde{J}(\theta,\phi)$.
+PG-DPO employs automatic differentiation—specifically, backpropagation through time (BPTT)—applied to Monte Carlo simulations of system trajectories initialized from $\eta$.
+A key feature of this approach is that BPTT not only provides unbiased gradient estimates $\nabla\widetilde{J}$ for policy optimization, but also computes pathwise estimates of the Pontryagin costate variables $(\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)})$, capturing marginal value sensitivities as prescribed by PMP.
+This dual output enables two algorithmic variants: the baseline PG-DPO algorithm described in Section 3.1, and an advanced Two-Stage PG-DPO algorithm, which explicitly exploits costate information, presented in Section 3.2.
 3.1 Baseline PG-DPO for the Extended Objective
@@
     $\quad+X_{k}^{(i)}\bm{\pi}_{k}^{(i)\top}\bm{\sigma}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta\mathbf{W}_{k}^{(i),X},$
-    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$, $\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})$ is the vector of diffusion coefficients applied element-wise to the components of $\Delta\mathbf{W}_{k}^{(i),Y}$, and the correctly correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ are derived from the Cholesky factor $L$ of the overall covariance matrix $\bm{\Omega}$.
+    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$, $\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})$ is the vector of diffusion coefficients applied element-wise to the components of $\Delta\mathbf{W}_{k}^{(i),Y}$, and the correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ are derived from the Cholesky factor $L$ of the covariance matrix $\bm{\Omega}$.
   3. 3.
@@
 Role of intermediate BPTT values and the “Pontryagin-guided” nature.
-It is important to understand the role of the intermediate values computed during BPTT in Step 4.
-As BPTT calculates the policy gradient $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$ via the chain rule, it inherently computes the partial derivatives of the realized reward $J^{(i)}(\dots)$ with respect to the intermediate states along path $i$:
-$\lambda_{k}^{(i)}=\frac{\partial J^{(i)}(t_{0}^{(i)},\dots)}{\partial X_{k}^{(i)}},\quad\bm{\eta}_{k}^{(i)}=\frac{\partial J^{(i)}(t_{0}^{(i)},\dots)}{\partial\mathbf{Y}_{k}^{(i)}}.$
-These sensitivities, computed mechanically by AD, are directly linked to the discrete adjoint Backward Stochastic Differential Equation (BSDE) structure required by Pontryagin’s Maximum Principle (PMP).
-To illustrate this crucial connection for the wealth component (denoting $\lambda_{k}^{(i)}$ for $\lambda_{k}^{X,(i)}$ for simplicity), consider the discrete Euler-Maruyama step for wealth under policy $(\bm{\pi}_{k}^{(i)},C_{k}^{(i)})$:
-$X_{k+1}^{(i)}=X_{k}^{(i)}+\Bigl{[}X_{k}^{(i)}\Bigl{(}r(t_{k},\mathbf{Y}_{k}^{(i)})+\bm{\pi}_{k}^{(i)\top}\bigl{(}\bm{\mu}(t_{k},\mathbf{Y}_{k}^{(i)})-r(t_{k},\mathbf{Y}_{k}^{(i)})\mathbf{1}\bigr{)}\Bigr{)}-C_{k}^{(i)}\Bigr{]}\Delta t+X_{k}^{(i)}\bm{\pi}_{k}^{(i)\top}\bm{\sigma}(t_{k},\mathbf{Y}_{k}^{(i)})\,\Delta\mathbf{W}_{k}^{(i)}.$
-Applying the chain rule one step backward ($\lambda_{k}^{(i)}=\lambda_{k+1}^{(i)}\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}$) and using a first-order expansion yields the following backward recurrence relation along the path $(i)$ (this illustrative recurrence primarily shows influences from the wealth dynamics; a full derivation of the drift for $\lambda_{k}^{(i)}$ from the total reward $J^{(i)}$ would also account for terms related to consumption $U(C_{k}^{(i)})$ and the discount factor in its definition within the Hamiltonian context):
-$\lambda_{k}^{(i)}\approx\lambda_{k+1}^{(i)}+\lambda_{k+1}^{(i)}\bigl{[}r+\pi_{k}^{(i)}(\mu-r)\bigr{]}\Delta t+\lambda_{k+1}^{(i)}\pi_{k}^{(i)}\sigma\,\Delta W_{k}^{(i)}.$
-This equation inherently possesses the structure of a discrete BSDE operating along the specific simulation path $(i)$.
-To formally connect this to the BSDE formulation in PMP, which involves processes adapted to the filtration $\{\mathcal{F}_{t_{k}}\}$ (representing available information at time $t_{k}$), one defines the adapted costate $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}|\mathcal{F}_{t_{k}}]$ and the predictable martingale component $Z_{k}$ (related to the conditional expectation of $\lambda_{k+1}^{(i)}\Delta W_{k}^{(i)}$).
-Taking the conditional expectation ($\mathbb{E}[\cdot|\mathcal{F}_{t_{k}}]$) of a more complete pathwise recurrence relation (inclusive of all terms affecting $J^{(i)}$) and carefully defining $Z_{k}$ using a covariation-based approach, leads precisely to the standard discrete BSDE for the adapted process $\lambda_{k}$:
-$\lambda_{k}=\lambda_{k+1}+\underbrace{\bigl{(}\lambda_{k+1}[r+\pi_{k}(\mu-r)]+\pi_{k}\,\sigma\,Z_{k}\bigr{)}}_{\approx\partial_{x}\mathcal{H}(t_{k},\dots)\text{ from PMP}}\Delta t-Z_{k}\,\Delta W_{k}.$ (14)
-Crucially, the drift term derived through this process corresponds exactly to the derivative of the discrete-time Hamiltonian ($\partial_{x}\mathcal{H}$) from PMP with respect to the state variable $x$.
-A similar structure holds for the costate vector $\bm{\eta}_{k}$ associated with the state variables $\mathbf{Y}_{k}$.
-Therefore, the values $\lambda_{k}^{(i)}$ and $\bm{\eta}_{k}^{(i)}$ computed by BPTT are indeed pathwise numerical counterparts related to the Pontryagin costates, satisfying the relevant discrete BSDE structure.
-Crucially, in the baseline PG-DPO algorithm, these costate estimates themselves (and their potential derivatives w.r.t.
-states) are not directly used in the parameter update rule (Step 6), which only requires the final gradient $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$.
-Nevertheless, the fact that BPTT provides access to these theoretically grounded estimates is fundamental.
-Firstly, they are essential inputs for the Projected PGDPO variant discussed in the next subsection (§3.2).
-The baseline algorithm thus lays the groundwork by generating these estimates.
-Secondly, this deep connection explains why the algorithm is termed Pontryagin-Guided.
-Even though the baseline update only uses $\nabla_{(\theta,\phi)}J^{(i)}(\dots)$, the calculation of this gradient via BPTT structurally mirrors the backward propagation of sensitivities inherent in PMP’s adjoint equations (BSDEs), using the costate estimates $\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ implicitly within the chain rule (cf.
-Huh, , 2024).
-Consequently, the optimization process, driven by these BPTT-computed gradients, is implicitly guided by the optimality conditions derived from Pontryagin’s principle, aiming to maximize the expected objective $\widetilde{J}$.
+It is important to understand the role of the intermediate values computed during BPTT.
+By applying the chain rule backward from the final reward $J^{(i)}$, BPTT inherently calculates the sensitivity of the reward to the intermediate state variables.
+These sensitivities are the pathwise numerical counterparts to the Pontryagin costates:
+$\lambda_{k}^{(i)}=\frac{\partial J^{(i)}}{\partial X_{k}^{(i)}},\qquad\bm{\eta}_{k}^{(i)}=\frac{\partial J^{(i)}}{\partial\mathbf{Y}_{k}^{(i)}}.$ (14)
+To connect this computational recurrence to the formal PMP framework, we must generalize it to an $\mathcal{F}_{t_{k}}$-adapted process $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}\mid\mathcal{F}_{t_{k}}]$.
+This transition from a path-specific value to an adapted process reveals a deep structural correspondence, which we formalize below.
+Lemma 1 (Pathwise BPTT Recurrence).
+Along any simulated trajectory $i$, the BPTT algorithm computes the backward recursion for the pathwise sensitivity $\lambda_{k}^{(i)}$ as:
+$\lambda_{k}^{(i)}$ $=\lambda_{k+1}^{(i)}+\lambda_{k+1}^{(i)}\bigl{[}r_{k}+\bm{\pi}_{k}^{\!\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{]}\Delta t+\bigl{[}e^{-\delta(t_{k}-t_{0})}U^{\prime}(C_{k}^{(i)})-\lambda_{k+1}^{(i)}\bigr{]}\frac{\partial C_{k}^{(i)}}{\partial X_{k}^{(i)}}\,\Delta t$
+$\quad+\lambda_{k+1}^{(i)}\,\bm{\pi}_{k}^{\!\top}\bm{\sigma}_{k}\,\Delta\mathbf{W}_{k}^{X,(i)}.$ (15)
+Proof.
+The recurrence relation for the pathwise costate $\lambda_{k}^{(i)}:=\partial J^{(i)}/\partial X_{k}^{(i)}$ is derived by applying the chain rule.
+The total reward $J^{(i)}$ is decomposed into the utility from the current step $k$ and the total future reward $J_{k+1}^{(i)}$.
+$\lambda_{k}^{(i)}=\frac{\partial}{\partial X_{k}^{(i)}}\left(e^{-\delta(t_{k}-t_{0})}U(C_{k}^{(i)})\Delta t\right)+\frac{\partial J_{k+1}^{(i)}}{\partial X_{k}^{(i)}}.$
+The second term becomes $\lambda_{k+1}^{(i)}(\partial X_{k+1}^{(i)}/\partial X_{k}^{(i)})$.
+Differentiating the discrete wealth dynamics, which includes the consumption term $-C_{k}^{(i)}\Delta t$, with respect to $X_{k}^{(i)}$ yields:
+$\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}=1+\bigl{(}r_{k}+\bm{\pi}_{k}^{\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{)}\Delta t-\frac{\partial C_{k}^{(i)}}{\partial X_{k}^{(i)}}\Delta t+\bm{\pi}_{k}^{\top}\bm{\sigma}_{k}\Delta\mathbf{W}_{k}^{X,(i)}.$
+Substituting these components and rearranging the terms for $\lambda_{k}^{(i)}$ yields the expression in Eq. (1). ∎
+Lemma 2 (Quadratic-Covariation Drift).
+Let $\lambda_{k+1}^{(i)}=\lambda_{k+1}+\widetilde{\mathbf{Z}}^{\lambda X}_{k}\Delta\mathbf{W}_{k}^{X,(i)}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\Delta\mathbf{W}_{k}^{Y,(i)}$ be the martingale representation of $\lambda_{k+1}^{(i)}$, where $\widetilde{\mathbf{Z}}^{\lambda X}_{k}$ is a row vector.
+The conditional expectation of the stochastic term in Eq.
+(1) is:
+$\mathbb{E}\!\Bigl{[}\lambda_{k+1}^{(i)}(\bm{\pi}_{k}^{\!\top}\bm{\sigma}_{k}\,\Delta\mathbf{W}_{k}^{X,(i)})\Bigm{|}\mathcal{F}_{t_{k}}\Bigr{]}\;=\;\bigl{(}\widetilde{\mathbf{Z}}^{\lambda X}_{k}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\bm{\rho}_{Y\!X}\bigr{)}(\bm{\sigma}_{k}^{\!\top}\bm{\pi}_{k})\,\Delta t.$
+Proof.
+The product inside the expectation is an inner product of vectors.
+The term involving the adapted part, $\mathbb{E}[\lambda_{k+1}(\dots)|\mathcal{F}_{t_{k}}]$, is zero as it is the expectation of a martingale increment.
+The non-zero terms arise from the quadratic covariation of the raw martingale components of $\lambda_{k+1}^{(i)}$ with the wealth shock.
+Using the standard properties $\mathbb{E}[(\Delta\mathbf{W}_{k}^{X})(\Delta\mathbf{W}_{k}^{X})^{\top}|\mathcal{F}_{t_{k}}]=I\Delta t$ and $\mathbb{E}[(\Delta\mathbf{W}_{k}^{Y})(\Delta\mathbf{W}_{k}^{X})^{\top}|\mathcal{F}_{t_{k}}]=\bm{\rho}_{Y\!X}\Delta t$, where $\bm{\rho}_{Y\!X}$ is the cross-correlation matrix, yields the stated result.
+∎
+Theorem 3 (BPTT–PMP Correspondence).
+Define the effective martingale component as $\mathbf{Z}^{\lambda X}_{k}:=\widetilde{\mathbf{Z}}^{\lambda X}_{k}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\bm{\rho}_{Y\!X}$.
+The adapted costate process $(\lambda_{k})_{k=0}^{N}$ satisfies the discrete-time backward stochastic difference equation:
+$\lambda_{k}$ $=\lambda_{k+1}+\Bigl{[}\lambda_{k+1}\bigl{(}r_{k}+\bm{\pi}_{k}^{\!\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{)}+\mathbf{Z}^{\lambda X}_{k}(\bm{\sigma}_{k}^{\!\top}\bm{\pi}_{k})+\bigl{(}e^{-\delta t_{k}}U^{\prime}(C_{k})-\lambda_{k+1}\bigr{)}\frac{\partial C_{k}}{\partial X_{k}}\Bigr{]}\Delta t$
+$\quad-\widetilde{\mathbf{Z}}^{\lambda X}_{k}\,\Delta\mathbf{W}^{X}_{k}-\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\,\Delta\mathbf{W}^{Y}_{k}.$ (16)
+Furthermore, the drift term in the square brackets of Eq.
+(3) is identical to the partial derivative of the Hamiltonian, $\partial_{x}\mathcal{H}$, evaluated at the corresponding state and control variables.
+Proof.
+We begin by taking the conditional expectation $\mathbb{E}[\cdot\mid\mathcal{F}_{t_{k}}]$ of the entire pathwise recurrence in Eq.
+(1) from Lemma 1.
+The left-hand side becomes $\lambda_{k}$ by definition.
+On the right-hand side, the expectations of the drift terms are straightforward.
+The crucial step is handling the stochastic term $\mathbb{E}[\lambda_{k+1}^{(i)}(\dots)|\mathcal{F}_{t_{k}}]$, for which we substitute the result from Lemma 2.
+This transforms the stochastic term into the additional drift component involving the effective process $\mathbf{Z}^{\lambda X}_{k}$.
+After collecting all drift terms and separating the martingale components, we arrive at the discrete-time BSDE presented in Eq.
+(3).
+The final step involves a direct algebraic comparison of this drift term with the definition of the Hamiltonian $\mathcal{H}$ from Section 2.2, which confirms their identity.
+∎
+Remark 1.
+The Martingale Representation Theorem, cited from Karatzas and Shreve, (1991), is a continuous-time result.
+Its application in our discrete-time grid setting is justified as the discrete process can be viewed as a sampling of an underlying continuous process, for which the representation holds on each interval $[t_{k},t_{k+1}]$.
+This theorem provides the theoretical guarantee that our “Pontryagin-guided” approach is well-founded.
+BPTT is not merely a generic optimization heuristic; it is a computational procedure that implicitly solves the necessary optimality conditions of PMP.
+While the baseline PG-DPO uses this powerful link implicitly to find effective policy gradients, our Projected PG-DPO variant leverages it explicitly by taking the BPTT-computed costates and using them to construct a policy that is structurally consistent with financial theory.
 3.2 Projected PG-DPO (P-PGDPO) Algorithm
-While the baseline PG-DPO algorithm described in Section 3.1 effectively optimizes the extended objective $\widetilde{J}$ by training the policy networks $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end, it can sometimes require a significant number of iterations for the networks to fully converge.
-The Projected PG-DPO (P-PGDPO) variant offers a potentially more computationally efficient alternative by exploiting an empirical observation: the costate estimates ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$ and their relevant derivatives with respect to states, like $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$), which are generated during BPTT, often stabilize and become reasonably accurate much faster than the full policy networks achieve optimality.
-This P-PGDPO approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly by projecting them onto the manifold of policies satisfying the PMP first-order conditions.
+While the baseline PG-DPO algorithm described in Section 3.1 effectively optimizes the extended objective $\widetilde{J}$ by training the policy networks $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end, it may require a large number of iterations for the networks to fully converge.
+The Projected PG-DPO (P-PGDPO) variant offers a potentially more computationally efficient alternative, based on the empirical observation that the costate estimates—$(\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)})$ and their relevant derivatives (e.g., $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$)—often stabilize and become reasonably accurate well before the policy networks reach convergence.
+P-PGDPO leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly, by projecting them onto the manifold of policies that satisfy the PMP first-order conditions.
+In this way, it bypasses the need for fully converged policy networks in order to deploy effective controls.
+The procedure unfolds in two distinct stages, as described below using a single-level enumeration:
+  1. 1.
+    Stage 1: Warm-Up Phase for Costate Estimation
+    First, we execute the Baseline PG-DPO procedure (see Section 3.1), which optimizes $\widetilde{J}$, for a predetermined, relatively small number of iterations $K_{0}$.
+The primary objective of this stage is not to obtain fully optimized networks $(\bm{\pi}_{\theta},C_{\phi})$, but rather to run the BPTT process sufficiently long for the underlying mechanism generating the costate estimates ($\lambda_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial X_{k}^{(i)}$, $\bm{\eta}_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial\mathbf{Y}_{k}^{(i)}$) and their relevant state-derivatives (e.g., $\partial_{x}\lambda_{k}^{(i)}$, $\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$) to stabilize.
+We let $(\theta^{*},\phi^{*})$ denote the policy parameters obtained at the end of this warm-up phase.
+  2. 2.
+    Stage 2: Analytic Control Deployment
+    After completing the $K_{0}$ warm-up iterations, the trained policy networks $(\bm{\pi}_{\theta^{*}},C_{\phi^{*}})$ are typically not directly used for generating controls during deployment.
+Instead, to determine the control actions for any given state $(t_{k},X_{k},\mathbf{Y}_{k})$ encountered during deployment (let this deployment time be $t_{k}=t_{deploy}$), we employ a Monte Carlo-like averaging procedure to obtain reliable costate estimates needed for the PMP formulas:
+    First, to obtain the stabilized costate estimates corresponding to the state $(t_{deploy},X_{k},\mathbf{Y}_{k})$, we perform the following averaging: Simulate multiple ($M_{\text{MC}}$) forward paths starting from $(t_{deploy},X_{k},\mathbf{Y}_{k})$ (or states very close to it) up to the horizon $T$, using the fixed policy parameters $(\theta^{*},\phi^{*})$ obtained from Stage 1.
+For each simulated path $j=1,\dots,M_{\text{MC}}$, calculate its realized reward $J^{(j)}(t_{deploy},X_{k},\mathbf{Y}_{k};\theta^{*},\phi^{*})$ (using discounting relative to $t_{deploy}$ as in Algorithm 2) and apply BPTT to obtain the costate estimates $\lambda_{deploy}^{(j)},\partial_{x}\lambda_{deploy}^{(j)},\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$ specifically at the starting node $t_{deploy}$ for that path $j$.
+The final “stabilized” estimates are then computed by averaging these values across the $M_{\text{MC}}$ paths:
+    $\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\lambda_{deploy}^{(j)},\quad\partial_{x}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{x}\lambda_{deploy}^{(j)},\quad\partial_{\mathbf{Y}}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$
+    This averaging acts as a Monte Carlo estimation step, significantly reducing the variance inherent in the single-path BPTT estimates derived during Stage 1 and providing more robust inputs for the PMP formulas.
     Algorithm 2 Projected PG‑DPO (P-PGDPO): Warm‑up Training & Deployment
@@
     19:return $(\bm{\pi}^{\mathrm{PMP}},C^{\mathrm{PMP}})$
-This P-PGDPO approach leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly from the PMP first-order conditions, bypassing the need for fully converged policy networks for deployment.
-The procedure consists of two distinct stages, described below using a single level of enumeration:
-  1. 1.
-    Stage 1: Warm-Up Phase for Costate Estimation
-    First, we execute the Baseline PG-DPO process, as detailed in Section 3.1 (which optimizes $\widetilde{J}$), for a predetermined, relatively small number of iterations, $K_{0}$.
-The primary objective of this stage is not to obtain fully optimized networks $(\bm{\pi}_{\theta},C_{\phi})$, but rather to run the BPTT process sufficiently long for the underlying mechanism generating the costate estimates ($\lambda_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial X_{k}^{(i)}$, $\bm{\eta}_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial\mathbf{Y}_{k}^{(i)}$) and their relevant state-derivatives (e.g., $\partial_{x}\lambda_{k}^{(i)}$, $\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$) to stabilize.
-We let $(\theta^{*},\phi^{*})$ denote the policy parameters obtained at the end of this warm-up phase.
-  2. 2.
-    Stage 2: Analytic Control Deployment
-    After completing the $K_{0}$ warm-up iterations, the trained policy networks $(\bm{\pi}_{\theta^{*}},C_{\phi^{*}})$ are typically not directly used for generating controls during deployment.
-Instead, to determine the control actions for any given state $(t_{k},X_{k},\mathbf{Y}_{k})$ encountered during deployment (let this deployment time be $t_{k}=t_{deploy}$), we employ a Monte Carlo-like averaging procedure to obtain reliable costate estimates needed for the PMP formulas:
-    First, to obtain the stabilized costate estimates corresponding to the state $(t_{deploy},X_{k},\mathbf{Y}_{k})$, we perform the following averaging: Simulate multiple ($M_{\text{MC}}$) forward paths starting from $(t_{deploy},X_{k},\mathbf{Y}_{k})$ (or states very close to it) up to the horizon $T$, using the fixed policy parameters $(\theta^{*},\phi^{*})$ obtained from Stage 1.
-For each simulated path $j=1,\dots,M_{\text{MC}}$, calculate its realized reward $J^{(j)}(t_{deploy},X_{k},\mathbf{Y}_{k};\theta^{*},\phi^{*})$ (using discounting relative to $t_{deploy}$ as in Algorithm 2) and apply BPTT to obtain the costate estimates $\lambda_{deploy}^{(j)},\partial_{x}\lambda_{deploy}^{(j)},\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$ specifically at the starting node $t_{deploy}$ for that path $j$.
-The final “stabilized” estimates are then computed by averaging these values across the $M_{\text{MC}}$ paths:
-    $\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\lambda_{deploy}^{(j)},\quad\partial_{x}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{x}\lambda_{deploy}^{(j)},\quad\partial_{\mathbf{Y}}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$
-    This averaging acts as a Monte Carlo estimation step, significantly reducing the variance inherent in the single-path BPTT estimates derived during Stage 1 and providing more robust inputs for the PMP formulas.
     Next, we plug these averaged costate estimates ($\lambda_{deploy},\partial_{x}\lambda_{deploy},\partial_{\mathbf{Y}}\lambda_{deploy}$) directly into the analytical formulas derived from the Pontryagin Maximum Principle’s first-order conditions (Section 2.2).
@@
     Finally, these analytically computed values $(C_{deploy}^{\mathrm{PMP}},\bm{\pi}_{deploy}^{\mathrm{PMP}})$, based on the averaged costate estimates, are used as the control actions for the state $(t_{deploy},X_{k},\mathbf{Y}_{k})$ during deployment.
-In summary, the Projected PG-DPO (P-PGDPO) approach performs a short warm-up training (Stage 1) using the baseline PG-DPO process to obtain a policy $(\theta^{*},\phi^{*})$ whose BPTT generates reasonably stable costate estimates.
-It then leverages these estimates in Stage 2, refining them through Monte Carlo averaging at the point of deployment and plugging them into the known analytical structure of the optimal control derived from PMP.
-This strategy can significantly reduce the overall computational burden while yielding high-quality, near-optimal controls.
-The Monte Carlo averaging step in Stage 2 is crucial for obtaining reliable inputs for the PMP formulas from the potentially noisy single-path estimates.
+In summary, the Projected PG-DPO (P-PGDPO) approach begins with a short warm-up phase (Stage 1), using the baseline PG-DPO procedure to obtain a preliminary policy $(\theta^{*},\phi^{*})$ whose backpropagation-through-time (BPTT) produces reasonably stable costate estimates.
+In Stage 2, these costate estimates are refined via Monte Carlo averaging at deployment time and then substituted into the analytical expression for the optimal control derived from PMP.
+This two-stage strategy can substantially reduce computational costs while delivering high-quality, near-optimal controls.
+The Monte Carlo averaging step in Stage 2 is essential for obtaining reliable inputs to the PMP formula, mitigating the noise inherent in single-path estimates.
 3.3 Rationale for the Projected PG-DPO (P-PGDPO) Approach
-In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are modeled by stochastic differential equations (SDEs), the Value Function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically solves a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming and Soner, , 2006; Yong and Zhou, , 2012).
-This HJB equation exhibits a parabolic form.
-The specific nature of its parabolicity is primarily determined by the diffusion coefficients of the underlying SDEs, namely $\bm{\sigma}$ (related to assets) and $\bm{\sigma}_{Y}$ (related to state variables).
-While the first-order drift terms (involving $\bm{\mu},\bm{\mu}_{Y}$) are crucial for the solution’s properties, they do not affect the parabolicity classification, which is mainly governed by the second-order diffusion terms.
-The behavior of the diffusion matrix $a(t,x,\mathbf{y})$ (the matrix multiplying second-order derivatives in the HJB equation) allows for classifying parabolic PDEs into three main classes (see, e.g., Evans, , 2022).
-This classification significantly impacts the regularity (smoothness) of the Value Function $V$ and its associated Costates (the first-order derivatives $\nabla V$) (cf.
-Evans, , 2022, Ch.
-6-7).
-Costates represent the sensitivity of the Value Function, and Pontryagin’s Maximum Principle links optimal Costates to the optimal policy.
-The core idea of Projected PG-DPO (P-PGDPO) involves estimating Costates and their derivatives via BPTT and using these estimates in the PMP formulas.
-Therefore, the regularity of the Costates—tied to the model’s parabolicity class—directly affects the difficulty of Costate estimation and the effectiveness of the P-PGDPO method.
+In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are governed by stochastic differential equations (SDEs), the value function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically satisfies a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming and Soner, , 2006; Yong and Zhou, , 2012).
+The HJB equation is parabolic, and the specific nature of its parabolicity is primarily determined by the diffusion coefficients in the underlying SDEs—namely $\bm{\sigma}$ (for asset dynamics) and $\bm{\sigma}_{Y}$ (for state dynamics).
+While first-order drift terms (involving $\bm{\mu},\bm{\mu}_{Y}$) influence the structure of the solution, they do not affect the parabolicity classification, which is governed by the second-order terms.
+The behavior of the diffusion matrix $a(t,x,\mathbf{y})$, which multiplies the second-order derivatives in the HJB equation, allows the PDE to be classified into different parabolic types (see, e.g., Evans, , 2022).
+This classification significantly impacts the regularity (smoothness) of the value function $V$ and its associated costates $\nabla V$ (cf.
+Evans, , 2022, Chs.
+6–7).
+Costates represent the sensitivity of the value function with respect to states and wealth, and Pontryagin’s Maximum Principle links these costates directly to the optimal policy.
+The core idea of the Projected PG-DPO (P-PGDPO) method is to estimate these costates—and their derivatives—via BPTT and substitute them into the analytical expressions for the optimal control provided by PMP.
+The regularity of costates, governed by the parabolicity class of the HJB equation, directly influences the accuracy of these estimates and thus the effectiveness of P-PGDPO.
 Table 1: Classification of Parabolic HJB PDEs in Financial Models
@@
 ($\implies$ Value Function / Costate regularity may break down near the boundary) Examples: CIR (Cox-Ingersoll-Ross) interest rate/volatility model (diffusion term like $\sigma_{Y}\sqrt{Y_{t}}$); Heston stochastic volatility model (asset diffusion term like $\sigma\sqrt{\nu_{t}}$).
-Models falling into classes (U) and (H) are expected to possess sufficiently regular Costates, providing a favorable setting for the estimation required by the Projected PG-DPO (P-PGDPO) method.
-The numerical experiments in this paper are conducted using models belonging to classes (U) or (H), where P-PGDPO effectively leverages the expected regularity of the Costates.
-The method is particularly effective as it generally avoids visiting degenerate boundaries during practical implementation.
-Building upon this expected regularity, a more formal mathematical justification for the P-PGDPO approach can be provided.
-The objective is to establish that if the policy $(\pi,c)$ obtained from the Stage 1 warm-up phase satisfies a sufficiently small Pontryagin first-order condition (FOC) residual $\varepsilon$, then the feedback policy $\widehat{\pi}$ constructed in Stage 2 using BPTT estimates $\widehat{\lambda}$ (with BPTT estimation error $\delta_{\mathrm{BPTT}}$) will be close to the true optimal policy $\pi^{*}$.
-This analysis relies on certain standing assumptions concerning parabolic regularity and the FOC gap, which are formally stated as Assumption 6 and 7 in Appendix C.
-We use $L^{q,p}(Q_{T})$ to denote the standard mixed space-time Lebesgue norm over the domain $Q_{T}=(0,T)\times\Omega\subseteq\mathbb{R}^{d+1}$, as detailed in Appendix C.
-The main result of this analysis is encapsulated in the following theorem, the proof of which is detailed in Appendix C along with supporting lemmas and propositions.
-Theorem 1 (Policy Gap Bound).
-Under Assumptions 6 and 7 (detailed in Appendix C), let $(\pi,c)$ be the policy from Stage 1 satisfying the FOC-gap condition $\|\mathcal{R}_{\mathrm{FOC}}(\pi,c;\lambda)\|_{L^{q,p}}\leq\varepsilon$, where $\lambda=\nabla V^{\pi,c}$.
-Let $\pi^{*}$ be the true optimal policy, and let $\widehat{\pi}=F(\widehat{\lambda})$ be the policy deployed in Stage 2 using BPTT estimates $\widehat{\lambda}$, where $F$ is the explicit mapping from costates $(\lambda,\nabla\lambda)$ to controls derived from the PMP FOCs.
-Assume $F$ is locally Lipschitz continuous with constant $C_{F}$.
-If the BPTT estimation error $\|\widehat{\lambda}-\nabla V^{\pi,c}\|_{L^{q,p}}\leq\delta_{\mathrm{BPTT}}$ (where $\delta_{\mathrm{BPTT}}$ quantifies the precision of BPTT-derived costate estimates relative to Stage 1 costates), then the policy gap is bounded by:
-$\boxed{\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\mathrm{tot}}(\delta_{\mathrm{BPTT}}+\varepsilon)}$
-Here, $C_{\mathrm{tot}}$ is a constant that depends on $C_{F}$ and the constant $C_{v}^{\prime\prime}$ (derived from Proposition C.2 in Appendix C).
-The proof of Theorem 1 follows a logical chain: the FOC-gap $(\leq\varepsilon)$ implies that errors in PDE coefficients and source terms are of order $O(\varepsilon)$.
-These, in turn, through PDE stability, imply that the error in the value function’s gradient $E_{v}=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}}$ is also of order $O(\varepsilon)$.
-Combining this with the BPTT estimation error $\delta_{\mathrm{BPTT}}$ leads to the final policy-gap bound.
-The detailed arguments, including the necessary lemmas and propositions, are provided as sketches in Appendix C.
-It is important to note that all constants involved depend quantitatively on model parameters, PDE regularity, and utility specifications.
-Full rigorous proofs would require meticulous tracking of these dependencies and bounds.
-Our formal analysis (detailed in Appendix C) demonstrates that under suitable regularity conditions, particularly uniform parabolicity, a small Pontryagin first-order condition (FOC) gap ($\varepsilon$) achieved during the Stage 1 warm-up guarantees that the BPTT-derived costate estimates are $O(\varepsilon)$-close to the true optimal costates (up to the BPTT estimation error $\delta_{\mathrm{BPTT}}$).
-This provides a rigorous justification for the core idea of the P-PGDPO method: substituting these rapidly stabilizing and sufficiently accurate costate estimates into the analytical PMP formulas yields a near-optimal policy $\widehat{\pi}$ with an error bound of $O(\varepsilon+\delta_{\mathrm{BPTT}})$ (Theorem 1).
+The reason this formal justification is crucial lies in a fundamental challenge facing many simulation-based optimization methods.
+A critical issue in direct policy optimization, including standard reinforcement learning, is that a learned policy can achieve a near-optimal objective value (a small value gap) while being structurally very different from the true optimal policy (a large policy gap).
+This occurs because the objective function is often locally “flat” near the optimum, meaning first-order policy deviations may only cause second-order utility losses.
+Our framework circumvents this issue by shifting focus from the value itself to the satisfaction of Pontryagin’s first-order conditions (FOCs).
+The success of this approach hinges on a key property of the HJB equations common in finance: their parabolic nature.
+Parabolic PDEs, akin to the heat equation, enforce a “smoothing” effect, ensuring that a well-behaved value function also has well-behaved derivatives (costates).
+This is not merely an intuitive notion; it is mathematically formalized by regularity theory in Sobolev spaces, which guarantees that convergence in value implies convergence of its derivatives.
+This regularity is precisely what allows us to trust the BPTT-derived costates and reconstruct a robust policy via the PMP formula.
+Building on this expected regularity, we provide a more formal justification for P-PGDPO.
+We aim to bound the policy gap, $\|\widehat{\pi}-\pi^{*}\|$, which is the error between our P-PGDPO policy $\widehat{\pi}$ and the true optimal policy $\pi^{*}$.
+The bound depends on two primary sources of error.
+The first is the FOC residual, denoted by $\varepsilon$, which measures how much the policy from the warm-up stage violates the optimality conditions of PMP.
+The second is the BPTT estimation error, $\delta_{\mathrm{BPTT}}$, in the costate estimate $\widehat{\lambda}$.
+This estimation error itself stems from two numerical sources: the time discretization of the SDEs (with time step $\Delta t$) and the Monte Carlo sampling used for the expectation (with batch size $M$).
+The following theorem encapsulates this relationship, with full technical details provided in Appendix C.
+{restatable}
+[Policy Gap Bound]theoremthm:policy_gap Let $\pi^{*}$ be the true optimal policy and $\widehat{\pi}$ be the policy generated by the P-PGDPO algorithm using a time step $\Delta t$ and batch size $M$.
+Under the regularity conditions specified in Assumption 6 in the Appendix, the gap between the P-PGDPO policy and the true optimum is bounded as follows:
+$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\text{tot}}\left(\varepsilon+\kappa_{1}\Delta t+\frac{\kappa_{2}}{\sqrt{M}}\right)$
+where:
+  • $\varepsilon$ is the $L^{q,p}$-norm of the Pontryagin FOC residual from the warm-up policy.
+  • $\kappa_{1},\kappa_{2}$ are positive constants governing the BPTT estimation error, which arises from time discretization (error proportional to $\Delta t$) and Monte Carlo sampling (error proportional to $1/\sqrt{M}$).
+  • $C_{\text{tot}}$ is a positive constant that depends on the model parameters but not on $\varepsilon$, $\Delta t$, or $M$.
+The proof of Theorem 1, detailed in Appendix C, quantifies the intuition discussed above.
+The logical chain proceeds as follows: a small FOC-gap ($\varepsilon$) ensures that the policy being evaluated is “close” to satisfying the PDE’s structural requirements.
+Due to the smoothing property of the uniformly parabolic PDE—guaranteed by regularity theory in Sobolev spaces (e.g., Dong and Kim, (2009))—this “closeness” is not confined to the value function alone but propagates to its derivatives.
+With the gradients (costates) being verifiably close to the optimal ones, the final step of projecting them through the PMP map becomes a well-posed, stable operation, leading to a policy that is also close to the true optimum.
+Our formal analysis demonstrates that under suitable regularity conditions, particularly uniform parabolicity, a small Pontryagin FOC gap ($\varepsilon$) from the warm-up stage guarantees that the P-PGDPO policy $\widehat{\pi}$ is provably close to the true optimum $\pi^{*}$, up to numerical errors from simulation.
+The bound explicitly shows that to achieve a highly accurate policy, one must not only ensure the warm-up is effective (small $\varepsilon$) but also use a sufficiently small time step ($\Delta t$) and a large batch size ($M$) to control the BPTT estimation error.
+When the objective function is relatively flat near the optimum, the FOC residual $\varepsilon$ for the warm-up policy tends to be small, making the numerical error terms the primary drivers of the final policy gap.
+This provides a rigorous justification for the core idea of the P-PGDPO method: substituting these numerically-controlled and rapidly stabilizing costate estimates into the analytical PMP formulas yields a near-optimal policy.
 However, extending this rigorous proof framework to settings beyond uniform parabolicity, such as hypo-elliptic (H-class, (Hörmander, , 1967)) or boundary-degenerate (D-class, (Oleinik, , 2012)) models, or to general non-affine systems, faces significant analytical challenges, as key assumptions (like global $L^{p}$ regularity or the costate floor) may no longer hold.
-Despite these formal hurdles, the fundamental algorithmic rationale underpinning the P-PGDPO approach remains persuasive for broader applicability.
-Neither Pontryagin’s Maximum Principle (PMP) (Pontryagin, , 2014; Yong and Zhou, , 2012) nor backpropagation-through-time (BPTT) (Goodfellow, , 2016) are inherently limited to uniformly parabolic or affine settings.
-The P-PGDPO strategy leverages the potentially advantageous decoupling of the problem into: (i) estimating potentially simpler or faster-converging costate functions via BPTT, and (ii) utilizing the robust analytical structure of PMP to construct the control.
-This rationale, strongly supported by the numerical success demonstrated even in high dimensions for the affine benchmark case (Section 4), suggests that PG-DPO (especially the P-PGDPO variant) holds promise as a practical tool for a wider class of challenging continuous-time control problems.
-Nonetheless, confirming its effectiveness and potentially refining the theoretical guarantees for models outside the strictly regular class analyzed here remains an important avenue for future research, likely requiring tailored analysis or different analytical techniques (e.g., focusing on regions predominantly visited by the state process).
+Despite these formal hurdles, the fundamental algorithmic rationale of P-PGDPO remains powerfully persuasive for broader applicability.
+The core innovation is the decoupling of costate estimation from direct policy construction.
+Standard end-to-end methods often conflate these two tasks, making it difficult to learn the policy’s fine structure in a ‘flat’ optimization landscape.
+P-PGDPO, by contrast, leverages BPTT for what it excels at—efficient gradient (costate) estimation—and then applies the robust, analytical structure of PMP to deterministically map these estimates to a high-fidelity policy.
+This principle, strongly supported by our numerical results, suggests that P-PGDPO is a practical and powerful tool for a wider class of challenging continuous-time control problems, even in cases where the conditions required for the proof may not strictly hold.
 4 Numerical Experiments
-In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework (both Baseline and Projected PG-DPO (P-PGDPO) variants).
-We focus on the multi-asset, multi-factor Ornstein-Uhlenbeck (OU) portfolio optimization problem introduced in Section 2.3.2.
-The objective is to maximize the expected utility of terminal wealth, $J=\mathbb{E}[U(X_{T})]$, using a CRRA utility function with risk aversion $\gamma=2.0$.
-For these experiments, we specialize the general model by assuming no intermediate consumption (i.e., $C_{t}\equiv 0$) and focus solely on the utility from terminal wealth (effectively setting the bequest weight $\kappa=1$ if $U(X_{T})$ is the direct objective, or adjusting $\kappa$ appropriately if a specific HARA utility form with $a\neq 0$ was being precisely mapped).
-This specific affine model setting allows for comparison against a semi-analytical benchmark solution derived from the HJB equation (details in Appendix B.2).
-Furthermore, to provide a comprehensive assessment against other contemporary deep learning techniques for high-dimensional control problems, we also compare our PG-DPO framework with the Deep BSDE method (Han et al., , 2018; Weinan, , 2017).
-The Deep BSDE approach reformulates the stochastic control problem in terms of a system of backward stochastic differential equations, approximating the unknown components of the solution (typically related to the value function gradient or control variables) using neural networks.
-These networks are trained by minimizing the error incurred in satisfying the BSDE dynamics and terminal conditions through Monte Carlo simulation.
-We compare the policies generated by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and the Deep BSDE method against the HJB-based benchmark using the Root Mean Squared Error (RMSE) metric across various dimensions, scaling up to n=50 assets and k=10 state factors.
-For complete details on the specific parameter values, the implementations of PG-DPO and Deep BSDE (including network architectures and training setups), benchmark computation, and evaluation procedures, please refer to Appendix D.
-The subsequent subsections present and analyze the results of these comparative experiments.
+In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework, including both the baseline and Projected PG-DPO (P-PGDPO) variants.
+The experiments are based on the multi-asset, multi-factor Ornstein–Uhlenbeck (OU) portfolio optimization problem introduced in Section 2.3.2.
+For this setting, we specialize the general model by assuming no intermediate consumption (i.e., $C_{t}=0$) and focus exclusively on utility from terminal wealth, with $\kappa=1$, so that the objective becomes $J=\mathbb{E}[U(X_{T})]$.
+This affine setting admits a semi-analytical benchmark solution, derived from the associated HJB equation, which we use as a reference for evaluation (see Appendix B.2).
+To further assess the effectiveness of our approach, we compare PG-DPO with a state-of-the-art deep learning method for high-dimensional control problems—the Deep BSDE framework (Han et al., , 2018; Weinan, , 2017).
+The Deep BSDE method reformulates the stochastic control problem as a system of backward stochastic differential equations (BSDEs).
+Neural networks are trained to approximate key solution components, typically related to the value function and its gradient.
+These approximations are then used to construct the control policy, often by enforcing or approximating first-order optimality conditions derived from the Hamilton–Jacobi–Bellman equation.
+The networks are trained by minimizing residuals in the BSDE dynamics and terminal condition using Monte Carlo simulation.
+We evaluate the control policies produced by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and the Deep BSDE method (when used for control) against the semi-analytical benchmark.
+The evaluation metric is Root Mean Squared Error (RMSE) in portfolio weights, assessed across a range of problem sizes, including up to $n=50$ assets and $k=10$ state factors.
+Complete details on model parameters, PG-DPO and Deep BSDE implementations (including neural network architectures and training protocols), benchmark computation, and evaluation methodology are provided in Appendix D.
+The following subsections present and analyze the results of these comparative experiments.
 4.1 Convergence Speed and Accuracy Analysis
@@
 (c) $n=50,\;k=10$
-Figure 1: Average policy RMSE comparison between Baseline PG-DPO (e.g., blue), Projected PG-DPO (P-PGDPO) (e.g., orange), and Deep BSDE (e.g., green) relative to the benchmark solution across different asset dimensions ($n$) with $k=10$ factors fixed.
-Values averaged over $k$ evaluation ‘slices.’ For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation or visualization) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$.
+Figure 1: Average policy RMSE comparison between Baseline PG-DPO (e.g., blue), Projected PG-DPO (P-PGDPO) (e.g., orange), and Deep BSDE (e.g., green) relative to the benchmark solution across different asset dimensions ($n$) with $k=10$ factors.
+Values averaged over $k=10$ evaluation ‘slices.’ For each slice $j=1,\dots,10$, the evaluation (e.g., for RMSE calculation or visualization) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$.
 Further details on the evaluation methodology are provided in Appendix D.
 Note the logarithmic scale on the $y$-axis.
-This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO framework (Baseline and Projected PG-DPO (P-PGDPO) variants) against the Deep BSDE method.
-We analyze their performance across various dimensional settings, focusing on how the number of assets ($n$) and the number of state factors ($k$) impact accuracy.
-Figure 1 (Note: Assuming you will update the figure label in your main document if it includes Deep BSDE results) displays the average Root Mean Squared Error (RMSE) between the learned/derived policies and the benchmark optimal policy.
-The RMSE is plotted against training epochs on a logarithmic scale for configurations with $k=10$ state variables and varying numbers of assets ($n=1,10,50$).
-The RMSE for P-PGDPO represents the error of the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ (Eq.
-(11)), which combines myopic and hedging components.
-As detailed in Section 4.2, the myopic component’s error is typically negligible ($O(10^{-7})$).
-Thus, the total RMSE for P-PGDPO is almost entirely driven by the accuracy of estimating the hedging component using BPTT-derived costate derivatives, $\partial_{\mathbf{Y}}\lambda_{t}^{*}$.
-Table 2 summarizes the minimum achieved policy RMSE and the epochs at which these minima were observed for the $k=10$ scenario with varying $n$.
-A consistent observation is the superior performance of the P-PGDPO method, which achieves significantly lower RMSE values compared to both Baseline PG-DPO and Deep BSDE across all tested asset dimensions.
-For instance, with $n=1,k=10$, P-PGDPO’s minimum RMSE is approximately an order of magnitude better than Deep BSDE and two orders of magnitude better than Baseline PG-DPO.
-Deep BSDE generally performs between the Baseline and P-PGDPO methods.
-While P-PGDPO maintains its lead as $n$ increases to 50, the performance gap between P-PGDPO and Deep BSDE tends to narrow, suggesting that the increasing dimensionality of $n$ poses challenges even for the PMP-based analytical solution, possibly due to the ill-conditioning of the covariance matrix $\Sigma$.
-All methods show an increase in RMSE as $n$ grows, indicating the inherent difficulty of high-dimensional asset allocation.
+This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO framework—including both the Baseline and Projected PG-DPO (P-PGDPO) variants—against the Deep BSDE method, when applied to derive control policies.
+We evaluate performance across various problem dimensions, focusing on how the number of assets ($n$) and the number of state factors ($k$) affect policy accuracy.
+Figure 1 shows the average Root Mean Squared Error (RMSE) between the learned or derived policies and the semi-analytical benchmark policy.
+The RMSE is plotted against the number of training iterations on a logarithmic scale for experiments with $k=10$ state variables and varying asset counts ($n=1,10,50$).
+For P-PGDPO, the reported RMSE corresponds to the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ (Eq.
+(11)), which combines the myopic and intertemporal hedging components.
+As discussed in Section 4.2, the myopic error is typically negligible—on the order of $O(10^{-7})$.
+Thus, the total RMSE for P-PGDPO is effectively governed by the accuracy of the hedging component, which depends on estimating the costate gradient $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ via BPTT.
+Table 2: Summary of Minimum Policy RMSE and Convergence Iterations
+Dimension Method Min. Policy RMSE Iterations at Min.
+($n$ assets, $k=10$ factors)
+n=1, k=10 Baseline PG-DPO $1.751\times 10^{-2}$ 2800
+P-PGDPO $4.710\times 10^{-4}$ 1200
+Deep BSDE $8.146\times 10^{-2}$ 1
+n=10, k=10 Baseline PG-DPO $4.996\times 10^{-2}$ 9800
+P-PGDPO $2.799\times 10^{-3}$ 4800
+Deep BSDE $7.623\times 10^{-2}$ 200
+n=50, k=10 Baseline PG-DPO $1.070\times 10^{-1}$ 8800
+P-PGDPO $1.522\times 10^{-2}$ 10000
+Deep BSDE $8.199\times 10^{-2}$ 1200
+Note: Minimum policy RMSE values are averaged across $k$ evaluation slices.
+For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation) focuses on state factor $Y_{j}$, while all other factors $Y_{\ell}$, for $\ell\neq j$, are fixed at their respective long-run means $\theta_{Y,\ell}$.
+Further details on the evaluation procedure are provided in Appendix D.
+The iteration numbers indicate when the minimum average RMSE was observed; evaluations and logging were performed every 200 iterations.
+Table 2 summarizes the minimum achieved policy RMSE values and the corresponding iteration counts at which these minima occur, across $k=10$ evaluation ‘slcies’ with varying numbers of assets $n$.
+In each evaluation slice $j=1,\dots,10$, the RMSE and visualization are computed with respect to state factor $Y_{j}$, while all other factors $Y_{\ell}$, for $\ell\neq j$, are fixed at their respective long-term means $\theta_{Y,\ell}$.
+Further details on the evaluation methodology are provided in Appendix D.
+A consistent and notable finding is the superior performance of the P-PGDPO method.
+It achieves substantially lower RMSEs than both Baseline PG-DPO and Deep BSDE across all tested asset dimensions.
+For example, in the $n=1$ case, P-PGDPO attains a minimum RMSE of $4.710\times 10^{-4}$, significantly outperforming Baseline PG-DPO ($1.751\times 10^{-2}$) and Deep BSDE ($8.146\times 10^{-2}$).
+This result highlights the effectiveness of P-PGDPO in delivering high-accuracy control policies.
+The comparison between Baseline PG-DPO and Deep BSDE is more nuanced.
+Deep BSDE often reaches its minimum RMSE in fewer iterations—e.g., 1 iteration for $n=1$, and 200 iterations for $n=10$—compared to 2800 and 9800 iterations, respectively, for Baseline PG-DPO (see Table 2).
+This suggests faster initial convergence for Deep BSDE.
+However, its final RMSE is not consistently lower than that of Baseline PG-DPO.
+For both $n=1$ and $n=10$, Baseline PG-DPO ultimately achieves better accuracy: $1.751\times 10^{-2}$ vs.
+$8.146\times 10^{-2}$, and $4.996\times 10^{-2}$ vs.
+$7.623\times 10^{-2}$, respectively.
+In the high-dimensional case of $n=50$, however, Deep BSDE achieves a lower final RMSE ($8.199\times 10^{-2}$) than Baseline PG-DPO ($1.070\times 10^{-1}$).
+As illustrated in Figure 1, after early convergence, Deep BSDE’s RMSE may plateau or exhibit greater variance, while Baseline PG-DPO often demonstrates more gradual but continued improvement in certain configurations.
+P-PGDPO maintains a significant accuracy advantage even as $n$ increases to 50.
+However, all methods exhibit rising RMSE with increasing $n$ (see Table 2), reflecting the inherent challenges of high-dimensional asset allocation—potentially exacerbated by issues such as the ill-conditioning of the covariance matrix $\Sigma$, as discussed later.
 To further investigate the distinct effects of asset dimensionality ($n$) versus factor dimensionality ($k$), we conducted two additional sets of experiments.
-Table 2: Summary of Minimum Policy RMSE and Convergence Epochs
-Dimension Method Min. Policy RMSE Epochs at Min.
-($n$ assets, $k=10$ factors)
-n=1, k=10 Baseline PG-DPO $1.75\times 10^{-2}$ 2800
-P-PGDPO $4.72\times 10^{-4}$ 1200
-Deep BSDE $4.78\times 10^{-3}$ 600
-n=10, k=10 Baseline PG-DPO $5.11\times 10^{-2}$ 9600
-P-PGDPO $5.26\times 10^{-3}$ 1000
-Deep BSDE $3.46\times 10^{-2}$ 1400
-n=50, k=10 Baseline PG-DPO $1.05\times 10^{-1}$ 8800
-P-PGDPO $1.58\times 10^{-2}$ 5400
-Deep BSDE $4.90\times 10^{-2}$ 9200
-Note: Minimum Policy RMSE values are averaged across $k$ evaluation ’slices.’ For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$.
-Further details on the evaluation methodology are provided in Appendix D.
-Epochs indicate the recorded iteration number when the minimum average RMSE was observed; evaluations and logging were performed every 200 iterations.
-First, Table 3 presents results where the number of assets is fixed at $n=1$, and the number of state factors $k$ is varied ($k=1,5,10$).
-In this scenario, the performance of all three methodologies remains relatively stable across different values of $k$.
-For P-PGDPO, the RMSE even shows a slight improvement as $k$ increases, possibly due to more information being available for costate estimation.
-Deep BSDE and Baseline PG-DPO also exhibit no significant degradation in performance.
-This suggests that, for a small number of assets, the ”curse of dimensionality” associated with an increasing number of state factors ($k$) is largely managed by these neural network-based approaches.
-The underlying PMP structure or BSDE formulation appears robust to the increase in $k$ when $n$ is small and fixed, with Monte Carlo sampling and function approximation being the primary challenges rather than ill-conditioned linear algebra.
-Table 3: Summary of Minimum Policy RMSE and Convergence Epochs for $n=1$ (Varying $k$)
-$k$ factors Method Min. Policy RMSE Epochs at Min.
+First, Table 3 reports results for the case where the number of assets is fixed at $n=1$, and the number of state factors is varied ($k=1,5,10$).
+In this setting, P-PGDPO consistently achieves the lowest RMSE across all configurations.
+Specifically, its RMSE is $1.120\times 10^{-3}$ for $k=1$, improves to $2.500\times 10^{-4}$ for $k=5$, and is $4.710\times 10^{-4}$ for $k=10$.
+While not strictly monotonic, P-PGDPO demonstrates strong performance across all $k$, with the best result observed at $k=5$.
+The performances of Baseline PG-DPO and Deep BSDE are more varied. For $n=1$:
+  • At $k=1$: Deep BSDE ($9.350\times 10^{-3}$) outperforms Baseline PG-DPO ($3.260\times 10^{-2}$).
+  • At $k=5$: Deep BSDE ($3.094\times 10^{-2}$) outperforms Baseline PG-DPO ($3.714\times 10^{-2}$).
+  • At $k=10$: Baseline PG-DPO ($1.751\times 10^{-2}$) outperforms Deep BSDE ($8.146\times 10^{-2}$); however, Deep BSDE converges to its minimum in just one iteration.
+These results suggest that, when the number of assets is small, the “curse of dimensionality” associated with increasing $k$ is largely mitigated by neural network–based approaches.
+Nevertheless, their relative performance can shift.
+Both the PMP and BSDE formulations appear robust to increases in factor dimensionality when $n$ is fixed and small.
+In this case, Monte Carlo sampling error and function approximation quality become the primary computational bottlenecks, rather than issues related to ill-conditioned linear algebra.
+Second, Table 4 reports results where the number of state factors is fixed at $k=2$, and the number of assets $n$ is varied ($n=1,10,50$).
+In stark contrast to the previous case, increasing $n$ leads to a substantial deterioration in performance across all three methods.
+Baseline PG-DPO shows the most pronounced increase in RMSE, rising from $4.770\times 10^{-2}$ for $n=1$ to $3.875\times 10^{-1}$ for $n=50$.
+While P-PGDPO still achieves the lowest RMSE in all cases, its error also increases with $n$—from $1.990\times 10^{-4}$ to $7.330\times 10^{-2}$.
+Deep BSDE’s performance degrades most dramatically, with RMSE rising from $2.675\times 10^{-3}$ at $n=1$ to $6.340\times 10^{-1}$ at $n=50$.
+Notably, in the $n=50,k=2$ case, Baseline PG-DPO ($3.875\times 10^{-1}$) outperforms Deep BSDE ($6.340\times 10^{-1}$).
+This pronounced degradation with increasing $n$, even for small $k$, points to a distinct source of difficulty—one not solely attributable to function approximation in high-dimensional state spaces.
+A likely culprit is the numerical instability introduced by the asset covariance matrix $\Sigma$.
+As $n$ increases, the $n\times n$ matrix $\Sigma$ becomes more prone to ill-conditioning (i.e., a high condition number).
+Since optimal portfolio policies in the PMP framework (Eq.
+(11)) involve computing $\Sigma^{-1}$, small estimation errors in the costates or model parameters can be significantly amplified, resulting in large policy errors.
+This observation suggests that the “curse” associated with large $n$ is not merely about learning a higher-dimensional function, but also reflects the increasing fragility of the optimization problem itself.
+The structure of the problem becomes numerically unstable, which adversely affects any method that explicitly or implicitly depends on $\Sigma$ or its inverse.
+These findings highlight that while neural network–based methods have made substantial progress in managing high-dimensional state spaces (associated with large $k$), challenges related to a large number of decision variables (i.e., assets $n$)—particularly in the presence of ill-conditioned covariance matrices—remain a significant concern.
+Addressing this “asset-dimension curse” may require incorporating techniques such as covariance matrix shrinkage, spectral regularization, or other forms of preconditioning within the optimization problem.
+Table 3: Summary of Minimum Policy RMSE and Convergence Iterations for $n=1$ (Varying $k$)
+$k$ factors Method Min. Policy RMSE Iterations at Min.
 ($n=1$ asset)
-1 Baseline PG-DPO $3.26\times 10^{-2}$ 1000
-P-PGDPO $1.20\times 10^{-3}$ 800
-Deep BSDE $7.39\times 10^{-3}$ 800
-5 Baseline PG-DPO $3.71\times 10^{-2}$ 1400
-P-PGDPO $2.64\times 10^{-4}$ 800
-Deep BSDE $1.02\times 10^{-2}$ 1400
-10 Baseline PG-DPO $1.75\times 10^{-2}$ 2800
-P-PGDPO $4.72\times 10^{-4}$ 1200
-Deep BSDE $4.78\times 10^{-3}$ 600
-Table 4: Summary of Minimum Policy RMSE and Convergence Epochs for $k=2$ (Varying $n$)
-$n$ assets Method Min. Policy RMSE Epochs at Min.
+1 Baseline PG-DPO $3.260\times 10^{-2}$ 1000
+P-PGDPO $1.120\times 10^{-3}$ 800
+Deep BSDE $9.350\times 10^{-3}$ 3600
+5 Baseline PG-DPO $3.714\times 10^{-2}$ 1400
+P-PGDPO $2.500\times 10^{-4}$ 800
+Deep BSDE $3.094\times 10^{-2}$ 9400
+10 Baseline PG-DPO $1.751\times 10^{-2}$ 2800
+P-PGDPO $4.710\times 10^{-4}$ 1200
+Deep BSDE $8.146\times 10^{-2}$ 1
+Table 4: Summary of Minimum Policy RMSE and Convergence Iterations for $k=2$ (Varying $n$)
+$n$ assets Method Min. Policy RMSE Iterations at Min.
 ($k=2$ factors)
-1 Baseline PG-DPO $4.77\times 10^{-2}$ 1200
-P-PGDPO $1.98\times 10^{-4}$ 800
-Deep BSDE $4.63\times 10^{-3}$ 3400
-10 Baseline PG-DPO $1.12\times 10^{-1}$ 4000
-P-PGDPO $1.22\times 10^{-2}$ 1400
-Deep BSDE $2.32\times 10^{-2}$ 2600
-50 Baseline PG-DPO $2.24\times 10^{-1}$ 7200
-P-PGDPO $5.55\times 10^{-2}$ 4400
-Deep BSDE $1.88\times 10^{-1}$ 4400
-Second, Table 4 shows results where the number of state factors is fixed at $k=2$, and the number of assets $n$ is varied ($n=1,10,50$).
-In stark contrast to the previous case, increasing $n$ leads to a significant deterioration in performance for all three methods.
-Baseline PG-DPO shows the most substantial increase in RMSE.
-While P-PGDPO still yields the lowest RMSE, its error also increases considerably with $n$.
-Deep BSDE’s performance also degrades, approaching that of Baseline PG-DPO for $n=50$.
-This pronounced degradation with increasing $n$ (even for a small $k$) points towards a different kind of challenge not solely related to the dimensionality of the state space for function approximation.
-A primary suspect is the numerical instability arising from the asset covariance matrix $\Sigma$.
-As $n$ increases, $\Sigma$ ($n\times n$) is more likely to become ill-conditioned (i.e., have a high condition number).
-The optimal policy calculation, particularly in the PMP framework (Eq.
-(11)), involves the inverse $\Sigma^{-1}$.
-If $\Sigma$ is ill-conditioned, any small errors in the estimated costates or market parameters can be greatly amplified by $\Sigma^{-1}$, leading to large errors in the computed policy.
-This suggests that the “curse” associated with large $n$ is not just about learning a higher-dimensional policy function but also about the increasing fragility of the underlying mathematical problem structure due to the properties of $\Sigma$.
-The problem itself becomes “damaged” or numerically challenging, impacting any method that relies on or implicitly reflects this structure.
-These findings underscore that while neural network approaches have made significant strides in handling high-dimensional state spaces (related to $k$), the challenges posed by a large number of decision variables (assets $n$), especially when intertwined with ill-conditioned covariance matrices, require further attention.
-Strategies such as covariance matrix shrinkage or other regularization techniques applied prior to or within the optimization might be necessary to effectively address this “asset-dimension curse”.
+1 Baseline PG-DPO $4.770\times 10^{-2}$ 1200
+P-PGDPO $1.990\times 10^{-4}$ 800
+Deep BSDE $2.675\times 10^{-3}$ 3600
+10 Baseline PG-DPO $1.145\times 10^{-1}$ 4000
+P-PGDPO $3.645\times 10^{-3}$ 4600
+Deep BSDE $3.835\times 10^{-2}$ 1
+50 Baseline PG-DPO $3.875\times 10^{-1}$ 5600
+P-PGDPO $7.330\times 10^{-2}$ 9800
+Deep BSDE $6.340\times 10^{-1}$ 3800
 4.2 High-Dimensional Performance Analysis ($n=50,k=10$)
 We now visually examine the performance of the three methodologies in the most challenging setting considered: $n=50$ assets and $k=10$ state factors.
-This high-dimensional scenario, where traditional DP methods are intractable, serves to highlight the distinct capabilities and limitations of each approach.
+This high-dimensional configuration, where traditional dynamic programming methods are computationally intractable, highlights the distinct capabilities and limitations of each approach.
 As a reference, Figure 2 displays the benchmark analytical solution for the first asset ($\pi_{1}^{\text{sol}}$), decomposed into its myopic ($\pi_{1,0}^{\text{sol}}$) and intertemporal hedging ($\pi_{1,1}^{\text{sol}}$) components.
-The hedging component is clearly non-trivial and exhibits complex dependencies on both the state variable $Y_{1}$ and time-to-maturity, underscoring its importance for optimal allocation.
+The hedging component is clearly non-trivial, exhibiting complex dependencies on both the state variable $Y_{1}$ and the remaining time to maturity.
+This underscores its importance for achieving optimal portfolio allocation in dynamic settings.
 (a) Benchmark Policy ($\pi_{1}^{\text{sol}}$)
@@
 These serve as the ground truth.
-Figure 3 provides a side-by-side comparison of the policies learned by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and Deep BSDE, alongside their respective errors, for the first asset.
-These plots correspond to iterations where each method achieved its (near) optimal performance in the $n=50,k=10$ setting, as indicated in Table 2 (e.g., iteration 5400 for P-PGDPO, which is the nearest logged checkpoint to the observed minimum RMSE at iteration 5399, as evaluations were performed every 200 iterations).
+Figure 3 presents a side-by-side comparison of the policies learned by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and Deep BSDE, along with their corresponding errors, for the first asset.
+These plots correspond to the iterations at which each method achieved its (near) optimal performance in the $n=50,k=10$ setting, as reported in Table 2.
+For instance, the policy shown for P-PGDPO corresponds to iteration 6400, which is the nearest logged evaluation point to the observed minimum RMSE at iteration 5399 (evaluation was performed every 200 iterations, and the global minimum RMSE for P-PGDPO in this setting was ultimately achieved at iteration 10000, as also noted in Table 2).
 (a) Baseline Policy ($\pi_{1,\theta}$)
@@
 Figure 3: Comparison of total policies (left column) and policy errors (right column) for the first asset ($i=1$) across Baseline PG-DPO (top row), Projected PG-DPO (P-PGDPO) (middle row), and Deep BSDE (bottom row) in the $n=50,k=10$ case.
-Results shown at (near) optimal iterations for each method (e.g., iteration 5400 for P-PGDPO), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2).
+Results shown at (near) optimal iterations for each method (e.g., iteration 6400 for P-PGDPO as plotted), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2).
 Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
@@
 Different color bar ranges reflect varying policy/error magnitudes.
-The Baseline PG-DPO method (Figure 3a,b), despite utilizing a standard neural network architecture (detailed in Appendix D.2), struggles significantly.
-While its learned policy exhibits some non-linear structure, it bears little resemblance to the benchmark, and the corresponding error is extremely large.
-This performance visually confirms the poor quantitative results in Table 2 and highlights the difficulty of directly approximating the high-dimensional optimal policy function using a neural network without leveraging more structural information from the problem.
+The Baseline PG-DPO method (Figure 3a,b), employing a standard neural network architecture (see Appendix D.2), performs poorly in this high-dimensional setting.
+Although the learned policy exhibits some non-linear structure, it bears little resemblance to the benchmark solution, and the corresponding error is substantial.
+This outcome visually corroborates the relatively high RMSE of $1.070\times 10^{-1}$ reported for Baseline PG-DPO in the $n=50,k=10$ case (Table 2), and underscores the difficulty of directly approximating a high-dimensional optimal policy function using generic neural architectures without incorporating structural insights from the problem’s analytic formulation.
 (a) P-PGDPO Myopic ($\pi_{1,0}^{\mathrm{PMP}}$)
@@
 (d) Hedging Error
-Figure 4: Decomposition of the Projected PG-DPO (P-PGDPO) policy for the first asset ($i=1$) in the $n=50,k=10$ case (results from iteration 5400), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2).
+Figure 4: Decomposition of the Projected PG-DPO (P-PGDPO) policy for the first asset ($i=1$) in the $n=50,k=10$ case (results from iteration 6400 as plotted), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2).
 Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
@@
 Note different color bar ranges.
-The Deep BSDE method (Figure 3e,f) shows a marked improvement over the Baseline.
+The Deep BSDE method (Figure 3e,f) shows a marked improvement over Baseline PG-DPO.
 Its learned policy qualitatively resembles the myopic component of the benchmark (cf.
-Figure 2b), primarily capturing dependencies on the state variable $Y_{1}$.
-However, the policy largely fails to capture the complex intertemporal hedging demands.
-The error plot for Deep BSDE (Figure 3f) shows significant discrepancies in regions where hedging is important, suggesting that the hedging component is largely absent from the Deep BSDE solution.
-This difficulty in capturing hedging demands can be attributed to several factors.
-The Deep BSDE method primarily focuses on matching the terminal condition and ensuring consistency of the value process $Y_{t}$ and the gradient-related process $Z_{t}$ (approximating first-order derivatives of the value function) through its loss function.
-Intertemporal hedging, however, critically depends on the mixed second-order derivative $V_{x\mathbf{y}}$, which is not explicitly represented or directly targeted by the standard two-network Deep BSDE architecture.
-Furthermore, the loss function penalizes inaccuracies in $Z_{t}$ primarily through a quadratic martingale term.
-Consequently, small errors in estimating the gradient components relevant to $V_{x\mathbf{y}}$ might have a limited impact on the overall training objective but can be significantly magnified when constructing the policy, especially after multiplication by the potentially ill-conditioned inverse covariance matrix $\Sigma^{-1}$ in high-dimensional settings.
-As a result, while the Deep BSDE approach can effectively learn the dominant myopic component of the policy, it may struggle to precisely fit the more nuanced intertemporal hedging patterns.
-In stark contrast, the Projected PG-DPO (P-PGDPO) method (Figure 3c,d) delivers a policy that is visually almost identical to the benchmark total policy (Figure 2a).
-The corresponding error is substantially smaller than that of the other two methods.
-This superior performance underscores the efficacy of its decoupled learning structure: first estimating costates via BPTT and stabilizing them through Monte Carlo averaging, then analytically constructing the policy using the PMP FOCs.
-The remarkable accuracy of P-PGDPO is further elucidated by decomposing its policy into myopic and hedging components, as shown in Figure 4 (results from iteration 5400).
-The myopic component (Figure 4(a)) is virtually identical to the benchmark myopic component, with its error (Figure 4(c)) being negligible ($O(10^{-7})$).
-The hedging component derived by P-PGDPO (Figure 4(b)) also successfully captures the complex shape of the benchmark hedging demand (Figure 2c).
-The remaining error in the hedging component (Figure 4(d)) dictates the overall error of the P-PGDPO policy.
-This residual hedging error is likely not due to a fundamental flaw in the P-PGDPO algorithm itself but rather a consequence of the numerical challenges inherent in high-dimensional asset allocation.
-Specifically, the hedging term in the PMP FOC involves $\Sigma^{-1}$:
-$\pi^{\text{hedge}}=-\frac{1}{X_{t}V_{xx}}\,\underbrace{\Sigma^{-1}}_{\text{cond num can be large for }n\uparrow}(\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y})\,V_{x\mathbf{Y}}.$
-For $n=50$, the $50\times 50$ covariance matrix $\Sigma$ can easily become ill-conditioned.
-Even if the BPTT-derived estimate for $V_{x\mathbf{Y}}$ is quite accurate, any small residual noise can be significantly amplified when multiplied by $\Sigma^{-1}$ if its condition number is large.
-This amplification effect is the primary suspect for the observed hedging error.
-In conclusion, the $n=50,k=10$ visualizations powerfully demonstrate that Projected PG-DPO (P-PGDPO) can effectively ”break the dimensional barrier” where Baseline PG-DPO fails and Deep BSDE offers only a partial (myopic-focused) solution.
-The remaining inaccuracies in the hedging component point towards the need for addressing the ill-conditioning of the covariance matrix in very high asset dimensions.
+Figure 2b), primarily capturing the dependency on the state variable $Y_{1}$.
+However, the policy fails to capture the more complex intertemporal hedging demands.
+The corresponding error plot (Figure 3f) reveals significant discrepancies in regions where hedging plays a critical role, suggesting that the hedging component is largely absent from the Deep BSDE solution.
+This shortcoming can be attributed to both architectural and training limitations.
+The Deep BSDE framework is primarily designed to match the terminal condition and ensure consistency of the forward value process $Y_{t}$ (approximating the value function) and the backward process $Z_{t}$ (approximating its gradient).
+However, intertemporal hedging depends crucially on the mixed second-order derivative $V_{x\mathbf{y}}$, which is neither explicitly represented nor directly supervised in the standard two-network Deep BSDE setup.
+Moreover, the loss function penalizes deviations in $Z_{t}$ through a quadratic martingale residual, which may assign relatively low weight to errors in specific gradient directions, particularly those influencing $V_{x\mathbf{y}}$.
+These small gradient estimation errors can nonetheless be magnified when constructing the control policy, especially through multiplication by the potentially ill-conditioned inverse covariance matrix $\Sigma^{-1}$ in high-dimensional settings.
+As a result, while the Deep BSDE method effectively learns the dominant myopic component of the policy, it tends to underperform in capturing the subtler hedging demands essential for dynamic portfolio optimization.
+In stark contrast, the Projected PG-DPO (P-PGDPO) method (Figure 3c,d) produces a policy that is visually nearly indistinguishable from the benchmark total policy (Figure 2a).
+The corresponding error is substantially smaller than that of the other two methods, consistent with its superior RMSE of $1.522\times 10^{-2}$ for the $n=50,k=10$ case (Table 2).
+This performance underscores the effectiveness of P-PGDPO’s decoupled learning structure: first estimating costates via BPTT and stabilizing them through Monte Carlo averaging, then analytically constructing the policy using the PMP first-order conditions.
+The exceptional accuracy of P-PGDPO is further illuminated by decomposing its policy into myopic and hedging components, as shown in Figure 4 (from iteration 6400).
+The myopic component (Figure 4(a)) closely matches the benchmark, with its error (Figure 4(c)) being negligible—on the order of $O(10^{-7})$.
+The hedging component (Figure 4(b)) also successfully captures the complex shape of the benchmark hedging demand (Figure 2c).
+The residual error (Figure 4(d)) in this component dominates the total policy error.
+Notably, this hedging error likely does not stem from a flaw in the P-PGDPO algorithm itself, but rather reflects the numerical challenges inherent to high-dimensional asset allocation.
+Specifically, the hedging term in the PMP solution involves the inverse of the asset return covariance matrix $\Sigma$:
+$\pi^{\text{hedge}}=-\frac{1}{X_{t}V_{xx}}\,\underbrace{\Sigma^{-1}}_{\text{canbe ill-conditioned for large }n}(\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y})\,V_{x\mathbf{Y}}.$
+For $n=50$, the $50\times 50$ matrix $\Sigma$ is prone to ill-conditioning.
+Thus, even small residual noise in the BPTT-estimated $V_{x\mathbf{Y}}$ may be significantly amplified by $\Sigma^{-1}$, especially when its condition number is large.
+This amplification effect is the most likely cause of the observed hedging error.
+In conclusion, the $n=50,k=10$ visualizations clearly demonstrate that Projected PG-DPO is capable of “breaking the dimensional barrier,” where Baseline PG-DPO fails and Deep BSDE provides only a partial (myopic-focused) solution.
+The remaining inaccuracies in the hedging component point to a pressing need for addressing the ill-conditioning of $\Sigma$ in very high asset dimensions—potentially through regularization, shrinkage techniques, or architecture-aware preconditioning.
+4.3 Non-affine Asset Return Dynamics
+In this section, we extend our numerical investigation to a setting where the asset prices evolve according to a non-affine process.
+We assume that the state factors state factors, $\mathbf{Y}_{t}=[Y_{1,t},\dots,Y_{k,t}]^{T}$, follow the same multivariate Ornstein-Uhlenbeck (OU) process as described in Section 2.1.
+The expected returns of risky assets, $\bm{\mu}_{S}(t,\mathbf{Y}_{t})$, are given by
+$\bm{\mu}_{S}(t,\mathbf{Y}_{t})=\text{diag}(\bm{\sigma})(\bm{\alpha}\mathbf{Y}_{\beta,t})+r\mathbf{1}=\text{diag}(\bm{\sigma})\left(\bm{\alpha}(\mathbf{Y}_{t}+\bm{\beta}\mathbf{Y}_{t}^{2})\right)+r\mathbf{1},$
+where $\bm{\alpha}$ is a $d\times k$ matrix of factor loadings and the multiplication and addition in $\bm{\alpha}(\mathbf{Y}_{t}+\bm{\beta}\mathbf{Y}_{t}^{2})$ is understood as element-wise operations.
+The vector $\bm{\beta}$ controls the strength and direction of the non-linear impact of $\mathbf{Y}_{t}$ on asset returns.
+Its overall magnitude is parameterized by $\beta_{\text{norm}}=||\bm{\beta}||_{2}$, and for our experiments, $\bm{\beta}$ is generated as $\bm{\beta}=\beta_{\text{norm}}\cdot\mathbf{u}$, where $\mathbf{u}$ is a randomly generated unit vector (with a fixed seed for consistency).
+When $\beta_{\text{norm}}=0$, $\bm{\beta}=\mathbf{0}$, which implies $\mathbf{Y}_{\beta,t}=\mathbf{Y}_{t}$.
+In this case, the model for $\bm{\mu}_{S}$ becomes affine in $\mathbf{Y}_{t}$, and the overall portfolio optimization problem reduces to the affine setting analyzed in Sections 4.1 and 4.2.
+The primary objective of this analysis is twofold: first, to validate the solutions obtained by our Projected PG-DPO (P-PGDPO) framework for this model by examining their behavior as $\beta_{\text{norm}}\to 0$, where we expect convergence to the known affine solution.
+Second, we investigate how the optimal policy systematically changes as the strength of this non-linear component in the risk premium, dictated by $\beta_{\text{norm}}$, increases.
+All experiments in this subsection are conducted for the most challenging setting, $n=50$ assets and $k=10$ state factors, utilizing the P-PGDPO method.
+To quantify the deviation of the P-PGDPO policy in this non-affine setting from the affine benchmark solution as $\beta_{\text{norm}}$ changes, Table 5 presents the Root Mean Squared Error (RMSE) of the P-PGDPO learned policy, $\bm{\pi}(\beta_{\text{norm}})$, relative to the affine analytical solution ($\bm{\pi}(\beta_{\text{norm}}=0)$).
+For the results presented in this table, all P-PGDPO models were trained for 10,000 iterations.
+In the non-affine case ($\beta_{\text{norm}}>0$), where an analytical solution to the control problem is unavailable, this fixed number of iterations is used for evaluation, as opposed to selecting the number of iterations based on the minimum error against a known benchmark in earlier sections.
+Table 5: Policy RMSE of P-PGDPO for the model with non-affine risk premia vs. affine benchmark ($n=50,k=10$). Results reported after 10,000 training iterations.
+$\beta_{\text{norm}}$ Min. Policy RMSE (P-PGDPO)
+0.0 0.04701
+0.1 0.05686
+0.5 0.06755
+1.0 0.11469
+2.0 0.20120
+3.0 0.27530
+4.0 0.33570
+The results in Table 5 clearly demonstrate two key points.
+First, for $\beta_{\text{norm}}=0.0$, the P-PGDPO method achieves a low RMSE of $0.04701$.
+This is consistent with its strong performance in the affine setting (reported in Section 4.1) and indicates an accurate approximation of the affine analytical solution, as expected when the non-linear component vanishes.
+Second, as $\beta_{\text{norm}}$ increases from $0.1$ to $4.0$, the difference of the P-PGDPO policy from the fixed affine benchmark, measured by RMSE, monotonically increases.
+This trend signifies that the optimal policy systematically diverges from the affine optimal policy as the magnitude of the non-linear component grows.
+The consistent increase suggests that P-PGDPO is stably adapting the policy to this increasing non-linearity in the market price of risk, rather than producing erratic or unstable solutions.
+(a) Analytical solution ($\beta_{\text{norm}}=0$).
+(b) P-PGDPO policy ($\beta_{\text{norm}}=0.0$).
+(c) P-PGDPO policy ($\beta_{\text{norm}}=0.1$).
+(d) P-PGDPO policy ($\beta_{\text{norm}}=1.0$).
+(e) P-PGDPO policy ($\beta_{\text{norm}}=2.0$).
+Figure 5: Visual comparison of optimal policies for the first asset ($\pi_{1}$) against the first state factor ($Y_{1}$) and time-to-maturity ($T-t$) in the $n=50,k=10$ setting.
+(a) shows the analytical solution for the affine model ($\beta_{\text{norm}}=0$).
+(b)-(e) display P-PGDPO learned policies for models with increasing strength of non-linearity in asset risk premia, corresponding to $\beta_{\text{norm}}$ values of $0.0,0.1,1.0,2.0$, respectively.
+Note the systematic change in policy structure and value ranges (indicated by colorbars) as $\beta_{\text{norm}}$ increases.
+To complement these quantitative findings, Figure 5 provides a visual comparison of the learned policies.
+Figure 5(a) displays the analytical solution for the affine model ($\beta_{\text{norm}}=0$).
+Figures 5(b)–5(e) show the P-PGDPO learned policies for models where $\beta_{\text{norm}}=0.0,0.1,1.0,$ and $2.0$, respectively.
+These plots depict the policy for the first asset ($\pi_{1}$) as a function of the first state factor ($Y_{1}$) and time-to-maturity ($T-t$), holding other state factors at their long-term means.
+As observed in Figure 5(b), the P-PGDPO policy for $\beta_{\text{norm}}=0.0$ closely matches the analytical affine solution shown in Figure 5(a), visually confirming the low RMSE reported in Table 5.
+As $\beta_{\text{norm}}$ increases to $0.1$ (Figure 5(c)), $1.0$ (Figure 5(d)), and $2.0$ (Figure 5(e)), a clear and systematic transformation of the policy surface is evident.
+The policy values tend to increase in certain regions (indicated by the shift towards warmer colors and the expanding range of the colorbars), and the overall shape of the policy function changes.
+This visual progression corroborates the quantitative results presented in Table 5, illustrating that the P-PGDPO framework captures the impact of the non-linear term in the risk premium in a consistent manner.
+Since the non-affine component $\bm{\beta}\,\mathbf{Y}_{t}^{2}$ directly alters how state variables $\mathbf{Y}_{t}$ influence the investment opportunity set (via $\bm{\mu}_{S}$), these changes in the overall policy are likely significantly driven by adjustments to the intertemporal hedging component.
+In summary, the combined quantitative and visual results from this analysis strongly suggest that the P-PGDPO method provides robust and consistent solutions for portfolio optimization problems where the asset prices follow general non-affine dynamics.
+The learned policies correctly converge to the affine solution as the specific non-linear component vanishes ($\beta_{\text{norm}}\to 0$), and they exhibit a systematic and stable response to increasing levels of this non-linearity.
+This provides confidence in applying PG-DPO to more complex financial models with non-affine dynamics where analytical solutions are typically unavailable.
 5 Conclusion
@@
 This decoupling strategy proved remarkably successful, enabling P-PGDPO to tackle problems of previously prohibitive scales.
-For instance, in a challenging $n=50$ asset, $k=10$ factor environment, P-PGDPO achieved a total policy RMSE of $\mathbf{1.58\times 10^{-2}}$ relative to the analytical benchmark, with the myopic component error being negligible ($O(10^{-7})$) and, critically, accurately reconstructing complex hedging terms (Section 4.2, Figure 4).
+For instance, in a challenging $n=50$ asset, $k=10$ factor environment, P-PGDPO achieved a total policy RMSE of $\mathbf{1.52\times 10^{-2}}$ relative to the analytical benchmark, with the myopic component error being negligible ($O(10^{-7})$) and, critically, accurately reconstructing complex hedging terms (Section 4.2, Figure 4).
 The theoretical underpinnings of P-PGDPO are strengthened by our Policy Gap Bound (Theorem 1), which guarantees near-optimal policies given accurate costate estimation and a small FOC gap.
@@
 Its demonstrated ability to recover complex, non-linear hedging strategies in high dimensions suggests its potential applicability to large institutional portfolios, such as those involving numerous ETFs or derivative hedging.
-Moreover, the core principle of P-PGDPO—“estimate costates first, then derive policy”—holds promise for a broader class of problems, where PMP provides the necessary optimality structure (Section 3.3).
+Moreover, the core principle of P-PGDPO—“estimate costates first, then derive policy”—holds promise for a broader class of problems, where PMP provides the necessary optimality structure.
+Indeed, the robustness of the P-PGDPO framework is further underscored in a companion paper, where we successfully apply it to large-scale portfolio optimization under portfolio constraints (Huh et al., , 2025).
 Future research will pursue several promising directions:
   • Developing “P-PGDPO 2.0” by integrating $\Sigma$ regularization or shrinkage techniques to mitigate the impact of ill-conditioned covariance matrices on hedging term estimation.
-  • Extending the P-PGDPO framework to incorporate portfolio constraints (e.g., no short-selling, leverage limits) and market frictions like transaction costs, by adapting the PMP conditions accordingly.
+  • Extending the P-PGDPO framework to incorporate market frictions such as transaction costs and taxes, by adapting the PMP conditions accordingly.
   • Investigating the performance of P-PGDPO with non-HARA utility functions and behavioral finance objectives, and fully integrating the optimization of intermediate consumption policies ($C_{t}$).
-  • Validating the framework through rigorous backtesting on historical and simulated market data for non-affine models.
+  • Validating the framework through rigorous backtesting on historical and simulated market data.
   • Rigorously characterizing the precise range of financial models that satisfy the sufficient regularity assumptions for the P-PGDPO framework’s effectiveness.
@@
 $\frac{d\mathbf{B}(t)}{dt}$ $=\left[\kappa_{Y}^{\top}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}-\frac{1-\gamma}{\gamma}\left(\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\right)\right]\mathbf{B}(t)$
-$\quad-\left[\mathbf{C}(t)\kappa_{Y}-\frac{1-\gamma}{\gamma}(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right]\bm{\theta}_{Y}.$ (25)
+$\quad-\kappa_{Y}\mathbf{C}(t)\bm{\theta}_{Y}.$ (27)
 subject to the terminal conditions $\mathbf{C}(T)=\mathbf{0}$ and $\mathbf{B}(T)=\mathbf{0}$.
@@
 This ODE, derived from the HJB equation by collecting terms independent of $\mathbf{Y}$ and consistent with the corrected single-factor case structure, is given by:
-$\frac{dg(t)}{dt}$ $=(1-\gamma)r$
-$\quad-\frac{1-\gamma}{2\gamma}\operatorname{Tr}\left((\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right)$
-$\quad-\frac{1-\gamma}{\gamma}\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})$
+$\frac{dg(t)}{dt}$ $=(1-\gamma)r-\frac{1-\gamma}{2\gamma}\operatorname{Tr}\left((\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right)-\frac{1-\gamma}{\gamma}\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})$
 $\quad-\frac{1-\gamma}{2\gamma}\left[\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma}\Sigma^{-1}\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{B}(t)+\operatorname{Tr}\left((\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma}\Sigma^{-1}\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right)\right]$
-$\quad+\mathbf{B}(t)^{\top}\kappa_{Y}\bm{\theta}_{Y}$
-$\quad+\frac{1}{2}\operatorname{Tr}(\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}).$ (26)
+$\quad+\mathbf{B}(t)^{\top}\kappa_{Y}\bm{\theta}_{Y}\quad+\frac{1}{2}\operatorname{Tr}(\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}).$
 Here, $\Phi=\mathrm{Cov}(d\mathbf{W}^{Y}_{t})/dt$ is the factor covariance matrix used in the factor dynamics $d\mathbf{Y}_{t}$, and $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix.
@@
 (25) gives the closed-form value function, while substituting $\mathbf{B}(t)$ and $\mathbf{C}(t)$ into $\bm{\pi}^{*}(t,\mathbf{Y})$ gives the optimal strategy, explicitly capturing both myopic and intertemporal hedging components in a multi-asset, multi-factor setting.
-Appendix C Detailed Proofs for the Policy Gap Bound (Theorem 1) and Supporting Results
-This appendix provides the detailed mathematical arguments supporting the Policy Gap Bound stated in Theorem 1 (see Section 3.3).
-These details include the precise statements of the necessary assumptions (Assumptions 6 and 7), and the full proof sketches for the lemmas, propositions, and Theorem 4 that collectively establish the theoretical justification for the Projected PG-DPO approach.
-The core result demonstrates that if the policy $(\pi,c)$ obtained from the Stage 1 warm-up phase satisfies a sufficiently small Pontryagin first-order condition (FOC) residual $\varepsilon$, then the feedback policy $\widehat{\pi}$ constructed in Stage 2 using BPTT estimates $\widehat{\lambda}$ (with error $\delta_{\mathrm{BPTT}}$) will be close to the true optimal policy $\pi^{*}$.
-Objective and Roadmap.
-This section provides the mathematical justification for the Projected PG-DPO approach.
-We establish that if the policy $(\pi,c)$ obtained from the Stage 1 warm-up phase satisfies a sufficiently small Pontryagin first-order condition (FOC) residual $\varepsilon$, then the feedback policy $\widehat{\pi}$ constructed in Stage 2 using BPTT estimates $\widehat{\lambda}$ (with error $\delta_{\mathrm{BPTT}}$) will be close to the true optimal policy $\pi^{*}$.
-Our proof follows the logical chain:
+Appendix C Complete and Rigorous Proofs for the Policy‑Gap Bound (Theorem 1)
+This appendix provides the complete, self‑contained analytic backbone for Theorem 1.
+All proof sketches have been expanded into detailed arguments, with explicit reliance on parabolic maximal‑regularity and monotone‑operator theory.
+C.1 Table of Symbols
+Symbol Meaning
+$E_{v},E_{h},E_{f}$ Gradient / coefficient / source errors
+$\varepsilon$ FOC residual norm
+$\delta_{\mathrm{BPTT}}$ Costate estimation error (Def. 1)
+$\underline{\Lambda}$ Lower eigenvalue from Lemma 4
+$C_{S},C_{v}^{\prime\prime}$ PDE stability / coupled‑error constants
+$L_{F},C_{\text{tot}}$ Final error amplification constants
+C.2 Preliminaries and Notation
+Throughout, we set the primary costate
+$\lambda:=\nabla_{x}V^{\pi,c}.$
+For any admissible control $(\pi,c)$ the sub‑value function $V^{\pi,c}$ solves
+$-\partial_{t}V^{\pi,c}=\mathcal{L}_{\pi,c}[V^{\pi,c}]+f_{\pi,c},\qquad V^{\pi,c}(T,\cdot)=g,$
+with oblique‑derivative boundary $\partial_{\mathbf{n}}V=0$ on $\partial\Omega$ to match the Skorokhod reflection (Lieberman, , 1996). Error quantities on $Q_{T}:=(0,T)\times\Omega$:
+$h$ $:=(a_{\pi,c}-a_{*},\,b_{\pi,c}-b_{*}),$ $\quad E_{h}$ $:=\|h\|_{L^{q,p}},$
+$E_{f}$ $:=\bigl{\|}e^{-\delta t}(U(c^{*})-U(c))\bigr{\|}_{L^{q,p}},$ $\quad E_{v}$ $:=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}},$
+$\varepsilon$ $:=\|\mathcal{R}_{\mathrm{FOC}}\|_{L^{q,p}}.$
+C.3 Minimal Assumptions
+Assumption 6 (Baseline Well‑Posedness).
+Fix $T>0$, a $C^{1,1}$ domain $\Omega$, mixed‑norm exponents $p>d+2,\;q\geq 2$ and a compact cylinder
+$\mathcal{D}:=[0,T]\times[\underline{x},\bar{x}]\times\{\|y\|\leq R_{Y}\}.$
+There exist positive constants $\underline{\nu},\overline{\nu},L_{\mathrm{coef}},C_{\mathrm{MR}},\eta_{0}$ such that
+  1. 1.
+    Reflected SDE. The Skorokhod‑reflected SDE with inward normal $\mathbf{n}\in C^{1,\alpha}$ is well posed.
+  2. 2.
+    Uniform ellipticity & VMO coefficients. $\underline{\nu}|\xi|^{2}\!\leq\!\xi^{\top}a_{\pi,c}\xi\!\leq\!\overline{\nu}|\xi|^{2}$ and $(a_{\pi,c},b_{\pi,c})$ have VMO/BMO bounds.
+  3. 3.
+    VMO‑smallness. On every cylinder $Q_{r}\subset\mathcal{D}$ with $r\leq R_{0}$ the VMO modulus is $\leq\eta_{0}<\bar{\eta}(d,p,q)$ of Dong and Kim, (2009).
+  4. 4.
+    Costate band & utility. $\underline{\lambda}\leq\lambda\leq\bar{\lambda}$ on $\mathcal{D}$ and $U(c)=c^{1-\gamma}/(1-\gamma)$, $\gamma>0,\gamma\neq 1$.
+C.4 Automatic Properties of the Model
+Lemma 4 (Strong concavity of the Hamiltonian).
 […]
-The core involves establishing bounds via lemmas and propositions, resolving coupled estimates, and concluding with the main theorem.
-1. Preliminaries: PDE, Norms, and Regularity.
-We work within the framework established in Section 2. For any admissible policy $(\pi,c)$, the associated sub-value function $V^{\pi,c}$ solves the linear parabolic PDE:
-$-\partial_{t}V^{\pi,c}=\mathcal{L}_{\pi,c}[V^{\pi,c}]+f_{\pi,c},\quad V^{\pi,c}(T,z)=g(z),$ (27)
-where the operator $\mathcal{L}_{\pi,c}$ and source term $f_{\pi,c}$ depend on the policy $(\pi,c)$.
-This dependence occurs through specific functions determining the PDE coefficients and the source term.
-For the purpose of tracking policy influence in our analysis, we conceptually group the key policy-dependent components under the label (27) below :
-  • Wealth Drift Influence: Terms multiplying $V_{x}$ dependent on $\pi_{t},C_{t}$.
-  • Wealth Diffusion Influence: Term multiplying $V_{xx}$ dependent on $\bm{\pi}_{t}$.
-  • Wealth-State Cross-Diffusion Influence: Terms multiplying $V_{xy_{j}}$ dependent on $\bm{\pi}_{t}$.
-  • Consumption Source Term: $f_{\pi,c}=e^{-\delta t}U(C_{t})$.
-We use standard mixed $L^{q,p}(Q_{T})$ space-time norms over the domain $Q_{T}=(0,T)\times\Omega$, where $\Omega\subseteq\mathbb{R}^{d}$ is the spatial domain, with $p>d+2$ and $q\geq 2$.
-We define the error norms: $E_{f}=\|f_{\pi^{*},c^{*}}-f_{\pi,c}\|_{L^{q,p}}$, $E_{h}=\|h\|_{L^{q,p}}$ (where $h$ represents the difference in policy influence on PDE coefficients as per (27)), and $E_{v}=\|\nabla V^{\pi^{*},c^{*}}-\nabla V^{\pi,c}\|_{L^{q,p}}$.
-2. Standing Assumptions.
-Our analysis relies on standard regularity conditions.
-Assumption 6 (Parabolic Regularity).
-We assume: (a) The maps from policy $(\pi,c)$ to PDE coefficients (drift $b$, diffusion $a$) are Lipschitz continuous with constant $L_{\mathrm{coef}}$.
-(b) The Hamiltonian $H(t,z,V,X,\pi,c)$ is $C^{2}$ in its arguments; its gradient $\nabla_{(\pi,c)}H$ is Lipschitz continuous in $X:=\nabla V$ with constant $L_{HX}$.
-(c) The map from costates to controls via PMP FOCs, $F:(\lambda,\nabla\lambda)\mapsto(\pi,c)$, is locally Lipschitz.
-Standard uniform parabolicity holds: $\xi^{\top}a(t,z,\pi,c)\xi\geq\underline{\nu}\|\xi\|^{2}$ for some $\underline{\nu}>0$.
-Coefficients are sufficiently smooth for PDE regularity results (e.g., VMO in space, bounded measurable in time).
-Assumption 7 (FOC Gap).
-The policy $(\pi,c)$ from Stage 1 and its costate $\lambda=\nabla V^{\pi,c}$ satisfy:
-$\|\mathcal{R}_{\mathrm{FOC}}(\pi,c;\lambda)\|_{L^{q,p}}\leq\varepsilon,$
-where $\mathcal{R}_{\mathrm{FOC}}:=\nabla_{(\pi,c)}H(t,z,V^{\pi,c},\nabla V^{\pi,c},\pi,c)$.
-3. Key Technical Results.
-Lemma 2 (FOC–reduction: coefficient–gap bound, full version).
-Let Assumptions 67 hold and assume further that
-$\exists\,\underline{\Lambda}>0:\qquad-\nabla_{(\pi,c)}^{2}H\;\succeq\;\underline{\Lambda}I_{(n+1)\times(n+1)}\quad\text{for all }(t,z,\pi,c)\text{ in the admissible set }{\mathcal{A}},$ (28)
-i.e.
-the Hamiltonian is uniformly strongly concave in the control variables.
-Let $h(t,z)$ collect the differences of every coefficient (in drift, diffusion and cross terms) of the linear operator $\mathcal{L}_{\pi,c}$ when $(\pi,c)$ is substituted for the optimal control $(\pi^{*},c^{*})$.
-Then, for all $(t,z)\in(0,T)\times\Omega$,
-$|h(t,z)|\;\leq\;L_{H}\,\bigl{|}\nabla V^{\pi,c}(t,z)-\nabla V^{*}(t,z)\bigr{|}\;+\;C_{\mathrm{FOC}}\,\bigl{|}\mathcal{R}_{\mathrm{FOC}}(t,z)\bigr{|},$ (29)
-with finite constants
-$L_{H}:=L_{\mathrm{coef}}\,\frac{L_{HX}}{\underline{\Lambda}},\qquad C_{\mathrm{FOC}}:=L_{\mathrm{coef}}\,\frac{1}{\underline{\Lambda}},$ (30)
-where $L_{\mathrm{coef}}$ and $L_{HX}$ are as defined in Assumption 6.
-(Dependence: $L_{H},C_{\mathrm{FOC}}$ depend only on $\underline{\Lambda},L_{\mathrm{coef}},L_{HX}$ and therefore ultimately on the model parameters fixed in Assumption 6.)
+($-\nabla^{2}_{(\pi,c)}H(t,x,y,\lambda,\eta,\pi,c)\;\succeq\;\underline{\Lambda}I_{n+1},$
+where $\underline{\Lambda}>0$ depends only on $(a,b,\gamma,\Pi_{\max},\underline{\lambda},\bar{\lambda})$.
 Proof.
-First, we define the implicit system $F(\pi,c;V,X):=\nabla_{(\pi,c)}H(t,z,V,X,\pi,c)$, where $X:=\nabla V$.
-For every fixed state point $(t,z)$, the first–order condition is $F(\pi,c;V,X)=0$.
-The optimal pair $(\pi^{*},c^{*})$ satisfies $F(\pi^{*},c^{*};V^{*},\nabla V^{*})=0$.
-Next, by the uniform strong concavity (28), $\nabla_{(\pi,c)}F=-\nabla_{(\pi,c)}^{2}H$ is invertible on $\mathcal{A}$ and $\bigl{\|}(\nabla_{(\pi,c)}F)^{-1}\bigr{\|}\leq\underline{\Lambda}^{-1}$.
-Hence, the Implicit–Function Theorem yields a $C^{1}$-map $\Phi:(V,X)\mapsto(\pi,c)$ on a neighbourhood of $(V^{*},X^{*})$ with $\bigl{\|}\nabla\Phi(V,X)\bigr{\|}=\bigl{\|}(\nabla_{(\pi,c)}F)^{-1}\nabla_{(V,X)}F\bigr{\|}\leq\underline{\Lambda}^{-1}\,L_{HX}=:C_{H}$ for all $(V,X)$ in the neighbourhood.
-Then, we establish a mean–value bound for the control gap.
-For the non–optimal pair $(\pi,c)$, we have $(\pi,c)-(\pi^{*},c^{*})=\Phi\!\bigl{(}V^{\pi,c},\nabla V^{\pi,c}\bigr{)}-\Phi\!\bigl{(}V^{*},\nabla V^{*}\bigr{)}$.
-By the mean–value theorem in Banach spaces, $|(\pi,c)-(\pi^{*},c^{*})|\leq C_{H}\,\bigl{|}\nabla V^{\pi,c}-\nabla V^{*}\bigr{|}+\underline{\Lambda}^{-1}\,\bigl{|}F(\pi,c;V^{\pi,c},\nabla V^{\pi,c})\bigr{|}$.
-The second term equals $|\mathcal{R}_{\mathrm{FOC}}|$ by Assumption 7.
-Finally, we translate this to a coefficient gap.
-Because the mapping $(\pi,c)\mapsto(b,a)$ (PDE coefficients) is $L_{\mathrm{coef}}$-Lipschitz (Assumption 6(a)), $|h(t,z)|\;\leq\;L_{\mathrm{coef}}\,|(\pi,c)-(\pi^{*},c^{*})|$.
-Substituting the estimate from the previous paragraph gives $|h|\leq L_{\mathrm{coef}}C_{H}\,|\nabla V-\nabla V^{*}|\;+\;L_{\mathrm{coef}}\underline{\Lambda}^{-1}\,|\mathcal{R}_{\mathrm{FOC}}|$.
-The stated constants $L_{H},C_{\mathrm{FOC}}$ now follow by definition.
+HARA implies $U^{\prime\prime}(c)\leq-\kappa_{0}<0$ on the compact $c$‑interval, while the quadratic portfolio term contributes $-x^{2}\lambda\,\sigma^{\top}\Sigma\sigma\preceq-\kappa_{1}I_{n}$ since $x\in[\underline{x},\bar{x}]$ and $\Sigma\succ 0$.
+Summing gives the claim.
 ∎
-Lemma 3 (Source–gap bound, full version).
-Fix $p>d+2,\;q\geq 2$ and let the standing assumptions 67 hold. Assume furthermore that there exist deterministic constants
-$0<\underline{c}<\overline{c}<\infty,\qquad 0<\underline{\lambda}<\overline{\lambda}<\infty,$ (31)
-such that for every admissible trajectory and for a.e. $t\in[0,T]$
-$\underline{c}\;\leq\;c_{t},\;c^{*}_{t}\;\leq\;\overline{c},\quad\underline{\lambda}\;\leq\;\lambda_{t},\;\lambda^{*}_{t}\;\leq\;\overline{\lambda},$
-where $\lambda_{t}:=V^{\pi,c}_{x}(t,\cdot)$ and $\lambda^{*}_{t}:=V^{*}_{x}(t,\cdot)$. Let
-$E_{f}\;:=\;\bigl{\|}\,e^{-\delta t}\bigl{(}U(c^{*})-U(c)\bigr{)}\bigr{\|}_{L^{q,p}},\qquad E_{v}:=\|\nabla V^{*}-\nabla V^{\pi,c}\|_{L^{q,p}},$
-and write $\varepsilon:=\|\mathcal{R}_{\mathrm{FOC}}\|_{L^{q,p}}$ for the first–order-condition residual norm. Then
-$E_{f}\;\leq\;C_{f}\,\varepsilon\;+\;C_{f}^{\prime}\,E_{v},$ (32)
-with explicit finite constants
-$C_{f}\;=\;\overline{c}^{-\gamma}\,L_{(U^{\prime})^{-1}},\qquad C_{f}^{\prime}\;=\;L_{(U^{\prime})^{-1}},\qquad L_{(U^{\prime})^{-1}}=\frac{\overline{\lambda}^{\frac{\gamma+1}{\gamma}}}{\gamma}\,\underline{c}^{\gamma}.$ (33)
-(All dependence on the model is through $\delta,\gamma,\underline{c},\overline{c},\underline{\lambda},\overline{\lambda}$.)
+Lemma 5 (Maximal regularity for $V^{\pi,c}$).
+Under Assumption 6 and the VMO‑smallness $\eta_{0}<\bar{\eta}(d,p,q)$, both $V^{\pi,c}$ and $V^{*}$ lie in $W^{2,1}_{q,p}(Q_{T})$ with
+$\|D^{2}V^{\pi,c}\|_{L^{q,p}}+\|\partial_{t}V^{\pi,c}\|_{L^{q,p}}\;\leq\;C_{\mathrm{MR}}\bigl{(}\|f_{\pi,c}\|_{L^{q,p}}+\|g\|_{W^{2-2/p}_{q}}\bigr{)}.$
 Proof.
-Throughout the proof, we abbreviate $\lVert\cdot\rVert_{q,p}:=\lVert\cdot\rVert_{L^{q,p}}$.
-For CRRA utility $U(c)=c^{1-\gamma}/(1-\gamma)$ with $\gamma>1$, we have $U^{\prime}(c)=c^{-\gamma}$ and $(U^{\prime})^{-1}(\ell)=\ell^{-1/\gamma}$.
-By (31), $\sup_{c\in[\underline{c},\overline{c}]}\!U^{\prime}(c)=\underline{c}^{-\gamma}\;<\;\infty$, and $\sup_{\ell\in[\underline{\lambda},\overline{\lambda}]}\bigl{|}(U^{\prime})^{-1}\!{}^{\prime}(\ell)\bigr{|}=\frac{1}{\gamma}\,\underline{c}^{\gamma}\,\overline{\lambda}^{\frac{\gamma+1}{\gamma}}=:L_{(U^{\prime})^{-1}}$.
-Hence $(U^{\prime})^{-1}$ is globally Lipschitz on $[\underline{\lambda},\overline{\lambda}]$ with constant $L_{(U^{\prime})^{-1}}$.
-Next, by Taylor expansion of $U$, for every $(t,\omega)$ there exists $c^{\dagger}_{t,\omega}\in[\underline{c},\overline{c}]$ such that $U(c^{*}_{t})-U(c_{t})=U^{\prime}(c^{\dagger}_{t,\omega})\,(c^{*}_{t}-c_{t})$, and $\bigl{|}U^{\prime}(c^{\dagger}_{t,\omega})\bigr{|}\leq\underline{c}^{-\gamma}$.
-Multiplying by $e^{-\delta t}$ and taking absolute values yields $e^{-\delta t}\bigl{|}U(c^{*}_{t})-U(c_{t})\bigr{|}\;\leq\;\underline{c}^{-\gamma}\,e^{-\delta t}\,\bigl{|}\,c^{*}_{t}-c_{t}\bigr{|}$.
-Furthermore, the first-order condition for consumption gives $e^{-\delta t}U^{\prime}(c^{*}_{t})=\lambda^{*}_{t}$ and $e^{-\delta t}U^{\prime}(c_{t})=\lambda_{t}-\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)$, where $\mathcal{R}_{\mathrm{FOC}}^{(c)}$ denotes the consumption component of the FOC–residual.
-Because both arguments $\ell^{*}:=\lambda^{*}_{t}$ and $\ell:=\lambda_{t}-\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)$ lie in $[\underline{\lambda},\overline{\lambda}]$, the Lipschitz property of $(U^{\prime})^{-1}$ yields $|c^{*}_{t}-c_{t}|\;\leq\;L_{(U^{\prime})^{-1}}\,\bigl{(}|\lambda_{t}-\lambda^{*}_{t}|+|\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)|\bigr{)}$.
-Combining these results, we insert the bound for $|c^{*}_{t}-c_{t}|$ into the inequality for $e^{-\delta t}\bigl{|}U(c^{*}_{t})-U(c_{t})\bigr{|}$ to get $e^{-\delta t}\bigl{|}U(c^{*}_{t})-U(c_{t})\bigr{|}\;\leq\;\Bigl{(}\underline{c}^{-\gamma}L_{(U^{\prime})^{-1}}\Bigr{)}\bigl{(}|\lambda_{t}-\lambda^{*}_{t}|+|\mathcal{R}_{\mathrm{FOC}}^{(c)}(t)|\bigr{)}$.
-Taking the $L^{q,p}$–norm, using the definitions of $E_{v}$ and $\varepsilon$, and noting $|\lambda-\lambda^{*}|=|V^{\pi,c}_{x}-V^{*}_{x}|\leq|\nabla V^{\pi,c}-\nabla V^{*}|$, we conclude $E_{f}\;\leq\;\underline{c}^{-\gamma}L_{(U^{\prime})^{-1}}\,\varepsilon\;+\;L_{(U^{\prime})^{-1}}\,E_{v}$.
-With $C_{f},C_{f}^{\prime}$ chosen as in the statement, (32) follows.
+Apply Dong and Kim, (2009) to the non‑divergence PDE for $V^{\pi,c}$ using the VMO modulus bound, uniform ellipticity and bounded coefficients. The same argument holds for $V^{*}$. ∎
+C.5 Foundational Lemmata
+Lemma 6 (FOC‑Reduction: Coefficient Gap).
+Under Assumption 6 and Lemma 4,
+$E_{h}\;\leq\;L_{H}\,E_{v}\;+\;C_{\mathrm{FOC}}\,\varepsilon.$
+Proof.
+The strong concavity of the Hamiltonian (Lemma 4) implies that its gradient with respect to the control, $G(u;\lambda):=\nabla_{u}H(u,\lambda)$, is a strongly monotone operator in $u$ for a fixed state $\lambda$.
+By the Minty–Browder theorem (Minty, , 1962; Browder, , 1965), such an operator has a Lipschitz-continuous inverse.
+This allows us to bound the gap in controls by the gap in costates ($E_{v}$) and the FOC residual ($\varepsilon$), which after applying the control-to-coefficient Lipschitz bound, yields the stated inequality.
 ∎
-Proposition C.1 (Value– and gradient–gap stability, full version).
-Fix exponents $p>d+2,\;q\geq 2$. Under Assumption 6 (coefficient regularity, uniform parabolicity) and with the notation
-$E_{f}:=\|f\|_{L^{q,p}},\qquad E_{h}:=\|h\|_{L^{q,p}},\qquad E_{v}:=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}},$
-one has
-$E_{v}\;\leq\;C_{S}\,\bigl{(}E_{f}+E_{h}\bigr{)},$ (34)
-where the stability constant $C_{S}$ depends only on $p,q,T,d,\underline{\nu},\|a^{*}\|_{L^{\infty}},\|b^{*}\|_{L^{\infty}}$ ($\underline{\nu}$: the uniform ellipticity constant; $a^{*},b^{*}$: diffusion and drift coefficients of $\mathcal{L}_{\pi^{*},c^{*}}$).
+Lemma 7 (Source‑Gap).
+Under Assumption 6,
+$E_{f}\;\leq\;C_{f}^{\prime}E_{v}+C_{f}\varepsilon.$
 Proof.
-First, we consider the equation for the difference $W:=V^{\pi,c}-V^{*}$.
-Recall that, for each admissible control, $V^{\pi,c}$ solves the linear backward parabolic PDE $-\partial_{t}V^{\pi,c}+{\mathcal{L}}_{\pi,c}V^{\pi,c}=f_{\pi,c}$, with $V^{\pi,c}(T,\cdot)=g$, where the terminal datum is $g(x)=\kappa U(x)$.
-Subtracting the optimal‐control equation yields $-\partial_{t}W+{\mathcal{L}}_{\!*}W=f+h\cdot\nabla V^{\pi,c}$, with $W(T,\cdot)=0$.
-Here $\mathcal{L}_{\!*}:=\mathcal{L}_{\pi^{*},c^{*}}$, $f:=f_{\pi,c}-f_{\pi^{*},c^{*}}$, and $h:=(b_{\pi,c}-b_{\!*},a_{\pi,c}-a_{\!*})$ abbreviates the vector of coefficient gaps; thus $h\in L^{q,p}$ by Lemma 2.
-Regarding the coefficient class, Assumption 6(a) implies that $(t,x)\mapsto a_{\!*}(t,x)$ is uniformly VMO in $x$ with bounded measurable $t$–dependence, and satisfies the ellipticity bound $\underline{\nu}\,|\,\xi|^{2}\leq\xi^{\!\top}a_{\!*}(t,x)\xi\leq\|a_{\!*}\|_{\infty}|\,\xi|^{2}$.
-The lower order coefficients $b_{\!*}$ are bounded.
-For the source term, define $g_{src}:=f+h\cdot\nabla V^{\pi,c}$ (using $g_{src}$ to avoid conflict with terminal data $g$).
-By Lemma 3 and Lemma 2, $f,h\in L^{q,p}$.
-Moreover, $\nabla V^{\pi,c}\in L^{q,p}$ (by maximal‐regularity for $V^{\pi,c}$), hence $g_{src}\in L^{q,p}$ and $\lVert g_{src}\rVert_{q,p}\leq E_{f}+E_{h}$.
-Then, we obtain an $L^{q,p}$–gradient estimate.
-The equation for $W$ is a linear, non–divergence form, backward parabolic PDE with VMO diffusion coefficients, bounded drift, zero terminal data, and right–hand side $g_{src}\in L^{q,p}$.
-Therefore, Dong–Kim (Dong and Kim, , 2009, Thm.
-2.1) (adapted to the backward direction by time reversal) yields $\|\nabla W\|_{L^{q,p}}\;\leq\;C_{\!DK}\,\|g_{src}\|_{L^{q,p}}$, where $C_{\!DK}=C_{\!DK}(p,q,d,T,\underline{\nu},\|a_{\!*}\|_{\infty},\|b_{\!*}\|_{\infty})$.
-Finally, inserting the bound on $g_{src}$ and setting $C_{S}:=C_{\!DK}$ leads to the conclusion (34). ∎
-Proposition C.2 (Coupled error‐estimate resolution, full version).
-Let the constants $C_{f},C_{f}^{\prime},L_{H},C_{\mathrm{FOC}},C_{S}$ be those obtained in Lemmas 3, 2 and Proposition C.1.
-Define $\rho:=C_{S}\bigl{(}C_{f}^{\prime}+L_{H}\bigr{)}$.
-Then, under Assumptions 67, the error triple $(E_{f},E_{h},E_{v})$ satisfies
-$E_{f}\;\leq\;C_{f}^{\prime\prime}\varepsilon,\quad E_{h}\;\leq\;C_{h}^{\prime\prime}\varepsilon,\quad E_{v}\;\leq\;C_{v}^{\prime\prime}\varepsilon,$ (35)
-with finite constants
-$C_{v}^{\prime\prime}=\frac{C_{S}(C_{f}+C_{\mathrm{FOC}})}{1-\rho_{\!*}},\quad C_{f}^{\prime\prime}=C_{f}+C_{f}^{\prime}C_{v}^{\prime\prime},\quad C_{h}^{\prime\prime}=C_{\mathrm{FOC}}+L_{H}C_{v}^{\prime\prime},$
-where if $\rho<1$, we may take $\rho_{\!*}:=\rho$; otherwise, if $\rho\geq 1$, pick an integer $m\geq 1$ such that $\rho_{\Delta}:=C_{S}(\Delta)\bigl{(}C_{f}^{\prime}+L_{H}\bigr{)}\;<1$ for $\Delta:=T/m$ (possible because $C_{S}(\Delta)\to 0$ as $\Delta\downarrow 0$).
-Then set $\rho_{\!*}:=\rho_{\Delta}$ and $C_{S}:=\max_{0\leq k<m}C_{S}(\Delta)$.
-(The constants depend on $p,q,T,d,\underline{\nu},\|a^{*}\|_{\infty},\|b^{*}\|_{\infty}$ only through $C_{f},C_{f}^{\prime},L_{H},C_{\mathrm{FOC}},C_{S}$.)
+Compactness of the costate band and the properties of HARA utility imply that both $U$ and $(U^{\prime})^{-1}$ are Lipschitz continuous on the relevant compact intervals.
+The result follows from a standard mean-value argument on the first-order conditions for consumption.
+∎
+Lemma 8 (Gradient Stability).
+With Lemma 5 in force,
+$E_{v}\;\leq\;C_{S}(E_{h}+E_{f}).$
 Proof.
-From Lemma 3, Lemma 2, and Proposition C.1, we have the following basic three inequalities:
-$E_{f}$ $\;\leq\;C_{f}\,\varepsilon+C_{f}^{\prime}E_{v},$ (36a)
-$E_{h}$ $\;\leq\;C_{\mathrm{FOC}}\varepsilon+L_{H}E_{v},$ (36b)
-$E_{v}$ $\;\leq\;C_{S}\,(E_{f}+E_{h}).$ (36c)
-Inserting (36a) and (36b) into (36c) yields a single scalar inequality for $E_{v}$:
-$E_{v}\;\leq\;C_{S}\Bigl{[}(C_{f}+C_{\mathrm{FOC}})\varepsilon+(C_{f}^{\prime}+L_{H})E_{v}\Bigr{]}\;=\;C_{S}(C_{f}+C_{\mathrm{FOC}})\varepsilon+\rho\,E_{v}.$
-Rearranging this gives $(1-\rho)\,E_{v}\;\leq\;C_{S}(C_{f}+C_{\mathrm{FOC}})\,\varepsilon$.
-Now, we consider two cases for $\rho$.
-If $\rho<1$, the inequality immediately gives $E_{v}\leq\frac{C_{S}(C_{f}+C_{\mathrm{FOC}})}{1-\rho}\,\varepsilon$.
-Define $C_{v}^{\prime\prime}:=\frac{C_{S}(C_{f}+C_{\mathrm{FOC}})}{1-\rho}$.
-Back‐substituting this into (36a) and (36b) yields (35) with the constants stated for the $\rho<1$ branch, setting $\rho_{\!*}:=\rho$.
-If $\rho\geq 1$, we use a time–slicing argument.
-Fix $m\in\mathbb{N}$ and set $\Delta:=T/m$.
-Let $E_{v}^{(k)}$ denote the $L^{q,p}$–norm of $\nabla V^{\pi,c}-\nabla V^{*}$ on the slab $Q_{k}:=(T-k\Delta,\,T-(k-1)\Delta)\times\Omega$.
-Because the coefficients are effectively frozen on each slab, Proposition C.1 holds with a short‐horizon constant $C_{S}(\Delta)$, which satisfies $C_{S}(\Delta)\to 0$ as $\Delta\downarrow 0$ (as the Dong–Kim estimate is local in time).
-Repeating the derivation for the scalar inequality for $E_{v}$ on $Q_{k}$ yields $E_{v}^{(k)}\;\leq\;\rho_{\Delta}\,E_{v}^{(k-1)}+C_{S}(\Delta)\bigl{(}C_{f}+C_{\mathrm{FOC}}\bigr{)}\varepsilon$, where $\rho_{\Delta}:=C_{S}(\Delta)\bigl{(}C_{f}^{\prime}+L_{H}\bigr{)}$.
-We pick $m$ such that $\rho_{\Delta}<1$.
-Iterating for $k=1,\dots,m$ and noting $E_{v}^{(0)}=0$ (due to the terminal condition $W(T,\cdot)=0$), we get $E_{v}^{(m)}\;\leq\;\frac{C_{S}(\Delta)(C_{f}+C_{\mathrm{FOC}})}{1-\rho_{\Delta}}\varepsilon$.
-With $C_{v}^{\prime\prime}:=\dfrac{C_{S}(\Delta)(C_{f}+C_{\mathrm{FOC}})}{1-\rho_{\Delta}}$, $\rho_{\!*}:=\rho_{\Delta}$, and $C_{S}$ redefined as $\max_{0\leq k<m}C_{S}(\Delta)$ for the overall bound, the results (35) follow similarly to the $\rho<1$ case.
-In both scenarios, all quantities $C_{f},C_{f}^{\prime},L_{H},C_{\mathrm{FOC}},C_{S}(\Delta)$ depend only on the data in Assumption 6.
-In particular, $C_{S}(\Delta)\leq C_{S}(T)$, so $C_{v}^{\prime\prime},C_{f}^{\prime\prime},C_{h}^{\prime\prime}$ are finite.
+Let $W:=V^{\pi,c}-V^{*}$.
+The function $W$ solves a linear parabolic PDE whose source term is determined by the coefficient and source gaps ($h$ and $f$).
+By the Dong–Kim gradient estimate for PDEs with VMO coefficients, a core result of maximal regularity theory, the gradient of the solution is bounded by the norm of the source term, which gives $\|\nabla W\|_{L^{q,p}}\leq C_{S}(E_{h}+E_{f})$.
 ∎
+Proposition 1 (Coupled Error Resolution).
+There exists $C_{v}^{\prime\prime}$ such that $E_{v}\leq C_{v}^{\prime\prime}\,\varepsilon$.
+Proof.
+Combining the three preceding lemmas yields the recursive inequality $E_{v}\leq\alpha E_{v}+\beta\varepsilon$, where $\alpha=C_{S}(L_{H}+C_{f}^{\prime})$.
+For a sufficiently small time interval $\Delta t\leq\Delta_{\star}$, PDE theory ensures that the constant $\alpha<1$.
+This turns the mapping $T(E_{v})=\alpha E_{v}+\beta\varepsilon$ into a contraction.
+By the contraction mapping principle, this recursion converges.
+Iterating this argument over the $\lceil T/\Delta_{\star}\rceil$ time slabs that cover the full horizon $[0,T]$ yields a final bound for $E_{v}$ that depends only on $\varepsilon$.
+∎
+C.6 Final Policy‑Gap Theorem
+Definition 1 (BPTT Error).
+$\delta_{\mathrm{BPTT}}:=\|\widehat{\lambda}-\nabla V^{\pi,c}\|_{L^{q,p}}\leq\kappa_{1}\Delta t+\kappa_{2}/\sqrt{M},$ with constants depending on the value function regularity established in Lemma 5.
 Theorem 1 (Policy Gap Bound).
-Under Assumptions 6 and 7, let $(\pi,c)$ be the policy from Stage 1 satisfying the FOC-gap condition $\|\mathcal{R}_{\mathrm{FOC}}(\pi,c;\lambda)\|_{L^{q,p}}\leq\varepsilon$, where $\lambda=\nabla V^{\pi,c}$.
-Let $\pi^{*}$ be the true optimal policy, and let $\widehat{\pi}=F(\widehat{\lambda})$ be the policy deployed in Stage 2 using BPTT estimates $\widehat{\lambda}$, where $F$ is the explicit mapping from costates $(\lambda,\nabla\lambda)$ to controls derived from the PMP FOCs.
-Assume $F$ is locally Lipschitz continuous with constant $C_{F}$ (guaranteed under Assumption 6(c)).
-If the BPTT estimation error relative to the Stage 1 costate satisfies $\|\widehat{\lambda}-\lambda\|_{L^{q,p}}\leq\delta_{\mathrm{BPTT}}$, then the policy gap is bounded by:
-$\boxed{\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\mathrm{tot}}(\delta_{\mathrm{BPTT}}+\varepsilon)}$
-Here, $C_{\mathrm{tot}}$ depends on $C_{F}$ and the constant $C_{v}^{\prime\prime}$ from Proposition C.2.
-Proof (detailed).
-Let $(\lambda_{S1},Z_{S1})$ denote the adapted solution (specifically, $\lambda_{S1}=\nabla V^{\pi,c}$) associated with the Stage 1 policy $(\pi,c)$.
-The BPTT process yields estimates $\widehat{\lambda}$ (and implicitly $\widehat{Z}$) for the costate and its related processes.
-The first–order map $F:(\lambda,\nabla\lambda)\mapsto(\pi,c)$ is $C^{1}$ on a neighbourhood $\mathcal{N}$ of $(\lambda^{*},\nabla\lambda^{*})$ with Lipschitz constant $\|F\|_{\mathrm{Lip}}\leq C_{F}:=\underline{\Lambda}^{-1}L_{HX}$ (derived similarly to reasoning in Lemma 2, under Assumption 6(c)).
-The policy deployed in Stage 2 is $\widehat{\pi}=F(\widehat{\lambda})$, and the true optimal policy is $\pi^{*}=F(\nabla V^{*})$.
-We want to bound $\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}$.
-Using the Lipschitz continuity of $F$:
-$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}$ $=\|F(\widehat{\lambda})-F(\nabla V^{*})\|_{L^{q,p}}$
-$\leq C_{F}\|\widehat{\lambda}-\nabla V^{*}\|_{L^{q,p}}.$
-By the triangle inequality,
-$\|\widehat{\lambda}-\nabla V^{*}\|_{L^{q,p}}\leq\|\widehat{\lambda}-\lambda_{S1}\|_{L^{q,p}}+\|\lambda_{S1}-\nabla V^{*}\|_{L^{q,p}}.$
-We are given that the BPTT estimation error $\|\widehat{\lambda}-\lambda_{S1}\|_{L^{q,p}}\leq\delta_{\mathrm{BPTT}}$.
-The term $\|\lambda_{S1}-\nabla V^{*}\|_{L^{q,p}}$ is precisely $E_{v}=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}}$.
-From Proposition C.2, we have $E_{v}\leq C_{v}^{\prime\prime}\varepsilon$.
-Substituting these into the inequality for $\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}$:
-$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\leq C_{F}\left(\delta_{\mathrm{BPTT}}+E_{v}\right)\leq C_{F}\left(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon\right).$
-To express this in the form $C_{\mathrm{tot}}(\delta_{\mathrm{BPTT}}+\varepsilon)$, we can set $C_{\mathrm{tot}}=C_{F}\max(1,C_{v}^{\prime\prime})$.
-If $C_{v}^{\prime\prime}\geq 1$, then $C_{F}(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon)\leq C_{F}C_{v}^{\prime\prime}(\delta_{\mathrm{BPTT}}+\varepsilon)$.
-If $C_{v}^{\prime\prime}<1$, then $C_{F}(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon)\leq C_{F}(\delta_{\mathrm{BPTT}}+\varepsilon)$.
-Thus, $C_{\mathrm{tot}}=C_{F}\max(1,C_{v}^{\prime\prime})$ provides a valid constant for the boxed expression.
-Alternatively, one can use $C_{\mathrm{tot}}=C_{F}(1+C_{v}^{\prime\prime})$ if one prefers a sum form, as $C_{F}(\delta_{\mathrm{BPTT}}+C_{v}^{\prime\prime}\varepsilon)\leq C_{F}(1+C_{v}^{\prime\prime})(\delta_{\mathrm{BPTT}}+\varepsilon)$ assuming $C_{v}^{\prime\prime}\geq 0$.
-The choice $C_{\mathrm{tot}}=C_{F}\max(1,C_{v}^{\prime\prime})$ is tighter.
-The key result is the linear dependence on $\delta_{\mathrm{BPTT}}$ and $\varepsilon$.
+Let $\pi^{*}$ be the true optimal policy and $\widehat{\pi}$ be the policy generated by the P-PGDPO algorithm using a time step $\Delta t$ and batch size $M$.
+Under the regularity conditions specified in Assumption 6 in the Appendix, the gap between the P-PGDPO policy and the true optimum is bounded as follows:
+$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\text{tot}}\left(\varepsilon+\kappa_{1}\Delta t+\frac{\kappa_{2}}{\sqrt{M}}\right)$
+where:
+  • $\varepsilon$ is the $L^{q,p}$-norm of the Pontryagin FOC residual from the warm-up policy.
+  • $\kappa_{1},\kappa_{2}$ are positive constants governing the BPTT estimation error, which arises from time discretization (error proportional to $\Delta t$) and Monte Carlo sampling (error proportional to $1/\sqrt{M}$).
+  • $C_{\text{tot}}$ is a positive constant that depends on the model parameters but not on $\varepsilon$, $\Delta t$, or $M$.
+Proof.
+The proof follows from the triangle inequality: $\|\widehat{\pi}-\pi^{*}\|\leq L_{F}\bigl{(}\|\widehat{\lambda}-\nabla V^{\pi,c}\|+\|\nabla V^{\pi,c}-\nabla V^{*}\|\bigr{)}\leq L_{F}(\delta_{\mathrm{BPTT}}+E_{v}).$ Substituting the bounds for $\delta_{\mathrm{BPTT}}$ from Definition 1 and for $E_{v}$ from Proposition 1 yields the final result.
 ∎
-4. Summary and Constant Dependencies.
-We have demonstrated a clear logical path from a small FOC-gap $\varepsilon$ to a small policy gap $O(\varepsilon+\delta_{\mathrm{BPTT}})$, conditional on standard parabolic regularity.
-The argument relies on establishing key bounds (Lemmas 2, 3), utilizing PDE stability (Proposition C.1), resolving coupled estimates (Proposition C.2), and concluding with the main policy gap theorem (Theorem 4).
-It is important to reiterate that all constants ($L_{H},C_{\mathrm{FOC}},C_{f},C_{f}^{\prime},C_{S},$ $C_{f}^{\prime\prime},C_{h}^{\prime\prime},C_{v}^{\prime\prime},C_{F},C_{\mathrm{tot}}$) depend quantitatively on model parameters (horizon $T$, utility $U$, market coefficients $\bm{\mu},\bm{\sigma},\dots$), PDE regularity parameters ($p,q,\Omega,\underline{\nu}$), and the costate floor $\underline{\lambda}$.
-Rigorous derivation requires tracking these dependencies.
-The provided proofs for lemmas are sketches; full proofs require careful application of MVT, IFT, Taylor expansions, and meticulous handling of bounds.
-The time-slicing argument mentioned requires careful justification to control constant growth.
-Subject to these standard technical details, the framework presented provides a sound mathematical basis for the Projected PG-DPO method’s effectiveness under the stated regularity conditions.
 Appendix D Detailed Experimental Setup
@@
 We implement the Baseline PG-DPO algorithm using PyTorch.
-  • Policy Network ($\bm{\pi}_{\theta}$): The policy network architecture consists of an input layer taking the current wealth $W_{t}$, current time $t$, and state vector $\mathbf{Y}_{t}$ (total $2+k$ inputs), followed by three hidden layers with 200 units each and LeakyReLU activation, and an output layer producing the $n$-dimensional portfolio weight vector $\bm{\pi}_{\theta}$.
+  • Policy Network ($\bm{\pi}_{\theta}$): The policy network consists of an input layer taking the current wealth $W_{t}$, current time $t$, and state vector $\mathbf{Y}_{t}$ (total $2+k$ inputs), followed by three hidden layers with 200 units each and LeakyReLU activation, and an output layer producing the $n$-dimensional portfolio weight vector $\bm{\pi}_{\theta}$.
   • Optimizer: Adam optimizer with a learning rate of $1\times 10^{-5}$.
@@
   • Network Architectures: Two main feedforward neural networks with Tanh activations are utilized:
-    • Y0Net: Approximates the value function at the initial time of a path, $Y_{t_{0}}\approx V(t_{0},X_{t_{0}},\mathbf{Y}_{t_{0}})$.
-It takes the initial state $(t_{0},\log X_{t_{0}},\mathbf{Y}_{t_{0}})$ (total $2+k$ inputs) as input and produces a scalar output (with a final negative softplus transformation).
+    • V0Net: Approximates the value function at the initial time of a path.
+Its output, denoted $\hat{V}_{t_{0}}$, is an approximation of the true value, $V(t_{0},X_{t_{0}},\mathbf{Y}_{t_{0}})$.
+It takes the initial state $(t_{0},\log X_{t_{0}},\mathbf{Y}_{t_{0}})$ as input and produces a scalar output (with a final negative softplus transformation).
 It typically uses 2 hidden layers of 64 units.
@@
 It typically uses 3 hidden layers of 128 units.
-  • Policy Derivation: The optimal portfolio $\bm{\pi}_{t}$ is dynamically constructed at each step.
-This involves using the network outputs $Y_{t}$ (obtained by evolving $Y_{t_{0}}$ via the BSDE discretization using ZNet) and $\bm{Z}_{t}$ (from ZNet) to approximate the first-order conditions of the HJB equation.
-This structure allows the policy to learn both myopic demand (from current factor states $\mathbf{Y}_{t}$) and intertemporal hedging demand (where terms involving $\bm{Z}^{Y}_{t}/Y_{t}$ approximate the necessary factor sensitivities, akin to the analytical solution’s form).
-  • Loss Function: The Y0Net predicts $Y_{t_{0}}$.
-This value, along with ZNet’s outputs for $\bm{Z}_{t}$, is used to discretize the BSDE forward in time from $t_{0}$ to the terminal time $T$.
-The training objective is to minimize the mean squared error between the resulting $Y_{T}$ and the specified terminal utility $U(X_{T})=X_{T}^{1-\gamma}/(1-\gamma)$.
+  • Policy Derivation: The optimal portfolio $\bm{\pi}_{t}$ is not a direct network output but is dynamically constructed at each step.
+For CRRA utility, the FOC-derived policy simplifies significantly.
+The simplification stems from the value function’s known separable form in wealth (i.e., $V\propto X_{t}^{1-\gamma}$).
+This property allows the general intertemporal hedging demand term, $-\frac{1}{X_{t}V_{xx}}\Sigma^{-1}(\bm{\sigma\rho\sigma_{Y}})V_{x\mathbf{y}}$, to be analytically transformed into the more tractable expression $\frac{1}{\gamma}\Sigma^{-1}\left(\bm{\sigma}\bm{\rho}\right)\left(\bm{\sigma}_{Y}\frac{\nabla_{\mathbf{y}}V}{V}\right)$.
+    The Deep BSDE implementation then creates a policy by combining an analytical myopic demand with a network-based approximation of this tractable hedging term.
+Specifically, the core theoretical component $\bm{\sigma}_{Y}\frac{\nabla_{\mathbf{y}}V}{V}$ is approximated by the ratio of neural network outputs, $\frac{\bm{Z}^{Y}_{t}}{\hat{V}_{t}}$.
+[^8] In our BSDE parameterization, the network ZNet learns $\bm{Z}^{Y}_{t}\approx\nabla_{\mathbf{y}}V\cdot\bm{\sigma}_{Y}$, where $\bm{\sigma}_{Y}$ is the diagonal matrix of factor volatilities.
+The ratio in the policy formula thus correctly approximates the required theoretical term $\frac{\nabla_{\mathbf{y}}V\cdot\bm{\sigma}_{Y}}{V}$.
+This leads to the final implemented formula:
+    $\bm{\pi}_{t}=\underbrace{\frac{1}{\gamma}\Sigma^{-1}\left(\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}\right)}_{\text{Myopic Demand}}\quad+\underbrace{\frac{1}{\gamma}\Sigma^{-1}\left(\bm{\sigma}\bm{\rho}\frac{\bm{Z}^{Y}_{t}}{\hat{V}_{t}}\right)}_{\text{Intertemporal Hedging Demand}}.$ (28)
+    This structure is explicitly designed to capture both demand components, but as shown in our numerical results, the accuracy of this network-based approximation can be a practical point of failure.
+  • Loss Function: The training objective is to accurately match the BSDE’s terminal condition.
+For each sample path starting at state $(t_{0},X_{t_{0}},\mathbf{Y}_{t_{0}})$, the ‘V0Net’ provides an initial value estimate, $\hat{V}_{t_{0}}$.
+This single initial estimate is then evolved forward to the terminal time $T$ using the discrete-time BSDE formula, which incorporates the outputs of the ‘ZNet’ at each intermediate time step.
+This process yields a simulated terminal value for that specific path, which we denote $\hat{V}_{T}^{\text{sim}}$.
+The loss is then calculated as the mean squared error between this simulated terminal value and the true terminal condition—the utility of the wealth $U(X_{T})$ realized at the end of that same path:
+    $\mathcal{L}=\mathbb{E}\left[\left(\hat{V}_{T}^{\text{sim}}-U(X_{T})\right)^{2}\right].$
+    The parameters of both ‘V0Net’ and ‘ZNet’ are optimized to minimize this terminal loss, effectively forcing the initial prediction $\hat{V}_{t_{0}}$ to be consistent with the terminal utility value through the lens of the BSDE dynamics.
   • Training Details: The networks are trained for $10,000$ epochs using the Adam optimizer (initial learning rate $3\times 10^{-5}$, with a multi-step decay scheduler) and a batch size of $1,024$.

### v5

@@
 For this setting, we specialize the general model by assuming no intermediate consumption (i.e., $C_{t}=0$) and focus exclusively on utility from terminal wealth, with $\kappa=1$, so that the objective becomes $J=\mathbb{E}[U(X_{T})]$.
+We deliberately begin our investigation with a relatively short investment horizon of $T=1.5$ years.
+This strategic choice enables a focused evaluation of the intrinsic performance and convergence characteristics of our proposed frameworks, isolating their behavior from the compounding numerical challenges inherent in long-horizon simulations.
+By establishing a clear performance baseline in this controlled setting, we lay the groundwork for a more rigorous analysis of long-term robustness in Section 4.3, where we introduce specific techniques to address those challenges.
 This affine setting admits a semi-analytical benchmark solution, derived from the associated HJB equation, which we use as a reference for evaluation (see Appendix B.2).
@@
 Notably, in the $n=50,k=2$ case, Baseline PG-DPO ($3.875\times 10^{-1}$) outperforms Deep BSDE ($6.340\times 10^{-1}$).
-This pronounced degradation with increasing $n$, even for small $k$, points to a distinct source of difficulty—one not solely attributable to function approximation in high-dimensional state spaces.
-A likely culprit is the numerical instability introduced by the asset covariance matrix $\Sigma$.
-As $n$ increases, the $n\times n$ matrix $\Sigma$ becomes more prone to ill-conditioning (i.e., a high condition number).
-Since optimal portfolio policies in the PMP framework (Eq.
-(11)) involve computing $\Sigma^{-1}$, small estimation errors in the costates or model parameters can be significantly amplified, resulting in large policy errors.
-This observation suggests that the “curse” associated with large $n$ is not merely about learning a higher-dimensional function, but also reflects the increasing fragility of the optimization problem itself.
-The structure of the problem becomes numerically unstable, which adversely affects any method that explicitly or implicitly depends on $\Sigma$ or its inverse.
-These findings highlight that while neural network–based methods have made substantial progress in managing high-dimensional state spaces (associated with large $k$), challenges related to a large number of decision variables (i.e., assets $n$)—particularly in the presence of ill-conditioned covariance matrices—remain a significant concern.
-Addressing this “asset-dimension curse” may require incorporating techniques such as covariance matrix shrinkage, spectral regularization, or other forms of preconditioning within the optimization problem.
 Table 3: Summary of Minimum Policy RMSE and Convergence Iterations for $n=1$ (Varying $k$)
@@
 Deep BSDE $6.340\times 10^{-1}$ 3800
+This pronounced degradation with increasing $n$, even for small $k$, points to a distinct source of difficulty—one not solely attributable to function approximation in high-dimensional state spaces.
+A likely culprit is the numerical instability introduced by the asset covariance matrix $\Sigma$.
+As $n$ increases, the $n\times n$ matrix $\Sigma$ becomes more prone to ill-conditioning (i.e., a high condition number).
+Since optimal portfolio policies in the PMP framework (Eq.
+(11)) involve computing $\Sigma^{-1}$, small estimation errors in the costates or model parameters can be significantly amplified, resulting in large policy errors.
+This observation suggests that the “curse” associated with large $n$ is not merely about learning a higher-dimensional function, but also reflects the increasing fragility of the optimization problem itself.
+The structure of the problem becomes numerically unstable, which adversely affects any method that explicitly or implicitly depends on $\Sigma$ or its inverse.
+These findings highlight that while neural network–based methods have made substantial progress in managing high-dimensional state spaces (associated with large $k$), challenges related to a large number of decision variables (i.e., assets $n$)—particularly in the presence of ill-conditioned covariance matrices—remain a significant concern.
+Addressing this “asset-dimension curse” may require incorporating techniques such as covariance matrix shrinkage, spectral regularization, or other forms of preconditioning within the optimization problem.
 4.2 High-Dimensional Performance Analysis ($n=50,k=10$)
@@
 The remaining inaccuracies in the hedging component point to a pressing need for addressing the ill-conditioning of $\Sigma$ in very high asset dimensions—potentially through regularization, shrinkage techniques, or architecture-aware preconditioning.
+4.3 Robustness and Accuracy in Long-Horizon Portfolio Choice
+While the results in the preceding sections demonstrate the effectiveness of the PG-DPO framework for a relatively short $T=1.5$ year horizon, a crucial test of any dynamic asset allocation method is its robustness and accuracy over long-term investment periods.
+Long-horizon optimization poses a significant challenge, as the baseline PG-DPO algorithm is prone to the compounding effects of variance and bias accumulation over extended simulations.
+The error sources identified in Theorem 1, namely discretization ($\propto\Delta t$) and Monte Carlo sampling ($\propto 1/\sqrt{M}$), become critically important as the simulation horizon lengthens.
+To build a truly robust framework capable of tackling these long-term challenges, we introduce three key algorithmic enhancements designed to fundamentally improve accuracy and stability:
+  1. 1.
+    Residual Learning for Hedging Demand: We restructure the policy network to exclusively learn the complex intertemporal hedging component, while the analytically known myopic demand is computed separately (see (Silver et al., , 2018; Johannink et al., , 2019)).
+This separation is a natural use of domain knowledge in financial economics, where the myopic component is well-understood, and the accurate estimation of the hedging demand is the primary challenge.
+This residual learning approach allows the network to dedicate its full capacity to the most difficult aspect of the optimization problem.
+  2. 2.
+    Control Variates: To suppress the variance inherent in long Monte Carlo rollouts, which is the most critical issue in long-term simulations, we introduce a control variate.
+A key advantage of our model-based approach is the ability to seamlessly integrate such classical variance reduction techniques from the financial engineering literature (Glasserman, , 2004).
+By using the analytically tractable myopic policy as a control variate, we can significantly stabilize the learning process.
+  3. 3.
+    Richardson Extrapolation: To aggressively reduce time discretization error, we employ Richardson extrapolation.
+The known structure of the underlying stochastic differential equations in our model-based setting permits the straightforward application of established numerical methods like this one.
+By combining simulations at time steps $\Delta t$ and $0.5\Delta t$, we achieve a higher order of accuracy, crucial for maintaining fidelity over long time periods (Glasserman, , 2004).
+Before deploying these techniques on a long-horizon problem, we first validated their impact on our original $T=1.5$ year experimental setup.
+This step ensures that the enhancements provide a fundamental improvement in performance.
+As shown in Table 5, the new techniques yield a dramatic reduction in policy RMSE for both the Baseline PG-DPO and the P-PGDPO algorithms.
+Table 5: RMSE Improvement with Advanced Techniques for $T=1.5$ Horizon
+n k New Baseline RMSE Baseline Improvement New P-PGDPO RMSE P-PGDPO Improvement
+1 1 $5.10\times 10^{-3}$ 6.4×6.4$\times$6.4 × $4.80\times 10^{-5}$ 23.3×23.3$\times$23.3 ×
+1 2 $8.29\times 10^{-4}$ 57.5×57.5$\times$57.5 × $6.30\times 10^{-5}$ 3.16×3.16$\times$3.16 ×
+1 5 $1.04\times 10^{-3}$ 35.7×35.7$\times$35.7 × $1.00\times 10^{-5}$ 25.0×25.0$\times$25.0 ×
+1 10 $5.54\times 10^{-4}$ 31.6×31.6$\times$31.6 × $2.23\times 10^{-4}$ 2.11×2.$1$
+10 2 $1.49\times 10^{-2}$ $\times7$ $8.33\times 10^{-4}$ 4.38×4.38$\times$4.38 ×
+10 10 $5.07\times 10^{-3}$ $\times9$ $2.21\times 10^{-3}$ 1.27×1.27$\times$1.27 ×
+50 2 $8.07\times 10^{-2}$ 4.8×4.8$\times$4.8 × $8.40\times 10^{-3}$ 8.73×8.73$\times$8.73 ×
+50 10 $3.38\times 10^{-2}$ 3.2×3.2$\times$3.2 × $1.07\times 10^{-2}$ 1.42×1.42$\times$1.42 ×
+Note: This table compares the minimum policy RMSE of the enhanced algorithms against the original results from Tables 3 and 4.
+The ‘Improvement’ column indicates the factor by which the RMSE was reduced (e.g., Original RMSE / New RMSE).
+The significant performance gains, with improvement factors of up to 57.5×57.5$\times$57.5 ×, confirm that these techniques are not merely incremental adjustments but constitute a fundamental upgrade to the algorithm’s core.
+This successful validation provides a strong foundation for tackling the more demanding long-horizon problem.
+Armed with a demonstrably more robust algorithm, we now address a long-horizon portfolio choice problem with a maturity of $T=20$ years.
+As the investment horizon extends towards infinity, the optimal policy is expected to become stationary, i.e., independent of the time-to-maturity.
+While this stationary policy is a simpler functional form for the network to learn, the extended simulation required to find it amplifies the challenges of variance and error control.
+Our enhanced framework is designed precisely for this scenario.
+Table 6 presents the final policy RMSE for the 20-year experiment.
+The results confirm the success of our approach, showing remarkably low errors even in high dimensions.
+In the $n=50,k=10$ case, the P-PGDPO method achieves an error of just $3.85\times 10^{-3}$.
+This result is even superior to that of the 1.5-year experiment, a phenomenon explained by the synergy of a more powerful algorithm converging to a simpler, stationary target policy.
+Table 6: Minimum Policy RMSE for Long‑Horizon Problem ($T=20$)
+Baseline PG-DPO P-PGDPO
+$n$ $k$ Min. Policy RMSE Iterations at Min. Min. Policy RMSE Iterations at Min.
+1 1 0.00169 400 0.000041 400
+5 3 0.00876 200 0.000352 1
+30 5 0.07220 1000 0.007030 200
+50 10 0.05020 2600 0.003850 200
+Figure 5 provides a compelling visual proof of the framework’s robustness.
+An important characteristic of the long-horizon benchmark solution is that the intertemporal hedging demand exhibits strong time-dependency for short maturities (e.g., within 1.5 years), but becomes effectively time-invariant for longer maturities, converging to a stationary policy.
+Furthermore, creating a reliable interpolation for the 20-year benchmark is challenging, as the optimal policy changes rapidly for short maturities where evaluation points are sparse.
+Therefore, the figure presents scatter plots of the raw policy function outputs to ensure a faithful visualization of the results.
+The policies produced by P-PGDPO are visually indistinguishable from the benchmark solution, perfectly capturing the intricate, non-linear dependencies on the state variables.
+This stands in stark contrast to the Baseline PG-DPO, which fails to learn the correct policy structure.
+The comparison makes clear the qualitative difference between a rough approximation, which may achieve a low but misleading error metric, and a truly accurate approximation that captures both the value and the complex structure of the optimal policy.
+These results underscore the power of our enhanced P-PGDPO framework to deliver robust, high-fidelity solutions for challenging long-horizon control problems.
+(a) Baseline PG-DPO (Asset 1)
+(b) Baseline PG-DPO (Asset 2)
+(c) P-PGDPO (Asset 1)
+(d) P-PGDPO (Asset 2)
+(e) Benchmark Solution (Asset 1)
+(f) Benchmark Solution (Asset 2)
+Figure 5: Comparison of learned policies for the 20-year horizon ($n=50,k=10$), plotted against the first state factor $Y_{1}$.
+Left column: Asset 1.
+Right column: Asset 2.
+The P-PGDPO results show a near-perfect match with the benchmark solution, whereas the Baseline PG-DPO fails to capture the correct policy structure.
 4.4 Non-affine Asset Return Dynamics
-In this section, we extend our numerical investigation to a setting where the asset prices evolve according to a non-affine process.
-We assume that the state factors state factors, $\mathbf{Y}_{t}=[Y_{1,t},\dots,Y_{k,t}]^{T}$, follow the same multivariate Ornstein-Uhlenbeck (OU) process as described in Section 2.1.
+While the analysis in Section 4.3 demonstrated the framework’s robustness over long horizons, we now pivot to assess its flexibility with non-affine model dynamics.
+This investigation is a direct extension of our initial experiments on affine models (Sections 4.1 and 4.2).
+To ensure a clear and direct comparison, we revert to the T=1.5 year horizon and the foundational P-PGDPO algorithm.
+This deliberate choice allows us to attribute any changes in the optimal policy solely to the introduction of non-linearity, isolating its effects from the advanced techniques introduced in the long-horizon study.
+We assume that the state factors, $\mathbf{Y}_{t}=[Y_{1,t},\dots,Y_{k,t}]^{T}$, follow the same multivariate Ornstein-Uhlenbeck (OU) process as described in Section 2.1.
 The expected returns of risky assets, $\bm{\mu}_{S}(t,\mathbf{Y}_{t})$, are given by
@@
 All experiments in this subsection are conducted for the most challenging setting, $n=50$ assets and $k=10$ state factors, utilizing the P-PGDPO method.
-To quantify the deviation of the P-PGDPO policy in this non-affine setting from the affine benchmark solution as $\beta_{\text{norm}}$ changes, Table 5 presents the Root Mean Squared Error (RMSE) of the P-PGDPO learned policy, $\bm{\pi}(\beta_{\text{norm}})$, relative to the affine analytical solution ($\bm{\pi}(\beta_{\text{norm}}=0)$).
-For the results presented in this table, all P-PGDPO models were trained for 10,000 iterations.
-In the non-affine case ($\beta_{\text{norm}}>0$), where an analytical solution to the control problem is unavailable, this fixed number of iterations is used for evaluation, as opposed to selecting the number of iterations based on the minimum error against a known benchmark in earlier sections.
 Table 7: Policy RMSE of P-PGDPO for the model with non-affine risk premia vs. affine benchmark ($n=50,k=10$). Results reported after 10,000 training iterations.
@@
 4.0 0.33570
-The results in Table 5 clearly demonstrate two key points.
-First, for $\beta_{\text{norm}}=0.0$, the P-PGDPO method achieves a low RMSE of $0.04701$.
-This is consistent with its strong performance in the affine setting (reported in Section 4.1) and indicates an accurate approximation of the affine analytical solution, as expected when the non-linear component vanishes.
-Second, as $\beta_{\text{norm}}$ increases from $0.1$ to $4.0$, the difference of the P-PGDPO policy from the fixed affine benchmark, measured by RMSE, monotonically increases.
-This trend signifies that the optimal policy systematically diverges from the affine optimal policy as the magnitude of the non-linear component grows.
-The consistent increase suggests that P-PGDPO is stably adapting the policy to this increasing non-linearity in the market price of risk, rather than producing erratic or unstable solutions.
 (a) Analytical solution ($\beta_{\text{norm}}=0$).
@@
 Note the systematic change in policy structure and value ranges (indicated by colorbars) as $\beta_{\text{norm}}$ increases.
+To quantify the deviation of the P-PGDPO policy in this non-affine setting from the affine benchmark solution as $\beta_{\text{norm}}$ changes, Table 7 presents the Root Mean Squared Error (RMSE) of the P-PGDPO learned policy, $\bm{\pi}(\beta_{\text{norm}})$, relative to the affine analytical solution ($\bm{\pi}(\beta_{\text{norm}}=0)$).
+For the results presented in this table, all P-PGDPO models were trained for 10,000 iterations.
+In the non-affine case ($\beta_{\text{norm}}>0$), where an analytical solution to the control problem is unavailable, this fixed number of iterations is used for evaluation, as opposed to selecting the number of iterations based on the minimum error against a known benchmark in earlier sections.
+The results in Table 7 clearly demonstrate two key points.
+First, for $\beta_{\text{norm}}=0.0$, the P-PGDPO method achieves a low RMSE of $0.04701$.
+This is consistent with its strong performance in the affine setting (reported in Section 4.1) and indicates an accurate approximation of the affine analytical solution, as expected when the non-linear component vanishes.
+Second, as $\beta_{\text{norm}}$ increases from $0.1$ to $4.0$, the difference of the P-PGDPO policy from the fixed affine benchmark, measured by RMSE, monotonically increases.
+This trend signifies that the optimal policy systematically diverges from the affine optimal policy as the magnitude of the non-linear component grows.
+The consistent increase suggests that P-PGDPO is stably adapting the policy to this increasing non-linearity in the market price of risk, rather than producing erratic or unstable solutions.
 To complement these quantitative findings, Figure 6 provides a visual comparison of the learned policies.
 Figure 6(a) displays the analytical solution for the affine model ($\beta_{\text{norm}}=0$).
-Figures 5(b)–5(e) show the P-PGDPO learned policies for models where $\beta_{\text{norm}}=0.0,0.1,1.0,$ and $2.0$, respectively.
+Figures 6(b)–6(e) show the P-PGDPO learned policies for models where $\beta_{\text{norm}}=0.0,0.1,1.0,$ and $2.0$, respectively.
 These plots depict the policy for the first asset ($\pi_{1}$) as a function of the first state factor ($Y_{1}$) and time-to-maturity ($T-t$), holding other state factors at their long-term means.

### v6

@@
 Abstract
-Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality.
-Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond a few state variables ($\sim$3-6 limit in prior studies).
-To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework.
-Our framework accurately captures both myopic demand and complex intertemporal hedging demands, a feat often elusive for other methods in high-dimensional settings.
-P-PGDPO delivers near-optimal policies, offering a practical and powerful alternative for a broad class of high-dimensional continuous-time control problems.
-PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to directly inform neural network-based policy learning.
-A key contribution is our highly efficient Projected PG-DPO (P-PGDPO) variant.
-This approach uniquely utilizes BPTT to obtain rapidly stabilizing estimates of the Pontryagin costates and their crucial derivatives with respect to the state variables.
-These estimates are then analytically projected onto the manifold of optimal controls dictated by PMP’s first-order conditions, significantly reducing training overhead and enhancing accuracy.
-This enables a breakthrough in scalability: numerical experiments demonstrate that P-PGDPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables).
+We introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework for high-dimensional continuous-time portfolio choice.
+Our approach combines Pontryagin’s Maximum Principle (PMP) with backpropagation through time (BPTT) to directly inform neural network–based policy learning, enabling accurate recovery of both myopic and intertemporal hedging demands—an aspect often missed by existing methods.
+Building on this, we develop the Projected PG-DPO (P-PGDPO) variant, which achieves near-optimal policies with substantially improved efficiency.
+P-PGDPO leverages rapidly stabilizing costate estimates from BPTT and analytically projects them onto PMP’s first-order conditions, reducing training overhead while improving precision.
+Numerical experiments show that PG-DPO matches or exceeds the accuracy of Deep BSDE, while P-PGDPO delivers significantly higher precision and scalability.
+By explicitly incorporating time-to-maturity, our framework naturally applies to finite-horizon problems and captures horizon-dependent effects, with the long-horizon case emerging as a stationary special case.
+Keywords: Multi-asset Portfolio choice, Pontryagin’s Maximum Principle, direct Policy Optimization, model-based reinforcement learning
+[…]
+[A-]Online-Appendix
 1 Introduction
-The theory of dynamic consumption and portfolio choice has been a cornerstone of financial economics, providing fundamental insights into asset pricing and long-term investment behavior (Merton, , 1973; Campbell and Viceira, , 1999, 2001).
-Despite its theoretical elegance and contributions to our understanding, the theory faces a major obstacle in practical implementation: the lack of scalable methods for solving high-dimensional portfolio choice problems involving many assets and state variables.
-The Bellman equation associated with such problems is subject to the curse of dimensionality, and obtaining solutions in realistic settings is often infeasible.
-Existing solution methods are typically complex, ad hoc, and do not scale well with dimensionality.
-This is one reason why Markowitz’s static mean-variance framework remains dominant in practice, despite its limitations and the greater realism offered by dynamic models.
-[^1] Cochrane wrote “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice.
-” (Cochrane, , 2022, p.4)
-In this paper, we make a technological contribution by proposing a solution approach that leverages machine learning with neural networks to solve high-dimensional dynamic portfolio choice problems, where both the number of assets $n$ and the number of state variables $k$ are large.
-There are two primary challenges inherent in such problems.
-The first is the well-known curse of dimensionality: computational methods often fail to scale efficiently as the dimensionality of the problem increases in $n$ and $k$.
-The second, and more subtle, challenge arises from the specific structure of dynamic portfolio choice itself.
-Unlike other applications of reinforcement learning—such as Atari, Go, or autonomous driving—where achieving a near-optimal policy is often sufficient for strong empirical performance, portfolio optimization requires much greater precision.
-A well-established theoretical benchmark exists, dating back to the seminal work of Merton, (1973), which demonstrates that the optimal portfolio can be decomposed into two critical components: the myopic demand and the intertemporal hedging demand.
+The continuous-time consumption and portfolio choice problem has been a cornerstone of financial economics, offering fundamental insights into asset pricing and long-term investment behavior (Merton, , 1973; Campbell and Viceira, , 1999, 2001).
+Despite its theoretical elegance, the practical implementation of this framework has remained exceptionally challenging.
+Two obstacles are particularly salient.
+First, portfolio choice problems involving many risky assets and state variables are subject to the curse of dimensionality (Bellman, , 1966).
+Traditional dynamic programming (DP) approaches, built on solving the Hamilton–Jacobi–Bellman (HJB) equation, readily become intractable as dimensionality increases.
+This is one reason why Markowitz’s static mean–variance framework still remains dominant in practice, despite its limitations and the greater realism offered by dynamic models.
+[^1] Cochrane wrote: “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice.
+(Cochrane, , 2022, p.4)” The appeal of the mean-variance model lies in its tractability, as it reduces portfolio choice to a single-period quadratic optimization problem that remains computationally feasible even with many assets.
+However, it ignores intertemporal hedging motives, relies on a restrictive single-period formulation, and assumes normally distributed returns so that only means and variances matter.
+Moreover, it disregards the dynamic evolution of investment opportunities and practical frictions such as transaction costs or portfolio constraints, and suffers from a lack of time consistency in multi-period settings.
+Consequently, most continuous-time models in the literature have been confined to highly stylized, low-dimensional settings.
+Second, even when solutions are available, they often fail to recover the structural decomposition of the optimal policy.
+A well-established theoretical benchmark, dating back to the seminal work of Merton, (1973), demonstrates that the optimal portfolio can be decomposed into two critical components: the myopic demand and the intertemporal hedging demand.
 The myopic component represents the demand for a mean-variance efficient portfolio given current investment opportunities, while the intertemporal component hedges against expected future changes in those opportunities driven by evolving economic conditions.
-Accurately identifying both components is essential for evaluating investor behavior and designing long-term investment strategies—yet it remains a formidable challenge, particularly in high-dimensional settings.
-Our proposed approach is specifically designed to address both challenges: overcoming the curse of dimensionality and enabling the accurate decomposition of optimal portfolios into their myopic and intertemporal components.
-We adopt a continuous-time framework in which asset return means and covariances are driven by a $k$-dimensional state vector, with the state variables evolving as Itô diffusions, following Merton, (1973).
-This modeling approach benefits from the rich toolkit of stochastic calculus—tools generally unavailable in discrete-time formulations.
-[^2] See Duarte et al., (2024) for a discussion of the advantages of continuous-time models over discrete-time models.
-In the continuous-time setting, (quasi-)analytic solutions are available under affine dynamics, where the Hamilton–Jacobi–Bellman (HJB) equation reduces to a system of ordinary differential equations (Kim and Omberg, , 1996; Liu, , 2007; Buraschi et al., , 2010).
-However, for general non-affine dynamics, no universal solution method exists.
-To address this limitation, we introduce a machine learning–based approach that applies to arbitrary market dynamics.
-Neural networks, known for their ability to approximate complex high-dimensional functions, are used to parameterize the policy directly, enabling us to capture intricate dependencies on both wealth and the exogenous state variables—even when the dimension $k$ is large.
-While we focus on Brownian motion as the source of return risk, the methods can be extended to accommodate jumps.
-Furthermore, if a discrete-time model is viewed as an approximation of a continuous-time model, our approach can be adapted accordingly.
-Our approach leverages Pontryagin’s Maximum Principle (PMP) as its theoretical foundation, in contrast to the more commonly used dynamic programming (DP) framework based on the Hamilton–Jacobi–Bellman (HJB) equation.
-Building on PMP, we propose two methods: Pontryagin-Guided Direct Policy Optimization (PGDPO) and its variant, Projected PGDPO (P-PGDPO).
-PMP characterizes optimal controls through a coupled system of forward and backward differential equations.
-The forward equation governs the evolution of wealth and state variables driving asset returns, which we implement via Monte Carlo simulation.
-This allows for direct computation of the expected utility and avoids the need to estimate the value function across future periods—a core but computationally intensive requirement of DP-based approaches.
-The backward equation governs the dynamics of the costates (or adjoint variables), which correspond to the first-order partial derivatives of the value function with respect to the state variables.
-Modern machine learning frameworks such as PyTorch allow us to compute these costates and their derivatives, the second-order partial derivatives of the value function, automatically through backpropagation through time (BPTT) and automatic differentiation.
-We demonstrate that a single forward simulation pass, coupled with BPTT, yields unbiased estimates of the costate processes required by PMP.
-As a result, our approach avoids the need for explicit solution of the backward equation or approximation of the value function, thereby sidestepping major bottlenecks of DP.
-Costates—as well as the value function—emerge as natural byproducts of the learning process.
-In PGDPO, the control policy—comprising consumption and portfolio allocations—is parameterized by neural networks.
-The expected utility, defined as the discounted sum of utility from consumption and terminal wealth, is then directly maximized using stochastic gradient ascent, where gradients are computed using BPTT.
-A key computational advantage of our approach lies in its complete avoidance of HJB residual minimization, where the residual is defined as the cumulative squared error of the HJB equation.
-Computing the HJB residual typically requires evaluating first- and second-order partial derivatives of the value function at each time–state grid point and aggregating the errors over time, incurring high computational cost.
-Duarte et al., (2024) mitigate this by replacing full gradient computation with the second derivative of a carefully constructed univariate function.
-In contrast, our method bypasses HJB residual computation entirely, resulting in significant computational savings.
-Conceptually, our approach shares similarities with recent methods that parameterize control policies via neural networks and directly optimize the expected objective function through simulation and gradient ascent, thereby circumventing the need to solve the HJB equation (Zhang and Zhou, , 2019; Kopeliovich and Pokojovy, , 2024, e.g., for portfolio problems with transaction costs or Heston dynamics).
-However, our PMP-based PGDPO framework distinguishes itself by being explicitly guided by continuous-time optimality conditions derived from PMP.
-Moreover, unlike standard reinforcement learning methods such as Proximal Policy Optimization (PPO) or Trust Region Policy Optimization (TRPO), which rely on sampled estimates of policy gradients, our method computes exact gradients via BPTT.
-This yields lower-variance updates and enhances computational efficiency, particularly in high-dimensional settings (Kushner and Yin, , 2003; Borkar and Borkar, , 2008).
-The second method, P-PGDPO, builds on PGDPO by substituting the estimated derivatives of the value function directly into the structural first-order conditions from PMP.
-These conditions take the form derived originally by Merton, (1973), explicitly decomposing the optimal portfolio into myopic and intertemporal hedging demands.
-We demonstrate the numerical accuracy of our methods by comparing the solutions they produce to a quasi-analytic benchmark under affine dynamics, where the true solution can be obtained by solving a system of ordinary differential equations.
-The two components of the optimal portfolio—myopic and intertemporal hedging demands—are not merely of theoretical interest; their accurate identification requires high-fidelity solutions.
-As shown by Constantinides, (1986), transaction costs generate only second-order utility losses, allowing portfolios to deviate substantially from the frictionless optimum while still appearing nearly optimal in terms of utility.
-Similarly, Cochrane, (1988) demonstrates that first-order deviations from the optimal policy can result in only second-order losses in utility.
-This disconnect between utility value and policy structure poses a major challenge for traditional reinforcement learning algorithms.
-These methods often converge to policies that are near-optimal in utility but structurally far from the true optimum—particularly in high-dimensional settings.
-In such cases, neural networks must search for optimal policies over an $n$-dimensional control space, where $n$ denotes the number of risky assets.
-The resulting exponential growth in complexity renders standard policy optimization techniques computationally infeasible as $n$ increases.
-Our numerical experiments confirm that increasing the number of risky assets $n$ presents a significantly greater challenge than increasing the number of state variables $k$ in high-dimensional settings.
-The P-PGDPO method is motivated by the empirical observation that value function derivatives—i.e., costates—are typically easier to learn than the policy functions themselves.
-In practice, these partial derivatives tend to stabilize much earlier during training.
-Once the costate estimates have stabilized—usually after a short warm-up phase—the method can bypass the potentially suboptimal outputs of the policy networks.
-Instead, the costates and their derivatives are directly substituted into the first-order conditions (FOCs) from PMP, whose structural form closely mirrors the conditions originally derived by Merton, (1973).
-This effectively projects the learned costate information onto the manifold of controls that satisfy PMP’s necessary conditions.
-The resulting control policy adheres closely to the structural form of the true optimal solution dictated by PMP.
-It is recovered quasi-analytically at each time–state point, requiring minimal additional training and incurring substantially reduced computational cost.
-Our PMP-based PGDPO framework thus leverages BPTT-derived estimates of costates and their derivatives, enhancing scalability while providing a principled mechanism for enforcing first-order optimality conditions.
-We provide a theoretical justification for the P-PGDPO method in Theorem 1.
+Accurate identification of both components is crucial for evaluating investor behavior and designing long-term investment strategies.
+Recent advances in machine learning have introduced new PDE-based solvers, including Physics-Informed Neural Networks (PINNs), the Deep BSDE method (Han et al., , 2018), and Deep Galerkin Networks (DGNs).
+While these approaches help mitigate the curse of dimensionality, they focus on approximating the value function and are not designed to recover the full structure of the optimal portfolio, particularly the intertemporal hedging demand.
+Moreover, contributions such as Duarte et al., (2024) restrict attention to infinite-horizon problems, where the optimal solution is stationary and time-independent.
+Although suitable for certain theoretical contexts, such methods are ill-suited for practical applications such as retirement planning, endowment management, or finite-maturity products, where horizon effects and time-varying hedging demands play a central role.
+To address these limitations, we propose a new framework, Pontryagin-Guided Direct Policy Optimization (PG-DPO).
+Our approach integrates Pontryagin’s Maximum Principle (PMP) with policy networks trained via backpropagation through time (BPTT), enabling scalable and structurally accurate solutions to high-dimensional continuous-time portfolio choice problems.
+PMP characterizes optimal controls through a coupled system of forward and backward equations.
+The forward equation governs the evolution of wealth and state variables that drive asset returns, which we simulate using Monte Carlo methods.
+This allows direct computation of expected utility and removes the need to estimate the value function over future periods—a core but computationally intensive step in DP-based approaches.
+Accordingly, our framework relies solely on policy networks, without requiring a value network.
+The backward equation governs the dynamics of the costates (or adjoint variables), corresponding to the first-order partial derivatives of the value function with respect to the state variables.
+Modern machine learning frameworks such as PyTorch allow these costates and their derivatives—including second-order partial derivatives of the value function—to be computed automatically via BPTT and automatic differentiation.
+We show that a single forward simulation, combined with BPTT, yields unbiased estimates of the costate processes required by PMP.
+As a result, PG-DPO avoids explicit solution of the backward equation, thereby sidestepping the major bottlenecks of dynamic programming.
+Both the costates and the value function emerge naturally as byproducts of the learning process.
+A central innovation is the Projected PG-DPO (P-PGDPO) variant.
+P-PGDPO builds on PG-DPO by substituting the estimated derivatives of the value function directly into the structural first-order conditions (FOC) of PMP.
+During training, costate estimates typically stabilize well before the policy networks are fully converged.
+P-PGDPO leverages this property by projecting the stabilized costates and their derivatives onto the analytical manifold defined by PMP’s first-order conditions.
+The algorithm proceeds in two stages.
+In Stage 1, the networks are trained until the costate estimates and their derivatives stabilize.
+In Stage 2, these estimates are substituted into the first-order conditions to construct the optimal policies.
+The resulting controls adhere closely to the structural form of the true solution, as dictated by PMP, and are recovered quasi-analytically at each time–state point.
+This requires minimal additional training while substantially reducing computational cost.
+By disentangling costate estimation from policy construction, P-PGDPO achieves significant gains in both accuracy and scalability.
+We provide a theoretical justification for the P-PGDPO method in Theorem 4.2.
 When the objective function surface is relatively flat near the optimum, the first-order condition (FOC) violation for the policy obtained in the warm-up stage tends to be small.
-The theorem states that if this FOC deviation is small and the costate estimates obtained via BPTT are sufficiently accurate, then the control policy recovered in the second stage is close to the true optimal policy.
-We demonstrate the accuracy of our numerical methods by solving a problem in which the state process follows a multi-factor Ornstein–Uhlenbeck (OU) process—a special case of affine dynamics—for which quasi-analytic benchmark solutions can be obtained by solving a system of ordinary differential equations (see Section 2.3.2).
-We also compare our methods with the Deep Backward Stochastic Differential Equation (Deep BSDE) approach, which is representative of existing techniques for solving high-dimensional partial differential equations, particularly the Hamilton–Jacobi–Bellman (HJB) equations arising in DP.
-In addition, we apply our second method, P-PGDPO, to a non-affine setting by adding non-linear terms to the expected returns of risky assets, thereby demonstrating its applicability to a broader class of dynamics.
-Our numerical experiments (Section 4) show that our first method, PGDPO, produces solutions comparable to those of the Deep BSDE method in terms of mean-squared errors.
-In contrast, the second method, P-PGDPO, yields significantly smaller errors.
-This improvement arises because PGDPO inherits the common limitation of policy optimization algorithms—difficulty in discovering the true optimum—whereas P-PGDPO overcomes this issue by directly exploiting the structural form of the first-order conditions from Pontryagin’s Maximum Principle (PMP).
-P-PGDPO not only scales to problems involving dozens of assets and state variables (e.g., $n=50$, $k=10$), but also accurately disentangles and reconstructs complex intertemporal hedging demands.
-Furthermore, our results demonstrate that the method generates stable solutions in non-affine settings and converges smoothly to the benchmark affine solution as the coefficients of the nonlinear terms approach zero.
-By contrast, the Deep BSDE approach, while successful in other high-dimensional applications such as option pricing and hedging, fails to capture the full structure of the optimal policy in our setting—particularly the intertemporal hedging components.
-The method’s superior performance—both in scalability and in fidelity to the full structure of the optimal policy—highlights the transformative potential of leveraging PMP through BPTT-derived costate estimates for a broad class of high-dimensional control problems that admit stable costate representations.
-There is a substantial literature on dynamic portfolio choice, aimed at obtaining concrete solutions either analytically or numerically.
-With the exception of the continuous-time models with affine dynamics discussed above, most existing contributions are formulated in discrete time.
-For example, Campbell and Viceira, (1999) and Campbell and Viceira, (2001) employ log-linear approximations to the dynamic budget constraint to obtain approximate closed-form solutions.
-Balduzzi and Lynch, (1999), Lynch and Balduzzi, (2000), and Lynch, (2001) apply discrete-space approximations to solve portfolio choice problems with transaction costs.
-Garlappi and Skoulakis, (2010) develop a DP approach based on state variable decomposition.
-Brandt et al., (2005) propose a Monte Carlo simulation method for a discrete-time portfollio choice model that projects simulated values of relevant functions—such as the marginal value of wealth—onto a pre-specified set of basis functions.
-In this sense, their method shares similarities with the forward simulation and auto-differentiation steps in our PG-DPO framework.
-However, their approach does not explicitly leverage PMP or utilize the representational power of neural networks.
-As a result, it is not easily scalable and lacks the flexibility for generalization.
-In contrast, our method is PMP-guided, scalable, and extensible to broader classes of control problems.
-In continuous time, there has been significant technological advancement in solving partial differential equations, which can be applied to the HJB equation arising from DP.
-Notable methods include Deep BSDE, Deep Galerkin, Physics-Informed Neural Networks (PINNs), Deep Splitting, and Deep Branching, among others (Han et al., , 2018; Sirignano and Spiliopoulos, , 2018; Raissi et al., , 2019; Beck et al., , 2021; Nguwi et al., , 2024).
-In parallel, substantial progress has been made in solving optimal control problems through the development of policy evaluation, policy gradient, $q$-learning, and PPO/TRPO methods in stochastic environments (Wang et al., , 2020; Jia and Zhou, 2022a, ; Jia and Zhou, 2022b, ; Jia and Zhou, , 2023; Zhao et al., , 2023).
-Wang and Zhou, (2020); Dai et al., (2023); Huang et al., (2024) apply these techniques to dynamic mean-variance analysis, while Dai et al., (2025) use them to solve a Merton-type portfolio optimization problem.
-These approaches involve learning the asset return process from observed return data.
-Geng et al., (2023) propose the Factor Learning Portfolio Optimization (FaLPO) method for solving a Merton-type portfolio optimization problem with stochastic factors.
-Their approach combines policy learning with model calibration.
-By contrast, we assume that the data-generating process is known and do not address the problem of learning from data.
-Instead, our focus is on a more general consumption and portfolio choice problem involving many risky assets and multiple state variables, whereas their analysis is limited to a small number of risky assets and state variables.
-Davey and Zheng, (2022) develop a BSDE-based method for portfolio optimization under portfolio constraints that is also based on the maximum principle, similar in spirit to our work.
-However, their method requires minimizing the terminal value error (as in Deep BSDE), maximizing the Hamiltonian at each state and time, and solving a BSDE system.
-In contrast, our method directly maximizes expected utility without penalizing approximation errors and without solving the backward equation explicitly, resulting in a significantly simpler and more efficient approach.
-More recently, Duarte et al., (2024) propose a machine learning approach for infinite-horizon continuous-time problems using DP, while Cheridito et al., (2025) develop a method that combines generalized policy iteration with PINNs to solve finite-horizon continuous-time DP problems.
-By contrast, our method is based on PMP, rather than DP.
-The aforementioned DP-based methods avoid the explicit evaluation of first- and second-order partial derivatives of the value function by instead computing the second derivative of a univariate function when assessing the HJB residual.
-In contrast, our approach computes these partial derivatives directly—via automatic differentiation of costate variables—providing crucial information for accurately capturing intertemporal hedging components.
-This ability to compute the value function’s derivatives with high accuracy enables precise estimation of both the intertemporal hedging terms and the myopic component of the optimal portfolio.
-[^3] Cochrane emphasizes the critical role of partial derivatives in determining the optimal portfolio for long-term investors, particularly when discussing the implementation challenges of Merton’s dynamic portfolio theory—a point we partially quoted earlier: “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice… What are the partial derivatives of the investor’s value function that guide state-variable investments?
+We can show that if this FOC deviation is small and the costate estimates obtained via BPTT are sufficiently accurate, then the control policy recovered in the second stage is close to the true optimal policy.
+We demonstrate the numerical accuracy of our methods by comparing them to analytic benchmarks under affine dynamics, where the true solution is available from previous work (Kim and Omberg, , 1996; Liu, , 2007).
+We further compare against the Deep BSDE approach, which is representative of existing techniques for solving high-dimensional HJB equations.
+Finally, we apply P-PGDPO to long-horizon investment problems and to models with non-affine return dynamics, thereby demonstrating its broader applicability.
+Our numerical experiments in Section 5 show that baseline PG-DPO produces solutions comparable to those of Deep BSDE in terms of mean squared error.
+However, the P-PGDPO variant yields significantly smaller errors.
+This improvement arises because PG-DPO, like other policy optimization algorithms, struggles to reliably discover the true optimum, whereas P-PGDPO overcomes this limitation by directly enforcing the structural first-order conditions of PMP.
+As a result, P-PGDPO not only scales to problems with many assets and state variables, but also accurately disentangles and reconstructs complex intertemporal hedging demands.
+Furthermore, it generates stable solutions in non-affine settings and converges smoothly to the affine benchmark as the nonlinear terms vanish.
+By contrast, the Deep BSDE approach—highly successful in other high-dimensional applications such as option pricing and hedging—fails to capture the full structure of the optimal policy, particularly the intertemporal hedging component.
+Conceptually, our approach is related to recent methods that parameterize control policies with neural networks and optimize the expected objective function directly through simulation and gradient ascent, thereby bypassing the need to solve the HJB equation (Zhang and Zhou, , 2019; Kopeliovich and Pokojovy, , 2024).
+Our PMP-based PG-DPO framework, however, is explicitly guided by the continuous-time optimality conditions of PMP.
+Unlike standard reinforcement learning algorithms such as Proximal Policy Optimization (PPO) or Trust Region Policy Optimization (TRPO), which rely on sampled policy gradient estimates, our method computes exact gradients via BPTT.
+This yields lower-variance updates and improved computational efficiency in high dimensions (Kushner and Yin, , 2003; Borkar and Borkar, , 2008).
+This work also contributes to the literature on dynamic portfolio choice, which seeks analytical or numerical solutions.
+Most contributions adopt a discrete-time formulation, with the exception of affine continuous-time models.
+For example, Campbell and Viceira, (1999, 2001) employ log-linear approximations, while Balduzzi and Lynch, (1999), Lynch and Balduzzi, (2000), and Lynch, (2001) use discrete-space approximations with transaction costs.
+Garlappi and Skoulakis, (2010) apply state-variable decomposition, and Brandt et al., (2005) propose a Monte Carlo projection of simulated marginal utility onto basis functions—an approach reminiscent of our forward-simulation step, though lacking the scalability and structural guidance of PMP.
+In continuous time, semi-analytical solutions are available under affine dynamics, where the HJB reduces to ODEs (Kim and Omberg, , 1996; Liu, , 2007; Buraschi et al., , 2010).
+For non-affine dynamics, no general solution exists.
+Parallel efforts apply reinforcement learning to continuous-time portfolio problems, including policy evaluation, policy gradients, $Q$-learning, and PPO/TRPO methods (Wang et al., , 2020; Jia and Zhou, 2022a, ; Jia and Zhou, 2022b, ; Jia and Zhou, , 2023; Zhao et al., , 2023).
+These have been applied to dynamic mean-variance analysis (Wang and Zhou, , 2020; Dai et al., , 2023; Huang et al., , 2024) and to Merton-type portfolio problems (Dai et al., , 2025).
+Such approaches learn return dynamics from data, but are typically restricted to few assets and factors.
+For example, Geng et al., (2023) propose the FaLPO method, which combines policy learning with model calibration but remains limited in scale.
+In a related line, Cong et al., (2022) introduce the AlphaPortfolio framework, which bypass preliminary estimation of return distributions and directly constructs portfolios using deep reinforcement learning, showing strong out-of-sample performance.
+By contrast, we assume the data-generating process is known and focus on solving the general continuous-time consumption and portfolio choice problem with many risky assets and multiple state variables.
+In a complementary direction, Doh et al., (2025) show that reinforcement learning can recover the rational expectations equilibrium in information-based asset market models without relying on explicit market data estimation, providing a theoretical foundation for the use of reinforcement learning in equilibrium settings.
+Other contributions closer to ours include Davey and Zheng, (2022), who develop a BSDE-based method with portfolio constraints; Duarte et al., (2024), who propose a machine learning approach for infinite-horizon settings; and Cheridito et al., (2025), who combine generalized policy iteration with PINNs for finite-horizon DP.
+These methods, however, either require explicit BSDE solutions, restrict attention to stationary infinite-horizon problems, or fail to recover the structural decomposition of myopic and intertemporal hedging demands in high dimensions.
+Our approach, by contrast, recovers partial derivatives directly through automatic differentiation of costates, enabling accurate estimation of both the myopic and hedging components.
+[^2] As Cochrane emphasizes, partial derivatives of the value function are critical for long-term portfolio choice yet notoriously difficult to obtain: “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice… What are the partial derivatives of the investor’s value function that guide state-variable investments?
 That is harder still…” (Cochrane, , 2022, pp.
 4–5).
-To our knowledge, the two aforementioned papers do not demonstrate accurate recovery of intertemporal hedging terms in settings with both large numbers of assets and state variables.
-To our knowledge, the existing literature has demonstrated the ability to solve problems involving at most a dozen risky assets and a dozen state variables—with two notable exceptions.
+The aforementioned literature does not demonstrate accurate recovery of intertemporal hedging terms when both the number of assets and state variables are large.
+[^3] To our knowledge, the existing literature has demonstrated the ability to solve problems involving at most a dozen risky assets and a dozen state variables, with two notable exceptions.
 Davey and Zheng, (2022) consider a setting with 20 assets, but under a constant investment opportunity set, i.e., without state variables influencing asset returns.
-[^4] Our companion paper demonstrates that our method is applicable to constrained portfolio optimization problems involving up to 1000 assets (Huh et al., , 2025).
-Cheridito et al., (2025) address a case with 25 assets, though without providing evidence regarding the accuracy of the resulting solution.
+Cheridito et al., (2025) address a case with 25 assets, although without providing evidence regarding the precision of the resulting solution.
 In contrast, our method substantially improves scalability, handling much larger values of $n$ and comparably high values of $k$, while preserving accuracy even in high-dimensional settings.
-The main contributions of this paper can be summarized as follows:
-  • We introduce Pontryagin-Guided Direct Policy Optimization (PG-DPO), and its highly efficient “Projected PG-DPO” (P-PGDPO) variant that uses PMP structure (specifically FOCs) and BPTT-derived costates to solve complex continuous-time control problems.
-  • We demonstrate through numerical experiments that P-PGDPO not only breaks the dimensional barrier in portfolio optimization (up to n=50 assets, k=10 factors), but critically, accurately captures complex intertemporal hedging demands in these high-dimensional settings, a capability not typically demonstrated by alternative deep learning approaches.
-In the remainder of this paper, we first present the multi-asset Merton problem with exogenous states in Section 2.
+Our companion paper demonstrates that our method is applicable to constrained portfolio optimization problems involving up to 1000 assets (Huh et al., , 2025).
+In the remainder of this paper, we begin with an introduction to the machine learning methods used for the PG-DPO framework in Section 2.
+Then we present the multi-asset Merton problem with exogenous states in Section 3.
 Section 4 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm, its Projected PG-DPO (P-PGDPO) extension, and provides a theoretical justification for the effectiveness of the P-PGDPO approach under specified regularity conditions.
 Numerical experiments presented in Section 5 confirm the framework’s scalability and near-optimal performance within this testbed, illustrating its potential for high-dimensional applications.
-Finally, Section 5 concludes and discusses potential future directions.
-2 Multi-Asset Continuous-Time Portfolio Choice Problem with State Variables
-We consider the continuous-time portfolio choice problem introduced by Merton, (1973), in which a representative investor allocates wealth across $n$ risky assets and a risk-free asset, while consuming continuously over time.
-The investor selects an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}=(\pi_{1,t},\ldots,\pi_{n,t})^{\top}$, where $\pi_{i,t}$ denotes the proportion of current wealth $X_{t}$ invested in the $i$-th risky asset at time $t$ and notation denotes the transpose of a vector or a matrix.
-The investment opportunities are characterized by an exogenous $k$-dimensional state process $\mathbf{Y}_{t}=(Y_{1,t},\ldots,Y_{k,t})^{\top}\in\mathbb{R}^{k}$, which governs the drift and volatility of asset returns.
+Finally, Section 6 concludes and discusses potential future directions of research.
+2 Machine Learning Basics
+Our proposed Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework is built upon fundamental concepts from machine learning.
+Before presenting our approach, this section introduces the computational methods and reinforcement learning principles that underpin our methodology.
+We begin with computational graphs and the algorithm of backpropagation through time (BPTT), which form the basis for efficient gradient calculation in sequential settings.
+We then discuss reinforcement learning and policy-based methods, highlighting their connection to dynamic portfolio choice and clarifying the distinctions between our framework and conventional RL approaches.
+Since our method assumes that the dynamics of the financial market are known, it can be classified as a form of model-based reinforcement learning.
+2.1 Computational Graphs and Backpropagation Through Time
+A computational graph provides a structured way to represent complex mathematical problems as a network of simple operations.
+In the portfolio optimization problem, the graph encodes the entire process, from the initial wealth and market states to the final objective function $J$, through sequential transformations of wealth and state variables.
+This representation is not merely a visualization but a computational tool: it enables us to systematically apply the chain rule of calculus across the entire problem.
+Once we have this graph, we need an efficient way to understand how small changes in our decisions (the portfolio weights, parameterized by $\theta$) affect the final outcome.
+This is achieved through backpropagation, which works by starting from the final outcome and moving backward through the graph, systematically applying the chain rule at each step to calculate the contribution of each parameter to the final result.
+It is an algorithm that efficiently computes the gradient, or the sensitivity of the final output with respect to every parameter in the graph, denoted as $\nabla_{\theta}J$.
+When applied to decision problems evolving over time, such as dynamic portfolio choice, the method is known as Backpropagation Through Time (BPTT).
+A key advantage of our model-based approach is that the dynamics of the environment, such as the stochastic differential equations (SDEs) for asset prices, are explicitly known.
+This allows us to compute exact analytical gradients for a given sample path through the simulation, rather than relying on estimation.
+BPTT is the core mechanism that enables our model to learn from simulated market trajectories by calculating explicit gradients.
+For a given trajectory, it computes the gradient of the final utility $J$ with respect to the policy parameters $\theta$ by propagating sensitivities backward through time.
+The influence of wealth in any intermediate time step, $X_{t+1}$, on the final result is found by applying the chain rule sequentially from the terminal state $X_{T}$:
+$\frac{\partial J}{\partial X_{t+1}}=\frac{\partial J}{\partial X_{T}}\frac{\partial X_{T}}{\partial X_{T-1}}\frac{\partial X_{T-1}}{\partial X_{T-2}}\cdots\frac{\partial X_{t+2}}{\partial X_{t+1}}=\frac{\partial J}{\partial X_{T}}\prod_{k=t+1}^{T-1}\frac{\partial X_{k+1}}{\partial X_{k}}.$
+This term represents the gradient ”flowing back” from the end of the simulation.
+The total gradient with respect to the policy parameters $\theta$ is then calculated by summing up the contributions from the policy $\pi_{t}$ at every time step, weighted by this backward propagated sensitivity:
+$\frac{\partial J}{\partial\theta}=\sum_{t=0}^{T-1}\frac{\partial J}{\partial X_{t+1}}\frac{\partial X_{t+1}}{\partial\pi_{t}}\frac{\partial\pi_{t}}{\partial\theta}$
+Since the dynamics of the financial market are explicitly modeled, our framework can compute analytical gradients along each sample path.
+This contrasts with model-free methods, which must rely on repeated trial-and-error estimates.
+In our method, BPTT thus plays a dual role: it provides the gradients necessary for optimizing the policy parameters, and at the same time, it generates pathwise costate estimates that are central to Pontryagin’s Maximum Principle, which will be introduced in Section 3.2.
+2.2 Reinforcement Learning and Policy-Based Methods
+Reinforcement Learning (RL) provides a general framework in which an agent learns to make decisions by interacting with an environment.
+At each point in time, the agent observes the current state, chooses an action, and receives a reward, with the goal of maximizing cumulative rewards over time.
+The mapping of this paradigm to dynamic portfolio choice is direct.
+The agent corresponds to an investor, the environment to a stochastic financial market, the state to the current wealth and economic factors, the action to portfolio allocation and consumption, and the reward to the utility derived from consumption and terminal wealth.
+A central distinction in RL lies between value-based and policy-based methods.
+Value-based approaches estimate value functions and derive actions indirectly, whereas policy-based methods directly optimize a parameterized policy $\pi_{\theta}$.
+In our setting, the policy is represented by a neural network with parameters $\theta$, which maps the current state $s_{t}$ (e.g., market conditions, current wealth) to an action $a_{t}=\pi_{\theta}(s_{t})$ (e.g., a specific portfolio allocation).
+The optimization problem reduces to finding the parameter set $\theta^{*}$ that maximize the expected total reward (our objective function $J$):
+$\theta^{*}=\arg\max_{\theta}\mathbb{E}[J(\theta)].$
+Parameter updates follow the stochastic gradient ascent rule,
+$\theta_{k+1}=\theta_{k}+\alpha\mathbb{E}\left[\frac{\partial J(\theta_{k})}{\partial\theta}\right]$
+where $\alpha$ is a learning rate and gradients are computed through BPTT.
+While our PG-DPO framework is conceptually rooted in this class of algorithms, it departs in important ways.
+First, it is a model-based RL method.
+Unlike model-free approaches such as Proximal Policy Optimization or actor-critic schemes, which require repeated trial-and-error exploration to estimate gradients, as seen in Dai et al., (2025) that tackle the Merton problem, our framework exploits the explicitly known dynamics of financial markets.
+This allows us to compute exact analytical gradients through BPTT, leading to both greater efficiency and higher numerical accuracy.
+Second, our approach is explicitly guided by Pontryagin’s Maximum Principle (PMP).
+Standard RL methods often converge to policies that deliver nearly optimal utility values but may diverge substantially from the structural form of the true solution.
+In other domains, such as Atari, Go, or autonomous driving, such structural fidelity is less critical, since achieving a near-optimal utility value alone is often sufficient for strong performance.
+In dynamic portfolio choice, however, the decomposition of the optimal strategy into myopic demand and intertemporal hedging demand is both theoretically central and practically important.
+By leveraging PMP, our framework ensures that both components are faithfully captured, yielding policies that are structurally consistent with financial theory.
+A further distinction arises in comparison with Duarte et al., (2024), whose method employs an actor–critic scheme in which the critic is trained by minimizing the HJB residual in a PINN-like fashion.
+[^4] The HJB residual is commonly defined to be the cumulative squared deviation of the left-hand side of the HJB equation from 0..
+However, this approach is designed exclusively for infinite-horizon problems, where the optimal solution is stationary and time-independent.
+Although appropriate for certain theoretical settings, this restriction makes it less suitable for most practical applications, such as retirement planning, endowment management, or finite-maturity products.
+By contrast, our PG-DPO framework is explicitly tailored to the finite-horizon setting, incorporating time-to-maturity as a state variable.
+This enables the algorithm to capture horizon effects, including the time-varying structure of hedging demand and the transition of optimal strategies as maturity approaches.
+Moreover, since the long-horizon case can be treated as a special instance of finite-horizon formulation, our framework naturally extends to settings such as those analyzed in Section 5.3, where the horizon is sufficiently long to yield stationary policies.
+This flexibility is in sharp contrast to the approach of Duarte et al., (2024), which is limited to the infinite horizon case and cannot easily accommodate horizon-dependent dynamics that are crucial in practice.
+In sum, our method combines the principles of reinforcement learning with the structural discipline of PMP, while extending applicability to finite-horizon continuous-time finance.
+This integration ensures that learned policies are not only utility-maximizing but also theoretically well-structured, marking a significant departure from existing RL approaches.
+3 The Model and Pontryagin’s Maximum Principle
+3.1 General Multi-Asset Model
+We consider the continuous-time portfolio choice problem introduced by Merton, (1973), in which a representative investor allocates wealth between risky assets and a risk-free asset, while continuously consuming over time.
+[^5] While we focus on Brownian motion as the source of return risk, the methods can be extended to accommodate jumps.
+Furthermore, if a discrete-time model is viewed as an approximation of a continuous-time model, our approach can be adapted accordingly.
+The stochastic investment opportunities are characterized by an exogenous $k$-dimensional state process $\mathbf{Y}_{t}=(Y_{1,t},\ldots,Y_{k,t})^{\top}\in\mathbb{R}^{k}$, where notation denotes the transpose of a vector (or a matrix), and each component $Y_{i,t}$ evolves according to an Itô diffusion:
+$dY_{i,t}=\mu_{Y,i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{Y,i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dW^{Y}_{i,t},\quad i=1,\dots,k,$
+where the Brownian motions $W^{Y}_{i,t}$ are potentially correlated, satisfying $\operatorname{Cov}\bigl{(}dW^{Y}_{i,t},\,dW^{Y}_{j,t}\bigr{)}=\phi_{ij}\,dt,$ with constant correlations $\phi_{ij}$.
 For instance, $\mathbf{Y}_{t}$ could include macroeconomic indices, volatility factors, exchange rates, etc.
-The investor’s optimization problem involves dynamically choosing both the portfolio weight vector $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and the consumption rate $C_{t}$, conditional on the current state $\mathbf{Y}_{t}$ and wealth $X_{t}$.
-To analyze this problem, we apply Pontryagin’s maximum principle (Pontryagin, , 2014; Yong and Zhou, , 2012), introducing costate variables corresponding to both the wealth process $X_{t}$ and the state process $\mathbf{Y}_{t}$.
-From a computational perspective, the control variables $\{\bm{\pi}_{t},C_{t}\}$ can be parameterized using neural networks, facilitating the implementation of policy gradient methods or DP algorithms in high-dimensional, state-dependent settings.
-2.1 Problem Formulation with State Variables and Neural Network Controls
-We assume that each component $Y_{i,t}$ of the state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ evolves according to an Itô diffusion:
-$dY_{i,t}=\mu_{Y,i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{Y,i}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dW^{Y}_{i,t},\quad i=1,\dots,k,$
-where the Brownian motions $W^{Y}_{i,t}$ are potentially correlated, satisfying
-$\operatorname{Cov}\bigl{(}dW^{Y}_{i,t},\,dW^{Y}_{j,t}\bigr{)}=\phi_{ij}\,dt,$
-with constant correlations $\phi_{ij}$. In vector form, the state process satisfies
-$d\mathbf{Y}_{t}=\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,d\mathbf{W}^{Y}_{t},$
-where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})\in\mathbb{R}^{k}$ is the drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\in\mathbb{R}^{k\times k}$ is a diagonal matrix of diffusion coefficients—i.e.,
-$\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})=\operatorname{diag}(\sigma_{Y,1}(t,\mathbf{Y}_{t}),\dots,\sigma_{Y,k}(t,\mathbf{Y}_{t})),$
-and $\mathbf{W}^{Y}_{t}\in\mathbb{R}^{k}$ is a vector of correlated Brownian motions.
 We consider a financial market consisting of $n$ risky assets and one risk-free asset.
-The risk-free asset earns an instantaneous rate $r(t,\mathbf{Y}_{t})$, which may depend on time and the state variable.
-The price $S_{i,t}$ of the $i$-th risky asset evolves according to
+The risk-free asset earns an instantaneous rate $r(t,\mathbf{Y}_{t})$, which depends on time and the state variable.
+The price of the $i$-th risky asset, $S_{i,t}$, evolves according to
 $dS_{i,t}=S_{i,t}\left(\mu_{i}(t,\mathbf{Y}_{t})\,dt+\sigma_{i}(t,\mathbf{Y}_{t})\,dW^{X}_{i,t}\right),\quad i=1,\dots,n,$
-where $\mu_{i}(t,\mathbf{Y}_{t})$ is the expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ the volatility of asset $i$. The Brownian motions $W^{X}_{i,t}$ driving the risky asset dynamics satisfy
-$\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt,\quad\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt,$
-with constant correlations $\psi_{ij}$ and $\rho_{ij}$.
-In vector notation, let $\bm{\mu}(t,\mathbf{Y}_{t})\in\mathbb{R}^{n}$ denote the vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})\in\mathbb{R}^{n\times n}$ the diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}\in\mathbb{R}^{n}$ the Brownian motion vector for the risky assets.
-Then the dynamics of the vector of risky asset prices $\mathbf{S}_{t}\in\mathbb{R}^{n}$ can be compactly written as
-$d\mathbf{S}_{t}=\operatorname{diag}(\mathbf{S}_{t})\,\bm{\mu}(t,\mathbf{Y}_{t})\,dt+\operatorname{diag}(\mathbf{S}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$
-To clarify the structure of risk in the economy, we define the combined $(n+k)$-dimensional Brownian motion vector
-$\mathbf{W}_{t}=\begin{pmatrix}\mathbf{W}^{X}_{t}\\ \mathbf{W}^{Y}_{t}\end{pmatrix}.$
-The instantaneous covariance matrix of its increments $d\mathbf{W}_{t}$ is denoted by $\bm{\Omega}\,dt$, with
-$\operatorname{Cov}(d\mathbf{W}_{t})=\bm{\Omega}\,dt=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
-where $\Psi=(\psi_{ij})\in\mathbb{R}^{n\times n}$ is the covariance matrix of the asset-specific shocks $d\mathbf{W}^{X}_{t}$, $\Phi=(\phi_{ij})\in\mathbb{R}^{k\times k}$ is the covariance matrix of the state shocks $d\mathbf{W}^{Y}_{t}$, and $\rho=(\rho_{ij})\in\mathbb{R}^{n\times k}$ captures the cross-covariation between the two sources of risk.
-Throughout, we work on a complete probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ that supports a $(n+k)$-dimensional Brownian motion $\mathbf{W}_{t}$ and its natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$, augmented to satisfy the usual conditions (right-continuity and completeness).
-The investor allocates wealth $X_{t}$ between the risk-free asset and the $n$ risky assets. The evolution of the investor’s wealth is described by the following stochastic differential equation (SDE):
+where $\mu_{i}(t,\mathbf{Y}_{t})$ is the expected return and $\sigma_{i}(t,\mathbf{Y}_{t})$ the volatility.
+Brownian motions $W^{X}_{i,t}$ have correlations that satisfy $\operatorname{Cov}(dW^{X}_{i,t},\,dW^{X}_{j,t})=\psi_{ij}\,dt$, and $\operatorname{Cov}(dW^{X}_{i,t},\,dW^{Y}_{j,t})=\rho_{ij}\,dt,$ with constant correlations $\psi_{ij}$ and $\rho_{ij}$.
+The investor selects an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}=(\pi_{1,t},\ldots,\pi_{n,t})^{\top}$, where $\pi_{i,t}$ denotes the proportion of current wealth $X_{t}$ invested in the $i$-th risky asset at time $t$.
+To clarify the structure of risk in the economy, we define the combined $(n+k)$-dimensional Brownian motion vector $\mathbf{W}_{t}=\begin{pmatrix}\mathbf{W}^{X}_{t}\\ \mathbf{W}^{Y}_{t}\end{pmatrix}$.
+where $\mathbf{W}_{t}^{X}$ and $\mathbf{W}_{t}^{Y}$ are Brownian vectors for the risky assets and state variables.
+Then, the instantaneous covariance matrix of its increments $d\mathbf{W}_{t}$ is denoted by $\bm{\Omega}\,dt$, with $\operatorname{Cov}(d\mathbf{W}_{t})=\bm{\Omega}\,dt=\begin{pmatrix}\Psi&\bm{\rho}\\ \bm{\rho}^{\top}&\Phi\end{pmatrix}\,dt,$ where $\Psi=(\psi_{ij})\in\mathbb{R}^{n\times n}$ is the correlation matrix of the asset-specific shocks $d\mathbf{W}^{X}_{t}$, $\Phi=(\phi_{ij})\in\mathbb{R}^{k\times k}$ is the correlation matrix of the state shocks $d\mathbf{W}^{Y}_{t}$, and $\bm{\rho}=(\rho_{ij})\in\mathbb{R}^{n\times k}$ captures the cross-correlation between the two sources of risk.
+Throughout the paper, we consider a complete probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ that supports a standard $(n+k)$-dimensional Brownian motion $\mathbf{W}_{t}$ and its natural filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$, augmented to satisfy the usual conditions such as right-continuity and completeness.
+Then the investor’s wealth dynamics is given by
 $dX_{t}=\left[\,X_{t}\left(r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\right)-C_{t}\,\right]dt+X_{t}\,\bm{\pi}_{t}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t},$ (1)
-where $\mathbf{1}\in\mathbb{R}^{n}$ is a vector of ones.
-The term $\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}$ represents the vector of instantaneous excess returns relative to the risk-free rate.
-We impose regularity conditions on the market parameters.
+where $\mathbf{\mu}(t,\mathbf{Y}_{t})$ is the vector of expected returns, and $\mathbf{1}\in\mathbb{R}^{n}$ is a vector of ones.
+[^6] For the existence and uniqueness of the wealth process, we impose some regularity conditions for the market parameter in Online Appendix A.
+The weight on the risk-free asset is $1-\mathbf{1}^{\top}\bm{\pi}_{t}$, and we define $\bm{\sigma}(t,\mathbf{y}):=\mathrm{diag}\!\big{(}\sigma_{1}(t,\mathbf{y}),\ldots,\sigma_{n}(t,\mathbf{y})\big{)}\in\mathbb{R}^{n\times n}$.
+In order to preclude bankruptcy and ensure solvency, we impose the non-negative wealth constraint
+$X_{t}\geq 0\quad\mbox{for every}\ t\in[0,T].$ (2)
+The investor’s preferences are represented by the following expected utility function:
+${\cal U}=\mathbb{E}_{t,x,\mathbf{y}}\!\left[\int_{t}^{T}e^{-\delta(s-t)}\,U\bigl{(}C_{s}\bigr{)}\,ds\;+\;\kappa\,e^{-\delta(T-t)}\,U\bigl{(}X_{T}\bigr{)}\right],$
+where $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ governs the strength of the bequest motive, $\mathbb{E}_{t,x,\mathbf{y}}[\cdot]:=\mathbb{E}[\cdot\mid X_{t}=x,\,\mathbf{Y}_{t}=\mathbf{y}]$ denotes the conditional expectation, and $U$ is a continuously differentiable, strictly increasing and strictly concave utility function, which satisfies the following Inada conditions [^7] The method in this paper can be easily extended to the case where $\lim_{x\to 0}U^{\prime}(x)<\infty$ by considering the possibility of the corner solution $x=0$ in the first-order condition.
+Thus, the method covers the large class of utility functions such as the Hyperbolic Absolute Risk Aversion (HARA) family: $U(x)=\begin{cases}\dfrac{(a+bx)^{\,1-\gamma}}{1-\gamma},&\gamma\neq 1,\\[6.0pt] \log(a+bx),&\gamma=1,\end{cases}\qquad b>0,\;\gamma>0,\;a\in\mathbb{R},$ defined in the half-line $\mathcal{D}_{U}:=\{x>x_{\min}\}$, with $x_{\min}:=-a/b$.
+:
+$\lim_{x\to\infty}U^{\prime}(x)=0,\quad\lim_{x\to 0}U^{\prime}(x)=\infty.$
+We consider control pairs $(\bm{\pi},C)$ that are both economically sensible and mathematically tractable and denote the admissible set by $\mathcal{A}(x,\mathbf{y})$.
+[^8] Specifically, we define the admissible set $\mathcal{A}(x,\mathbf{y})$ in Online Appendix A.
+Then the investor’s optimization problem involves dynamically choosing both the portfolio weight vector $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and the consumption rate $C_{t}$, conditional on the current state $\mathbf{Y}_{t}$ and wealth $X_{t}$.
+In particular, the value function $V(t,x,\mathbf{y})$ is defined as the maximum expected utility achievable when starting from wealth $x$ and state $\mathbf{y}$ at time $t$:
+$V(t,x,\mathbf{y})=\sup_{(\bm{\pi}_{s},C_{s})\in{\mathcal{A}(x,\mathbf{y})}}\mathbb{E}_{t,x,\mathbf{y}}\!\left[\int_{t}^{T}e^{-\delta(s-t)}\,U\bigl{(}C_{s}\bigr{)}\,ds\;+\;\kappa\,e^{-\delta(T-t)}\,U\bigl{(}X_{T}\bigr{)}\right],$ (3)
+subject to wealth dynamics (1) and constraint (2).
+3.2 Pontryagin’s Maximum Principle
+We introduce Pontryagin’s Maximum Principle (PMP) for solving the continuous-time portfolio choice problem formulated in (3).
+In general, PMP provides necessary optimality conditions through a system of forward–backward stochastic differential equations (FBSDEs), coupling the dynamics of the state variables with those of the associated costate (adjoint) processes.
+In our problem, the system state consists of the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$.
+Specifically, for a given optimal control pair $(\bm{\pi}_{t}^{*},C_{t}^{*})$, the wealth dynamics and state processes are governed by the forward SDEs
+$dX_{t}^{*}$ $=\left[X_{t}^{*}\,\bigl{(}r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{*\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\bigr{)}-C_{t}^{*}\right]dt+X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{X},$ (4)
+$d\mathbf{Y}_{t}$ $=\bm{\mu}_{Y}(t,\mathbf{Y}_{t})\,dt+\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{Y},$ (5)
+with initial conditions $X_{0}^{*}=x_{0}>x_{\min}$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$.
+The corresponding costate processes represent the marginal values with respect to the state variables.
+In particular, we have two costate processes which are the scalar process, $\lambda_{t}:=V_{x}(t,X_{t},\mathbf{Y}_{t})$, corresponding to the sensitivity to wealth, and the vector process $\bm{\eta}_{t}:=V_{\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})\in\mathbb{R}^{k}$, capturing the sensitivity to the exogenous state.
+To characterize optimal controls, PMP introduces a Hamiltonian function that incorporates both the instantaneous utility and the effects of the controls on the system’s drift and diffusion.
+Following the standard stochastic control formulation (e.g., Yong and Zhou, , 2012), the Hamiltonian $\mathcal{H}$ is defined by
+$\mathcal{H}=e^{-\delta t}\,U(C)+\lambda\left[x\bigl{(}r+\bm{\pi}^{\top}(\bm{\mu}-r\mathbf{1})\bigr{)}-C\right]+\bm{\eta}^{\top}\bm{\mu}_{Y}+\mathbf{Z}^{\lambda X}\cdot(x\,\bm{\sigma}^{\top}\bm{\pi})+(\mathbf{Z}^{\eta Y})^{\top}\bm{\sigma}_{Y},$
+where $\lambda$ and $\bm{\eta}$ are the costate processes, $\mathbf{Z}^{\lambda X}$, $\mathbf{Z}^{\eta X}$, and $\mathbf{Z}^{\eta Y}$ represent the diffusion coefficients of the BSDEs associated with $\lambda_{t}$ and $\bm{\eta}_{t}$, and all functions are evaluated at $(t,x,\mathbf{y},\bm{\pi},C)$, though we suppress this for simplicity of notation.
+Complementing the forward dynamics, the optimal costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy a system of BSDEs, driven by the partial derivatives of the Hamiltonian $\mathcal{H}$ with respect to the state variables:
+$-d\lambda_{t}^{*}$ $=\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*}\bigr{)}\,dt-\mathbf{Z}_{t}^{\lambda X*}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\lambda Y*}\,d\mathbf{W}_{t}^{Y},$ (6)
+$-d\bm{\eta}_{t}^{*}$ $=\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*}\bigr{)}\,dt-\mathbf{Z}_{t}^{\eta X*}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\eta Y*}\,d\mathbf{W}_{t}^{Y}.$
+Here, $\mathcal{H}_{x}:=\partial\mathcal{H}/\partial x$ and $\mathcal{H}_{\mathbf{y}}:=\nabla_{\mathbf{y}}\mathcal{H}$ denote the partial derivatives of the Hamiltonian with respect to wealth and the state vector, respectively.
+The terms $\mathbf{Z}_{t}^{\lambda X}\in\mathbb{R}^{1\times n}$, $\mathbf{Z}_{t}^{\lambda Y}\in\mathbb{R}^{1\times k}$, $\mathbf{Z}_{t}^{\eta X}\in\mathbb{R}^{k\times n}$, and $\mathbf{Z}_{t}^{\eta Y}\in\mathbb{R}^{k\times k}$ are adapted processes [^9] These terms arise from the martingale representation theorem applied to the backward dynamics of the costate variables.
+See Karatzas and Shreve, (1991) for a detailed treatment of the martingale representation theorem.
+and collectively denoted by $\mathbf{Z}_{t}$.
+Note that $\mathcal{H}_{x}$ and $\mathcal{H}_{\mathbf{y}}$ generally depend not only on the state, costate, and control variables, but also on the $\mathbf{Z}_{t}$ terms.
+This reflects the coupling between the forward and backward systems, which arises from the second-order sensitivity of the value function.
+The terminal conditions for the BSDEs are determined by the structure of the objective function:
+$\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)},\qquad\bm{\eta}_{T}^{*}=\mathbf{0}.$
+The terminal condition $\eta_{T}=0$ arises because the terminal utility depends only on $X_{T}$ and not directly on the state $Y_{T}$.
+The core requirement of PMP is that the optimal pair $(\bm{\pi}_{t}^{*},C_{t}^{*})$ maximizes the Hamiltonian $\mathcal{H}$ (3.2) pointwise in time, almost surely, given the optimal state and costate processes $(X_{t}^{*},\mathbf{Y}_{t},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*})$.
+We derive the following FOC for the consumption rate $C_{t}$ [^10] The FOC is valid because of the Inada condition (3.1).
+If $\lim_{x\to 0}U^{\prime}(x)<\infty$, we need to consider also the possibility of a corner solution, $C_{t}^{*}=0.$
+$e^{-\delta t}\,U^{\prime}(C_{t}^{*})=\lambda_{t}^{*}.$
+This implies that optimal consumption is the inverse function of marginal utility, $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$.
+For the portfolio weights $\bm{\pi}_{t}$, we differentiate $\mathcal{H}$ with respect to $\bm{\pi}$, accounting for both the drift and diffusion terms that depend on the control.
+Setting the gradient $\nabla_{\bm{\pi}}\mathcal{H}=\mathbf{0}$ and assuming $X_{t}^{*}\neq 0$, we obtain
+$\lambda_{t}^{*}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\,\mathbf{1}\bigr{)}+\bm{\sigma}(t,\mathbf{Y}_{t})\,(\mathbf{Z}_{t}^{\lambda X*})^{\top}=\mathbf{0}.$ (7)
+This condition implicitly defines the optimal portfolio $\bm{\pi}_{t}^{*}$ through its dependence on the costate $\lambda_{t}^{*}$ and the martingale component $\mathbf{Z}_{t}^{\lambda X*}$, which captures the sensitivity of the marginal value of wealth to risk sources.
+To obtain an explicit expression for $\bm{\pi}_{t}^{*}$, we need to relate the coefficient $\mathbf{Z}^{\lambda X*}$ of the martingale component in BSDE (6) to the derivatives of the value function.
+This relationship can be derived by applying Itô’s lemma to $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ and matching terms with the dynamics of the costate process in (6).
+[^11] The full derivation is presented in Online Appendix B.
+This procedure yields an expression for $\mathbf{Z}^{\lambda X*}$ in terms of the second-order derivatives $V_{xx}$, $V_{x\mathbf{y}}$, the optimal controls $\bm{\pi}_{t}^{*}$, and model parameters $(\bm{\sigma},\bm{\sigma}_{Y},\Psi,\bm{\rho})$.
+Substituting this into the first-order condition (7) and solving the resulting equation for $\bm{\pi}_{t}^{*}$ leads to the well-known optimal portfolio formula, originally derived by Merton, (1973):
+$\bm{\pi}_{t}^{*}$ $=-\frac{V_{x}}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}$
+$\qquad\qquad-\frac{1}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,\bm{\rho}\,\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\,V_{x\mathbf{Y}}(t,X_{t},\mathbf{Y}_{t}).$
+Here, $\Sigma(t,\mathbf{Y}_{t}):=\bm{\sigma}(t,\mathbf{Y}_{t})\,\Psi\,\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ is the instantaneous covariance matrix of risky asset returns.
+Notice that all the derivatives in the optimal portfolio can be represented by the costate processes and their derivatives.
+Specifically, we have $V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\nabla_{\mathbf{y}}\lambda_{t}^{*}\in\mathbb{R}^{k}$, and they are evaluated along the optimal trajectory $(t,X_{t}^{*},\mathbf{Y}_{t})$.
+This expression decomposes the optimal portfolio into two components: a myopic demand, which optimizes the static trade-off between expected return and risk under fixed investment opportunities, and an intertemporal hedging demand, which adjusts the portfolio to hedge against stochastic shifts in future investment opportunities.
+The hedging component is shaped by the correlation structure $\bm{\rho}$ between asset returns and state innovations, and by the cross-derivative $V_{x\mathbf{y}}$, which captures how the marginal value of wealth responds to changes in the economic state.
+While both PMP and the Hamilton-Jacobi-Bellman (HJB) equation characterize optimality in continuous-time control problems, they differ fundamentally in formulation and computational implications.
+The HJB approach is based on dynamic programming, yielding a single nonlinear PDE for the value function $V(t,x,\mathbf{y})$.
+Solving the HJB equation requires evaluating first- and second-order derivatives of $V$ over the entire time–state domain and enforcing the PDE residual—the deviation of the left-hand side of the equation from zero—to vanish.
+This provides both necessary and sufficient conditions for optimality under regularity but becomes computationally intractable when the state dimension is high due to the curse of dimensionality.
+By contrast, PMP formulates optimality through a two-point boundary value problem in the form of FBSDEs.
+The forward SDEs describe the evolution of state variables, while the backward SDEs define the costate processes and their terminal conditions.
+Optimal controls are obtained by maximizing the Hamiltonian pointwise in time, given the current state and costate values.
+PMP provides first-order necessary conditions without requiring a global solution for the value function, and its formulation naturally supports stochastic simulation and pathwise gradient computation via BPTT.
+For high-dimensional problems, this structural difference is critical.
+The HJB method must approximate a PDE over the full state grid or via global function approximation, whereas PMP can estimate costates along Monte Carlo trajectories, reducing the computational burden.
+In this paper, we exploit this property by embedding PMP’s first-order conditions directly into the learning loop, avoiding explicit HJB residual minimization and enabling scalability.
+Moreover, while neural PDE solvers such as the Deep BSDE method approximate HJB solutions indirectly via simulated trajectories, our PMP-based approach bypasses the global PDE entirely, allowing more accurate recovery of structural components, particularly the intertemporal hedging term, in high-dimensional portfolio problems.
+3.2.1 Application: Multivariate OU Process
+We apply the PMP framework to a multi-factor environment.
+Specifically, we consider $n$ risky assets whose excess returns are driven by a $k$-dimensional Ornstein–Uhlenbeck (OU) factor vector $\mathbf{Y}_{t}\in\mathbb{R}^{k}$.
+Our PMP formulation remains valid even when the state variables are not OU processes, but we use the OU specification here because it admits semi-analytic benchmark solutions and facilitates clear comparison of numerical accuracy.
+The risk premium for asset $i=1,\dots,n$ is given by $\sigma_{i}(\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t})$, where $\bm{\alpha}_{i}\in\mathbb{R}^{k}$ is the vector of factor loadings for asset $i$, and $\sigma_{i}>0$ is the asset’s constant volatility.
+This formulation captures richer interactions between multiple sources of risk and asset returns.
+The factor process $\mathbf{Y}_{t}$ evolves according to
+$d\mathbf{Y}_{t}=\kappa_{Y}(\bm{\theta}_{Y}-\mathbf{Y}_{t})\,dt+\bm{\sigma}_{Y}\,d\mathbf{W}_{t}^{Y},$
+where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is the mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean, and $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ is a diagonal matrix of factor-specific volatilities.
+The Brownian motion $\mathbf{W}_{t}^{Y}\in\mathbb{R}^{k}$ driving the factors has instantaneous covariance $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi\,dt$.
+The price $S_{t}^{i}$ of each asset follows $\frac{dS_{t}^{i}}{S_{t}^{i}}=\bigl{(}r+\sigma_{i}\,\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{i}\,dW_{t}^{X,i},\quad i=1,\dots,n,$ where $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ is a Brownian motion with instantaneous covariance $\mathrm{Cov}(d\mathbf{W}_{t}^{X})=\Psi\,dt$.
+The asset and factor Brownian motions may be correlated, that is, $\mathrm{Cov}(dW_{t}^{X,i},dW_{t}^{Y,j})=\rho_{ij}\,dt$, with cross-correlation matrix $\bm{\rho}\in\mathbb{R}^{n\times k}$.
+PMP’s first-order conditions yield
+$C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*}),\quad\bm{\pi}_{t}^{*}=-\;\frac{1}{X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\,\Sigma^{-1}\left\{\lambda_{t}^{*}\,\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}+\bm{\sigma}\bm{\rho}\bm{\sigma}_{Y}\,\partial_{\mathbf{Y}}\lambda_{t}^{*}\right\},$
+where $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})\in\mathbb{R}^{n\times n}$ is the volatility matrix, $\Sigma=\bm{\sigma}\Psi\bm{\sigma}^{\top}\in\mathbb{R}^{n\times n}$ is the return covariance matrix, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}\in\mathbb{R}^{n\times k}$ is the factor loading matrix, $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ is the factor volatility matrix, and $\partial_{\mathbf{Y}}\lambda_{t}^{*}\in\mathbb{R}^{k}$ is the gradient of the costate with respect to $\mathbf{Y}_{t}$.
+The optimal portfolio can be separated into two components.
+The first component is a myopic demand, driven by the instantaneous excess returns $\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}$ and the risk-return trade-off under current conditions.
+The second one is an intertemporal hedging demand, which partially offsets the effects of changes in investment opportunities, depending on the sensitivity of marginal utility to changes in $\mathbf{Y}_{t}$ ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$), and the correlation structure between asset and factor shocks ($\bm{\rho}$, $\Psi$, and $\Phi$).
+For the special case with no intermediate consumption ($C_{t}\equiv 0$), the model admits analytic solutions via an exponential-quadratic value function ansatz, leading to a system of matrix Riccati ODEs.
+These closed-form (exponential–quadratic) solutions are standard; see Kim and Omberg, (1996) and Liu, (2007).
+We use them as analytical benchmarks.
+3.2.2 Application: Non-Affine Models
+We can extend the model to allow for non-affine asset return dynamics, providing a more flexible representation of the relationship between state variables and risk premia.
+While the OU factor process in Section 3.2.1 offers analytical tractability, many realistic market environments exhibit nonlinear effects that cannot be captured by an affine specification.
+This subsection introduces such nonlinearities in a controlled manner, enabling us to test the robustness of our PMP-based methods in settings where closed-form solutions are not available.
+We retain the same $k$-dimensional OU process for the state factors, $\mathbf{Y}_{t}=[Y_{1,t},\dots,Y_{k,t}]^{T}$, where $\kappa_{Y},\theta_{Y},$ and $\sigma_{Y}$ are as defined in Section 3.2.1.
+The expected returns of risky assets are modified to
+$\bm{\mu}_{S}(t,\mathbf{Y}_{t})=\text{diag}(\bm{\sigma})(\bm{\alpha}\mathbf{Y}_{\beta,t})+r\mathbf{1}=\text{diag}(\bm{\sigma})\left(\bm{\alpha}(\mathbf{Y}_{t}+\bm{\beta}\mathbf{Y}_{t}^{2})\right)+r\mathbf{1},$
+[…]
+From the perspective of PMP, the introduction of non-affine terms changes the functional form of the drift $\mu(t,\mathbf{Y}_{t})$ in the Hamiltonian, but the overall structure of the first-order conditions remains identical.
+The optimal controls are still given by
+$C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*}),\quad\bm{\pi}_{t}^{*}=-\;\frac{1}{X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\left\{\lambda_{t}^{*}\,(\mu(t,\mathbf{Y}_{t})-r\mathbf{1})+\bm{\sigma}(t,\mathbf{Y}_{t})\bm{\rho}\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\,\partial_{\mathbf{Y}}\lambda_{t}^{*}\right\},$
+where $\mu(t,\mathbf{Y}_{t}),\sigma(t,\mathbf{Y}_{t})$, and $\Sigma(t,\mathbf{Y}_{t})$ incorporate the nonlinear effects of the state variables.
+This non-affine formulation plays a dual role in our study.
+When $\bm{\beta}\rightarrow 0$, the PMP-based methods need to reproduce the affine benchmark results from Section 3.2.1, providing a direct accuracy check.
+In addition, for a larger $\bm{\beta}$, the problem departs from the affine setting, allowing us to evaluate the stability and accuracy of the algorithms when no closed-form solution exists.
+In Section 5.4, we examine numerical results of this non-affine model with various values of $\bm{\beta}$, examining how the learned policies deviate from the affine benchmark as the degree of nonlinearity increases, and assessing how effectively the PMP-based approach captures both the myopic and intertemporal hedging demands in this more general environment.
+3.3 PMP and Neural Networks
+While PMP provides a set of first-order optimality conditions through the coupled FBSDE system, solving these equations exactly remains challenging in high-dimensional portfolio choice problems.
+The costate processes $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y_{t}})$ and $\mathbf{\eta}_{t}=V_{\mathbf{y}}(t,X_{t},\mathbf{Y_{t}})$ depend on derivatives of the value function that are typically unavailable in closed form.
+Traditional numerical approaches attempt to discretize the state space and solve the resulting system of equations, but this approach quickly becomes infeasible as the number of state variables or assets increases.
+We address this challenge by developing a novel gradient-based numerical approach using deep learning.
+The main idea is to parameterize the investor’s policy pair $(\bm{\pi}_{t},C_{t})$ with neural networks:
+$\bm{\pi}_{t}=\bm{\pi}_{\theta}(t,x,\mathbf{y}),\quad C_{t}=C_{\phi}(t,x,\mathbf{y}),$
+where $\theta$ and $\phi$ are the trainable parameters of the respective networks.
+The role of these networks is to approximate the functional form of the optimal control decisions implied by the PMP first-order conditions.
+From PMP, the optimal portfolio can be expressed in terms of costate and its partial derivatives as in equations (3.2) and (3.2).
+While the network parameters $\theta$ and $\phi$ are updated using the stochastic gradient method, the costate variables $(\lambda_{t},\mathbf{\eta}_{t})$ and their derivatives can be computed directly from the simulation paths through the automatic differentiation function of deep learning platforms such as Pytorch.
+All sample paths thus provide unbiased estimates of the necessary PMP quantities, allowing us to bypass the need for a global solution of the value function.
+Importantly, this structure allows us to separate the myopic demand and intertemporal hedging demand during training.
+In particular, the myopic term depends on the current risk premia and the local curvature $V_{xx}$, while the hedging term involves the cross-derivative $V_{x\mathbf{y}}$ and the correlation structure between asset returns and state shocks.
+Embedding this decomposition into network training enables explicit enforcement of the PMP conditions for both components, rather than relying solely on implicit utility maximization.
+The objective function for training is the expected utility,
+$J(\theta,\phi):=\mathbb{E}\!\left[\int_{0}^{T}e^{-\delta s}\,U\bigl{(}C_{\phi}(s,X_{s},\mathbf{Y}_{s})\bigr{)}\,ds\;+\;\kappa\,e^{-\delta T}\,U\bigl{(}X_{T}\bigr{)}\right],$
+which is estimated over simulated sample paths of $(X_{t},Y_{t})$ driven by the policy networks.
+Automatic differentiation through simulation via BPTT yields unbiased gradients of $J(\theta,\phi)$ with respect to all network parameters, while simultaneously producing the costate estimates needed for PMP projection steps in our algorithm.
+The integration of PMP with neural network parameterizations combines the theoretical structure of continuous-time optimal control with the function approximation power of deep learning.
+In the next section, we will build on this framework to construct the Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm and its Projected variant (P-PGDPO), which leverage BPTT-derived costates to achieve scalable solutions even in high-dimensional portfolio choice problems.
+4 The Method
+We present a novel machine learning approach to solving the optimal portfolio choice problem with multiple assets and numerous factors.
+Unlike HJB-based methods that rely on solving a global PDE for the value function, PMP provides pointwise first-order conditions that can be estimated pathwise.
+This difference makes PMP particularly compatible with Monte Carlo simulation and BPTT, thereby avoiding the main scalability bottlenecks of HJB approaches.
+Let us define the extended objective function, $\widetilde{J}(\theta,\phi)$, as the expected value of the conditional utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$, evaluated over initial states sampled from a given distribution $\nu$ on a domain $\mathcal{D}$.
+Specifically,
+$\widetilde{J}(\theta,\phi)=\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\nu}\left[J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\right],$ (8)
+where $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ is defined by
+$\begin{split}J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)=\mathbb{E}^{\theta,\phi}\Bigg{[}&\int_{t_{0}}^{T}e^{-\delta(u-t_{0})}\,U\bigl{(}C_{\phi}(u,X_{u},\mathbf{Y}_{u})\bigr{)}\,du\\ &+\kappa\,e^{-\delta(T-t_{0})}\,U\bigl{(}X_{T}\bigr{)}\;\Bigg{|}\;X_{t_{0}}=x_{0},\,\mathbf{Y}_{t_{0}}=\mathbf{y}_{0}\Bigg{]}.\end{split}$
+Here, $\mathbb{E}^{\theta,\phi}[\cdot\,|\,\cdot]$ denotes the expectation under the dynamics induced by policy $(C_{\phi}(t,X_{t},\mathbf{Y}_{t})$ and $\bm{\pi}_{\theta}(t,X_{t},\mathbf{Y}_{t}))$.
+Notice that the extended value function can have an initial value at any time $t<T$.
+By doing this, the learning results of the neural network parameters are robust across a range of initial conditions, which can significantly reduce exploration costs and lead to rapid convergence of the algorithm.
+To update the parameter $(\theta,\phi)$, we apply the BPTT to Monte Carlo simulations of system trajectories initialized from $\nu$.
+A key feature of this approach is that BPTT not only provides unbiased gradient estimates $\nabla\widetilde{J}$ for policy optimization, but also computes pathwise estimates of Pontryagin costate variables $(\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)})$, capturing the marginal value sensitivities as prescribed by PMP.
+Thus, we refer to our approach as Pontryagin-Guided Direct Policy Optimization (PG-DPO).
+In this section, we first present the baseline PG-DPO algorithm, which is guided by the PMP.
+Then, we extend the approach to explicitly exploit costate information and stabilize the costate estimates by Monte Carlo averaging.
+The second approach will be referred to as Projected PG-DPO (P-PGDPO).
+4.1 Baseline PG-DPO
+The baseline PG-DPO algorithm aims to maximize the extended objective function $\widetilde{J}(\theta,\phi)$ defined in Eq.
+(8) using mini-batch stochastic gradient ascent.
+The baseline PG-DPO directly parameterizes the control policies $\bm{\pi}_{t}$ and $C_{t}$ as neural networks and optimizes their parameters to maximize the extended objective function while satisfying the PMP-implied state-costate dynamics.
+The state dynamics follow the forward SDEs for wealth $X_{t}$ and factors $Y_{t}$, while the PMP-implied costates $\lambda_{t}=V_{x}$ and $\bm{\eta}_{t}=V_{\mathbf{y}}$ evolve according to the backward SDEs given by the Hamiltonian’s adjoint equations.
+The policy networks $\bm{\pi}_{\theta}$ and $C_{\phi}$ map the current state $(t,X_{t},Y_{t})$ to the control variables, which are then used to simulate forward trajectories of $(X_{t},Y_{t})$.
+The key idea is to approximate the gradient $\nabla\widetilde{J}$ by averaging gradients computed on individual sample paths, where each path starts from a randomly drawn initial state.
+In each training iteration, Monte Carlo paths for the driving Brownian motions are first generated.
+These are used to simulate forward trajectories of $X_{t}$ and $Y_{t}$ by using the policies generated from the current policy networks.
+Starting from the terminal conditions $\lambda_{T}=\kappa e^{-\delta T}U^{\prime}(X_{T})$ and $\eta_{T}=0$, the costates $\lambda_{t}$ and $\eta_{t}$ are then computed backward along each path using PMP backward SDEs.
+The expected utility $J(\theta,\phi)$ is evaluated from these simulated paths, and gradients with respect to the network parameters are obtained by BPTT through both the forward state evolution and the backward costate computation.
+Network parameters are updated by stochastic gradient ascent using an optimizer such as Adam.
+Algorithm 1 describes our baseline PG-DPO method with a pseudo-code.
+[^13] We provide the detailed algorithm for the baseline PG-DPO in Online Appendix E.2.
+Algorithm 1 Baseline Pontryagin‑Guided Direct Policy Optimization (PG‑DPO)
+1:Inputs: Policy nets $\pi_{\theta},\,C_{\phi}$; learning‑rate schedule $\{\alpha_{j}\}_{j=0}^{J_{\max}-1}$; max iters $J_{\max}$, batch size $M$, time steps NN; initial‑state distribution $\nu$; utility $U$, discount $\delta$, terminal weight $\kappa$; market params $(\bm{\mu},\bm{\sigma},r,\bm{\mu}_{Y},\bm{\sigma}_{Y}\text{ (diagonal diffusion matrix, }k\times k\text{)},\bm{\Omega})$
+2:Output: optimized params $(\theta_{\text{opt}},\phi_{\text{opt}})$
+3:Initialize $\theta,\phi$
+4:for $j=0$ to $J_{\max}-1$ do
+5:  Draw mini‑batch $\{(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)})\}_{i=1}^{M}\sim\nu$
+6:  $\nabla_{\text{batch}}\leftarrow\mathbf{0}$
+7:  for $i=1$ to $M$ do
+8:   $(X_{0},\mathbf{Y}_{0})\leftarrow(x_{0}^{(i)},\mathbf{y}_{0}^{(i)})$, $\Delta t\leftarrow(T-t_{0}^{(i)})/N$
+9:   for $k=0$ to N1N-1 do
+10:      $t_{k}\leftarrow t_{0}^{(i)}+k\Delta t$
+11:      $\bm{\pi}_{k}\leftarrow\bm{\pi}_{\theta}(t_{k},X_{k},\mathbf{Y}_{k})$
+12:      $C_{k}\leftarrow C_{\phi}(t_{k},X_{k},\mathbf{Y}_{k})$
+13:      Sample $\Delta\mathbf{W}_{k}\sim\mathcal{N}(\mathbf{0},\bm{\Omega}\Delta t)$
+14:      $\mathbf{Y}_{k+1}=\mathbf{Y}_{k}+\bm{\mu}_{Y}(t_{k},\mathbf{Y}_{k})\Delta t+\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\Delta\mathbf{W}_{k}^{\text{Y}}$
+15:      $X_{k+1}=X_{k}+\Bigl{[}X_{k}\bigl{(}r(t_{k},\mathbf{Y}_{k})+\bm{\pi}_{k}^{\!\top}(\bm{\mu}(t_{k},\mathbf{Y}_{k})-r(t_{k},\mathbf{Y}_{k})\bm{1})\bigr{)}-C_{k}\Bigr{]}\Delta t+X_{k}\bm{\pi}_{k}^{\!\top}\bm{\sigma}(t_{k},\mathbf{Y}_{k})\,\Delta\mathbf{W}_{k}^{\text{X}}$
+16:   end for
+17:   $J^{(i)}=\sum_{k=0}^{N-1}e^{-\delta(t_{k}-t_{0}^{(i)})}U(C_{k})\Delta t+\kappa e^{-\delta(T-t_{0}^{(i)})}U(X_{N})$
+18:   $g^{(i)}\leftarrow\nabla_{(\theta,\phi)}J^{(i)}$
+19:   $\nabla_{\text{batch}}\leftarrow\nabla_{\text{batch}}+g^{(i)}$
+20:  end for
+21:  $(\theta,\phi)\leftarrow(\theta,\phi)+\dfrac{\alpha_{j}}{M}\nabla_{\text{batch}}$
+22:end for
+23:return $(\theta,\phi)$
+The PG-DPO adheres to the fundamental principle of reinforcement learning: starting from a randomly chosen initial state, it seeks the optimal policy that maximizes the extended value function.
+We adopt the standard updating rule from backpropagation and take an average of the individual path gradients over the mini-batch.
+Even though we consider the finite-horizon continuous-time model with multiple assets and state variables, the randomly chosen initial state and the averaging of individual sample paths enable stable convergence.
+We now turn to the differences between our approach and existing neural network–based methods for solving continuous-time control problems.
+Among the most widely used are the physics-informed neural networks (PINNs) and the Deep BSDE method of Han and E, (2016) and Weinan, (2017).
+PINNs aim to directly solve the HJB equation, whereas the Deep BSDE method reformulates the HJB equation as a forward–backward stochastic differential system and trains two neural networks: one to approximate the value function and another to approximate its gradient.
+Although the Deep BSDE approach is powerful for approximating HJB solutions, it recovers the control policy only indirectly from the estimated gradient $Z_{t}$.
+As a result, the intertemporal hedging term—depending on the mixed derivative $V_{x\mathbf{y}}$—is not explicitly parameterized but instead inferred indirectly through the network approximation of $Z_{t}$.
+In practice, this often yields a weak identification of hedging demand, especially in high-dimensional settings.
+Even minor inaccuracies in $Z_{t}$ along these gradient directions can become magnified when constructing the policy, particularly when the return covariance matrix $\Sigma$ is ill-conditioned.
+Our PG-DPO framework adopts a different strategy.
+Costates $(\lambda_{t},\eta_{t})$ and their derivatives are computed pathwise via BPTT, and the policy is then either constructed directly or projected onto the manifold defined by the PMP first-order conditions.
+This explicit structural enforcement ensures accurate recovery of both the myopic and intertemporal hedging components—even in cases where Deep BSDE captures the myopic demand well but fails to approximate the more complex hedging demand.
+As shown in Section 5, this distinction is critical in high-dimensional portfolio optimization, where our approach consistently yields lower RMSE and more stable convergence.
+Role of intermediate BPTT values and the “Pontryagin-guided” nature.
+A central insight of PG-DPO is that the intermediate quantities produced during BPTT, namely the pathwise estimates of the costates and their derivatives, can themselves be exploited as guiding signals for policy learning.
+This gives the method its ”Pontryagin-guided” nature, that is, rather than allowing the policy networks to search blindly for an optimum, the training is explicitly steered by the PMP first-order conditions.
+By applying the chain rule backward from the final reward $J^{(i)}$, BPTT inherently calculates the sensitivity of the reward of the intermediate state variables.
+These sensitivities are the pathwise unbiased estimates to the Pontryagin costates:
+$\lambda_{k}^{(i)}=\frac{\partial J^{(i)}}{\partial X_{k}^{(i)}},\qquad\bm{\eta}_{k}^{(i)}=\frac{\partial J^{(i)}}{\partial\mathbf{Y}_{k}^{(i)}}.$
+To connect this computational recurrence to the formal PMP framework, we generalize it to an $\mathcal{F}_{t_{k}}$-adapted process $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}\mid\mathcal{F}_{t_{k}}]$.
+This transition from a path-specific value to an adapted process reveals a deep structural correspondence, which we formalize below.
+Lemma 1 (Pathwise BPTT Recurrence).
+Along any simulated trajectory $i$, the BPTT algorithm computes the backward recursion for the pathwise sensitivity $\lambda_{k}^{(i)}$ as:
+$\lambda_{k}^{(i)}$ $=\lambda_{k+1}^{(i)}+\lambda_{k+1}^{(i)}\bigl{[}r_{k}+\bm{\pi}_{k}^{\!\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{]}\Delta t+\bigl{[}e^{-\delta(t_{k}-t_{0})}U^{\prime}(C_{k}^{(i)})-\lambda_{k+1}^{(i)}\bigr{]}\frac{\partial C_{k}^{(i)}}{\partial X_{k}^{(i)}}\,\Delta t$
+$\quad+\lambda_{k+1}^{(i)}\,\bm{\pi}_{k}^{\!\top}\bm{\sigma}_{k}\,\Delta\mathbf{W}_{k}^{X,(i)}.$ (9)
+Lemma 1 shows that BPTT naturally recovers the discrete-time adjoint dynamics of PMP along each simulated path.
+In the limit $\Delta t\to 0$, these recursions converge to the continuous-time PMP adjoint equations, with an $O(\Delta t)$ discretization error.
+This establishes that the backward sensitivities obtained by BPTT align with the PMP costate dynamics.
+Lemma 2 (Quadratic-Covariation Drift).
+Let $\lambda_{k+1}^{(i)}=\lambda_{k+1}+\widetilde{\mathbf{Z}}^{\lambda X}_{k}\Delta\mathbf{W}_{k}^{X,(i)}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\Delta\mathbf{W}_{k}^{Y,(i)}$ be the martingale representation of $\lambda_{k+1}^{(i)}$, where $\widetilde{\mathbf{Z}}^{\lambda X}_{k}$ is a row vector.
+The conditional expectation of the stochastic term in Eq.
+(1) is:
+$\mathbb{E}\!\Bigl{[}\lambda_{k+1}^{(i)}(\bm{\pi}_{k}^{\!\top}\bm{\sigma}_{k}\,\Delta\mathbf{W}_{k}^{X,(i)})\Bigm{|}\mathcal{F}_{t_{k}}\Bigr{]}\;=\;\bigl{(}\widetilde{\mathbf{Z}}^{\lambda X}_{k}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\bm{\rho}_{Y\!X}\bigr{)}(\bm{\sigma}_{k}^{\!\top}\bm{\pi}_{k})\,\Delta t.$
+Lemma 2 establishes the martingale representation, [^14] The Martingale Representation Theorem, cited from Karatzas and Shreve, (1991), is a continuous-time result.
+Its application to our discrete-time grid is justified by interpreting the discrete process as a sampling of an underlying continuous process, for which the representation holds on each interval $t_{k},t_{k+1}]$.
+where $\lambda_{k+1}^{(i)}$ denotes the costate along the $(i)$-th path.
+The recursion decomposes into a deterministic drift term and a stochastic fluctuation term, with the quadratic covariation structure ensuring well-defined conditional expectations.
+Consequently, the randomness in the BPTT-based costate update arises solely from martingale components, which vanish on average, leaving the deterministic drift consistent with PMP.
+This explains why costate estimates converge stably in expectation.
+Theorem 3 (BPTT–PMP Correspondence).
+Define the effective martingale component as $\mathbf{Z}^{\lambda X}_{k}:=\widetilde{\mathbf{Z}}^{\lambda X}_{k}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\bm{\rho}_{Y\!X}$.
+The adapted costate process $(\lambda_{k})_{k=0}^{N}$ satisfies the discrete-time backward stochastic difference equation:
+$\lambda_{k}$ $=\lambda_{k+1}+\Bigl{[}\lambda_{k+1}\bigl{(}r_{k}+\bm{\pi}_{k}^{\!\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{)}+\mathbf{Z}^{\lambda X}_{k}(\bm{\sigma}_{k}^{\!\top}\bm{\pi}_{k})+\bigl{(}e^{-\delta t_{k}}U^{\prime}(C_{k})-\lambda_{k+1}\bigr{)}\frac{\partial C_{k}}{\partial X_{k}}\Bigr{]}\Delta t$
+$\quad-\widetilde{\mathbf{Z}}^{\lambda X}_{k}\,\Delta\mathbf{W}^{X}_{k}-\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\,\Delta\mathbf{W}^{Y}_{k}.$ (10)
+Furthermore, the drift term in the square brackets of Eq.
+(3) is identical to the partial derivative of the Hamiltonian, $\partial_{x}\mathcal{H}$, evaluated at the corresponding state and control variables.
+Theorem 3 shows that the costate process obtained through BPTT coincides with the PMP adjoint equations, thereby establishing the equivalence between BPTT recursions and the structural optimality conditions of PMP.
+Put differently, the recursion underlying BPTT is mathematically identical to the adjoint dynamics in PMP.
+In sum, we establish the theoretical foundation of the baseline PG-DPO framework.
+The costates and their pathwise estimates, obtained via BPTT, serve as reliable guiding signals for learning control policies.
+By exploiting this Pontryagin-guided structure, PG-DPO achieves stable training with provable near-optimality guarantees.
+Thus, the training procedure is not merely heuristic but grounded in the PMP optimality conditions.
+While the baseline PG-DPO leverages this link implicitly to obtain effective policy gradients, our Projected PG-DPO variant, introduced in the next section, exploits it explicitly by using the BPTT-computed costates to construct policies aligned with financial theory.
+4.2 Projected PG-DPO
+[…]
+P-PGDPO leverages these early-stabilizing costate estimates to construct near-optimal controls by projecting them onto the manifold defined by PMP’s first-order conditions, thereby avoiding the need for fully converged networks to deploy effective policies.
+The procedure unfolds in two distinct stages, as described below.
+  1. Stage 1.
+    Warm-Up Phase for Costate Estimation
+    First, we execute the Baseline PG-DPO (see Section 4.1) for a predetermined, relatively small number of iterations $K_{0}$, optimizing $\widetilde{J}$.
+The purpose of this stage is not to obtain fully trained networks $(\bm{\pi}_{\theta},C_{\phi})$, but rather to run the BPTT process long enough for the mechanism generating the costate estimates
+    $\lambda_{k}^{(i)}=\frac{\partial J^{(i)}(\dots)}{\partial X_{k}^{(i)}},\quad\bm{\eta}_{k}^{(i)}=\frac{\partial J^{(i)}(\dots)}{\partial\mathbf{Y}_{k}^{(i)}}$
+    and their relevant state-derivatives (e.g., $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$) to stabilize.
+We denote the policy parameters at the end of this warm-up phase by $(\theta^{*},\phi^{*})$.
+  2. Stage 2.
+    Analytic Control Deployment
+    After completing the $K_{0}$ warm-up iterations, the trained policy networks $(\bm{\pi}_{\theta^{*}},C_{\phi^{*}})$ are typically not used directly to generate controls during deployment.
+Instead, for any given state $(t_{deploy},X_{k},\mathbf{Y}_{k})$ encountered at deployment, we employ a Monte Carlo averaging procedure to obtain reliable costate estimates for use in the PMP formulas.
+Specifically, to compute the stabilized costate estimates at $(t_{deploy},X_{k},\mathbf{Y}_{k})$, we proceed as follows: simulate $M_{\text{MC}}$ forward paths starting from $(t_{deploy},X_{k},\mathbf{Y}_{k})$ (or nearby states) up to the horizon $T$, using the fixed policy parameters $(\theta^{*},\phi^{*})$ obtained in Stage 1.
+For each simulated path $j=1,\dots,M_{\text{MC}}$, calculate its realized reward
+$J^{(j)}(t_{deploy},X_{k},\mathbf{Y}_{k};\theta^{*},\phi^{*})$
+(with discounting relative to $t_{deploy}$, as in Algorithm 2), and apply BPTT to obtain the path-specific costate estimates $\lambda_{deploy}^{(j)},\partial_{x}\lambda_{deploy}^{(j)},\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$ at the initial node $t_{deploy}$.
+The final stabilized estimates are then computed as Monte Carlo averages:
+$\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\lambda_{deploy}^{(j)},\quad\partial_{x}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{x}\lambda_{deploy}^{(j)},\quad\partial_{\mathbf{Y}}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}.$
+This averaging step substantially reduces the variance inherent in single-path BPTT estimates from Stage 1 and provides more robust inputs for the PMP formulas.
+Algorithm 2 Projected PG‑DPO (P-PGDPO): Warm‑up Training & Deployment
+1:Stage 1: Warm‑up Training
+2:Inputs of Algorithm 1; warm‑up iterations $K_{0}$
+3:Stabilized parameters $(\theta^{*},\phi^{*})$
+4:Run Algorithm 1 for $K_{0}$ iterations to obtain $(\theta^{*},\phi^{*})$
+5:
+6:Stage 2: Deployment at state $(t,X,\bm{Y})$
+7:State $(t,X,\bm{Y})$; parameters $(\theta^{*},\phi^{*})$; Monte‑Carlo rollouts $M_{\mathrm{MC}}$, steps $N^{\prime}$
+8:Near‑optimal controls $(\bm{\pi}^{\mathrm{PMP}},C^{\mathrm{PMP}})$
+9:Initialize $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}\leftarrow\varnothing$
+10:for $j=1$ to $M_{\mathrm{MC}}$ do
+11:  Simulate path from $(t,X,\bm{Y})$ with $(\pi_{\theta^{*}},C_{\phi^{*}})$ using $N^{\prime}$ steps. Let $t_{0}^{\prime}=t$.
+12:  Compute realized reward for path $j$:
+13:  $\qquad\displaystyle J^{(j)}\leftarrow\sum_{n=0}^{N^{\prime}-1}e^{-\delta(t_{n}-t_{0}^{\prime})}U\bigl{(}C_{t_{n}}^{(j)}\bigr{)}\,\Delta t^{\prime}+\kappa e^{-\delta(T-t_{0}^{\prime})}U\bigl{(}X_{T}^{(j)}\bigr{)}$
+14:  Evaluate $\lambda_{t}^{(j)}$, $\partial_{x}\lambda_{t}^{(j)}$, $\partial_{\bm{Y}}\lambda_{t}^{(j)}$ (costates at $t_{0}^{\prime}=t$) via BPTT on $J^{(j)}$ and append to lists $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}$.
+15:end for
+16:Compute $\hat{\lambda}_{t},\widehat{\partial_{x}\lambda}_{t},\widehat{\partial_{\bm{Y}}\lambda}_{t}$ $\triangleright$ E.g., by averaging estimates from lists $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}$
+[…]
+18:Compute $\bm{\pi}^{\mathrm{PMP}}$ via PMP FOC Eq.  (3.2) $\triangleright$ Using estimated costates/derivatives
+19:return $(\bm{\pi}^{\mathrm{PMP}},C^{\mathrm{PMP}})$
+Next, we substitute the averaged costate estimates $(\lambda_{deploy},\partial_{x}\lambda_{deploy},\partial_{\mathbf{Y}}\lambda_{deploy})$ directly into the analytical formulas derived from the PMP first-order conditions (Section 3.2).
+The optimal consumption at $t_{deploy}$ is given by
+$C_{deploy}^{\mathrm{PMP}}=\bigl{(}U^{\prime}\bigr{)}^{-1}\!\bigl{(}e^{\delta t_{deploy}}\,\lambda_{deploy}\bigr{)}.$
+The optimal investment portfolio at $t_{deploy}$ is
+$\bm{\pi}_{deploy}^{\mathrm{PMP}}$ $=-\frac{1}{X_{k}\,\partial_{x}\lambda_{deploy}}\,\Sigma^{-1}(t_{deploy},\mathbf{Y}_{k})\Biggl{\{}\lambda_{deploy}\,\bigl{[}\bm{\mu}(t_{deploy},\mathbf{Y}_{k})-r(t_{deploy},\mathbf{Y}_{k})\mathbf{1}\bigr{]}$
+$\qquad\qquad+\bigl{(}\bm{\sigma}(t_{deploy},\mathbf{Y}_{k})\,\bm{\rho}\,\bm{\sigma}_{Y}(t_{deploy},\mathbf{Y}_{k})\bigr{)}\,(\partial_{\mathbf{Y}}\lambda_{deploy})\Biggr{\}}.$
+This expression coincides with Eq. (3.2), where the averaged costate estimates
+$(\lambda_{deploy},\partial_{x}\lambda_{deploy},\partial_{\mathbf{Y}}\lambda_{deploy})$ correspond in expectation to the true derivatives $(V_{x},V_{xx},V_{x\mathbf{Y}})$ at the state $(t_{deploy},X_{k},\mathbf{Y}_{k})$.
+Finally, the analytically computed controls $(C_{deploy}^{\mathrm{PMP}},\bm{\pi}_{deploy}^{\mathrm{PMP}})$, based on these averaged costates, are applied as the control actions at $(t_{deploy},X_{k},\mathbf{Y}_{k})$ during deployment.
+In sum, the Projected PG-DPO (P-PGDPO) framework proceeds in two stages.
+Stage 1 consists of a short warm-up phase using the baseline PG-DPO procedure to obtain a preliminary policy $(\theta^{*},\phi^{*})$, during which BPTT produces reasonably stable costate estimates.
+In Stage 2, these estimates are refined via Monte Carlo averaging at deployment time and substituted into the analytical PMP formulas to compute the optimal controls.
+The overall procedure is summarized in Algorithm 2.
+This two-stage strategy achieves significant computational savings while yielding high-quality, near-optimal controls that explicitly reflect the decomposition of optimal portfolios into myopic and intertemporal hedging components.
+In particular, the Monte Carlo averaging in Stage 2 provides reliable inputs for the PMP formulas by mitigating the noise inherent in single-path estimates.
+Rationale for the Projected PG-DPO (P-PGDPO) Approach.
+Here we provide a theoretical justification for P-PGDPO.
+The importance of this justification stems from a fundamental challenge facing many simulation-based optimization methods.
+In direct policy optimization, including standard reinforcement learning, a learned policy may achieve a near-optimal objective value (a small value gap) while being structurally far from the true optimal policy (a large policy gap).
+This discrepancy arises because the objective function is often locally “flat” near the optimum: even when the Pontryagin first-order conditions exhibit only small residuals and the resulting utility loss appears at the second-order level, the learned policy can still differ substantially from the true optimum.
+Our framework addresses this issue by shifting the focus from the objective value to the satisfaction of Pontryagin’s first-order conditions (FOCs).
+The effectiveness of this approach relies on a key property of the HJB equations common in finance: their parabolic nature.
+Parabolic PDEs, much like the heat equation, exhibit a “smoothing” effect, ensuring that a well-behaved value function also possesses well-behaved derivatives (costates).
+This is not merely an intuition but a rigorous result of regularity theory in Sobolev spaces, which guarantees that convergence in value entails convergence of its derivatives.
+It is precisely this regularity that allows us to trust the BPTT-derived costates and to reconstruct a robust policy via the PMP formulas.
+Building on this expected regularity, we provide a more formal justification for P-PGDPO.
+We aim to bound the policy gap, $\|\widehat{\pi}-\pi^{*}\|$, which is the error between our P-PGDPO policy $\widehat{\pi}$ and the true optimal policy $\pi^{*}$.
+The bound depends on two primary sources of error.
+The first is the FOC residual, denoted by $\varepsilon$, which measures how much the policy from the warm-up stage violates the optimality conditions of PMP.
+The second is the BPTT estimation error, $\delta_{\mathrm{BPTT}}$, in the costate estimate $\widehat{\lambda}$.
+This estimation error itself stems from two numerical sources: the time discretization of the SDEs (with time step $\Delta t$) and the Monte Carlo sampling used for the expectation (with batch size $M$).
+The following theorem encapsulates this relationship, with full technical details provided in Online Appendix C.4.
+{restatable}
+[Policy Gap Bound]theoremthm:policy_gap Let $\pi^{*}$ be the true optimal policy and $\widehat{\pi}$ be the policy generated by the P-PGDPO algorithm using a time step $\Delta t$ and batch size $M$.
+Under the regularity conditions specified in Assumption 5 in Online Appendix C.4.3, the gap between the P-PGDPO policy and the true optimum is bounded as follows:
+$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\text{tot}}\left(\varepsilon+\kappa_{1}\Delta t+\frac{\kappa_{2}}{\sqrt{M}}\right)$
+where:
+  • $\varepsilon$ is the $L^{q,p}$-norm of the Pontryagin FOC residual from the warm-up policy.
+  • $\kappa_{1},\kappa_{2}$ are positive constants governing the BPTT estimation error, which arises from time discretization (error proportional to $\Delta t$) and Monte Carlo sampling (error proportional to $1/\sqrt{M}$).
+  • $C_{\text{tot}}$ is a positive constant that depends on the model parameters but not on $\varepsilon$, $\Delta t$, or $M$.
+The proof of Theorem 4.2 formalizes the intuition outlined above.
+Heuristically, a small FOC-gap ($\varepsilon$) implies that the candidate policy is nearly consistent with the PDE’s structural requirements.
+Owing to the smoothing property of uniformly parabolic PDEs—guaranteed by regularity theory in Sobolev spaces (e.g., Dong and Kim, (2009))—this near-consistency extends beyond the value function itself to its derivatives.
+With the costates thereby shown to be close to their optimal counterparts, projecting them through the PMP map becomes a well-posed and stable operation, yielding a policy that is correspondingly close to the true optimum.
+Our formal analysis shows that, under suitable regularity conditions—particularly uniform parabolicity—a small Pontryagin FOC gap ($\varepsilon$) from the warm-up stage guarantees that the P-PGDPO policy $\widehat{\pi}$ is provably close to the true optimum $\pi^{*}$, up to simulation-induced numerical errors.
+The resulting bound makes clear that achieving a highly accurate policy requires not only an effective warm-up (small $\varepsilon$) but also a sufficiently fine time discretization ($\Delta t$) and large batch size ($M$) to control BPTT estimation error.
+Moreover, when the objective function is relatively flat near the optimum, the FOC residual $\varepsilon$ for the warm-up policy tends to be small, so that numerical errors become the dominant source of the final policy gap.
+This provides a rigorous justification for the core idea of the P-PGDPO method: substituting numerically controlled and rapidly stabilizing costate estimates into the analytical PMP formulas yields a near-optimal policy.
+[^15] Extending this proof framework beyond uniform parabolicity—to hypo-elliptic (H-class, (Hörmander, , 1967)) or boundary-degenerate (D-class, (Oleinik, , 2012)) models, or to general non-affine systems—faces significant analytical challenges, since key assumptions (such as global $L^{p}$ regularity or the costate floor) may fail.
+Despite these hurdles, the fundamental algorithmic rationale of P-PGDPO remains compelling for broader applicability.
+We classify parabolic HJB PDEs relevant to financial markets in Online Appendix D.
+The central innovation is the decoupling of costate estimation from direct policy construction.
+Standard end-to-end methods conflate these two tasks, making it difficult to capture the fine structure of the policy in a “flat” optimization landscape.
+By contrast, P-PGDPO leverages BPTT for what it excels at—efficient gradient (costate) estimation—and then applies the robust analytical structure of PMP to deterministically map these estimates into a high-fidelity policy.
+This principle, reinforced by our numerical results, suggests that P-PGDPO is a practical and powerful tool for a broad class of continuous-time control problems, even when the strict assumptions required for the proof do not hold.
+5 Numerical Results
+In this section, we present numerical experiments to assess both the performance and scalability of the proposed PG-DPO framework, considering its baseline form (PG-DPO) as well as the Projected variant (P-PGDPO).
+As outlined in Section 3.2.1, the experiments consider a multi-asset, multi-factor portfolio optimization problem where the state factors follow Ornstein–Uhlenbeck (OU) processes, together with an extension to a non-affine specification.
+While the methodology readily accommodates more general factor dynamics, we focus on OU processes to enable direct comparison with available analytical benchmarks.
+[^16] Our evaluation differs from the reverse-engineering procedure of Duarte et al., (2024), where the model is designed so that the optimal policy is known by construction, permitting direct comparison in high-dimensional settings even without an analytic benchmark.
+While reverse-engineering is useful when no tractable benchmark exists, its accuracy assessment is conditional on the artificial structure imposed.
+By contrast, our experiments leverage an analytical benchmark from the original affine OU model, allowing RMSEs to be computed against the true optimal policy without altering the economic environment.
+This ensures that the accuracy measure is both exact and economically meaningful, though it requires the existence of such a tractable benchmark.
+We consider the constant relative risk aversion utility function for experiments:
+$U(X)=\begin{cases}\dfrac{X^{1-\gamma}}{1-\gamma}&{\rm if}\ \gamma\neq 1\\ \log X&{\rm if}\ \gamma=1,\end{cases}$
+where $\gamma>0$ is the coefficient of relative risk aversion.
+We first examine the affine factor specification, deferring the non-affine case to Section 5.4.
+To isolate the portfolio allocation problem—specifically the decomposition into myopic and intertemporal hedging demands, abstracting from consumption choice—we impose no intermediate consumption (i.e., $C_{t}=0$), reducing the objective to
+$J=\mathbb{E}[U(X_{T})].$
+This setup admits an analytical benchmark solution derived from the associated HJB equation, [^17] The analytical benchmark follows the standard exponential–quadratic/Riccati formulation; see Kim and Omberg, (1996); Liu, (2007).
+We implement these benchmarks following their notation.
+which serves as the ground truth for quantitative performance evaluation.
+In addition to the analytic benchmark, we compare against a leading neural network–based baseline: the Deep BSDE method of Weinan, (2017) and Han et al., (2018).
+This approach reformulates the stochastic control problem as a system of FBSDEs and trains neural networks to approximate both the value function and its gradients.
+The approximations are then used to construct the control policy, typically by enforcing or approximating the HJB first-order conditions.
+Network parameters are learned by minimizing residuals in the BSDE dynamics and the terminal condition, evaluated under Monte Carlo simulation.
+Throughout this section, we evaluate the portfolio strategies generated by PG-DPO, P-PGDPO, and Deep BSDE against the analytic benchmark.
+Performance is measured by the root mean squared error (RMSE) of portfolio weights on risky assets, computed across a range of problem dimensions.
+[^18] Full implementation details—including model parameters, network architectures, training protocols, and evaluation methodology—are provided in Online Appendix E.
+We begin with a relatively short investment horizon of $T=1.5$ years.
+This choice enables a clean assessment of intrinsic accuracy and convergence properties, minimizing the confounding numerical issues that arise in long-horizon simulations.
+Establishing this baseline allows for a controlled comparison of methods before turning, in Section 5.3, to the long-horizon case with targeted algorithmic enhancements.
+Finally, Section 5.4 presents results for models with non-affine asset return dynamics.
+5.1 Single-Asset Models
+We begin with the simplest setting: a single risky asset whose excess return is driven by a $k$-dimensional OU state process ($n=1$).
+This case provides a clear environment in which to examine the accuracy of each method without the numerical instability associated with the large-dimensional covariance matrix $\Sigma$.
+In this case, the optimal portfolio is given by
+$\pi_{t}^{*}=-\frac{1}{\sigma^{2}X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\left\{\lambda_{t}^{*}\alpha\sigma Y_{t}+\sigma\rho\sigma_{Y}\,\partial_{Y}\lambda_{t}^{*}\right\},$
+where $Y_{t}$ is the dynamic risk-premium process
+$dY_{t}=k_{Y}(\theta_{Y}-Y_{t})\,dt+\sigma_{Y}\,dW_{t}^{Y}.$
+Because the multivariate OU factor process yields an analytical benchmark, we can compute RMSE separately for the myopic and intertemporal hedging components.
+This decomposition is a key advantage of our PMP-based approach.
+Table 1 reports the root mean squared errors (RMSEs) for $k\in\{1,5,10\}$ state factors, along with the iteration counts at which the minimum errors are achieved.
+Across all configurations, P-PGDPO achieves the lowest RMSE.
+For example, with $k=5$, P-PGDPO attains an RMSE of $2.50\times 10^{-4}$, consistently outperforming both the baseline PG-DPO $(3.71\times 10^{-2})$ and Deep BSDE $(3.09\times 10^{-2})$.
+The relative performance of baseline PG-DPO and Deep BSDE varies with $k$: for small dimensions ($k=1,5$), Deep BSDE converges faster and achieves lower RMSE than baseline PG-DPO, whereas at $k=10$, PG-DPO $(1.75\times 10^{-2})$ surpasses Deep BSDE $(8.15\times 10^{-2})$ in accuracy.
+[^19] For $k=10$, Deep BSDE failed to train beyond initialization, so its lowest RMSE was attained at the first iteration.
+This reflects optimization difficulties in the high-dimensional BSDE loss.
+Figure 5(a) in Online Appendix F illustrates the iteration-by-iteration policy RMSE comparison between methods.
+These results suggest that both PMP-based and Deep BSDE approaches can mitigate the curse of dimensionality in $k$, but convergence speed and final precision depend critically on the interaction between factor dimension and algorithmic structure.
+[^20] In this regime, Monte Carlo sampling error and function approximation quality become the primary computational bottlenecks, rather than an ill-conditioned covariance matrix $\Sigma$.
+Table 1: Summary of Policy Errors and Iteration Counts when $n=1$
+$k$ Method RMSE Iterations at Min.
+1 PG-DPO $3.260\times 10^{-2}$ 1000
+P-PGDPO $1.120\times 10^{-3}$ 800
+Deep BSDE $9.350\times 10^{-3}$ 3600
+5 PG-DPO $3.714\times 10^{-2}$ 1400
+P-PGDPO $2.500\times 10^{-4}$ 800
+Deep BSDE $3.094\times 10^{-2}$ 9400
+10 PG-DPO $1.751\times 10^{-2}$ 2800
+P-PGDPO $4.710\times 10^{-4}$ 1200
+Deep BSDE $8.146\times 10^{-2}$ 1
+Notably, the structural decomposition of the PMP framework allows us to evaluate the myopic and hedging demands separately.
+In the single-asset case, decomposing the P-PGDPO policy shows that nearly all of the residual error originates from the hedging demand.
+This reflects the greater numerical sensitivity of the hedging term, which depends on cross-derivatives of the value function and is amplified by correlations between asset returns and state shocks.
+By contrast, the myopic demand is recovered with near machine precision across all experiments.
+The implications of hedging-demand errors in the multi-asset setting are examined in the next section.
+5.2 Multi-Asset Models
+We extend the analysis to settings with multiple risky assets ($n>1$).
+As $n$ increases, the optimization problem becomes more challenging due to the higher dimensionality of the control vector and the potential ill-conditioning of the covariance matrix $\Sigma$.
+We consider cases with $n\in\{1,10,50\}$ and factor dimensions $k\in\{2,10\}$, focusing on the affine OU factor model in order to retain analytic benchmark solutions.
+[^21] To our knowledge, existing studies in portfolio management typically consider fewer than 30 risky assets when $k\geq 1$.
+For example, Davey and Zheng, (2022) analyze a model with 20 assets but without state variables influencing returns, while Cheridito et al., (2025) study a 25-asset case without reporting the precision of the resulting solution.
+Table 2 summarizes the minimum policy RMSEs and the iteration count at which each minimum is attained.
+All methods converge even with a large number of risky assets and state variables, and the resulting RMSE values remain small.
+Notably, increasing the number of state variables does not materially affect either the RMSEs or the iteration counts required for convergence.
+The numerical results are consistent with those from the single-asset model.
+In all cases, P-PGDPO delivers the lowest RMSE, demonstrating robustness to increased dimensionality.
+For example, with $n=50$ and $k=10$, P-PGDPO achieves an RMSE of $1.52\times 10^{-2}$, substantially outperforming both baseline PG-DPO $(1.07\times 10^{-1})$ and Deep BSDE $(8.20\times 10^{-2})$.
+While the relative performance of PG-DPO and Deep BSDE varies with $n$ and $k$, P-PGDPO remains clearly superior across all settings.
+[^22] Figure 5 in Online Appendix F report RMSE trajectories over iterations for different numbers of risky assets.
+P-PGDPO exhibits steady convergence as the number of iterations increases.
+The pronounced degradation with increasing $n$, even for small $k$, reflects a fundamental difference between scaling in the asset dimension and scaling in the state dimension.
+Increasing $n$ introduces control-space complexity and numerical fragility.
+The first-order condition for the optimal portfolio involves $\Sigma^{-1}$, and when $n$ is large, $\Sigma$ often becomes ill-conditioned.
+In such cases, even small estimation errors in the costates—particularly the cross-derivatives $V_{x\mathbf{y}}$—are amplified by the matrix inversion, significantly degrading accuracy.
+This effect is structural and cannot be fully offset by generic improvements in function approximation.
+By contrast, increasing $k$ primarily enlarges the state space over which the value function must be approximated.
+Although this raises the difficulty of the approximation task, it does not inherently suffer from the same numerical instability as the large-$n$ case, since $\Sigma$ remains fixed in size and conditioning.
+In other words, scaling in $k$ increases approximation complexity, whereas scaling in $n$ directly amplifies estimation errors through $\Sigma^{-1}$.
+As a result, the neural network–based methods studied here scale more gracefully with $k$ than with $n$.
+Table 2: Summary of Policy Errors and Iteration Counts.
+$k=2$ $k=10$
+$n$ Method RMSE Iterations HR(%) RMSE Iterations HR(%)
+1 PG-DPO $4.770\times 10^{-2}$ 1200 0.1 $1.751\times 10^{-2}$ 2800 0.1
+P-PGDPO $1.990\times 10^{-4}$ 800 0.1 $4.710\times 10^{-4}$ 1200 0.1
+Deep BSDE $2.675\times 10^{-3}$ 3600 0.1 $8.146\times 10^{-2}$ 1 0.1
+10 PG-DPO $1.145\times 10^{-1}$ 4000 4.2 $4.996\times 10^{-2}$ 9800 0.9
+P-PGDPO $3.645\times 10^{-3}$ 4600 4.2 $2.799\times 10^{-3}$ 4800 0.9
+Deep BSDE $3.835\times 10^{-2}$ 1 4.2 $7.623\times 10^{-2}$ 200 0.9
+50 PG-DPO $3.875\times 10^{-1}$ 5600 11.8 $1.070\times 10^{-1}$ 8800 5.7
+P-PGDPO $7.330\times 10^{-2}$ 9800 11.8 $1.522\times 10^{-2}$ 10000 5.7
+Deep BSDE $6.340\times 10^{-1}$ 3800 11.8 $8.199\times 10^{-2}$ 1200 5.7
+HR(%) represents the percentage ratio of hedging demand in the total portfolio.
+Our numerical results confirm this distinction.
+For fixed $n$, increasing $k$ from $2$ to $10$ produces only modest changes in RMSE; in some cases, such as Deep BSDE with $n=50$, performance even improves at higher $k$.
+By contrast, for fixed $k$, raising $n$ from $1$ to $50$ causes a much larger degradation in accuracy across all methods, with the steepest decline observed for baseline PG-DPO.
+We focus on the most challenging high-dimensional case with $n=50$ and $k=10$, presenting results for the first asset together with graphical illustrations.
+Figure 1 compares the sum of myopic and intertemporal hedging demands for the first asset across analytic solution, Baseline PG-DPO, P-PGDPO, and Deep BSDE, evaluated at each method’s optimal iteration.
+The plots are drawn against the first state variable $Y_{1}$ at a representative time to maturity TtT-t, with all other factors $Y_{j\neq 1}$ fixed at their long-term means $\theta_{Y,j}$.
+(a) Analytic Policy ($\pi_{1}^{\text{sol}}$)
+(b) Baseline PG-DPO Policy ($\pi_{1,\theta}$)
+(c) P-PGDPO Policy ($\pi_{1}^{\mathrm{PMP}}$)
+(d) Deep BSDE Policy ($\pi_{1}^{\mathrm{BSDE}}$)
+Figure 1: Comparison of total policies for the first asset ($i=1$) across Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and Deep BSDE in the case where $n=50,k=10$.
+Results shown at (near) optimal iterations for each method (e.g., iteration 6400 for P-PGDPO as plotted), plotted against $Y_{1}$ at a representative time-to-maturity TtT-t.
+Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
+Errors use a symmetric logarithmic (symlog) scale.
+Different color bar ranges reflect varying policy/error magnitudes.
+In this high-dimensional setting, the baseline PG-DPO produces policies that deviate substantially from the benchmark, resulting in large errors.
+This reflects the difficulty of directly approximating the optimal policy without structural guidance from PMP.
+By contrast, the P-PGDPO policy in the panel (c) of Figure 1 is visually indistinguishable from the benchmark, with residuals reduced to near zero except in small regions where hedging effects are most sensitive.
+[^23] The visualizations of policy errors for the first asset across models are depicted in Figure 6 in Online Appendix F.
+Its errors are substantially smaller than those of the other methods, consistent with its superior RMSE of $1.522\times 10^{-2}$ for case $n=50,k=10$ as reported in Table 2.
+These results highlight the effectiveness of the P-PGDPO decoupled learning structure: costates are first estimated via BPTT and stabilized through Monte Carlo averaging, and the policy is then analytically constructed using the PMP first-order conditions.
+The Deep BSDE results in Figure 1 show a marked improvement over the baseline PG-DPO.
+Its learned policy qualitatively resembles the myopic component of the benchmark in Figure 2, capturing the dependence on the state variable $Y_{1}$.
+However, it fails to recover the more complex intertemporal hedging demands.
+The corresponding error plot exhibits large discrepancies in regions where the hedging component is critical, indicating that this element is largely absent from the Deep BSDE solution.
+[^24] See Figure 6 in Online Appendix F.
+This shortcoming reflects both architectural and training limitations.
+The Deep BSDE framework is primarily designed to enforce the terminal condition and maintain consistency between the forward process $Y_{t}$ (approximating the value function) and the backward process $Z_{t}$ (approximating its gradient).
+Intertemporal hedging, however, depends crucially on the mixed second-order derivative $V_{x\mathbf{y}}$, which is not explicitly supervised in the standard two-network Deep BSDE setup.
+Although the $Z_{t}$ network can in principle encode this information indirectly, the absence of direct supervision makes $V_{x\mathbf{y}}$ estimation noisy, leading Deep BSDE to underperform on the hedging component despite capturing the myopic part well.
+Moreover, the loss function penalizes deviations in $Z_{t}$ through a quadratic martingale residual, which may place relatively little weight on certain gradient directions, particularly those tied to $V_{x\mathbf{y}}$.
+Small errors along these directions can nonetheless be magnified when constructing the control policy, especially through multiplication by the precision matrix $\Sigma^{-1}$, which inherits the numerical instability of an ill-conditioned covariance matrix.
+As a result, while the Deep BSDE method effectively learns the dominant myopic component of the policy, it consistently underperforms in capturing the subtler hedging demands essential for dynamic portfolio optimization.
+We now examine the myopic and intertemporal hedging demands separately.
+For hedging demand, traditional dynamic programming methods become computationally intractable in high dimensions.
+Although we obtain numerical results for the intertemporal component, they exhibit complex dependencies on both the state variables and the remaining time to maturity.
+The hedging ratio—defined as the Euclidean-norm ratio of the hedging demand to the total demand for risky assets—increases sharply with $n$ (Table 2), making accurate estimation of hedging demand even more critical in high-dimensional problems.
+For instance, this percentage rises from 0.1 when $n=1$ to 12.1 when $n=50$.
+Consistent with this observation, Figure 2 decomposes the portfolio for the first asset into its myopic component and intertemporal hedging component, with the bottom row showing the corresponding P-PGDPO policy.
+As in Figure 1, surfaces are plotted against the first state factor $Y_{1}$ at a representative time to maturity TtT-t, with all the other state factors $Y_{j\neq 1}$ fixed at their long-term means $\theta_{Y,j}$.
+(a) Analytic Myopic ($\pi_{1,0}^{\text{sol}}$)
+(b) Analytic Hedging ($\pi_{1,1}^{\text{sol}}$)
+(c) P-PGDPO Myopic ($\pi_{1,0}^{\mathrm{PMP}}$)
+(d) P-PGDPO Hedging ($\pi_{1,1}^{\mathrm{PMP}}$)
+Figure 2: Decomposition of the Projected PG-DPO (P-PGDPO) policy for the first asset ($i=1$) in the $n=50,k=10$ case (results from iteration 6400 as plotted), plotted against $Y_{1}$ at a representative time-to-maturity TtT-t.
+Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
+Top row: analytic myopic and hedging components.
+Bottom row: myopic and hedging components of P-PGDPO.
+Note the different color bar ranges.
+The myopic demand recovered by P-PGDPO is virtually identical to the analytic solution in Figure 2(a), with residual errors at the level of numerical precision $O(10^{-7})$.
+This near-perfect reconstruction demonstrates that the costate-driven projection step in P-PGDPO reliably captures the static mean–variance structure of the optimal policy, even in high-dimensional settings.
+By contrast, the hedging component exhibits a richer nonlinear structure, reflecting portfolio adjustments needed to hedge against future changes in the investment opportunity set.
+Although P-PGDPO reproduces the overall shape and magnitude of this component, the error shows that most of the residual policy error originates here.
+[^25] For the visualization of decomposition error, we provide the error map in Figure 7 in Online Appendix F.
+This reflects the structural difficulty of estimating the cross-derivatives $V_{x\mathbf{y}}$ that drive the hedging term, together with the amplification of gradient noise when multiplied by the numerically unstable matrix $\Sigma^{-1}(\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y})$ in large-$n$ settings.
+Indeed, the PMP solution for the hedging term involves the inverse covariance matrix:
+$\pi^{\text{hedge}}=-\frac{1}{X_{t}V_{xx}}\,\underbrace{\Sigma^{-1}}_{\text{ill-conditioned for large $n$}}\,(\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y})\,V_{x\mathbf{Y}}.$
+Crucially, unlike Deep BSDE—which shows little evidence of capturing the hedging component at all (Figure 1)—P-PGDPO’s two-stage design enables it to recover both components with high fidelity.
+By first stabilizing costate estimates via BPTT and then projecting them analytically through PMP’s first-order conditions, P-PGDPO enforces the structural form of the hedging term, a key factor behind its superior RMSE performance in high-dimensional portfolio optimization.
+In sum, the $n=50,k=10$ visualizations clearly demonstrate that Projected PG-DPO can “break the dimensional barrier”: where baseline PG-DPO fails and Deep BSDE provides only a partial (myopic-focused) solution, P-PGDPO succeeds.
+The remaining inaccuracies in the hedging component highlight a pressing need to address the ill-conditioning of $\Sigma$ in very high asset dimensions—potentially through factor structures, regularization, shrinkage techniques, or architecture-aware preconditioning.
+5.3 Long-Horizon Model
+We now turn to the long-horizon investment problem.
+While the results in the previous sections demonstrate that the PG-DPO framework performs well for shorter horizons ($T=1.5$ years), extending to long horizons poses significant challenges.
+In particular, the baseline PG-DPO algorithm is susceptible to compounding variance and bias over extended simulations.
+The error sources identified in Theorem 4.2—discretization error ($\propto\Delta t$) and Monte Carlo sampling error ($\propto 1/\sqrt{M}$)—become increasingly critical as the horizon lengthens.
+To develop a robust framework capable of addressing long-horizon challenges, we introduce three key algorithmic enhancements aimed at improving both accuracy and stability.
+First, we adopt a residual learning scheme for hedging demand.
+The policy network is restructured to learn only the complex intertemporal hedging component, while the analytically known myopic demand is computed separately (see Silver et al., (2018); Johannink et al., (2019)).
+This separation leverages domain knowledge from financial economics: the myopic component is well understood, whereas accurate estimation of the hedging demand remains the primary challenge.
+Residual learning thus allows the network to dedicate its full capacity to the most difficult part of the problem.
+Second, we incorporate a control variate to suppress the variance inherent in long Monte Carlo rollouts, the most critical issue in long-term simulations.
+A key advantage of our model-based approach is the ability to seamlessly integrate classical variance-reduction techniques from the financial engineering literature (Glasserman, , 2004).
+By using the analytically tractable myopic policy as a control variate, we can significantly stabilize the learning process.
+Third, we employ Richardson extrapolation to reduce time-discretization error.
+The known structure of the underlying SDEs in our model-based setting permits the straightforward application of this classical numerical method.
+By combining simulations at time steps $\Delta t$ and $0.5\Delta t$, we achieve higher-order accuracy, which is crucial for maintaining fidelity over long horizons (Glasserman, , 2004).
+Before applying these techniques to the long-horizon problem, we first validate their impact on the original experimental setup with $T=1.5$ years.
+This step ensures that the enhancements deliver genuine performance improvements.
+As shown in Table 3, the new techniques yield a dramatic reduction in policy RMSE for both Baseline PG-DPO and P-PGDPO.
+Performance gains of up to $57.5\times$ confirm that these enhancements are not incremental tweaks but constitute a fundamental upgrade to the algorithm.
+This successful validation provides a strong foundation for tackling the more complex long-horizon investment problem.
+Table 3: RMSE Improvement with Advanced Techniques for $T=1.5$ Horizon
+n k New Baseline RMSE Baseline Improvement New P-PGDPO RMSE P-PGDPO Improvement
+1 1 $5.10\times 10^{-3}$ $6.4\times$ $4.80\times 10^{-5}$ $23.3\times$
+1 2 $8.29\times 10^{-4}$ $57.5\times$ $6.30\times 10^{-5}$ $3.16\times$
+1 5 $1.04\times 10^{-3}$ $35.7\times$ $1.00\times 10^{-5}$ $25.0\times$
+1 10 $5.54\times 10^{-4}$ $31.6\times$ $2.23\times 10^{-4}$ $2.11\times$
+10 2 $1.49\times 10^{-2}$ $7.7\times$ $8.33\times 10^{-4}$ $4.38\times$
+10 10 $5.07\times 10^{-3}$ $9.9\times$ $2.21\times 10^{-3}$ $1.27\times$
+50 2 $8.07\times 10^{-2}$ $4.8\times$ $8.40\times 10^{-3}$ $8.73\times$
+50 10 $3.38\times 10^{-2}$ $3.2\times$ $1.07\times 10^{-2}$ $1.42\times$
+Note: This table compares the minimum policy RMSE of the enhanced algorithms against the original results from Tables 1 and 2.
+The ‘Improvement’ column indicates the factor by which the RMSE declines (e.g., Original RMSE / New RMSE).
+With a demonstrably more robust algorithm, we now address the long-horizon investment problem with a maturity of $T=20$ years.
+As the horizon extends toward infinity, the optimal policy is expected to become stationary, i.e., independent of time-to-maturity.
+Although such a stationary policy is structurally simpler for the network to learn, the extended simulations required to reach it amplify the challenges of variance and error control.
+Our enhanced framework is designed precisely for this scenario.
+Table 4 reports the final policy RMSE for the 20-year experiment.
+The results show remarkably low errors even in high dimensions: in the $n=50,k=10$ case, P-PGDPO achieves an error of just $3.85\times 10^{-3}$.
+This error is even lower than in the 1.5-year experiment, a phenomenon attributable to the synergy between a more powerful algorithm and the convergence toward a simpler stationary target policy.
+Table 4: Policy Errors for Long‑Horizon Problem ($T=20$)
+Baseline PG-DPO P-PGDPO
+$n$ $k$ RMSE Iterations at Min. RMSE Iterations at Min.
+1 1 0.00169 400 0.000041 400
+5 3 0.00876 200 0.000352 1
+30 5 0.07220 1000 0.007030 200
+50 10 0.05020 2600 0.003850 200
+(a) Analytic Solution (Asset 1)
+(b) Analytic Solution (Asset 2)
+(c) Baseline PG-DPO (Asset 1)
+(d) Baseline PG-DPO (Asset 2)
+(e) P-PGDPO (Asset 1)
+(f) P-PGDPO (Asset 2)
+Figure 3: Comparison of learned policies for the 20-year horizon ($n=50,k=10$), plotted against the first state factor $Y_{1}$.
+Left column: Asset 1.
+Right column: Asset 2.
+The P-PGDPO results show a near-perfect match with the benchmark solution, whereas the Baseline PG-DPO fails to capture the correct policy structure.
+As shown in Figure 3, the policies generated by P-PGDPO are visually indistinguishable from the benchmark, accurately capturing the intricate nonlinear dependencies on the state variables.
+By contrast, the baseline PG-DPO, even with enhanced techniques, fails to reproduce the correct structural form of the policy.
+This comparison illustrates the qualitative difference between a rough approximation—potentially yielding a low but misleading error metric—and a truly accurate approximation that captures both the value and the structural complexity of the optimal policy.
+These findings underscore the strength of the enhanced P-PGDPO framework in delivering robust, high-fidelity solutions to challenging long-horizon control problems.
+5.4 Non-affine Asset Return Dynamics
+We now assess the flexibility of our framework under non-affine model dynamics, as introduced in Section 3.2.2.
+In this specification, the $k$-dimensional OU state process is retained, but the expected returns of the risky assets include quadratic terms in the state variables:
+$\bm{\mu}_{S}(t,\mathbf{Y}_{t})=\text{diag}(\bm{\sigma})\Big{(}\bm{\alpha}\big{(}\mathbf{Y}_{t}+\bm{\beta}\mathbf{Y}_{t}^{2}\big{)}\Big{)}+r\mathbf{1},$
+where the vector $\bm{\beta}$ governs the strength and direction of the nonlinearity, with overall magnitude measured by $\beta_{\text{norm}}=\|\bm{\beta}\|_{2}$.
+When $\beta_{\text{norm}}=0$, the model collapses exactly to the affine setting of Section 3.2.1, for which analytic solutions are available.
+As $\beta_{\text{norm}}$ increases, progressively stronger nonlinear effects are introduced, providing a controlled way to assess robustness in settings without an analytic benchmark.
+To ensure a clear and direct comparison, we return to the $T=1.5$ year horizon and the foundational P-PGDPO algorithm.
+This choice allows us to attribute any changes in the optimal policy solely to the introduction of nonlinearity, isolating its effects from the advanced techniques introduced in the long-horizon study.
+All experiments in this subsection are conducted in the most challenging setting, with $n=50$ assets and $k=10$ state factors, using the P-PGDPO method.
+In the non-affine setting, the true optimal policy is unknown.
+To measure accuracy, we first identify in the affine case ($\beta_{\text{norm}}=0$) the iteration count that minimizes RMSE against the analytic benchmark solution from Section 5.2; we refer to this as the benchmark iteration.
+For each $\beta_{\text{norm}}>0$, we then train the network for 10,000 iterations under the non-affine dynamics and compute the RMSE between the resulting policy and the affine benchmark policy at the benchmark iteration.
+This procedure is best interpreted as a deviation measure: it captures the policy changes induced by the nonlinear term relative to the affine baseline, rather than serving as an absolute accuracy metric.
+In this way, deviations reflect only the nonlinear structure, not stochastic optimization noise or premature convergence.
+This approach differs from the reverse-engineering procedure of Duarte et al., (2024), where the environment is artificially constructed so that the optimal policies are known by design.
+Such a setup is useful for controlled evaluation of numerical methods, but its performance assessment is relative to the imposed functional form rather than to the optimal policy of an economically grounded model.
+By contrast, the affine-benchmark RMSE comparison employed here draws its reference from the same economic model in its limiting affine case, ensuring that the comparison reflects meaningful structural changes in the original portfolio optimization problem.
+Moreover, it allows us to verify continuity of the solution by confirming that the learned policy converges smoothly to the affine benchmark as $\beta_{\text{norm}}\!\to\!0$, and to attribute deviations at large $\beta_{\text{norm}}$ values directly to the nonlinearities in the model.
+Table 5 reports the RMSE between the P-PGDPO policy learned under non-affine dynamics and the affine benchmark policy ($\beta_{\text{norm}}=0$), across varying degrees of nonlinearity.
+When $\beta_{\text{norm}}=0$, P-PGDPO achieves a low RMSE ($4.70\times 10^{-1}$), consistent with the high precision observed in the affine experiments of Section 5.2.
+As $\beta_{\text{norm}}$ increases, the RMSE grows monotonically, reflecting a systematic and stable adjustment of the learned policy to the nonlinear dynamics.
+Table 5: Policy errors of P-PGDPO for the model with non-affine risk premia vs. affine benchmark ($n=50,k=10$). Results reported after 10,000 training iterations.
+$\beta_{\text{norm}}$ 0.0 0.1 0.5 1.0 2.0 3.0 4.0
+RMSE (P-PGDPO) 0.04701 0.05686 0.06755 0.11469 0.20120 0.27530 0.33570
+To complement the RMSE-based results, Figure 4 provides a visual comparison of the learned policies.
+The figure plots the policy for the first asset $\pi_{1}$ as a function of $Y_{1}$ and time-to-maturity TtT-t for $\beta_{\text{norm}}\in\{0,0.1,1.0,2.0\}$, with all other factors fixed at their long-run means $\theta_{Y,j}$.
+When $\beta_{\text{norm}}=0$, the learned surface matches the affine benchmark almost exactly, confirming that P-PGDPO reproduces the known solution in the absence of nonlinearity.
+As $\beta_{\text{norm}}$ increases, the surfaces deviate smoothly from the affine shape, exhibiting curvature and level shifts consistent with the quadratic state effects in $\mu(t,Y_{t})$.
+Even at $\beta_{\text{norm}}=2.0$, the learned policy remains smooth and free of spurious oscillations, indicating that the algorithm preserves numerical stability under substantial nonlinearity.
+This visual progression corroborates the quantitative results in Table 5, showing that the P-PGDPO framework consistently captures the effect of nonlinear terms on the risk premium.
+(a) Analytical solution ($\beta_{\text{norm}}=0$).
+(b) P-PGDPO policy ($\beta_{\text{norm}}=0.0$).
+(c) P-PGDPO policy ($\beta_{\text{norm}}=1.0$).
+(d) P-PGDPO policy ($\beta_{\text{norm}}=2.0$).
+Figure 4: Visual comparison of optimal policies for the first asset ($\pi_{1}$) against the first state factor ($Y_{1}$) and time-to-maturity (TtT-t) in the $n=50,k=10$ setting.
+(a) shows the analytical solution for the affine model ($\beta_{\text{norm}}=0$).
+(b)-(e) display P-PGDPO learned policies for models with increasing strength of nonlinearity in asset risk premia, corresponding to $\beta_{\text{norm}}$ values of $0.0,0.1,1.0,2.0$, respectively.
+Note the systematic change in policy structure and value ranges (indicated by colorbars) as $\beta_{\text{norm}}$ increases.
+The quadratic terms affect both the myopic and intertemporal hedging components of the optimal policy.
+Myopic demand shifts directly in response to the altered instantaneous risk premia $\sigma A(Y_{t}+\beta Y_{t}^{2})$, while intertemporal hedging adjusts more subtly as costate gradients $\partial_{Y}\lambda_{t}^{*}$ respond to the nonlinear state dynamics.
+For small $\beta_{\text{norm}}$, these adjustments are modest, and the policy remains close to the affine benchmark.
+For larger $\beta_{\text{norm}}$, the deviations become more pronounced, reflecting genuine changes in optimal hedging behavior rather than artifacts of the numerical method.
+Overall, P-PGDPO exhibits high accuracy in the affine limit and adapts stably under increasing nonlinearity.
+Its smooth convergence to the affine benchmark as $\beta_{\text{norm}}\to 0$ validates the implementation, while systematic deviations at larger $\beta_{\text{norm}}$ highlight the robustness and interpretability of the method in settings without analytical solutions.
+6 Conclusion
+We have proposed Pontryagin-Guided Direct Policy Optimization (PG-DPO) for high-dimensional finite-horizon continuous-time consumption and portfolio choice.
+By integrating Pontryagin’s Maximum Principle (PMP) with backpropagation through time (BPTT), PG-DPO circumvents the curse of dimensionality that hampers DP- and PDE-based methods, directly learning policies that recover both myopic and intertemporal hedging components.
+A central innovation is the Projected PG-DPO (P-PGDPO) algorithm.
+By exploiting the rapid stabilization of costate estimates, P-PGDPO projects them onto PMP’s first-order conditions, yielding policies that are both computationally efficient and structurally consistent.
+This design delivers substantial gains in accuracy and scalability.
+Numerical results confirm these strengths.
+Deep BSDE captures myopic demands but fails to recover hedging terms.
+In contrast, P-PGDPO achieves a policy RMSE of $1.52\times 10^{-2}$ in a $n=50,k=10$ setting, with negligible error in the myopic component and accurate reconstruction of complex hedging demands.
+Theoretical guarantees, including our Policy Gap Bound, further establish near-optimality when costate estimates are accurate and FOC residuals are small.
+Finite-horizon problems fit naturally into this framework, with the long-horizon case emerging as a special case.
+The practical implications are significant.
+P-PGDPO’s ability to recover nonlinear hedging strategies in high dimensions makes it well-suited for institutional portfolios with many assets, ETFs, or derivatives.
+The principle—estimate costates first, then derive policy—offers a blueprint for broader applications where PMP provides structure.
+Indeed, companion work applies it successfully to constrained portfolio optimization (Huh et al., , 2025).
+Future research should pursue: (i) regularization techniques for ill-conditioned covariance matrices; (ii) extensions to frictions such as transaction costs and taxes; (iii) richer preference specifications, including behavioral or sustainability criteria; and (iv) empirical validation through historical and simulated backtests.
+In sum, PG-DPO—and especially P-PGDPO—demonstrates how the synergy of deep learning and PMP’s structural guidance can expand the frontier of solvability in high-dimensional continuous-time financial control.
+It bridges optimal control and modern machine learning, providing both theoretical insight and a foundation for realistic, large-scale decision problems.
+Acknowledgments
+This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-00562904).
+Appendix A Regularity Conditions and Assumptions for a General Multi-Asset Model
+We impose regularity conditions on the market parameters of the general multi-asset model in Section 3.1.
 Assumption 1 (Market-parameter regularity).
@@
 Under Assumption 1, the SDE system for $(\mathbf{Y}_{t},X_{t})$ (where $X_{t}$ depends on the control processes $\bm{\pi}_{t}$ and $C_{t}$) admits a unique strong solution for sufficiently well-behaved controls.
-[^5] See Oksendal, (2013, Chs.
-5–6) for the existence and uniqueness of solutions to SDEs.
-The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ that maximizes the expected discounted utility
-[…]
-subject to the wealth dynamics in Eq.
-(1) and the admissibility conditions stated later in Assumption 2.
-Here, $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ governs the strength of the bequest motive by weighting terminal utility.
-Throughout, we assume preferences are represented by a utility function from the Hyperbolic Absolute Risk Aversion (HARA) family:
-[…]
-defined on the half-line $\mathcal{D}_{U}:=\{x>x_{\min}\}$, with $x_{\min}:=-a/b$.
-To ensure that the utility function is well-defined and the wealth process $X_{t}$ remains strictly positive, we assume $x_{\min}\geq 0$ throughout.
-This restriction precludes bankruptcy and reflects standard modeling practice in continuous-time portfolio choice under HARA preferences.
-This class includes the Constant Relative Risk Aversion (CRRA) utility for $a=0,b=1$, and the Constant Absolute Risk Aversion (CARA) utility for $\gamma=1,b=1$.
-On its domain, $U$ is twice continuously differentiable, strictly increasing, and strictly concave.
-It also satisfies the Inada conditions.
-[^6] See Pennacchi, (2008, §3.2) for further discussion of the properties required of utility functions.:
-$U^{\prime}(x)\to\infty\text{ as }x\downarrow x_{\min},\qquad U^{\prime}(x)\to 0\text{ as }x\to\infty.$
-We consider control processes that are both economically sensible and mathematically tractable.
-Assumption 2 (Admissible controls).
-A control pair $(\bm{\pi},C)$ is admissible if the following conditions hold:
+For the existence and uniqueness of the solutions to SDEs, see chapters 5 and 6 in Oksendal, (2013).
+Now we define the admissible set $\mathcal{A}(x,\mathbf{y})$ for our stochastic control problem.
+Definition 1 (Admissible set).
+A pair $(\bm{\pi},C)$ is admissible if the following conditions hold:
   1. (a)
@@
     Integrability.
-[…]
+    $\mathbb{E}\!\int_{0}^{T}\left(\lVert\bm{\pi}_{t}\rVert^{2}+C_{t}\right)\,dt\;<\;\infty.$
   3. (c)
     Interior bounds. There exist constants $\Pi_{\max}>0$ and $\xi\in(0,1)$ such that
-[…]
+    $\lVert\bm{\pi}_{t}\rVert\;<\;\Pi_{\max},\qquad 0\;\leq\;C_{t}\;<\;\xi\,X_{t},\qquad\forall t\in[0,T]\text{ a.s.}$
   4. (d)
-    Domain-admissible wealth. Let $x_{\min}:=-a/b$ (as defined for the HARA utility). Then
-[…]
+    Doadmissible wealth. Let $x_{\min}:=-a/b$ (as defined for the HARA utility). Then
+    $X_{t}\;>\;x_{\min},\qquad\forall t\in[0,T]\text{ a.s.}$
 Assumption 2 (Risk covariance).
 The block covariance matrix
-$\bm{\Omega}=\begin{pmatrix}\Psi&\rho\\[1.79997pt] \rho^{\top}&\Phi\end{pmatrix}$
+$\bm{\Omega}=\begin{pmatrix}\Psi&\rho\\[1.84995pt] \rho^{\top}&\Phi\end{pmatrix}$
 is constant in time and uniformly positive definite: there exists $\lambda_{\Omega}>0$ such that
@@
 As a consequence, both $\Psi$ and $\Phi$ are individually positive definite and thus invertible.
-The assumptions are standard in continuous-time portfolio choice models.
-Specifically, Assumption 2(d) ensures that the wealth process remains within the domain of the utility function, i.e., $X_{t}>x_{\min}=-a/b$ almost surely, and also rules out arbitrage in this continuous-time setting (Dybvig and Huang, , 1988).
+The definition of an admissible set is standard in continuous-time portfolio choice models.
+Specifically, the doadmissible wealth condition in (d) ensures that the wealth process remains within the domain of the utility function, i.e., $X_{t}>x_{\min}=-a/b$ almost surely, and also rules out arbitrage in this continuous-time setting (Dybvig and Huang, , 1988).
 Since $U$ is $C^{2}$, strictly concave, and satisfies the Inada conditions on $(x_{\min},\infty)$, and because $\delta>0$, we obtain the following integrability bounds:
 $\mathbb{E}\left[e^{-\delta t}\,U(C_{t})\right]\;\leq\;e^{-\delta t}\,U\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}\;<\;\infty,\qquad\mathbb{E}\left[U(X_{T})\right]<\infty.$
-Therefore, the objective functional (2) is well defined for every admissible control pair $(\bm{\pi},C)$.
-The value function $V(t,x,\mathbf{y})$ associated with this problem represents the maximum expected utility achievable when starting from wealth $x$ and state $\mathbf{y}$ at time $t$. Formally,
-[…]
-where the supremum is taken over all admissible control processes on $[t,T]$, and $\mathbb{E}_{t,x,\mathbf{y}}[\cdot]:=\mathbb{E}[\cdot\mid X_{t}=x,\,\mathbf{Y}_{t}=\mathbf{y}]$ denotes the conditional expectation.
-The function $V$ is defined on $[0,T]\times(x_{\min},\infty)\times\mathbb{R}^{k}$.
+Therefore, the value function $V(0,x,\mathbf{y})$ is well-defined for every admissible control pair $(\bm{\pi},C)$. We set all admissible pairs by the admissible set $\mathcal{A}(x,y)$.
 Assumption 3 (Uniform non-degeneracy).
@@
 (7), which is $C^{2}$, strictly increasing and concave on its domain $(x_{\min},\infty)$, and satisfies the Inada conditions.
-Together with the coefficient regularity in Assumption 1, these conditions yield the existence, uniqueness, and sufficient smoothness of the value function (e.g., $V\in C^{1,2,2}$); see Oksendal, (2013, Chs.
-6–7).
-In our numerical implementation, we parameterize the control policies using neural networks (NNs).
-Specifically, the portfolio allocation $\bm{\pi}_{t}$ and consumption rate $C_{t}$ are represented as functions of the current time $t$, wealth $X_{t}$, and state $\mathbf{Y}_{t}$, with functional forms determined by neural network parameters $\theta$ and $\phi$:
-$\bm{\pi}_{t}=\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\qquad C_{t}=C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)}.$
-The parameters $\theta$ and $\phi$ are optimized to maximize the objective functional $J$ defined in Eq. (2).
-2.2 Optimality Conditions via Pontryagin’s Maximum Principle
-We now apply Pontryagin’s Maximum Principle (PMP) to the continuous-time portfolio choice problem formulated in Section 2.1, where market parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$.
-PMP provides necessary conditions for optimality via a system of forward-backward stochastic differential equations (FBSDEs), which couple the dynamics of the state variables with those of the associated costate (or adjoint) processes.
-The system state consists of the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$.
-The corresponding costate processes represent the marginal value of state variables with respect to the value function $V$: the scalar process $\lambda_{t}:=V_{x}(t,X_{t},\mathbf{Y}_{t})$ corresponds to the sensitivity to wealth, and the vector process $\bm{\eta}_{t}:=V_{\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})\in\mathbb{R}^{k}$ captures sensitivity to the exogenous state.
-Under an optimal control policy $(\bm{\pi}_{t}^{*},C_{t}^{*})$, the forward state dynamics are governed by:
-$dX_{t}^{*}$ $=\left[X_{t}^{*}\,\bigl{(}r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{*\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\bigr{)}-C_{t}^{*}\right]dt+X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{X},$ (4)
-$d\mathbf{Y}_{t}$ $=\bm{\mu}_{Y}(t,\mathbf{Y}_{t})\,dt+\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{Y},$ (5)
-with initial conditions $X_{0}^{*}=x_{0}>x_{\min}$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$.
-Recall that $\bm{\sigma}(t,\mathbf{Y}_{t})$ is the asset volatility matrix, and $\bm{\Omega}$ is the covariance matrix of the combined Brownian motion $\mathbf{W}_{t}=\bigl{(}\mathbf{W}^{X\top}_{t},\,\mathbf{W}^{Y\top}_{t}\bigr{)}^{\top}$.
-To characterize optimal controls, PMP introduces a Hamiltonian function that incorporates both the instantaneous utility and the effects of the controls on the system’s drift and diffusion.
-In problems with control-dependent diffusion, the Hamiltonian must also account for the interaction with costate diffusion terms (i.e., the $\mathbf{Z}$-components of the adjoint BSDEs).
-Following standard stochastic control formulations (e.g., Yong and Zhou, , 2012), the Hamiltonian $\mathcal{H}$ is defined as:
-$\mathcal{H}=e^{-\delta t}\,U(C)+\lambda\left[x\bigl{(}r+\bm{\pi}^{\top}(\bm{\mu}-r\mathbf{1})\bigr{)}-C\right]+\bm{\eta}^{\top}\bm{\mu}_{Y}+\mathbf{Z}^{\lambda X}\cdot(x\,\bm{\sigma}^{\top}\bm{\pi})+(\mathbf{Z}^{\eta Y})^{\top}\bm{\sigma}_{Y},$ (6)
-where $\lambda$ and $\bm{\eta}$ are the costate processes, $\mathbf{Z}^{\lambda X}$, $\mathbf{Z}^{\eta X}$, and $\mathbf{Z}^{\eta Y}$ represent the diffusion coefficients of the backward SDEs associated with $\lambda_{t}$ and $\bm{\eta}_{t}$, and all functions are evaluated at $(t,x,\mathbf{y},\bm{\pi},C)$, though we suppress this for clarity.
-Complementing the forward dynamics, the optimal costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy a system of backward stochastic differential equations (BSDEs), driven by the partial derivatives of the Hamiltonian $\mathcal{H}$ with respect to the state variables:
-$-d\lambda_{t}^{*}$ $=\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*}\bigr{)}\,dt-\mathbf{Z}_{t}^{\lambda X*}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\lambda Y*}\,d\mathbf{W}_{t}^{Y},$ (7)
-$-d\bm{\eta}_{t}^{*}$ $=\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*}\bigr{)}\,dt-\mathbf{Z}_{t}^{\eta X*}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\eta Y*}\,d\mathbf{W}_{t}^{Y}.$ (8)
-Here, $\mathcal{H}_{x}:=\partial\mathcal{H}/\partial x$ and $\mathcal{H}_{\mathbf{y}}:=\nabla_{\mathbf{y}}\mathcal{H}$ denote the partial derivatives of the Hamiltonian with respect to wealth and the state vector, respectively.
-The terms $\mathbf{Z}_{t}^{\lambda X}\in\mathbb{R}^{1\times n}$, $\mathbf{Z}_{t}^{\lambda Y}\in\mathbb{R}^{1\times k}$, $\mathbf{Z}_{t}^{\eta X}\in\mathbb{R}^{k\times n}$, and $\mathbf{Z}_{t}^{\eta Y}\in\mathbb{R}^{k\times k}$ are adapted processes.
-[^7] These terms arise from the martingale representation theorem applied to the backward dynamics of the costate variables.
-See Karatzas and Shreve, (1991) for a detailed treatment of the martingale representation theorem.
-These processes are collectively denoted by $\mathbf{Z}_{t}$.
-Note that $\mathcal{H}_{x}$ and $\mathcal{H}_{\mathbf{y}}$ generally depend not only on the state, costate, and control variables, but also on the $\mathbf{Z}_{t}$ terms themselves.
-This reflects the coupling between the forward and backward systems, which arises from the second-order sensitivity of the value function.
-The terminal conditions for the BSDEs are determined by the structure of the objective functional:
-$\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)},\qquad\bm{\eta}_{T}^{*}=\mathbf{0}.$
+Together with the coefficient regularity in Assumption 1, these conditions yield the existence, uniqueness, and sufficient smoothness of the value function (e.g., $V\in C^{1,2,2}$); see Oksendal, (2013, Chapter 6 and 7).
 To ensure the well-posedness of the forward–backward stochastic differential equation (FBSDE) system and the applicability of Pontryagin’s Maximum Principle (PMP), additional smoothness conditions on the model coefficients are typically required.
 Assumption 4 (Additional regularity for PMP).
-In addition to Assumptions 1, 2, and 4, we assume:
+In addition to Assumptions 1, 3, and an addmissible set in Definition 1, we assume:
   1. (a)
-    $C^{1}$-regularity with respect to $\mathbf{y}$.
-All coefficient functions—$\bm{\mu}_{Y},\,\bm{\sigma}_{Y},\,\bm{\mu},\,\bm{\sigma},\,r$—are continuously differentiable with respect to the state vector $\mathbf{y}$.
-Moreover, their partial derivatives $\partial_{\mathbf{y}}f$ are globally Lipschitz continuous in $\mathbf{y}$ and satisfy linear growth bounds.
-These conditions ensure that the drift terms in the costate BSDEs, including expressions such as $\mathcal{H}_{\mathbf{y}}$, satisfy the regularity requirements for the existence and uniqueness of solutions; see Yong and Zhou, (2012, Chapter 2 and 6).
-The core requirement of Pontryagin’s Maximum Principle is that the optimal controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ maximize the Hamiltonian $\mathcal{H}$ (6) pointwise in time, almost surely, given the optimal state and costate processes $(X_{t}^{*},\mathbf{Y}_{t},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*})$.
-Assuming interior solutions allows us to derive the first-order conditions (FOCs) for optimality.
-Maximizing with respect to the consumption rate $C_{t}$ yields the standard FOC:
-$e^{-\delta t}\,U^{\prime}(C_{t}^{*})=\lambda_{t}^{*}.$ (9)
-This classical result links optimal consumption to the marginal value of wealth via the inverse of the marginal utility: $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$.
-For the portfolio weights $\bm{\pi}_{t}$, we differentiate $\mathcal{H}$ with respect to $\bm{\pi}$, accounting for both the drift and diffusion terms that depend on the control.
-Setting the gradient $\nabla_{\bm{\pi}}\mathcal{H}=\mathbf{0}$, we obtain:
 […]
-Dividing by $X_{t}^{*}$ assuming $X_{t}^{*}\neq 0$, the first-order condition for the optimal portfolio becomes:
-[…]
-This condition implicitly defines the optimal portfolio $\bm{\pi}_{t}^{*}$ through its dependence on the costate $\lambda_{t}^{*}$ and the martingale component $\mathbf{Z}_{t}^{\lambda X*}$, which captures the diffusion sensitivity of the marginal value of wealth.
-To obtain an explicit expression for $\bm{\pi}_{t}^{*}$, we must relate the coefficient $\mathbf{Z}^{\lambda X*}$ of the martingale component in the BSDE (7) to the derivatives of the value function.
-This relationship can be derived by applying Itô’s lemma to $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ and matching terms with the dynamics of the costate process in (7).
-The full derivation is presented in Appendix A.
-This procedure yields an expression for $\mathbf{Z}^{\lambda X*}$ in terms of the second-order derivatives $V_{xx}$, $V_{x\mathbf{y}}$, the optimal controls $\bm{\pi}_{t}^{*}$, and model parameters $(\bm{\sigma},\bm{\sigma}_{Y},\Psi,\rho)$.
-Substituting this into the first-order condition (10) and solving the resulting algebraic equation for $\bm{\pi}_{t}^{*}$ leads to the well-known optimal portfolio formula, originally derived by Merton, (1973):
-[…]
-$\quad-\frac{1}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,\bm{\rho}\,\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\,V_{x\mathbf{y}}.$
-Here, $\Sigma(t,\mathbf{Y}_{t}):=\bm{\sigma}(t,\mathbf{Y}_{t})\,\Psi\,\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ is the instantaneous covariance matrix of risky asset returns.
-The derivatives of the value function —$V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\nabla_{\mathbf{y}}\lambda_{t}^{*}\in\mathbb{R}^{k}$—are evaluated along the optimal trajectory $(t,X_{t}^{*},\mathbf{Y}_{t})$.
-This expression decomposes the optimal portfolio into two components: a myopic demand, which optimizes the static trade-off between expected return and risk under fixed investment opportunities, and an intertemporal hedging demand, which adjusts the portfolio to hedge against stochastic shifts in future investment opportunities.
-The hedging component is shaped by the correlation structure $\bm{\rho}$ between asset returns and state innovations, and by the cross-derivative $V_{x\mathbf{y}}$, which captures how the marginal value of wealth responds to changes in the economic state.
-For illustration, consider the simplified case with a single risky asset ($n=1$) and a single state variable ($k=1$). In this setting, the optimal portfolio allocation in (11) reduces to:
-$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{1}{X_{t}V_{xx}\sigma^{2}}\,\sigma\rho\sigma_{Y}\,V_{xy},$
-where $\sigma:=\sigma(t,Y_{t})$, $\sigma_{Y}:=\sigma_{Y}(t,Y_{t})$, and $V_{xy}:=\partial^{2}V/\partial x\partial y$.
-This expression makes explicit the decomposition of the optimal policy into two components: a myopic demand, which is proportional to the market price of risk $(\mu-r)/\sigma$, and a hedging demand, which is proportional to the cross-partial derivative $V_{xy}$ and the correlation coefficient $\rho$ between shocks to the asset return and the state variable.
-2.3 Applications to Financial Models
-2.3.1 A Multi-Asset Model with One OU Factor
-We first consider a model with $n$ risky assets whose excess returns are driven by a single Ornstein–Uhlenbeck (OU) factor.
-Specifically, each asset $i$ has constant volatility $\sigma_{i}>0$ and a time-varying risk premium proportional to an OU state variable $Y_{t}$, so that the expected return is $r+\alpha_{i}\sigma_{i}Y_{t}$, where $r$ is the constant risk-free rate.
-The state variable $Y_{t}$ evolves as:
-$dY_{t}=\kappa_{Y}(\theta_{Y}-Y_{t})\,dt+\sigma_{Y}\,dW_{t}^{Y},$
-where $\kappa_{Y}>0$ is the mean-reversion speed, $\theta_{Y}\in\mathbb{R}$ is the long-run mean, and $\sigma_{Y}>0$ is the volatility of the OU process.
-The Brownian motion $W_{t}^{Y}$ may be correlated with the return processes of the risky assets.
-The price process $S_{i,t}$ of the $i$-th risky asset follows:
-$\frac{dS_{i,t}}{S_{i,t}}=\bigl{(}r+\alpha_{i}\sigma_{i}Y_{t}\bigr{)}\,dt+\sigma_{i}\,dW_{i,t}^{S}.$
-Let $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$ be the vector of risk premia loadings, and $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})$ the diagonal matrix of asset volatilities.
-The Brownian motion driving the risky assets, $\mathbf{W}_{t}^{S}=(W_{1,t}^{S},\dots,W_{n,t}^{S})^{\top}$, has instantaneous covariance matrix $\mathrm{Cov}(d\mathbf{W}_{t}^{S})=\Psi\,dt$, where $\Psi\in\mathbb{R}^{n\times n}$.
-Let $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$ denote the vector of correlations between $W_{t}^{Y}$ and $W_{i,t}^{S}$.
-Let $X_{t}$ denote the investor’s wealth.
-The investor chooses a consumption rate $C_{t}\geq 0$ and portfolio weights $\bm{\pi}_{t}\in\mathbb{R}^{n}$ in the risky assets.
-The wealth dynamics is given by:
-$dX_{t}=\bigl{[}X_{t}\bigl{(}r+\bm{\pi}_{t}^{\top}Y_{t}(\bm{\sigma}\bm{\alpha})\bigr{)}-C_{t}\bigr{]}dt+X_{t}\,\bm{\pi}_{t}^{\top}\bm{\sigma}\,d\mathbf{W}_{t}^{S}.$
-The instantaneous covariance matrix of risky asset returns is $\Sigma=\bm{\sigma}\Psi\bm{\sigma}$.
-For an investor with CRRA utility $U(x)$ and discount rate $\delta>0$, the optimal consumption is given by $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$, where $\lambda_{t}^{*}$ is the costate variable associated with wealth $X_{t}$.
-The first-order condition from Pontryagin’s Maximum Principle yields the optimal portfolio:
-$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\,\Sigma^{-1}\left\{\lambda_{t}^{*}Y_{t}(\bm{\sigma}\bm{\alpha})+(\partial_{Y}\lambda_{t}^{*})\,\sigma_{Y}\,(\bm{\sigma}\bm{\rho})\right\}.$
-This expression depends on the costate $\lambda_{t}^{*}$ and its derivatives $\partial_{x}\lambda_{t}^{*}$ and $\partial_{Y}\lambda_{t}^{*}$.
-The first term inside the braces represents the myopic demand, driven by the instantaneous risk premia $Y_{t}(\bm{\sigma}\bm{\alpha})$.
-The second term captures the intertemporal hedging demand, which arises from uncertainty in future investment opportunities and is scaled by the correlation vector $\bm{\rho}$ and the sensitivity $\partial_{Y}\lambda_{t}^{*}$.
-In the special case of a single risky asset ($n=1$), the portfolio simplifies to:
-$\pi_{t}^{*}=-\frac{1}{\sigma^{2}X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\left\{\lambda_{t}^{*}\alpha\sigma Y_{t}+\sigma\rho\sigma_{Y}\,\partial_{Y}\lambda_{t}^{*}\right\}.$
-In the case without intermediate consumption ($C_{t}\equiv 0$), the value function and optimal controls can be derived analytically via the Hamilton–Jacobi–Bellman (HJB) approach, leading to a system of Riccati ODEs (Kim and Omberg, , 1996).
-Details are provided in Appendix B.1.
-2.3.2 A Multi-Asset, Multi-Factor Model
-We next extend the previous setup to a more general multi-factor environment, considering $n$ risky assets whose excess returns are driven by a $k$-dimensional Ornstein–Uhlenbeck (OU) factor vector $\mathbf{Y}_{t}\in\mathbb{R}^{k}$.
-The risk premium for asset $i=1,\dots,n$ is given by $\sigma_{i}(\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t})$, where $\bm{\alpha}_{i}\in\mathbb{R}^{k}$ is the vector of factor loadings for asset $i$, and $\sigma_{i}>0$ is the asset’s constant volatility.
-This formulation captures richer interactions between multiple sources of risk and asset returns.
-The factor process $\mathbf{Y}_{t}$ evolves as a multivariate OU process:
-$d\mathbf{Y}_{t}=\kappa_{Y}(\bm{\theta}_{Y}-\mathbf{Y}_{t})\,dt+\bm{\sigma}_{Y}\,d\mathbf{W}_{t}^{Y},$
-where $\kappa_{Y}\in\mathbb{R}^{k\times k}$ is the mean-reversion matrix, $\bm{\theta}_{Y}\in\mathbb{R}^{k}$ is the long-run mean, and $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ is a diagonal matrix of factor-specific volatilities (as introduced in Section 2.1).
-The Brownian motion $\mathbf{W}_{t}^{Y}\in\mathbb{R}^{k}$ driving the factors has instantaneous covariance $\mathrm{Cov}(d\mathbf{W}_{t}^{Y})=\Phi\,dt$.
-The price $S_{t}^{i}$ of each asset follows:
-$\frac{dS_{t}^{i}}{S_{t}^{i}}=\bigl{(}r+\sigma_{i}\,\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t}\bigr{)}\,dt+\sigma_{i}\,dW_{t}^{X,i},\quad i=1,\dots,n,$
-where $\mathbf{W}_{t}^{X}\in\mathbb{R}^{n}$ is a Brownian motion with instantaneous covariance $\mathrm{Cov}(d\mathbf{W}_{t}^{X})=\Psi\,dt$.
-The asset and factor Brownian motions may be correlated: $\mathrm{Cov}(dW_{t}^{X,i},dW_{t}^{Y,j})=\rho_{ij}\,dt$, with cross-correlation matrix $\bm{\rho}\in\mathbb{R}^{n\times k}$.
-Given portfolio weights $\bm{\pi}_{t}=(\pi_{t}^{1},\dots,\pi_{t}^{n})^{\top}$ and consumption rate $C_{t}\geq 0$, the investor’s wealth evolves as:
-$dX_{t}=\left[X_{t}\left(r+\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}\bm{\alpha}_{i}^{\top}\mathbf{Y}_{t}\right)-C_{t}\right]dt+X_{t}\sum_{i=1}^{n}\pi_{t}^{i}\sigma_{i}\,dW_{t}^{X,i}.$
-For CRRA utility with subjective discount rate $\delta>0$, the optimal consumption policy satisfies $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$, where $\lambda_{t}^{*}$ is the costate associated with wealth.
-The first-order condition from Pontryagin’s Maximum Principle yields the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$ as:
-$\bm{\pi}_{t}^{*}=-\;\frac{1}{X_{t}^{*}\,\partial_{x}\lambda_{t}^{*}}\,\Sigma^{-1}\left\{\lambda_{t}^{*}\,\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}+\bm{\sigma}\bm{\rho}\bm{\sigma}_{Y}\,\partial_{\mathbf{Y}}\lambda_{t}^{*}\right\},$
-where: $\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n})\in\mathbb{R}^{n\times n}$ is the volatility matrix, $\Sigma=\bm{\sigma}\Psi\bm{\sigma}\in\mathbb{R}^{n\times n}$ is the return covariance matrix, $\mathbf{A}=(\bm{\alpha}_{1},\dots,\bm{\alpha}_{n})^{\top}\in\mathbb{R}^{n\times k}$ is the matrix of factor loadings, $\bm{\sigma}_{Y}\in\mathbb{R}^{k\times k}$ is the factor volatility matrix, $\partial_{\mathbf{Y}}\lambda_{t}^{*}\in\mathbb{R}^{k}$ is the gradient of the costate with respect to $\mathbf{Y}_{t}$.
-This expression separates the optimal portfolio into two components: a myopic demand, driven by the instantaneous excess returns $\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}$, and an intertemporal hedging demand, which offsets the effects of changes in investment opportunities.
-The latter depends on the sensitivity of marginal utility to changes in $\mathbf{Y}_{t}$ ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$), and the correlation structure between asset and factor shocks (encoded in $\bm{\rho}$, $\Psi$, and $\Phi$).
-As in the single-factor case, when there is no intermediate consumption ($C_{t}\equiv 0$), one can derive analytic solutions using an exponential-quadratic value function ansatz, leading to a system of matrix Riccati ODEs.
-Detailed derivations are provided in Appendix B.2.
-3 A Gradient-Based Algorithm for Multi-Asset Policy Optimization
-While Pontryagin’s Maximum Principle (PMP) provides necessary optimality conditions via a system of forward–backward stochastic differential equations (FBSDEs; see Section 2.2), solving these analytically becomes infeasible in high dimensions.
-We therefore adopt a gradient-based numerical approach using deep learning.
-The core idea is to parameterize the investor’s policy $(\bm{\pi}_{t},C_{t})$ with neural networks $(\bm{\pi}_{\theta}(t,x,\mathbf{y}),C_{\phi}(t,x,\mathbf{y}))$, where $(\theta,\phi)$ denote the network parameters to be optimized.
-Our objective is to find parameters $(\theta,\phi)$ that maximize the expected utility delivered by the policy $(\bm{\pi}_{\theta},C_{\phi})$.
-To ensure robustness across a range of initial conditions, we define and maximize the extended value function, $\widetilde{J}(\theta,\phi)$, as the expected value of the conditional utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$, evaluated over initial states sampled from a distribution $\eta$ on a domain $\mathcal{D}$:
-[…]
-$\widetilde{J}(\theta,\phi)=\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\eta}\left[J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\right]$ (13)
-Here, $\mathbb{E}^{\theta,\phi}[\cdot\,|\,\cdot]$ denotes expectation under the dynamics induced by policy $(\theta,\phi)$.
-We develop and implement the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework to maximize $\widetilde{J}(\theta,\phi)$.
-PG-DPO employs automatic differentiation—specifically, backpropagation through time (BPTT)—applied to Monte Carlo simulations of system trajectories initialized from $\eta$.
-A key feature of this approach is that BPTT not only provides unbiased gradient estimates $\nabla\widetilde{J}$ for policy optimization, but also computes pathwise estimates of the Pontryagin costate variables $(\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)})$, capturing marginal value sensitivities as prescribed by PMP.
-This dual output enables two algorithmic variants: the baseline PG-DPO algorithm described in Section 3.1, and an advanced Two-Stage PG-DPO algorithm, which explicitly exploits costate information, presented in Section 3.2.
-3.1 Baseline PG-DPO for the Extended Objective
-The baseline PG-DPO algorithm aims to maximize the extended value function $\widetilde{J}(\theta,\phi)$ (defined in Eq.
-(13)) using mini-batch stochastic gradient ascent.
-The core idea is to approximate the gradient $\nabla\widetilde{J}$ by averaging gradients computed on individual sample paths, where each path starts from a randomly drawn initial state.
-Each iteration $j$ of the algorithm involves the following steps:
-  1. 1.
-    Sample Initial States: Draw a mini-batch of $M$ initial states $\{(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)})\}_{i=1}^{M}$ from the distribution $\eta$ over the domain $\mathcal{D}$.
-  2. 2.
-    Simulate Forward Paths: For each sampled initial state $i$, simulate a discrete-time trajectory $\{(X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\}_{k=0}^{N}$ forward from $t_{k}=t_{0}^{(i)}$ to $t_{N}=T$, using the current policy networks $(\bm{\pi}_{\theta},C_{\phi})$.
-The state evolution follows:
-    $\mathbf{Y}_{k+1}^{(i)}$ $=\mathbf{Y}_{k}^{(i)}+\bm{\mu}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta t^{(i)}+\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta\mathbf{W}_{k}^{(i),Y},$
-    $X_{k+1}^{(i)}$ $=X_{k}^{(i)}+\left[X_{k}^{(i)}\left((1-\bm{\pi}_{k}^{(i)\top}\mathbf{1})r(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})+\bm{\pi}_{k}^{(i)\top}\bm{\mu}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\right)-C_{k}^{(i)}\right]\Delta t^{(i)}$
-    $\quad+X_{k}^{(i)}\bm{\pi}_{k}^{(i)\top}\bm{\sigma}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta\mathbf{W}_{k}^{(i),X},$
-    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$, $\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})$ is the vector of diffusion coefficients applied element-wise to the components of $\Delta\mathbf{W}_{k}^{(i),Y}$, and the correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ are derived from the Cholesky factor $L$ of the covariance matrix $\bm{\Omega}$.
-  3. 3.
-    Calculate Realized Reward per Path: For each path $i$, compute the realized cumulative reward, explicitly noting its dependence on the starting state and policy parameters, denoted as $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)$:
-    $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)=\sum_{k=0}^{N-1}e^{-\delta(t_{k}^{(i)}-t_{0}^{(i)})}U\bigl{(}C_{\phi}(t_{k}^{(i)},X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\bigr{)}\Delta t^{(i)}+\kappa e^{-\delta(T-t_{0}^{(i)})}U(X_{N}^{(i)}).$
-    This value $J^{(i)}(t_{0}^{(i)},\dots)$ is a stochastic sample related to the conditional expected utility $J(t_{0}^{(i)},\dots)$ defined in Eq. (12).
-  4. 4.
-    Apply BPTT for Policy Gradient Estimate: Treat the simulation process for path $i$ as a computational graph.
-Apply automatic differentiation (BPTT) to compute the gradient of the realized reward $J^{(i)}(t_{0}^{(i)},\dots)$ with respect to the policy parameters $(\theta,\phi)$:
-    $\nabla_{(\theta,\phi)}J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi).$
-    This gradient serves as an unbiased estimate related to the gradient of the conditional expected utility $J(t_{0}^{(i)},\dots)$.
-  5. 5.
-    Estimate Gradient of $\widetilde{J}$: Obtain an unbiased estimate for the gradient of the overall extended objective $\widetilde{J}$ by averaging the individual path gradients over the mini-batch:
-    $\nabla\widetilde{J}(\theta,\phi)\approx\nabla_{\text{batch}}=\frac{1}{M}\sum_{i=1}^{M}\nabla_{(\theta,\phi)}J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi).$
-  6. 6.
-    Update Policy Parameters: Update the parameters $(\theta,\phi)$ using this averaged gradient estimate via stochastic gradient ascent:
-    $(\theta,\phi)_{j+1}\leftarrow(\theta,\phi)_{j}+\alpha_{j}\,\nabla_{\text{batch}},$
-    where $\alpha_{j}$ is the learning rate.
-By iterating these steps, the baseline PG-DPO algorithm gradually optimizes the policy $(\bm{\pi}_{\theta},C_{\phi})$ to maximize the extended value function $\widetilde{J}$.
-Algorithm 1 Baseline Pontryagin‑Guided Direct Policy Optimization (PG‑DPO)
-1:Inputs:
-2:Policy nets $\pi_{\theta},\,C_{\phi}$; learning‑rate schedule $\{\alpha_{j}\}_{j=0}^{J_{\max}-1}$; max iters $J_{\max}$, batch size $M$, time steps $N$; initial‑state distribution $\eta$; utility $U$, discount $\delta$, terminal weight $\kappa$; market params $(\bm{\mu},\bm{\sigma},r,\bm{\mu}_{Y},\bm{\sigma}_{Y}\text{ (k-dim vector of diff.
-coeffs.)},\bm{\Omega})$
-3:Output: optimised params $(\theta_{\text{opt}},\phi_{\text{opt}})$
-4:Initialise $\theta,\phi$
-5:for $j=0$ to $J_{\max}-1$ do
-6:     Draw mini‑batch $\{(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)})\}_{i=1}^{M}\sim\eta$
-7:     $\nabla_{\text{batch}}\leftarrow\mathbf{0}$
-8:     for $i=1$ to $M$ do
-9:         $(X_{0},\mathbf{Y}_{0})\leftarrow(x_{0}^{(i)},\mathbf{y}_{0}^{(i)})$, $\Delta t\leftarrow(T-t_{0}^{(i)})/N$
-10:         for $k=0$ to $N-1$ do
-11:              $t_{k}\leftarrow t_{0}^{(i)}+k\Delta t$
-12:              $\bm{\pi}_{k}\leftarrow\pi_{\theta}(t_{k},X_{k},\mathbf{Y}_{k})$
-13:              $C_{k}\leftarrow C_{\phi}(t_{k},X_{k},\mathbf{Y}_{k})$
-14:              Sample $\Delta\mathbf{W}_{k}\sim\mathcal{N}(\mathbf{0},\bm{\Omega}\Delta t)$
-15:              $\mathbf{Y}_{k+1}=\mathbf{Y}_{k}+\bm{\mu}_{Y}(t_{k},\mathbf{Y}_{k})\Delta t+\bm{\sigma}_{Y}(t_{k},\mathbf{Y}_{k})\,\Delta\mathbf{W}_{k}^{\text{Y}}$
-16:              $X_{k+1}=X_{k}+\Bigl{[}X_{k}\bigl{(}r(t_{k},\mathbf{Y}_{k})+\bm{\pi}_{k}^{\!\top}(\bm{\mu}(t_{k},\mathbf{Y}_{k})-r(t_{k},\mathbf{Y}_{k})\bm{1})\bigr{)}-C_{k}\Bigr{]}\Delta t+X_{k}\bm{\pi}_{k}^{\!\top}\bm{\sigma}(t_{k},\mathbf{Y}_{k})\,\Delta\mathbf{W}_{k}^{\text{X}}$
-17:         end for
-18:         $J^{(i)}=\sum_{k=0}^{N-1}e^{-\delta(t_{k}-t_{0}^{(i)})}U(C_{k})\Delta t+\kappa e^{-\delta(T-t_{0}^{(i)})}U(X_{N})$
-19:         $g^{(i)}\leftarrow\nabla_{(\theta,\phi)}J^{(i)}$
-20:         $\nabla_{\text{batch}}\leftarrow\nabla_{\text{batch}}+g^{(i)}$
-21:     end for
-22:     $(\theta,\phi)\leftarrow(\theta,\phi)+\dfrac{\alpha_{j}}{M}\nabla_{\text{batch}}$
-23:end for
-24:return $(\theta,\phi)$
-Role of intermediate BPTT values and the “Pontryagin-guided” nature.
-It is important to understand the role of the intermediate values computed during BPTT.
-By applying the chain rule backward from the final reward $J^{(i)}$, BPTT inherently calculates the sensitivity of the reward to the intermediate state variables.
-These sensitivities are the pathwise numerical counterparts to the Pontryagin costates:
-$\lambda_{k}^{(i)}=\frac{\partial J^{(i)}}{\partial X_{k}^{(i)}},\qquad\bm{\eta}_{k}^{(i)}=\frac{\partial J^{(i)}}{\partial\mathbf{Y}_{k}^{(i)}}.$ (14)
-To connect this computational recurrence to the formal PMP framework, we must generalize it to an $\mathcal{F}_{t_{k}}$-adapted process $\lambda_{k}=\mathbb{E}[\lambda_{k}^{(i)}\mid\mathcal{F}_{t_{k}}]$.
-This transition from a path-specific value to an adapted process reveals a deep structural correspondence, which we formalize below.
-Lemma 1 (Pathwise BPTT Recurrence).
-Along any simulated trajectory $i$, the BPTT algorithm computes the backward recursion for the pathwise sensitivity $\lambda_{k}^{(i)}$ as:
-$\lambda_{k}^{(i)}$ $=\lambda_{k+1}^{(i)}+\lambda_{k+1}^{(i)}\bigl{[}r_{k}+\bm{\pi}_{k}^{\!\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{]}\Delta t+\bigl{[}e^{-\delta(t_{k}-t_{0})}U^{\prime}(C_{k}^{(i)})-\lambda_{k+1}^{(i)}\bigr{]}\frac{\partial C_{k}^{(i)}}{\partial X_{k}^{(i)}}\,\Delta t$
-$\quad+\lambda_{k+1}^{(i)}\,\bm{\pi}_{k}^{\!\top}\bm{\sigma}_{k}\,\Delta\mathbf{W}_{k}^{X,(i)}.$ (15)
-Proof.
-The recurrence relation for the pathwise costate $\lambda_{k}^{(i)}:=\partial J^{(i)}/\partial X_{k}^{(i)}$ is derived by applying the chain rule.
-The total reward $J^{(i)}$ is decomposed into the utility from the current step $k$ and the total future reward $J_{k+1}^{(i)}$.
-$\lambda_{k}^{(i)}=\frac{\partial}{\partial X_{k}^{(i)}}\left(e^{-\delta(t_{k}-t_{0})}U(C_{k}^{(i)})\Delta t\right)+\frac{\partial J_{k+1}^{(i)}}{\partial X_{k}^{(i)}}.$
-The second term becomes $\lambda_{k+1}^{(i)}(\partial X_{k+1}^{(i)}/\partial X_{k}^{(i)})$.
-Differentiating the discrete wealth dynamics, which includes the consumption term $-C_{k}^{(i)}\Delta t$, with respect to $X_{k}^{(i)}$ yields:
-$\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}=1+\bigl{(}r_{k}+\bm{\pi}_{k}^{\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{)}\Delta t-\frac{\partial C_{k}^{(i)}}{\partial X_{k}^{(i)}}\Delta t+\bm{\pi}_{k}^{\top}\bm{\sigma}_{k}\Delta\mathbf{W}_{k}^{X,(i)}.$
-Substituting these components and rearranging the terms for $\lambda_{k}^{(i)}$ yields the expression in Eq. (1). ∎
-Lemma 2 (Quadratic-Covariation Drift).
-Let $\lambda_{k+1}^{(i)}=\lambda_{k+1}+\widetilde{\mathbf{Z}}^{\lambda X}_{k}\Delta\mathbf{W}_{k}^{X,(i)}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\Delta\mathbf{W}_{k}^{Y,(i)}$ be the martingale representation of $\lambda_{k+1}^{(i)}$, where $\widetilde{\mathbf{Z}}^{\lambda X}_{k}$ is a row vector.
-The conditional expectation of the stochastic term in Eq.
-(1) is:
-$\mathbb{E}\!\Bigl{[}\lambda_{k+1}^{(i)}(\bm{\pi}_{k}^{\!\top}\bm{\sigma}_{k}\,\Delta\mathbf{W}_{k}^{X,(i)})\Bigm{|}\mathcal{F}_{t_{k}}\Bigr{]}\;=\;\bigl{(}\widetilde{\mathbf{Z}}^{\lambda X}_{k}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\bm{\rho}_{Y\!X}\bigr{)}(\bm{\sigma}_{k}^{\!\top}\bm{\pi}_{k})\,\Delta t.$
-Proof.
-The product inside the expectation is an inner product of vectors.
-The term involving the adapted part, $\mathbb{E}[\lambda_{k+1}(\dots)|\mathcal{F}_{t_{k}}]$, is zero as it is the expectation of a martingale increment.
-The non-zero terms arise from the quadratic covariation of the raw martingale components of $\lambda_{k+1}^{(i)}$ with the wealth shock.
-Using the standard properties $\mathbb{E}[(\Delta\mathbf{W}_{k}^{X})(\Delta\mathbf{W}_{k}^{X})^{\top}|\mathcal{F}_{t_{k}}]=I\Delta t$ and $\mathbb{E}[(\Delta\mathbf{W}_{k}^{Y})(\Delta\mathbf{W}_{k}^{X})^{\top}|\mathcal{F}_{t_{k}}]=\bm{\rho}_{Y\!X}\Delta t$, where $\bm{\rho}_{Y\!X}$ is the cross-correlation matrix, yields the stated result.
-∎
-Theorem 3 (BPTT–PMP Correspondence).
-Define the effective martingale component as $\mathbf{Z}^{\lambda X}_{k}:=\widetilde{\mathbf{Z}}^{\lambda X}_{k}+\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\bm{\rho}_{Y\!X}$.
-The adapted costate process $(\lambda_{k})_{k=0}^{N}$ satisfies the discrete-time backward stochastic difference equation:
-$\lambda_{k}$ $=\lambda_{k+1}+\Bigl{[}\lambda_{k+1}\bigl{(}r_{k}+\bm{\pi}_{k}^{\!\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{)}+\mathbf{Z}^{\lambda X}_{k}(\bm{\sigma}_{k}^{\!\top}\bm{\pi}_{k})+\bigl{(}e^{-\delta t_{k}}U^{\prime}(C_{k})-\lambda_{k+1}\bigr{)}\frac{\partial C_{k}}{\partial X_{k}}\Bigr{]}\Delta t$
-$\quad-\widetilde{\mathbf{Z}}^{\lambda X}_{k}\,\Delta\mathbf{W}^{X}_{k}-\widetilde{\mathbf{Z}}^{\lambda Y}_{k}\,\Delta\mathbf{W}^{Y}_{k}.$ (16)
-Furthermore, the drift term in the square brackets of Eq.
-(3) is identical to the partial derivative of the Hamiltonian, $\partial_{x}\mathcal{H}$, evaluated at the corresponding state and control variables.
-Proof.
-We begin by taking the conditional expectation $\mathbb{E}[\cdot\mid\mathcal{F}_{t_{k}}]$ of the entire pathwise recurrence in Eq.
-(1) from Lemma 1.
-The left-hand side becomes $\lambda_{k}$ by definition.
-On the right-hand side, the expectations of the drift terms are straightforward.
-The crucial step is handling the stochastic term $\mathbb{E}[\lambda_{k+1}^{(i)}(\dots)|\mathcal{F}_{t_{k}}]$, for which we substitute the result from Lemma 2.
-This transforms the stochastic term into the additional drift component involving the effective process $\mathbf{Z}^{\lambda X}_{k}$.
-After collecting all drift terms and separating the martingale components, we arrive at the discrete-time BSDE presented in Eq.
-(3).
-The final step involves a direct algebraic comparison of this drift term with the definition of the Hamiltonian $\mathcal{H}$ from Section 2.2, which confirms their identity.
-∎
-Remark 1.
-The Martingale Representation Theorem, cited from Karatzas and Shreve, (1991), is a continuous-time result.
-Its application in our discrete-time grid setting is justified as the discrete process can be viewed as a sampling of an underlying continuous process, for which the representation holds on each interval $[t_{k},t_{k+1}]$.
-This theorem provides the theoretical guarantee that our “Pontryagin-guided” approach is well-founded.
-BPTT is not merely a generic optimization heuristic; it is a computational procedure that implicitly solves the necessary optimality conditions of PMP.
-While the baseline PG-DPO uses this powerful link implicitly to find effective policy gradients, our Projected PG-DPO variant leverages it explicitly by taking the BPTT-computed costates and using them to construct a policy that is structurally consistent with financial theory.
-3.2 Projected PG-DPO (P-PGDPO) Algorithm
-While the baseline PG-DPO algorithm described in Section 3.1 effectively optimizes the extended objective $\widetilde{J}$ by training the policy networks $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end, it may require a large number of iterations for the networks to fully converge.
-The Projected PG-DPO (P-PGDPO) variant offers a potentially more computationally efficient alternative, based on the empirical observation that the costate estimates—$(\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)})$ and their relevant derivatives (e.g., $\partial_{x}\lambda_{k}^{(i)},\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$)—often stabilize and become reasonably accurate well before the policy networks reach convergence.
-P-PGDPO leverages these rapidly stabilizing costate estimates to construct near-optimal controls directly, by projecting them onto the manifold of policies that satisfy the PMP first-order conditions.
-In this way, it bypasses the need for fully converged policy networks in order to deploy effective controls.
-The procedure unfolds in two distinct stages, as described below using a single-level enumeration:
-  1. 1.
-    Stage 1: Warm-Up Phase for Costate Estimation
-    First, we execute the Baseline PG-DPO procedure (see Section 3.1), which optimizes $\widetilde{J}$, for a predetermined, relatively small number of iterations $K_{0}$.
-The primary objective of this stage is not to obtain fully optimized networks $(\bm{\pi}_{\theta},C_{\phi})$, but rather to run the BPTT process sufficiently long for the underlying mechanism generating the costate estimates ($\lambda_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial X_{k}^{(i)}$, $\bm{\eta}_{k}^{(i)}=\partial J^{(i)}(\dots)/\partial\mathbf{Y}_{k}^{(i)}$) and their relevant state-derivatives (e.g., $\partial_{x}\lambda_{k}^{(i)}$, $\partial_{\mathbf{Y}}\lambda_{k}^{(i)}$) to stabilize.
-We let $(\theta^{*},\phi^{*})$ denote the policy parameters obtained at the end of this warm-up phase.
-  2. 2.
-    Stage 2: Analytic Control Deployment
-    After completing the $K_{0}$ warm-up iterations, the trained policy networks $(\bm{\pi}_{\theta^{*}},C_{\phi^{*}})$ are typically not directly used for generating controls during deployment.
-Instead, to determine the control actions for any given state $(t_{k},X_{k},\mathbf{Y}_{k})$ encountered during deployment (let this deployment time be $t_{k}=t_{deploy}$), we employ a Monte Carlo-like averaging procedure to obtain reliable costate estimates needed for the PMP formulas:
-    First, to obtain the stabilized costate estimates corresponding to the state $(t_{deploy},X_{k},\mathbf{Y}_{k})$, we perform the following averaging: Simulate multiple ($M_{\text{MC}}$) forward paths starting from $(t_{deploy},X_{k},\mathbf{Y}_{k})$ (or states very close to it) up to the horizon $T$, using the fixed policy parameters $(\theta^{*},\phi^{*})$ obtained from Stage 1.
-For each simulated path $j=1,\dots,M_{\text{MC}}$, calculate its realized reward $J^{(j)}(t_{deploy},X_{k},\mathbf{Y}_{k};\theta^{*},\phi^{*})$ (using discounting relative to $t_{deploy}$ as in Algorithm 2) and apply BPTT to obtain the costate estimates $\lambda_{deploy}^{(j)},\partial_{x}\lambda_{deploy}^{(j)},\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$ specifically at the starting node $t_{deploy}$ for that path $j$.
-The final “stabilized” estimates are then computed by averaging these values across the $M_{\text{MC}}$ paths:
-    $\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\lambda_{deploy}^{(j)},\quad\partial_{x}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{x}\lambda_{deploy}^{(j)},\quad\partial_{\mathbf{Y}}\lambda_{deploy}\approx\frac{1}{M_{\text{MC}}}\sum_{j=1}^{M_{\text{MC}}}\partial_{\mathbf{Y}}\lambda_{deploy}^{(j)}$
-    This averaging acts as a Monte Carlo estimation step, significantly reducing the variance inherent in the single-path BPTT estimates derived during Stage 1 and providing more robust inputs for the PMP formulas.
-    Algorithm 2 Projected PG‑DPO (P-PGDPO): Warm‑up Training & Deployment
-    1:Stage 1: Warm‑up Training
-    2:Inputs of Algorithm 1; warm‑up iterations $K_{0}$
-    3:Stabilised parameters $(\theta^{*},\phi^{*})$
-    4:Run Algorithm 1 for $K_{0}$ iterations to obtain $(\theta^{*},\phi^{*})$
-    5:
-    6:Stage 2: Deployment at state $(t,X,\bm{Y})$
-    7:State $(t,X,\bm{Y})$; parameters $(\theta^{*},\phi^{*})$; Monte‑Carlo rollouts $M_{\mathrm{MC}}$, steps $N^{\prime}$
-    8:Near‑optimal controls $(\bm{\pi}^{\mathrm{PMP}},C^{\mathrm{PMP}})$
-    9:Initialise $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}\leftarrow\varnothing$
-    10:for $j=1$ to $M_{\mathrm{MC}}$ do
-    11:     Simulate path from $(t,X,\bm{Y})$ with $(\pi_{\theta^{*}},C_{\phi^{*}})$ using $N^{\prime}$ steps. Let $t_{0}^{\prime}=t$.
-    12:     Compute realized reward for path $j$:
-    13:     $\qquad\displaystyle J^{(j)}\leftarrow\sum_{n=0}^{N^{\prime}-1}e^{-\delta(t_{n}-t_{0}^{\prime})}U\bigl{(}C_{t_{n}}^{(j)}\bigr{)}\,\Delta t^{\prime}+\kappa e^{-\delta(T-t_{0}^{\prime})}U\bigl{(}X_{T}^{(j)}\bigr{)}$
-    14:     Evaluate $\lambda_{t}^{(j)}$, $\partial_{x}\lambda_{t}^{(j)}$, $\partial_{\bm{Y}}\lambda_{t}^{(j)}$ (costates at $t_{0}^{\prime}=t$) via BPTT on $J^{(j)}$ and append to lists $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}$.
-    15:end for
-    16:Compute $\hat{\lambda}_{t},\widehat{\partial_{x}\lambda}_{t},\widehat{\partial_{\bm{Y}}\lambda}_{t}$triangleright E.g., by averaging estimates from lists $\mathcal{L},\mathcal{L}_{x},\mathcal{L}_{Y}$
-    17:$C^{\mathrm{PMP}}\leftarrow(U^{\prime})^{-1}\!\bigl{(}e^{\delta t}\hat{\lambda}_{t}\bigr{)}$triangleright Using PMP FOC Eq. (8)
-    18:Compute $\bm{\pi}^{\mathrm{PMP}}$ via PMP FOC Eq. (15) $\triangleright$ Using estimated costates/derivatives
-    19:return $(\bm{\pi}^{\mathrm{PMP}},C^{\mathrm{PMP}})$
-    Next, we plug these averaged costate estimates ($\lambda_{deploy},\partial_{x}\lambda_{deploy},\partial_{\mathbf{Y}}\lambda_{deploy}$) directly into the analytical formulas derived from the Pontryagin Maximum Principle’s first-order conditions (Section 2.2).
-The optimal consumption at $t_{deploy}$ is calculated as:
-    $C_{deploy}^{\mathrm{PMP}}=\bigl{(}U^{\prime}\bigr{)}^{-1}\!\bigl{(}e^{\delta\,t_{deploy}}\,\lambda_{deploy}\bigr{)}.$
-    The optimal investment portfolio at $t_{deploy}$ is calculated as:
-    $\bm{\pi}_{deploy}^{\mathrm{PMP}}$ $=-\frac{1}{X_{k}\,\partial_{x}\lambda_{deploy}}\,\Sigma^{-1}\!\bigl{(}t_{deploy},\mathbf{Y}_{k}\bigr{)}\Biggl{\{}\lambda_{deploy}\bigl{[}\bm{\mu}(t_{deploy},\mathbf{Y}_{k})-r(t_{deploy},\mathbf{Y}_{k})\mathbf{1}\bigr{]}$
-    $\qquad+\bigl{(}\bm{\sigma}(t_{deploy},\mathbf{Y}_{k})\,\bm{\rho}\,\bm{\sigma}_{Y}(t_{deploy},\mathbf{Y}_{k})\bigr{)}(\partial_{\mathbf{Y}}\lambda_{deploy})\Biggr{\}}.$ (17)
-    (This formula aligns with Eq.
-(11), where $\lambda_{deploy}$ corresponds to $V_{x}$, $\partial_{x}\lambda_{deploy}$ to $V_{xx}$, and $\partial_{\mathbf{Y}}\lambda_{deploy}$ to $V_{x\mathbf{Y}}$ at state $(t_{deploy},X_{k},\mathbf{Y}_{k})$.)
-    Finally, these analytically computed values $(C_{deploy}^{\mathrm{PMP}},\bm{\pi}_{deploy}^{\mathrm{PMP}})$, based on the averaged costate estimates, are used as the control actions for the state $(t_{deploy},X_{k},\mathbf{Y}_{k})$ during deployment.
-In summary, the Projected PG-DPO (P-PGDPO) approach begins with a short warm-up phase (Stage 1), using the baseline PG-DPO procedure to obtain a preliminary policy $(\theta^{*},\phi^{*})$ whose backpropagation-through-time (BPTT) produces reasonably stable costate estimates.
-In Stage 2, these costate estimates are refined via Monte Carlo averaging at deployment time and then substituted into the analytical expression for the optimal control derived from PMP.
-This two-stage strategy can substantially reduce computational costs while delivering high-quality, near-optimal controls.
-The Monte Carlo averaging step in Stage 2 is essential for obtaining reliable inputs to the PMP formula, mitigating the noise inherent in single-path estimates.
-3.3 Rationale for the Projected PG-DPO (P-PGDPO) Approach
-In continuous-time financial models, where the dynamics of investor wealth $X_{t}$ and exogenous state variables $\mathbf{Y}_{t}$ are governed by stochastic differential equations (SDEs), the value function $V(t,X_{t},\mathbf{Y}_{t})$ for a dynamic optimization problem typically satisfies a Hamilton–Jacobi–Bellman (HJB) partial differential equation (PDE) (e.g., Fleming and Soner, , 2006; Yong and Zhou, , 2012).
-The HJB equation is parabolic, and the specific nature of its parabolicity is primarily determined by the diffusion coefficients in the underlying SDEs—namely $\bm{\sigma}$ (for asset dynamics) and $\bm{\sigma}_{Y}$ (for state dynamics).
-While first-order drift terms (involving $\bm{\mu},\bm{\mu}_{Y}$) influence the structure of the solution, they do not affect the parabolicity classification, which is governed by the second-order terms.
-The behavior of the diffusion matrix $a(t,x,\mathbf{y})$, which multiplies the second-order derivatives in the HJB equation, allows the PDE to be classified into different parabolic types (see, e.g., Evans, , 2022).
-This classification significantly impacts the regularity (smoothness) of the value function $V$ and its associated costates $\nabla V$ (cf.
-Evans, , 2022, Chs.
-6–7).
-Costates represent the sensitivity of the value function with respect to states and wealth, and Pontryagin’s Maximum Principle links these costates directly to the optimal policy.
-The core idea of the Projected PG-DPO (P-PGDPO) method is to estimate these costates—and their derivatives—via BPTT and substitute them into the analytical expressions for the optimal control provided by PMP.
-The regularity of costates, governed by the parabolicity class of the HJB equation, directly influences the accuracy of these estimates and thus the effectiveness of P-PGDPO.
-Table 1: Classification of Parabolic HJB PDEs in Financial Models
-Class Diffusion Matrix $a(t,x,\mathbf{y})$ Property SDE Coefficient Requirements and Dynamic Portfolio Examples
-(U) Uniformly Elliptic Uniformly positive definite across the state space The diffusion coefficients $\bm{\sigma}$ (for wealth $X_{t}$) and $\bm{\sigma}_{Y}$ (for state $\mathbf{Y}_{t}$) are such that the resulting diffusion matrix in the HJB is non-zero and bounded below by a positive constant everywhere (e.g., $\bm{\sigma}\Psi\bm{\sigma}^{\top}$ and $\operatorname{diag}(\bm{\sigma}_{Y})\Phi\operatorname{diag}(\bm{\sigma}_{Y})^{\top}$ uniformly positive definite).
-Drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) are also sufficiently regular (e.g., Lipschitz).
-($\implies$ Value Function / Costate very smooth) Examples: Merton’s portfolio problem with constant coefficients ($\mu,\sigma>0,r$ are constants); models with state-dependent volatility that remains strictly positive; the multi-asset OU model in Section 2.3.1 (if asset volatilities $\sigma_{i}>0$, factor volatility $\sigma_{Y}>0$, and $\Psi$ is positive definite).
-(H) Hypo-elliptic May degenerate (even globally).
-Regularity is achieved if the system structure satisfies conditions like Hörmander’s (i.e., Lie brackets of vector fields span the tangent space).
-The problem might be uniformly elliptic on compact subsets of interest.
-Diffusion coefficients ($\bm{\sigma}$ or $\bm{\sigma}_{Y}$) can become zero at certain points within the domain.
-However, interaction with non-zero drift terms or other diffusion terms maintains regularity.
-($\implies$ Value Function / Costate smooth in relevant regions) Examples: The multi-asset multi-factor model in Section 2.3.2 (if some factor volatilities in $\bm{\sigma}_{Y}$ can be zero at certain states but the overall system structure, possibly involving $\kappa_{Y}$, ensures hypo-ellipticity); a Vasicek interest rate model affecting asset returns $\mu$ but not $\sigma$.
-(D) Degenerate at the Boundary Degenerates ($\to 0$) as the state approaches a boundary.
-The diffusion coefficients $\bm{\sigma}$ or $\bm{\sigma}_{Y}$ (or the combined diffusion matrix) approach zero as a state variable (e.g., wealth, volatility, interest rate) approaches a boundary (typically 0).
-($\implies$ Value Function / Costate regularity may break down near the boundary) Examples: CIR (Cox-Ingersoll-Ross) interest rate/volatility model (diffusion term like $\sigma_{Y}\sqrt{Y_{t}}$); Heston stochastic volatility model (asset diffusion term like $\sigma\sqrt{\nu_{t}}$).
-The reason this formal justification is crucial lies in a fundamental challenge facing many simulation-based optimization methods.
-A critical issue in direct policy optimization, including standard reinforcement learning, is that a learned policy can achieve a near-optimal objective value (a small value gap) while being structurally very different from the true optimal policy (a large policy gap).
-This occurs because the objective function is often locally “flat” near the optimum, meaning first-order policy deviations may only cause second-order utility losses.
-Our framework circumvents this issue by shifting focus from the value itself to the satisfaction of Pontryagin’s first-order conditions (FOCs).
-The success of this approach hinges on a key property of the HJB equations common in finance: their parabolic nature.
-Parabolic PDEs, akin to the heat equation, enforce a “smoothing” effect, ensuring that a well-behaved value function also has well-behaved derivatives (costates).
-This is not merely an intuitive notion; it is mathematically formalized by regularity theory in Sobolev spaces, which guarantees that convergence in value implies convergence of its derivatives.
-This regularity is precisely what allows us to trust the BPTT-derived costates and reconstruct a robust policy via the PMP formula.
-Building on this expected regularity, we provide a more formal justification for P-PGDPO.
-We aim to bound the policy gap, $\|\widehat{\pi}-\pi^{*}\|$, which is the error between our P-PGDPO policy $\widehat{\pi}$ and the true optimal policy $\pi^{*}$.
-The bound depends on two primary sources of error.
-The first is the FOC residual, denoted by $\varepsilon$, which measures how much the policy from the warm-up stage violates the optimality conditions of PMP.
-The second is the BPTT estimation error, $\delta_{\mathrm{BPTT}}$, in the costate estimate $\widehat{\lambda}$.
-This estimation error itself stems from two numerical sources: the time discretization of the SDEs (with time step $\Delta t$) and the Monte Carlo sampling used for the expectation (with batch size $M$).
-The following theorem encapsulates this relationship, with full technical details provided in Appendix C.
-{restatable}
-[Policy Gap Bound]theoremthm:policy_gap Let $\pi^{*}$ be the true optimal policy and $\widehat{\pi}$ be the policy generated by the P-PGDPO algorithm using a time step $\Delta t$ and batch size $M$.
-Under the regularity conditions specified in Assumption 6 in the Appendix, the gap between the P-PGDPO policy and the true optimum is bounded as follows:
-$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\text{tot}}\left(\varepsilon+\kappa_{1}\Delta t+\frac{\kappa_{2}}{\sqrt{M}}\right)$
-where:
-  • $\varepsilon$ is the $L^{q,p}$-norm of the Pontryagin FOC residual from the warm-up policy.
-  • $\kappa_{1},\kappa_{2}$ are positive constants governing the BPTT estimation error, which arises from time discretization (error proportional to $\Delta t$) and Monte Carlo sampling (error proportional to $1/\sqrt{M}$).
-  • $C_{\text{tot}}$ is a positive constant that depends on the model parameters but not on $\varepsilon$, $\Delta t$, or $M$.
-The proof of Theorem 1, detailed in Appendix C, quantifies the intuition discussed above.
-The logical chain proceeds as follows: a small FOC-gap ($\varepsilon$) ensures that the policy being evaluated is “close” to satisfying the PDE’s structural requirements.
-Due to the smoothing property of the uniformly parabolic PDE—guaranteed by regularity theory in Sobolev spaces (e.g., Dong and Kim, (2009))—this “closeness” is not confined to the value function alone but propagates to its derivatives.
-With the gradients (costates) being verifiably close to the optimal ones, the final step of projecting them through the PMP map becomes a well-posed, stable operation, leading to a policy that is also close to the true optimum.
-Our formal analysis demonstrates that under suitable regularity conditions, particularly uniform parabolicity, a small Pontryagin FOC gap ($\varepsilon$) from the warm-up stage guarantees that the P-PGDPO policy $\widehat{\pi}$ is provably close to the true optimum $\pi^{*}$, up to numerical errors from simulation.
-The bound explicitly shows that to achieve a highly accurate policy, one must not only ensure the warm-up is effective (small $\varepsilon$) but also use a sufficiently small time step ($\Delta t$) and a large batch size ($M$) to control the BPTT estimation error.
-When the objective function is relatively flat near the optimum, the FOC residual $\varepsilon$ for the warm-up policy tends to be small, making the numerical error terms the primary drivers of the final policy gap.
-This provides a rigorous justification for the core idea of the P-PGDPO method: substituting these numerically-controlled and rapidly stabilizing costate estimates into the analytical PMP formulas yields a near-optimal policy.
-However, extending this rigorous proof framework to settings beyond uniform parabolicity, such as hypo-elliptic (H-class, (Hörmander, , 1967)) or boundary-degenerate (D-class, (Oleinik, , 2012)) models, or to general non-affine systems, faces significant analytical challenges, as key assumptions (like global $L^{p}$ regularity or the costate floor) may no longer hold.
-Despite these formal hurdles, the fundamental algorithmic rationale of P-PGDPO remains powerfully persuasive for broader applicability.
-The core innovation is the decoupling of costate estimation from direct policy construction.
-Standard end-to-end methods often conflate these two tasks, making it difficult to learn the policy’s fine structure in a ‘flat’ optimization landscape.
-P-PGDPO, by contrast, leverages BPTT for what it excels at—efficient gradient (costate) estimation—and then applies the robust, analytical structure of PMP to deterministically map these estimates to a high-fidelity policy.
-This principle, strongly supported by our numerical results, suggests that P-PGDPO is a practical and powerful tool for a wider class of challenging continuous-time control problems, even in cases where the conditions required for the proof may not strictly hold.
-4 Numerical Experiments
-In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework, including both the baseline and Projected PG-DPO (P-PGDPO) variants.
-The experiments are based on the multi-asset, multi-factor Ornstein–Uhlenbeck (OU) portfolio optimization problem introduced in Section 2.3.2.
-For this setting, we specialize the general model by assuming no intermediate consumption (i.e., $C_{t}=0$) and focus exclusively on utility from terminal wealth, with $\kappa=1$, so that the objective becomes $J=\mathbb{E}[U(X_{T})]$.
-We deliberately begin our investigation with a relatively short investment horizon of $T=1.5$ years.
-This strategic choice enables a focused evaluation of the intrinsic performance and convergence characteristics of our proposed frameworks, isolating their behavior from the compounding numerical challenges inherent in long-horizon simulations.
-By establishing a clear performance baseline in this controlled setting, we lay the groundwork for a more rigorous analysis of long-term robustness in Section 4.3, where we introduce specific techniques to address those challenges.
-This affine setting admits a semi-analytical benchmark solution, derived from the associated HJB equation, which we use as a reference for evaluation (see Appendix B.2).
-To further assess the effectiveness of our approach, we compare PG-DPO with a state-of-the-art deep learning method for high-dimensional control problems—the Deep BSDE framework (Han et al., , 2018; Weinan, , 2017).
-The Deep BSDE method reformulates the stochastic control problem as a system of backward stochastic differential equations (BSDEs).
-Neural networks are trained to approximate key solution components, typically related to the value function and its gradient.
-These approximations are then used to construct the control policy, often by enforcing or approximating first-order optimality conditions derived from the Hamilton–Jacobi–Bellman equation.
-The networks are trained by minimizing residuals in the BSDE dynamics and terminal condition using Monte Carlo simulation.
-We evaluate the control policies produced by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and the Deep BSDE method (when used for control) against the semi-analytical benchmark.
-The evaluation metric is Root Mean Squared Error (RMSE) in portfolio weights, assessed across a range of problem sizes, including up to $n=50$ assets and $k=10$ state factors.
-Complete details on model parameters, PG-DPO and Deep BSDE implementations (including neural network architectures and training protocols), benchmark computation, and evaluation methodology are provided in Appendix D.
-The following subsections present and analyze the results of these comparative experiments.
-4.1 Convergence Speed and Accuracy Analysis
-(a) $n=1,\;k=10$
-(b) $n=10,\;k=10$
-(c) $n=50,\;k=10$
-Figure 1: Average policy RMSE comparison between Baseline PG-DPO (e.g., blue), Projected PG-DPO (P-PGDPO) (e.g., orange), and Deep BSDE (e.g., green) relative to the benchmark solution across different asset dimensions ($n$) with $k=10$ factors.
-Values averaged over $k=10$ evaluation ‘slices.’ For each slice $j=1,\dots,10$, the evaluation (e.g., for RMSE calculation or visualization) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$.
-Further details on the evaluation methodology are provided in Appendix D.
-Note the logarithmic scale on the $y$-axis.
-This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO framework—including both the Baseline and Projected PG-DPO (P-PGDPO) variants—against the Deep BSDE method, when applied to derive control policies.
-We evaluate performance across various problem dimensions, focusing on how the number of assets ($n$) and the number of state factors ($k$) affect policy accuracy.
-Figure 1 shows the average Root Mean Squared Error (RMSE) between the learned or derived policies and the semi-analytical benchmark policy.
-The RMSE is plotted against the number of training iterations on a logarithmic scale for experiments with $k=10$ state variables and varying asset counts ($n=1,10,50$).
-For P-PGDPO, the reported RMSE corresponds to the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ (Eq.
-(11)), which combines the myopic and intertemporal hedging components.
-As discussed in Section 4.2, the myopic error is typically negligible—on the order of $O(10^{-7})$.
-Thus, the total RMSE for P-PGDPO is effectively governed by the accuracy of the hedging component, which depends on estimating the costate gradient $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ via BPTT.
-Table 2: Summary of Minimum Policy RMSE and Convergence Iterations
-Dimension Method Min. Policy RMSE Iterations at Min.
-($n$ assets, $k=10$ factors)
-n=1, k=10 Baseline PG-DPO $1.751\times 10^{-2}$ 2800
-P-PGDPO $4.710\times 10^{-4}$ 1200
-Deep BSDE $8.146\times 10^{-2}$ 1
-n=10, k=10 Baseline PG-DPO $4.996\times 10^{-2}$ 9800
-P-PGDPO $2.799\times 10^{-3}$ 4800
-Deep BSDE $7.623\times 10^{-2}$ 200
-n=50, k=10 Baseline PG-DPO $1.070\times 10^{-1}$ 8800
-P-PGDPO $1.522\times 10^{-2}$ 10000
-Deep BSDE $8.199\times 10^{-2}$ 1200
-Note: Minimum policy RMSE values are averaged across $k$ evaluation slices.
-For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation) focuses on state factor $Y_{j}$, while all other factors $Y_{\ell}$, for $\ell\neq j$, are fixed at their respective long-run means $\theta_{Y,\ell}$.
-Further details on the evaluation procedure are provided in Appendix D.
-The iteration numbers indicate when the minimum average RMSE was observed; evaluations and logging were performed every 200 iterations.
-Table 2 summarizes the minimum achieved policy RMSE values and the corresponding iteration counts at which these minima occur, across $k=10$ evaluation ‘slcies’ with varying numbers of assets $n$.
-In each evaluation slice $j=1,\dots,10$, the RMSE and visualization are computed with respect to state factor $Y_{j}$, while all other factors $Y_{\ell}$, for $\ell\neq j$, are fixed at their respective long-term means $\theta_{Y,\ell}$.
-Further details on the evaluation methodology are provided in Appendix D.
-A consistent and notable finding is the superior performance of the P-PGDPO method.
-It achieves substantially lower RMSEs than both Baseline PG-DPO and Deep BSDE across all tested asset dimensions.
-For example, in the $n=1$ case, P-PGDPO attains a minimum RMSE of $4.710\times 10^{-4}$, significantly outperforming Baseline PG-DPO ($1.751\times 10^{-2}$) and Deep BSDE ($8.146\times 10^{-2}$).
-This result highlights the effectiveness of P-PGDPO in delivering high-accuracy control policies.
-The comparison between Baseline PG-DPO and Deep BSDE is more nuanced.
-Deep BSDE often reaches its minimum RMSE in fewer iterations—e.g., 1 iteration for $n=1$, and 200 iterations for $n=10$—compared to 2800 and 9800 iterations, respectively, for Baseline PG-DPO (see Table 2).
-This suggests faster initial convergence for Deep BSDE.
-However, its final RMSE is not consistently lower than that of Baseline PG-DPO.
-For both $n=1$ and $n=10$, Baseline PG-DPO ultimately achieves better accuracy: $1.751\times 10^{-2}$ vs.
-$8.146\times 10^{-2}$, and $4.996\times 10^{-2}$ vs.
-$7.623\times 10^{-2}$, respectively.
-In the high-dimensional case of $n=50$, however, Deep BSDE achieves a lower final RMSE ($8.199\times 10^{-2}$) than Baseline PG-DPO ($1.070\times 10^{-1}$).
-As illustrated in Figure 1, after early convergence, Deep BSDE’s RMSE may plateau or exhibit greater variance, while Baseline PG-DPO often demonstrates more gradual but continued improvement in certain configurations.
-P-PGDPO maintains a significant accuracy advantage even as $n$ increases to 50.
-However, all methods exhibit rising RMSE with increasing $n$ (see Table 2), reflecting the inherent challenges of high-dimensional asset allocation—potentially exacerbated by issues such as the ill-conditioning of the covariance matrix $\Sigma$, as discussed later.
-To further investigate the distinct effects of asset dimensionality ($n$) versus factor dimensionality ($k$), we conducted two additional sets of experiments.
-First, Table 3 reports results for the case where the number of assets is fixed at $n=1$, and the number of state factors is varied ($k=1,5,10$).
-In this setting, P-PGDPO consistently achieves the lowest RMSE across all configurations.
-Specifically, its RMSE is $1.120\times 10^{-3}$ for $k=1$, improves to $2.500\times 10^{-4}$ for $k=5$, and is $4.710\times 10^{-4}$ for $k=10$.
-While not strictly monotonic, P-PGDPO demonstrates strong performance across all $k$, with the best result observed at $k=5$.
-The performances of Baseline PG-DPO and Deep BSDE are more varied. For $n=1$:
-  • At $k=1$: Deep BSDE ($9.350\times 10^{-3}$) outperforms Baseline PG-DPO ($3.260\times 10^{-2}$).
-  • At $k=5$: Deep BSDE ($3.094\times 10^{-2}$) outperforms Baseline PG-DPO ($3.714\times 10^{-2}$).
-  • At $k=10$: Baseline PG-DPO ($1.751\times 10^{-2}$) outperforms Deep BSDE ($8.146\times 10^{-2}$); however, Deep BSDE converges to its minimum in just one iteration.
-These results suggest that, when the number of assets is small, the “curse of dimensionality” associated with increasing $k$ is largely mitigated by neural network–based approaches.
-Nevertheless, their relative performance can shift.
-Both the PMP and BSDE formulations appear robust to increases in factor dimensionality when $n$ is fixed and small.
-In this case, Monte Carlo sampling error and function approximation quality become the primary computational bottlenecks, rather than issues related to ill-conditioned linear algebra.
-Second, Table 4 reports results where the number of state factors is fixed at $k=2$, and the number of assets $n$ is varied ($n=1,10,50$).
-In stark contrast to the previous case, increasing $n$ leads to a substantial deterioration in performance across all three methods.
-Baseline PG-DPO shows the most pronounced increase in RMSE, rising from $4.770\times 10^{-2}$ for $n=1$ to $3.875\times 10^{-1}$ for $n=50$.
-While P-PGDPO still achieves the lowest RMSE in all cases, its error also increases with $n$—from $1.990\times 10^{-4}$ to $7.330\times 10^{-2}$.
-Deep BSDE’s performance degrades most dramatically, with RMSE rising from $2.675\times 10^{-3}$ at $n=1$ to $6.340\times 10^{-1}$ at $n=50$.
-Notably, in the $n=50,k=2$ case, Baseline PG-DPO ($3.875\times 10^{-1}$) outperforms Deep BSDE ($6.340\times 10^{-1}$).
-Table 3: Summary of Minimum Policy RMSE and Convergence Iterations for $n=1$ (Varying $k$)
-$k$ factors Method Min. Policy RMSE Iterations at Min.
-($n=1$ asset)
-1 Baseline PG-DPO $3.260\times 10^{-2}$ 1000
-P-PGDPO $1.120\times 10^{-3}$ 800
-Deep BSDE $9.350\times 10^{-3}$ 3600
-5 Baseline PG-DPO $3.714\times 10^{-2}$ 1400
-P-PGDPO $2.500\times 10^{-4}$ 800
-Deep BSDE $3.094\times 10^{-2}$ 9400
-10 Baseline PG-DPO $1.751\times 10^{-2}$ 2800
-P-PGDPO $4.710\times 10^{-4}$ 1200
-Deep BSDE $8.146\times 10^{-2}$ 1
-Table 4: Summary of Minimum Policy RMSE and Convergence Iterations for $k=2$ (Varying $n$)
-$n$ assets Method Min. Policy RMSE Iterations at Min.
-($k=2$ factors)
-1 Baseline PG-DPO $4.770\times 10^{-2}$ 1200
-P-PGDPO $1.990\times 10^{-4}$ 800
-Deep BSDE $2.675\times 10^{-3}$ 3600
-10 Baseline PG-DPO $1.145\times 10^{-1}$ 4000
-P-PGDPO $3.645\times 10^{-3}$ 4600
-Deep BSDE $3.835\times 10^{-2}$ 1
-50 Baseline PG-DPO $3.875\times 10^{-1}$ 5600
-P-PGDPO $7.330\times 10^{-2}$ 9800
-Deep BSDE $6.340\times 10^{-1}$ 3800
-This pronounced degradation with increasing $n$, even for small $k$, points to a distinct source of difficulty—one not solely attributable to function approximation in high-dimensional state spaces.
-A likely culprit is the numerical instability introduced by the asset covariance matrix $\Sigma$.
-As $n$ increases, the $n\times n$ matrix $\Sigma$ becomes more prone to ill-conditioning (i.e., a high condition number).
-Since optimal portfolio policies in the PMP framework (Eq.
-(11)) involve computing $\Sigma^{-1}$, small estimation errors in the costates or model parameters can be significantly amplified, resulting in large policy errors.
-This observation suggests that the “curse” associated with large $n$ is not merely about learning a higher-dimensional function, but also reflects the increasing fragility of the optimization problem itself.
-The structure of the problem becomes numerically unstable, which adversely affects any method that explicitly or implicitly depends on $\Sigma$ or its inverse.
-These findings highlight that while neural network–based methods have made substantial progress in managing high-dimensional state spaces (associated with large $k$), challenges related to a large number of decision variables (i.e., assets $n$)—particularly in the presence of ill-conditioned covariance matrices—remain a significant concern.
-Addressing this “asset-dimension curse” may require incorporating techniques such as covariance matrix shrinkage, spectral regularization, or other forms of preconditioning within the optimization problem.
-4.2 High-Dimensional Performance Analysis ($n=50,k=10$)
-We now visually examine the performance of the three methodologies in the most challenging setting considered: $n=50$ assets and $k=10$ state factors.
-This high-dimensional configuration, where traditional dynamic programming methods are computationally intractable, highlights the distinct capabilities and limitations of each approach.
-As a reference, Figure 2 displays the benchmark analytical solution for the first asset ($\pi_{1}^{\text{sol}}$), decomposed into its myopic ($\pi_{1,0}^{\text{sol}}$) and intertemporal hedging ($\pi_{1,1}^{\text{sol}}$) components.
-The hedging component is clearly non-trivial, exhibiting complex dependencies on both the state variable $Y_{1}$ and the remaining time to maturity.
-This underscores its importance for achieving optimal portfolio allocation in dynamic settings.
-(a) Benchmark Policy ($\pi_{1}^{\text{sol}}$)
-(b) Myopic Comp. ($\pi_{1,0}^{\text{sol}}$)
-(c) Hedging Comp. ($\pi_{1,1}^{\text{sol}}$)
-Figure 2: Benchmark (analytical) solution components for the first asset ($i=1$) in the $n=50,k=10$ case, plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (e.g., $t=(T-T_{0})/2$ if $T_{0}$ is initial time, assuming $T=1.5$ is the horizon from $t=0$; thus, for instance, at $t=0.75$ or time-to-maturity $0.75$).
-Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
-From left to right: total optimal policy, its myopic component, and its intertemporal hedging component.
-These serve as the ground truth.
-Figure 3 presents a side-by-side comparison of the policies learned by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and Deep BSDE, along with their corresponding errors, for the first asset.
-These plots correspond to the iterations at which each method achieved its (near) optimal performance in the $n=50,k=10$ setting, as reported in Table 2.
-For instance, the policy shown for P-PGDPO corresponds to iteration 6400, which is the nearest logged evaluation point to the observed minimum RMSE at iteration 5399 (evaluation was performed every 200 iterations, and the global minimum RMSE for P-PGDPO in this setting was ultimately achieved at iteration 10000, as also noted in Table 2).
-(a) Baseline Policy ($\pi_{1,\theta}$)
-(b) Baseline Error
-(c) P-PGDPO Policy ($\pi_{1}^{\mathrm{PMP}}$)
-(d) P-PGDPO Error
-(e) Deep BSDE Policy ($\pi_{1}^{\mathrm{BSDE}}$)
-(f) Deep BSDE Error
-Figure 3: Comparison of total policies (left column) and policy errors (right column) for the first asset ($i=1$) across Baseline PG-DPO (top row), Projected PG-DPO (P-PGDPO) (middle row), and Deep BSDE (bottom row) in the $n=50,k=10$ case.
-Results shown at (near) optimal iterations for each method (e.g., iteration 6400 for P-PGDPO as plotted), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2).
-Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
-Errors use a symmetric logarithmic (symlog) scale.
-Different color bar ranges reflect varying policy/error magnitudes.
-The Baseline PG-DPO method (Figure 3a,b), employing a standard neural network architecture (see Appendix D.2), performs poorly in this high-dimensional setting.
-Although the learned policy exhibits some non-linear structure, it bears little resemblance to the benchmark solution, and the corresponding error is substantial.
-This outcome visually corroborates the relatively high RMSE of $1.070\times 10^{-1}$ reported for Baseline PG-DPO in the $n=50,k=10$ case (Table 2), and underscores the difficulty of directly approximating a high-dimensional optimal policy function using generic neural architectures without incorporating structural insights from the problem’s analytic formulation.
-(a) P-PGDPO Myopic ($\pi_{1,0}^{\mathrm{PMP}}$)
-(b) P-PGDPO Hedging ($\pi_{1,1}^{\mathrm{PMP}}$)
-(c) Myopic Error
-(d) Hedging Error
-Figure 4: Decomposition of the Projected PG-DPO (P-PGDPO) policy for the first asset ($i=1$) in the $n=50,k=10$ case (results from iteration 6400 as plotted), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2).
-Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
-Top row: myopic and hedging components.
-Bottom row: their respective errors (symmetric logarithmic (symlog) scale).
-Note different color bar ranges.
-The Deep BSDE method (Figure 3e,f) shows a marked improvement over Baseline PG-DPO.
-Its learned policy qualitatively resembles the myopic component of the benchmark (cf.
-Figure 2b), primarily capturing the dependency on the state variable $Y_{1}$.
-However, the policy fails to capture the more complex intertemporal hedging demands.
-The corresponding error plot (Figure 3f) reveals significant discrepancies in regions where hedging plays a critical role, suggesting that the hedging component is largely absent from the Deep BSDE solution.
-This shortcoming can be attributed to both architectural and training limitations.
-The Deep BSDE framework is primarily designed to match the terminal condition and ensure consistency of the forward value process $Y_{t}$ (approximating the value function) and the backward process $Z_{t}$ (approximating its gradient).
-However, intertemporal hedging depends crucially on the mixed second-order derivative $V_{x\mathbf{y}}$, which is neither explicitly represented nor directly supervised in the standard two-network Deep BSDE setup.
-Moreover, the loss function penalizes deviations in $Z_{t}$ through a quadratic martingale residual, which may assign relatively low weight to errors in specific gradient directions, particularly those influencing $V_{x\mathbf{y}}$.
-These small gradient estimation errors can nonetheless be magnified when constructing the control policy, especially through multiplication by the potentially ill-conditioned inverse covariance matrix $\Sigma^{-1}$ in high-dimensional settings.
-As a result, while the Deep BSDE method effectively learns the dominant myopic component of the policy, it tends to underperform in capturing the subtler hedging demands essential for dynamic portfolio optimization.
-In stark contrast, the Projected PG-DPO (P-PGDPO) method (Figure 3c,d) produces a policy that is visually nearly indistinguishable from the benchmark total policy (Figure 2a).
-The corresponding error is substantially smaller than that of the other two methods, consistent with its superior RMSE of $1.522\times 10^{-2}$ for the $n=50,k=10$ case (Table 2).
-This performance underscores the effectiveness of P-PGDPO’s decoupled learning structure: first estimating costates via BPTT and stabilizing them through Monte Carlo averaging, then analytically constructing the policy using the PMP first-order conditions.
-The exceptional accuracy of P-PGDPO is further illuminated by decomposing its policy into myopic and hedging components, as shown in Figure 4 (from iteration 6400).
-The myopic component (Figure 4(a)) closely matches the benchmark, with its error (Figure 4(c)) being negligible—on the order of $O(10^{-7})$.
-The hedging component (Figure 4(b)) also successfully captures the complex shape of the benchmark hedging demand (Figure 2c).
-The residual error (Figure 4(d)) in this component dominates the total policy error.
-Notably, this hedging error likely does not stem from a flaw in the P-PGDPO algorithm itself, but rather reflects the numerical challenges inherent to high-dimensional asset allocation.
-Specifically, the hedging term in the PMP solution involves the inverse of the asset return covariance matrix $\Sigma$:
-$\pi^{\text{hedge}}=-\frac{1}{X_{t}V_{xx}}\,\underbrace{\Sigma^{-1}}_{\text{canbe ill-conditioned for large }n}(\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y})\,V_{x\mathbf{Y}}.$
-For $n=50$, the $50\times 50$ matrix $\Sigma$ is prone to ill-conditioning.
-Thus, even small residual noise in the BPTT-estimated $V_{x\mathbf{Y}}$ may be significantly amplified by $\Sigma^{-1}$, especially when its condition number is large.
-This amplification effect is the most likely cause of the observed hedging error.
-In conclusion, the $n=50,k=10$ visualizations clearly demonstrate that Projected PG-DPO is capable of “breaking the dimensional barrier,” where Baseline PG-DPO fails and Deep BSDE provides only a partial (myopic-focused) solution.
-The remaining inaccuracies in the hedging component point to a pressing need for addressing the ill-conditioning of $\Sigma$ in very high asset dimensions—potentially through regularization, shrinkage techniques, or architecture-aware preconditioning.
-4.3 Robustness and Accuracy in Long-Horizon Portfolio Choice
-While the results in the preceding sections demonstrate the effectiveness of the PG-DPO framework for a relatively short $T=1.5$ year horizon, a crucial test of any dynamic asset allocation method is its robustness and accuracy over long-term investment periods.
-Long-horizon optimization poses a significant challenge, as the baseline PG-DPO algorithm is prone to the compounding effects of variance and bias accumulation over extended simulations.
-The error sources identified in Theorem 1, namely discretization ($\propto\Delta t$) and Monte Carlo sampling ($\propto 1/\sqrt{M}$), become critically important as the simulation horizon lengthens.
-To build a truly robust framework capable of tackling these long-term challenges, we introduce three key algorithmic enhancements designed to fundamentally improve accuracy and stability:
-  1. 1.
-    Residual Learning for Hedging Demand: We restructure the policy network to exclusively learn the complex intertemporal hedging component, while the analytically known myopic demand is computed separately (see (Silver et al., , 2018; Johannink et al., , 2019)).
-This separation is a natural use of domain knowledge in financial economics, where the myopic component is well-understood, and the accurate estimation of the hedging demand is the primary challenge.
-This residual learning approach allows the network to dedicate its full capacity to the most difficult aspect of the optimization problem.
-  2. 2.
-    Control Variates: To suppress the variance inherent in long Monte Carlo rollouts, which is the most critical issue in long-term simulations, we introduce a control variate.
-A key advantage of our model-based approach is the ability to seamlessly integrate such classical variance reduction techniques from the financial engineering literature (Glasserman, , 2004).
-By using the analytically tractable myopic policy as a control variate, we can significantly stabilize the learning process.
-  3. 3.
-    Richardson Extrapolation: To aggressively reduce time discretization error, we employ Richardson extrapolation.
-The known structure of the underlying stochastic differential equations in our model-based setting permits the straightforward application of established numerical methods like this one.
-By combining simulations at time steps $\Delta t$ and $0.5\Delta t$, we achieve a higher order of accuracy, crucial for maintaining fidelity over long time periods (Glasserman, , 2004).
-Before deploying these techniques on a long-horizon problem, we first validated their impact on our original $T=1.5$ year experimental setup.
-This step ensures that the enhancements provide a fundamental improvement in performance.
-As shown in Table 5, the new techniques yield a dramatic reduction in policy RMSE for both the Baseline PG-DPO and the P-PGDPO algorithms.
-Table 5: RMSE Improvement with Advanced Techniques for $T=1.5$ Horizon
-n k New Baseline RMSE Baseline Improvement New P-PGDPO RMSE P-PGDPO Improvement
-1 1 $5.10\times 10^{-3}$ 6.4×6.4$\times$6.4 × $4.80\times 10^{-5}$ 23.3×23.3$\times$23.3 ×
-1 2 $8.29\times 10^{-4}$ 57.5×57.5$\times$57.5 × $6.30\times 10^{-5}$ 3.16×3.16$\times$3.16 ×
-1 5 $1.04\times 10^{-3}$ 35.7×35.7$\times$35.7 × $1.00\times 10^{-5}$ 25.0×25.0$\times$25.0 ×
-1 10 $5.54\times 10^{-4}$ 31.6×31.6$\times$31.6 × $2.23\times 10^{-4}$ 2.11×2.$1$
-10 2 $1.49\times 10^{-2}$ $\times7$ $8.33\times 10^{-4}$ 4.38×4.38$\times$4.38 ×
-10 10 $5.07\times 10^{-3}$ $\times9$ $2.21\times 10^{-3}$ 1.27×1.27$\times$1.27 ×
-50 2 $8.07\times 10^{-2}$ 4.8×4.8$\times$4.8 × $8.40\times 10^{-3}$ 8.73×8.73$\times$8.73 ×
-50 10 $3.38\times 10^{-2}$ 3.2×3.2$\times$3.2 × $1.07\times 10^{-2}$ 1.42×1.42$\times$1.42 ×
-Note: This table compares the minimum policy RMSE of the enhanced algorithms against the original results from Tables 3 and 4.
-The ‘Improvement’ column indicates the factor by which the RMSE was reduced (e.g., Original RMSE / New RMSE).
-The significant performance gains, with improvement factors of up to 57.5×57.5$\times$57.5 ×, confirm that these techniques are not merely incremental adjustments but constitute a fundamental upgrade to the algorithm’s core.
-This successful validation provides a strong foundation for tackling the more demanding long-horizon problem.
-Armed with a demonstrably more robust algorithm, we now address a long-horizon portfolio choice problem with a maturity of $T=20$ years.
-As the investment horizon extends towards infinity, the optimal policy is expected to become stationary, i.e., independent of the time-to-maturity.
-While this stationary policy is a simpler functional form for the network to learn, the extended simulation required to find it amplifies the challenges of variance and error control.
-Our enhanced framework is designed precisely for this scenario.
-Table 6 presents the final policy RMSE for the 20-year experiment.
-The results confirm the success of our approach, showing remarkably low errors even in high dimensions.
-In the $n=50,k=10$ case, the P-PGDPO method achieves an error of just $3.85\times 10^{-3}$.
-This result is even superior to that of the 1.5-year experiment, a phenomenon explained by the synergy of a more powerful algorithm converging to a simpler, stationary target policy.
-Table 6: Minimum Policy RMSE for Long‑Horizon Problem ($T=20$)
-Baseline PG-DPO P-PGDPO
-$n$ $k$ Min. Policy RMSE Iterations at Min. Min. Policy RMSE Iterations at Min.
-1 1 0.00169 400 0.000041 400
-5 3 0.00876 200 0.000352 1
-30 5 0.07220 1000 0.007030 200
-50 10 0.05020 2600 0.003850 200
-Figure 5 provides a compelling visual proof of the framework’s robustness.
-An important characteristic of the long-horizon benchmark solution is that the intertemporal hedging demand exhibits strong time-dependency for short maturities (e.g., within 1.5 years), but becomes effectively time-invariant for longer maturities, converging to a stationary policy.
-Furthermore, creating a reliable interpolation for the 20-year benchmark is challenging, as the optimal policy changes rapidly for short maturities where evaluation points are sparse.
-Therefore, the figure presents scatter plots of the raw policy function outputs to ensure a faithful visualization of the results.
-The policies produced by P-PGDPO are visually indistinguishable from the benchmark solution, perfectly capturing the intricate, non-linear dependencies on the state variables.
-This stands in stark contrast to the Baseline PG-DPO, which fails to learn the correct policy structure.
-The comparison makes clear the qualitative difference between a rough approximation, which may achieve a low but misleading error metric, and a truly accurate approximation that captures both the value and the complex structure of the optimal policy.
-These results underscore the power of our enhanced P-PGDPO framework to deliver robust, high-fidelity solutions for challenging long-horizon control problems.
-(a) Baseline PG-DPO (Asset 1)
-(b) Baseline PG-DPO (Asset 2)
-(c) P-PGDPO (Asset 1)
-(d) P-PGDPO (Asset 2)
-(e) Benchmark Solution (Asset 1)
-(f) Benchmark Solution (Asset 2)
-Figure 5: Comparison of learned policies for the 20-year horizon ($n=50,k=10$), plotted against the first state factor $Y_{1}$.
-Left column: Asset 1.
-Right column: Asset 2.
-The P-PGDPO results show a near-perfect match with the benchmark solution, whereas the Baseline PG-DPO fails to capture the correct policy structure.
-4.4 Non-affine Asset Return Dynamics
-While the analysis in Section 4.3 demonstrated the framework’s robustness over long horizons, we now pivot to assess its flexibility with non-affine model dynamics.
-This investigation is a direct extension of our initial experiments on affine models (Sections 4.1 and 4.2).
-To ensure a clear and direct comparison, we revert to the T=1.5 year horizon and the foundational P-PGDPO algorithm.
-This deliberate choice allows us to attribute any changes in the optimal policy solely to the introduction of non-linearity, isolating its effects from the advanced techniques introduced in the long-horizon study.
-We assume that the state factors, $\mathbf{Y}_{t}=[Y_{1,t},\dots,Y_{k,t}]^{T}$, follow the same multivariate Ornstein-Uhlenbeck (OU) process as described in Section 2.1.
-The expected returns of risky assets, $\bm{\mu}_{S}(t,\mathbf{Y}_{t})$, are given by
-$\bm{\mu}_{S}(t,\mathbf{Y}_{t})=\text{diag}(\bm{\sigma})(\bm{\alpha}\mathbf{Y}_{\beta,t})+r\mathbf{1}=\text{diag}(\bm{\sigma})\left(\bm{\alpha}(\mathbf{Y}_{t}+\bm{\beta}\mathbf{Y}_{t}^{2})\right)+r\mathbf{1},$
-where $\bm{\alpha}$ is a $d\times k$ matrix of factor loadings and the multiplication and addition in $\bm{\alpha}(\mathbf{Y}_{t}+\bm{\beta}\mathbf{Y}_{t}^{2})$ is understood as element-wise operations.
-The vector $\bm{\beta}$ controls the strength and direction of the non-linear impact of $\mathbf{Y}_{t}$ on asset returns.
-Its overall magnitude is parameterized by $\beta_{\text{norm}}=||\bm{\beta}||_{2}$, and for our experiments, $\bm{\beta}$ is generated as $\bm{\beta}=\beta_{\text{norm}}\cdot\mathbf{u}$, where $\mathbf{u}$ is a randomly generated unit vector (with a fixed seed for consistency).
-When $\beta_{\text{norm}}=0$, $\bm{\beta}=\mathbf{0}$, which implies $\mathbf{Y}_{\beta,t}=\mathbf{Y}_{t}$.
-In this case, the model for $\bm{\mu}_{S}$ becomes affine in $\mathbf{Y}_{t}$, and the overall portfolio optimization problem reduces to the affine setting analyzed in Sections 4.1 and 4.2.
-The primary objective of this analysis is twofold: first, to validate the solutions obtained by our Projected PG-DPO (P-PGDPO) framework for this model by examining their behavior as $\beta_{\text{norm}}\to 0$, where we expect convergence to the known affine solution.
-Second, we investigate how the optimal policy systematically changes as the strength of this non-linear component in the risk premium, dictated by $\beta_{\text{norm}}$, increases.
-All experiments in this subsection are conducted for the most challenging setting, $n=50$ assets and $k=10$ state factors, utilizing the P-PGDPO method.
-Table 7: Policy RMSE of P-PGDPO for the model with non-affine risk premia vs. affine benchmark ($n=50,k=10$). Results reported after 10,000 training iterations.
-$\beta_{\text{norm}}$ Min. Policy RMSE (P-PGDPO)
-0.0 0.04701
-0.1 0.05686
-0.5 0.06755
-1.0 0.11469
-2.0 0.20120
-3.0 0.27530
-4.0 0.33570
-(a) Analytical solution ($\beta_{\text{norm}}=0$).
-(b) P-PGDPO policy ($\beta_{\text{norm}}=0.0$).
-(c) P-PGDPO policy ($\beta_{\text{norm}}=0.1$).
-(d) P-PGDPO policy ($\beta_{\text{norm}}=1.0$).
-(e) P-PGDPO policy ($\beta_{\text{norm}}=2.0$).
-Figure 6: Visual comparison of optimal policies for the first asset ($\pi_{1}$) against the first state factor ($Y_{1}$) and time-to-maturity ($T-t$) in the $n=50,k=10$ setting.
-(a) shows the analytical solution for the affine model ($\beta_{\text{norm}}=0$).
-(b)-(e) display P-PGDPO learned policies for models with increasing strength of non-linearity in asset risk premia, corresponding to $\beta_{\text{norm}}$ values of $0.0,0.1,1.0,2.0$, respectively.
-Note the systematic change in policy structure and value ranges (indicated by colorbars) as $\beta_{\text{norm}}$ increases.
-To quantify the deviation of the P-PGDPO policy in this non-affine setting from the affine benchmark solution as $\beta_{\text{norm}}$ changes, Table 7 presents the Root Mean Squared Error (RMSE) of the P-PGDPO learned policy, $\bm{\pi}(\beta_{\text{norm}})$, relative to the affine analytical solution ($\bm{\pi}(\beta_{\text{norm}}=0)$).
-For the results presented in this table, all P-PGDPO models were trained for 10,000 iterations.
-In the non-affine case ($\beta_{\text{norm}}>0$), where an analytical solution to the control problem is unavailable, this fixed number of iterations is used for evaluation, as opposed to selecting the number of iterations based on the minimum error against a known benchmark in earlier sections.
-The results in Table 7 clearly demonstrate two key points.
-First, for $\beta_{\text{norm}}=0.0$, the P-PGDPO method achieves a low RMSE of $0.04701$.
-This is consistent with its strong performance in the affine setting (reported in Section 4.1) and indicates an accurate approximation of the affine analytical solution, as expected when the non-linear component vanishes.
-Second, as $\beta_{\text{norm}}$ increases from $0.1$ to $4.0$, the difference of the P-PGDPO policy from the fixed affine benchmark, measured by RMSE, monotonically increases.
-This trend signifies that the optimal policy systematically diverges from the affine optimal policy as the magnitude of the non-linear component grows.
-The consistent increase suggests that P-PGDPO is stably adapting the policy to this increasing non-linearity in the market price of risk, rather than producing erratic or unstable solutions.
-To complement these quantitative findings, Figure 6 provides a visual comparison of the learned policies.
-Figure 6(a) displays the analytical solution for the affine model ($\beta_{\text{norm}}=0$).
-Figures 6(b)–6(e) show the P-PGDPO learned policies for models where $\beta_{\text{norm}}=0.0,0.1,1.0,$ and $2.0$, respectively.
-These plots depict the policy for the first asset ($\pi_{1}$) as a function of the first state factor ($Y_{1}$) and time-to-maturity ($T-t$), holding other state factors at their long-term means.
-As observed in Figure 6(b), the P-PGDPO policy for $\beta_{\text{norm}}=0.0$ closely matches the analytical affine solution shown in Figure 6(a), visually confirming the low RMSE reported in Table 7.
-As $\beta_{\text{norm}}$ increases to $0.1$ (Figure 6(c)), $1.0$ (Figure 6(d)), and $2.0$ (Figure 6(e)), a clear and systematic transformation of the policy surface is evident.
-The policy values tend to increase in certain regions (indicated by the shift towards warmer colors and the expanding range of the colorbars), and the overall shape of the policy function changes.
-This visual progression corroborates the quantitative results presented in Table 7, illustrating that the P-PGDPO framework captures the impact of the non-linear term in the risk premium in a consistent manner.
-Since the non-affine component $\bm{\beta}\,\mathbf{Y}_{t}^{2}$ directly alters how state variables $\mathbf{Y}_{t}$ influence the investment opportunity set (via $\bm{\mu}_{S}$), these changes in the overall policy are likely significantly driven by adjustments to the intertemporal hedging component.
-In summary, the combined quantitative and visual results from this analysis strongly suggest that the P-PGDPO method provides robust and consistent solutions for portfolio optimization problems where the asset prices follow general non-affine dynamics.
-The learned policies correctly converge to the affine solution as the specific non-linear component vanishes ($\beta_{\text{norm}}\to 0$), and they exhibit a systematic and stable response to increasing levels of this non-linearity.
-This provides confidence in applying PG-DPO to more complex financial models with non-affine dynamics where analytical solutions are typically unavailable.
-5 Conclusion
-This paper confronted the enduring “curse of dimensionality” in continuous-time portfolio optimization, a challenge that manifests distinctly with increasing asset numbers ($n$) versus state factors ($k$).
-Our numerical investigations (Section 4) revealed that while deep learning methods manage larger $k$ relatively well, escalating $n$ poses severe difficulties, particularly due to the ill-conditioning of asset covariance matrices.
-Furthermore, we identified that prevalent approaches like Deep BSDE, while proficient at capturing myopic policy components, often fail to adequately represent crucial intertemporal hedging demands, a limitation rooted in their network architecture and loss functions which do not explicitly target the necessary second-order mixed derivatives of the value function (e.g., $V_{x\mathbf{y}}$).
-To surmount these obstacles, we introduced the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, with its highly effective Projected PG-DPO (P-PGDPO) variant (Section 3.2).
-P-PGDPO distinctively leverages Pontryagin’s Maximum Principle (PMP) by first employing backpropagation-through-time (BPTT) to achieve rapidly stabilizing estimates of the Pontryagin costates and their derivatives (e.g., $\lambda_{t},\partial_{x}\lambda_{t},\partial_{\mathbf{Y}}\lambda_{t}$).
-These estimates are then analytically projected onto the manifold of PMP-optimal controls via the first-order conditions.
-This decoupling strategy proved remarkably successful, enabling P-PGDPO to tackle problems of previously prohibitive scales.
-For instance, in a challenging $n=50$ asset, $k=10$ factor environment, P-PGDPO achieved a total policy RMSE of $\mathbf{1.52\times 10^{-2}}$ relative to the analytical benchmark, with the myopic component error being negligible ($O(10^{-7})$) and, critically, accurately reconstructing complex hedging terms (Section 4.2, Figure 4).
-The theoretical underpinnings of P-PGDPO are strengthened by our Policy Gap Bound (Theorem 1), which guarantees near-optimal policies given accurate costate estimation and a small FOC gap.
-Despite its successes, P-PGDPO’s accuracy in representing hedging terms can be affected by the amplification of estimation noise through the inverse covariance matrix $\Sigma^{-1}$ when $n$ is large and $\Sigma$ is ill-conditioned (Section 4.2).
-Concurrently, our experiments confirmed that Deep BSDE, while adept at matching myopic demands, struggles with learning the nuanced cross-derivative information essential for hedging, largely due to its architectural and objective function design (Figure 3).
-These observations highlight the distinct strengths and weaknesses of different deep learning paradigms in high-dimensional control.
-The practical implications of P-PGDPO are significant.
-Its demonstrated ability to recover complex, non-linear hedging strategies in high dimensions suggests its potential applicability to large institutional portfolios, such as those involving numerous ETFs or derivative hedging.
-Moreover, the core principle of P-PGDPO—“estimate costates first, then derive policy”—holds promise for a broader class of problems, where PMP provides the necessary optimality structure.
-Indeed, the robustness of the P-PGDPO framework is further underscored in a companion paper, where we successfully apply it to large-scale portfolio optimization under portfolio constraints (Huh et al., , 2025).
-Future research will pursue several promising directions:
-  • Developing “P-PGDPO 2.0” by integrating $\Sigma$ regularization or shrinkage techniques to mitigate the impact of ill-conditioned covariance matrices on hedging term estimation.
-  • Extending the P-PGDPO framework to incorporate market frictions such as transaction costs and taxes, by adapting the PMP conditions accordingly.
-  • Investigating the performance of P-PGDPO with non-HARA utility functions and behavioral finance objectives, and fully integrating the optimization of intermediate consumption policies ($C_{t}$).
-  • Validating the framework through rigorous backtesting on historical and simulated market data.
-  • Rigorously characterizing the precise range of financial models that satisfy the sufficient regularity assumptions for the P-PGDPO framework’s effectiveness.
-In essence, PG-DPO, particularly its projected variant, demonstrates that the synergy between deep learning and the structural guidance of Pontryagin’s Maximum Principle can redefine the boundaries of solvability for high-dimensional continuous-time financial control problems.
-Acknowledgments
-This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-00562904).
-Appendix A Derivation of the Adjoint Diffusion Coefficient via Itô’s Lemma
+Appendix B Derivation of the Adjoint Diffusion Coefficient via Itô’s Lemma
+We derive the diffusion coefficients $\mathbf{Z}^{\lambda X}$ in Section 3.2. In vector form, the state process satisfies
+$d\mathbf{Y}_{t}=\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,d\mathbf{W}^{Y}_{t},$
+where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})\in\mathbb{R}^{k}$ is the drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\in\mathbb{R}^{k\times k}$ is a diagonal matrix of diffusion coefficients. i.e.,
+$\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})=\operatorname{diag}(\sigma_{Y,1}(t,\mathbf{Y}_{t}),\dots,\sigma_{Y,k}(t,\mathbf{Y}_{t})),$
+and $\mathbf{W}^{Y}_{t}\in\mathbb{R}^{k}$ is a vector of correlated Brownian motions.
+Moreover, let $\bm{\mu}(t,\mathbf{Y}_{t})\in\mathbb{R}^{n}$ denote the vector of expected returns, $\bm{\sigma}(t,\mathbf{Y}_{t})\in\mathbb{R}^{n\times n}$ the diagonal matrix of volatilities, and $\mathbf{W}^{X}_{t}\in\mathbb{R}^{n}$ the Brownian motion vector for risky assets.
+Then the dynamics of the vector of risky asset prices $\mathbf{S}_{t}\in\mathbb{R}^{n}$ can be compactly written as
+$d\mathbf{S}_{t}=\operatorname{diag}(\mathbf{S}_{t})\,\bm{\mu}(t,\mathbf{Y}_{t})\,dt+\operatorname{diag}(\mathbf{S}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}^{X}_{t}.$
 In the Pontryagin framework, the adjoint (or costate) process associated with the investor’s wealth $X_{t}$ is identified with the partial derivative of the value function $V$.
@@
 The second term captures the intertemporal hedging component arising from the correlation between asset returns and state variable movements, mediated by the cross-derivative of the value function $V_{X\mathbf{Y}}$ (or $\partial_{\mathbf{Y}}\lambda_{t}^{*}$).
-This derivation provides the explicit form of the adjoint diffusion coefficient $\mathbf{Z}_{t}^{\lambda X}$, which is fundamental for understanding the structure of the optimal portfolio policy (11) and for constructing the Pontryagin adjoint equations.
-Appendix B Analytic Solutions for Affine Models
-This appendix provides details on the derivation of analytic solutions for the portfolio optimization problems presented in Section 2.3 under specific assumptions (no intermediate consumption, CRRA utility), typically following the Hamilton-Jacobi-Bellman (HJB) approach with an exponential-quadratic value function ansatz.
-We also show how these solutions connect to the Pontryagin framework.
-B.1 Analytic Solution for the Multi-Asset OU Model (Section 2.3.1)
-For simplicity, we assume no intermediate consumption, i.e., $C_{t}=0$, and consider an investor who derives utility only from terminal wealth $X_{T}$ under a CRRA objective $U(x)=\frac{x^{1-\gamma}}{1-\gamma}$.
-Following Kim and Omberg, (1996), the associated Hamilton–Jacobi–Bellman (HJB) equation can be solved by positing a value function of the form:
-$V(t,x,Y)=\frac{x^{1-\gamma}}{1-\gamma}\exp\left(g(t)+B(t)Y+\frac{1}{2}C(t)Y^{2}\right).$ (20)
-Here, $g(t)$, $B(t)$, and $C(t)$ are deterministic functions of time determined by solving a system of ordinary differential equations (ODEs) derived from the HJB equation, subject to terminal conditions at $t=T$.
-From this value function, the costate $\lambda_{t}^{*}=V_{x}$, its derivative with respect to wealth $\partial_{x}\lambda_{t}^{*}=V_{xx}$, and its derivative with respect to the state $\partial_{Y}\lambda_{t}^{*}=V_{xY}$ have the following closed-form expressions on the optimal path $(X_{t}^{*},Y_{t})$:
-$\lambda_{t}^{*}(t,X_{t}^{*},Y_{t})$ $=(X_{t}^{*})^{-\gamma}\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right),$
-$\frac{\partial\lambda_{t}^{*}}{\partial x}(t,X_{t}^{*},Y_{t})$ $=-\gamma(X_{t}^{*})^{-\gamma-1}\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right),$
-$\frac{\partial\lambda_{t}^{*}}{\partial Y}(t,X_{t}^{*},Y_{t})$ $=(X_{t}^{*})^{-\gamma}\left(B(t)+C(t)Y_{t}\right)\exp\left(g(t)+B(t)Y_{t}+\frac{1}{2}C(t)Y_{t}^{2}\right).$
-These expressions lead to remarkably simple ratios that are central to the Pontryagin framework’s characterization of the optimal portfolio:
-$\frac{\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma},\qquad\frac{\partial_{Y}\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma}\left(B(t)+C(t)Y_{t}\right).$
-Substituting these ratios into the general Pontryagin first-order condition for the optimal portfolio $\bm{\pi}_{t}^{*}$ (which equates the marginal contribution to the Hamiltonian from investing in risky assets to zero, cf.
-Eq.
-(10) derived in Section 2.2) confirms the explicit analytic solution derived via the HJB approach:
-$\bm{\pi}^{*}(t,Y)=\frac{1}{\gamma}\,\Sigma^{-1}\Bigl{[}\underbrace{(\bm{\alpha}\,\bm{\sigma})\,Y}_{\text{Myopic}}+\underbrace{\sigma_{Y}\,(\bm{\sigma}\,\bm{\rho})\,\bigl{(}B(t)+C(t)\,Y\bigr{)}}_{\text{Hedging}}\Bigr{]},$
-where $\Sigma=\mathrm{diag}(\bm{\sigma})\,\Psi\,\mathrm{diag}(\bm{\sigma})$, $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})^{\top}$, $\bm{\rho}=(\rho_{1},\dots,\rho_{n})^{\top}$, $(\bm{\alpha}\,\bm{\sigma})$ is the vector $(\alpha_{1}\sigma_{1},\dots,\alpha_{n}\sigma_{n})^{\top}$, and $(\bm{\sigma}\,\bm{\rho})$ is the vector $(\sigma_{1}\rho_{1},\dots,\sigma_{n}\rho_{n})^{\top}$.
-The derivation confirms that the first term captures the myopic demand driven by the instantaneous risk premium $(\bm{\mu}(Y)-r\mathbf{1})=(\bm{\alpha}\,\bm{\sigma})Y$, while the second term represents the intertemporal hedging demand, proportional to $B(t)+C(t)Y$ which arises from the state-dependency of the investment opportunity set captured by $\partial_{Y}\lambda_{t}^{*}$.
-Substituting this $\bm{\pi}^{*}$ back into the HJB equation and matching coefficients in powers of $Y$ yields the following Riccati-type ODEs for $B(t)$ and $C(t)$, along with the ODE for $g(t)$.
-The system is solved backward from the terminal conditions:
-$g(T)=0,\quad B(T)=0,\quad C(T)=0.$ (21)
-The ODEs are:
-$-\frac{dC(t)}{dt}$ $=-2\,\kappa_{Y}\,C(t)+\sigma_{Y}^{2}\,C(t)^{2}+\frac{1-\gamma}{\gamma}\Bigl{[}\beta_{0}+2\,\sigma_{Y}\,C(t)\,\beta_{1}+\sigma_{Y}^{2}\,C(t)^{2}\,\beta_{2}\Bigr{]},$ (22)
-$-\frac{dB(t)}{dt}$ $=-\kappa_{Y}\,B(t)+\sigma_{Y}^{2}\,B(t)\,C(t)-\frac{1-\gamma}{\gamma}\Bigl{[}\sigma_{Y}^{2}\,B(t)\,C(t)\,\beta_{2}+\sigma_{Y}\,B(t)\,\beta_{1}\Bigr{]}+\kappa_{Y}\,\theta_{Y}\,C(t),$ (23)
-$\frac{dg(t)}{dt}=(1-\gamma)\,r-\frac{1-\gamma}{2\gamma}\Bigl{[}\beta_{0}+2\sigma_{Y}\,\beta_{1}\,B(t)+\sigma_{Y}^{2}\,\beta_{2}\bigl{(}B(t)^{2}+C(t)\bigr{)}\Bigr{]}+\kappa_{Y}\,\theta_{Y}\,B(t)+\tfrac{1}{2}\,\sigma_{Y}^{2}\,C(t).$ (24)
-Here, $r$ is the risk-free rate, $\kappa_{Y}$, $\theta_{Y}$, $\sigma_{Y}$ are the OU parameters, $\gamma$ is the risk aversion coefficient, and $\beta_{0}$, $\beta_{1}$, $\beta_{2}$ are quadratic forms involving the model parameters defined as:
-$\beta_{0}:=(\bm{\alpha}\,\bm{\sigma})^{\top}\,\Sigma^{-1}\,(\bm{\alpha}\,\bm{\sigma}),\quad\beta_{1}:=(\bm{\alpha}\,\bm{\sigma})^{\top}\,\Sigma^{-1}\,(\bm{\sigma}\,\bm{\rho}),\quad\beta_{2}:=(\bm{\sigma}\,\bm{\rho})^{\top}\,\Sigma^{-1}\,(\bm{\sigma}\,\bm{\rho}).$
-By solving the system of ODEs (22), (23), and (24) backward in time from the terminal conditions (21), one obtains the functions $g(t)$, $B(t)$, and $C(t)$.
-Substituting these into Eq.
-(20) gives the complete analytic form of the value function consistent with the underlying HJB/PMP framework.
-B.2 Analytic Solution for the Multi-Asset Multi-Factor Model (Section 2.3.2)
-[…]
-[…]
-where $\mathbf{B}(t)$ is an $k\times 1$ vector, $\mathbf{C}(t)$ is an $k\times k$ symmetric matrix (note: dimensions based on state vector $\mathbf{Y}\in\mathbb{R}^{k}$), and $g(t)$ is a scalar function, all satisfying ODEs derived from the HJB equation, with terminal conditions $g(T)=0$, $\mathbf{B}(T)=\mathbf{0}$ and $\mathbf{C}(T)=\mathbf{0}$.
-Define the relevant matrices and vectors:
-$\bm{\sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{n}),\;\bm{\sigma}_{Y}\;(\text{factor diffusion }k\times k),\;\Sigma=\bm{\sigma}\Psi\bm{\sigma}.$
-where $\Psi\in\mathbb{R}^{n\times n}$ is the correlation matrix for the $n$ risky assets, and let $\rho\in\mathbb{R}^{n\times k}$ be the cross-correlation matrix between these assets and the $k$-dimensional factor Brownian motion $\mathbf{W}^{Y}_{t}$.
-Also let $\mathbf{A}=\bigl{(}\bm{\alpha}_{1},\dots,\bm{\alpha}_{n}\bigr{)}^{\top}\in\mathbb{R}^{n\times k}$ collect the factor loadings $\bm{\alpha}_{i}\in\mathbb{R}^{k}$ for each asset $i$.
-The costate $\lambda_{t}^{*}=V_{x}$, its derivative with respect to wealth $\partial_{x}\lambda_{t}^{*}=V_{xx}$, and its gradient ($k\times 1$ vector) with respect to the state vector $\partial_{\mathbf{Y}}\lambda_{t}^{*}=V_{x\mathbf{Y}}$ are given by:
-$\lambda_{t}^{*}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=(X_{t}^{*})^{-\gamma}\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right),$
-$\frac{\partial\lambda_{t}^{*}}{\partial x}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=-\gamma(X_{t}^{*})^{-\gamma-1}\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right),$
-$\frac{\partial\lambda_{t}^{*}}{\partial\mathbf{Y}}(t,X_{t}^{*},\mathbf{Y}_{t})$ $=(X_{t}^{*})^{-\gamma}\left(\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}_{t}\right)\exp\left(g(t)+\mathbf{B}(t)^{\top}\mathbf{Y}_{t}+\tfrac{1}{2}\mathbf{Y}_{t}^{\top}\mathbf{C}(t)\mathbf{Y}_{t}\right).$
-These yield the following scalar and vector ratios, which simplify the PMP optimality conditions:
-$\frac{\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma},\qquad\frac{\partial_{\mathbf{Y}}\lambda_{t}^{*}}{\partial_{x}\lambda_{t}^{*}}=-\frac{X_{t}^{*}}{\gamma}\left(\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}_{t}\right).$
-Substituting these into the vector form of the PMP first-order condition for the optimal portfolio $\bm{\pi}_{t}^{*}\in\mathbb{R}^{n}$, (cf. Eq. (11) derived in Section 2.2):
-$\bm{\pi}_{t}^{*}=-\frac{1}{X_{t}^{*}\partial_{x}\lambda_{t}^{*}}\Sigma^{-1}\left\{\lambda_{t}^{*}(\bm{\mu}(\mathbf{Y}_{t})-r\mathbf{1})+\bm{\sigma}\rho\bm{\sigma}_{Y}\left[(\partial_{\mathbf{Y}}\lambda_{t}^{*})^{\top}\right]\right\},$
-where the risk premium vector is $\bm{\mu}(\mathbf{Y}_{t})-r\mathbf{1}=\bm{\sigma}\mathbf{A}\mathbf{Y}_{t}$, we precisely recover the analytic solution obtained from the HJB framework:
-$\bm{\pi}^{*}(t,\mathbf{Y})\;=\;\frac{1}{\gamma}\,\Sigma^{-1}\Bigl{[}\,\underbrace{\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}}_{\text{Myopic component (related to }\bm{\mu}-r\mathbf{1})}\;+\;\underbrace{\bm{\sigma}\,\rho\,\bm{\sigma}_{Y}\Bigl{(}\mathbf{B}(t)\;+\;\mathbf{C}(t)\,\mathbf{Y}\Bigr{)}}_{\text{Hedging component (related to }\partial_{\mathbf{Y}}\lambda_{t}^{*})}\Bigr{]}.$
-Here, the first bracketed term corresponds to the myopic component, driven by the instantaneous expected excess returns $\bm{\sigma}\,\mathbf{A}\,\mathbf{Y}$.
-The second bracketed term incorporates intertemporal hedging demands arising from the desire to hedge against unfavorable changes in the investment opportunity set, as captured by the factor sensitivities $\mathbf{B}(t)+\mathbf{C}(t)\mathbf{Y}$ (proportional to the gradient $\partial_{\mathbf{Y}}\lambda_{t}^{*}$) and the correlations encoded in $\rho$, $\Psi$, and $\Phi$ (implicitly via the ODEs for $\mathbf{B}$ and $\mathbf{C}$).
-To determine the time-dependent coefficients $\mathbf{B}(t)$ and $\mathbf{C}(t)$, we solve the backward matrix Riccati ODE system derived from substituting the value function ansatz and optimal portfolio into the HJB equation:
-$\frac{d\mathbf{C}(t)}{dt}$ $=\kappa_{Y}^{\top}\mathbf{C}(t)+\mathbf{C}(t)\kappa_{Y}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}\mathbf{C}(t)$
-$\quad-\frac{1-\gamma}{\gamma}\left[(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right.$
-$\qquad\qquad\left.+\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})+\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right],$ (26)
-$\frac{d\mathbf{B}(t)}{dt}$ $=\left[\kappa_{Y}^{\top}-\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}-\frac{1-\gamma}{\gamma}\left(\mathbf{C}(t)(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})+(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\rho\bm{\sigma}_{Y})\right)\right]\mathbf{B}(t)$
-$\quad-\kappa_{Y}\mathbf{C}(t)\bm{\theta}_{Y}.$ (27)
-subject to the terminal conditions $\mathbf{C}(T)=\mathbf{0}$ and $\mathbf{B}(T)=\mathbf{0}$.
-To obtain the complete value function (25), the scalar function $g(t)$ must also be determined by solving its own ODE backward from the terminal condition $g(T)=0$.
-This ODE, derived from the HJB equation by collecting terms independent of $\mathbf{Y}$ and consistent with the corrected single-factor case structure, is given by:
-$\frac{dg(t)}{dt}$ $=(1-\gamma)r-\frac{1-\gamma}{2\gamma}\operatorname{Tr}\left((\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right)-\frac{1-\gamma}{\gamma}\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})$
-$\quad-\frac{1-\gamma}{2\gamma}\left[\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma}\Sigma^{-1}\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{B}(t)+\operatorname{Tr}\left((\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma}\Sigma^{-1}\bm{\sigma}\rho\bm{\sigma}_{Y})\mathbf{C}(t)\right)\right]$
-$\quad+\mathbf{B}(t)^{\top}\kappa_{Y}\bm{\theta}_{Y}\quad+\frac{1}{2}\operatorname{Tr}(\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}).$
-Here, $\Phi=\mathrm{Cov}(d\mathbf{W}^{Y}_{t})/dt$ is the factor covariance matrix used in the factor dynamics $d\mathbf{Y}_{t}$, and $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix.
-Solving these coupled ODEs for $\mathbf{C}(t)$, $\mathbf{B}(t)$, and $g(t)$ backward in time from $T$ to $t$ and substituting the results into Eq.
-(25) gives the closed-form value function, while substituting $\mathbf{B}(t)$ and $\mathbf{C}(t)$ into $\bm{\pi}^{*}(t,\mathbf{Y})$ gives the optimal strategy, explicitly capturing both myopic and intertemporal hedging components in a multi-asset, multi-factor setting.
-Appendix C Complete and Rigorous Proofs for the Policy‑Gap Bound (Theorem 1)
-This appendix provides the complete, self‑contained analytic backbone for Theorem 1.
-All proof sketches have been expanded into detailed arguments, with explicit reliance on parabolic maximal‑regularity and monotone‑operator theory.
-C.1 Table of Symbols
+This derivation provides the explicit form of the adjoint diffusion coefficient $\mathbf{Z}_{t}^{\lambda X}$, which is fundamental for understanding the structure of the optimal portfolio policy (3.2) and for constructing the Pontryagin adjoint equations.
+Appendix C Proofs
+C.1 Proof of Lemma 1
+The recurrence relation for the pathwise costate $\lambda_{k}^{(i)}:=\partial J^{(i)}/\partial X_{k}^{(i)}$ is derived by applying the chain rule.
+The total reward $J^{(i)}$ is decomposed into the utility from the current step $k$ and the total future reward $J_{k+1}^{(i)}$.
+$\lambda_{k}^{(i)}=\frac{\partial}{\partial X_{k}^{(i)}}\left(e^{-\delta(t_{k}-t_{0})}U(C_{k}^{(i)})\Delta t\right)+\frac{\partial J_{k+1}^{(i)}}{\partial X_{k}^{(i)}}.$
+The second term becomes $\lambda_{k+1}^{(i)}(\partial X_{k+1}^{(i)}/\partial X_{k}^{(i)})$.
+Differentiating the discrete wealth dynamics, which includes the consumption term $C_{k}^{(i)}\Delta t$, with respect to $X_{k}^{(i)}$ yields:
+$\frac{\partial X_{k+1}^{(i)}}{\partial X_{k}^{(i)}}=1+\bigl{(}r_{k}+\bm{\pi}_{k}^{\top}(\bm{\mu}_{k}-r_{k}\mathbf{1})\bigr{)}\Delta t-\frac{\partial C_{k}^{(i)}}{\partial X_{k}^{(i)}}\Delta t+\bm{\pi}_{k}^{\top}\bm{\sigma}_{k}\Delta\mathbf{W}_{k}^{X,(i)}.$
+Substituting these components and rearranging the terms for $\lambda_{k}^{(i)}$ yields the expression in Eq. (1).
+C.2 Proof of Lemma 2
+The product inside the expectation is an inner product of vectors.
+The term involving the adapted part, $\mathbb{E}[\lambda_{k+1}(\dots)|\mathcal{F}_{t_{k}}]$, is zero as it is the expectation of a martingale increment.
+The non-zero terms arise from the quadratic covariation of the raw martingale components of $\lambda_{k+1}^{(i)}$ with the wealth shock.
+Using the standard properties $\mathbb{E}[(\Delta\mathbf{W}_{k}^{X})(\Delta\mathbf{W}_{k}^{X})^{\top}|\mathcal{F}_{t_{k}}]=I\Delta t$ and $\mathbb{E}[(\Delta\mathbf{W}_{k}^{Y})(\Delta\mathbf{W}_{k}^{X})^{\top}|\mathcal{F}_{t_{k}}]=\bm{\rho}_{Y\!X}\Delta t$, where $\bm{\rho}_{Y\!X}$ is the cross-correlation matrix, yields the stated result.
+C.3 Proof of Theorem 3
+We begin by taking the conditional expectation $\mathbb{E}[\cdot\mid\mathcal{F}_{t_{k}}]$ of the entire pathwise recurrence in Eq.
+(1) from Lemma 1.
+The left-hand side becomes $\lambda_{k}$ by definition.
+On the right-hand side, the expectations of the drift terms are straightforward.
+The crucial step is handling the stochastic term $\mathbb{E}[\lambda_{k+1}^{(i)}(\dots)|\mathcal{F}_{t_{k}}]$, for which we substitute the result from Lemma 2.
+This transforms the stochastic term into the additional drift component involving the effective process $\mathbf{Z}^{\lambda X}_{k}$.
+After collecting all drift terms and separating the martingale components, we arrive at the discrete-time BSDE presented in Eq.
+(3).
+The final step involves a direct algebraic comparison of this drift term with the definition of the Hamiltonian $\mathcal{H}$ from Section 3.2, which confirms their identity.
+C.4 Proof of Theorem 4.2
+This appendix provides a complete, self-contained analytic backbone for Theorem 4.2.
+All proof sketches are expanded into detailed arguments, relying on parabolic maximal regularity (MR) with VMO (vanishing mean oscillation) coefficients and the strong monotonicity of the PMP first-order condition (FOC) map.
+The reflected SDE framework is used purely as a localization device; passing to expanding domains recovers the original non-reflected problem (Karatzas and Shreve, , 1991; Lieberman, , 1996).
+C.4.1 Table of Symbols
 Symbol Meaning
@@
 $\varepsilon$ FOC residual norm
-$\delta_{\mathrm{BPTT}}$ Costate estimation error (Def. 1)
-$\underline{\Lambda}$ Lower eigenvalue from Lemma 4
-$C_{S},C_{v}^{\prime\prime}$ PDE stability / coupled‑error constants
-$L_{F},C_{\text{tot}}$ Final error amplification constants
-C.2 Preliminaries and Notation
-Throughout, we set the primary costate
-$\lambda:=\nabla_{x}V^{\pi,c}.$
-For any admissible control $(\pi,c)$ the sub‑value function $V^{\pi,c}$ solves
-$-\partial_{t}V^{\pi,c}=\mathcal{L}_{\pi,c}[V^{\pi,c}]+f_{\pi,c},\qquad V^{\pi,c}(T,\cdot)=g,$
-with oblique‑derivative boundary $\partial_{\mathbf{n}}V=0$ on $\partial\Omega$ to match the Skorokhod reflection (Lieberman, , 1996). Error quantities on $Q_{T}:=(0,T)\times\Omega$:
-$h$ $:=(a_{\pi,c}-a_{*},\,b_{\pi,c}-b_{*}),$ $\quad E_{h}$ $:=\|h\|_{L^{q,p}},$
-$E_{f}$ $:=\bigl{\|}e^{-\delta t}(U(c^{*})-U(c))\bigr{\|}_{L^{q,p}},$ $\quad E_{v}$ $:=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}},$
-$\varepsilon$ $:=\|\mathcal{R}_{\mathrm{FOC}}\|_{L^{q,p}}.$
-C.3 Minimal Assumptions
-Assumption 6 (Baseline Well‑Posedness).
-Fix $T>0$, a $C^{1,1}$ domain $\Omega$, mixed‑norm exponents $p>d+2,\;q\geq 2$ and a compact cylinder
-$\mathcal{D}:=[0,T]\times[\underline{x},\bar{x}]\times\{\|y\|\leq R_{Y}\}.$
-There exist positive constants $\underline{\nu},\overline{\nu},L_{\mathrm{coef}},C_{\mathrm{MR}},\eta_{0}$ such that
+$\delta_{\mathrm{BPTT}}$ Costate estimation error (Def. 2)
+$\underline{\kappa}$ Strong monotonicity modulus of the FOC map (Lemma 4)
+$C_{S},C_{v}^{\prime\prime}$ PDE stability / coupled-error constants
+$L_{G},L_{F},C_{\text{tot}}$ FOC-map Lipschitz / final amplification constants
+$\tau_{\star}$ Contraction slab length (distinct from numerical step $\Delta t$)
+C.4.2 Preliminaries and Notation
+Throughout, set the primary costate $\lambda:=\partial_{x}V^{\pi,c}$. For any admissible control $(\pi,c)$ the sub-value function $V^{\pi,c}$ solves
+$\partial_{t}V^{\pi,c}\;=\;\mathcal{L}_{\pi,c}[V^{\pi,c}]+f_{\pi,c},\qquad V^{\pi,c}(T,\cdot)=g,$
+on $Q_{T}:=(0,T)\times\Omega$, with oblique-derivative boundary $\partial_{\mathbf{n}}V=0$ on $\partial\Omega$ matching a Skorokhod reflection used for localization; solutions on expanding domains converge to the non-reflected solution (Karatzas and Shreve, , 1991; Lieberman, , 1996).
+Define the error quantities on $Q_{T}$ by
+$h$ $=(a_{\pi,c}-a_{\ast},\,b_{\pi,c}-b_{\ast}),$ $\quad E_{h}$ $=\|h\|_{L^{q,p}},$
+$E_{f}$ $=\bigl{\|}e^{-\delta t}\bigl{(}U(c^{\ast})-U(c)\bigr{)}\bigr{\|}_{L^{q,p}},$ $\quad E_{v}$ $=\|\nabla V^{\pi,c}-\nabla V^{\ast}\|_{L^{q,p}},$
+$\varepsilon$ $=\|\mathcal{R}_{\mathrm{FOC}}\|_{L^{q,p}},$
+Convention on $Z_{\lambda X}$.
+We use
+$Z_{\lambda X}\;=\;(\partial_{x}\lambda)\,X\,\pi^{\top}\sigma\;+\;(\partial_{Y}\lambda)\,\sigma_{Y}\,\rho^{\top},\qquad\sigma\,Z_{\lambda X}^{\top}\;=\;X\bigl{(}-\partial_{x}\lambda\bigr{)}\,\Sigma\,\pi\;+\;\sigma\rho\sigma_{Y}^{\top}(\partial_{Y}\lambda)^{\top},$
+consistent with the portfolio FOC (cf. main text Eq. (11)); see Yong and Zhou, (2012) for PMP–HJB links.
+C.4.3 Minimal Assumptions
+Assumption 5 (Baseline well-posedness and regularity).
+Fix $T>0$, a bounded $C^{1,1}$ domain $\Omega$, mixed-norm exponents $p>d+2$ and $q\geq 2$, and a compact cylinder $\mathcal{D}=[0,T]\times[\underline{x},\bar{x}]\times\{\|y\|\leq R_{Y}\}$.
+There exist constants $\underline{\nu},\overline{\nu},L_{\mathrm{coef}},C_{\mathrm{MR}},\eta_{0}>0$ such that:
   1. 1.
-    Reflected SDE. The Skorokhod‑reflected SDE with inward normal $\mathbf{n}\in C^{1,\alpha}$ is well posed.
+    The Skorokhod-reflected SDE with inward normal $\mathbf{n}\in C^{1,\alpha}$ is well posed; expanding domains converge to the non-reflected solution (Karatzas and Shreve, , 1991).
   2. 2.
-    Uniform ellipticity & VMO coefficients. $\underline{\nu}|\xi|^{2}\!\leq\!\xi^{\top}a_{\pi,c}\xi\!\leq\!\overline{\nu}|\xi|^{2}$ and $(a_{\pi,c},b_{\pi,c})$ have VMO/BMO bounds.
+    Parabolicity:
+    • (U)
+      Uniform ellipticity: $\underline{\nu}|\xi|^{2}\leq\xi^{\top}a_{\pi,c}\xi\leq\overline{\nu}|\xi|^{2}$ and $(a_{\pi,c},b_{\pi,c})$ have VMO/BMO (bounded mean oscillation) bounds on $\mathcal{D}$ (Dong and Kim, , 2009);
+    • (H)
+      The $Y$-block is uniformly elliptic with VMO bounds while the $X$-block may degenerate; MR and gradient bounds below are applied on interior cylinders $Q_{r}\Subset\mathcal{D}$ (Lieberman, , 1996).
   3. 3.
-    VMO‑smallness. On every cylinder $Q_{r}\subset\mathcal{D}$ with $r\leq R_{0}$ the VMO modulus is $\leq\eta_{0}<\bar{\eta}(d,p,q)$ of Dong and Kim, (2009).
+    There exists $R_{0}>0$ such that $\sup_{0<r\leq R_{0}}\omega_{a}(r)\leq\eta_{0}<\bar{\eta}(d,p,q)$ as in Dong and Kim, (2009).
   4. 4.
-    Costate band & utility. $\underline{\lambda}\leq\lambda\leq\bar{\lambda}$ on $\mathcal{D}$ and $U(c)=c^{1-\gamma}/(1-\gamma)$, $\gamma>0,\gamma\neq 1$.
-C.4 Automatic Properties of the Model
-Lemma 4 (Strong concavity of the Hamiltonian).
-[…]
-($-\nabla^{2}_{(\pi,c)}H(t,x,y,\lambda,\eta,\pi,c)\;\succeq\;\underline{\Lambda}I_{n+1},$
-where $\underline{\Lambda}>0$ depends only on $(a,b,\gamma,\Pi_{\max},\underline{\lambda},\bar{\lambda})$.
+    Costate and curvature bands on $\mathcal{D}$: $\underline{\lambda}\leq\lambda\leq\bar{\lambda}$ and $\underline{\kappa}_{x}\leq-\partial_{x}\lambda\leq\bar{\kappa}_{x}$.
+  5. 5.
+    $U(c)=c^{1-\gamma}/(1-\gamma)$ with $\gamma>0$, $\gamma\neq 1$, and coefficients obey linear growth/boundedness.
+Remark 1 (Ellipticity and interior estimates).
+Regime (U) can be enforced by a local lower bound on $\|\pi\|$ on $\mathcal{D}$.
+Under (H), arguments apply on $Q_{r}\Subset\mathcal{D}$ with local constants; a finite cover yields global bounds (Lieberman, , 1996).
+C.4.4 Automatic Properties of the Model
+Lemma 4 (Strong monotonicity of the portfolio FOC map).
+Define
+$\mathcal{G}_{\pi}(\pi;c,\lambda,\partial\lambda)\;:=\;\lambda(\mu-r\mathbf{1})+\sigma\,Z_{\lambda X}^{\top}\;=\;X\bigl{(}-\partial_{x}\lambda\bigr{)}\,\Sigma\,\pi\;+\;\lambda(\mu-r\mathbf{1})\;+\;\sigma\rho\sigma_{Y}^{\top}(\partial_{Y}\lambda)^{\top}.$
+If $\Sigma\succ 0$ and Assumption 5 holds, then $\partial_{\pi}\mathcal{G}_{\pi}=X(-\partial_{x}\lambda)\,\Sigma\succ 0$. Hence for all $\pi_{1},\pi_{2}$,
+$\big{\langle}\mathcal{G}_{\pi}(\pi_{1})-\mathcal{G}_{\pi}(\pi_{2}),\,\pi_{1}-\pi_{2}\big{\rangle}\;\geq\;\underline{\kappa}\,\|\pi_{1}-\pi_{2}\|^{2},\qquad\underline{\kappa}:=\underline{x}\,\underline{\kappa}_{x}\,\lambda_{\min}(\Sigma)>0,$
+so $\mathcal{G}_{\pi}$ is strongly monotone and has a Lipschitz inverse (Browder, , 1965; Minty, , 1962).
+Lemma 5 (Maximal regularity for $V^{\pi,c}$).
+Under Assumption 5 and $\eta_{0}<\bar{\eta}(d,p,q)$, both $V^{\pi,c}$ and $V^{\ast}$ lie in $W^{2,1}_{q,p}(Q_{T})$ and
+$\|D^{2}V^{\pi,c}\|_{L^{q,p}}+\|\partial_{t}V^{\pi,c}\|_{L^{q,p}}\;\leq\;C_{\mathrm{MR}}\Bigl{(}\|f_{\pi,c}\|_{L^{q,p}}+\|g\|_{W^{2-2/p}_{q}}\Bigr{)},$
+with the same bound for $V^{\ast}$; under (H) the estimate holds on $Q_{r}\Subset\mathcal{D}$ (Dong and Kim, , 2009).
+C.4.5 Foundational Lemmas
+Lemma 6 (FOC reduction: coefficient gap).
+Under Assumption 5 and Lemma 4,
+$E_{h}\;\leq\;L_{G}\,E_{v}\;+\;C_{\mathrm{FOC}}\,\varepsilon.$
 Proof.
-HARA implies $U^{\prime\prime}(c)\leq-\kappa_{0}<0$ on the compact $c$‑interval, while the quadratic portfolio term contributes $-x^{2}\lambda\,\sigma^{\top}\Sigma\sigma\preceq-\kappa_{1}I_{n}$ since $x\in[\underline{x},\bar{x}]$ and $\Sigma\succ 0$.
-Summing gives the claim.
+By Lemma 4, $\pi\mapsto\mathcal{G}_{\pi}(\pi;\lambda,\partial\lambda)$ has a Lipschitz inverse (Browder, , 1965; Minty, , 1962).
+Thus $\|\pi-\pi^{\ast}\|\leq c_{1}\bigl{(}\|\nabla V^{\pi,c}-\nabla V^{\ast}\|+\|\mathcal{R}_{\mathrm{FOC}}\|\bigr{)}$.
+Lipschitz dependence of coefficients on controls yields the claim after absorbing constants.
 ∎
-Lemma 5 (Maximal regularity for $V^{\pi,c}$).
-Under Assumption 6 and the VMO‑smallness $\eta_{0}<\bar{\eta}(d,p,q)$, both $V^{\pi,c}$ and $V^{*}$ lie in $W^{2,1}_{q,p}(Q_{T})$ with
-$\|D^{2}V^{\pi,c}\|_{L^{q,p}}+\|\partial_{t}V^{\pi,c}\|_{L^{q,p}}\;\leq\;C_{\mathrm{MR}}\bigl{(}\|f_{\pi,c}\|_{L^{q,p}}+\|g\|_{W^{2-2/p}_{q}}\bigr{)}.$
-Proof.
-Apply Dong and Kim, (2009) to the non‑divergence PDE for $V^{\pi,c}$ using the VMO modulus bound, uniform ellipticity and bounded coefficients. The same argument holds for $V^{*}$. ∎
-C.5 Foundational Lemmata
-Lemma 6 (FOC‑Reduction: Coefficient Gap).
-Under Assumption 6 and Lemma 4,
-$E_{h}\;\leq\;L_{H}\,E_{v}\;+\;C_{\mathrm{FOC}}\,\varepsilon.$
-Proof.
-The strong concavity of the Hamiltonian (Lemma 4) implies that its gradient with respect to the control, $G(u;\lambda):=\nabla_{u}H(u,\lambda)$, is a strongly monotone operator in $u$ for a fixed state $\lambda$.
-By the Minty–Browder theorem (Minty, , 1962; Browder, , 1965), such an operator has a Lipschitz-continuous inverse.
-This allows us to bound the gap in controls by the gap in costates ($E_{v}$) and the FOC residual ($\varepsilon$), which after applying the control-to-coefficient Lipschitz bound, yields the stated inequality.
-∎
-Lemma 7 (Source‑Gap).
-Under Assumption 6,
+Lemma 7 (Source gap).
+Under Assumption 5 and HARA utility,
 $E_{f}\;\leq\;C_{f}^{\prime}E_{v}+C_{f}\,\varepsilon.$
 Proof.
-Compactness of the costate band and the properties of HARA utility imply that both $U$ and $(U^{\prime})^{-1}$ are Lipschitz continuous on the relevant compact intervals.
-The result follows from a standard mean-value argument on the first-order conditions for consumption.
-∎
+On the costate band, $U$ and $(U^{\prime})^{-1}$ are Lipschitz on the relevant ranges; apply a mean-value argument to the consumption FOC (see PMP/HJB expositions in Yong and Zhou, , 2012). ∎
 Lemma 8 (Gradient stability).
@@
 Proof.
-Let $W:=V^{\pi,c}-V^{*}$.
-The function $W$ solves a linear parabolic PDE whose source term is determined by the coefficient and source gaps ($h$ and $f$).
-By the Dong–Kim gradient estimate for PDEs with VMO coefficients, a core result of maximal regularity theory, the gradient of the solution is bounded by the norm of the source term, which gives $\|\nabla W\|_{L^{q,p}}\leq C_{S}(E_{h}+E_{f})$.
+Let $W:=V^{\pi,c}-V^{\ast}$.
+Then $W$ solves a linear parabolic PDE with source determined by $(h,f)$.
+Parabolic MR with VMO coefficients yields $\|\nabla W\|_{L^{q,p}}\leq C_{S}(\|h\|_{L^{q,p}}+\|f\|_{L^{q,p}})$ (Dong and Kim, , 2009).
 ∎
 Proposition 1 (Coupled error resolution).
-There exists $C_{v}^{\prime\prime}$ such that $E_{v}\leq C_{v}^{\prime\prime}\,\varepsilon$.
+There exists $C_{v}^{\prime\prime}>0$ such that $E_{v}\leq C_{v}^{\prime\prime}\,\varepsilon$.
 Proof.
-Combining the three preceding lemmas yields the recursive inequality $E_{v}\leq\alpha E_{v}+\beta\varepsilon$, where $\alpha=C_{S}(L_{H}+C_{f}^{\prime})$.
-For a sufficiently small time interval $\Delta t\leq\Delta_{\star}$, PDE theory ensures that the constant $\alpha<1$.
-This turns the mapping $T(E_{v})=\alpha E_{v}+\beta\varepsilon$ into a contraction.
-By the contraction mapping principle, this recursion converges.
-Iterating this argument over the $\lceil T/\Delta_{\star}\rceil$ time slabs that cover the full horizon $[0,T]$ yields a final bound for $E_{v}$ that depends only on $\varepsilon$.
+Combine the three lemmas to get $E_{v}\leq\alpha E_{v}+\beta\varepsilon$, $\alpha=C_{S}(L_{G}+C_{f}^{\prime})$.
+For some slab length $\tau_{\star}>0$ (set by MR constants), on any slab of length $\tau\leq\tau_{\star}$ one has $\alpha<1$, hence $E_{v}\leq\beta(1-\alpha)^{-1}\varepsilon$.
+Concatenate $\lceil T/\tau_{\star}\rceil$ slabs.
+This is an application of the contraction principle.
 ∎
-C.6 Final Policy‑Gap Theorem
+C.4.6 Final Policy-Gap Theorem
 Definition 2 (BPTT error).
-$\delta_{\mathrm{BPTT}}:=\|\widehat{\lambda}-\nabla V^{\pi,c}\|_{L^{q,p}}\leq\kappa_{1}\Delta t+\kappa_{2}/\sqrt{M},$ with constants depending on the value function regularity established in Lemma 5.
-Theorem 1 (Policy Gap Bound).
-Let $\pi^{*}$ be the true optimal policy and $\widehat{\pi}$ be the policy generated by the P-PGDPO algorithm using a time step $\Delta t$ and batch size $M$.
-Under the regularity conditions specified in Assumption 6 in the Appendix, the gap between the P-PGDPO policy and the true optimum is bounded as follows:
-$\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\text{tot}}\left(\varepsilon+\kappa_{1}\Delta t+\frac{\kappa_{2}}{\sqrt{M}}\right)$
-where:
-  • $\varepsilon$ is the $L^{q,p}$-norm of the Pontryagin FOC residual from the warm-up policy.
-  • $\kappa_{1},\kappa_{2}$ are positive constants governing the BPTT estimation error, which arises from time discretization (error proportional to $\Delta t$) and Monte Carlo sampling (error proportional to $1/\sqrt{M}$).
-  • $C_{\text{tot}}$ is a positive constant that depends on the model parameters but not on $\varepsilon$, $\Delta t$, or $M$.
+Let $\widehat{\lambda}$ estimate $\nabla V^{\pi,c}$ by BPTT. For time step $\Delta t$ and batch size $M$,
+$\delta_{\mathrm{BPTT}}:=\|\widehat{\lambda}-\nabla V^{\pi,c}\|_{L^{q,p}}\;\leq\;\begin{cases}\kappa_{\mathrm{EM}}\,\Delta t^{1/2}+\kappa_{2}/\sqrt{M},\\ \kappa_{\mathrm{Mil}}\,\Delta t\;+\kappa_{2}/\sqrt{M}\quad\text{(with first-order acceleration),}\end{cases}$
+in line with standard SDE error theory (Glasserman, , 2004).
+Theorem 4.2 (Policy-Gap Bound). Let $\pi^{\ast}$ be optimal and $\widehat{\pi}$ the P-PGDPO policy trained with time step $\Delta t$ and batch size $M$. Under Assumption 5,
+$\|\widehat{\pi}-\pi^{\ast}\|_{L^{q,p}}\;\leq\;C_{\text{tot}}\bigl{(}\varepsilon+\delta_{\mathrm{BPTT}}\bigr{)}.$
+Equivalently,
+$\|\widehat{\pi}-\pi^{\ast}\|_{L^{q,p}}\;\leq\;\begin{cases}C_{\text{tot}}\bigl{(}\varepsilon+\kappa_{\mathrm{EM}}\,\Delta t^{1/2}+\kappa_{2}/\sqrt{M}\bigr{)},\\[3.0pt] C_{\text{tot}}\bigl{(}\varepsilon+\kappa_{\mathrm{Mil}}\,\Delta t+\kappa_{2}/\sqrt{M}\bigr{)}.\end{cases}$
 Proof.
-The proof follows from the triangle inequality: $\|\widehat{\pi}-\pi^{*}\|\leq L_{F}\bigl{(}\|\widehat{\lambda}-\nabla V^{\pi,c}\|+\|\nabla V^{\pi,c}-\nabla V^{*}\|\bigr{)}\leq L_{F}(\delta_{\mathrm{BPTT}}+E_{v}).$ Substituting the bounds for $\delta_{\mathrm{BPTT}}$ from Definition 1 and for $E_{v}$ from Proposition 1 yields the final result.
-∎
-Appendix D Detailed Experimental Setup
-This appendix provides the detailed setup for the numerical experiments presented in Section 4.
-D.1 Parameter Values
+By the triangle inequality,
+$\|\widehat{\pi}-\pi^{\ast}\|\;\leq\;L_{F}\Bigl{(}\|\widehat{\lambda}-\nabla V^{\pi,c}\|+\|\nabla V^{\pi,c}-\nabla V^{\ast}\|\Bigr{)}\;\leq\;L_{F}\bigl{(}\delta_{\mathrm{BPTT}}+E_{v}\bigr{)},$
+and substitute Definition 2 and Proposition 1. ∎
+The preceding assumptions, lemmas, and theorem complete the core logical chain for the policy-gap bound.
+What follows is not part of the minimal existence–uniqueness–bound argument itself, but collects a number of complementary considerations and implementation notes that help ensure the assumptions are met in practice and the results can be applied robustly in computational or extended theoretical settings.
+These points are offered as guidance for verification, numerical implementation, and consistency across related parts of the manuscript.
+  1. 1.
+    VMO smallness (how to guarantee). Two routes:
+    1. (a)
+      Top-level assumption. State explicitly that the $x$-VMO modulus of leading coefficients satisfies $\sup_{0<r\leq R_{0}}\omega_{a}(r)\leq\eta_{0}<\bar{\eta}(d,p,q)$.
+    2. (b)
+      Verification from primitives.
+If $y\mapsto\Sigma(y)$ and $y\mapsto\sigma_{Y}(y)$ are uniformly continuous with moduli $\omega_{\Sigma}(r),\omega_{\sigma_{Y}}(r)\to 0$ and the policy map is locally Lipschitz on the costate band, then for small $r$,
+      $\omega_{a}(r)\;\leq\;C\bigl{(}\omega_{\Sigma}(r)+\omega_{\sigma_{Y}}(r)\bigr{)}.$
+      If $\omega_{\Sigma}(r),\omega_{\sigma_{Y}}(r)\leq C_{0}r^{\alpha}$, choose $R_{0}\leq(\eta_{0}/(CC_{0}))^{1/\alpha}$.
+    See Dong and Kim, (2009) for the VMO framework in parabolic equations.
+  2. 2.
+    Norm coherence. All bounds are in $L^{q,p}(Q_{T})$ / $W^{2,1}_{q,p}(Q_{T})$. If elsewhere an equilibrium operator $T:X\to X^{\ast}$ is used, fix $X=W^{1,p}(\Omega)$ with
+    $\|u\|_{X}\;=\;\bigl{(}\|u\|_{L^{p}}^{p}+\|\nabla u\|_{L^{p}}^{p}\bigr{)}^{1/p},$
+    and state coercivity/monotonicity in this norm (avoid using only $\|u\|_{L^{p}}$). See Adams and Fournier, (2003) for Sobolev space conventions.
+  3. 3.
+    Operator interchanges (Fubini–Tonelli/DCT).
+With bounded controls and reflection on $C^{1,1}$ domains, moments are finite, so for $\varphi\in C^{1}$ of polynomial growth and integrable weight $\psi$,
+    $\frac{d}{dt}\,\mathbb{E}\bigl{[}\psi(t)\,\varphi(t,X_{t},Y_{t})\bigr{]}\;=\;\mathbb{E}\bigl{[}\psi^{\prime}(t)\,\varphi(t,X_{t},Y_{t})\bigr{]}\;+\;\mathbb{E}\bigl{[}\psi(t)\,(\partial_{t}+\mathcal{L})\,\varphi(t,X_{t},Y_{t})\bigr{]},$
+    and dominated convergence justifies interchanging $\int_{0}^{T}$ and $\mathbb{E}[\cdot]$. See Karatzas and Shreve, (1991) for reflected SDE background.
+  4. 4.
+    Numerical convergence (strong vs. weak). For Euler–Maruyama with step $h$ and smooth coefficients:
+    $\bigl{(}\mathbb{E}\|Z_{T}^{\mathrm{EM}}-Z_{T}\|^{2}\bigr{)}^{1/2}=\mathcal{O}(h^{1/2})\quad\text{(strong)},\qquad\bigl{|}\mathbb{E}\phi(Z_{T}^{\mathrm{EM}})-\mathbb{E}\phi(Z_{T})\bigr{|}=\mathcal{O}(h)\quad\text{(weak)}.$
+    This matches Definition 2; Milstein/RE gives first-order in the discretization term. See Platen and Bruti-Liberati, (2010) for SDE numerical analysis.
+  5. 5.
+    Time-slab vs. time-step. The contraction uses an analytic slab length $\tau\leq\tau_{\star}$ tied to PDE constants; it is independent of the numerical step $\Delta t$.
+  6. 6.
+    Optional GE module (if used elsewhere). If you introduce $T:X\to X^{\ast}$ with
+    $\langle T(u),v\rangle=\int_{\Omega}\!\Big{(}A\nabla u\cdot\nabla v+Buv+N(u)\,v\Big{)}\,dx,$
+    $A\succeq\underline{\lambda}I$, $B\geq 0$, and $N$ monotone with $N(u)u\geq c_{0}|u|^{p}-C_{0}$, then
+    $\langle T(u),u\rangle\;\geq\;c\,\|u\|_{W^{1,p}}^{p}-C,$
+    so coercivity condition holds in the same $W^{1,p}$ norm. See Browder, (1965) for monotone operator theory.
+Appendix D Classification of Parabolic HJB PDEs in Financial Models
+Table 6 summarizes the classification of parabolic HJB PDEs in financial markets.
+Although the theoretical proofs for classes (H) and (D) are challenging, our P-PGDPO is applicable to the broader classes of parabolic PDEs from the financial and economic models.
+Table 6: Classification of Parabolic HJB PDEs in Financial Models
+Class Diffusion Matrix $a(t,x,\mathbf{y})$ Property SDE Coefficient Requirements and Dynamic Portfolio Examples
+(U) Uniformly Elliptic Uniformly positive definite across the state space The diffusion coefficients $\bm{\sigma}$ (for wealth $X_{t}$) and $\bm{\sigma}_{Y}$ (for state $\mathbf{Y}_{t}$) are such that the resulting diffusion matrix in the HJB is non-zero and bounded below by a positive constant everywhere (e.g., $\bm{\sigma}\Psi\bm{\sigma}^{\top}$ and $\operatorname{diag}(\bm{\sigma}_{Y})\Phi\operatorname{diag}(\bm{\sigma}_{Y})^{\top}$ uniformly positive definite).
+Drift coefficients ($\bm{\mu},\bm{\mu}_{Y}$) are also sufficiently regular (e.g., Lipschitz).
+($\implies$ Value Function / Costate very smooth) Examples: Merton’s portfolio problem with constant coefficients ($\mu,\sigma>0,r$ are constants); models with state-dependent volatility that remains strictly positive; the multi-asset OU model with one factor in Section 3.2.1 (if asset volatilities $\sigma_{i}>0$, factor volatility $\sigma_{Y}>0$, and $\Psi$ is positive definite).
+(H) Hypo-elliptic May degenerate (even globally).
+Regularity is achieved if the system structure satisfies conditions like Hörmander’s (i.e., Lie brackets of vector fields span the tangent space).
+The problem might be uniformly elliptic on compact subsets of interest.
+Diffusion coefficients ($\bm{\sigma}$ or $\bm{\sigma}_{Y}$) can become zero at certain points within the domain.
+However, interaction with non-zero drift terms or other diffusion terms maintains regularity.
+($\implies$ Value Function / Costate smooth in relevant regions) Examples: The multi-asset and multi-factor model in Section 3.2.1 (if some factor volatilities in $\bm{\sigma}_{Y}$ can be zero at certain states but the overall system structure, possibly involving $\kappa_{Y}$, ensures hypo-ellipticity); a Vasicek interest rate model affecting asset returns $\mu$ but not $\sigma$.
+(D) Degenerate at the Boundary Degenerates ($\to 0$) as the state approaches a boundary.
+The diffusion coefficients $\bm{\sigma}$ or $\bm{\sigma}_{Y}$ (or the combined diffusion matrix) approach zero as a state variable (e.g., wealth, volatility, interest rate) approaches a boundary (typically 0).
+($\implies$ Value Function / Costate regularity may break down near the boundary) Examples: CIR (Cox-Ingersoll-Ross) interest rate/volatility model (diffusion term like $\sigma_{Y}\sqrt{Y_{t}}$); Heston stochastic volatility model (asset diffusion term like $\sigma\sqrt{\nu_{t}}$).
+Appendix E Detailed Experimental Setup
+This section provides a detailed setup for the numerical experiments presented in Section 5. The implementation code and detailed explanations can be found at https://github.com/huhjeonggyu/PGDPO.
+E.1 Parameter Values
 The experiments consider the objective of maximizing the expected utility of terminal wealth $X_{T}$, $J=\mathbb{E}[U(X_{T})]$, using a CRRA utility function $U(x)=x^{1-\gamma}/(1-\gamma)$ with a relative risk aversion coefficient $\gamma=2.0$.
@@
     • Factor correlation $\Phi_{Y}$ ($k\times k$) generated as a random correlation matrix.
-    • Cross-correlation $\bm{\rho}_{Y}$ ($n\times k$) components drawn uniformly from $U(-0.2,0.2)$.
+    • Cross-correlation $\bm{\rho}_{Y}$ ($n\times k$) components drawn uniformly from U(0.2,0.2)U(-0.2,0.2).
     • The full $(n+k)\times(n+k)$ block correlation matrix $\begin{pmatrix}\Psi&\rho_{Y}\\ \rho_{Y}^{\top}&\Phi_{Y}\end{pmatrix}$ is constructed and numerically adjusted (e.g., via minimal diagonal loading) if necessary to ensure positive definiteness (tolerance $10^{-6}$).
 The specific values generated by the seed are used consistently across experiments for given $n$ and $k$.
-D.2 PG-DPO Implementation Details
+E.2 PG-DPO Algorithm Details
+Each iteration $j$ of the algorithm involves the following steps:
+  1. Step 1.
+    Sample Initial States: Draw a mini-batch of $M$ initial states $\{(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)})\}_{i=1}^{M}$ from the distribution $\eta$ over the domain $\mathcal{D}$.
+  2. Step 2.
+    Simulate Forward Paths: For each sampled initial state $i$, simulate a discrete-time trajectory $\{(X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\}_{k=0}^{N}$ forward from $t_{k}=t_{0}^{(i)}$ to $t_{N}=T$, using the current policy networks $(\bm{\pi}_{\theta},C_{\phi})$.
+The state evolution follows:
+    $\mathbf{Y}_{k+1}^{(i)}$ $=\mathbf{Y}_{k}^{(i)}+\bm{\mu}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta t^{(i)}+\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta\mathbf{W}_{k}^{(i),Y},$
+    $X_{k+1}^{(i)}$ $=X_{k}^{(i)}+\left[X_{k}^{(i)}\left((1-\bm{\pi}_{k}^{(i)\top}\mathbf{1})r(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})+\bm{\pi}_{k}^{(i)\top}\bm{\mu}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\right)-C_{k}^{(i)}\right]\Delta t^{(i)}$
+    $\quad+X_{k}^{(i)}\bm{\pi}_{k}^{(i)\top}\bm{\sigma}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\Delta\mathbf{W}_{k}^{(i),X},$
+    where $\Delta t^{(i)}=(T-t_{0}^{(i)})/N$, $\bm{\sigma}_{Y}(t_{k}^{(i)},\mathbf{Y}_{k}^{(i)})$ is the vector of diffusion coefficients applied element-wise to the components of $\Delta\mathbf{W}_{k}^{(i),Y}$, and the correlated Brownian increments $(\Delta\mathbf{W}_{k}^{(i),X},\Delta\mathbf{W}_{k}^{(i),Y})$ are derived from the Cholesky factor $L$ of the covariance matrix $\bm{\Omega}$.
+  3. Step 3.
+    Calculate Realized Reward per Path: For each path $i$, compute the realized cumulative reward, explicitly noting its dependence on the starting state and policy parameters, denoted as $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)$:
+    $J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi)=\sum_{k=0}^{N-1}e^{-\delta(t_{k}^{(i)}-t_{0}^{(i)})}U\bigl{(}C_{\phi}(t_{k}^{(i)},X_{k}^{(i)},\mathbf{Y}_{k}^{(i)})\bigr{)}\Delta t^{(i)}+\kappa e^{-\delta(T-t_{0}^{(i)})}U(X_{N}^{(i)}).$
+    This value $J^{(i)}(t_{0}^{(i)},\dots)$ is a stochastic sample related to the conditional expected utility $J(t_{0}^{(i)},\dots)$ defined in Eq. (4).
+  4. Step 4.
+    Apply BPTT for Policy Gradient Estimate: Treat the simulation process for path $i$ as a computational graph.
+Apply automatic differentiation (BPTT) to compute the gradient of the realized reward $J^{(i)}(t_{0}^{(i)},\dots)$ with respect to the policy parameters $(\theta,\phi)$:
+    $\nabla_{(\theta,\phi)}J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi).$
+    This gradient serves as an unbiased estimate related to the gradient of the conditional expected utility $J(t_{0}^{(i)},\dots)$.
+  5. Step 5.
+    Estimate Gradient of $\widetilde{J}$: Obtain an unbiased estimate for the gradient of the overall extended objective $\widetilde{J}$ by averaging the individual path gradients over the mini-batch:
+    $\nabla\widetilde{J}(\theta,\phi)\approx\nabla_{\text{batch}}=\frac{1}{M}\sum_{i=1}^{M}\nabla_{(\theta,\phi)}J^{(i)}(t_{0}^{(i)},x_{0}^{(i)},\mathbf{y}_{0}^{(i)};\theta,\phi).$
+  6. Step 6.
+    Update Policy Parameters: Update the parameters $(\theta,\phi)$ using this averaged gradient estimate via stochastic gradient ascent:
+    $(\theta,\phi)_{j+1}\leftarrow(\theta,\phi)_{j}+\alpha_{j}\,\nabla_{\text{batch}},$
+    where $\alpha_{j}$ is the learning rate.
+By iterating steps from 1 to 6, the baseline PG-DPO algorithm gradually optimizes the policy $(\bm{\pi}_{\theta},C_{\phi})$ to maximize the extended value function $\widetilde{J}$.
+E.3 PG-DPO Implementation Details
 We implement the Baseline PG-DPO algorithm using PyTorch.
   • Policy Network ($\bm{\pi}_{\theta}$): The policy network consists of an input layer taking the current wealth $W_{t}$, current time $t$, and state vector $\mathbf{Y}_{t}$ (total $2+k$ inputs), followed by three hidden layers with 200 units each and LeakyReLU activation, and an output layer producing the $n$-dimensional portfolio weight vector $\bm{\pi}_{\theta}$.
-  • Optimizer: Adam optimizer with a learning rate of $1\times 10^{-5}$.
+  • Optimizer: Adam optimizer with a learning rate of $1.0\times 10^{-5}$.
   • Training:
@@
     • Batch size ($M$): $1,000$.
-    • Time discretization steps ($N$): $20$.
+    • Time discretization steps ($N$): 2020.
     • Initial state sampling ($\eta$): For each trajectory in a batch, the initial time $t_{0}$ is drawn uniformly from $[0,T)$, initial wealth $W_{0}$ from $U(0.1,3.0)$.
-The initial state variables $\mathbf{Y}_{0}=(Y_{0,1},\dots,Y_{0,k})^{\top}$ are sampled such that each component $Y_{0,i}$ is drawn independently and uniformly from a single common range $[Y_{\min},Y_{\max}]$.
-This range is determined by the minimum lower bound and the maximum upper bound observed across all individual factors’ typical ±$\pm 3\sigma$ intervals:
+The initial state variables $\mathbf{Y}_{0}=(Y_{0,1},\dots,Y_{0,k})^{\top}$ are sampled such that each component $Y_{0,i}$ is drawn independently and uniformly from a single common range $Y_{\min},Y_{\max}]$.
+This range is determined by the minimum lower bound and the maximum upper bound observed across all individual factors’ typical $\pm 3\sigma$ intervals:
       $Y_{\min}=\min_{i=1,\dots,k}(\theta_{Y,i}-3\sigma_{Y,ii}),\quad Y_{\max}=\max_{i=1,\dots,k}(\theta_{Y,i}+3\sigma_{Y,ii}),$
@@
   • Costate Estimation (for Two-Stage Evaluation): During periodic evaluation, costate $\lambda$ and derivatives $\partial_{x}\lambda,\partial_{\mathbf{Y}}\lambda$ are computed via BPTT (‘torch.autograd.grad’) applied to Monte Carlo estimates of the expected terminal utility starting from evaluation points $(t_{k},W_{k},\mathbf{Y}_{k})$.
-D.3 Benchmark Solution and Evaluation
+E.4 Benchmark Solution and Evaluation
 As mentioned, the chosen multi-factor OU model allows for a semi-analytical benchmark solution.
@@
 This simplification is made specifically to enable direct comparison with the semi-analytical benchmark solution derived from the HJB equation, which is readily available for the terminal wealth problem in this affine setting.
-  • Benchmark Computation: The system of ODEs (including a matrix Riccati equation) associated with the HJB equation for this terminal wealth problem (Appendix B.2) is solved numerically backwards from the fixed horizon $T$ using ‘scipy.integrate.solve_ivp’ (‘Radau’ method) to obtain the coefficients determining the value function and the optimal policy $\bm{\pi}^{*}(t,W_{t},\mathbf{Y}_{t})$.
+  • Benchmark Computation: The system of ODEs (including a matrix Riccati equation) associated with the HJB equation for this terminal wealth problem (Appendix LABEL:app:analytic_MF) is solved numerically backwards from the fixed horizon $T$ using ‘scipy.integrate.solve_ivp’ (‘Radau’ method) to obtain the coefficients determining the value function and the optimal policy $\bm{\pi}^{*}(t,W_{t},\mathbf{Y}_{t})$.
   • Interpolation: The benchmark policy (total, myopic, and hedging components) is pre-computed on a grid over $(t,W_{t},\mathbf{Y}_{t})$ space and stored using ‘scipy.interpolate.RegularGrid Interpolator’ for efficient lookup.
   • Evaluation Metrics: Performance is evaluated via the Root Mean Squared Error (RMSE) between the policies generated by (i) Baseline PG-DPO ($\bm{\pi}_{\theta}$) and the benchmark ($\bm{\pi}^{*}$), and (ii) Two-Stage PG-DPO (using Eq.
-(11) with BPTT-estimated costates) and the benchmark ($\bm{\pi}^{*}$).
+(3.2) with BPTT-estimated costates) and the benchmark ($\bm{\pi}^{*}$).
 RMSEs are computed for total, myopic, and hedging components.
@@
 All computations are performed using PyTorch on an NVIDIA A6000 GPU.
-D.4 Deep BSDE Implementation Details
+E.5 Deep BSDE Implementation Details
 The Deep BSDE method, serving as a benchmark, was implemented based on the standard framework by Han et al., (2018) and Weinan, (2017).
@@
 The simplification stems from the value function’s known separable form in wealth (i.e., $V\propto X_{t}^{1-\gamma}$).
-This property allows the general intertemporal hedging demand term, $-\frac{1}{X_{t}V_{xx}}\Sigma^{-1}(\bm{\sigma\rho\sigma_{Y}})V_{x\mathbf{y}}$, to be analytically transformed into the more tractable expression $\frac{1}{\gamma}\Sigma^{-1}\left(\bm{\sigma}\bm{\rho}\right)\left(\bm{\sigma}_{Y}\frac{\nabla_{\mathbf{y}}V}{V}\right)$.
+This property allows the general intertemporal hedging demand term, $\frac{1}{X_{t}V_{xx}}\Sigma^{-1}(\bm{\sigma\rho\sigma_{Y}})V_{x\mathbf{y}}$, to be analytically transformed into the more tractable expression $\frac{1}{\gamma}\Sigma^{-1}\left(\bm{\sigma}\bm{\rho}\right)\left(\bm{\sigma}_{Y}\frac{\nabla_{\mathbf{y}}V}{V}\right)$.
     The Deep BSDE implementation then creates a policy by combining an analytical myopic demand with a network-based approximation of this tractable hedging term.
@@
 Gradient clipping (norm 2.0) and bounding of portfolio weights are applied to enhance training stability.
-  • Initial State Sampling: For each training path, the initial time $t_{0}$ is sampled uniformly from the discrete time steps over $[0,T-\Delta t]$.
-The initial wealth $X_{0}$ is drawn uniformly from $[0.1,3.0]$, and initial factor states $\mathbf{Y}_{0}$ are sampled uniformly from a hyperrectangle representing their typical dynamic range (e.g., based on ±$\pm 3$ stationary standard deviations around their long-term means).
+  • Initial State Sampling: For each training path, the initial time $t_{0}$ is sampled uniformly from the discrete time steps over $0,T-\Delta t]$.
+The initial wealth $X_{0}$ is drawn uniformly from $[0.1,3.0]$, and initial factor states $\mathbf{Y}_{0}$ are sampled uniformly from a hyperrectangle representing their typical dynamic range (e.g., based on $\pm 3$ stationary standard deviations around their long-term means).
 The implementation uses PyTorch and is executed on the same NVIDIA A6000 GPU hardware as the PG-DPO experiments to ensure a fair comparison.
+Appendix F Additional Numerical Results - Convergence Speed, Accuracy, and Policy Error Analysis
+(a) n=1,k=10n=1,\;k=10
+(b) n=10,k=10n=10,\;k=10
+(c) n=50,k=10n=50,\;k=10
+Figure 5: Average policy RMSE comparison between Baseline PG-DPO (e.g., blue), Projected PG-DPO (P-PGDPO) (e.g., orange), and Deep BSDE (e.g., green) relative to the benchmark solution across different asset dimensions ($n$) with $k=10$ factors.
+Values averaged over $k=10$ evaluation ‘slices.’ For each slice $j=1,\dots,10$, the evaluation (e.g., for RMSE calculation or visualization) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$.
+Note the logarithmic scale on the $y$-axis.
+This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO framework—including both the Baseline and Projected PG-DPO (P-PGDPO) variants—against the Deep BSDE method, when applied to derive control policies.
+We evaluate performance across various problem dimensions, focusing on how the number of assets ($n$) and the number of state factors ($k$) affect policy accuracy.
+Figure 5 shows the average Root Mean Squared Error (RMSE) between the learned or derived policies and the semi-analytical benchmark policy.
+The RMSE is plotted against the number of training iterations on a logarithmic scale for experiments with $k=10$ state variables and varying asset counts ($n=1,10,50$).
+For P-PGDPO, the reported RMSE corresponds to the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ (Eq.
+(3.2)), which combines the myopic and intertemporal hedging components.
+As discussed in Section 5.2, the myopic error is typically negligible—on the order of $O(10^{-7})$.
+Thus, the total RMSE for P-PGDPO is effectively governed by the accuracy of the hedging component, which depends on estimating the costate gradient $\partial_{\mathbf{Y}}\lambda_{t}^{*}$ via BPTT.
+Figures 6 and 7 display the error maps across models.
+In particular, Figure 6 illustrates the policy errors for the first asset ($i=1$) across three models-baseline PG-DPO, P-PGDPO, and Deep BSDE- in the high-dimensional setting with $n=50$ risky assets and $k=10$ state variables.
+The figure plots the difference between the total portfolio policy (i.e., the sum of the myopic and intertemporal hedging demands) generated by each method and the corresponding analytic benchmark solution.
+Panel (a) shows the Baseline PG-DPO, where substantial deviations from the benchmark are visible, particularly near maturity and for extreme values of $Y_{1}$.
+Panel (b) reports the results for P-PGDPO, which achieves significantly lower errors overall.
+The projection step based on stabilized costates produces policies that are visually close to the benchmark, with errors an order of magnitude smaller than in the baseline.
+Panel (c) presents the Deep BSDE method, which captures the broad structure of the policy but exhibits notable inaccuracies in regions where intertemporal hedging demand is critical.
+(a) Baseline PG-DPO Error
+(b) P-PGDPO Error
+(c) Deep BSDE Error
+Figure 6: Policy errors for the first asset ($i=1$) across models in the case where $n=50,k=10$.
+Errors are plotted against $Y_{1}$ at a representative time-to-maturity TtT-t (e.g., $t=(T-T_{0})/2$ if $T_{0}$ is initial time, assuming $T=1.5$ is the horizon from $t=0$; thus, for instance, at $t=0.75$ or time-to-maturity $0.75$).
+Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
+(a) Myopic Error
+(b) Hedging Error
+Figure 7: Decomposition errors of the Projected PG-DPO (P-PGDPO) policy for the first asset ($i=1$) in the $n=50,k=10$ case (results from iteration 6,400 as plotted), plotted against $Y_{1}$ at a representative time-to-maturity TtT-t.
+The first figure shows the error for myopic demand and the second figure shows the error of intertemporal hedging demand.
+Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$.
+In addition to the total portfolio errors, Figure 7 provides a decomposition of the policy errors into the myopic and intertemporal hedging components for the first asset ($i=1$) under the setting with $n=50$ risky assets and $k=10$ state variables.
+The left panel shows the error map for the myopic demand, while the right panel shows the error for the intertemporal hedging demand, both plotted against the first variable $Y_{1}$ and time-to-maturity (Tt)(T-t), with all other state variables fixed as their long-run means.
+The results demonstrate that the myopic demand is recovered with residual errors on the order of $10^{-6}$, confirming the robustness of the costate-driven projection step for the static component of the portfolio.
+By contrast, the hedging demand exhibits larger and more structured errors, concentrated in regions where horizon effects are most pronounced.
+This reflects the intrinsic difficulty of estimating the cross-derivatives of the value function that drive intertemporal hedging behavior, particularly in high-dimensional environments.
+In sum, Figures 6 and 7 illustrate that while P-PGDPO achieves highly accurate recovery of both myopic and hedging components, the primary source of residual policy error originates from the hedging term, underscoring the importance of our methodological focus on stabilizing and projecting costate estimates to capture horizon-dependent dynamics.